A Complete Bibliography of Publications in Conservation Biology: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.06

Title word cross-reference

th [Mef06a]. ♡ [Din03b].

0Pacific [BMP07].

10 [Joh09a]. 13 [Koh09]. 16-Year [New06]. 1990s [CAR05]. 1996-1998 [JML00].

2 [Mor00a]. 2000 [MFGB07]. 2003 [Ano03f, Ano04i, Ano04k]. 2004 [Ano03a, Ano04a, Ano04h, Ano04j, Ano04g, Ano04l, Ano05j]. 2005 [Ano05a, Ano05b, Ano05i, Ano05h, Ano05l, Ano06j]. 2006 [Ano06a, Ano06b, Ano06h, Ano06i, Ano06g, Ano06k, Ano07o, Ano07r]. 2007 [Ano07a, Ano07d, Ano07p, Ano07q, Ano07n, Ano08n, Ano08q, LD07]. 2008 [Ano08p, Ano08o, Ano08m, Ano09k, Ano09m]. 2009 [Ano09b, Ano09l, Ano09j, Ano09n]. 2010 [MPE^{+}07, RP07]. 21st [Whi08, SHB08]. 22-Year [LLV^{+}02]. 28-Year [CBG^{+}00].

300th [SGLB^{+}07].

9-11 [Orr02a].

Abajo [AAO^{+}09, EK08, MO04]. Abajo-Arriba [AAO^{+}09]. Abalón [RBHKK02]. Abalone [RBP01, RBHKK02]. Abalones [FVS03]. Abandon [ML00, Vit04]. Abandonados [SZBR03, SRH^{+}07]. Abandonar [CDM09]. Abandoned [SZBR03, SRH^{+}07]. Abbreviated [MA05]. Abejas [EBAW01, FMPB07, KSDBT02, KT02, PZG^{+}05, SD03, WGK07]. Abejorros [WCX09]. Abierto [BVH^{+}02, GNEE07]. Abiertos [HSR05, PV09a]. Abilities [LLW02]. Ability [CHFR09, SWK05, SBDD05]. Abogacía [NV09]. Abortion [Led06]. Abroad [CHMP05]. Absence [BHW01, JFWP06, Pol06]. Absolutas [LWV04]. Absolute [LWV04]. Abstractions [Say09]. Abulón [RBP01, FVS03]. Abundance [AGBP09, BHB08, BSW^{+}06, BVH^{+}02, CWH07, CLW^{+}09, DOGN08, DWN01, EM08, FSC02, For00b, GV02, GWS^{+}06, GLMN00, GM00b, GM09, Gus02, HCRM09, Hew04, JFWP06, KHHK03, LS09, LBPBG07, LFR^{+}08, LF00, LB02, MNB09, MR00b, MEA00, MCM^{+}00, MP07b, RBK^{+}05, Roy04, SRH^{+}07, SLS05, SBSL06, WHJJ04, WZD^{+}00]. Abundance-Area [MP07b]. Abundances [Bal02]. Abundancia [AGBP09, BHB08, BSW^{+}06, BVH^{+}02, CWH07, CLW^{+}09, DOGN08, DWN01, EM08, FSC02, For00b, GV02, GWS^{+}06, GLMN00, GM00b, GM09, Gus02, HCRM09, JFWP06, KHHK03, LS09, LFR^{+}08, LF00, LB02, MNB09, MR00b, MEA00, MCM^{+}00, MP07b, RBK^{+}05, Roy04, SLS05, SBSL06, WZD^{+}00, SRH^{+}07, WHJJ04]. Abundancia-Área [MP07b]. Academia [Du 01, Fle01, WHM01]. Academic [Far04, GOL^{+}07, MGB^{+}07, MS09]. Académica [MS09]. Academically [ID00]. Académico [MS09]. Academy [EYZ09]. Acaena [LLW02]. Acantilado [MNL03]. Acantilados [KL06]. Acceleration [EB09]. Acceptability [SS06]. Acceptable [Ber03]. Acceptance [MM03b, Ste08a]. Acceso [DC06b, PL03]. Access [DC06b]. Accessibility [PL03]. Accessible [Coc04]. Accidental [DLCV00, MBR^{+}02]. Accidentally [You00]. Acción [AHR^{+}09, HLG03]. Acciones [BWS09, Dun02, KCC06, MCG^{+}09, Pyk05, SSS^{+}08, The03]. According [GOL^{+}07]. Accounting [BMP^{+}06, Pra05, SFT00]. Accounts [PDS^{+}07]. Accumulation [DFS05, ZGR02]. Accuracy [AM05, MPDT01]. Accurate [SS08a]. Aceleración [EB09]. Acelerado [Nos01a]. Aceptabilidad [SS06]. Acercamiento [BRV06, MPE^{+}07]. Acervos [ASF^{+}09]. Achievable [Coc04]. Achieved [WPS08]. Achieving [KB06]. Acknowledgments [Ano01b, Ano01c, Ano02c, Ano04c, Ano05d, Ano06d, Ano07i, Ano08i, Ano09i, Fie09, Mef03a]. Acopio [PCME09]. Acquisition [Cze02b, TW06, TWS06]. across [Bur04, CVS^{+}09, DSE^{+}08, DWW04, Far04, GCV03, HKN01, KS05, LH04, MKG03, RG02, TAJ04, VMM07, WCX09]. Act [Lov06, Pau08, ALH^{+}04, BZD03, CK05, CHBK02, DZM08, DP01, Fal07, FEGB09, Gre09, HBC^{+}06, HP01, RM01, SDRO01, VNP06]. Acta [RM01, ALH^{+}04, BZD03, CHBK02, DP01, Fal07, FEGB09, HBC^{+}06, HP01, SDRO01, VNP06]. Acting [HLG03]. Action [AHR^{+}09, Bre02a, DTCC03, Dun02, GMCP09, JLRBB^{+}07, KCC06, Mar02b, Mer08, Orr07b, Pyk05, The03]. Actions [BWS09, MCG^{+}09, SSS^{+}08]. Actitudes [AJF03, GEJK05, JSJV03, NTGT03]. Active [MP07a]. Actividad [BHB08, CNM01, Ric04b, CFH^{+}02]. Actividades [BJM04, Bor08, Fer01, GBH^{+}09, HLB^{+}03, IMD01, RS02b]. Activities [BJM04, Bor08, HLB^{+}03]. Activity [BHB08, CNM01, CFH^{+}02, GBH^{+}09, PCHN04, Ric04b]. Activo [MP07a]. Actual [EG07, GHL08, MRAF^{+}07]. Actuales [Gut05, MT08, PTP^{+}08, Web01]. Actually [Gro05b]. Acuacultura [Fra09]. Acuario [KB03, TH03]. Acuarios [MCR^{+}04]. Acuática [ML04, RWBG06]. Acuáticas [BTK08, HH08, LRO^{+}08, RO08, RBT08, RS02b, TF00]. Acuáticos [Sem02]. Acuerdos [WMPC08]. Acuicultura [VTRG00]. acuminata [BK02]. Acumulación [DFS05, ZGR02]. Adaptabilidad [EKL^{+}00, For02, GNEE07, Kel01, LM05, NT01, RF03, Ree05, WG06, FDBM04]. Adaptación [BNP^{+}07, Bro08b, CWNW01, LRS^{+}09b, MOO09, PRL03, VSO04, WRL02]. Adaptation [Bro08b, LRS^{+}09b, MOO09, McK09, PRL03, PAB^{+}08, SA04, VSO04]. Adaptativa [DA07, GGA01, VSAG06]. Adaptativas [DW05a, TW06]. Adaptativo [GWS07, JMS^{+}06a, MP07a]. Adaptive [Bei03, Blu07, BNP^{+}07, DW05a, DGG^{+}01, DA07, GGA01, GWS07, IF02, JMS^{+}06a, MP07a, TW06, VSAG06, Wil02a]. Adaptivo [IF02, Wil02a]. Addition [Sal02]. Additional [LW02a]. Additions [O’C05]. Additive [GCV03]. Addressing [BC08, GZK06, VO03]. Adds [ZR01]. Adecuación [KLH08, LARR03]. Adecuado [PD03a]. Adelante [Rod03]. Adequacy [LARR03]. Adequate [BHW01]. Adicional [LW02a]. Aditiva [GCV03]. Adjacent [Ant08, FB00, SS02a]. Adjustments [Pow06]. Administración [BMP^{+}06]. Admixture [KGS^{+}07]. ADN [BST^{+}05, COW03, JGL^{+}02, PBSR06, RLC^{+}00, UWB04, WCD^{+}08]. ADNm [AMF^{+}01]. ADNmt [LCS04, TCSD00]. Adoption [Pro04]. Adoration [Bye04]. Adquisición [Cze02b, TW06, TWS06]. Adrift [Bei01, Woo01]. Adult [GMW08, MMLM01, WBW06]. Adultas [GMW08]. Adultos [MMLM01, RSCL04, SL00a, WBW06]. Adults [RSCL04, SL00a]. Advanced [Mar04]. Advances [Cla09]. Advancing [QR06]. Advantage [ME00]. Advantages [SHJ^{+}09]. Advent [PHW05]. Advocacy [BT07b, Cha08, Cze07a, GC09, Lac07, Mef07a, ML00, NV09, SRL^{+}07, Wri07]. Adyacentes [FB00, SS02a]. Aegla [PLBBJC09]. AES [LME08]. aesculifolia [HDQSL06]. Aesthetics [AFK^{+}09]. Afecta [Vee00]. Afectada [For00a]. Afectan [CSB07, CDM09, ET03, FBCG00, Rue05]. Affairs [Fie00b]. Affect [CSB07, CDM09]. Affected [For00a]. Affecting [ET03, FBCG00, Rue05]. Affects [Vee00]. Africa [AAO^{+}09, Ant08, Awo08, BBB^{+}07, BH04, Bra03a, Büs08, CCS^{+}03, DNZK02, Fje07, GCF^{+}07, Inf01, Joh01, KTN^{+}09, KDC^{+}06, LMP00, LADM05, LFA^{+}07, MB00, MR00a, ND04, OALM^{+}00, Pim07b, TB03, Wes03, WSA^{+}05, Ant08, CCS^{+}03, WSA^{+}05, AAO^{+}09, BVK02, Bra03a, CPME09, Fje07, GCF^{+}07, HVR08, KTN^{+}09, KDC^{+}06, LCT^{+}06, LADM05, TB03]. African [BVK02, BRV06, CBG^{+}00, CPME09, COW03, CH01, Cre01a, CB01b, CW08, FPW01, FBL^{+}09b, Ful00, GMW08, HVR08, LCT^{+}06, LCG^{+}08, MR00a, OALM^{+}00, OWOF03, RMSG^{+}03, RH09, RSB05, RB06, VGB^{+}04, WCD^{+}08, WSQP07]. Africano [COW03, FBL^{+}09b, RH09, VGB^{+}04, WCD^{+}08, RMSG^{+}03]. Africanos [CBG^{+}00, CH01, Cre01a, CB01b, CW08, FPW01, GMW08, LCG^{+}08, OWOF03, RSB05, RB06, WSQP07]. Afromontane [LMP00]. Afromontañoso [LMP00]. Afrontar [RB07]. Afrotropical [KFBG08]. after [BB06, GV02, GMHS00, Gri00, Kni06, LRS^{+}09b, MS02, MRM^{+}05, NL06]. Again [Dix07, Pri01a]. against [APSB02, Fol06, Hod04b]. Agalleras [DLCV00, Dav07, SFT00]. Age [Ano02b, Ara09a, Bec09, HY03, Ros06]. Agencias [KB08]. Agencies [KB08]. Agency [Far04]. Agenda [Cor04a, RRCOM^{+}06, SMRR02, VBRNP08]. Agendas [CPR^{+}07]. Agent [Mcc05]. Agente [Mcc05, VK07]. Agentes [Rud06]. Agents [Mer08, Rud06]. Aggregations [SCD^{+}08]. Agotamiento [Rom03]. Agraria [CAR05]. Agrarios [MHGF04]. Agreements [WMPC08]. Agregaciones [SCD^{+}08]. Agri [KBS^{+}04]. Agri-Ambientes [KBS^{+}04]. Agri-Environment [KBS^{+}04]. Agrícola [BBH06, Fau01, FVMG03, FCL00, JP09, PGP04, SLB^{+}07, SOADMR08, WHJJ04, WMS^{+}01]. Agrícolas [CLC^{+}08, DCP^{+}03, GLF^{+}08, GSB04, Her09, JSHM06, JSJ05, KBS^{+}04, RDEF01]. Agricultores [JSJV03]. Agricultura [CCC04, GH02, HBV02, NTMT^{+}03, PHW05, RL06]. Agricultural [BBH06, DCP^{+}03, Don04, Fau01, FVMG03, FCL00, GLF^{+}08, GSB04, HKC^{+}08, JP09, JSHM06, KBS^{+}04, MV06, PGP04, PJB^{+}06, RDEF01, SLB^{+}07, SOADMR08, VP07, WA00, WHJJ04, WMS^{+}01]. Agriculturally [RL06]. Agriculture [CCC04, GH02, HBV02, Jac08, Jor01, NTMT^{+}03, PHW05, San06, You02]. Agroecología [BJB^{+}02]. Agroecológica [PV02]. Agroecological [PV02]. Agroecology [BJB^{+}02]. Agroecosistemas [Lai05, RCP00]. Agroecosystems [Lai05, PMEH05, RCP00]. Agroforestales [KSDBT02]. Agroforestería [BMV09]. Agroforestry [BW08b, BMV09, Gri00, Gro05b, KSDBT02]. Agropecuarios [FBK02]. Agrosistemas [PMEH05]. Agrupados [LKG06]. Agrupamientos [JGL^{+}02]. Agua [CMC^{+}01, EHJ^{+}04, HBKF05, HGH^{+}03, PHW05, SMV02, XGS^{+}09]. Aguas [ATG05, RFD01]. Águilas [FNP09]. Ahead [RS06]. Ahus [CMM05a, CMM05a]. Aid [Nel09]. Aiding [Wil08b]. AIDS [Awo08]. Ain’t [Woo00]. Aislada [Bos07, TRSR00, War05]. Aisladas [FBL09a, NT01, Pan02]. Aislado [GWS^{+}06]. Aislados [GGGS00, RTN01, SKBB00]. Aislamiento [CHFR09, FRD^{+}09, FRSG^{+}02, GGW09, NR03, SSY06, TM04, WH01]. Ajustes [Pow06]. al [ADD^{+}07, BCK^{+}09, BS01a, BB00, BRS^{+}05, BZD03, Bro08b, BC09, CNHT09, DH00a, Dal05, DC06b, DBML07, EVD01, EKL^{+}00, FRD^{+}09, FF08, FBL^{+}09b, GMMPP08, Gha08, GM00b, Gor08, HM09, HLB^{+}03, KO01, Lam02, Lin03a, MB07, MBFT09, MBMLJ09, MP09, MOO09, MCG^{+}09, Mcs06, MPR^{+}01, MPE^{+}07, MDW01, NVP07, NB00, OHC08, Old03, Pan09, PJYO04, Pol06, PMEC04, RJW09, SBS^{+}09, SRW06, Sch10, SSY06, SHM04, SH02, SK02b, SCC^{+}09, Tay03b, Tem02, Tol07, TSKD08, UKC^{+}09, VSAG06, WABT08, WS03, WP03, Wil00a, You01, ZR01]. Alabama [IF02]. Alaka’i [FTC^{+}04b, FTC^{+}04b]. Alaska [Bei08, BHW01, Urb06b]. Alberta [CW00]. Alcance [PL03]. Alcanzando [DS08]. Alcelafo [FSS^{+}00]. Alces [CFH^{+}02]. alcon [Wal04]. Aldea [Mad05]. Aleatoria [MBQ05]. Aleatorio [HW08]. Alélica [FTC04a]. Alélicas [SPFF^{+}08]. Alerta [LVM02]. Aletas [CMA^{+}06]. Algodón [CCC04]. Algorithms [CM03, PC01]. Algoritmos [CM03]. Algunas [EG07, MB09b]. Algunos [LARR03]. Alianzas [Joh03b, SZ05]. Alien [ACMA04, CWL^{+}00, DOGN08, HCRM09, KS03, MS08, PA04, Pim07a, PJK03, STT05, Tal04]. Alimentarios [GBMH07]. Alimenticia [HCRM09]. Alimenticias [VSC06]. Alimenticio [CWNW01]. Alimento [AP00, RSW05]. Allá [Dal05, DWWF04, Gui07, SDJK04, ZSVL07, RH01]. Allee [ARB^{+}07, HM01, Ste08b]. Allegheny [Log03]. Allele [SPFF^{+}08]. Allelic [FTC04a]. Alliances [Joh03b, SZ05]. Allocate [MMBP08]. Allocation [JMP09, MDH^{+}07]. Allometry [DAR08]. Allozyme [Pod01]. Almeja [DSE^{+}08]. Alometría [DAR08]. Along [Col09, BMA^{+}08, GFTB04, PJ00, PA04, RR00, SRH^{+}07]. Alosa [BGWE00]. Alozímica [Pod01]. Alpinismo [MNL03]. Already [Bro05]. Alrededor [Ant08, LSZ^{+}08, SB03]. Alrededores [HR02, MYFF^{+}07]. Alta [BRYK01, GGGS00, Kel01, WMP05]. Altamente [FPL^{+}04]. Altas [PJYO04]. Alter [WZD^{+}00]. Alteración [GV05, LCG^{+}08]. Alteraciones [CGR^{+}05]. Alteran [WZD^{+}00]. Alteration [GV05, LCG^{+}08]. Alterations [CGR^{+}05]. Alternativa [CCC04, FO06]. Alternativas [AW02, BCS04, FMLL06, KHHK03, HM08]. Alternative [AW02, CWM08, CCC04, FMLL06, FO06, HN09, KHHK03, Kri04, O’K01b, SRSP01]. Alternatives [BCS04, HM08, Wil09a]. Alternativos [CWM08, HN09, Kri04]. Altiplano [LO06, SMC00]. Altitud [PA04]. Altitudinales [LZD^{+}03, SSFL08]. Alto [CCH03]. Alumbrado [BBM04]. Alumni [MS09]. Aman [Bat04]. Amarillas [KM00]. Amarillo [KSB01]. amarus [AT05]. Amazon [Bru02, CN06, Gav04, LW01, NARL^{+}04, NSA06, NSB^{+}06, PL03, PMZ06, PRMV06, RW02, SZ05, UMT^{+}01, ZSVL07]. Amazonas [UMT^{+}01, BMMR03, LW01, LF00, NSA06, PL03, RW02, SZ05, WLO^{+}00]. Amazonia [MRM^{+}05, BMMR03, Fea05, GV02, HR04c, LSL04, LF00, NSB^{+}06, Per05a, PPPV08, SBSL06, DS01, Fea05, Gav04, HR04c, LSL04, MRM^{+}05, Per01, PMZ06, PRMV06, SS01, ZSVL07]. Amazonian [DS01, BK02, CRD05, GV02, LLV^{+}02, LP08, MP07b, NAC^{+}06, Per00, PZ01, Per01, PDS^{+}07, SHM04, SBSL06, SS01, VCHBP08, WLO^{+}00]. Amazónica [MP07b, SHM04]. Amazónicas [CRD05, Per05a, VCHBP08]. Amazónico [BK02, LLV^{+}02, PDS^{+}07]. Amazónicos [LP08, NAC^{+}06, Per00]. Ambient [BOC00, SKP00]. Ambiental [BOC00, CHW^{+}01, Che09, CW09, DH00a, HSDS01, LME08, LOM03, LSF03, PS06, SRW06, SKP00, TK06, TRCC^{+}05, WSQP07]. Ambientales [CHW^{+}01, Ehr05a, FNFM01, FNF03, For00b, GGA01, LME08, LU09, LMP02, MCG^{+}09, NV09, SAPRB07, SJF^{+}05, SBDD05, SDL07, Sta01, Wun07]. Ambiente [DS08, SSB^{+}08]. Ambientes [FHB^{+}08, KBS^{+}04, RWJRR06, WMD^{+}04, ZG05]. Ambystoma [GGW09, Pyk05, TSG06]. Amenaza [Ant08, CGR^{+}05, Hay09b, MP09, MRDC06, ND04, NKW09, RL06, SL08, VCHBP08, KCS04]. Amenazada [GN00, GBC^{+}04, GBH^{+}09, PGP04, TR05, WABT07a]. Amenazadas [dLBBG^{+}08, BV01b, Cze05, DAB06, Fai08, GLMN00, HY03, JFWP06, JMP09, Kal05, LLWCW09, MCG^{+}08, McK02, MP08, MRAF^{+}07, MSSM09, NA01, NWRD03, OBK^{+}04, PCL^{+}03, TWS06, VMV07, YE07]. Amenazado [ZBABR00]. Amenazados [CV08, LS09]. Amenazantes [Hay09b]. Amenazas [BKV^{+}05, DGSS03, HSMK07, HTM^{+}05, RAW07, RSW05, SSS^{+}08, TFG08, The03, SRHS09, TSKD07]. Amendment [Pau08]. América [EM03, BHMC01, CVB^{+}09, Chr08, EM03, Fie00b, GHL08, Hit07, JG08, LRW03, MGB^{+}07, MSOM^{+}06, MHC^{+}01, MMB^{+}03, NFD^{+}09, PEC01, Rod03, RSPM05, Rot01, Rou00, SBVC03, VJ07, VMM07, Web04, YLR^{+}01, Yul06, CVB^{+}09, GK06, Hut06, JG08, LRW03, MHC^{+}01, MMB^{+}03, NMB09, OMP^{+}05, PEC01, SRW^{+}08, SBVC03, VJ07, VMM07]. American [Ank04, BGWE00, BMMW07, CW08, DC06a, DeN07, FMDK09, GK06, GCBF09, GNR01, HM01, HL07b, Hut06, KFK^{+}09, Mal03, Nug08, NMB09, Orr01c, OMP^{+}05, PLBBJC09, PAA^{+}08, Rob00, SRW^{+}08, VS05b, WD01]. Americana [BGWE00, JGL^{+}02]. Americanas [HL07b]. Americano [FMDK09, HM01, Rob00, SFS00, KFK^{+}09]. Americanos [CW08]. Americans [McB00]. Amerindian [WG01]. Amerindias [WG01]. Amiotrófica [MBC03]. among [ALR05, BJ05, BSKH^{+}06, BS00, GEJK05, Gib07, GSB^{+}06, Hob01, JMP09, KF02, LW01, LM07, LCS04, MMBP08, MA08, OBK^{+}04, Olu00, RS02a, RSC03, RIBGG08, SKBB00]. Amortiguamiento [CS07b, FPSD09, HRS08, HNG04, RS02b, SB03, WD03]. Amount [McK02, WSLW00]. Amphibian [AE01, ACC^{+}05, BBB08, BR01, Bee02, BFMM02, BBO^{+}01, BRK^{+}02, BRK^{+}03, BS01b, BSM08, BS00, CHW^{+}01, CF01, Cor03, DSJ02, GWS^{+}06, GH02, LRW03, MT01, MSBGP03, MHC^{+}01, MSSM09, OMP^{+}05, PC08, Pop08, RO04, RK05, SGG07, SM07a, Sch10, SF03, SKBB00, Sta01, TM02, YLR^{+}01]. Amphibians [CMCG07, FPSD09, GB09, GW05, GSB04, GP08, HRS08, JLZW03, KBM01, MDJP04, PC02, RGF^{+}04, RBK^{+}05, RB06, RS02c, SF04, SD02, Sem02, SB03, Sha04, SS05, SYWR03, SKP00]. Amphibolite [KU02, KU02]. Amplia [FNFM01, VVMV09]. Ampliación [FE07]. Amplificación [COW03]. Amplifying [COW03]. Amplio [BMA^{+}08]. Amplitud [HP08]. Amur [KGM^{+}02]. Amyotrophic [MBC03]. Anacapa [PLA00]. anacapae [PLA00]. Análisis [ABE^{+}02b, BBB08, BMP^{+}06, BFMM02, CTR02, Che09, CV08, DC05, Fai08, GH02, Har06a, HLSZ02, JGL^{+}02, KS05, MPDT01, MAP03, MM04, NLCAR09, PGP04, RMD^{+}02, RWJRR06, SK04, SH03, SHM04, SLCW08, Sta01, TFG08, TVB^{+}06, TWGH00, TK06, VMM07, VTA02, WGGS07, FP02, HPG09, KGM01, LVBS^{+}04, MP09, MWEB06, NLCAR09, PPBW08, QAMW03, RWS02, Rob00, Rue05, STWH00, ST05, TRSR00, WBSG07, Wil03, WSKH08, PCFC01]. Analíticas [TCSD00]. Analizar [CB01b, LRG^{+}02]. Analyses [Bal00, BMP^{+}06, JGL^{+}02, NLCAR09, WGGS07]. Analysis [ABE^{+}02b, BBB08, BFMM02, Bro00, CTR02, Che09, Cri00, CV08, DC05, EVD01, EFLW02, Fai08, FP02, GRG^{+}05, Har06a, HLSZ02, HPG09, KS05, KGM01, LVBS^{+}04, MP09, MBQAM06, MPDT01, MAP03, MM04, MDW01, MWEB06, NLCAR09, PGP04, PPBW08, Pro04, QAMW03, RT00, RMD^{+}02, RWJRR06, RWS02, Rob00, Rue05, SK04, SH03, STWH00, SHM04, SLCW08, ST05, Sta01, TFG08, TVB^{+}06, TWGH00, TRSR00, TK06, TSKD08, VW05, VMM07, VTA02, WBSG07, WSKH08]. Analytical [TCSD00]. Analyze [CB01b]. Analyzing [LRG^{+}02]. Anathemas [Dal09]. Ancestral [MTB00, Mef09a]. Ancestrally [JRL03]. Ancestralmente [JRL03]. Anchas [BMBM06]. Anchos [WE07]. Ancient [MTB00]. Andean [KFNB09]. Andes [ABE^{+}02b, KFSR^{+}06]. Andino [KFNB09]. Andrews [PJ00]. Anfibio [MSSM09, SM07a]. Anfibios [AE01, ACC^{+}05, BBB08, BR01, BFMM02, BBO^{+}01, BS01b, BSM08, BS00, CHW^{+}01, CF01, CMCG07, DSJ02, FPSD09, GB09, GWS^{+}06, GW05, GSB04, GP08, GH02, HRS08, JLZW03, KBM01, LRW03, MT01, MSBGP03, MHC^{+}01, OMP^{+}05, PC02, PC08, Pop08, RGF^{+}04, RBK^{+}05, RB06, RS02c, RK05, SGG07, SF03, Sem02, SB03, SS05, SYWR03, SKBB00, Sta01, SKP00, YLR^{+}01]. Angles [BBK02]. Anguila [CTNR02]. Ángulos [BBK02]. Anidación [DD04b, FCG^{+}05, FN02a, GPC^{+}03, MM02b, Rem03, Wel06]. Anidamiento [KSP00]. Anidan [Koe03, SP00]. Animal [AGBP09, Bat04, BTGS05, Cri00, Cze05, ED07, Hut08, MT08, PP08b, PP08a, Pyk00, Qia07, RS01, Rua00, SHM04]. Animal-Distribution [Cri00]. Animales [Bat04, Bau05, BST^{+}05, CH01, ED07, GBE00, MT08, PP08a, Pyk00, Qia07, RSW05, Rua00, SHM04]. Animals [Bat04, Bau05, BST^{+}05, CH01, GBE00, RSW05]. Anne [APS07, APS07]. Annealing [CA05]. Anniversary [Mef06a, SGLB^{+}07]. Announcement [Ano03d]. Annual [Ano03a, Büs08, HH03, HH06b, Pod01]. Anolis [KGS^{+}07]. Años [CBG^{+}00, GV02, LLV^{+}02, New06, PDS^{+}07, SPM05, SSY06, TNTH^{+}04]. Answers [Gos07]. Ant [ARP05, DBL06, DOGN08, HHSC00, SMM09]. Antarctica [ANW^{+}08]. Antártida [ANW^{+}08]. ante [RBT08]. Antelope [FSS^{+}00]. Anteproyecto [GBH^{+}05]. Anteriormente [PN05, SM09]. Antes [APCS05, PDF^{+}00, EM03]. Anthropochorous [GA02]. Anthropogenic [BHB08, FES01, FBL^{+}09b, HSMK07, Kes01, Lus03, MMF^{+}05, PRMV06]. Anthropogenically [VTR^{+}09]. Anthropological [WB06]. Anthropology [Bro06]. Anticipating [Ada08]. antillarum [PMKDS01]. Antílope [FSS^{+}00]. Antimonopolio [TAJ04]. Antitrust [TAJ04]. Antropogénica [FES01, MMF^{+}05]. Antropogénicamente [VTR^{+}09]. Antropogénicas [HSMK07, Kes01]. Antropogénico [BHB08, FBL^{+}09b]. Antropogénicos [Lus03, PRMV06]. Ants [PV02, PAA^{+}08, RCP00, TS03]. Anual [HH06b, Pod01]. Anuales [HH03]. Anuran [Roy04]. Anuros [Roy04]. Anyway [HTU^{+}06]. Apacentamiento [Mar05, PM05]. Apalaches [Car07, Has00, SRH^{+}07, SKP00]. Apareamiento [EBAW01, FLQ03, LM05]. Aphanomyces [KFK^{+}09]. Aphelocoma [DW05a]. Aplicación [Cor02, KZSV05, Lin03a, RMSG^{+}03, SSB^{+}08, Slo02, Esl00, FBK00, Gil01b, Kot00, PBN^{+}04, RWJRR06, SC00, Wan04]. Aplicaciones [FN02a, WM01, BR01]. Aplicado [EGN^{+}05]. Apocalypse [Mar00b]. Apoyo [For02, WR01, Wed02]. Appalachian [Has00, SRH^{+}07, SKP00]. Appalachians [Car07]. appendiculata [CNC04]. Appendix [GB05b]. Applicability [BL07]. Application [BR01, Cor02, FBK00, GHTMR01, Gil01b, Kot00, Lin03a, RMSG^{+}03, SC00, SSB^{+}08, Slo02, Wan04]. Applications [FN02a, PMEC04, RLK^{+}01, WM01]. Applied [EGN^{+}05]. Applying [Esl00, Hal02, KZSV05, MBH07, PBN^{+}04, RWJRR06, WP03]. Appraisal [CMC^{+}00]. Appreciating [FLF04, Sch05]. Approach [AKM07, BKWB05, BRV06, Bra05b, CM06, Cor06, Cua00, Cze02b, GBC^{+}04, GB05a, HCRM09, HM02, HBKF05, HCZ00, HDE00, JLF09, JHHL04, JAH^{+}05, LW06, LTC^{+}08, LMS^{+}02, MDCK07, MC09, MH01a, MPR^{+}01, MPE^{+}07, Old03, OIK^{+}06, PRL03, RMSG^{+}03, RAG03, Rot01, SL00a, TMR^{+}07, VMV07, Wal04, WFP01, WM05, ZR00, ZG05]. Approaches [ARMH04, Ano04d, BPV09, BBC^{+}04, Cur07, DBB^{+}09b, HSDS01, JMH01, Les05, RT00, SLC^{+}04, SMSB05, TCSD00, TWGH00]. Approximate [NS09]. Approximating [MW07]. Aprendidas [WL04]. Aprendizaje [CW08]. April [Ano06i, Ano07q, Ano08o, Ano09l, Ano05h, Ano06g, Ano07n, Ano08m, Ano09j]. Aproximación [PRL03, RAG03, WFP01, SL00a]. Aproximaciones [HSDS01, JMH01]. Aproximada [NS09]. Aproximarse [MW07]. Aptitud [Pyk05, RSB05, GB09]. Aqua [PFL01]. Aquaculture [Fra09, GL03, VTRG00]. Aquaculture-Escaped [VTRG00]. Aquarium [KB03, TH03]. Aquariums [MCR^{+}04]. Aquatic [BTK08, HH08, LRO^{+}08, ML04, RO08, RBT08, RWBG06, Sem02, TF00]. Aquatic-Breeding [Sem02]. Aranae [DCD^{+}02]. Araneae [DCD^{+}02]. Araucarian [BCS00]. Araucario [BCS00]. Arbol [CQLF02, MCA^{+}00, PS00a, SHM00, BFB^{+}09, FJ07, FLQ03, HDQSL06, KFBG08, MRP07]. Arboladas [GG02]. Arbolados [Cun00]. Arboles [CBH01, GGGS00, MR00a, WLO^{+}00, APSB02, CH01, CLC^{+}08, GRB02, GGA01, GLF^{+}08, HN09, HKM06, LVBS^{+}04, MM07, NH09, NAC^{+}06, NFB05, PAA^{+}08, RWJRR06, SH09]. Arbórea [WBBA03]. Arbóreas [GFCP03, KHP04]. Arborícola [PGP04]. Arborícolas [BOJ09]. Arbustiva [VDP00]. Arbustivas [KF01a]. Arbusto [BBK02, EBAW01, MBQAM06, ME00, YE07]. Arbustos [CBH01, SK02b]. Archetypal [Woo01]. Archipelago [Che09, DGSS03, APS07, SBK^{+}06]. Archipiélago [DGSS03, APS07, Che09]. Archives [Sta01]. Archivos [Sta01]. Arcoiris [MBQ05]. Arctic [BHW01, FES01]. Ardilla [BV04]. Ardillas [RMN00]. Area [AKM07, BBBM02, BCP^{+}02, Che09, DGC05, Fat06, FRSG^{+}02, For00a, GM08a, HDK^{+}00, JFS08, Kri04, LSZ^{+}08, McK02, MK02a, MP07b, NHHG^{+}05, NNM00, Ova03, PSBG05, PM07, RRK02, RB06, SJF^{+}05, SH02, SD03, TB01, VT09, Vee00, Wal09, WMP05, WSBP09, Wol00, YKIO08, KTN^{+}09, Che09, KTN^{+}09, WMP05, AKM07, BBBM02, BCP^{+}02, DMDFV08, Fat06, FRSG^{+}02, GM08a, JFS08, LSZ^{+}08, MP07b, PSBG05, PSGA09, RB06, SH02, VT09, WSBP09]. Area-Based [AKM07]. Areas [ALM07, BPV09, BMBM06, BSKH^{+}06, BDR04b, Bro04, BPL^{+}01, CRD05, CWC00, CNPS04, DK06, DHZM09, FVS03, FDL^{+}05, FMT^{+}04, FRR07, FNS^{+}03, GMMPP08, GCF^{+}07, GHL08, HDK^{+}00, Inf01, IIG05, Kal05, KCK^{+}08, KM00, KC07, KCR^{+}08, Lai00, LBR07, LSN07, LB01, LARK01, Lyn06, Lyn08, MB00, MP09, Mas03, MC09, MPR^{+}01, MA08, MBB^{+}03, PH02, PA04, PMC01, Pol06, RBRB06, RBP01, RBHKK02, RCMV06, RLXL01, RB05, SMV02, Sil05, SBB^{+}08, Spe02, SD01, TVD03, Tur06, TKB^{+}06, UKC^{+}09, VTR^{+}09, VTA02, WGW01, WB06, WMD^{+}06, XM07, ZG05, ZGR02, Dal05, GC09, LGB05, MC09, MN03, Dal05, DK06, GC09, RB05, SMV02, AW01, ALM07, BPV09, BMBM06, BSKH^{+}06, BMM^{+}02, BPL^{+}01, CRD05, CWC00, CNPS04, CS07b, DHZM09, FVS03, FDL^{+}05, FMT^{+}04, FRR07, FNS^{+}03]. Áreas [GMMPP08, GCF^{+}07, GSR^{+}04, GBC^{+}04, GHL08, JFS08, Kal05, KCK^{+}08, KM00, KCR^{+}08, Kom03, Kri04, Lai00, LBR07, LSN07, Lyn06, MP09, MC09, McK02, MYFF^{+}07, MK02a, MA08, MBB^{+}03, PH02, Pol06, RRK02, RBRB06, RBP01, RBHKK02, RCMV06, SJF^{+}05, SBB^{+}08, Spe02, SD01, Tur06, TKB^{+}06, UKC^{+}09, VTR^{+}09, VTA02, Wal09, XM07, ZG05, ZGR02]. Argentina [HHSC00]. Argentine [HHSC00]. Arguing [NB00, SNM00]. Argument [Gro05a]. Argumentos [DN01]. Arguments [DN01, Fie04]. Arid [ACMA04, BMB^{+}08, Cur02b, VMBC02]. Árida [BMB^{+}08]. Áridas [ACMA04]. Áridos [VMBC02]. Arizona [KBR03, LSC07, MBQAM06, WSKH08]. Ark [Bow09, Web04]. Armageddon [Joh05b, Orr05a]. Armiños [Cho06]. Army [RCP00]. Arnica [LDG^{+}00]. Arrecifal [FDL^{+}05, UWB04]. Arrecifales [WBSG07]. Arrecife [DAKB08, KB03]. Arrecifes [BMC^{+}09, BMP07, CMM05a, DDN01, FMP^{+}05, GWJ^{+}07, GGC03, HR04b, MCG^{+}09, MA08, MBB^{+}08, SCD^{+}08, TH03, WS03]. Arrendamiento [BMRB04]. Arriba [AAO^{+}09, EK08, FRD^{+}09]. Arriba-Abajo [AAO^{+}09]. Arrieras [RCP00]. Arrives [Doa01]. Arrogance [Web04]. Arroyo [MO04, WD03]. Arroyos [CS07b, HLG03, JSWS00, LB02, MY02, MK02b, NR03, PJ00, PW01, RBK^{+}05, Tow03]. Arrozales [Elp00]. Arruinará [MF00]. Art [Whe03]. Artemisa [HWH^{+}02, WRW^{+}02]. Artes [KSH00]. Artesanal [HR04b, Rut01]. Artesanales [CDM09, DLCV00, GEJK05, HAG^{+}09]. Arthropod [BTM^{+}05, BMA^{+}08, BRA^{+}03b, CCC04, DCD^{+}02]. Arthropods [JP09]. article [NL03b]. Ártico [FES01]. Articulación [CAR05]. Articulation [CAR05]. Artifact [Bal00]. Artifically [LW02a]. Artificial [BEM^{+}04, Faa04, HSB01, KWvH08, LCZ^{+}05, NR03, TB04]. Artificiales [BEM^{+}04, HSB01, KWvH08, LCZ^{+}05, NR03, TB04]. Artificialmente [LW02a]. Artisanal [CDM09, DLCV00, GEJK05, HR04b, HAG^{+}09, Rut01]. Artists [Bag09, Tol07]. Artrópodos [BTM^{+}05, BMA^{+}08, BRA^{+}03b, CCC04, DCD^{+}02, JP09]. Arts [JMM07]. Ascárido [Log03]. Ascendance [WDL^{+}02]. Ash [LO06]. Asia [BH04, CSS00, HPK^{+}04, MKVZ^{+}09, PCD^{+}03, SL00b, SCS06]. Asiatic [SRR00, SRW06]. Asiáticos [SRR00]. Aside [Her09, KB05, VW04, VW05]. Asignación [MBL02, MDH^{+}07, JMP09]. Asignar [MMBP08]. Asilvestrados [NMT^{+}04]. Asimetría [Ber05, DAR08, LVM02]. Asinara [VTA02]. Asistencia [Nel09, Wil08b]. Asistida [MH08]. Ask [BBA^{+}02]. Asnos [SRR00]. Asociación [FH09]. Asociada [BGWE00, MSWF00]. Asociadas [GG02, MNB09, MML06, OK01a]. Asociados [BRA08, CML02, HRS08, MS08, WSKH08]. Asociar [SGG07]. asper [LG05a]. Ass [SRR00, SRW06]. Assemblage [ARP05, LLH08, LHPH06, MSBGP03, MMF^{+}05, PP01]. Assemblage-Level [MMF^{+}05]. Assemblages [BMA^{+}08, CNM01, FN02a, GFTB04, SOADMR08, VS09, WA05]. Assembling [Pre04]. Assess [But03, RGGS02, Sch00, TS06, Wal09]. Assessed [OBK^{+}04]. Assessing [APCS05, AM05, Asc02, CJ08, FRSG^{+}02, HGH^{+}00, JSHM06, JAH^{+}05, MK02b, NKW09, Ols00, PBS03, Pop08, PK05, Pyk05, RO08, REC^{+}00, SM09, TG04, TWS06, WTEK08, WSCP05, WMD^{+}06, WSBP09]. Assessment [ACC^{+}05, ACMA04, ARA^{+}09b, AHR^{+}09, Baw05, BD05, CLBC03, DDAK04, DD01, EBAW01, FRH04, FM09, FMP^{+}05, FRR07, GM08a, Gre05, HCE^{+}01, JSJV03, JHHL04, KDL07, KSP00, KDKS08, LMG06, LRS09a, LARR03, MDF05, MP09, MCM^{+}00, MH08, Nor08, NCVBW02, OGH04, PPLS01, PLBBJC09, PCFC01, PA06b, Ros02, SHM04, SFT00, SCM^{+}07, The03, TKB^{+}06, WFP01, WWR^{+}02, ZSG01]. Assessments [Ant08, DAB06, FBCG00, KDC^{+}06, MDCK07]. Assignment [MBL02]. Assisted [BGWE00, Hun07b, MHS07, MH08, SAS^{+}09]. Associated [BGWE00, BRA08, GG02, HRS08, MNB09, MS08, MML06, OK01a, WSKH08]. Association [FH09]. Assumptions [FB08, Pow06]. astaci [KFK^{+}09]. Astronaut [RLK^{+}01]. Astronautas [RLK^{+}01]. Asymmetry [Ber05, DAR08, LVM02]. Atendiendo [BC08]. Atlantic [CT01, Chi00a, Dav07, Gal01a, Gal01b, HLG03, MSBOF^{+}09, RSD03, REC^{+}00, SAA^{+}01, SK08, TPS^{+}05, VTRG00]. Atlantica [Ibi08]. Atlántico [Chi00a, Dav07, HLG03, MSBOF^{+}09, RSD03, REC^{+}00, SAA^{+}01, TPS^{+}05, VTRG00]. Atmósfera [WCP01]. Atmosphere [WCP01]. Atracción [WS04]. Atravesar [BV04]. Atributos [BNH02, CVS^{+}09, DW05b, GC06, LLF08, MKK^{+}09, MHM09, LRW03]. atricapilla [Rem03]. Attaining [DS08]. Attending [CHMP05]. Attitude [LW06]. Attitudes [AJF03, JSJV03, NTGT03]. Attitudinal [GEJK05]. Attraction [WS04]. Attributes [BNH02, GC06, MHM09]. Atún [WELCH09]. Atwood [Orr05b]. Auctions [JLF09]. Audiences [Nad04]. August [Ano06b, Ano07d, Ano09b, Ano06b, Ano07d, Ano07o, Ano08c, Ano08n, Ano09b, Ano09k]. Aumentada [GLMN00]. Ausencia [JFWP06, Pol06]. Austral [CVB^{+}09, RSPM05]. Australia [BIK^{+}06, Cra02, FDBM04, Hob01, KPH07, MK02a, Pat00, PBK09, PK05, RWJRR06, RLK^{+}01, RSW05, VS09, WFP01]. Australian [BS09, BOC00, CB01a, CT07, CW02, FH09, Gil01a, GLV05, GN00, PBS03, RS02a, VS09, Whe02]. Australiana [CW02, GN00]. Australianas [BS09, BOC00, FH09, PBS03, RS02a]. Australianos [CB01a, CT07]. australiensis [HGH^{+}03]. australis [SB04]. Authentic [Ehr09]. Author [Ano03b, Ano08b, Ano09a, Ano04a, Ano05a, Ano06a, Ano07a]. Authorship [GS01, Har06a]. Auto [MRP07]. Autocompatibilidad [VJ07]. Autocorrelación [FP02]. Autocorrelacionada [KKHS05]. Autocorrelated [KKHS05]. Autocorrelation [FP02]. Autóctonos [Kal05]. Autodeterminación [SP09]. Autoincompatible [GPMM08, LDG^{+}00]. Autonomía [DS08]. Autonomy [DS08]. Autonomy-Supportive [DS08]. Autopista [FD00]. Autoría [GS01]. Autorías [Har06a]. Autoridad [BGHM06]. Availability [CMC^{+}01, ESSGL00, GPMM08, HSR05, MGB^{+}07, PJ00]. Ave [BB06, GNEE07, NS01, PV09a]. Averaging [WMVK03]. Aves [Álv02, AM05, ASD01, BNH02, BMM^{+}01, BB04, BHB08, BM05a, BS09, BJB^{+}02, BFD07, BGD09, Bor08, BV01a, BF03, Bra06, dLBBG^{+}08, BTS09, CWH07, CMP^{+}00, Cas03, CCC04, CKL^{+}00, CCM00c, CNM01, CSBS01, CLC^{+}08, DMSP00, DOGN08, Dav07, DS01, DJC^{+}07, DM04b, DWN01, Dun02, EGN^{+}05, EAA^{+}09, ERB01, EBAW01, Esl00, ES08, FJ00, FTC^{+}04b, GRB02, GGGS00, Gil01b, GM01, GPC^{+}03, GBC^{+}04, GBMH07, GG02, GV05, HP04, HP08, HRW^{+}03, Her09, HL07b, IMD01, JSJV03, JSHM06, JTZ^{+}08, KS05, KFSR^{+}06, KJSH08, Koe03, KBR03, KA03, Lai05, LMD05, Lat00, LF09, LSL04, LP08, LBPBG07, LFR^{+}08, LGK02, LCS04, MBR^{+}02, MG05, MBMLJ09, MTB00, MM07, MR00b, MGC01, MKDT01, New06, NL05, NFB05, OK01a, OHC08, PSBG05, PP03, PAA^{+}08, PMBN01, QRF07, RJW09, Ram01]. Aves [RM01, RSD03, RR01, RAR^{+}04, RS02b, Sch03b, SPM05, SSM03, SLB^{+}07, SSFL08, SBDD05, SMC00, SWD00, SP00, SK02b, SBSL06, SS01, SGH02, TB04, TNFH07, TKB^{+}06, TUE06, VCHBP08, VDP00, VMM07, Vee06, WMM04, WS04, War05, WFP01, WMS^{+}01, FR07]. Avian [BB04, CCC04, DWN01, GBMH07, Hen04, JSJ05, Lat00, LFR^{+}08, LGK02, PMBN01, Poo00, RM01, RR01, San00, SSM03, SSY06, SK02b, SGH02]. Aviar [JSJ05, SSY06]. Avifauna [BVK02, BCD^{+}08, CSS00, Rot01, SLKP06, TSKD07, WBBA03]. Avifaunal [BVK02, FJ00, SLKP06]. Avifaunas [Mal03, TSKD08]. Avis [Kir08].

Avispas [FBP^{+}09, KSDBT02, KT02, PAHM04, SD03]. Avistamiento [MRDC06]. Avoid [GBE00]. Avoidance [PPW09]. Avoided [VT09]. Avoiding [LHG^{+}03, Rai00]. AVP [CB01b, MBQAM06]. Avutarda [GBH^{+}09]. Awaited [Doa01]. Awareness [PKM00]. Awful [Ran00]. Ayuda [LV04]. Ayudar [BNP^{+}07]. Azul [SH03]. Azuleja [MSWF00]. Azules [GLMN00].

B [BBB08, BOC00, DSJ02, PRL^{+}02, SWR00, SKP00, WG06]. B-UV [PRL^{+}02]. B-UV/pH [PRL^{+}02]. B/pH [PRL^{+}02]. Babirusa [CMGSM00]. Baby [BBA^{+}02]. Back [Bre02a, Gru04, McC03, PV09a]. Backed [GK06, TM04]. Background [TR01]. Bacterial [RIBGG08]. Bacterias [RIBGG08]. Bad [Bat04, Bel02, Bru02, BV01b, Ghi06, Say09]. Badgers [RPD01]. Baggage [Orr09a]. Bahía [SAA^{+}01, VDHD01]. Bahian [SAA^{+}01]. Baiji [Kle06, RG06, WZW^{+}06, YBWZ06]. Baja [BBBM02, FKU05, GPMM08, MO04]. Bajas [GAF00, Lam04, Mcs05, STE^{+}01, WG01]. Bajo [ADF04, CH01, DP01, FPL^{+}04, GGGS00, KKHS05, KLH08, LB01, ML04, RS01, VVMV09, WCPB03, WRL02, AFB^{+}00, NB06, RM01, SDRO01]. Bajos [MML06, PZG^{+}05, LDG^{+}00, Wol00]. Balaena [SDRO01]. Balance [Hod04a, Kle03, KCK^{+}08, MP08]. Balancear [MP08]. Balancing [Wil08b]. Báldi [VW05]. Baleares [TR05]. Balearic [TR05]. Ballena [REC^{+}00, SDRO01]. Ballenas [VT09, WMPC08]. Ballistics [Cob03]. Balmford [SBS^{+}09, War00]. Bancos [HGH^{+}00, Rom03, NH09]. Bandicoot [SLCW08]. Banding [JGL^{+}02]. Banff [CW00, Cas03, CW00]. Bank [BW08a, GM00a]. Banking [FNM05, RH08]. Banks [Col09, NH09]. Banteng [BIK^{+}06]. Barbary [CMC^{+}01]. Barcodes [Rub06b]. Barcoding [Des06, Rub06a]. Barcos [VT09]. Bark [BFB^{+}09, CMC^{+}01]. Bark-Stripping [CMC^{+}01]. Barometers [GSB^{+}06]. Barómetros [GSB^{+}06]. Barras [Rub06b]. Barred [HMF^{+}04, PR03]. Barrera [FDL^{+}05, IML^{+}05, UWB04]. Barreras [FRD^{+}09, JMS^{+}06a, KTN^{+}09, MMGB05, MMB^{+}03, NR03]. Barrier [FDL^{+}05, Har03, IML^{+}05, UWB04]. Barriers [FRD^{+}09, FCN^{+}06, Gha01, JMS^{+}06a, JP08, MMGB05, MMB^{+}03, NR03]. Basada [AOGMS09, Ber04c, Cur02a, DS08, LMG06, SCB^{+}01]. Basadas [LWV04, SHM00]. Basado [AKM07, LMP02, NF04, SHJ^{+}09, TNTH^{+}04]. Basados [ARMH04, BPV09, CWCG02, CB01b, FNF03, GBE^{+}06]. Base [APS07, BMMW07, CMA^{+}06, FNS^{+}03, GC06, HR04a, KDL07, MRAF^{+}07, PLBBJC09, PMC01, PK05, QRF07, RB06, SF05, STS05, SCM^{+}07, TVB^{+}06, WBSG07, WMP05]. Based [APS07, AKM07, BPV09, BS01a, Ber04c, BMMW07, BAKHT06, CPSS03, CMA^{+}06, CWCG02, CB01b, Cur02a, Cur07, DS08, Fai08, FNF03, FNS^{+}03, GL03, GC06, GBE^{+}06, HR04a, HR04c, JH05, Kar09, LW06, Lea06, LMG06, LMP02, LWV04, MSL06, MRAF^{+}07, NF04, PMC01, QR06, QRF07, RB06, SCB^{+}01, SHJ^{+}09, Sem02, SF05, STS05, SCM^{+}07, TVB^{+}06, WBSG07, WMP05, AOGMS09]. Baseline [SCD^{+}08, Wil02b]. Bases [HCE^{+}01, HLJV07, MD04, Sem02]. Basic [HC09, Mag01, PGD^{+}07]. Básica [HC09]. Basics [Par08]. Basin [BRV06, FMT^{+}04, HWH^{+}02, UMT^{+}01, WRW^{+}02, WWR^{+}02, BGD09, BRYK01, DMDFV08, FKU05, FNFM01, FN02a, MK02b, PL03, WSR^{+}00]. Basis [KDL07, PLBBJC09]. Bat [CPR05, GGGS00, Hol08, KWV08a, MEA00, PMZ06, PMEH05, WHJJ04]. Bat- [GGGS00]. Bath [BBA^{+}02]. Batrachochytrium [BRS^{+}05, KPH07, MSSM09]. Bats [AW02, KWvH08, SK04, TRCC^{+}05]. Battle [Hod04b]. Baulas [TSP^{+}08]. Bay [AJF03, ZKNL02, VDHD01, ZKNL02]. Bayesian [BOJ09, BBC^{+}04, CJ08, MBL02, PPBW08, Wad00, WSBP09, WMVK03]. Bayesiana [MBL02]. Bayesiano [PPBW08, WMVK03]. Bayesianos [BOJ09, CJ08, Wad00, WSBP09]. Be [DP01, Law06, PBC^{+}09, Ste08b, VT09, ALH^{+}04, MV06]. Beach [BBM04]. Bear [Côt05, DN01, DOW^{+}06]. Bearing [DDMMZD06]. Bears [BST^{+}05, MM02a, NWRD03]. Beaufort [Tay03b]. Becoming [KL05, Swa09]. Beds [Rom03]. Bee [ABE01, WGK07]. Beef [NSA06]. Bees [FMPB07, KSDBT02, KT02, PZG^{+}05, SD03]. Beetle [BRA08, DW05b, FH05, LMP02, PMEH05]. Beetles [Fat06, HW08, HKN01, HKM06, LLF08, MMLM01, TK07]. before [APCS05, MS02]. Beginning [Fle08]. Behavior [Bek04, BBK02, BBM04, BFJ04, CMC^{+}01, DOGN08, GVTM07, Jac09, Lus03, PPW09, SFD09, SS02a, WBW06, WHM01]. Behavioral [Bra03a, FB08, JSJV03, MS04, WBW05, ZSSM09]. Behaviorally [GSG08]. Behind [NMB09, Joh09b, Kun03, Tas09]. Beier [HRN00]. Being [ADD^{+}07, Bei05, CJ08, SBM06]. Beissinger [Ris02]. Believe [Ebe04]. Believers [Win04]. Bellbirds [PB04]. Bellflower [EKL^{+}00]. Belts [Ric04a]. Ben [Sha01]. Beneficial [NT01]. Beneficio [AHTM05, AKM07, BSM08, Cla01]. Beneficios [KFL08, MP08, PKB^{+}08, PBRG07, WTEK08, BTM^{+}05, BSKH^{+}06, CMM05a, LF09, RBP01]. Benéficos [NT01]. Benefit [AHTM05, AKM07, GALW09, Tho00]. Benefit-Function [AKM07]. Benefits [BTM^{+}05, BSKH^{+}06, CHMP05, CMM05a, Hol08, KFL08, KWV08a, LF09, MP08, PKB^{+}08, PBRG07, RBP01, WTEK08, WLS^{+}08]. Benjuí [GFCP03]. Benthic [KSH00]. Béntico [KSH00]. Benzoin [GFCP03]. Beres [MSB^{+}01]. Berg [ML00]. Berger [OS04]. Bering [Tay03b]. Best [Bra05b, KK08, KDC^{+}06, Pre04]. Beta [WU05, WMP05]. Better [Caf06, Cla09, GGT08, Had07, JGM03b, RT00, WMVK03]. Between [GT03, Van05, AS01, BRTW07, BGHM06, BGHMb07, Ben02, dLBBG^{+}08, Bro06, CGR01, Coc01, CW08, Cze08, DHZM09, Elp00, HPF^{+}06, HBLW09, Jam02, JP09, KCK^{+}08, KK09, LCZ^{+}05, LM07, LS03, Lyn08, Mad05, MNRD^{+}08, MKK^{+}09, NFM04, PFL01, PP08a, Qia07, RF03, Ree05, RSCL04, SMG00, SVBR07, SMW^{+}07, WCP01, WG06, ZSMW06]. Beyond [Ada06, Caf00b, Dal05, DWWF04, Fle01, HRP^{+}04, Joh05a, NTMT^{+}03, Nos01a, RH01, SDJK04, ZSVL07, Gui07, Urb06a]. Bhutan [NSW08]. Bias [AW01, Bro00, TB04]. Biased [MBQ05]. Biases [MF00, MPFBH07, SGG07]. Bien [Kot00, MKH07]. Bienestar [CPR^{+}07]. Bifurcaciones [BMS08]. Big [GB05b, JP08, Kar01, LC07b, VB06]. Big-Leaf [GB05b]. Bighorn [ODE01, RBG06, SZS01]. Bignoniaceae [GZK06]. Bignonieae [SLCJ08]. Bilateral [DAR08]. Billed [Gra01]. Billion [Koh09]. Bio [Fle04b, Mei03]. Biocombustible [DBB^{+}09a, GGT08]. Biocombustibles [GGT08]. Biocontrol [GLV05]. Biodiesel [Koh07]. Biodiversidad [ATG05, AJ05, AW01, ALM07, AHTM05, BHMC01, BGHM06, BADF^{+}05, BMV09, BMM^{+}02, BWS09, BBW^{+}07, BV00, BTS09, CS05, CPD^{+}03, CY08, CS07b, CGR^{+}05, CAR05, Cze08, DBB^{+}09a, ESB05, Fat06, FB00, Fon09, GRJB^{+}07, GCV03, Gib00, GBC^{+}05, GGT08, GM09, HBD^{+}09, HJP05, HBKF05, HM03, HPG09, HLJV07, HW02, JGM03a, JB06, KDD^{+}04, KC01, KSP00, KWL^{+}09, KCK^{+}08, Law08, LCTL05, LOM03, LMB00, LN06, LO06, LSZ^{+}08, MB07, MKG03, MFGB07, MLN^{+}06, MDF05, MLMB08, MOO09, Mcs05, MN03, MK02a, MLB08, MGH^{+}09, MC05a, MYH09, MNH01, NH09, OGH04, OB00, PA06a, PCL08, PAA^{+}08, PCFC01, PB04, PS04, Pyk00, QR05, Rau02, RDEF01, RLC^{+}07, Rob09, Ros06, Ros08, SCB^{+}01, SSS^{+}08, SHJ^{+}09, SP09, Spe02, SD01, SGH02, TPS^{+}05, TG05, TKB^{+}06, VDHD01, VW04]. Biodiversidad [VIMMO05, WPS08, WDA^{+}06, WD01, XGS^{+}09, ZR00, GHP^{+}05, OK00]. Biodiversity [Abe02a, ATG05, AJ05, Ano01j, Ara02, ALM07, AHTM05, Bal01, BFB^{+}09, BHMC01, BGHM06, BGHMb07, BW08a, BKWB05, BADF^{+}05, BMV09, BMM^{+}02, BWS09, BBW^{+}07, BL03, BNM^{+}09, BV00, BTS09, BAKHT06, CSS00, CS05, CPD^{+}03, CY08, CS07b, CGR^{+}05, CAR05, Cze03, Cze08, DBB^{+}09a, Don04, ESB05, Fai02, Fat06, FB00, Fon09, GCV03, Gha01, Gib00, Gil01a, GHP^{+}05, GBC^{+}05, Gri00, GGT08, GM09, Gro05b, GM08b, HML^{+}02, HBD^{+}09, HJP05, HKC^{+}08, HBKF05, HM03, HPG09, HLJV07, HW02, JGM03a, JB06, Jor01, KDD^{+}04, KC01, KSP00, KWL^{+}09, KB05, KCK^{+}08, Koh07, KK09, LMR^{+}07, Law08, LCTL05, LOM03, LS07, LMB00, LN06, LO06, LSZ^{+}08, MB07, MKG03, MFGB07, MLN^{+}06, MDF05, MLMB08, MOO09, McS03, Mcs05, Mef08, MN03, MK02a, MLB08, MGH^{+}09, MC05a, MDRB05, MYH09]. Biodiversity [Mor00a, MNH01, NH09, NSW08, NPT05, NC03b, OGH04, Ode09, ODPW02, OK00, OB00, PA06a, PDLS08, PAB^{+}08, PJB^{+}06, PCL08, PAA^{+}08, PCFC01, Poo00, PB04, Pre04, PS04, Put07, Pyk00, QR05, QR06, RDEF01, RLC^{+}07, Rob09, Ros06, Ros08, Rud03, SCB^{+}01, SSS^{+}08, SBM06, SHJ^{+}09, SP09, Sch06, She01, SSK^{+}05, SMW^{+}07, SBK^{+}06, Sod06, Spe00, Spe02, Ste06a, SS02b, SD01, SGH02, TPS^{+}05, TG05, TKB^{+}06, VP06, VDHD01, VW04, VW05, VJ09, VIMMO05, WPS08, WDA^{+}06, WK09, WD01, WP03, WD09, WMD^{+}06, XGS^{+}09, You01, ZR00, Rau02]. Bioeconomic [Cho06, LMG06]. Bioeconómico [Cho06]. Bioeconómicos [LMG06]. Biofísicas [PMC^{+}07]. Biofísicos [HR02]. Biofuel [DBB^{+}09a, GGT08]. Biofuels [GGT08]. Biogeografía [Dal05, LV04, CML02, DCP^{+}03, HDDEB03, RDEF01].

Biogeográfica [RTCR03, SSB^{+}08]. Biogeográficas [Spe02]. Biogeográficos [SGG07]. Biogeographic [SGG07, SSB^{+}08, Spe02]. Biogeographical [RTCR03]. Biogeography [CML02, DCP^{+}03, Dal05, HDDEB03, LV04, RDEF01]. Bioko [FYC00]. Biologí [Rod03]. Biología [BR05, CES05, CP00, Cla01, Ehr00, FBC^{+}06, HH06a, LG02, LSOBA04, PRL03, RH01, SAM07, Sha01, Sim03, SM01, TS03, YL05, HS06, MKVZ^{+}09]. Biológica [Ang00, Asq01, CPR^{+}07, FBK00, FNS^{+}03, GS01, HM08, HC09, KLH08, LLS00, MCK06, Sag07, SVV07, SSY06, SAA^{+}09, TMR^{+}07, TAP^{+}01, Wad00]. Biological [Ang00, BCMA^{+}01, BEB00, BMB^{+}08, Car04, CPR^{+}07, Dav07, FNS^{+}03, GR03, Gay02, GS01, HM08, HC09, Hod04b, LLS00, LH04, LD03, LSF04, LO02, LARR03, LS04, MCK06, NS09, Old03, Pem01, PL06, Pro04, Ric04a, Ric07, Sag07, SVV07, SB03, SSY06, SAA^{+}09, TMR^{+}07, WABT07a, WSS^{+}04, WWH03, Wil02b]. Biological-Economic [HM08]. Biologically [SS08a, Sem02]. Biológicamente [Dav07, SS08a]. Biológicas [BCMA^{+}01, LD03, McC01, Old03, Ric07, Sem02]. Biológico [BEB00, BMB^{+}08, LH04, LO02, LARR03, WABT07a]. Biológicos [GR03, NS09, SB03, WSS^{+}04]. Biologist [Ben00, Ben01b, Bru02]. Biologists [Ano05c, Bea09b, Bei05, Fie00a, Flo05, PP08a, PR01b, SRL^{+}07, Sto04, Tho03]. Biólogos [PP08a]. Biology [Abe02a, AO06, AL03, Ano03a, Ano05k, Ano07e, Asq01, Awo08, BC06, Baw06, Bea09a, Bee00, BH01, BR05, BBM07, Bre06, Bro06, BT07b, CCM00a, CES05, CP00, Cha08, Cla01, Cob03, Du 01, Ehr00, FBC^{+}06, FBK00, Fre06, GW05, GOL^{+}07, Haf08, HH06a, HS06, IB03, Jes03, Kaw04, Ker09, KT06, Kni09, LG02, LK04, LSOBA04, Lov06, Mag01, Man02, MKVZ^{+}09, Mef01a, MEN06, Mef06b, MSN06, MGB^{+}07, Nas06, NL03a, Nor00, Nos07, ODPW02, OS01, PRL03, Pér05b, Pri03, RC04, RF00, RM03, RH01, Rod03, Sar06, SAM07, Sha01, Sim03, Smi04a, SL00b, SM01, Sto04, TSH06, Tem02, TOR^{+}04, TS03, TAP^{+}01, VS05a, Wad00, WHM01, Wit03, YL05, Ano05g, Ano06f, TOR^{+}04, Edi03, Har06a, Mef06a, Rob06b, SGLB^{+}07, Wil00b]. Bioma [UKC^{+}09]. Biomasa [HCRM09, GWS^{+}06]. Biomass [GWS^{+}06, HCRM09, Tal04]. Biome [UKC^{+}09]. Biophysical [Car02b, HR02, PMC^{+}07]. Biosafety [HS06]. Bioseguridad [HS06]. Biosfera [BTBC^{+}03, TVD03, FNS^{+}03]. Biosphere [FNS^{+}03, TVD03, BTBC^{+}03]. Biotic [Fox07, Hob01, IF02, ODD05, RWRW04]. Biótica [Fox07, IF02]. Bióticas [Hob01]. Biótico [RWRW04]. Bird [Álv02, ASD01, BNH02, BV01a, BF03, Bra06, dLBBG^{+}08, BTS09, CMP^{+}00, CS06, CCM00c, CNM01, DMSP00, DJC^{+}07, DM04b, Dun02, EBAW01, GGGS00, GPC^{+}03, GBC^{+}04, GNEE07, GG02, GV05, HP04, Her09, HL07b, JSJV03, KS05, KFSR^{+}06, Lai05, LME08, LF09, LBPBG07, MG05, MBMLJ09, MR00b, MGC01, New06, NL05, OHC08, Pat00, Pau08, PV09a, PP03, QRF07, RJW09, RSD03, RAR^{+}04, SPM05, SSFL08, SBDD05, SMC00, SBSL06, TNFH07, TKB^{+}06, TUE06, VDP00, VMM07, VS05b, WMM04, War05, WMS^{+}01]. Bird-Generated [GGGS00]. Bird-Plant [Pat00]. Bird-Pollinated [EBAW01]. Birds [BMM^{+}01, BS09, BJB^{+}02, BGD09, Bor08, CWH07, CKL^{+}00, CSBS01, CLC^{+}08, DOGN08, Dav07, DeS03, DS01, EGN^{+}05, EAA^{+}09, ERB01, EBAW01, Esl00, ES08, FTC^{+}04b, FR07, GRB02, Gil01b, GM01, HP08, HRW^{+}03, IMD01, KJSH08, Koe03, KBR03, KA03, LMD05, LF09, LSL04, LP08, LCS04, MBR^{+}02, MTB00, MM07, MKDT01, NFB05, PAA^{+}08, Rab09, Ram01, SLB^{+}07, SWD00, SP00, SS01, VCHBP08, Vee06, WFP01]. Birth [SGLB^{+}07]. Bison [LCC^{+}07, SRW^{+}08]. Bisonte [LCC^{+}07, SRW^{+}08]. Bivalve [RFD01]. Bivalvos [RFD01]. Black [Côt05, DOW^{+}06, MTB00, Put05, Shi06, TRSR00, WBDP02, Wie07]. Black-and-White [Wie07]. Black-Tailed [Shi06]. Blackburn [TSKD08]. Blackcap [Rem03]. Blanca [HPK^{+}04]. Blancas [Ols07]. Blanco [BEB00]. Blanqueado [GWJ^{+}07, WS03]. Blanqueamiento [RPT00]. Blast [FMP^{+}05]. Blazing [Vic02]. Bleaching [GMHS00, GWJ^{+}07, RPT00, WS03]. Blind [Wes05]. Blossey [HH09]. Blue [GLMN00, KJSH08, LC07b, SH03, KJSH08]. Bluefin [SK08]. Blueprint [GBH^{+}05]. Bluff [Koh09]. Boat [GCBF09]. Boaters [AJF03]. Boats [Lus03, RS02b]. Bobcats [RSF^{+}03]. Body [CB01a, DAR08, LP07, SFD09, SCP^{+}02]. Bog [ND03]. Boiga [PV09b]. Bold [Flo05]. Bolivia [FND06, STE^{+}01]. Bolivian [RW02, HGH^{+}00, Kes01, RS01]. Bolivianas [HGH^{+}00]. Boliviano [RW02]. Bolivianos [Kes01, RS01]. Book [Ano00a, Fie03b, Fle04a, O’C05, Sch02]. Books [Ano01a, Ano02a, Ano03c, Ano03f, Ano04i, Ano04k, Ano04h, Ano04j, Ano04g, Ano05b, Ano05j, Ano05i, Ano05h, Ano06b, Ano06j, Ano06h, Ano06i, Ano06g, Ano07d, Ano07o, Ano07r, Ano07p, Ano07q, Ano07n, Ano08c, Ano08n, Ano08q, Ano08p, Ano08o, Ano08m, Ano09b, Ano09k, Ano09m, Ano09l, Ano09j, Kri05]. Boom [GL03]. Boot [SB06]. Borda [RS02b]. Borde [BB04, CFC02, DD04b, HMB^{+}05, MMLM01, MJ03, OBM09, PSBG05, PRMV06, RPD01, SG01, TM04, WGGS07]. Borderea [Gar03a]. Borderlands [Cur02a]. Bordes [BBK02, BVH^{+}02, CP01, CM06, DMSP00, ED07, FTR01, SRH^{+}07, SLS05, WCP01]. Boreal [BHB08, BJ05, BOJ09, GJD05, HW08, HKN01, HBV02, HNG04, HKM06, IMD01, JMB01, JMS06b, Kom03, LSN07, Mar02a, MJ03, MM03a, MYH09, PBL00, RMN00, SSSK06, TK07]. Boreal-Forest [MYH09]. Boreales [BJ05, BOJ09, GSR^{+}04, HBV02, HNG04, HKM06, IMD01, JMS06b, Kom03, MM03a, MYH09, RMN00, SSSK06, TK07]. Boring [DSE^{+}08]. Born [TAP^{+}01]. Bornean [Lam04, WLCS00]. Borneo [Lam04, MN03, SLELA05, WLCS00]. Borrego [EMW^{+}04, SZS01]. Bos [BIK^{+}06]. Bosawas [SMT07]. Boscoso [EB09, SLKP06, WGK07]. Boscosos [BRK^{+}04, DBB^{+}09a, MM07]. Bosque [ASM05, ARPEBM09, BV04, BHB08, BVK02, BCS00, Bor08, BCP^{+}02, CFC02, CWL^{+}00, CMP^{+}00, CQLF02, CSS00, Chi00a, CH01, CCM00c, DMSP00, Dun04, EAA^{+}09, FTR01, FBL^{+}09b, FLQ03, Gav04, GV02, Gil01b, GG02, GV05, GJD05, HCBK00, HP04, HP08, HKN01, HDQSL06, HBV02, HAJ^{+}04, HN09, JHHL04, Kau04, KWvH08, LLV^{+}02, LCT^{+}06, LLH08, LMP00, LVBS^{+}04, LP07, MR00a, MB01, MJ03, MK09, NAC^{+}06, NFB05, NS01, OFS^{+}04, PJYO04, PB09, PMEH05, Ram01, RTN01, RH09, RSD03, RWS02, Ric04b, RCP00, RR01, RS02c, SAA^{+}01, SLELA05, SG01, SPM05, SWD00, SLS05, STE^{+}01, SBSL06, SOADMR08, TPS^{+}05, TO02, TUE06, WCP01, WLCS00, WSLW00, WZD^{+}00, Bei08, PJ00, Rig01]. Bosque-Pastura [KWvH08]. Bosques [AKO03, ALR05, ARA^{+}09b, BMBM06, BMM^{+}01, BD06, BJ05, BOJ09, BKWB05, Bor05, BWS09, BAF04, BEM^{+}04, Bur04, CP01, CPD^{+}09, CPME09, Coc01, CSM^{+}06, CAFL03, Cra02, Cun00, DWN01, EM03, FPM02, FH05, FO06, FR07, FLQ03, GRB02, GRJB^{+}07, GK06, GC06, GSR^{+}04, GPC^{+}03, GBMH07, GAF00, GH02, Gus02, HH00, HMB^{+}05, Has00, HVR08, HFR07, Hut06, HNG04, HKM06, IMD01, JMS06b, Kau04, Kes01, KFBG08, KD05, Kom03, KIAH08, KA03, Lam04, LMD05, LLF08, LW01, LCG^{+}08, LMB00, LO06, LGK02, MMGB05, MRP07, MRM^{+}05, MML06, MM03a, MYH09, MSBOF^{+}09, New06, NMB09, Olu00, PBP09, PBK09, Per00, Per01, PL03, PRMV06, RHS05, RGF^{+}04, RMN00, RRH^{+}09, RS01, RWRW04, Rud06, SSSK06, SMN00, SSM03, SRH^{+}07, SH09, STL02, SBB^{+}08, SHYH06, SCC^{+}09, SDJ06]. Bosques [SBVC03, TA08, TK07, WCP01, WD01, WBV^{+}05, BMPNC^{+}03]. Botanical [MCR^{+}04]. Botánico [PM07]. Botánicos [MCR^{+}04]. Botanist [PM07]. Botanists [LS03]. Botella [BR01, BJTD03]. Botrychium [Gun02]. Botswana [RLK^{+}01]. Bottleneck [BR01, BJTD03]. Bottlenose [BSW^{+}06, Lus03, Wad02]. Bottom [AAO^{+}09, EK08]. Bottom-Up [AAO^{+}09, EK08]. Bottrill [KC08]. Boundaries [Gui07, KTN^{+}09, NTMT^{+}03, PW01, RCMV06]. Boundary [DGC05, GDAV00, MW07, NHHG^{+}05]. Boundary-Quality [MW07]. Bowhead [SDRO01, SRDG03, Tay03b]. Brachyramphus [PBN^{+}04, Rap06]. Brachyrampus [CLBC03]. Brasil [LP05, ATG05, AJ05, CLMD05, CAR05, DS01, GHP^{+}05, LCTL05, LFP05, MG05, MRM^{+}05, MSBOF^{+}09, RC05, RSD03, SAA^{+}01, You05]. Brasileña [Fea05, ZSVL07]. Brasileñas [RB05]. Brasileño [BF03, Chi00a, KM05, TPS^{+}05, LF00, UMT^{+}01]. Brasileños [Rod05, SS05]. Brassica [GPMM08, WRL02]. Brazil [ATG05, AJ05, AR09, CT01, CLMD05, CAR05, DS01, GHP^{+}05, Ibi08, LCTL05, LFP05, LP05, Lov05, MG05, MRM^{+}05, MDRB05, MSBOF^{+}09, RC05, RSD03, SAA^{+}01, You05]. Brazilian [LF00, UMT^{+}01, BF03, BDRC05, Chi00a, Fea05, KM05, Rod05, RB05, Sil05, SS05, TPS^{+}05, ZSVL07]. Breaking [Sad05, SC09]. Brecha [KCR^{+}08]. Brechas [HPF^{+}06]. Bred [AAO^{+}07, FPF00, GBE00]. Breed [CMCG07, RMSG^{+}03]. Breeding [ASD01, BBO^{+}01, BRK^{+}03, Bow09, Cor03, CS07b, FPL^{+}04, For02, GP08, HRS08, HRW^{+}03, HL07b, KHM00, KBR03, LCS04, MR00b, MBQ05, MKDT01, ODE01, PC02, QRF07, RG04a, RK05, SM07a, SPM05, Sem02, SP00, SS03b, Tem02, WR01, Wed02, ZSSM09].

Breeding- [ASD01]. Breeds [RMSG^{+}03, Rua00, TVB^{+}06]. Breñal [QAMW03]. Bretaña [DSTPW07, KZA09]. Bridge [GT03, Mar04, Nos00]. Bridging [Jam02, Pae05]. Brief [MDRB05]. Brieuc [VDHD01, VDHD01]. Brighter [Mul07]. Brincando [Lat00]. Bringing [GG06, KK08]. Brisbane [RWJRR06]. Britain [DSTPW07, KZA09]. Británica [QR05, QRF07, WSS^{+}04]. British [VTRG00, WSS^{+}04, QR05, QRF07, VTRG00]. Broad [Arm05, Cha06, Fje07, FNFM01, PAHM04]. Broad-Scale [Arm05, Cha06, Fje07, FNFM01]. Broadleaved [BMBM06]. Broken [Dix07]. Brown [ASF^{+}09, BST^{+}05, NWRD03, PV09b, SLCW08, WBBA03, BRTW07]. Brummitt [MM03b]. Bubble [Hay04]. Buceo [PIH00]. Buena [LG02]. Buenos [Bat04]. Buffel [ACMA04]. Buffer [GRG^{+}05, HRS08, HNG04, RS02b, SB03, WD03]. Buffer-Strip [HNG04]. Buffer-Zone [RS02b]. Buffers [CS07b, FPSD09]. Bufo [BS09, BR01, SB08]. Búho [CJ08]. Búhos [SS03b, TG04]. Build [Haa09]. Building [FMPB07, Inf01, Joh05b, KNH07, MC05a, RSPM05, RRCOM^{+}06, Sad05]. Buildup [JMB01]. Bull [PD03a]. Bullfrogs [BSM08]. Bumblebee [WCX09]. Bunting [Wel06]. Burden [CGR01, Nos09]. Bureau [KB08]. Bureaucratic [RVNM08]. Bureaucratically [PPRB06]. Burials [SSG01]. Burn [NTMT^{+}03, Ram01]. Burned [Cze03, VSK01]. Burning [ASD01, Cze06, HLB^{+}03, Pan02, PA06a, RO04, SF03, TK07, VS09]. Buró [KB08]. Busca [LMP02]. Bush [BBK02, Ber05, RSW05]. Bush-Cricket [BBK02]. Bush-Crickets [Ber05]. Bushland [RAW07]. Bushmeat [BBB^{+}07, CMR05, Dav02, DC06b, FYC00, FPM02, LMG06, PCME09, TB03, WG01, WSA^{+}05]. Buskirk [KB05]. Búsqueda [SS02a]. Bustard [GBH^{+}09]. Butterflies [BKK03, FNFM01, FM09, HHD^{+}08, HDDEB03, KSB04, KT02, Lew01, NFM04, VVMV09, VMV07, WLCS00]. Butterfly [BE00, BTBC^{+}03, BCP^{+}02, BTS09, CWNW01, CPO03, FN02a, FNF03, Gut05, Had00, HJ05, MK09, RDW01, SH03, SC05, Wal04]. Buzos [GPZ04]. Buzz [NC03b]. Bwindi [HCBK00, OBM09, RIBGG08, HCBK00, OBM09, RIBGG08]. Bycatch [CLŽ^{+}07, DLCV00, LC07a, MBR^{+}02, RDN06, UCAGG08, WELCH09, WD09, ŽWGW09].

C. [RNL^{+}04]. Caatinga [LCTL05]. Cabalga [Kau04]. Caballitos [FV05, GPZ04]. Caballos [LEPH02]. Cabecera [NR03]. Cabildeo [GC09]. Cabo [GPZ04, WPS08, WHA^{+}05]. Cabras [CD05]. Cacao [BMV09]. Cacería [CWC00, ESSGL00, FJ07, GMW08, HMF03, LCT^{+}06, LCG^{+}08, LMG06, MNA07, NTMT^{+}03, NRB00, OSSK^{+}07, Per00, Per01, RS01, SK02a, SH09, SHM04, PBP09]. Cacti [TVD03]. Cactos [TVD03]. Cadena [CMR05]. Cadenas [Lus03]. Caducifolio [GG02]. Caecilian [GW05]. Café [ARP05, Hie05, KSDBT02, Mad05, MK09, PV02, PBRG07, PMEH05, Ric04b]. Cafetaleras [KJSH08, SMFPCA05]. Cafetaleros [PAA^{+}08, RCP00]. Cage [Fra09]. Calakmul [ESSGL00]. calamita [BR01, SB08]. Cálaos [SH09]. Calculating [SS08a]. Calculation [MS04]. Cálculo [MS04, SS08a]. Caledonia [PDLS08]. Calentamiento [CBH01, GK06, GD05, GB01, HJ05, MLN^{+}06]. Calidad [BCP^{+}02, CMC^{+}01, EKL^{+}00, FRSG^{+}02, JSHM06, LP08, MW07, PV09a, SB08, VO03, PV02]. Caliente [WMLS09]. California [Bei01, BS01a, CKL^{+}00, DSJ02, Dav05, EMW^{+}04, GSG08, HH03, HM04, JC07, KF01a, Lei09, LM07, MSB^{+}00, MSB^{+}01, MB09b, OFS^{+}04, Pyk05, RPM09, RSF^{+}03, RBK^{+}05, Ris02, RBHKK02, Rub01, SS08a, TG04, UCAGG08, ZBABR00]. californicus [TM04]. californiense [Pyk05]. Caliza [BKK03]. Call [Bre02a, Mar02b]. Calles [FJ00]. Callicott [Wil00a]. Calling [GB01, Koh09, Roy04]. Calls [Mei06]. Calystegia [WH01]. Camaleón [FSC02]. Camarón [HGH^{+}03]. Cambiando [LCTL05]. Cambiante [DC06b, HH08, LRO^{+}08]. Cambio [ARA^{+}09b, BTK08, BBO^{+}01, Bro08b, CHW^{+}01, CPSS03, Car07, CLW^{+}09, DSJ02, EMW^{+}04, FBL^{+}09b, GMA07, HBBD08, HL07b, LW01, LBG03, LBPBG07, MBMLJ09, MBQAM06, MKK^{+}09, MOO09, MCG^{+}09, MPR^{+}01, NRBM05, Nos01a, OHC08, PM05, PTP^{+}08, RO08, RBT08, RPT00, Ric07, Ros08, Rud06, SRW06, SSM03, SYWR03, SBP07, SCC^{+}09, SBSL06, TVD03, TFGJ06, TS06, WHA^{+}05, XGS^{+}09, LSL^{+}06, MGPM04, RDAL09, SM09, SSFL08]. Cambios [Cua00, DDC^{+}00, DAR08, DMG^{+}00, GSB^{+}06, Gui07, LM07, McC01, NNM00, NMB09, PBK09, PMBN01, Pop08, REC^{+}00, SFD09, SPM05, SSY06, VGB^{+}04, WM01, BMA^{+}08, BCP^{+}02, FTC^{+}04b, Gav04, GV02, HMF03, LHPH06, MBMLJ09, RPP^{+}00, ZSSM09]. Camino [BMS08]. Caminos [Cas03, DS01, FB00, GS05, HR04a, JF04, KGM^{+}02, LSL04, PJ00, SRH^{+}07, SG04, SD01, Tur06, Urb06b, WCPB03, GB03, HR04a]. Camp [SB06]. Campanula [EKL^{+}00]. Campion [VSAG06]. Campo [DCP^{+}03, GM00a, Hay09a, JLZW03, PBRG07, VO03]. Campos [BSM08, RS02c]. camschatcense [TO02]. Can [Bre03, CPME09, Fje07, GT03, Gro05b, Jac09, Pér05b, Sch06]. Canada [CW00, DK06, FEGB09, GOL^{+}07, JML00, LME08, LSN07, LO02, ML02, MNL03, Ped03, QRF07, WE07, WU05, DK06, FEGB09, JML00, LME08, LSN07, LO02, ML02, MNL03, Ped03, SMC00, WE07, WU05]. Canadian [IMD01, MPFBH07, RPP^{+}00]. Canadienses [IMD01, RPP^{+}00]. Canal [VDHD01]. Canarias [HLJV07]. Canary [HLJV07]. Cañas [Sie04]. Cane [BS09, PBS03, Sie04]. cangrejo [KFK^{+}09]. Cangrejos [EHJ^{+}04, HBA09, JAH^{+}05, PLBBJC09]. Canine [HLSZ02]. Canino [HLSZ02]. Canoas [BBW^{+}07]. Canoe [BBW^{+}07, DGC05, NHHG^{+}05]. Canoe-Making [BBW^{+}07]. Canopies [WCP01]. Canopy [NH09]. Canora [NS01]. Canoras [BHB08, BFD07, Cas03, DWN01, JSHM06, OK01a, PSBG05, Sch03b, SS01, TB04, WS04]. Can’t [GBB^{+}08]. Canteras [BKK03]. Cantidad [McK02, WSLW00]. Capacidad [Cze05, SWK05, SBDD05, WS03, BTK08, CHFR09]. Capacitación [VO03]. Capacity [BTK08, Cze05, KNH07, LF09, RSPM05, RRCOM^{+}06]. Capas [BMB^{+}08]. Cape [WHA^{+}05]. Capercaillie [SGH02, SGH02]. Capital [Far08, PJB^{+}06, PS04, Sch06, PS04]. Capstone [BH01]. Cápsulas [NH09]. Capsules [NH09]. Captación [PSGA09]. Caption [Ano06c, Ano07f, Ano07g, Ano07h, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, BSKH^{+}07, BJPL07a, VSH^{+}08]. Captive [AAO^{+}07, Bow09, FPF00, FHB^{+}08, GBE00, GP08, KHM00, KLH08, ODE01, TTSP04, Wed02]. Captive-Bred [AAO^{+}07, FPF00, GBE00]. Captive-Breeding [Bow09]. Captivity [For02]. Captura [CLBC03, FDL^{+}05, JAH^{+}05, LG05a, LB05, LC07a, LRS09a, NC03a, PD03a, PV06, RDN06, UCAGG08, WELCH09, ŽWGW09, DLCV00, MBR^{+}02]. Captura-Marca-Recaptura [LG05a, PV06]. Capturas [CLŽ^{+}07, PBSR06]. Capture [CLBC03, LG05a, PD03a, PV06]. Capture-Mark-Recapture [LG05a, PV06]. Capturing [BCK^{+}09, NC03a]. Carabid [HKN01]. Carabidae [HKN01]. Carábidos [HKN01]. Caracaras [MH01b]. Caracoles [HNG04, MNL03, MMF^{+}05, SLELA05, SF05, Urb06b]. Caracteres [CWNW01, GGA01]. Características [BFB^{+}09, CPO03, FDBM04, JP09, KD05, ML04, MTB00, SFD09, SKBY03, BGD09, RSJE06]. Caracterización [DBB^{+}09b, SS00]. Carapa [FJ07]. Carbon [BW08a, Mef07b, Orr07a]. Carcharhinus [WELCH09]. Cardo [LO02]. Careers [MSL06]. Carey [LCS03, PLB09]. Carga [CGR01, JGP09, LRS^{+}09b]. Caribbean [FCG^{+}05, GHL08, HR04b]. Caribe [FCG^{+}05, GHL08, HR04b]. Caribou [Sch03a]. Caribú [Sch03a]. Carismáticos [FM09]. Carl [SGLB^{+}07]. Carne [CMR05, DC06b, FYC00, FPM02, PCME09, TB03, WSA^{+}05]. Carnero [ODE01]. Carnivore [Bei02a, Ber06, Ber07, Car07, CMK^{+}07, Cre01a, TK03a, TNTH^{+}04, ZHM08]. Carnivore-Livestock [Ber06]. Carnivores [Cro02, FVS03, MNA07, OWOF03, RB07, Smi05]. Carnívoro [TNTH^{+}04]. Carnívoros [Ber06, Ber07, Car07, CMK^{+}07, Cre01a, Cro02, MNA07, OWOF03, RB07, TK03a, FVS03]. Carolina [KU02]. Carpinteros [Lam04, MGC01]. Carretera [MSWF00]. Carreteras [CM06, GS02, Has00, JSWS00, LCT^{+}06, LCG^{+}08, MMGB05, TF00, WSR^{+}00]. Carretero [CF01, For00a, FD00]. Carricero [FCL00]. Carried [GPZ04]. Carroll [MBFT09]. Carson [Pre09]. Cartografiar [CMS09]. Cascada [XGS^{+}09]. Cascadas [PJYO04]. Cascades [PJYO04]. Cascading [XGS^{+}09]. Case [APS07, APSB02, ACMA04, BT07a, BMBM06, BR01, BH04, DCD^{+}02, DA07, DBML07, FVS03, Fat06, Gay02, GM01, GZK06, GMB^{+}08, HDA01, HW08, HGH^{+}00, HY03, HLJV07, HW02, Ibi08, IB03, JHHL04, JMH01, KZSV05, KU02, KK09, Lin03b, LZD^{+}03, MBM04, MKH07, MC09, MCA^{+}00, MS09, MCM07, NWRD03, Orr09a, ODE01, OK00, PBK09, PV09a, PLL^{+}05, RAL^{+}04, RWJRR06, RWM^{+}06, RG04a, RMK07, RCMV06, Rom03, Rua00, Sch03b, SH03, SDRO01, SFT00, SD01, TVD03, TPW03, VEST02, VTA02, WSS^{+}04, WD01, WMPC08, Wil02c, Wil08b]. Caso [APS07, APSB02, ACMA04, BMBM06, BR01, DCD^{+}02, DBML07, FVS03, Fat06, GM01, GZK06, GMB^{+}08, HDA01, HW08, HGH^{+}00, HY03, HLJV07, HW02, JHHL04, JMH01, KZSV05, KU02, Lin03b, LZD^{+}03, MBM04, MKH07, MCA^{+}00, MS09, MCM07, NWRD03, ODE01, OK00, PBK09, PV09a, PLL^{+}05, RAL^{+}04, RWJRR06, RWM^{+}06, RG04a, RMK07, RCMV06, Rom03, Rua00, SH03, SDRO01, SFT00, SD01, TVD03, TPW03, VEST02, VTA02, WSS^{+}04, WMPC08, Wil02c, Wil08b, DA07, WD01]. Cassandra [DD04a, RS03]. Casual [MS08]. Casuales [MS08]. Casualties [McN03]. Cat [NMT^{+}04]. Cataloguing [Mic08, Poo00]. Catastrófica [ES04]. Catastrófico [TSG06]. Catastrophic [ES04, TSG06]. Catch [MBR^{+}02]. Catchments [Mag01, PSGA09]. Categorías [dLBBG^{+}08]. Categories [dLBBG^{+}08]. Categorización [OBK^{+}04]. Categorization [CD07, DC06a, OBK^{+}04, SM07b]. Cats [Din03b, LRS09a]. Cattle [HH03, KS03, KBR03, Mar05, PM05, RMSG^{+}03, TVB^{+}06]. Cauliflower [AYKM04]. caurina [HMF^{+}04, HODB09, NB06]. Causa [GM00a, HGH^{+}03, MBC03]. Causadas [SMG00]. Causado [MMLM01]. Causados [SRHS09]. Causal [TSKD08]. causante [KFK^{+}09]. Causas [APS07, EHJ^{+}04, Fox07, PBN^{+}04, Wol00]. Cause [GM00a, Gui07, MBC03]. Caused [GFTB04, MMLM01, SRHS09]. Causes [APS07, EHJ^{+}04, Fox07, HGH^{+}03, PCD^{+}03, PBN^{+}04, Wol00]. Causing [CW02]. Cautionary [BG04, EC01]. Cautivas [KLH08, TTSP04, Wed02]. Cautiverio [AAO^{+}07, FPF00, For02, GP08, KHM00, ODE01]. Cautivos [FHB^{+}08]. Cave [CMCH00]. Cavernícola [CMCH00]. Cavidades [Koe03]. Cavity [Koe03]. Cavity-Nesting [Koe03]. Cayos [Mcc09]. Caza [FYC00, JML00, LW02b, MBL02, PL03]. Cazadas [SFD09]. Cazadores [BST^{+}05, PL03, WZD^{+}00]. CBD [LJI^{+}08]. Cebra [BJPL07b]. Cecilianos [GW05]. Cedars [Moo07]. Ceiba [HDQSL06]. Cemeteries [BB01]. Cenchrus [ACMA04]. Censal [KW02]. Census [KW02]. Centaurea [LS01]. Centenaria [BB06]. Centennial [BB06]. Central [BBB^{+}07, DS01, DDBA^{+}07, LO06, MHC^{+}01, SMFPCA05, WSA^{+}05, BE00, BMMW07, CPME09, CNM01, DS01, DDMMZD06, DD04b, JG08, JSJ05, LSL04, LCT^{+}06, LBPBG07, LRW03, LAN05, MR00a, PA04, PJK03, SMFPCA05, SS01, TB03, WSA^{+}05]. Centro [HRW^{+}03, MHC^{+}01, DDMMZD06, JSJ05, LRW03, LAN05]. Century [Par01, Whi08, Rom03, SHB08]. cerca [GB01]. Cercado [JF04]. Cercanos [Ric04b]. Cerdo [TRSR00]. Cerdos [RPM09]. Cerradas [BMC^{+}09]. Cerrado [KM05]. Certhia [HAJ^{+}04]. Certificación [PBRG07]. Certification [Ben00, Ben01a, Ben01b, CPDO01, Gha01, PBRG07, PR01b]. Cerulean [Whi06]. cervicaria [EKL^{+}00]. Cetacean [BEH06, For00b]. Cetaceans [TWGH00]. Cetáceos [For00b, TWGH00]. Chaco [FFHW03, FFHW03]. Chain [CMR05]. Chains [Lus03]. Chairs [PS01b]. Challenge [Ada08, Ano02h, GB05b, JGM03b, MPE^{+}07, RA09, RP07]. Challenges [Ano02f, Baw06, Bra05a, BDRC05, CVB^{+}09, FC05, Gha07, HH06b, LCC^{+}07, MB07, MKVZ^{+}09, Mef06b, Rig01, Rod05, RSPM05, RS06, TPS^{+}05]. Challenging [Bra05b]. Chamaecrista [FG00]. Chamaedorea [EGPS04].

Chameleon [Sua02]. Chamorro [MBC03]. Chamorros [MBC03]. Chance [O’N07, Rol00]. Change [Asc02, ARA^{+}09b, Bee02, BTK08, BBO^{+}01, BRK^{+}02, BRK^{+}03, Bro08b, CHW^{+}01, CPSS03, Car07, CLW^{+}09, Cor03, DSJ02, DDC^{+}00, EMW^{+}04, FBL^{+}09b, GVS09, GSB^{+}06, GMA07, Har07, HBBD08, HL07b, Hol00, Hun07b, JGM03b, Kum08, LW06, LW01, LSL^{+}06, LBG03, LBPBG07, MBQAM06, MKK^{+}09, MOO09, McC01, McK09, MCG^{+}09, MGPM04, MHS07, MPR^{+}01, Mer08, NRBM05, Nos01a, NMB09, OHC08, Orr08a, Par08, PL06, PMBN01, PM05, PTP^{+}08, RO08, RBT08, RPT00, Ric05, Ric07, Rom09, RS06, Ros08, Rud06, SRW06, SM09, SPM05, SSM03, SSFL08, SSY06, SYWR03, SJP^{+}07, Spe07, SCC^{+}09, SBSL06, TVD03, TFGJ06, WHA^{+}05, XGS^{+}09]. Changes [BMA^{+}08, BCP^{+}02, Cua00, DAR08, FTC^{+}04b, Gav04, GV02, Gui07, HMF03, LM07, MBMLJ09, Mef09b, NNM00, PBK09, Pop08, RP07, REC^{+}00, SFD09, WM01, ZSSM09]. Changing [DC06b, HML^{+}02, HH08, HH09, LCTL05, LRO^{+}08, RPP^{+}00, RDAL09, Spe00, SBP07, TS06, VGB^{+}04]. Channel [Dav05]. Chapulines [KT02]. Characteristics [BFB^{+}09, BGD09, CPO03, MTB00, RSJE06]. Characterization [DBB^{+}09b]. Characterizing [SS00]. Characters [CWNW01]. Charcas [BSM08, BBH06, HRS08]. Charcos [CMCG07]. Charismatic [FM09]. Charr [MY02]. Checa [KFK^{+}09]. Check [Car02b, ZZ05]. Checklist [Nos03]. Cheetah [DBML07, KD00]. Cheetahs [Kel01, MMM03]. Cheetas [Kel01]. Chelonioidae [Car06]. Chesapeake [ZKNL02, ZKNL02]. Chewing [FMLL06, NBM^{+}04]. Chiapas [RCMV06]. Chicago [RH08, WM01]. Chickens [BJTD03]. Chicks [WBW05]. Chihuahua [MNH01, PMBN01]. Chihuahuan [LCC^{+}07, MNH01, PMBN01]. Chihuahuense [LCC^{+}07]. Children [You02, Fie03a]. Chile [ABE01, CCM00c, CNM01, DGSS03, FC05, LAN05, PA04]. Chilean [GL03, WMS^{+}01]. Chileno [WMS^{+}01]. Chimpancé [PPW^{+}07]. Chimpancés [TAP^{+}01]. Chimpanzee [PPW^{+}07]. Chimpanzees [TAP^{+}01]. China [Che09, EYZ09, Gru08, GX09, GJLL08, HPLTH07, HPK^{+}04, JHHL04, Lee09, LCZX09, LZD^{+}03, Qia07, Sey04, SFX^{+}08, XM07, ZJ04]. Chinese [EYZ09, HRS^{+}07, PPPV08]. Chinook [BPB07, HM08, HM09, Wil03, ZSMW06]. Chionochloa [LLW02]. Chipas [RCMV06]. Chipe [FTR01]. Chocan [CPR^{+}07]. Choice [BCS04, O’N07, RG04a, Vee00]. Choices [Bec09, BMS08, PCL08]. Choir [Joh05a, Nad04]. chouardii [Gar03a]. Chromosome [SL08]. Chronic [BHB08, TG04]. Chukchi [Tay03b]. Chytrid [KPH07, MSSM09, OMP^{+}05, Sch10]. Chytridiomycosis [Mcc05, WA05]. Cichlid [FPF00]. Cíclido [FPF00]. Cien [SAA^{+}09]. Ciencia [Cle06b, Cur02a, DWWF04, FN02b, MB07, MCF^{+}09, NFD^{+}09, PBC^{+}09, Sag07, SMRR02, SAM07, Wol06, ZMS^{+}08, Dav05, DN01, GC09, Lin03b, SM01, Tur06]. Ciencias [Mcs05, BV00]. Ciénegas [HBSD04]. Científica [RPT00]. Cientificas [BPV09, HCE^{+}01]. Científico [DWW04, Van06].

Científicos [LG05b, NV09]. Ciento [SPM05]. Cierre [LCS03]. Cierres [WELCH09]. Ciervo [BDSD^{+}05]. Cigüeñas [Ols07]. ciliaris [ACMA04]. Cimarrón [EMW^{+}04, SZS01]. Cimarrones [RPM09, RSW05]. Cinco [DAKB08, HMH00, HBBD08]. Cincuenta [SPM05]. Cinegética [FND06]. Cinegéticas [HMF03]. cinereus [GK06]. Circumpolar [FES01]. Circundante [NAC^{+}06]. CITES [CV08, Gin02, GB05b]. Citizen [EAR^{+}05, NTGT03]. Citizen-Science [EAR^{+}05]. City [Ano03a]. Ciudadanos [NTGT03]. Civil [Dud00, DGP^{+}02]. Civiles [DGP^{+}02]. Claims [LRS09a]. Clam [DSE^{+}08]. Clarification [Gor08, Pau08]. Clarifying [Hei03]. Clarity [Fla00]. Clark [MS02]. clarki [ALH^{+}04, NR03]. Clarkia [Pod01]. Claros [BV04, LSL04, LVBS^{+}04]. Clasificación [GZK06, HSMK07, HBKF05, JHHL04, LOM03, MCG^{+}08, RBM^{+}05, SSB^{+}08, SDRO01, TS09]. Clasificaciones [AFB^{+}00, SSS^{+}08, SDL07]. Clasificando [RCS^{+}03]. Clasificar [PD03a]. Classic [LDC06]. Classification [Col01, HBKF05, JHHL04, LOM03, RC00, RBM^{+}05, SSB^{+}08, SDRO01]. Classifications [AFB^{+}00, SSS^{+}08, SDL07]. Classify [PD03a]. Classifying [GZK06, MCG^{+}08, SBS^{+}09]. Classroom [Bal01, Bre03]. Clave [CHBK02, DMDFV08, Vee00]. Claves [Jan07]. clavium [PLL^{+}05]. Cleanup [LT05]. Clear [WD09]. Clearance [MRM^{+}05, STE^{+}01]. Clearcut [HKN01, PBL00]. Clearcutting [HNG04, KHHK03]. Clearing [KC08, MM07]. Clearings [LSL04]. Clethrionomys [TM04]. Cliff [KL06]. Cliff-Face [KL06]. Cliffrose [MBQAM06]. Clima [DBB^{+}09a, GB01, LRO^{+}08, LBPBG07, SRHS09, DDC^{+}00]. Climate [AE01, Bec09, Bee02, BTK08, BBO^{+}01, BRK^{+}02, BRK^{+}03, Bro08b, Car07, Cor03, DBB^{+}09a, DSJ02, DDC^{+}00, EMW^{+}04, GB01, GVS09, GMA07, HML^{+}02, Har07, HBBD08, HL07b, Hun07b, LW01, LRO^{+}08, LBG03, LBPBG07, MBMLJ09, MBQAM06, MOO09, McC01, McK09, MCG^{+}09, MHS07, Nos01a, Orr07b, Orr08a, Orr09a, Par08, PM05, PTP^{+}08, RO08, RBT08, RPT00, Rom09, RS06, Ros08, SRW06, SSFL08, SJP^{+}07, Spe07, SRHS09, TVD03, WHA^{+}05, XGS^{+}09, ZSMW06]. Climate-Change [DSJ02, MOO09, Par08, Rom09]. Climate-Driven [MBMLJ09]. Climatic [CBH01, FDBM04, Fox07, MGPM04]. Climática [AE01, FDBM04, Fox07, ZSMW06]. Climático [BTK08, BBO^{+}01, Bro08b, Car07, DSJ02, EMW^{+}04, GMA07, HBBD08, HL07b, LW01, LBG03, MBMLJ09, MBQAM06, MOO09, MCG^{+}09, MGPM04, Nos01a, PM05, PTP^{+}08, RO08, RBT08, RPT00, Ros08, SRW06, SSFL08, TVD03, WHA^{+}05, XGS^{+}09]. Climáticos [DDC^{+}00, McC01]. Climbing [KL06, ML02, MNL03]. Clode [Fra01]. Clonación [Ehr06]. Clonal [CNC04, YE07]. Cloning [Ehr06]. Closed [BMC^{+}09, TRSR00]. Closing [LS03, PR08]. Closures [LCS03, WELCH09]. Cloud [CFC02, MK09, PMEH05]. Cloud-Forest [CFC02, MK09]. Clouds [Mig00]. Clues [Jan07]. Cluster [Che09, TK06, Che09]. Clustered [LKG06]. Co [GEJK05, GB05a]. Co-Management [GEJK05, GB05a]. Coalition [Joh05b]. Coarse [Hun05, LGK02, SDJK04]. Coarse-Filter [SDJK04].

Coast [WSLW00, DD01]. Coastal [AJ05, BBM04, BRA^{+}03b, CKL^{+}00, CFO^{+}09, CNM01, DMS03, FSC02, MSBOF^{+}09, PAHM09, Rub01, SW07, SSB^{+}08, SB06, TWGH00, VTRG00]. Coastal-Marine [SB06]. Coasts [GFTB04]. Cobertura [Cua00, FJS04, GGW09, WM01]. coccineum [MRP07]. Cocoa [BMV09]. Coda [Got06]. Code [Ano05k]. Coded [Ano08l, FF08]. Códigos [Rub06b]. Codorniz [RAL^{+}04]. Coefficient [CGR01]. Coeficiente [CGR01]. coenia [Had00]. Coenonympha [CWNW01]. Coffee [ARP05, Hie05, KJSH08, KSDBT02, Mad05, MK09, PV02, PD03b, PBRG07, PAA^{+}08, PMEH05, RKR03a, RKR03b, Ric04b, RCP00, SMFPCA05]. Cogestión [GEJK05, GB05a]. Cognition [Wil08b]. Coherent [BVL^{+}03]. Coherente [BVL^{+}03]. Coincident [PHW05]. Coincidente [PHW05]. Colaboración [HSDS01, SCS06]. Colapso [BB06, BVK02]. Colaterales [LSOBA04]. Cold [Bar03]. Colecciones [HGH^{+}00, OGH04, PCFC01]. Colectadas [BST^{+}05]. Colectores [TH03]. Colegiadas [HP01]. Coleoptera [DCD^{+}02, HKN01, DCD^{+}02]. Coliflor [AYKM04]. Colinas [Lam04, aARLG03]. Colisión [TAJ04]. Collaborating [SCS06]. Collaboration [GPZ04, Mul07, Wes03]. Collaborative [Cur07, DTCC03, GALW09, HSDS01, KB06, GCBF09]. Collapse [BVK02, SK08]. Collateral [LSOBA04]. Collected [BST^{+}05]. Collecting [AGBP09, KMGB07, Sie04]. Collection [OGH04, PCFC01, VO03]. Collections [HGH^{+}00]. Collectors [TH03]. Collide [CPR^{+}07]. collina [WH01]. Collision [TAJ04]. Colobus [OALM^{+}00]. Colombia [Álv02, FNS^{+}03, KFSR^{+}06]. Colombian [ARP05]. Colombianas [ARP05]. Colonia [BF03, TS03]. Colonial [MTB00]. Colonias [HHSC00]. Colonies [HHSC00]. Colonist [SMT07]. Colonización [DMS03, EM03, FHL^{+}04, Had00, LF00, BPB07]. Colonizaciones [CSBS01]. Colonizadores [SMT07]. Colonization [BPB07, CSBS01, DMS03, FHL^{+}04, Had00, Hun07b, LF00, SAS^{+}09]. Colony [BF03, TS03, WBW05]. Color [DAKB08]. Coloración [FH09]. Colorado [BRYK01, FKU05, FKU05, NHHG^{+}05, OK01a, RFD01, WTEK08, BRYK01]. coloradoensis [RFD01]. Coloration [FH09]. Colossoma [RW02]. colubrina [SS02a]. Columbia [QRF07, WSS^{+}04, HWH^{+}02, QR05, QRF07, VTRG00, WSS^{+}04, WM05, WRW^{+}02, WWR^{+}02]. Colyvan [Tod00]. Combat [WWH03]. Combate [WCD^{+}08]. Combating [WCD^{+}08]. Combinación [FEKH09, Che09, PD03a, RBG^{+}04, TTSP04, TWGH00]. Combinados [CLC^{+}08]. Combined [CLC^{+}08]. Combining [Ano07e, Che09, PD03a, RBG^{+}04, TWGH00]. Combustibles [Bur04]. Comerciables [Ros06]. Comercial [GJLL08, KB03, WMPC08]. Comerciales [CMA^{+}06]. Comercialización [CMR05, SCP^{+}02]. Comercio [BM05b, COW03, CV08, DC06b, DSTPW07, FV05, LD03, PC08, RGW09, WCD^{+}08, WSCP05, ZJ04, CMM^{+}05b]. Comes [Ara09a]. Coming [Ano02b]. Comisiones [GBH^{+}05]. Commandments [Ehr01, Orr01c]. Comment [BCK^{+}09, Sch10]. Comments [BB00, Jes03, KB05]. Commerce [PC08].

Commercial [GJLL08, JML00, WMPC08]. Commission [GBH^{+}05]. Commitment [Ano01l]. Committee [TOR^{+}04]. Commodity [CMR05, Don04]. Common [Bro06, BJE^{+}05, Che02b, EN07, HJ07, HBSD04, Law06, LLW02, ML04, PRL^{+}02, Pas00, PN05, Roo00, SM09, SS01, Sua02, TMR^{+}07, VDHD01, VVMV09, WCX09]. Common-Ground [TMR^{+}07]. Common-Pool [Che02b]. Commonness [VEST02]. Communicating [Nad04, WG08]. Communities [ASD01, BJM04, BE00, BCS00, Bey04, BJB06, BS01b, Bra06, BRA^{+}03b, CMP^{+}00, CBG^{+}00, CFO^{+}09, DJC^{+}07, Dun04, Kra01, KL06, Lai05, Lam04, LEPH02, LGK02, MR00a, MRM^{+}05, NKW09, OK01a, OSSK^{+}07, Old03, PMZ06, PMBN01, Put03, RDW01, RWRW04, Rut01, SRSP01, SMM09, SLELA05, SSM03, SMC00, TF00, VDHD01, WM01, WGK07]. Community [BPV09, BD06, Ber04c, BF03, BMPNC^{+}03, CMM05a, Cur02a, DMSP00, DS08, DMG^{+}00, DDN01, EM08, FH05, GBC^{+}04, Gui07, GS08, Inf01, JC07, JH05, Kai03, Kes01, KU02, LLF08, LBPBG07, MH00, MSBGP03, MNL03, MO04, New06, NMB09, Per00, PP01, RTN01, SCB^{+}01, SSY06, SHM04, Ste08a, SDJK04, Tow03, Urb06a, Ano05g, Ano06f]. Community-Based [BPV09, Ber04c, Cur02a, DS08, JH05, SCB^{+}01]. Community-Composition [EM08]. Community-Managed [BMPNC^{+}03]. Community-Wide [Gui07]. Como [Ber04a, CMP^{+}00, HDA01, LOM03, MKDT01, PBSR06, PW01, SDRO01, MMBP08, AGBP09, BKK03, BSS^{+}05, BFMR07, BF03, BMB^{+}08, BPB^{+}05, BMPNC^{+}03, BRA08, CMC03, Cha06, CKL^{+}00, CDD^{+}05, DGC05, DDMMZD06, ESB05, Ehr06, FYC00, FVS03, FLF04, GB03, GM00a, GSB^{+}06, GV05, HHSC00, HL07b, HN09, Jan07, JLZW03, JMS06b, KDD^{+}04, KDL07, KTN^{+}09, KU02, KB03, KPJ06, LME08, LVM02, LM07, LV04, MMGB05, MB01, Mcc05, MEA00, MGC01, MBC03, MBB^{+}03, MBB^{+}08, NH09, NHHG^{+}05, NS09, NR03, NMB09, OHC08, OIK^{+}06, PA08, PBK09, PC08, QR05, RL06, RA04a, RAW07, Rub01, SRC^{+}02, SFS00, SBB^{+}08, SMFPCA05, Spe02, SDJK04, TB03, Tho05b, TRCC^{+}05, Van06, VSH^{+}07, VEST02, VIMMO05, VK07, WPS08, WD06, WCP01, Fje07, HPG09, NV09]. Comoros [GB05a, GB05a]. Companies [Put03]. Company [Ste08b]. Comparación [AOGMS09, CBG^{+}00, CPR05, CB01b, FPM02, HR04a, HR04c, IMD01, Lam04, TCSD00, Vee06, BPV09, BMC^{+}09, CLŽ^{+}07, HSB01]. Comparaciones [GS08, LLW02, SHM00, LB01]. Comparada [Cas03]. Comparar [TWGH00]. Comparativa [Cra02]. Comparative [ABE^{+}02b, BOC00, Cas03, Cra02, Rob00, SK04]. Comparativo [ABE^{+}02b, Rob00, SK04]. Comparativos [BOC00]. Compare [TWGH00]. Comparing [BPV09, CLŽ^{+}07, SHM00, TCSD00]. Comparison [AOGMS09, BMC^{+}09, CBG^{+}00, CPR05, CB01b, FPM02, HR04a, HR04c, HSB01, IMD01, Lam04, LB01, Vee06]. Comparisons [GS08, LLW02]. Compartimentos [Olu00]. Compartments [Cor00, Olu00]. Compatibilidad [Sag07, Pan02]. Compatibility [Pan02, Sag07, VJ07]. Compendio [Gra03]. Compendium [Gra03, Sur02]. Compensation [App03, NTGT03]. Compensatoria [BCS04]. Compensatorios [NTGT03].

Compensatory [BCS04]. Competencia [Cre01a, PAHM09, SMG00]. Competing [RHA^{+}05]. Competition [Cre01a, PAHM09, SMG00]. Competitivas [LLW02]. Competitive [LLW02]. Competitively [TTB03]. Competitividad [TTB03]. Compiladores [PK05]. Compilers [PK05]. Complejidad [ARP05, FLF04, KSH00]. Complejo [MBC03]. Complejos [BGHM06]. Complement [Hun05]. Complementar [Hun05]. Complementaridad [AW01]. Complementarios [HN09]. Complementarity [AW01, FRH04, Moi08]. Complementary [HN09]. Completas [GM01]. Complete [PA06b]. Completeness [GM01, NS09]. Completitud [NS09]. Complex [BGHM06, Ewi09, MBC03]. Complexities [KB04b]. Complexity [ARP05, FLF04, KSH00, PR01a]. Component [TRCC^{+}05]. Componente [TRCC^{+}05]. Componentes [KFSR^{+}06, TFGJ06]. Components [KFSR^{+}06, TFGJ06]. Comportamiento [BBM04, GVTM07, Lus03, WBW06]. Composición [BSS^{+}05, CKL^{+}00, CMA^{+}06, DDN01, EM03, EM08, GH02, HMB^{+}05, HH03, JC07, MDCK07, ML04, New06, RTN01, RPP^{+}00, SHT^{+}05, SBVC03]. Composition [BSS^{+}05, CKL^{+}00, CMA^{+}06, DDN01, EM03, EM08, GH02, HMB^{+}05, HH03, JC07, MDCK07, ML04, New06, RTN01, RPP^{+}00, SHT^{+}05, SBVC03]. Comprehensive [HM02]. Comprendiendo [BC08]. Comprensión [BFMM02, FCL00]. Compromise [Bei08]. Compromiso [Bei08, ZMS^{+}08]. Compute [Hol02]. Computers [Bre03]. Común [ASF^{+}09, BJE^{+}05, FCL00, ML04, PRL^{+}02, PN05, SM09, TMR^{+}07]. Comunalmente [BMPNC^{+}03]. Comunes [HJ07, HBSD04, LLW02, VVMV09, VEST02, WCX09]. Comunicando [WG08]. Comunidad [BD06, CMM05a, Cur02a, EM08, FH05, GBC^{+}04, Gui07, JC07, LLF08, LBPBG07, MSBGP03, MNL03, MO04, New06, NMB09, Per00, RTN01, SCB^{+}01, SSY06, SHM04]. Comunidades [ASD01, BJM04, BE00, BCS00, Ber04c, Bey04, BJB06, BS01b, Bra06, BRA^{+}03b, CBG^{+}00, DMSP00, DS08, DJC^{+}07, DMG^{+}00, Dun04, DDN01, GS08, Kai03, Kes01, KL06, Lai05, Lam04, LEPH02, LGK02, MH00, MR00a, MRM^{+}05, NKW09, OSSK^{+}07, Old03, PMZ06, PMBN01, RDW01, RWRW04, Rut01, SMM09, SLELA05, SSM03, SDJK04, Tow03, TF00, VDHD01, WM01, WGK07, CMP^{+}00, OK01a, SMC00]. Comunitaria [KU02]. Comunitarias [BPV09]. Comunitario [BF03]. Con [CMA^{+}06, APCS05, APS07, AAO^{+}09, AKM07, BBB08, BHMC01, BR01, BST^{+}05, BFD07, BVH^{+}02, BGD09, BS01b, BPB^{+}05, BGWE00, BMMW07, BRA08, CLBC03, CWC00, CJ08, CWNW01, Cla01, CWCG02, DCD^{+}02, DA07, DWW04, DN00, DBML07, ES04, EKL^{+}00, ES08, FE07, FVS03, Fat06, FRD^{+}09, FJ00, FMPB07, FNFM01, FH09, FV05, FMP^{+}05, GR03, GWS07, GLMN00, GC06, GPZ04, GG02, GAF00, HSR05, HRS08, HR04a, HMF03, HY03, HLJV07, IMD01, JW02, JHHL04, Kes01, KG03, KFBG08, KLH08, LME08, LP08, LM05, LD03, Lin03b, LRS09a, MBM04, MNB09, MBL02, MP08, Mcc09, MHM09, MHC^{+}01, MS08, MRAF^{+}07, NR03, OK01a, OBM09, Pan02, PHW05, PLBBJC09, PCHN04, PV06, PIH00, Pod01, PMC01, QRF07, QAMW03, RHS05, RRK02, RAL^{+}04, RG04a, RB06, RB07, RNL^{+}04]. con [Rua00, Sag07, SKBY03, SGSA07, SSSK06, SZ05, SSM03, Sem02, SH09, SMC00, SH02, SF05, SK02b, STS05, SCM^{+}07, TVB^{+}06, TMR^{+}07, TTB03, TAJ04, VVMV09, VBOV05, WPS08, WCD^{+}08, WBSG07, WGGS07, WMP05, WRL02, WSKH08, ZMS^{+}08, ZGR02]. Concatenación [TFG08]. Conceiving [SRW^{+}08]. Concentradores [WCP01]. Concentrators [WCP01]. Concentric [PGP04]. Concéntrico [PGP04]. Concepción [SRW^{+}08]. Concept [Bow09, Kot00, Pae05, PMC01, RA04a, Slo02, SGH02]. Concepto [Kot00, PMC01, RA04a, Slo02, SGH02]. Conceptos [PAHM04]. Concepts [CCM00a, Fla00, Hun00, Moi08, PS05, PAHM04, Wil00a]. Conceptual [BFMR07, FLF04, SMRR02]. Conceptuales [BFMR07]. Concern [KB04a, RBT08]. Concerns [DWW04]. Concesión [PCME09]. Concesiones [CPME09, RLC^{+}07]. Concession [PCME09]. Concessions [CPME09]. Conclusiones [ACC^{+}05]. Conclusions [ACC^{+}05]. Concluyentes [Mcc05]. Concordance [CWM08, LS07]. Concordancia [CWM08]. Concurrence [ES04]. Concurrencia [ES04]. Condado [PM07, OK01a]. Condición [AP00, CMP^{+}00, DW05b, TR01]. Condiciones [For02, GK06, MP07b, OFS^{+}04, PB09, WRL02]. Condition [AP00, CMP^{+}00, DW05b]. Conditions [GK06, OFS^{+}04, PB09, WRL02]. Condor [BS01a, MSB^{+}00, Bei01, Bei02b, MSB^{+}00, MSB^{+}01, Ris02]. Conducido [BS08]. Conducta [BBK02, CMC^{+}01, SFD09, SS02a, PPW09]. Conductora [NMB09]. Conductores [LM07]. Conductos [GB03]. Conductual [MS04]. Conductuales [Bra03a, FB08, JSJV03, WBW05, ZSSM09]. Conductualmente [GSG08]. Conduits [GB03]. Conectado [HCZ00]. Conectividad [CMCG07, CHFR09, DSE^{+}08, DOW^{+}06, LCS04, MA08, MU08, MB09b, RWS02, SC05, SB08, BBH06, DMS03]. Conejos [Cal06, DMDFV08]. Conespecífica [WS04]. Conexiones [CWCG02, Mad05, Pyk05, PB04]. Conference [CHMP05]. Confessions [Wil09b]. Confiabilidad [LWV04, MDCK07, MAP03]. Confiables [MD04]. Configuración [GH02]. Configuraciones [AW02]. Configuration [GH02]. Configurations [AW02]. Confinadas [HMH00]. Confinamiento [ZGR02]. Confined [HMH00]. Confinement [ZGR02]. Confiscation [And01b]. Conflict [BC08, CAR05, Cze08, MMM03, RAL^{+}04, RW03, TK03a, TNTH^{+}04]. Conflicting [Pem01]. Conflicto [BC08, Cze08, MMM03, RW03, TNTH^{+}04, TK03a]. Conflictos [BPB^{+}05, CAR05, DGP^{+}02, MBMLJ09, RAL^{+}04, RB07, TRNH00, TWW09, Ber06]. Conflicts [Ber06, BPB^{+}05, KK09, MBMLJ09, RB07, RP07, TRNH00, TWW09]. confluentus [PD03a]. Congo [DC06b, IIG05, PCME09, TAP^{+}01, WSR^{+}00]. Congress [And05, Bro04, Ter04]. Congressional [PD00]. Congruence [SDJK04]. Congruencia [SDJK04]. Congruencias [DC05]. Conifer [Hut06, KHP04]. Conífera [KHP04]. Coníferas [Hut06]. Conífero [SHM00]. Coniferous [SHM00]. Conjuntos [LRS^{+}09b]. Conkouati [TAP^{+}01, TAP^{+}01]. Connecting [DOW^{+}06, Ped07]. Connection [Orr07a]. Connections [Jes03]. Connectivity [BBH06, CMCG07, CHFR09, DSE^{+}08, DMS03, Joh05a, LCS04, MA08, MU08, MB09b, RWS02, SC05, SB08]. Connoisseur [Whi02]. Conocer [RFB09]. Conocidas [SS06]. Conocidos [HW02]. Conocimiento [AJF03, AGBP09, BTBC^{+}03, BBW^{+}07, DWW04, Dre05, RNP^{+}07, Wil08b]. Conquistó [LS01]. Consanguinidad [CGR01]. Consecuencias [Fea05, GFCP03, HBBD08, HJ07, Kel01, MTB00, MCM^{+}00, MSWF00, PLB09, RLC^{+}00, SGG07, STWH00, Tow03, BE00, BFB^{+}09, CW09, GPMM08, HRS08, JRL03, JGM03a, McC01, SWR00, VBOV05]. Consenso [PPP05, BRYK01]. Consensus [BBB^{+}07, BRYK01, Lea06, PPP05]. Consensus-Based [Lea06]. Consequences [BE00, BFB^{+}09, CW09, Fea05, GPMM08, HRS08, HBBD08, HJ07, JRL03, JGM03a, Kel01, MTB00, McC01, McD09, MSWF00, PLB09, RLC^{+}00, SGG07, SWR00, Tow03, VSK01, VBOV05, WB06, Wie06]. Conservación [AAO^{+}09, ADF04, Álv02, AMF^{+}01, AW02, Ang00, Ant08, ALM07, ARPEBM09, Asq01, BNH02, BE00, BTM^{+}05, BMBM06, Bat04, BMRB04, Bei08, Ber04c, BM04, BKWB05, BMV09, BSKH^{+}06, BFMR07, BNP^{+}07, BR05, BMB^{+}08, BIK^{+}06, BZD03, BFH^{+}06, BWS09, BRYK01, BD05, BPL^{+}01, CWL^{+}00, CWM08, CES05, CP00, CSS00, CVB^{+}09, CPR^{+}07, CPD^{+}09, Che09, CS05, Cho06, CMM05a, CMBC00, CMC^{+}00, Cla01, CPME09, CMGSM00, Cle06b, CMK^{+}07, CHF00, Cor02, CBH01, CAR05, Cur02a, CW08, CMS09, Cze02b, Cze08, DCP^{+}03, DN01, DS08, DKW01, DK06, DW05a, DMDFV08, DC05, DDMMZD06, DM04b, DPB^{+}09, DJOW07, Dre05, Dun02, DBML07, Ehr00, Ehr06, Esl00, FKU05, Fai08, Far08, FP02, Fer01, FPSD09, FBC^{+}06, FMT^{+}04, FBL09a, FB08, FMPB07, Fje07, FBK00, FN02a, FBK02, GCF^{+}07, GGA01]. Conservación [GK08, GEJK05, GCV03, GWS^{+}06, Gib00, Gil01b, GHP^{+}05, GPZ04, GDAV00, GM01, GBH^{+}05, GB05a, GMB^{+}08, GBH^{+}09, GP08, GV05, HPF^{+}06, HMF^{+}01, HCE^{+}01, HJP05, HTM^{+}05, HK03, Hay09a, HBKF05, HPK^{+}04, HS06, HCM03, HW02, Hun05, HKM06, JSJV03, JSW08, Joh03b, JAH08, JMH01, JFS08, KSC^{+}06, KSH00, KEJ06, Kal05, KC01, KB04a, KWL^{+}09, KU02, KRM^{+}00, KCK^{+}08, KDC^{+}06, KCC06, KCR^{+}08, KCS04, Kri04, LG02, LLS00, LBR07, Lat00, LWSM03, Law08, LCTL05, LP08, LSL^{+}06, LKG06, Les05, LW02a, LSOBA04, LN06, LADM05, LSZ^{+}08, Lin03b, Log03, LHG^{+}03, LP01, LZD^{+}03, LV04, LG05b, LMMF00, MBM04, MB00, MFGB07, MCF^{+}09, MBMLJ09, MT01, MLMB08, MKH07, MOO09, McK02, MWBP07, MMBP08, MYFF^{+}07, MGPM04, MKVZ^{+}09, Mcs05, MCM^{+}00, MK02a, MPR^{+}01, MHGF04, MHM09]. Conservación [MPE^{+}07, MCR^{+}04, MU08, MAM^{+}03, MYH09, MBC03, MBB^{+}03, MH01b, MS09, MCM07, MSBOF^{+}09, NA06, NLCAR09, NC03a, NSA06, NRBM05, NP06, NFD^{+}09, NR03, Old03, OIK^{+}06, PZG^{+}05, Pan02, PCL^{+}03, PA06a, PC02, PSF05, PMC^{+}07, PSG^{+}00, PLBBJC09, PCHN04, PKB^{+}08, PD03a, PCC03, PR03, PCL08, PP03, Phi00, PC07, PSGA09, PMC01, PEC01, PB04, PS04, PK01, PK05, PS06, PBC^{+}09, PPW^{+}07, Pyk05, QR05, RBT08, RRK02, RM06, RCS^{+}03, RS02a, RWBG06, RC05, RMSG^{+}03, RPD01, RWJRR06, RWM^{+}06, RDW01, RLC^{+}07, RA04a, RG04a, RH01, RLK^{+}01, Rob06b, RG02, Rod05, Rod03, RSB05, RB07, RAG03, RCL^{+}06, Rua00, Rub01, Rub06b, Rud06, RDAL09, SC09, Sag07, SCB^{+}01, SMRR02, SSS^{+}08, SRW^{+}08, SC00, SJF^{+}05, SMV02, SPFF^{+}08, Say05, SGB^{+}08, SMN00, SZ05, SL08, Sha01]. Conservación [SSB^{+}08, SFX^{+}08, STWH00, SCP^{+}02, SMH^{+}07, SBB^{+}08, Slo02, SF05, SM01, SEBM03, Spe02, SHT^{+}05, SCS06, SMSB05, SCM^{+}07, SD01, SDJK04, SAA^{+}09, SLH00, SRHS09, TPS^{+}05, TG05, TFG08, TVB^{+}06, The03, TRNH00, TWGH00, TTB03, TRCC^{+}05, Tur06, TAP^{+}01, Van06, VSH^{+}07, VO03, VTR^{+}09, Wad00, WPS08, Wal03, WTEK08, Wal04, Wal00, WM01, Wan04, WS04, WFP01, WS03, WMD^{+}04, WMPC08, Wil02a, WSR^{+}00, Wil02c, Wil08b, WSBP09, WWR^{+}02, WRL02, Wol06, WG08, Wun07, YLR^{+}01, You05, ZMS^{+}08, ZBABR00, ZKNL02, MH02, ATG05, AJF03, AJ05, Bro08b, CLMD05, FC05, Fre04, GW05, HCBK00, Hie05, HDE00, JTS06, KM05, KFNB09, KHP04, LFP05, MG05, MCA^{+}00, MLB08, MH02, MGH^{+}09, MB09b, Nel09, NB06, PLA00, PPP05, RNP^{+}07, Rap06, RM01, Rob09, RW03, SP09, SS05, SHYH06]. Conservación [ŽWGW09]. Conservadas [FTC04a]. Conservados [LVBS^{+}04]. Conservancy [DS08]. Conservando [MHGF04]. Conservar [Cze05, IF02, LF09, MP07a, ZR00]. Conservation [Abe02a, AAO^{+}09, Aco00, ATG05, AO06, ADF04, Álv02, AJ05, AMF^{+}01, AW02, ABE01, And01a, And01b, Ang00, Ang01, Ank04, Ano00d, Ano01j, Ano02b, Ano02j, Ano03e, Ano03a, Ano04d, Ano05c, Ano05g, Ano05k, Ano06f, Ant08, Ara09a, ACD^{+}07, ARPEBM09, Asc02, Asq01, Awo08, BNH02, BT07a, BE00, Bal02, BGRJ00, BC06, BTM^{+}05, BMBM06, Bat04, BG04, BSR04, Baw05, Baw06, Bea07, Bea09a, Bea09b, BMRB04, Bee00, Bei05, Bei08, Bek04, BBB^{+}07, BB00, Ber01, Ber04c, BH01, BBC^{+}04, BM04, BKWB05, BW09, BMV09, BFJ04, BSKH^{+}06, BFMR07, BNP^{+}07, BRD03, BR05, BMB^{+}08, Bow09, BIK^{+}06, BBM07, BDRC05, Bre01, Bre02a, Bre02b, Bre03, Bre06, Bri06, BISS06, BZD03, Bro08b, BDR04a, BFH^{+}06, BWS09, Bro06, BRYK01, BL07, BT07b, BD05, BPL^{+}01, Büs08, BAKHT06]. Conservation [Caf00a, CCM00a, CWM08, Cam08, CES05, CP00, CSS00, CVB^{+}09, CS07a, CPR^{+}07, Cha08, CPD^{+}09, Che09, CS05, Chi00b, Chi09b, Chi09a, Cho06, CRGG01, CMM05a, CMBC00, CMC^{+}00, Cla01, CPME09, CMGSM00, Cle06b, CMK^{+}07, CFO^{+}09, Cob03, Col00, CHF00, CCM00b, Cor04b, Cor02, CLMD05, CKF^{+}04, CBH01, CAR05, Cur02a, Cur07, CW08, CMS09, Cze02b, Cze03, Cze08, Da 03, DCP^{+}03, DN01, DS08, DKW01, DK06, DW05a, DMDFV08, DAB^{+}04, DC05, DDMMZD06, DM04a, DM04b, DPB^{+}09, DJOW07, Dre05, Du 01, Dud02, DHW01, Dun02, DGSS06, DBML07, Ehr00, Ehr06, Ehr08, EYZ09, Esl00, FKU05, FRH04, Fai08, FP02, FC05, Fer01, FPSD09, FBC^{+}06, Fie00a, FMT^{+}04, FBL09a, FB08, FBB^{+}09, FMPB07, Fje07, FBK00, FN02a, Flo05, Fol06, FNM05, FCN^{+}06, FBK02, Fre03, Fre04, Fre06]. Conservation [GL03, Gal01b, GMCP09, GCF^{+}07, GK08, GEJK05, GCV03, Gha01, Gha07, Gha08, GWS^{+}06, Gib00, GVS09, GLMN00, Gil01a, Gil01b, GA02, GHP^{+}05, GPZ04, GT03, GG06, GDAV00, GM01, GMHS00, GW05, GBH^{+}05, GB05a, GMB^{+}08, GOL^{+}07, GBH^{+}09, Gre09, GP08, GV05, GNR01, Had07, Haf08, Hal02, HPF^{+}06, HCBK00, HMF^{+}01, HML^{+}02, HJP05, HTM^{+}05, HK03, HKC^{+}08, Hay03, Hay09a, Hie05, HBKF05, HTU^{+}06, HPK^{+}04, Hil03, HS06, HDE00, HCM03, Hun00, HW02, Hun05, Hut08, HKM06, IB03, JSJV03, JMM07, JTS06, Jes03, JSW08, Joh01, Joh03b, JG04, Jon04, JH05, JAH08, JMH01, Juk01, JFS08, KSC^{+}06, KSH00, KEJ06, Kal05, KC01, Kaw04, KB04a, Ker09, KT06, KWL^{+}09, KU02, KRM^{+}00, KRW^{+}00, Kle06, KCK^{+}08, KM05, KDC^{+}06, KCC06, KC07, KC08, KCR^{+}08, Kni09, KCS04]. Conservation [KFNB09, KK09, KDKS08, Kri04, LG02, Laf03, LNP^{+}09, LMR^{+}07, LLS00, LBR07, Lat00, LF09, LWSM03, Law08, Lea06, LCTL05, LP08, LKG06, Les05, LW02a, LK04, LFP05, LSOBA04, LN06, LADM05, LFA^{+}07, LDC06, LSZ^{+}08, Lin03b, Log03, LHG^{+}03, LP01, LZD^{+}03, LS04, LV04, Lov06, Lov02, LG05b, LMMF00, MBM04, MB00, Mae02, Mag01, Mai04, MFGB07, Man02, MCF^{+}09, MG05, MBMLJ09, MT01, MLMB08, MKH07, MSL06, Mar02b, MBD^{+}03, MCA^{+}00, MOO09, McK02, McS03, MWBP07, MP07a, MMBP08, MYFF^{+}07, MGPM04, MBH07, MKVZ^{+}09, Mcs05, Mef01a, MEN06, Mef06b, Mef07a, Mei01, MSN06, MCM^{+}00, MK02a, MGB^{+}07, MPR^{+}01, MHGF04, MHM09, MPE^{+}07, MH02, MCR^{+}04, MGH^{+}09, MDW01, MU08, MDRB05, Moi08, MYH09, MBC03, MBB^{+}03, MH01b, MS09, MCM07, Mur00, MN07, Nad04, NA06]. Conservation [NSW08, Nas06, NLCAR09, NC03a, Nel09, NSA06, NRBM05, New08, NP06, NL03a, NPT05, NB06, Nor00, Nos07, Nos09, NFD^{+}09, NR03, O’C05, Old03, ODPW02, OS01, Orr09b, Oso00, OIK^{+}06, PZG^{+}05, Pan03, Pan02, PCL^{+}03, PA06a, PC02, Pat08, PE01, Pea01, PSF05, PMC^{+}07, PSG^{+}00, PZ01, PLBBJC09, PS01a, Pér05b, PLA00, PCHN04, PKB^{+}08, PP08b, PP08a, PD03a, PCC03, PR03, PPP05, PPP06, PCL08, PP03, PFR^{+}09, Phi00, PD03b, PC07, PSGA09, PMC01, PEC01, Pov01, PB04, Pre04, PB08, PS04, PKM00, Pri03, PK01, PK05, PS06, PBC^{+}09, PPW^{+}07, Pyk05, QR05, QR06, RBT08, Rai03, RT00, RNP^{+}07, RRK02, Rap06, RKR03a, RKR03b, Ray06a, Ray06b, RC04, RM06, RCS^{+}03, RS02a, RWBG06, RG06, RF00, RM03, Rei05a, RC05, RMSG^{+}03, RM01, RPD01, RWJRR06, RWM^{+}06]. Conservation [RDW01, RLC^{+}07, RA04a, RG04a, RH01, RLK^{+}01, Rob06b, Rob09, RG02, Rod05, Rod03, RSPM05, RRCOM^{+}06, RW03, RA04b, RBG06, RSB05, RB07, RP07, RAG03, RS06, RCL^{+}06, RANM08, Rua00, Rub01, Rub06b, Rud06, RDAL09, RVNM08, RVVNH09, Rya06, SC09, Sag07, SCB^{+}01, SMRR02, SSS^{+}08, SRSP01, SRW^{+}08, San06, SC00, SJF^{+}05, Sar06, SCJ^{+}04, SMV02, SPFF^{+}08, Say05, SGB^{+}08, SP09, Sch06, SNM00, SMN00, SZ05, SRL^{+}07, SL08, Sey04, Sha01, SSB^{+}08, She01, SMI^{+}07, SFX^{+}08, STWH00, SSK^{+}05, SCP^{+}02, SS05, SMH^{+}07, SB06, SBB^{+}08, Slo02, SVBR07, Smi04a, Smi09, SL00b, SF05, SM01, SEBM03, Spe00, Spe02, SHT^{+}05, SMSB05, Ste08b, Ste07, Sto04, SA04, SD01, SDJK04, Sua02, SAA^{+}09, Swa09, SLH00, SRHS09, TPS^{+}05, TG05, TSH06, TFG08, TVB^{+}06, Tem02, The03]. Conservation [TRNH00, Tho00, TWGH00, TTB03, Tho05a, Tho03, Tre04, TRCC^{+}05, TOR^{+}04, Tur06, TAP^{+}01, Urb06a, VP06, VL01, Van05, VS05a, Van06, VSH^{+}07, VBRNP08, VP07, VO03, VTR^{+}09, VBB^{+}08, VS05b, Wad00, WPS08, Wal03, WTEK08, Wal06, Wal04, Wal00, WM01, Wan04, WZW^{+}06, WS04, WK09, Wes05, WS03, Wes03, Whe03, Whe02, WE07, WHM01, Wie07, WMD^{+}04, WMPC08, Wil02a, WSR^{+}00, WMD^{+}06, Wil02c, Wil08b, WSBP09, WWR^{+}02, WRL02, Wit03, Wol06, WG08, Wun07, YBWZ06, You01, YLR^{+}01, You05, ZHM08, ZR01, ZMS^{+}08, ZBABR00, ZKNL02, ŽWGW09, Mef07a, Mef08, Wil00a, Edi03, Har06a, Mef06a, Rob06b, SGLB^{+}07, Wil00b]. Conservation-Area [JFS08]. Conservation-Priority [MPR^{+}01, PMC01, SBB^{+}08]. Conservation-Related [MHM09]. Conservation-Worthy [MB00]. Conservationist [Law06]. Conservationists [Bus02, CN06, Hen05, Jac09, Jam02, LS03, Woo00]. Conservative [CTNR02]. Conservativo [CTNR02]. Conserve [Ber04b, Cze05, DZM08, Gro05b, Hol02, IF02, LA00, SCS06]. Conserved [FTC04a, LVBS^{+}04]. Conserving [AJF03, ALM07, Cor00, Far08, FB08, GGA01, Gru08, Kre07, KHP04, MHGF04, MLB08, MAM^{+}03, MB09b, PS05, Rot01, SHYH06, WFP01, ZR00]. Consevación [PP08a]. Consideración [HH08, SFT00]. Consideraciones [AW02, HBC^{+}06]. Consideration [HH08]. Considerations [AW02, HBC^{+}06]. Considering [DZM08]. Consistencia [RBM^{+}05]. Consistency [RBM^{+}05]. Consistent [AFB^{+}00, GM09]. Consistente [GM09]. Consistentes [AFB^{+}00]. Conspecific [WS04]. Constitution [Lav03, Orr03a]. Constitutional [Raf03]. Constrained [NWRD03]. Constraints [KL06]. Constrasting [GAF00]. Construcción [FB00, MC05a, MY02]. Construction [FB00]. Constructive [JH05]. Consumer [BS08, BCHS08, WSA^{+}05]. Consumidores [BS08, BCHS08]. Consuming [DCT^{+}07]. Consumo [EG07, MBC03, PCME09, WSA^{+}05]. Consumption [ADD^{+}07, EG07, MBC03, PCME09]. Contaminación [PC08, SGG07, UMT^{+}01]. Contaminantes [BBB08, WCP01]. Contaminants [BBB08]. Contamination [UMT^{+}01]. Contemporánea [ARA^{+}09b]. Contemporary [ARA^{+}09b, EN07]. Content [LW06, VS05a]. Contentious [DW06]. Conteo [Pol06, TS06]. Context [Bor05, BAF04, CPO03, CA05, DM04b, DD04b, Hei01, JW02, JG04, Mef01a, PA04, PD03b, SD03, YE07]. Contexto [Bor05, BAF04, CPO03, CA05, DM04b, DD04b, FEGB09, JW02, PA04, SD03, YE07]. Contextualización [MCG^{+}09]. Contextualizing [MCG^{+}09]. Contiguo [MM02a]. Contiguos [CMCH00]. Contiguous [CMCH00, MM02a]. Continent [DC06a, Ped03, Pim07b]. Continental [DC05]. Continente [Ped03]. Continentes [WCX09]. Continents [WCX09]. Contingent [BCS04]. Contingente [BCS04]. Continua [GN00]. Continuas [FP02]. Continuidad [DBL06]. Continuing [GN00]. Continuity [DBL06, RGGS02]. Continuo [PW01]. Continuous [FP02]. Continuum [PW01]. Contornos [FLF04]. Contours [FLF04]. Contra [APSB02, LE02, SL00a, Mor00b]. Contracción [Rud06]. Contradiction [Ano02i, Hil03]. Contras [FRD^{+}09]. Contrastantes [GAF00]. Contribución [DAR08, MFGB07, OGH04, PPW^{+}07, RGW09, VSAG06]. Contribuciones [EBAW01, FTC04a, MM04]. Contribution [DAR08, MFGB07, OGH04, PPW^{+}07, RGW09, VSAG06]. Contributions [EBAW01, FTC04a, MM04]. Control [AL03, BCMA^{+}01, Ber06, BS08, BCHS08, Bey04, BEB00, CWL^{+}00, Car04, Cho06, CAFL03, HDA01, Hod04a, Hod04b, JLF09, KG02, LH04, LO02, LARR03, LS04, Nos03, Pem01, Pro04, PW01, BS08]. Controlados [DWW04]. Controlar [Hod04a]. Controversy [Tre04, Owe08]. Conundrum [Ang01, Bri06, Jor01]. Convencionales [CSB07, JSJV03, WHJJ04]. Conventional [JSJV03, WHJJ04]. Convergent [Fox07, KS05]. Convergentes [Fox07, KS05]. Conversion [Car07, HBV02, SWB^{+}08, SOADMR08, Car07, HBV02, SWB^{+}08, SOADMR08]. Conveyor [Ric04a]. Convolvulaceae [WH01]. Cool [Bar03, Per02]. Cooperation [RANM08]. Cooperativa [KB06]. Cooperative [RSW05]. Cooperativo [RSW05]. Copas [NH09]. Coprófagos [FH05]. Coprófilos [LMP02]. Coral [AYKM04, BMC^{+}09, BMP07, CMM05a, DSE^{+}08, DDN01, FMP^{+}05, GMHS00, GWJ^{+}07, GGC03, HR04b, Mas03, MA08, MBB^{+}08, O’C00, RPT00, TH03, WS03]. Corales [GWJ^{+}07, PIH00, WS03]. Coralinos [GWJ^{+}07, GGC03]. Corals [MA08, PIH00, Sur02]. Cordillera [WSLW00]. Core [CS07b, RPD01, SS00]. Corn [MMLM01, BRK^{+}03]. Corn-Rootworm [MMLM01]. Corona [Ped03]. Corporal [DAR08, Olu00, SFD09]. Corrections [Jes03]. Correctos [KMGB07]. Corredor [DOW^{+}06, SWD00]. Corredores [BBK02, CS06, CMS09, DGC05, FE07, Had00, HM04, KFA^{+}03, LP08, MH01a, MM03a, NHHG^{+}05, PW01, RCL^{+}06, SS00, Wel06, WE07, WHA^{+}05, PSF05]. Correlación [RF03]. Correlaciones [Ber06, CHW^{+}01, ES08, RS02a, VSGS02, WCX09, Bra03a, ERB01, JSW08, OBK^{+}04]. Correlates [Ber06, Bra03a, CHW^{+}01, ERB01, GNEE07, JSW08, KSB04, RS02a, WCX09]. Correlation [RF03]. Correlations [ES08, OBK^{+}04, VSGS02]. Correlativos [KSB04]. Corridor [CS06, DOW^{+}06, Had00, SS00, SWD00]. Corridors [BBK02, CMS09, DGC05, FE07, HRN00, HM04, KFA^{+}03, LP08, MH01a, MM03a, PSF05, PW01, RCL^{+}06, Wel06, WE07, WHA^{+}05]. Corruption [Orr04a, Put07]. Corte [BV01a, FMLL06, GV02]. Corteza [BFB^{+}09]. Corticoesterona [WBW06]. Corticosterona [TG04]. Corticosterone [TG04, WBW06]. Cosecha [BV01a, CWH07, Car07, EGPS04, FMDK09, Hut06, KTN^{+}09, LO06, SSSK06, Sie04, TNRJ02, TA08, TSP^{+}08, VMM07, ZSVL07, DA07, FO06, LN06, RBRB06]. Cosechadas [BMC^{+}09, RW02]. Cosechado [RR01]. Cosmopolitan [DAKB08]. Cosmopolitas [DAKB08]. Cost [BSW^{+}08, DJOW07, HM08, HM09, JMS06b, LADM05, Mcd04, RB07, WSPL09]. Cost-Effective [BSW^{+}08, DJOW07, HM08, HM09]. Cost-Effectiveness [WSPL09].

Cost-Efficiency [JMS06b]. Costa [MSBOF^{+}09, SSY06, TSP^{+}08, BD06, CWC00, DCP^{+}03, LLS00, SAPRB07, SLB^{+}07, TSP^{+}08, VTRG00]. Costas [DDC^{+}00, GFTB04]. Costeño [FSC02]. Costera [AJ05, DD01, WSLW00]. Costeras [CNM01, DMS03, PAHM09]. Costero [BBM04, BRA^{+}03b, CKL^{+}00, Rub01, SSB^{+}08, SB04]. Costo [RB07]. Costos [CWM08, NA06, PKB^{+}08]. Costs [BGRJ00, BMP^{+}06, CWM08, NA06, PKB^{+}08]. Cotton [CCC04, CW09]. Cougar [SLH00]. Council [CPDO01]. Count [Pol06, TS06]. Countdown [RP07, MPE^{+}07]. Counting [Din03a]. Countries [VP06]. Country [Rod05, VIMMO05]. Countryside [BG04, DCP^{+}03, HDDEB03, RDEF01, SLB^{+}07]. County [OK01a, PM07]. Course [BH01, BSM08, LCTL05, Ray06a]. Courses [Mag01]. Cover [Ano06c, Ano07f, Ano07g, Ano07h, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, BSKH^{+}07, BJPL07a, Cua00, FJS04, GGW09, VSH^{+}08, WM01, Whi06]. Cow [Pea01]. Cowbird [GM00b]. Coyotes [FH06, RSF^{+}03]. Crab [HBA09]. Crabs [HRS^{+}07, PLBBJC09]. Cracids [Bro05]. Cracks [SVBR07]. Crane [JGL^{+}02]. Cranes [HPK^{+}04]. Crataegus [GC07]. Crayfish [EHJ^{+}04, JAH^{+}05, KFK^{+}09]. Crayfishes [KFK^{+}09, PFL01]. Creación [GGT08]. Create [Had07]. Creating [Ano02h, GGT08, ID00]. Crecientes [Vee06]. Crecimiento [Cze08, HRS00, KHRS01, KHP04, Law08, MSN06, SFS05, SS01, CY08, JB06, Joh03b, Mcs05, Ros06, Ros08, UWB04]. Credibility [Ala08]. Créditos [RH08, SS08a]. Credits [RH08, SS08a]. Creencias [DHZM09]. Cremastra [CNC04]. Crepidula [VDHD01]. Crested [MH01b]. Criadas [AAO^{+}07]. Criaderos [AAO^{+}07, ZSSM09]. Crianza [BD05, OWOF03]. Crías [EKL^{+}00, FNP09, FPL^{+}04]. Cricket [BBK02]. Crickets [Ber05]. Crises [SBK^{+}06]. Crisis [Abe02a, BADF^{+}05, CNHT09, Cha08, FP07, Laf03, Leo04, Lin03b, Mef01b, ODPW02, SC09, SHB08, TB03, ZJ04]. Crist [SBH00]. Criteria [But03, DGG^{+}01, GHTMR01, MRAF^{+}07, ODE01, SH03, SB03, SDRO01, VTA02]. Criterio [KG03]. Criterion [KG03]. Criterios [KG03, MRAF^{+}07, NCVBW02, SH03, SDRO01, VTA02, ODE01, SB03]. Critica [JW02, FBC^{+}06, GSR^{+}04, GBC^{+}04, LMS^{+}02, LRS09a, MCM^{+}00, PA06a, YH01]. Critical [Büs08, CT07, Far08, HH06a, LRS09a, MCM^{+}00, Sem02]. Critically [AMF^{+}01, Gar03a, LVM02, TRCC^{+}05]. Críticamente [AMF^{+}01, LVM02]. Críticas [BMM^{+}02, RLXL01]. Criticism [FBC^{+}06, JH05]. Crítico [CT07, Far08, Gar03a, TRCC^{+}05]. Críticos [HH06a, Sem02]. Critique [JW02, LMS^{+}02, PA06a, RGGS02, YH01]. Critiques [BBA^{+}02]. crocea [DSE^{+}08]. Crocodile [Paq07]. Cromosomas [SL08]. Crónico [BHB08, TG04]. Crops [Álv02, Ric04b]. Cross [GS08, HPG09, RANM08, SDJK04]. Cross-Community [GS08]. Cross-National [HPG09]. Cross-Sector [RANM08]. Cross-Taxon [SDJK04]. Crossing [BV04, KTN^{+}09]. Crossroads [Ano01g, Hit07, Spe02].

Crowd [Fox08, Ste08b]. Crown [KF01b, Ped03]. Crown-Fire [KF01b]. Crowns [NH09]. Crusoe [ABE01, DGSS03]. Crustacean [JGM03a]. Crustáceos [JGM03a]. Crusts [BMB^{+}08]. Cruz [RPM09, STE^{+}01]. Cuando [FO06, HRS00, BSM08, Bat04, CPR^{+}07]. Cuánta [Sim03]. Cuantas [LP05]. Cuantificación [GWM^{+}07, Rau02, SBB^{+}08, BOJ09, FKU05, HMK^{+}09, LLS08, MCG^{+}08, OB00, PCVC06, RPM09]. Cuantificando [ND04]. Cuantitativa [HCE^{+}01]. Cuantitativo [MW07, Wal04]. Cuantitativos [BCP^{+}02, GV05]. Cuasi [FN04]. Cuatro [BST^{+}05, Cra02, Cun00, HR04a, Cre01a]. Cuello [BR01, BJTD03]. Cuenca [BRV06, BRYK01, FN02a, HWH^{+}02, SH02, WWR^{+}02, FMT^{+}04, BGD09, FKU05, MK02b, PL03, UMT^{+}01, WSR^{+}00, WRW^{+}02]. Cuencas [ABE^{+}02b, DDBA^{+}07, WD03]. Cuenta [Pra05, MPE^{+}07]. Cuento [RW03]. Cuerpo [CB01a, LP07, SCP^{+}02]. Cuestionable [Cle06b]. Cuestiones [KWL^{+}09]. Cuicatlán [TVD03]. Culebra [WBBA03]. Culebras [WBDP02]. Cultivadas [VW04]. Cultivating [Bre01]. Cultivation [GSB04, NSW08, Ram01]. Cultivo [BE00, Jan07, MBQ05, Ram01]. Cultivos [GSB04, Ric04b, Álv02]. Cultura [FFHW03]. Cultural [BB01, BBW^{+}07, Inf01, Pan03, RNP^{+}07]. Culture [Ano00d, Cur03, FFHW03, Fre03, Lam06, Rya06]. Cultures [Chi09b, FBB^{+}09]. Cumbre [GBC^{+}05]. cumulicola [QAMW03]. cupido [BJTD03]. Currency [EN07]. Current [EG07, GHL08, MT08, MRAF^{+}07, PTP^{+}08, Web01, Wil00a]. Curricula [Cla01]. Curriculum [BH01]. Curse [DD04a]. Curso [LCTL05]. Curva [LME08]. Curvas [JLF09, WMP05]. Curve [LME08]. Curves [JLF09, WMP05]. Customary [CSB07]. Cutthroat [ALH^{+}04, CK05, NR03, YH01]. Cutting [BV01a, Cre01b, GV02]. cyanea [Wel06]. Cyprinodon [RNL^{+}04]. Czech [KK09, KFK^{+}09].

Daisy [Mor00b]. Daly [ADD^{+}07, Tro00]. Dam [PCL^{+}03]. Dama [BDSD^{+}05, ZSSM09]. Damage [AYKM04]. Damming [MY02]. Dando [Pra05]. Dangers [BBM07, GLV05, Wie07]. Daño [AYKM04]. Dark [BT07a, Pim07b]. Darwinian [Blu07]. Darwinism [Dal00b, Tro00]. Dasmann [Jar03]. Data [AGBP09, Ano08l, BM06, Bre06, BDR04a, CLBC03, CHW^{+}01, EM08, Fal07, FF08, GR03, GCM^{+}08, GZK06, Gor08, HR04a, HR04c, JLRBB^{+}07, JLZW03, JFWP06, KMGB07, LSOBA04, LE02, MD04, MHC^{+}01, NL05, OGH04, PV06, PCFC01, Pre04, RKEJP06, RBHKK02, Roy04, SGSA07, Sta01, Ste08a, SCM^{+}07, TVB^{+}06, TCSD00, TA06, TS06, TNTH^{+}04, VO03, WM01, WSS^{+}04, WGGS07]. Databases [HLJV07, MD04]. Datos [AGBP09, CLBC03, CHW^{+}01, EM08, GR03, GZK06, HR04a, HR04c, HLJV07, JLZW03, JFWP06, KMGB07, LSOBA04, LE02, Mcc05, MD04, MHC^{+}01, NL05, OGH04, PV06, PCFC01, RKEJP06, RBHKK02, Roy04, SGSA07, Sta01, SCM^{+}07, TVB^{+}06, TCSD00, TS06, TNTH^{+}04, VO03, WM01, WSS^{+}04, WGGS07, Fal07]. Dave [Arc09]. David [Ano05c]. Days [Orr07b]. Deadwood [HKM06]. Deadwood-Dependent [HKM06]. Deaf [AO06]. Death [Kau04, MSWF00, Rom03]. Debate [BW08b, Che02a, Hun07b, KC08, MHS07, Phi00, Say03]. Debating [Gha08]. Debemos [GMMPP08]. Debida [SRW06]. Debido [GGI^{+}05, MBMLJ09, MY02, SFT00, SRHS09]. Débil [GMMPP08]. Debris [LGK02]. Debt [BJ05, CNPS04, HO02, Pur03, RRH^{+}09]. Década [QR05, SCM^{+}07]. Decade [QR05, SCM^{+}07]. Decay [LLV^{+}02]. Decaying [JMS06b]. December [Ano06j, Ano07r, Ano08n, Ano08q, Ano09m, Ano05i, Ano06h, Ano07p]. Decentralization [PB09]. Deciding [DZM08]. Deciduo [MB01, RR01, STE^{+}01]. Deciduos [HFR07, KA03, SBVC03]. Deciduous [GG02, KA03, MB01, RR01, STE^{+}01, SBVC03]. Decision [Blu07, Caf07, GWS07, MDW01, PPBW08, RAL^{+}04, RM03, Roo00, TW06, TUE06, WP03, Wil08b, PPBW08, RAL^{+}04, TW06]. Decision-Making [RM03]. Decision-Theory [WP03]. Decisiones [GWS07, Pra05, TUE06, WSS^{+}04, Wed02, Wil08b, BV04]. Decisions [BV04, BS01a, Pra05, RT00, WSS^{+}04, Wed02]. Deck [PS01b]. Declinación [AS01, Arm05, CDM09, CY08, DSJ02, Dun02, HBA09, LDG^{+}00, Mur01, Sch03b, SFS05, SWK05, SYWR03, TSP^{+}08, War05, WBBA03, BB06, BSW^{+}06, ES04, KDD02, PHW05]. Declinaciones [AE01, DKS^{+}09, FH09, Fra09, LRW03, MNB09, Mcc05, MHC^{+}01, Pol06, SGG07, CHW^{+}01, VVMV09]. Decline [BB06, BSW^{+}06, BFMM02, BS00, CY08, CW02, Dun02, ES04, KM00, KDD02, MKK^{+}06, PHW05, SFS05, SYWR03, Sta01, TSP^{+}08, War05, WBBA03]. Declines [AE01, Arm05, BR01, CHW^{+}01, DSJ02, DKS^{+}09, FH09, Fra09, LRW03, MNB09, Mcc05, MHC^{+}01, PCD^{+}03, Pol06, SGG07, Sch03b, VVMV09, YLR^{+}01]. Declining [AS01, But03, CDM09, LDG^{+}00, Men03, Mur01, PBN^{+}04, SWK05, SHM00]. Declive [SHM00]. Decreasing [Vee06]. Decrecientes [Vee06]. Dedication [Tut03]. Deep [Fle03, LSF04, Rob09]. Deep-History [LSF04]. Deer [ASM05, BDSD^{+}05, CAFL03, Côt05, FMDK09, MR00b, PLA00, PLL^{+}05, ZSSM09]. Deficiency [GZK06]. Deficiente [PBN^{+}04, RSB05]. Deficientes [PDH00]. Deficit [Ber07, Ber07, DC05]. Deficits [DC05]. Definición [CS07b, HCM03, MWBP07, RM06, MBM04, SWD00, TVB^{+}06]. Defining [DAB^{+}04, MBM04, SWD00, You01]. Definir [Che09, LV04, PEC01]. Definition [Tod00]. Deflating [Hay04]. Deforestación [JB06, KSO^{+}03, NSB^{+}06, RDAL09, CSM^{+}06, Fea05]. Deforestation [BS03, CSM^{+}06, Fea05, JB06, KSO^{+}03, NSB^{+}06, RDAL09]. Deforesting [Leo04]. Deformes [JLZW03]. Deformities [JLZW03]. Degradación [HVR08, SLELA05]. Degradado [MBM04]. Degradation [HVR08, SLELA05]. Degraded [MBM04, Sod06]. Degree [GS08]. del [ABE^{+}02b, ASM05, AT05, AGBP09, AMF^{+}01, AW02, AM05, ACC^{+}05, Ant08, ARP05, AOGMS09, ARA^{+}09b, ANW^{+}08, BNH02, BMM^{+}01, BCS04, BGC^{+}04, BCMA^{+}01, BHB08, BB06, BPB07, BST^{+}05, BMMR03, Ber06, Ber07, BBK02, Ber05, BJ05, BFD07, BFMM02, BBM04, BMV09, BEB00, BTBC^{+}03, BMB^{+}08, BWS09, BRYK01, BCP^{+}02, BAF04, BMMW07, BK02, BJE^{+}05, BRA08, Cal06, CMP^{+}00, CB01a, CF01, CPSS03, Car07, CJ08, CQLF02, CWNW01, Cha06, Chi00a, Cho06, CPD^{+}03, CMC^{+}01, CMBC00, Cla01, CHBK02, Cle06b, CWCG02, Coc01, CLW^{+}09, CPO03, CAFL03, CHF00, CCM00c, Cou02, CS07b, Cro02, CBH01, Cua00, Cun00, CV08, Cze05, Cze08, DDAK04, DCP^{+}03, DMSP00, DDC^{+}00, Dav07, DN01, DMDFV08, DWWF04, DNZK02, DM04b, Dre05, DD04b, DW05b, Dun04, DBML07, EKL^{+}00, EMW^{+}04, EK08, ED07, FKU05]. del [FVS03, Far08, FMDK09, Fat06, FRD^{+}09, Fer01, FPSD09, FMT^{+}04, FEGB09, FLF04, FCG^{+}05, FSC02, FPF00, FTR01, FRSG^{+}02, FM09, FFHW03, FCL00, FES01, FMLL06, For00a, FH09, FV05, FRR07, FBL^{+}09b, FBCG00, Fre04, Gav04, GM00a, GV02, GB01, GMW08, GN00, GDAV00, GMA07, GWJ^{+}07, GPC^{+}03, GSB04, GBH^{+}09, GD05, Gre05, GGW09, GVTM07, GV05, GH02, GFTB04, Gun02, GS08, GWM^{+}07, HHD^{+}08, HMK^{+}09, HH00, HODB09, HDA01, HCE^{+}01, HMB^{+}05, HTM^{+}05, HIS03, Has00, HR04a, HLSZ02, HH03, HP01, HBBD08, HLG03, HDQSL06, HH08, HR04c, HVR08, HL07b, Hob01, HBV02, HAG^{+}09, HFR07, HJ07, HJ05, HAJ^{+}04, HN09, IMD01, IF02, IML^{+}05, JLF09, JF04, JWB06, JGM03a, JMLG01, JSJ05, KSH00, Kai03, KEJ06, Kau04, KHRS01, KF01a, Kel01, KSP00, KTN^{+}09, KDG^{+}06]. del [KS03, KU02, KSDBT02, KM05, KDC^{+}06, KCS04, KSB01, KB08, Kot00, KFK^{+}09, Kri04, KL06, KA03, Lai00, LB05, LBR07, LLF08, LF09, LLV^{+}02, LCG^{+}08, LMP00, LCTL05, LBG03, LBPBG07, LLWCW09, LSN07, LH04, LW02a, LHPH06, LD03, LCS03, LCZX09, LE02, LO06, LARK01, LADM05, Lin03a, LCC^{+}07, aARLG03, LHG^{+}03, LP07, LAN05, LF00, LO02, LMMF00, LW02b, Lyn06, LBTS05, MCG^{+}08, Mad05, MR00a, MBL02, ML04, MM07, Mar05, MBQAM06, MRM^{+}05, MKK^{+}06, MCA^{+}00, MPDT01, McK02, MB01, ML02, MNL03, MP08, MYFF^{+}07, MM04, MS04, MCM^{+}00, MPR^{+}01, MSB^{+}00, MHM09, MLB08, MU08, MAM^{+}03, MW07, MML06, MBC03, ME00, MBB^{+}03, MY02, MH08, MNH01, MCM07, MK09, MSBOF^{+}09, NTMT^{+}03, NWRD03, NSA06, NT01, NNM00, NKW09, NB06, NRB00, NMB09, OALM^{+}00, OFS^{+}04]. del [ODE01, PRL^{+}02, PA06a, PA04, PM07, Ped03, PBN^{+}04, PMC^{+}07, PGP04, PSG^{+}00, Per01, PL03, PLBBJC09, PLA00, PN05, PV09a, PJYO04, PV09b, PB09, PPP05, PV06, Phi00, PBS03, PRMV06, PMBN01, PMEH05, PSGA09, PCFC01, PEC01, Pow06, PPW^{+}07, PM05, QRF07, RHA^{+}05, RO08, Ram01, RPM09, RRK02, Rap06, RMD^{+}02, RWBG06, RBM^{+}05, RW02, RWJRR06, RKEJP06, Rig01, RSF^{+}03, RBK^{+}05, Rob00, RA04a, RFD01, RRH^{+}09, RGW09, Rom03, RB06, REC^{+}00, Rub01, SAA^{+}01, SVV07, SRR00, SRW06, SRW^{+}08, SMM09, Say05, SLELA05, SM09, SM07a, SHJ^{+}09, Sch03b, SH03, SC05, SHM00, SZ05, STWH00, SLCJ08, SS02a, Shi06, SCP^{+}02, SS00, SWD00, SFS00, SMH^{+}07, SZS01, Slo02, SFT00, SSL06, SP00, SLCW08, ST05, Sta01, SKP00, SD03, STE^{+}01, SWB^{+}08, SB08, SBP07]. del [SMLB00, SS03b, TS09, TFG08, TVB^{+}06, TFGJ06, TS06, TO02, TRSR00, TPT^{+}07, TK06, Tur06, TKB^{+}06, TUE06, Urb06b, UMT^{+}01, VCHBP08, VSC06, VDHD01, VSH^{+}07, VDP00, Vic02, VTRG00, Wal09, WGW01, WMM04, WSS^{+}04, WCD^{+}08, WCPB03, WELCH09, WCP01, WSCP05, WM05, WD01, WRM05, WU05, WSR^{+}00, WHA^{+}05, WLO^{+}00, WLCS00, WRW^{+}02, WHH00, WH01, WTR01, XGS^{+}09, ZMS^{+}08, ZJ04, ZSG01, BK02]. Deletéreos [JGP09]. Deleterious [Cou02, JGP09]. Delfín [SFT00]. Delfines [BSW^{+}06, Lus03]. Delfinio [KSB01]. D’Elia [WABT08]. Delineation [CS07b]. Delisting [DP01, Ger03]. Delivering [BNM^{+}09]. Delta [NHHG^{+}05]. deltoides [SMG00]. Demand [Cua02, WG01, WSA^{+}05]. Demanda [WG01, WSA^{+}05]. Demandas [LRS09a]. Demencia [MBC03]. Dementia [MBC03]. Demise [Ril00]. Democracy [Mei01]. Democratic [DC06b, IIG05]. Democrática [DC06b]. Demografía [FTR01, GSB04, KHRS01, KHP04, LSL^{+}06, RSCL04, SM09, VBOV05, FND06, MSB^{+}00]. Demografías [Fox07]. Demográfica [BK02, CLBC03, FBL09a, Gar03a, KF02, PSG^{+}00, SL00a, WBSG07]. Demográficas [HRS08, LMD05, MSBGP03, MSWF00, NL05, RSJE06, SM07a, SWK05]. Demográfico [CM06, Cor02, GSG08]. Demográficos [BDSD^{+}05, LE02, MNA07, SFS05, Sie04]. Demographic [Asc02, BDSD^{+}05, BK02, CLBC03, CM06, Cor02, FBL09a, Gar03a, GSG08, HCK^{+}02, HRS08, KF02, KF03, LMD05, LE02, MSBGP03, MNA07, MSWF00, NL05, PSG^{+}00, RSJE06, RSC03, SL00a, SM07a, SFS05, SWK05, Sie04, WBSG07]. Demographics [GSB04]. Demographies [Fox07]. Demography [FND06, FTR01, GVTM07, KHRS01, KHP04, LSL^{+}06, MSB^{+}00, RSCL04, SM09, VBOV05, BS01a]. Demon [Sch07]. Demonstration [HM01, TPW03]. Demostración [HM01, TPW03]. dendrobatidis [BRS^{+}05, KPH07, MSSM09]. Denominador [TMR^{+}07]. Densidad [BS01b, GPC^{+}03, Kal05, LLF08, PH02, PV09a, Rem03, Chi00a, SBB^{+}08]. Density [BS01b, Chi00a, GPC^{+}03, Kal05, LLF08, PH02, PV09a, Rem03, SBB^{+}08, ZSMW06]. Denso [ZSMW06]. Dentro [Ant08, KM00, PIH00, BS00, GSB^{+}06, HLSZ02, RCMV06, VW04]. Depend [DW05b]. Dependen [DW05b]. Dependence [CP00, DGSS06, Ste08a, ZSMW06]. Dependencia [CP00, DH00a, ZSMW06, HK03]. Dependency [DH00a, HK03]. Dependent [BV04, GC07, HH00, HKM06, LLF08]. Dependiente [BV04, LLF08]. Dependientes [GC07, HH00, HKM06]. Depleted [PPBW08]. Depletion [Rom03]. Depredación [BS01b, BEM^{+}04, GC07, HAJ^{+}04, MMB^{+}03, NTGT03, OWOF03, STC03, TNTH^{+}04, Vic02, WMS^{+}01, WZD^{+}00, Sch03b]. Depredadores [BSW^{+}08, Ber06, FN04, FMSL08, GBE00, HDA01, HM04, LS09, RL06, TPW03, Sch03b, CTR02]. Depredation [MMB^{+}03, NTGT03, OWOF03]. Depresión [AHR^{+}09, DH00a, ET03, FNFM01, GGI^{+}05, VSAG06, FG00, KHM00]. Depression [AHR^{+}09, DH00a, ET03, FG00, GGI^{+}05, KHM00, VSAG06]. Deprived [PS00a]. Derechos [PP08a, MC09]. Deriva [CNC04, DSJ02]. Derivados [CMS09, Fai08, MNH01, PBSR06]. Derived [CMS09, MNH01, PBSR06, TNTH^{+}04]. Derriten [XGS^{+}09]. Desafíos [FC05, Rig01]. Desapariciones [WSKH08]. Desarrollando [SAM07]. Desarrollar [PC02, SH03]. Desarrollo [BL06, CNPS04, CMBC00, Fer01, GBC^{+}05, JB06, KSC^{+}06, LB05, LF09, ML04, MYFF^{+}07, MLB08, OK01a, PMC^{+}07, PPW09, RMK07, SMM09, SB04, VCHBP08, WSR^{+}00, BL01, Cla01, SDRO01]. Desarrollos [LKG06]. Desatendida [Dun05, SL08]. Descanso [RS02b]. Descentralización [PB09]. Descomposición [JMS06b, LLV^{+}02]. Descortezadora [CMC^{+}01]. desde [FJS04, PV09b, SC09]. Desechos [LGK02]. Desempeño [LBR07]. Desensamble [LLF08]. Desert [Bei02a, Cal02, EMW^{+}04, FBCG00, LCC^{+}07, MNH01, PMBN01, Put05, SS00]. Desert-Dwelling [EMW^{+}04]. Desérticos [VMBC02]. Desertificación [GFTB04]. Desertification [GFTB04]. Desertified [VMBC02]. Deserts [Lei09]. Desfasada [JWB06]. Desierto [FBCG00, LCC^{+}07, MNH01, PMBN01, SS00]. Design [BMP07, CDD^{+}05, Dav05, KCK^{+}08, LARK01, Lov02, Lyn06, RKEJP06, Ste08a, SBP07, WELCH09, WU05]. Designación [HH06a]. Designatable [Gre05]. Designatables [Gre05]. Designation [HH06a]. Designed [MKH07]. Designer [Web04]. Designing [BBBM02, BMS08, BVL^{+}03, DJOW07, FNS^{+}03, HRS00, KFSR^{+}06, KDC^{+}06, LZD^{+}03, Nel09, RCL^{+}06, WMD^{+}04]. Designs [Kri04, Nos03]. Desove [SCD^{+}08]. Desplazamiento [SSFL08, MC09]. Desplazamientos [DDN01]. Después [BJTD03, CAFL03, FMP^{+}05, GV02, MRM^{+}05, PDF^{+}00, ZSSM09, BB06, LRS^{+}09b, Pop08, Nos01a]. Destination [CMS09]. Destino [CRD05, CMS09, FR07, GAF00, PAHM09]. Destinos [JTS06]. Destrucción [DSJ02]. Destruction [DSJ02]. Destructive [Gil01a]. Desventajas [WELCH09]. Desviación [RFD01]. Desviaciones [FNP09]. Detail [McS03]. Detailed [SS08b]. Details [Bei02a]. Detección [DDN01, NFM04, NL05, RLC^{+}00, CP00, Fat06, HW08, Kai03, MBL02, Pol06]. Detect [HY03, NL05, ZLM03]. Detectabilidad [Kér04, PD03a, SLCJ08]. Detectability [Kér04, PD03a, SLCJ08]. Detectadas [RFB09]. Detectar [HY03, LTC^{+}08, ZLM03]. Detecting [CP00, Fat06, HW08, Kai03, LTC^{+}08, MBL02, Pol06, RLC^{+}00]. Detection [DDN01, NFM04]. Deterioration [BJE^{+}05, Tay03a]. Deterioro [BJE^{+}05]. Determinación [AOGMS09, CPD^{+}03, LCS04, HHD^{+}08]. Determinantes [MKK^{+}06, FN02a]. Determinants [FN02a, MKK^{+}06]. Determinar [BPL^{+}01, GM01, Hay09b]. Determinari [AP00]. Determination [Col00, SP09]. Determine [Hay09b]. Determining [AOGMS09, CPD^{+}03, Don06, HHD^{+}08, LCS04]. Determinism [Col00]. detrás [NMB09]. Detrimento [FVS03]. Detriments [FVS03]. Deuda [BJ05, RRH^{+}09, CNPS04, HO02]. Develop [PC02, SH03, Sch06]. Developing [Ano00d, AFK^{+}09, CNPS04, CMBC00, Cla01, LB05, RMK07, SAM07, SDRO01, VP06, WWH03]. Development [BBB^{+}07, BL01, BL06, Dav02, DR03, Fer01, Fol06, IB03, JB06, KSC^{+}06, LF09, ML04, MYFF^{+}07, MLB08, O’C00, OK01a, PMC^{+}07, PPW09, SMM09, Sey04, SB04, VCHBP08, VP06, WSR^{+}00, GBC^{+}05]. Developments [LKG06]. Deviations [FNP09]. Devices [LCS03]. Devil [McS03]. Devotion [SS08b]. Diadema [PMKDS01]. Diagnoses [VJ09]. Diagnosing [PBN^{+}04]. Diagnosis [PBN^{+}04, Wal00]. Diagnóstico [Wal00]. Dialog [AO06]. Dialogues [Ste05]. Dice [BADF^{+}05]. Dichotomy [San06]. Dictating [WG08]. Didham [ED07]. Die-Off [HBA09]. Dietsch [RKR03b]. Diferenciación [DW05a, Ber05]. Diferencias [AS01, DAKB08, GEJK05, Olu00, TB04, LCZ^{+}05, RB06, WBW05]. Diferente [CF01]. Diferentes [ADF04, BMM^{+}01, BFH^{+}06, CWC00, DWN01, KS03, KBS^{+}04, KLH08, OBK^{+}04, VSGS02]. Difference [Kar01, Lyn08]. Differences [AS01, DAKB08, GEJK05, LCZ^{+}05, Olu00, RB06, TB04, WBW05]. Different [AFK^{+}09, BMM^{+}01, BFH^{+}06, CWC00, DWN01, KBS^{+}04, KLH08, OBK^{+}04, RS01, Sal07, VSGS02]. Differentiation [Ber05]. Differing [CF01, KS03, Pem01]. Dificultades [LHG^{+}03]. Dilemas [Wal09]. Dilemmas [Wal09]. Dimension [Mas03, VNP06, VNP06]. Dimensions [BCK^{+}09, Har05, ODD05, Tre04]. Dimmick [You01]. Dinámica [BJE^{+}05, Cal06, Cre01a, FMDK09, FRSG^{+}02, FES01, GC06, GSG08, HWH^{+}02, LG05a, MSSM09, Pop08, SRW06, SGSA07, SM01, SLH00, WRM05, AE02, Bor05, FH06, LLH08, LP01, Mur01, SMH^{+}07, RLXL01]. Dinámicas [LMP00, MT01]. Dinámico [ARMH04]. Dinámicos [CPSS03, JTS06, LMG06, WBV^{+}05]. Dirección [Asq01, BRR^{+}02]. Directing [BRR^{+}02]. Directrices [Hut06, MRAF^{+}07, RMK07, TG05, KZSV05, PS06]. Dirigida [MSN06]. Disappearances [WSKH08]. Disappearing [Ewi09]. Disassembly [LLF08]. Disaster [WLS^{+}08]. Disciplina [MSN06]. Discipline [Cha08, Laf03, Mef01b, MSN06]. Disciplines [KT06]. Disconnects [SCJ^{+}04]. Discontinuity [Sta09]. Discourse [EN07]. Discovered [ABE01]. Discovery [Des06]. Discrepancias [BM05b]. Discrepancies [BM05b]. Disease [AKO03, BBB08, Cal06, EHJ^{+}04, GGI^{+}05, JLZW03, MSSM09, OK00, Sch10, SZS01, SSL06]. Diseases [APSB02, CMK^{+}07, PJNA07, RHA^{+}05]. Diseñados [MKH07]. Disenchanting [Rol06a]. Diseño [BBBM02, BMP07, BMS08, CDD^{+}05, Dav05, HRS00, KFSR^{+}06, KCK^{+}08, LARK01, LZD^{+}03, Lyn06, Nel09, Nos03, RKEJP06, SBP07, WELCH09, WU05, BVL^{+}03, DJOW07, FNS^{+}03, KDC^{+}06, RCL^{+}06, WMD^{+}04]. Diseños [Kri04]. Disentimiento [NF05b]. Disminución [BFMM02, BS00, Sta01, PBN^{+}04, YLR^{+}01]. Disminuciones [BR01, KM00]. Disminuida [PPBW08]. Disparate [Fox07]. Dispares [Fox07]. Dispersados [CH01, SH09]. Dispersal [BBK02, BTGS05, CW02, FJ07, KWvH08, KFA^{+}03, MK09, PDF^{+}00, PJ00, PV09b, SHM04, SLH00, TCSD00, TR05, VK07, WHA^{+}05, WZD^{+}00]. Dispersed [Baw06, CH01, SH09]. Disperser [PS00a]. Dispersión [BBK02, BTGS05, CW02, FJ07, KWvH08, KDG^{+}06, KFA^{+}03, MSOM^{+}06, MK09, PDF^{+}00, PJ00, PV09b, SHM04, SLH00, TCSD00, TR05, VK07, WHA^{+}05, WZD^{+}00, GGGS00]. Dispersor [PS00a]. Dispersos [GLF^{+}08]. Displacement [MC09, WHM01]. Disponibilidad [CMC^{+}01, CDM09, ESSGL00, GPMM08, HSR05, PJ00]. Dispositivos [LCS03]. Disrupción [CW02]. Disrupted [CW02]. Disruption [Pat00, TR05]. Disruptions [KR00]. Dissected [CS07a]. Dissent [NF05b, PPP06]. Distance [Ber04a, LBG03, NFB05, VK07]. Distances [RS02b]. Distancia [Ber04a, LBG03, NFB05, VK07]. Distancias [RS02b]. Distant [DMS03]. Distantes [DMS03]. Distemper [HLSZ02]. Distinct [ADF04, PD00]. Distinction [DW05a]. Distinguiendo [LM07]. Distinguishing [LM07, Mcs06]. Distintas [RS01]. Distractions [She01]. Distribución [Car07, GPZ04, GMA07, GBH^{+}09, HP08, HRS^{+}07, LM05, LHG^{+}03, LO02, MBMLJ09, MKK^{+}06, MKK^{+}09, MCM^{+}00, PLBBJC09, PSGA09, RBK^{+}05, SPM05, SBDD05, SLS05, SF05, TNFH07, VVMV09, VJ07, VMV07, VO03, WABT07a, ACMA04, GH02, HKN01, HL07b, LCC^{+}07, RAR^{+}04, RK05]. Distribucional [SYWR03]. Distribuciones [RWM^{+}06]. Distributing [LO02]. Distribution [ACMA04, BRTW07, Cri00, GPZ04, GMA07, GBH^{+}09, HKN01, Hew04, LM05, LCC^{+}07, LHG^{+}03, MBMLJ09, MKK^{+}06, MKK^{+}09, MCM^{+}00, PLBBJC09, RBK^{+}05, RAR^{+}04, RK05, SBDD05, SLS05, SF05, VJ07, VMV07, VO03]. Distributional [SYWR03]. Distributions [GH02, HL07b, RWM^{+}06, SPM05, TNFH07]. Distributiva [Ros08]. Distributive [Ros08]. Disturbance [Ano00b, BMA^{+}08, BSW^{+}06, BS08, FES01, GN00, GBMH07, HH00, HH06b, JSWS00, KKHS05, KGM^{+}02, Kes01, KFBG08, KL06, Lai00, MNRD^{+}08, MP09, MEA00, MP07b, New06, PAHM09, RS02b, SSSK06, SMG00, SBB^{+}08, SH02, SHYH06, WSQP07, WD03, WGK07, NL06]. Disturbance-Mediated [MP07b]. Disturbance-Prone [SHYH06]. Disturbance-Response [MP09]. Disturbances [GGC03]. Disturbed [MP09]. Disturbio [MP09]. Divergent [JTS06]. Divergentes [JTS06]. Divers [GPZ04]. Diverse [Bur04]. Diversidad [ALR05, ARP05, ARPEBM09, BTM^{+}05, BNP^{+}07, BGWE00, BMMW07, CPR05, CPO03, CSM^{+}06, DNZK02, ES04, ES08, Fai08, FTC04a, FPF00, FPL^{+}04, GFCP03, GRH05, GCV03, GV02, GWM^{+}07, HIS03, Hie05, HLJV07, HF04, HDE00, JG08, JHHL04, KFSR^{+}06, Kom03, KFNB09, KT02, LFR^{+}08, LH04, LM07, LF00, LJMR^{+}01, MR00a, Mar05, MR00b, MD04, MGC01, MA08, MK09, MSBOF^{+}09, NFM04, NC03a, PLBBJC09, PEC01, RF03, RMSG^{+}03, RAR^{+}04, REC^{+}00, RTCR03, SKBY03, SF03, SB04, SHYH06, SBVC03, SAA^{+}09, TMR^{+}07, TK06, VSAG06, WG06, WHH00, YKIO08, ADF04, CTNR02, GSR^{+}04, JSJ05, KSB01, MEA00, PMEH05, SH02, WU05]. Diversion [RFD01]. Diversity [ADF04, ALR05, ARP05, ARPEBM09, BTM^{+}05, BB01, BNP^{+}07, BGWE00, BMMW07, Cal02, CPR05, CLH^{+}04, CPO03, CTNR02, CSM^{+}06, DNZK02, ES04, ES08, FRH04, Fai08, FTC04a, FPF00, FPL^{+}04, Fle03, GMCP09, GFCP03, GRH05, GCV03, GV02, GSR^{+}04, GWM^{+}07, HIS03, Hie05, HLJV07, HF04, HDE00, JG08, JHHL04, JSJ05, Kai03, KFSR^{+}06, Kom03, KSB01, KFNB09, KT02, LFR^{+}08, LH04, LM07, LF00, LJMR^{+}01, MR00a, Mar05, MR00b, MEA00, MD04, Mei03, MGC01, MA08, MK09, MSBOF^{+}09, NFM04, NC03a, Ode09, Orr03b, PLBBJC09, PMEH05, PEC01, RF03, RMSG^{+}03, Rob06a, RO04, RAR^{+}04, REC^{+}00, RTCR03, SKBY03, SF03, SB04, SH02, SHYH06, SBVC03, SAA^{+}09, TMR^{+}07, Tem02, TK06, VSAG06, WG06, WU05, Wil02b, WHH00, YKIO08]. Diversos [Bur04]. Divide [Cze07a, Far04, Wri07]. División [GCV03]. Dixon [Sha01]. DNA [BST^{+}05, COW03, Des06, JGL^{+}02, PBSR06, RLC^{+}00, Rub06a, Rub06b, UWB04, WCD^{+}08]. Do [Ebe04, Pér05b, Tal04]. Doble [DBB^{+}09a]. Dobles [ARB^{+}07]. Documentación [Mcc09]. Documenting [Mar00b, Mcc09]. Documentos [PDS^{+}07]. Dodd [TM02]. Does [Dal00b, Kot00]. Dog [CB01b, Lei05, LADM05]. Dogs [Cre01a, FND06, Kot00, MRH^{+}00, Shi06]. Doing [FO06, GMCP09, KCR^{+}08]. Dollar [Cze03]. Dolphin [Kle06, SFT00]. Dolphins [BSW^{+}06, Lus03, Wad02]. Domain [LOM03]. Domestic [FND06, Rua00]. Domesticación [PLL^{+}05]. Domesticated [McN03]. Domestication [PLL^{+}05]. Domésticos [FND06, HN09, RSJE06, Rua00]. Dominados [LLS08, WMD^{+}04]. Dominant [GSB^{+}06]. Dominantes [GSB^{+}06]. Dominated [LLS08, WMD^{+}04]. Dominican [Lat00]. Dominicana [Lat00]. Doñana [RPD01]. donde [MH02]. Donlan [LSF04]. Don’t [Hun07a, VP04]. Doomed [PLB09]. Doomed-Egg [PLB09]. Dorsinegro [Mur01]. Dorso [YH01]. Dosel [NH09, WCP01]. Double [ARB^{+}07, DBB^{+}09a, LS04]. Double-Edged [LS04]. Down [AAO^{+}09, Bre01]. Downsizing [Haf08]. Downstream [MO04]. Dragonfly [PKM00]. Drainage [JGM03a]. Drainages [DDBA^{+}07]. Dread [Caf00b]. Dream [Swa09]. Dreaming [Lov04]. Drenado [JGM03a]. Drift [CNC04, DSJ02]. Drink [Per02]. Driven [BS08, MBMLJ09, MSN06]. Driver [VK07]. Drivers [LM07, RDAL09]. Drives [SB04]. Driving [DLCV00, NMB09]. Drosophila [DH00a, KLH08]. Drought [GSB^{+}06, LW01, LSC07, WLO^{+}00]. Dry [ANW^{+}08, CWL^{+}00, CQLF02, FLQ03, HDQSL06, LK04, SAKQ^{+}05, SOADMR08]. dudleyana [Pod01]. Due [Lee03, LC07b, SRW06, SRHS09, SFT00]. Dulce [EHJ^{+}04, HBKF05, HGH^{+}03, PHW05, PFL01, SMV02]. Dulceacuícola [BBH06]. Dulceacuícolas [DDMMZD06, KDL07, KZA09, LM07, PLBBJC09, Rue05, ST05]. Dulceaucícolas [SG04]. Dung [FH05, LLF08, LMP02, PMEH05]. Durant [DBML07]. Durante [Dun04, For02, PDF^{+}00, WLO^{+}00, GSB^{+}06, LSC07, PDS^{+}07]. During [Dun04, GSB^{+}06, For02, LSC07, WLO^{+}00]. Dwelling [EMW^{+}04, HLG03, MY02]. Dynamic [ARMH04, CPSS03, JTS06, LMG06, WBV^{+}05]. Dynamic-Landscape [WBV^{+}05]. Dynamics [AE02, Bor05, BJE^{+}05, Cal06, Cre01a, FMDK09, FRSG^{+}02, FES01, FH06, GC06, GSG08, HWH^{+}02, LG05a, LLH08, LMP00, LP01, MT01, Mur01, MSSM09, Pop08, RLXL01, SRW06, SGSA07, SMH^{+}07, SM01, SLH00, WRM05].

E.U.A [IF02, NHHG^{+}05]. E.U.A. [ALH^{+}04, BRK^{+}04, BJB06, CHBK02, CMCH00, Dal05, DGC05, EG07, Fal07, GSB04, HBC^{+}06, HWH^{+}02, KU02, KB08, PCHN04, RHS05, Rap06, RDN06, RWS02, SF03, SD01, VDP00, VNP06, WM05, WRW^{+}02, WWR^{+}02, ZMS^{+}08]. Eagles [FNP09]. Early [LVM02, LW02b, MRP07]. Earth [Kni07, Leo04, Put05, RLK^{+}01, Van05]. Earthquake [WLS^{+}08]. Earthworm [GJD05, HFR07, MNB09, NMB09]. Earthworms [Gun02, SZBR03]. Easements [MHGF04, RLC^{+}07]. Easier [Lim06]. East [CCS^{+}03, GCF^{+}07, RLC^{+}00, Sey04, VSK01]. Eastern [Mur01, SH09, SBVC03, CGR^{+}05, ES04, EM03, GV02, GK06, GNR01, IMD01, KPH07, LF00, MRM^{+}05, PJYO04, Sta01, VDP00, WELCH09]. Eat [Bea09b, GBB^{+}08]. Eating [WBSB09]. Echoes [Par01]. Eco [Caf00b]. Eco-Dread [Caf00b]. Ecolabeling [KEJ06]. Ecología [BHMC01, Cun00, FND06, Gra03, JMH01, LF09, RH01, Wit02, YL05, Cra02, WA05]. Ecológica [ACMA04, Ber07, BFMM02, BVL^{+}03, BC08, FLF04, FD00, Gav04, HIS03, HCZ00, Kau04, KBS^{+}04, LB05, LSF03, MC05a, MYH09, PPLS01, RAG03, SEBM03, ZR00]. Ecological [AGBP09, Ara02, ACMA04, Bat04, Bee09, Ber07, BFMM02, BFMR07, Bra03a, BVL^{+}03, BC08, CMP^{+}00, DDMMZD06, DM04a, Dre05, DW05b, EAR^{+}05, FLF04, FPW01, FP07, FD00, Gav04, HIS03, Hay04, HCZ00, JP09, JSW08, Kau04, KFL08, KJSH08, KBS^{+}04, KSB04, LB05, LRW03, LSF03, LO02, LARR03, MDF05, MKK^{+}06, MKK^{+}09, McB02, McC01, MCG^{+}09, MC05a, MYH09, MBB^{+}08, NKW09, NL06, PPLS01, Ped07, PA08, PR03, PBRG07, PAHM04, PW01, RNP^{+}07, RS02a, RAG03, SRSP01, SRW^{+}08, Sar06, SEBM03, TR05, Tre04, TF00, VSK01, WTEK08, WDA^{+}06, Wie06, WMVK03, ZR00]. Ecological-Economic [WDA^{+}06]. Ecological-Niche [DDMMZD06, PR03]. Ecologically [For00a, LMB00]. Ecológicamente [For00a, LMB00]. Ecológicas [Bat04, BFMR07, Bra03a, FPW01, JP09, JSW08, KFL08, MBB^{+}08, NKW09, PA08, RS02a]. Ecológico [AGBP09, DDMMZD06, Dre05, LARR03, MDF05, PW01, RNP^{+}07, SRW^{+}08, WDA^{+}06, PR03]. Ecológicos [CMP^{+}00, DW05b, KJSH08, KSB04, LN06, LRW03, LO02, MKK^{+}09, MCG^{+}09, PBRG07, PAHM04, TR05, WTEK08, WMVK03, MKK^{+}06]. Ecologistas [MC05a]. Ecologists [MC05a]. Ecology [Ano07e, Bal02, BHMC01, Cla09, Cra02, Cun00, Cur03, FND06, Gra03, HMH03, Jon04, JMH01, Lam02, LF09, Par01, RH01, RD05, Ste08b, Whi02, Wit02, WA05, YL05, Yul06]. Economía [DC06b, HCM03, PCHN04, SM01, NRBM05]. Economías [Pow06]. Economic [Ber06, BMV09, BV00, BD05, CY08, Cze08, Dal09, Fol06, HM08, HPG09, KFL08, KJSH08, MLMB08, Mef08, NA01, PPLS01, PMC^{+}07, PBRG07, Pow06, Ros06, Ros08, Tas09, VP06, WDA^{+}06, WL04]. Económica [BD05, HM08, HPG09, KCK^{+}08, MLMB08, NA01, PPLS01]. Economically [Gil01a]. Económicas [Ber06, BV00, PMC^{+}07, Pow06]. Económico [CY08, Cze08, Ros06, Ros08, WDA^{+}06]. Económicos [BMV09, KFL08, KJSH08, PBRG07, WL04]. Economics [And01c, HCM03, Led06, NRBM05]. Economies [Pow06]. Economists [Car02b]. Economy [DC06b, FP07, PCHN04, SM01]. Ecoregion [Mag04, MNH01, SD01, MNH01, SD01]. Ecoregiones [JW02, PLBBJC09]. Ecoregions [JW02, PLBBJC09, WDL^{+}02]. Ecorotulado [KEJ06]. Ecosistema [BMV09, BMB^{+}08, BRA08, EB09, GSB^{+}06, GM08a, HH08, HAG^{+}09, JLF09, KSP00, LLV^{+}02, LBTS05, MBM04, MLB08, MCM07, Ped03, SMM09, SMH^{+}07, Tow03, WD01, WGK07, ZSG01, NCVBW02]. Ecosistemas [Cha06, CA06, DC05, EK08, FES01, GWJ^{+}07, HSMK07, KS05, KB06, KB08, MBB^{+}08, Nor08, Ped03, RR00, Rig01, SD01, SRHS09, TAJ04]. Ecosistémicos [BRV06]. Ecosystem [ACD^{+}07, Ber04b, BMV09, BRV06, BMB^{+}08, BRA08, Cha06, DC05, EB09, EK08, Gha07, Gha08, GSB^{+}06, GM08a, Hal02, HH08, HH09, HAG^{+}09, JLF09, Kar09, KB06, KSP00, Kit07, KB08, KDKS08, LLV^{+}02, LN06, LBTS05, MBM04, MB09a, MLB08, MCM07, MBB^{+}08, Nor08, Ped03, RA09, Rig01, RD05, Sal05, SMM09, SMH^{+}07, TAJ04, Tow03, WD01, WGK07, ZSG01, Cor06, NCVBW02]. Ecosystem-Based [Kar09]. Ecosystem-Scale [GM08a]. Ecosystem-Service [ACD^{+}07]. Ecosystems [BHW01, CA06, CLH^{+}04, Cur07, FES01, GWJ^{+}07, HSMK07, HBLW09, KS05, KF01b, MDJP04, OS07, RR00, SD01, SRHS09]. Ecotone [HKN01]. Ecotono [HKN01]. Ecuación [PV09a]. Ecuador [CFC02]. Ecuatonano [CFC02]. Ed [SC09]. Edades [HY03]. Edge [BB04, CFC02, CP01, DMSP00, DD04b, GVS09, HMB^{+}05, MMLM01, MJ03, OBM09, Orr05b, PSBG05, PRMV06, RPD01, SG01, TM04, WGGS07]. Edge-Core [RPD01]. Edge-Related [OBM09]. Edged [LS04]. Edges [BBK02, BVH^{+}02, ED07, FTR01, SRH^{+}07, SLS05, WCP01]. Editor [Ano01b, Ano01c, Ano02c, Ano04c, Ano05d, Ano05e, Ano05f, Ano06d, Ano07i, Ano08i, Ano09i, Fie09, Mef03a]. Educación [Cla01, HSDS01, Hay09a, KSC^{+}06, ZKNL02, TRCC^{+}05]. Education [Aco00, BRD03, Bre01, Bre02a, Bre02b, Bre03, Bre06, Bri06, Cla01, HSDS01, Hay09a, KSC^{+}06, Mai04, Pér05b, Ste07, TRCC^{+}05, ZKNL02, TOR^{+}04]. Educational [WMCB02]. EE.UU. [BJTD03, SDRO01]. Eels [CTNR02]. Efectas [NRB00]. Efectiva [KDC^{+}06]. Efectivadad [CW00]. Efectivamente [MCG^{+}09]. Efectivas [FNS^{+}03, Nel09]. Efectividad [ALM07, Bey04, Cle06b, CLŽ^{+}07, FE07, HDA01, KTN^{+}09, MH01a, MBB^{+}03, RPD01, SCM^{+}07, TWS06, XM07, BMP07, DOW^{+}06, FBP^{+}09, Gut05, KBS^{+}04, NC03a, PK01, RCMV06, SJF^{+}05, SEBM03, TK06, VTR^{+}09, WMPC08]. Efectivo [AT05, CNC04, CMBC00, FPF00, KW02, SRR00, SM07a]. Efecto [BB04, Cal06, Cho06, Cre01a, ED07, FD00, GMA07, HM01, IML^{+}05, Koe03, MMLM01, MYFF^{+}07, PM07, RB06, Sch03b, SBP07, AKO03, BGC^{+}04, BBK02, Ber05, BTGS05, Bor08, BJE^{+}05, CP01, DJC^{+}07, FTR01, HF04, JMLG01, Ram01, SRR00, SBDD05, Shi06, VDHD01, WCPB03, WGK07]. Efectos [ASD01, BBB08, BM05a, Ber06, BRR^{+}02, CVS^{+}09, Car07, DBL06, DDBA^{+}07, DD04b, FBL^{+}09b, GS02, HP04, HIS03, JGP09, JTZ^{+}08, Kai03, KLH08, LH04, LARK01, MBQAM06, MRP07, MWBP07, MN03, NT01, PRMV06, PC07, Pyk00, RO08, RLK^{+}01, Say05, TTB03, TS06, TR05, TF00, WSQP07, WHJJ04, XGS^{+}09, AT05, ARB^{+}07, ANW^{+}08, BCS00, BBM04, BEB00, BRBMS06, BV01a, Bra06, BOC00, CWL^{+}00, CWH07, CF01, CQLF02, CWNW01, CCC04, CBG^{+}00, CMC^{+}01, CMGSM00, CPO03, CCM00c, CNM01, Cou02, Cua00, Cun00, CLC^{+}08, DMSP00, DOGN08, Dav07, DDK05, DS01, DNZK02, DGP^{+}02, EM03, EGPS04, EMW^{+}04, FMDK09, FH05, FFHW03, FLQ03, GC07, GK06, GWJ^{+}07, GPC^{+}03, GSB04, GVTM07, GFTB04, HH00, Has00, HR04b, HCRM09, HRW^{+}03, HDQSL06, HVR08, HAG^{+}09, HFR07, HBSD04, HNG04]. Efectos [IMD01, JF04, JSWS00, KHRS01, KWvH08, KGM^{+}02, KS03, KFBG08, KSDBT02, KHHK03, KD05, Kri04, KA03, Lai00, LSL04, LBPBG07, LLWCW09, LW02a, LEPH02, Lew01, LAN05, LF00, LRS^{+}09b, LO02, Lus03, ML04, Mar05, McK02, ML02, MNL03, MMF^{+}05, MNA07, MJ03, MO04, MY02, MK09, NA01, NAC^{+}06, PRL^{+}02, PSBG05, Per00, Per01, PMZ06, PIH00, Pyk04, Rem03, RGF^{+}04, RPD01, RSF^{+}03, RBK^{+}05, RFD01, Rut01, SMM09, SLELA05, SG01, SF03, SLCJ08, Sie04, SLS05, SG04, SOADMR08, SS03b, TM04, TNRJ02, TH03, TSP^{+}08, TO02, VCHBP08, VS09, VSC06, VJ07, Vic02, WMM04, WR01, WG06, WRM05, WGGS07, WLCS00, WD03, WRW^{+}02, WH01]. Effect [AKO03, BB04, BGC^{+}04, BBK02, Ber05, BTGS05, BJE^{+}05, CP01, Cal06, CF01, Cho06, Cre01a, DJC^{+}07, ED07, FD00, GMA07, HM01, HF04, IML^{+}05, Koe03, Lew01, MMLM01, MYFF^{+}07, NRB00, PM07, Ram01, SRR00, SBDD05, Shi06, SBP07, VDHD01, WGK07]. Effective [AT05, BSW^{+}08, CNC04, CMBC00, DJOW07, FPF00, FNS^{+}03, HM08, HM09, KW02, Kel01, KDC^{+}06, Nel09, SRR00, SM07a, Ste08a]. Effectively [MCG^{+}09]. Effectiveness [ALM07, BMP07, Ber07, Bey04, Cle06b, CW00, CLŽ^{+}07, DOW^{+}06, FE07, FBP^{+}09, Gut05, HDA01, KTN^{+}09, KBS^{+}04, MH01a, NC03a, PK01, RPD01, RCMV06, SJF^{+}05, SCJ^{+}04, SEBM03, SCM^{+}07, TK06, TWS06, VTR^{+}09, WMPC08, WSPL09, XM07]. Effects [AT05, ARB^{+}07, ASD01, ANW^{+}08, BBB08, BM05a, BCS00, Ber06, BBM04, BEB00, Bor08, BRBMS06, BV01a, Bra06, BOC00, CWL^{+}00, CVS^{+}09, CWH07, Car07, CQLF02, CWNW01, CCC04, CBG^{+}00, CMC^{+}01, CMGSM00, CPO03, CCM00c, CNM01, Cou02, Cua00, Cun00, CLC^{+}08, DMSP00, DBL06, DOGN08, Dav07, DDK05, DS01, DDBA^{+}07, DNZK02, DD04b, DGP^{+}02, EM03, EGPS04, EMW^{+}04, FMDK09, FH05, FTR01, FFHW03, FBL^{+}09b, FLQ03, GFCP03, GC07, GS02, GK06, GWJ^{+}07, GPC^{+}03, GSB04, Gre09, GVTM07, GFTB04, HH00, HP04, HIS03, Has00, HR04b, HCRM09, HRW^{+}03, HDQSL06, HVR08, HAG^{+}09, HFR07, HBSD04, HNG04, IMD01, JF04, JGP09, JMLG01, JSWS00, JTZ^{+}08, Kai03, KHRS01, KWvH08, KGM^{+}02, KS03, KFBG08, KB05, KSDBT02, KHHK03, KD05, KLH08, Kri04, KA03, Lai00, LW06, LSL04]. Effects [LBPBG07, LLWCW09, LH04, LW02a, LEPH02, LARK01, LAN05, LF00, LRS^{+}09b, LO02, Lus03, ML04, Mar05, MBQAM06, MRP07, McK02, ML02, MNL03, MWBP07, MMF^{+}05, MN03, MNA07, MJ03, MO04, MY02, MK09, NA01, NAC^{+}06, NT01, NL06, PRL^{+}02, PCD^{+}03, PSBG05, Per00, Per01, PMZ06, PRMV06, PC07, PIH00, Pyk00, Pyk04, RO08, Rem03, RGF^{+}04, RPD01, RSF^{+}03, RBK^{+}05, RO04, RLK^{+}01, RFD01, RB06, Rut01, SMM09, Say05, SLELA05, SG01, Sch03b, SF03, SLCJ08, Sie04, SLS05, SG04, Ste08b, SOADMR08, SS03b, TM04, TTB03, TS06, TNRJ02, TH03, TSP^{+}08, TO02, TR05, TF00, VCHBP08, VS09, VSC06, VJ07, Vic02, WMM04, WR01, WCPB03, WG06, WSQP07, WHJJ04, WRM05, WMD^{+}06, WGGS07, WLCS00, WD03, WRW^{+}02, WH01, XGS^{+}09]. Efficacy [GGW09, VT09]. Efficiency [CWM08, JMS06b, LADM05, MYH09, PD03a, SBP07, Wun07]. Effort [Ben01a, BMB^{+}08, CCM00b, EYZ09, FDBM04, Lyn06, MMBP08]. Efforts [Des06, DN00, Kun03, Rai03, Wil08b, ZR01]. Eficacia [Ber07, GGW09, VT09]. eficaz [MH02]. Eficiencia [PD03a, SBP07, Wun07, CWM08, MYH09]. Efímeros [Mar05, PM05]. Egg [PLB09]. Eggs [TSP^{+}08, VP04]. Ehrlén [MDW01]. Eight [GV02]. Ejecución [Fer01]. Ejemplo [ESB05, FNFM01, KFA^{+}03, PEC01]. Ejemplos [APCS05]. Elapid [RS02a]. Elápidas [RS02a]. Elasticidades [WG01]. Elasticities [WG01]. Elasticity [EVD01, MDW01]. elatior [VEST02]. Elder [Par03]. Elección [BCS04]. Electronic [Ano03d]. Elefante [COW03, WCD^{+}08]. Elefantes [FPW01, GMW08, RLK^{+}01]. Elegant [Fie04]. Element [GCM^{+}08]. Elemental [WG08]. Elementales [WG08]. Elementos [FPSD09, Sem02]. Elements [FPSD09, Sem02]. Elephant [COW03, McD03a, RLK^{+}01, WCD^{+}08]. Elephants [BH04, FPW01, GMW08]. Eleutherodactylus [DAR08]. Elevados [CHFR09]. Elevar [SGSA07]. Elevation [PA04]. Elevational [LZD^{+}03, SSFL08]. Elgar [Fra01]. Eliminar [DP01]. Elk [CFH^{+}02]. Eloquence [Lei09]. Elton [Par01]. Elusiva [PPW09]. Elusive [BST^{+}05, Fla00]. Elusivos [BST^{+}05]. Embarcaciones [RS02b]. Embodied [HAG^{+}09]. Embothrium [MRP07]. Embrace [Sei04]. Embracing [KC07]. Embriones [PRL^{+}02, SWR00, SKP00]. Embriónica [SWR00]. Embryonic [SWR00]. Embryos [PRL^{+}02, SWR00, SKP00]. Emergence [BFJ04]. Emergente [JLZW03]. Emergentes [RHA^{+}05, RMD^{+}02]. Emerging [JLZW03, KK09, RHA^{+}05, RMD^{+}02]. Emigrantes [RS02c]. Emigrating [RS02c]. Emoción [Wil08b]. Emotion [Wil08b]. Empírica [MLB08, SHM04]. Empirical [HL07a, MLB08, SHM04]. Empirically [CMS09]. Empíricamente [CMS09]. Empiricism [BEH06]. Empobrecimiento [SLKP06, RWRW04]. Empresariales [SCB^{+}01]. Encanto [Fre04]. Encima [MHM09]. Encino [ASD01, GMA07, GBC^{+}04]. Encontrando [Nor08]. Encounter [HMF03, Pol06]. Encouraging [BZD03]. Encuentro [HMF03, Pol06]. Encyclopedia [Kun03]. Encyclopedic [Kun03]. End [Ano08a, Mef09c, Meh05, Say03]. Endangered [AS01, ALH^{+}04, AT05, AMF^{+}01, BB06, BTGS05, Bre02b, BV00, BV01b, BRA08, CK05, CGR01, CHBK02, CMGSM00, Cze05, DZM08, DSE^{+}08, DK06, DDBA^{+}07, DP01, ERB01, Fal07, FSS^{+}00, Gar03a, GS01, GN00, Gre05, Gre09, HBC^{+}06, HPLTH07, HH06b, JWB06, KSP00, KSO^{+}03, KFA^{+}03, KSB01, KDD02, KGM01, LTC^{+}08, LVM02, MKH07, MBQAM06, MPFBH07, NWRD03, PZG^{+}05, PLL^{+}05, RW03, RLXL01, Rya06, Say05, SH03, SDRO01, SLCW08, SB08, Tay03a, Tay03b, Tho00, TS06, TRCC^{+}05, VT09, VNP06, Wal04, WABT07a, WL04, WMPC08, BZD03, HP01, NVP07, RM01, VNP06]. Endangerment [FH09, SSL06]. Endemic [CS06, Dal02, DOGN08, MLN^{+}06, MBQAM06, Mic08, PCL^{+}03, Pod01, RHA^{+}05, SWD00, TR05, WHH00]. Endémica [CS06, Pod01, TR05, WHH00]. Endémicas [DOGN08, MLN^{+}06, PCL^{+}03, SWD00]. Endémico [MBQAM06, RHA^{+}05]. Endemism [CRD05, DAKB08, FRH04, Kes01, LS07, MHM09]. Endemismo [CRD05, MHM09, DAKB08, Kes01]. Endogamia [AHR^{+}09, DH00a, FG00, JRL03, KLH08, LRS^{+}09b, MRP07, VSH^{+}07, ENN^{+}08, JWB06]. Endogámica [JRL03, VSAG06]. Endogenous [GLV05]. Enemies [LB00]. Energético [BHB08]. Energía [HAG^{+}09, PPW09]. Energy [BHB08, Hit07, HAG^{+}09, PPW09]. Energy-Sector [BHB08]. Enfermedad [AKO03, Cal06, EHJ^{+}04, JLZW03, MSSM09, OK00]. Enfermedades [APSB02, BBB08, CMK^{+}07, GGI^{+}05, OK00, RHA^{+}05, SZS01, SSL06, PJNA07]. Enfocado [MPR^{+}01]. Enfocando [VO03]. Enfogue [Cze02b]. Enfoque [BKWB05, CM06, HCRM09, MDCK07, Old03, VMV07, PBC^{+}09]. Enfoques [RCS^{+}03]. Enfrentando [GZK06]. Engagement [Van05, ZMS^{+}08]. Engaging [GMB^{+}08]. Engelman [Mcs06]. Engineer [BRA08]. Engineering [APSB02, Web05]. England [BCHS08, FO06, HBA09, SB04]. Enhance [Ric04b, SGSA07]. Enhancement [VW04]. Enhancing [FV05, ZMS^{+}08]. Enlaces [BMS08, DHZM09]. Enlargement [FE07]. Enlistado [Fal07, MRAF^{+}07]. Enlistar [HBC^{+}06, QR05, RFB09]. Enlisting [HW02]. Enmalles [SFT00]. Enough [Bre01, HRP^{+}04, Wil08a, Woo03]. Ensamblajes [CNM01, FN02a]. Ensamble [ARP05, LLH08, LHPH06, MMF^{+}05]. Ensambles [BMA^{+}08, GFTB04, SOADMR08, VS09, WA05]. Entanglement [SFT00]. Entender [Arm05, LF09, PRL03, Tho05b]. Entendiendo [EHJ^{+}04, JMS^{+}06a, LLF08]. Enterprise [SCB^{+}01]. Entomologists [KK09]. Entre [BHMC01, BBH06, Cur02a, RM01, SMG00, AS01, ALR05, BGHM06, BJ05, BSKH^{+}06, BS00, dLBBG^{+}08, CGR01, Coc01, CW08, Cze08, DHZM09, Elp00, GEJK05, GSB^{+}06, GS08, HPF^{+}06, Hob01, JP09, JMP09, KW02, KF02, KCK^{+}08, LW01, LCZ^{+}05, LM07, LCS04, Mad05, MNRD^{+}08, MKK^{+}09, MMBP08, MA08, NFM04, OBK^{+}04, Olu00, PP08a, Qia07, RS02a, RF03, Ree05, RSCL04, RIBGG08, SSM03, SKBB00, TB04, Tho05b, TK03a, TWW09, WCP01, WG06, Wil08b, ZSMW06]. Entrecruzas [ALH^{+}04]. Entrenamiento [Bau05, GBE00]. Entrepreneurial [RH08]. Environment [DS08, Dud00, ESB05, KBS^{+}04, Ols00, Raf03, SSB^{+}08]. Environmental [CHW^{+}01, Che09, Col00, CW09, DH00a, Ehr05a, EYZ09, Ewi09, FNFM01, FNF03, FP07, For00b, Fra04, Ful02, GGA01, HSDS01, Het04, HBLW09, LME08, LU09, LOM03, LMP02, LSF03, MCG^{+}09, NV09, O’C00, Pal05, PS06, RM03, RP07, SRW06, SAPRB07, SJF^{+}05, SBDD05, SDL07, Sta01, TK06, TRCC^{+}05, WSQP07, Wun07, LME08]. Environmentalism [Vit04]. Environmentalist [Moo02]. Environments [FHB^{+}08, RWJRR06, WMD^{+}04, ZG05]. Eólica [PPW09]. Ephemeral [Mar05, PM05]. Epidemiología [HLSZ02, LG02]. Epidemiologic [WSKH08]. Epidemiológica [CMK^{+}07]. Epidemiological [CMK^{+}07]. Epidemiológico [WSKH08]. Epidemiology [HLSZ02, LG02]. Epifitas [Hie05, KFNB09]. Epifíticos [Pyk04]. Epiphyte [Hie05, KFNB09]. Epiphytic [Pyk04]. Episteme [Kit07]. Equal [RSPM05]. Equating [Smi09]. Equation [PV09a]. Equilibrio [Wil08b]. Equipos [HSDS01]. Equivalencia [Elp00]. Equivalency [Elp00]. Equus [SRW06]. Era [MHS07, SP09]. Eradication [APCS05, APS07, BSW^{+}08, DTCC03, HDG^{+}07, NMT^{+}04, RPM09]. Eradications [CD05]. Eran [BEB00]. Eriocheir [HRS^{+}07]. Eroded [VSAG06]. Erosion [Bey04, BBW^{+}07, JLF09, MO04, Bey04, JLF09, MO04, BBW^{+}07]. Erosionada [VSAG06]. Erradicación [APCS05, APS07, BSW^{+}08, NMT^{+}04, RPM09, HDG^{+}07]. Erradicaciones [CD05]. Errata [Ano01d, Ano01e, Ano07j, Ano07k, RSJE07]. Erratum [Ano01f, Ano02d, Ano06e, Ano08j, Ano04b, Ano07b, Ano07c]. Errónea [Asq01]. Error [Ano02e, MM04, MF00]. Errores [MF00]. ESA [Ger03, Sal02]. Escala [AW02, Arm05, BGD09, BV01a, Cha06, FDL^{+}05, Fje07, FNFM01, GC07, GSR^{+}04, GM08a, HH00, HK03, HBV02, Kai03, KDD^{+}04, LB02, MBR^{+}02, MDCK07, MWEB06, MO04, NFM04, Qia07, RCL^{+}06, SRW^{+}08, SP00, UWB04, WSS^{+}04, VMBC02]. Escala-Dependientes [HH00]. Escalado [KL06, ML02]. Escalares [DAB06]. Escalas [CPD^{+}03, GCV03]. Escapados [VTRG00]. Escaped [VTRG00]. Escarabajo [BRA08]. Escarabajos [Fat06, FH05, HW08, HKN01, HKM06, LLF08, LMP02, MMLM01, PMEH05, TK07]. Escarbajos [DW05b]. Escarpado [ML02]. Escarpment [ML02, MNL03]. Escenario [PCC03]. Escenarios [BRV06, Fai08]. Esclerosis [MBC03]. Escollos [WDA^{+}06]. Esencial [CA05]. Esfuerzo [BMB^{+}08, FDBM04, Lyn06]. Esfuerzos [DN00, MMBP08, Wil08b]. Esnorquel [PIH00]. Espacial [AW02, CP00, CJ08, FP02, HLB^{+}03, LSN07, RKEJP06, SHM04, SBVC03, TNTH^{+}04, WHS02, WHH00, ZGR02]. Espaciales [CPSS03, DSJ02, FSC02, GCV03, KFSR^{+}06, Lai05, LMP02, NLCAR09, SK02a, VTA02]. Espacialmente [KKHS05, RBHKK02, TUE06]. Espacio [BGC^{+}04, BVH^{+}02, SRR00]. Espacios [HSR05]. Espaciotemporal [DJOW07, RLXL01]. España [NWRD03]. Especial [JW02]. Especialistas [FN04, PZG^{+}05]. Especialización [SK04]. Especie [AMF^{+}01, BFB^{+}09, BB06, Bos07, CS06, CMGSM00, CW02, DDBA^{+}07, EKL^{+}00, FJ07, GWS07, GPMM08, GDAV00, GNEE07, HMK^{+}09, KFA^{+}03, Kot00, LMS^{+}02, LDG^{+}00, MSOM^{+}06, NNM00, Old03, Pod01, PMC01, RA04a, Rub01, SRC^{+}02, SMG00, SGH02, Vee00, Vic02, WABT07a, WFP01, WMP05]. Especie-Area [Vee00, WMP05]. Especies [AS01, APCS05, ALH^{+}04, Arm05, AHTM05, ACMA04, ARPEBM09, BBBM02, BMBM06, BMMR03, BSS^{+}05, BOJ09, BTK08, BSKH^{+}06, BEB00, BV01a, BS00, BZD03, dLBBG^{+}08, BOC00, BL06, BMMW07, BV00, BV01b, BRR^{+}02, CWL^{+}00, CP01, CVS^{+}09, CES05, CF01, CKL^{+}00, CPD^{+}09, CDD^{+}05, CHBK02, CMA^{+}06, CCM00c, Cra02, CCH03, Cun00, Cze05, DBL06, DC06b, DK06, DCD^{+}02, DMS03, DFS05, DDMMZD06, DP01, DAKB08, DKS^{+}09, DAB06, ESB05, Ehr06, EM08, ED07, Fal07, Fat06, FEGB09, FMPB07, FPL^{+}04, FNFM01, FTN^{+}05, FBK02, GRB02, GFCP03, GGA01, GS01, GCV03, GV02, GLMN00, GMA07, GM01, GPC^{+}03, GBH^{+}09, Gre05, GG02, GBE^{+}06, Gus02, HBC^{+}06, HRS00, HH00, HR02, HP04, HP01, HW08, HBBD08, HMF03, HL07b, HC09, HFR07, HJ07, HBSD04, HF04, JTS06]. Especies [JWB06, JSW08, JFWP06, JMP09, JMH01, JMS06b, KB04a, KFL08, KTN^{+}09, KD05, KPJ06, KHP04, LBR07, LU09, LTC^{+}08, LRO^{+}08, LBG03, LBPBG07, LLWCW09, LP05, LCZ^{+}05, aARLG03, LMP02, LHG^{+}03, LW02b, LBTS05, MCG^{+}08, MLN^{+}06, MKH07, MP09, McK02, MP08, MD04, MHM09, MRAF^{+}07, MW07, MML06, MBB^{+}08, MSSM09, NA01, NFM04, NF04, NLCAR09, NWRD03, NP06, ND03, NBM^{+}04, OBK^{+}04, OIK^{+}06, PCL^{+}03, PM07, Pra05, PTP^{+}08, PJK03, Qia07, QR05, RO08, RBT08, RNP^{+}07, RM01, RWM^{+}06, RBK^{+}05, RPP^{+}00, RW03, Rub06b, RLXL01, Say05, SGB^{+}08, SPM05, SDRO01, SWD00, Sim03, SSL06, SKBB00, SEBM03, SS06, SD03, SHT^{+}05, SCC^{+}09, SDJK04, SBVC03, TTSP04, TTB03, TNFH07, TNRJ02, TK06, TWS06, UMT^{+}01, VMV07, VTR^{+}09, VSGS02, VNP06, WMM04, Web01, WSCP05, Wen08, WHJJ04, WL04]. Especies [WBBA03, WMLS09, ZJ04, ZLM03, ZGR02, KFK^{+}09, CMC03, CKL^{+}00, Che09, Fai08, KU02, LWSM03, LS01, LSF03, RH09]. Especies-Área [Fat06, Che09]. Específica [ØDEA00]. Específicas [JRL03]. Especificidad [ØDEA00]. Específico [LME08, Old03]. Especimenes [JLZW03, MD04, REC^{+}00]. Espejos [DWWF04]. Esperado [MC05b]. Esperando [RS02a]. Esperanza [MCG^{+}09, Nor08]. Esperanzadoras [MCG^{+}09]. Espontánea [SKBY03]. Esquemas [KBS^{+}04, SSB^{+}08]. Esquincos [BTGS05]. Essays [Kaw04]. Essential [CA05]. está [WABT07a, MHGF04, JWB06]. Estabilidad [FBL09a, MSBGP03, DMG^{+}00]. Estableciendo [KCK^{+}08]. Establecimiento [AFB^{+}00, FDBM04, KS05, PV09b, Rem03, APS07, FDL^{+}05, NKBD02]. Estables [WRL02]. Established [KL05]. Establishing [APS07, FDL^{+}05, NKBD02, WMD^{+}06]. Establishment [FDBM04, KS05, PV09b, YL05]. Estación [SSY06]. Estacionales [Lai05, PB04]. Estadística [DFS05]. Estadístico [FBC^{+}06]. Estado [CPME09, GHL08, RW00]. Estados [CMCH00, DP01, Kee06, LW02b, Ped03, BMM^{+}01, BJPL07b, BPB^{+}05, BL06, CTNR02, CML02, Cur02a, Cze05, DC05, Fau01, For00a, GS05, HRW^{+}03, LB01, MM02a, PH02, PM07, RW00, RLXL01, Sta01, Wal03]. Estamos [KMGB07]. Estándar [SSS^{+}08]. Estanques [JMLG01, PC02]. Estatal [BV01b]. Estate [CPME09]. Estáticos [CPSS03]. Estatus [AM05, Gre05, Hay09b, MRAF^{+}07, NKW09, QR05, QRF07, RRK02, RS02a, VCHBP08, Wil02c, ZJ04, SDJ06]. Este [IMD01, PJYO04, VDP00, WELCH09]. Estepa [HWH^{+}02, VDP00, WRW^{+}02]. Estepas [CGR^{+}05]. Esterilización [LRS09a]. Estiércol [LLF08, PMEH05]. Estimación [JLF09, MD04, ØDEA00, REC^{+}00, SHM04, APCS05, BST^{+}05, CS07b, DFS05, For00a, GS02, HLG03, MYFF^{+}07, PV06, RTCR03, SPM05, SYWR03, UCAGG08]. Estimaciones [AM05, SLCJ08, TCSD00, FHB^{+}08, Kér04, PJW^{+}02, SM07a]. Estimadas [Fai08]. Estimadores [RB06]. Estimar [Ant08, LS09]. Estimasciones [SGSA07]. Estimate [Ant08, For00a, LS09]. Estimated [Fai08]. Estimates [AM05, Ano04e, FHB^{+}08, Kér04, ØDEA00, PJW^{+}02, ROB^{+}04, SGSA07, SM07a, SPM05, SLCJ08, TCSD00]. Estimating [BST^{+}05, GS02, HLG03, JLF09, MYFF^{+}07, MD04, PV06, RTCR03, Sch10, SYWR03, UCAGG08]. Estimation [CS07b, DFS05]. Estimators [RB06]. Estocasticidad [KF02, SK02a, SRW06, Fox07]. Estorninos [Koe03]. Estrategia [Cua00, HCZ00, HDE00, Hun05, MC09, MH01a, NR03, RWBG06, GB05a, ZR00]. Estrategias [BFH^{+}06, FPL^{+}04, HRS00, KU02, Les05, MBM04, MOO09, MCM07, PC02, RBG^{+}04, SCB^{+}01, SMV02, SCM^{+}07, TCSD00, TTSP04, TWGH00, LG05b, PCL^{+}03, VSO04]. Estratégica [HH08]. Estratégicos [KC01]. Estrés [CWNW01, DH00a, GMW08, TG04, ZSSM09, FPW01]. Estresantes [Wen08]. Estructura [APCS05, BD06, Ber05, BFD07, BK02, CP01, CQLF02, CNC04, CY08, CNM01, DMG^{+}00, EM03, HMB^{+}05, HY03, KGM01, MR00a, ML04, New06, Per00, RTN01, SG04, WHS02, WHH00, CMR05, MRP07, PS00a, PJYO04, TS03]. Estructuradas [KMGB07]. Estuarina [Wol00]. Estuarine [LEPH02, RFD01, Wol00]. Estuarinos [RFD01]. Estuario [LEPH02]. Estudio [APS07, ACMA04, BMBM06, BR01, BTGS05, Che09, Cla01, DCD^{+}02, DA07, FVS03, Fat06, GPZ04, GM01, GZK06, HDA01, HW08, HM03, HY03, HLJV07, HW02, JHHL04, JMH01, KZSV05, KU02, LB05, Lin03b, LZD^{+}03, MBM04, MKH07, MCA^{+}00, MS09, New06, ODE01, Ped03, PBK09, PV09a, PJK03, RAL^{+}04, RWJRR06, RWM^{+}06, RG04a, RMK07, RCMV06, Rua00, SH03, SDRO01, SFT00, SD01, TVD03, TPW03, VEST02, WSS^{+}04, WMPC08]. Estudios [DBML07, Fje07, GMB^{+}08, Kér04, LF09, TB04]. Ethic [AFK^{+}09, Sag07]. Ethical [Bei05, DN01, McB02]. Ethics [Ano05k, Caf04, Fra04, Het04, LSF03, MC05a, Pal05, Rol06b, RBG06, Van06]. Ethiopia [FSS^{+}00]. Ethiopian [HLSZ02, HN09]. Ética [MC05a, LSF03, Sag07, Van06]. Éticos [DN01]. Etíope [HN09]. Etíope. [HLSZ02]. Etiopía [FSS^{+}00]. Eubalaena [REC^{+}00]. Eucalipto [LO06].

Eucalypt [LO06]. Eurasian [HAJ^{+}04, RPD01]. Eurasiático [RPD01]. Europa [BE00, BVL^{+}03, CWNW01, CGR^{+}05, DDK05, HRS^{+}07, LBPBG07, PBC^{+}09, PJK03, RLC^{+}00, SHJ^{+}09, TVB^{+}06, Web01]. Europe [BVL^{+}03, CWNW01, CGR^{+}05, DDK05, HRS^{+}07, Ols00, PBC^{+}09, PJK03, SHJ^{+}09, Web01]. Europea [PGP04, Pyk00]. European [RP07, BE00, BKK03, DMDFV08, EHJ^{+}04, FN04, Koe03, LBPBG07, PGP04, PCVC06, Pyk00, RLC^{+}00, SGB^{+}08, TVB^{+}06, VJ07, Wes05]. Europeas [BKK03, SGB^{+}08, VJ07]. Europeos [EHJ^{+}04, FN04, Koe03, PCVC06, DMDFV08]. Eutrofización [BCHS08]. Eutrophication [BCHS08]. Evaluación [AM05, ACC^{+}05, AAO^{+}07, ACMA04, ARA^{+}09b, AHR^{+}09, BD05, CJ08, CMC^{+}00, DDAK04, DD01, FM09, FBCG00, Gre05, HCE^{+}01, HCM03, JHHL04, KDL07, KSP00, KRM^{+}00, LARR03, MDF05, MP09, MCM^{+}00, MLB08, MH08, MCM07, Nor08, OGH04, ODE01, OIK^{+}06, PPBW08, PCFC01, SK02a, SM09, SFT00, SMSB05, TG04, The03, TKB^{+}06, WWR^{+}02, ZSG01, AGBP09, BPB^{+}05, BWS09, BGWE00, BMMW07, CLBC03, DBL06, DN00, EBAW01, FRSG^{+}02, FMP^{+}05, FRR07, GM01, GM08a, HSMK07, HDA01, HGH^{+}00, JSJV03, JSHM06, JML00, LE02, LMG06, LSZ^{+}08, LRS09a, MH01a, MCR^{+}04, MK02b, NKW09, NCVBW02, PPLS01, PMC^{+}07, PLBBJC09, PN05, PP03, PBS03, Pop08, PK05, Pyk05, RO08, SSB^{+}08, SFS00, SCM^{+}07, TCSD00, TWS06, WTEK08, WFP01, WSCP05, ZKNL02, MHC^{+}01, RH08]. Evaluaciones [Ant08, BPV09, DAB06, KDC^{+}06, MDCK07]. Evaluados [OBK^{+}04]. Evaluando [Rub01]. Evaluar [CV08, JAH^{+}05, MU08, RBHKK02, SHM04, TS06, Wal09, WABT07a, WSBP09]. Evaluate [CV08, RBHKK02, Wil03]. Evaluating [BPB^{+}05, BWS09, BGWE00, DBL06, GM01, HSMK07, HDA01, LE02, LSZ^{+}08, MH01a, MHC^{+}01, MCR^{+}04, MU08, MCM07, PMC^{+}07, PN05, PPBW08, PP03, Pri03, Rub01, SCJ^{+}04, SFS00, SHM04, TCSD00, WABT07a]. Evaluation [AGBP09, AAO^{+}07, Bei03, BMMW07, DC06a, DN00, HCM03, JML00, KRM^{+}00, MLB08, ODE01, OIK^{+}06, PCFC01, RH08, SK02a, SSB^{+}08, SRDG03, SMSB05, WD09, ZKNL02]. Evangelical [Van05]. Evangelicals [Hen05]. Even [Pim06]. Event [GMHS00]. Events [Orr02a]. Ever [Bro05, Wad02]. Evergreen [Coc01]. Everything [Bro05, Smi04a, Wad02]. Evidence [BB04, BTGS05, CW02, IML^{+}05, JLZW03, LME08, NT01, OMP^{+}05, PAA^{+}08, PK05, SSL06, TB04, UWB04, VTRG00]. Evidencia [BTGS05, JLZW03, LME08, PAA^{+}08, TB04, UWB04, BB04, CW02, IML^{+}05, NT01, OMP^{+}05, SSL06, VTRG00]. Evidencias [PK05]. Evitando [LHG^{+}03]. Evitar [GBE00]. Evolución [ADF04, BHMC01, LRS^{+}09b, SFD09]. Evolution [ADF04, AL03, BHMC01, LRS^{+}09b, PB08, SFD09]. Evolutionarily [DGG^{+}01]. Evolutionary [HMH03, PRL03, RM06, Ric05, SJP^{+}07, Yul06]. Evolutiva [PRL03]. Evolutivos [RM06]. Evolves [Rub06a]. Evolving [Cri00, Men03, Pea01]. Ewers [ED07]. Ex [CMGSM00, GG06, SPFF^{+}08].

Exactly [Tho09]. Exageración [PC07]. Exaggeration [PC07]. Exámen [SAA^{+}01]. Examinando [AS01]. Examination [GH02]. Examining [AS01, SAA^{+}01]. Example [BC09, ESB05, FNFM01, KFA^{+}03, PEC01]. Examples [APCS05]. Excedente [RL06]. Excel [Wit03]. Excesivo [EG07]. Excessive [EG07]. Exchanges [CW08]. Excluder [LCS03]. Exclusion [CWL^{+}00, LAN05, LB00, CWL^{+}00, LAN05, VT09]. Excluyentes [LCS03]. Excursionismo [SS03b]. Exhaustion [Jac02]. Existencia [Pow06]. Existing [Gut05]. Exit [CDM09]. Exito [BV01a, CPSS03, LDG^{+}00, TR01, Mor00b, AAO^{+}07, BMM^{+}01, RR01, Ant08, BB04, BFH^{+}06, BWS09, CQLF02, DOGN08, DNZK02, DD04b, FLQ03, GPC^{+}03, HHSC00, HRW^{+}03, HDQSL06, Kri04, Ols07, RPM09, Rem03, Shi06, VSH^{+}07, Wal09, Wel06, WH01]. Exitosa [LG05b]. Exitoso [NF04, KS05]. Exogamia [FG00]. Exogámica [ET03, GGI^{+}05]. Exotic [BPB07, CCH03, DN00, DHW01, EB09, GL03, GB03, Gun02, GJD05, Hod04a, HN09, KB04b, LO02, PJ00, PE01, Rem03, Web01, WSCP05]. Exóticas [ACMA04, BRR^{+}02, CWL^{+}00, CCH03, DOGN08, DN00, EB09, Gun02, GJD05, HCRM09, Hod04a, KS03, MS08, PJ00, PA04, PJK03, Web01, WSCP05]. Exótico [BPB07, LO02, Rem03]. Exóticos [Hod04a, HN09]. Expanding [UKC^{+}09]. Expansion [BJPL07b, Tol07, BJPL07b, RHS05, UKC^{+}09]. Expected [MC05b]. Expenditures [RM01]. Experience [TFGJ06]. Experiences [Juk01]. Experiencia [TFGJ06]. Experiment [ASM05, AP00]. Experimental [AP00, BS01b, CS06, CW02, CLŽ^{+}07, FMP^{+}05, GN00, GPC^{+}03, HM01, Lew01, MHJ^{+}09, NT01, PJ00, Pop08, RS02c, SD02, ST05, SS03b, TPW03]. Experimentales [BS01b, CLŽ^{+}07, GN00, Pop08, SS03b]. Experimentally [Ber05]. Experimentalmente [Ber05]. Experimentation [HRN00]. Experimento [ASM05]. Experimentos [DH00b]. Experiments [DH00b, Raf03, Tho05a]. Expert [BTBC^{+}03, CWCG02, O’C00, SBDD05]. Expert-Based [CWCG02]. Experto [BTBC^{+}03, SBDD05]. Expertos [CWCG02]. Explain [Dal00b, MKK^{+}09]. Explaining [Owe08]. Explanation [LSF03]. Explanations [RHA^{+}05]. Explicación [LSF03]. Explicaciones [RHA^{+}05]. Explican [MKK^{+}09]. Explicit [BPP05, GSG08, QAMW03, RBHKK02, TUE06, RT00]. Explícito [GSG08, TUE06]. Explícitos [QAMW03, RBHKK02]. Exploitation [FPM02, LHPH06]. Exploited [GWS07]. Exploiting [LS09]. Exploración [LG05a]. Explored [Bro08a]. Exploring [Hop07, LG05a, McB02]. Explosivos [FMP^{+}05]. Explotación [CPR05, GJLL08, GGC03, LS09, LHPH06, FPM02]. Explotada [GWS07]. Exposed [BSW^{+}06]. Exposición [FND06, MHC^{+}01, SWR00]. Exposure [FND06, MHC^{+}01, SWR00]. Expresada [WBW06]. Expressed [WBW06]. Expuestos [BSW^{+}06]. Extend [CPME09]. Extender [CPME09]. Extending [CHMP05]. Extensa [LSZ^{+}08]. Extensas [Pol06]. Extensión [FDBM04, Had00]. Extensiones [Wit02]. Extensive [AGBP09]. Extensively [BE00]. Extensivo [BE00]. Extensivos [AGBP09]. Extenso [LCZX09, RJW09]. Extent [PL03]. Extinción [ACC^{+}05, ARB^{+}07, AOGMS09, BJ05, dLBBG^{+}08, BJE^{+}05, CB01a, CNPS04, CGR01, Cho06, CMGSM00, DLCV00, DDBA^{+}07, DR02, ES08, Fai08, FN04, Fon09, FH09, Gil01b, GZK06, HO02, HRS08, HP04, HP08, HK03, HGH^{+}03, KKHS05, Kér04, KSB04, KSB01, LE02, LRG^{+}02, Log03, MCG^{+}08, MBQAM06, MKK^{+}06, MRDC06, MF00, Mor00b, OBK^{+}04, PH02, PJNA07, PN05, PCVC06, PJW^{+}02, RS02a, RBM^{+}05, RSJE06, RRH^{+}09, SK04, SM09, SZS01, SSL06, VSAG06, VWQI00, Wil02c, WMP05, WTR01, WG08, ZBABR00, ZGR02, ESB05, OALM^{+}00, WHS02]. Extinciones [Bra03a, BMM^{+}02, DW05a, DMG^{+}00, LU09, MLN^{+}06, RFB09, SSFL08, CSS00, CSBS01, Dun05, GWM^{+}07, RSD03]. Extinct [Ber03, Gil08, MCA^{+}00, Whi06, SMI^{+}07]. Extinction [ACC^{+}05, ARB^{+}07, AOGMS09, BJ05, dLBBG^{+}08, BMM^{+}02, BJE^{+}05, CB01a, CNPS04, Cho06, CE00, CSBS01, DLCV00, DR02, ESB05, EC01, ES08, Fai08, FN04, Fon09, Fra01, Gil01b, GZK06, HPLTH07, HO02, HRS08, HP04, HP08, HK03, HGH^{+}03, KKHS05, Kér04, KSB04, LE02, LRG^{+}02, MCG^{+}08, MBQAM06, MKK^{+}06, MRDC06, MF00, OBK^{+}04, OALM^{+}00, Orr05a, Par03, PJNA07, PN05, PCVC06, PJW^{+}02, RS02a, RBM^{+}05, RSJE06, RSC03, RRH^{+}09, SK04, STT05, SM09, SZS01, SSL06, VSAG06, VWQI00, WBSB09, WHS02, Wil02c, WMP05, WTR01, WG08, ZGR02]. Extinction-Rate [PJW^{+}02]. Extinctions [Bra03a, CSS00, DW05a, DMG^{+}00, Dun05, GWM^{+}07, LU09, MLN^{+}06, OS04, PH02, RFB09, RSD03, RM02, Sam06, San00, SSFL08]. Extinta [MCA^{+}00]. Extirpaciones [MM02a, Wol00]. Extirpadas [HLG03]. Extirpated [HLG03]. Extirpating [San06]. Extirpation [Côt05]. Extirpations [MM02a, Wol00]. Extracción [MR00a, PL03]. Extracting [RS00]. Extraction [MR00a, PL03]. Extraordinaria [GWS^{+}06]. Extrapolación [KG03]. Extrapolation [KG03]. Extraterritorial [NS01]. Extraterritoriales [NS01]. Extrema [GSB^{+}06]. Extreme [GSB^{+}06]. Extremidades [JLZW03]. Exurban [OK01a, The03]. Exurbanas [The03]. Exurbanizing [BJB06]. Exurbano [BJB06, OK01a]. Eye [Wes05].

F [Jar03]. F. [LO02]. Faaborg [VP04]. Fabaceae [FG00]. Face [Ber01, KL06, RBT08, RVVNH09]. Facilitan [DGC05, NHHG^{+}05, Urb06b]. Facilitate [AR09]. Facilitating [DGC05]. Facilitation [Urb06b]. Facilitator [Lak01]. Facilities [PA08, ZSSM09]. Fact [Bal00]. Factor [HHSC00, TB03]. Factores [CMM^{+}05b, Cre01a, LG05b, NWRD03, RDAL09, Rue05, Sta01, Wen08, WSKH08, CDM09, CW00, CAFL03, ET03, FBCG00, Gra01, HR02, MMM03, MS08, SK02a, TG04]. Factors [CMM^{+}05b, CDM09, CW00, CAFL03, Cre01a, ET03, FBCG00, Gra01, HR02, LG05b, MMM03, MS08, NWRD03, Pan09, Rue05, SK02a, Sta01, TG04, WSKH08]. Fail [Ano00d, Kni09]. Failing [Kni06]. Fails [Jac02]. Failure [APS07, BC06, BM05a, Lov02, TSG06]. Failures [HH06a]. Fair [WD09]. Faith [PMEC04]. Faiths [DHZM09]. falciformis [WELCH09]. Falco [ENN^{+}08]. Falcons [TR01]. Fall [Lei05]. Fallacies [Ros02]. Fallacy [Ful00, Wil02d]. Fallas [HH06a]. Fallen [SVBR07]. Fallida [APS07]. Fallow [BDSD^{+}05, ZSSM09]. Falso [FSC02]. Falta [Mcc05, UWB04]. Familia [BS00]. Familiar [Rub06a, Shi06]. Familiares [FB08]. familiaris [HAJ^{+}04]. Family [FB08, Shi06]. Family-Forest [FB08]. Fantasy [McD09]. Farallones [GN00]. Farm [MKDT01]. Farmer [JSJV03, Law06, PD03b]. Farmers [Gro05b]. Farming [BD05, HPLTH07, MV06, Sch00, SMC00, STP09]. Farmland [BJB^{+}02, CLC^{+}08, FBK02, Her09, KB05, SP00, VW04, VW05]. Farmlands [Jan07, JSJ05, MMM03]. Farms [JSJV03, KJSH08, WHJJ04]. fasciculata [FG00]. Fast [MSWF00]. Fate [CRD05, FR07, GK09, GAF00, PAHM09, Rai00, RA04b, Ter00]. Fates [JTS06]. Faulty [DR03]. Fauna [ANW^{+}08, BD05, CPD^{+}03, CML02, CMCH00, Dun04, GLV05, Has00, NTMT^{+}03, QRF07, WD06, BHMC01]. Faunaíctica [CCS^{+}03]. Faunal [CCS^{+}03, CML02, Dun04, FN02a, WD06]. Faunas [BHMC01]. Faunística [FN02a]. Faustian [Bec09]. Fauth [MDJP04, RO04]. Favourable [LNP^{+}09]. Fearful [Sei07]. Feasible [KNH07]. Feather [SS01]. Features [ML04, SKBY03]. February [Ano03f, Ano06h, Ano07p, Ano09n, Ano06i, Ano07q, Ano08o]. Fecal [TG04]. Feces [BST^{+}05]. Fecund [FPL^{+}04]. Fecundidad [EM08, KDL07]. Fecundity [EM08, KDL07]. Fecundos [FPL^{+}04]. Federal [LB01]. Feedback [Edi03]. Feedbacks [LW01]. Feeding [GBMH07, NBM^{+}04]. Feedstock [Koh07]. Female [GMW08]. Fen [HBSD04]. Fencing [JF04]. Fender [SH03, SH03]. Fennoscandia [IMD01, MPE^{+}07]. Fenología [FLQ03, GB01, PC02]. Fenológicos [HDQSL06]. Fenotípica [ADF04, OB00]. Feral [CD05, LEPH02, LRS09a, NMT^{+}04, RPM09, RSW05]. Ferales [CD05, LRS09a]. Fern [Gun02]. Fernandez [DGSS03, DGSS03]. Feromona [LS09]. Feromona-Kairomona [LS09]. Feromonas [LS09]. Ferriss [GCM^{+}08]. Ferrocarril [IML^{+}05]. Fertility [EBDM06]. Fiddling [Cze06]. Field [GOL^{+}07, Hay09a, PBRG07, RS02c, VO03]. Field-Testing [PBRG07]. Fields [Ano07e, Elp00]. Fifty [SPM05]. Fighting [Spe07]. Fijar [JLF09]. Fiji [OFN^{+}06, SS02a]. Fijian [OFN^{+}06]. Filling [Kel08]. Filogenética [ADF04, Fai08, PEC01]. Filogenético [Fai08, HMF^{+}01]. Filogeografía [DSE^{+}08]. Filopatría [SS02a]. Filosofías [SC09]. Filter [SDJK04]. Filters [Hun05]. Filtros [Hun05]. Fin [BGC^{+}04, CMA^{+}06]. Fina [GSR^{+}04]. Final [SSG01]. Financial [GK09, You05]. Financiamiento [HRS00]. Financieros [You05]. Fincas [KJSH08]. Finding [Kit07, Law06, Nor08]. Findings [BL07, CHBK02, RPT00]. Fine [GSR^{+}04, Hun05]. Fine-Scale [GSR^{+}04]. fines [RG04a]. Fingerprints [UWB04]. Finite [Rol06b]. Finlandeses [HKM06]. Finnish [HKM06]. Finos [Hun05]. Fire [BJM04, BM04, BAF04, Coc01, DWWF04, FEKH09, Gri00, HODB09, HIS03, HKM06, Kau04, KF01a, KF01b, Kee06, KCS04, MDJP04, Min01, NSB^{+}06, ND03, OFS^{+}04, PJYO04, QAMW03, RCMV06, SF04, SRHS09, VGB^{+}04, Vic02, Whe02]. Fire-Explicit [QAMW03].

Fire-Management [VGB^{+}04]. Fire-Suppression [BJM04]. Fires [Gil01a]. First [Joh08, Kun03, Par03, Par01, Whi06, ZSVL07]. First-of-Its-Kind [Kun03]. Fish [Bar03, BSKH^{+}06, CCS^{+}03, CA05, DDK05, DDBA^{+}07, FMT^{+}04, Fra09, GCBF09, GWJ^{+}07, GFTB04, KTN^{+}09, KBM01, KM00, KB03, LG05a, LW02a, LHPH06, LARK01, MHJ^{+}09, Mcc09, MO04, PCL^{+}03, PA08, Pop08, Rue05, Tow03, UMT^{+}01, WBSG07]. Fish-Passage [PA08]. Fisher [CDM09]. Fisheries [Bea09a, BB06, CSB07, Cor06, DLCV00, ES04, GWJ^{+}07, HAG^{+}09, JP08, KEJ06, Ker09, Kri04, LCS03, Lyn08, RDN06, WD09, ŽWGW09]. Fishers [GEJK05, GMB^{+}08]. Fishery [CDM09, DMG^{+}00, FVS03, GFTB04, KB03, MBR^{+}02, PR08, WELCH09]. Fishes [CV08, DDMMZD06, DAKB08, FKU05, Lei00, LM07, SCD^{+}08, TH03]. Fishing [AYKM04, Dav07, FMP^{+}05, HR04b, KSH00, Kai03, LARK01, Lyn06, Rut01]. Fishing-Gear [KSH00]. Fishless [DN00]. Física [AP00]. Físicos [KU02]. Fisiología [GMW08]. Fisiológicas [WBW05]. Fitness [CWNW01, EKL^{+}00, For02, GNEE07, Kel01, KLH08, LM05, NT01, RF03, Ree05, WG06, WRL02]. Fitness-Related [CWNW01]. Fitófago [ØDEA00]. Fitting [Est04]. Five [DAKB08, HMH00, HBBD08]. Fladry [MMB^{+}03]. Flagship [Bei01, BH04, Woo01]. Flames [Pyn04]. Flat [RLK^{+}01]. Flaws [MA05]. Flexible [BBC^{+}04]. Flocks [DS01, EM03]. Flood [RR00]. Flooding [SMG00]. Floodplain [GRB02, SK02b]. Flora [Mic08, PCVC06, PJW^{+}02, RVVNH09, ZGR02]. Floración [FLQ03]. Floral [SH02]. Flores [KSDBT02]. Florida [AJF03, CBH01, DOW^{+}06, HCZ00, JSJV03, JSJ05, KC01, aARLG03, Mai04, Mcc09, MH01b, MSWF00, QAMW03, SHM00]. floridana [KHP04]. Florística [PBK09]. Floristics [PBK09]. Flow [Cou02, NT01, PW01, TS06]. Flower [KSDBT02]. Flower-Visiting [KSDBT02]. Flowering [FLQ03]. Fluctuación [HMH00, LVM02]. Fluctuantes [KW02]. Fluctuating [KW02, LVM02]. Fluctuation [HMH00]. Flujo [CP01, Cou02, NT01, PW01, TS06]. Flux [CP01]. Flycatcher [Whi06]. Flying [MBC03, RMN00]. Foca [JML00, KGM01]. Focal [Arm02, BBBM02, Car00, KU02, Lam02, LMS^{+}02, WFP01]. Focal-Species [LMS^{+}02, WFP01]. Focales [BBBM02, KU02]. Focus [Kle06, Mef07a, Mef08, Nor00, PBC^{+}09]. Foe [SF04]. Folk [JHHL04]. Follow [Abe02a, Pan09, WP03]. Follow-Up [Pan09, WP03, Abe02a]. Following [BJTD03, CAFL03, FMP^{+}05, Pop08, STT05, VMBC02, ZSSM09]. Fondo [Kai03, LARK01]. Food [Ano02j, AP00, Ben02, CVS^{+}09, CWNW01, HCRM09, LAN05, STP09, VSC06]. Food-Web [CVS^{+}09, HCRM09]. Foraging [BBM04, GRB02, RS02b, WHJJ04]. Forb [HH06b, LS01]. Forbs [HH03]. Force [NMB09]. Forces [KORB02, Tho03]. Forecasted [MBMLJ09]. Forecasting [BJPL07b, LD03, MSOM^{+}06, PSG^{+}00, PV09b]. Forense [WCD^{+}08]. Forensics [WCD^{+}08]. Forest [AKO03, ASM05, And01a, Ano01j, Ano02b, Ano04d, BV04, BHB08, BVK02, Bei08, BD06, BMMR03, BCS00, BH04, Bor08, Bor05, BWS09, BCP^{+}02, BAF04, Bru05, CFC02, CWL^{+}00, CP01, CWH07, CMP^{+}00, CT01, CQLF02, CSS00, CPDO01, Cha06, Chi00a, Chr08, Col00, CH01, CCM00c, DMSP00, DW06, Dun04, ESSGL00, FH05, FB08, FTR01, FJ07, FLQ03, FJS04, GRB02, Gal01a, Gal01b, GMCP09, Gav04, GK08, GV02, Gha01, Gil01b, GC06, GPC^{+}03, Gra03, GBMH07, GG02, Gro05b, GV05, GJD05, HCBK00, HMB^{+}05, HP04, HP08, Har06b, Has00, HKN01, HDQSL06, HR04c, HVR08, HBV02, Hol08, HDK^{+}00, HAJ^{+}04, JMB01, Kau04, KWvH08, KWV08a, KD05, KIAH08, LMD05, LLF08, LW01, LLV^{+}02, LMP00, LP08, LMB00, LVBS^{+}04, MR00a, MMGB05, Mar02a, MRM^{+}05, MB01, MP07b]. Forest [MGC01, MPE^{+}07, MJ03, MML06, MM03a, MYH09, MK09, Nad04, NAC^{+}06, New06, NFB05, NPT05, NB06, NS01, Nos01a, OFS^{+}04, OFN^{+}06, Olu00, PPLS01, PJ00, PSBG05, PBK09, Per01, PV09a, PJYO04, PB09, PMEH05, Pow06, RHS05, RTN01, Rap06, RWBG06, RH09, RGF^{+}04, RMN00, RSD03, RWS02, Ric04b, RCP00, RR01, RRH^{+}09, RGGS02, RA04b, RWRW04, RS02c, Rud03, SAA^{+}01, SLELA05, SG01, SPM05, SNM00, SMN00, SLB^{+}07, STL02, SLS05, Smi09, SHYH06, STE^{+}01, SCC^{+}09, SS01, SOADMR08, SBVC03, TPS^{+}05, TFGJ06, TNRJ02, TA08, TO02, Tur06, TUE06, WCPB03, WCP01, WBV^{+}05, WZD^{+}00, CPDO01, KB08, Rig01, SF03]. Forest-Associated [MML06]. Forest-Dependent [BV04]. Forestal [Cha06, CPR05, FJS04, Gra03, GJLL08, HR04c, KB08, MML06, NB06, Nos01a, PPLS01, PV09a, Pow06, Rap06, RWBG06, TFGJ06, Tur06, WCPB03]. Forestales [CWH07, FB08, GK08, GNEE07, HDK^{+}00, IMD01, LP08, MGC01, MPE^{+}07, OFN^{+}06, SD01, TNRJ02]. Forested [BRK^{+}04, DBB^{+}09a, EB09, SD01, WGK07]. Forestry [Fra04, GNEE07, IMD01, Pyk04, SS02b]. Forests [ALR05, ARA^{+}09b, BMBM06, BMM^{+}01, Bee09, BJ05, BOJ09, BMPNC^{+}03, BEM^{+}04, Bur04, CPD^{+}09, CPME09, Coc01, CSM^{+}06, CAFL03, Cra02, DWN01, EM03, FPM02, FO06, GMCP09, GRJB^{+}07, GK06, GSR^{+}04, GPC^{+}03, GAF00, GH02, Gus02, HH00, HFR07, Hut06, HNG04, HKM06, IMD01, JMS06b, Kau04, Kes01, KFBG08, Kom03, KA03, Lam04, LO06, LGK02, MRP07, Moo07, MSBOF^{+}09, NMB09, Par02, PBP09, Per00, PL03, PRMV06, RGF^{+}04, RS01, Rud06, SAKQ^{+}05, SSSK06, SSM03, SRH^{+}07, SH09, Sie00, SBB^{+}08, SDJ06, SBVC03, Ter00, TK07, VSK01, WD01, Wes05, WSLW00]. Forever [DP01]. Forging [Far04]. Forgotten [Inf01]. Forks [BMS08]. Form [BMMW07, Ric07]. Forma [ED07, Ric07, YKIO08]. Formación [MS09, SKBY03]. Formal [MGB^{+}07]. Formas [BMMW07]. Formation [SKBY03, SS06]. Formerly [SM09]. fornicata [VDHD01]. Forraje [RS02b]. Forrajeo [BBM04, WHJJ04]. Forrajeras [GRB02]. Forward [ACD^{+}07, Edw04, Rod03]. Fosters [GALW09]. Fotografías [Mcc09, RLK^{+}01]. Foundation [BFB^{+}09, NFD^{+}09, ZKNL02]. Foundations [HCE^{+}01, ZMS^{+}08]. Four [Ano02f, BST^{+}05, BT07b, Cra02, Cre01a, Cun00, HR04a, Rol00, Sch02]. Fox [ARB^{+}07, HDA01, RSC03]. Foxes [MBC03, ST05]. Fracaso [BM05a, TSG06]. Fragmemtación [BD06]. Fragment [ED07]. Fragmentacion [Hob01, AKO03, AT05, BMMR03, BCS00, CQLF02, CTR02, Coc01, CCM00c, Cro02, Cun00, DH00b, DNZK02, DW05b, FH05, FLQ03, GC07, GC06, HRW^{+}03, HDQSL06, HVR08, HJ07, HBSD04, HAJ^{+}04, KD05, KIAH08, KA03, LMD05, LLF08, LW01, LRS^{+}09b, MHJ^{+}09, MW07, MY02, Per01, RHS05, RSF^{+}03, SAA^{+}01, STE^{+}01, TG05, TO02, VSC06, WM01, WRM05, WHH00, BK02, CVS^{+}09, GM00a, WCP01]. Fragmentada [Bos07, LM05]. Fragmentadas [Cou02, GG02, JGP09, SD03, WSBP09]. Fragmentado [ARPEBM09, CW02, DPTT^{+}05, Gra01, GGW09, LLH08, LMP00, LP07, NS01, PDF^{+}00, PMC01, RDEF01, SWD00, YL05]. Fragmentados [BOJ09, DWN01, HMB^{+}05, MRP07, MM02b, RL06, RWRW04, RBG^{+}04, VEST02]. Fragmentation [AKO03, AT05, BD06, BMMR03, BCS00, BK02, CVS^{+}09, CQLF02, CTR02, Coc01, CCM00c, Cro02, Cun00, DH00b, DNZK02, DW05b, FH05, FLQ03, GC07, GM00a, GC06, HRW^{+}03, HDQSL06, HVR08, Hob01, HJ07, HBSD04, HAJ^{+}04, KD05, KIAH08, KA03, LMD05, LLF08, LW01, LRS^{+}09b, MHJ^{+}09, MW07, MY02, Per01, RHS05, RSF^{+}03, SAA^{+}01, STE^{+}01, TG05, TO02, VSC06, WM01, WCP01, WRM05]. Fragmented [AE02, ARPEBM09, BOJ09, Bos07, CW02, Cou02, DPTT^{+}05, DWN01, Gra01, GGW09, GG02, HMB^{+}05, JGP09, LLH08, LMP00, LM05, LP07, MRP07, MM02b, NS01, PDF^{+}00, PMC01, RL06, RDEF01, RBG^{+}04, Sch02, SWD00, SD03, VEST02, YL05]. Fragmenting [Bru02]. Fragmento [AE02, ED07, MK09, SC05]. Fragmentos [BVK02, CH01, CCH03, DMS03, LLV^{+}02, MNRD^{+}08, MP07b, MJ03, MK09, NAC^{+}06, PBK09, RTN01, RSD03, SG01, SBSL06, SS01, WD06, WMP05, YKIO08, Ric04b]. Fragments [BVK02, CH01, CCH03, LLV^{+}02, MNRD^{+}08, MP07b, MK09, NAC^{+}06, RSD03, Ric04b, SG01, SBSL06, SS01, WD06, WMP05]. Fragoso [GCM^{+}08]. Frame [Sad05]. Framework [Dal05, EG07, GWS07, GBC^{+}05, KSC^{+}06, MB09a, MHS07, MU08, Rua00, SMRR02, TAP^{+}01, WABT07a, WP03]. Frameworks [BFMR07]. Framing [MF08, Orr06a]. Franca [REC^{+}00, SDRO01]. Francesa [FH05]. Franja [HNG04]. Frankel [SF00]. Frecuencia [SRHS09, YE07]. Frecuencias [KCS04, SPFF^{+}08]. Free [FPW01, Orr05b]. Free-Ranging [FPW01]. French [FH05]. Frequencies [KCS04, SPFF^{+}08]. Frequency [SRHS09, YE07]. Freshwater [Abe02a, BB00, BBH06, DDMMZD06, EHJ^{+}04, HBKF05, HGH^{+}03, JAH^{+}05, KDL07, KZA09, Kle06, LM07, PHW05, PLBBJC09, RANM08, Rue05, SMV02, ST05, SG04, VBRNP08]. Friend [SF04]. Friends [Law06]. Fringe [Mar02b]. Fritilaria [Wil02c]. Fritillary [Wil02c]. Frog [BRS^{+}05, KM00, Mcc05, PRL^{+}02, PGP04, PMEH05, SWR00, WM05, WG06, WA05]. Frogs [BD06, BRBMS06, BOC00, CFC02, DAR08, FH09, GB01, HVR08, Joh09b, SG01, WBSB09, WSKH08]. Frontera [Cur02a, MCF^{+}09]. Frontier [MCF^{+}09]. Frugivores [KFBG08]. Frugivoría [GC07]. Frugívoros [KWvH08, KFBG08, TRCC^{+}05]. Frugivorous [KWvH08]. Frugivory [GC07]. Fruit [TRCC^{+}05]. Fuego [BJM04, BM04, BAF04, DWWF04, FEKH09, HODB09, Kau04, Kee06, OFS^{+}04, PJYO04, RBRB06, RCMV06, VGB^{+}04, Vic02, HKM06]. Fuegos [DWW04, VGB^{+}04]. Fuel [Bur04, JMB01]. Fuente [CMS09, DMS03, DA07, FCL00, Mur01, NFB05, NRB00, PC08, SFS00]. Fuente-Destino [CMS09]. Fuente-Sumidero [FCL00, NRB00]. Fuente-Vertedero [Mur01]. Fuentes [HR04a, KGS^{+}07]. Fuera [NTMT^{+}03, RS02b]. Fuerte [GMMPP08]. Fuerza [NMB09]. Full [Whi03]. Full-Time [Whi03]. Función [AKM07, AHR^{+}09, Vee00]. Funciona [Kot00]. Funcionado [GK08]. Funcional [DJC^{+}07, Elp00, JSJ05, VMV07]. Funcionamiento [BSS^{+}05, BMV09, BMB^{+}08, FVS03, LBTS05, MLB08, WWR^{+}02]. Funciones [AHTM05, DFS05, LLF08, MBB^{+}08, SWD00]. Function [AKM07, AHR^{+}09, BMB^{+}08, MLB08, RD05, Vee00]. Functional [DJC^{+}07, Elp00, JSJ05, Ros02, VMV07]. Functioning [BMV09, LBTS05]. Functions [AHTM05, DFS05, LLF08, MBB^{+}08, SWD00]. Fundación [ZKNL02]. Fundaciones [ZMS^{+}08]. Fundadora [BFB^{+}09]. Fundamentals [Sta03]. Funding [HRS00, JP08]. Fungal [CDD^{+}05]. Fungi [BJ05, SLS05]. Fungus [BRS^{+}05, MSSM09, Sch10]. Further [Che02a, Mef03b]. Furthering [Hun07b]. Furtiva [GMW08]. Fusion [BC06]. Futuras [DGSS03, MRAF^{+}07, PTP^{+}08]. Future [ADD^{+}07, BC06, BRV06, CT01, DDC^{+}00, DGSS03, Gal01a, GLF^{+}08, HRS00, Hay09a, McC03, MRAF^{+}07, PTP^{+}08, Raf03, Ray06a, SAKQ^{+}05, SRW^{+}08, VCHBP08, VP07, Wil00b, WMLS09, You02]. Futuro [BRV06, GLF^{+}08, HRS00, Hay09a, SRW^{+}08, VCHBP08, WMLS09]. Futuros [DDC^{+}00]. Fuzzy [O’C00, RC00, Tod00].

Gabon [BMA^{+}08, TB01, WSA^{+}05, BMA^{+}08, TB01, WSA^{+}05]. Gacela [HMF^{+}01, KHM00]. Gacelas [IML^{+}05]. Gaharu [PPLS01]. Gaia [BR05]. Gaining [DPB^{+}09]. Galápagos [KDG^{+}06, Rut01]. Gall [PAHM04]. Game [And01a, HMF03, LW02b, MS02, PL03]. Ganado [Ber06, Bor05, BPB^{+}05, FFHW03, GM00b, HH03, Hob01, KS03, KBR03, Mad05, MM07, Mar05, MMB^{+}03, NSA06, OWOF03, PM05, RMSG^{+}03, RIBGG08, TVB^{+}06, TNTH^{+}04, VMBC02]. Ganadores [WCX09]. Gap [BV04, Cri00, DC05, GT03, Jam02, KCR^{+}08, LVBS^{+}04]. Gap-Crossing [BV04]. Gaps [HPF^{+}06]. Garden [Fre04, GG06]. Gardening [WK09]. Gardens [GFCP03, HN09, MCR^{+}04]. Gases [Jan07]. Gasto [Cle06b, HPF^{+}06, McK02]. Gastos [BMP^{+}06, RM01]. Gastrointestinal [RIBGG08]. Gastrointestinales [RIBGG08]. Gatos [LRS09a, NMT^{+}04, RSF^{+}03]. Gazelle [HMF^{+}01, KHM00, Tem02]. Gazelles [CGR01, IML^{+}05]. Gear [KSH00]. Geese [Gib07]. Gem [Har06b]. Gene [ASF^{+}09, AHR^{+}09, Cou02, NT01]. Genealogical [TRSR00]. Genealógico [JGL^{+}02, TRSR00]. Genebank [HGH^{+}00]. Genera [BSS^{+}05]. Generación [SS06, Wan04]. Generaciones [WR01]. Generada [GGGS00]. Generados [CWCG02]. General [EG07, MN03, MCM^{+}00, PRL03]. General-Purpose [PRL03]. Generales [PAHM04]. Generalidad [GBH^{+}09]. Generalist [EB09, PZG^{+}05]. Generalista [EB09]. Generalistas [PZG^{+}05]. Generality [GBH^{+}09]. Generalizada [OMP^{+}05]. Generalized [Moi08]. Generar [DJOW07]. Generate [DJOW07]. Generated [GGGS00, CWCG02]. Generation [Wan04]. Generations [Raf03, WR01]. Género [PLBBJC09]. Géneros [BSS^{+}05]. Generous [Lak01]. Genes [AHR^{+}09, Cou02, HGH^{+}00, NT01, Van04]. Genesis [Van04]. Genetic [APCS05, APSB02, AMF^{+}01, ASF^{+}09, BTM^{+}05, BFB^{+}09, BR01, BJTD03, BTGS05, BNP^{+}07, Bos07, BGWE00, CQLF02, CNC04, CTNR02, CHFR09, DSE^{+}08, DDBA^{+}07, DAKB08, EBAW01, ES08, ENN^{+}08, Fal07, FPF00, FPL^{+}04, FSS^{+}00, GMCP09, GRH05, GGA01, GK08, GM00a, HMF^{+}04, HJ07, HBSD04, KGS^{+}07, KSB01, KDD02, LTC^{+}08, LCZ^{+}05, Lin03a, LVBS^{+}04, LRS^{+}09b, LJMR^{+}01, LDG^{+}00, MRP07, MH01a, MA08, NC03a, Ols07, PS00a, PZG^{+}05, PLBBJC09, Pod01, RF03, RAW07, REC^{+}00, SC00, SKBY03, SM07a, SCP^{+}02, Tay03a, Tem02, TR01, WR01, WG06, WRL02, WHH00]. Genética [APCS05, APS07, APSB02, AMF^{+}01, BTM^{+}05, BFB^{+}09, BJTD03, BTGS05, BNP^{+}07, BGWE00, CQLF02, CMA^{+}06, CTNR02, DSE^{+}08, EBAW01, ES08, ENN^{+}08, FPF00, FPL^{+}04, FSS^{+}00, GRH05, GGA01, GM00a, HMF^{+}04, KGS^{+}07, KSB01, KIAH08, KDD02, LRS^{+}09b, LJMR^{+}01, LDG^{+}00, MRP07, MH01a, MA08, NC03a, PS00a, PZG^{+}05, PLBBJC09, Pod01, RF03, RSCL04, REC^{+}00, SKBY03, STWH00, SCP^{+}02, TR01, VBOV05, WG06, WRL02, WHH00, WRL02, GDAV00, PZG^{+}05, Wil02c, ZBABR00]. Genetically [HS06]. Genéticamente [HS06]. Genéticas [DAKB08, HJ07, LCZ^{+}05, SM07a]. Geneticist [Rom07]. Genético [BR01, Bos07, CHFR09, Lin03a, Ols07]. Genéticos [ASF^{+}09, DDBA^{+}07, Fal07, GK08, HBSD04, LVBS^{+}04, RAW07, SC00, WR01]. Genetics [APS07, BB00, BBC^{+}04, CMA^{+}06, GDAV00, KIAH08, LNP^{+}09, PZG^{+}05, RSCL04, STWH00, VBB^{+}08, VBOV05, Wil02c, WRL02, ZBABR00]. Génica [CNC04]. Genoma [AHR^{+}09]. Genomewide [AHR^{+}09]. Genotipos [MBL02, PRL03]. Genotypes [MBL02, PRL03]. Gente [CW08, MH02]. Gentiana [VBOV05]. Genus [PLBBJC09]. Geografía [BHMC01]. Geográfica [HGH^{+}00, YE07]. Geográficas [SKBY03]. Geográfico [SLCJ08]. Geográficos [JSW08]. Geographic [HGH^{+}00, JSW08, SLCJ08]. Geographical [SKBY03, YE07]. Geography [BHMC01, Put07]. Geomórfico [SSM03]. Geomorfología [JSWS00]. Geomorphic [SSM03]. Geomorphology [JSWS00]. Geopolitical [RG02]. Geopolíticas [RG02]. George [Sha01]. Gerson [FF08]. Gestión [Cha06, Cho06, CMM05a, GFCP03, Gav04, HH06b, HCM03, Hut06, JMS^{+}06a, JWB06, KEJ06, Kee06, KB06, KB08, Kri04, MDH^{+}07, MML06, Nel09, PSF05, PS04, PK05, PS06, RGF^{+}04, SSSK06, STC03, SS06, TG05, TK03a, VGB^{+}04, VBOV05, Wan04, WDA^{+}06, BRK^{+}04, TAJ04]. Gestores [MC05a]. Get [Ehr03]. Gets [LC07b]. Getting [Arm02, JGM03b]. Ghana [CMR05]. Ghats [BKWB05, KCS04, BKWB05, KCS04]. Ghazoul [KDKS08]. Ghettoization [NL03a]. Giant [DSE^{+}08, LZD^{+}03, LJMR^{+}01, SFX^{+}08]. Giants [Sei04]. Gibbon [AMF^{+}01]. Gibón [AMF^{+}01]. Gift [Pim06]. Gigante [DSE^{+}08, LJMR^{+}01]. Gigantes [LZD^{+}03, SFX^{+}08]. Gill [Dav07]. Gill-Net [Dav07]. Gillnet [DLCV00, SFT00]. Ginseng [FMDK09, HM01, Rob00]. Girasoles [MMLM01]. GIS [AOGMS09, CWCG02, WGGS07]. GIS-Based [AOGMS09]. GIS-Generated [CWCG02]. glacialis [REC^{+}00]. Glimpse [McC00]. Glimpses [Lov06]. Global [ACC^{+}05, Ano00b, Bow09, BMPNC^{+}03, BNM^{+}09, CBH01, DW06, DGSS06, Fer01, GMHS00, GK09, GMB^{+}08, GD05, HPF^{+}06, HSMK07, HJ05, JGM03b, JTZ^{+}08, KR00, LBG03, Leo04, MLN^{+}06, MCK06, MD04, ØDEA00, Pas00, PL06, PD03b, RDN06, RPT00, Ric07, Rue05, RVVNH09, SCD^{+}08, SAA^{+}09, UKC^{+}09, Ano05g, Ano06f, Mad05, Ros08]. Globales [GMB^{+}08, HSMK07, Mcc05, RDN06]. Globalización [Ehr05a, NSA06]. Globalization [Caf04, Ehr05a, NSA06]. Globally [Baw06]. Glory [WH01]. Glucocorticoid [CFH^{+}02]. Glucocorticoide [CFH^{+}02]. glyptostroboides [LCZ^{+}05]. Gnatcatcher [Rub01, ZBABR00]. Go [Ber03]. Goals [DS08, ERB01, PKB^{+}08, SEBM03, SBP07]. Goat [CD05]. Goats [Swe08]. Gobierno [Sta01, SMT07]. God [Van04]. Going [Gru04, Mef02]. Gold [MO04]. Goldenrods [Web01]. Goldenseal [Rob00]. Golet [BC09]. Golf [BSM08]. Golfo [LCS03, UCAGG08]. Golpes [VT09]. Gombe [PPW^{+}07]. Good [Bat04, Ghi06, Jes03, Kum08, LG02, Nos07, Par02, SMI^{+}07, Tut03]. Goods [Nor03]. Gordian [Cre01b]. Gorges [PCL^{+}03, PCL^{+}03]. Gorgojo [LO02]. Gorilas [RIBGG08]. Gorillas [RIBGG08]. Gorriones [RSJE06]. Got [Lim06]. Governance [BGHM06, BGHMb07, Nel09, SMW^{+}07]. Government [Sta01, SMT07]. Gradient [BMA^{+}08, MKG03]. Gradiente [BMA^{+}08, MKG03]. Gradientes [GM00b, GSB^{+}06]. Gradients [GSB^{+}06, GM00b]. Grado [GMW08, GS08]. Graduados [MS09]. Graduate [ID00, KSC^{+}06, MS09, Pér05b, VS05a, WMCB02]. Gráfico [MU08]. Gran [Arm05, Cha06, FDL^{+}05, Fje07, FNFM01, FN02a, HBV02, RCL^{+}06, SRW^{+}08, UWB04, DSTPW07, FDL^{+}05, KZA09, UWB04]. Grande [AT05]. Grandes [BGD09, DDC^{+}00, FN04, FMSL08, KSO^{+}03, PA08, MSOM^{+}06]. Granja [MKDT01]. Granjas [BJB^{+}02, MMM03, WHJJ04]. Grano [SDJK04]. Graph [MU08]. Graph-Theory [MU08]. graphical [Mei03]. Grass [ACMA04, Mai04, VMBC02]. Grass-Roots [Mai04]. Grasshoppers [KT02]. Grassland [BNH02, CLH^{+}04, CPO03, DBL06, HIS03, HH06b, JC07, LSC07, MBM04, Mor00b, MNH01, RJW09, SM09, Vee06, Vic02]. Grasslands [BE00, Bee09, BVH^{+}02, CGR^{+}05, CW08, HH03, LH04, ND04, SWB^{+}08, VMBC02, VSGS02]. Grazed [BE00]. Grazing [Bor05, Cur02b, FFHW03, FEKH09, GM00b, HIS03, HH03, Hob01, KHRS01, KS03, KT02, LSC07, Mad05, MM07, Mar05, PM05, Say03, SSM03, SK02b, Sul09]. Great [DDC^{+}00, MSOM^{+}06, SFS00, WGGS07, BGD09, FDL^{+}05, FNFM01, FN02a, KZA09, SFS00, UWB04]. Greater [BJTD03, MMK04, WRW^{+}02, WWR^{+}02, NCVBW02]. Greed [Nor03]. Green [HKM06, Put05]. Green-Tree [HKM06]. Greenback [YH01]. Greenhouse [Jan07, PL06]. Greening [Lav03, Sul09]. Gremio [TA08]. Gremios [GBMH07]. Grevillea [KDD02]. Grillo [BBK02]. Grillos [Ber05]. Gris [DN01]. Grizzly [DN01, MM02a]. groenlandicus [JML00]. Ground [Bro06, Du 01, Law06, TMR^{+}07]. Groundwater [Rai03]. Group [CWH07, GJLL08]. Group-Selection [CWH07]. Groups [ALR05, FTN^{+}05, FM09, KDD^{+}04, LBR07, LWSM03, MBB^{+}03, PP08a]. Grouse [ARMH04, PPW09, RAL^{+}04, TRNH00, WRW^{+}02, WWR^{+}02]. Groves [BKWB05]. Grow [BW08a]. Growing [KHP04]. Growth [CY08, Cze08, Dal09, HRS00, JB06, Joh03b, Joh03a, KHRS01, Lam06, Law08, Mcs05, Mef08, MSN06, MML06, Moo07, PR01a, PS00b, Ros06, Ros08, SFS05, SHYH06, SS01, SDJ06, UWB04]. Grueso [SDJK04]. Gruesos [Hun05, LGK02]. Grulla [JGL^{+}02]. Grullas [HPK^{+}04]. Grupo [GJLL08]. Grupos [ALR05, FTN^{+}05, FM09, KDD^{+}04, LBR07, LWSM03, MBB^{+}03, PP08a, RAL^{+}04]. Guadiana [FMT^{+}04, FMT^{+}04]. Guam [MBC03, PV09b, WBBA03]. Guard [DD01]. Guardia [DD01]. Guatemala [FVMG03]. Guayana [RVNM08, RVVNH09]. Guepardo [DBML07]. Guepardos [MMM03]. Guerra [DGP^{+}02, Ehr00, HBD^{+}09]. Guerrillas [Álv05]. Guest [Mae02]. Guiada [KFL08]. Guiana [FH05]. Guiar [Fje07]. Guías [CMBC00]. Guide [Bru02, Fje07, Gro05a]. Guided [KFL08, PIH00]. Guidelines [CMBC00, Hut06, KZSV05, MRAF^{+}07, MML06, PS06, RMK07, TG05, TOR^{+}04]. Guild [TA08]. Guilds [GBMH07]. Guinea [CMM05a, Hel08]. Gulf [LCS03, LS03, UCAGG08]. Gulo [FHL^{+}04]. guntheri [NKBD02]. Gunung [PPLS01, PPLS01]. Guppy [VSH^{+}07]. Guyana [FH05]. Gyps [PCD^{+}03].

H [PJ00]. H. [PJ00]. Habilidades [LLW02]. Habita [MY02]. Habitantes [GGC03]. Habitat [AT05, Ano00b, AE02, Arm05, BBBM02, BBK02, BOJ09, BMM^{+}02, BK02, BJE^{+}05, CM03, CVS^{+}09, CMP^{+}00, CJ08, Che09, CMC^{+}01, CWCG02, Coc01, CPO03, CA05, CW02, CHF00, CCM00c, CS07b, CSBS01, Cro02, Cun00, DBL06, DSJ02, DH00b, DAR08, DMS03, DPTT^{+}05, DNZK02, DJOW07, DW05b, EKL^{+}00, ESSGL00, ED07, FE07, Fer01, FLF04, FB08, FCG^{+}05, FBL^{+}09b, GC07, GN00, GSG08, GV05, HH06a, HCK^{+}02, HH00, HRS08, HVR08, Hob01, HC09, HJ07, HBSD04, HDE00, JSHM06, JMLG01, KSH00, Koh07, KA03, Lai00, LME08, LCG^{+}08, LSL^{+}06, LM07, LARK01, LMMF00, MHJ^{+}09, ML04, MNRD^{+}08, MO04, MY02, NLCAR09, NNM00, OHC08, Per01, PV09a, PD03a, PMBN01, PBL00, PB04, Rem03, RWJRR06, RSF^{+}03, RSB05, SM09, SWK05, SPM05, SC05, SSK^{+}05, SLCW08, SD03]. Habitat [SWB^{+}08, SB08, Tho00, TUE06, VSC06, VDP00, VMV07, VBOV05, WD06, WM05, WRM05, Wie06, Wil02a, WD03, WHH00, WH01, AT05, AE02, Arm05, BBBM02, BBK02, BOJ09, BMM^{+}02, BK02, BJE^{+}05, CM03, CVS^{+}09, CMP^{+}00, CJ08, CMC^{+}01, CWCG02, Coc01, CPO03, CA05, CW02, CHF00, CCM00c, CS07b, Cro02, Cun00, DBL06, DSJ02, DH00b, DAR08, DMS03, DPTT^{+}05, DNZK02, DJOW07, DW05b, EKL^{+}00, ED07, FE07, Fer01, FLF04, FB08, FCG^{+}05, FBL^{+}09b, GN00, GSG08, GV05, HCK^{+}02, HH00, HCE^{+}01, HRS08, HVR08, HC09, HJ07, HBSD04, HDE00, JSHM06, JMLG01, KSH00, KA03, Lai00, LCG^{+}08, LSL^{+}06, LM07, LARK01, LMMF00, MHJ^{+}09, ML04, MNRD^{+}08, MO04, MY02, NLCAR09, NNM00, OHC08, Per01, PV09a, PD03a, PMBN01, PBL00, PB04, Rem03, RSF^{+}03, RSB05]. Hábitat [SM09, SWK05, SPM05, SC05, SLCW08, SD03, SWB^{+}08, SB08, TUE06, VSC06, VDP00, VMV07, VBOV05, WM05, WRM05, Wil02a, WD03, WHH00, WH01, Che09]. Habitat-Quality [PV09a]. Habitat-Specific [LME08]. Habitats [Bat04, DGP^{+}02, Elp00, FBK02, HP04, HN09, JTS06, JSHM06, KC01, LFP05, MM07, MBB^{+}08, Mur01, PDF^{+}00, RDEF01, RTCR03, RS02c, SKBY03, SB03, VEST02, ZGR02, Bat04, CSBS01, DGP^{+}02, Elp00, ESSGL00, FBK02, HH06a, HP04, HN09, JTS06, JSHM06, LME08, LFP05, MM07, Mur01, PDF^{+}00, RDEF01, RTCR03, RS02c, SKBY03, SB03, WD06, ZGR02, KC01, MBB^{+}08]. Habits [Bel02]. Habituación [WBW06]. Habituation [WBW06]. Hacer [FO06]. Hacia [AW01, CP01, HL07b, JSJV03, RH01, dLBBG^{+}08, DLCV00, EK08, NTGT03, Hay09a, Hut06]. Haciendo [KCR^{+}08]. Haddad [NB00]. Hairless [TRSR00]. Halcones [TR01]. Half [Bel02]. Hallazgos [CHBK02]. Halsing [WSPL09]. Hampshire [LB02]. Handmaidens [BR05]. Happens [Kar09]. Hard [Orr04b, Tut03]. Hardwood [GJD05, HFR07, WCPB03]. Harp [JML00]. Hartebeest [FSS^{+}00]. Harvest [BBB^{+}07, Car07, EGPS04, FMDK09, FMLL06, KTN^{+}09, RG04a, TNRJ02, TSP^{+}08, ZSVL07]. Harvested [BMC^{+}09, RW02, RR01]. Harvesting [Bea09a, BV01a, CWH07, DA07, FO06, JAH^{+}05, Ker09, Sie04, TA08, VMM07]. Hatchery [AAO^{+}07, MBQ05]. Hawai [AYKM04, FR07, KGM01, TH03]. Hawaiano [CWL^{+}00]. Hawai’i [DDAK04, AYKM04, DDAK04, TH03]. Hawaiian [CWL^{+}00, FR07, KGM01]. Hawthorn [GC07]. Hay [LP05]. Head [Lei05]. Headwater [LB02, NR03]. Health [DKW01, MK02b, PK01, SW07]. Healthy [OS07]. Hearts [JMM07]. Heat [MBH07]. Heath [CWNW01]. Heaven [Van05]. Heavy [CSS00, Nos09]. Heces [BST^{+}05]. Hectare [NBM^{+}04]. Hectárea [NBM^{+}04]. Hector [SFT00]. Helecho [Gun02]. Heliconia [BK02]. Hell [Put05]. Help [BNP^{+}07, LV04, Sch06]. Helping [Had07, TSH06]. Hembras [GMW08]. hemionus [SRW06]. Hemisphere [Cam08]. Hemorrágica [Cal06]. Hemorrhagic [Cal06]. Henry [Caf01]. Hepáticas [MJ03]. Herb [BK02, LM05, LE02, TO02]. Herbáceo [MB01]. Herbivore [EB09, JC07]. Herbivores [Fon09, FMLL06, RL06, SCM^{+}07]. Herbivoría [FMDK09, HBA09, RTN01, FEKH09]. Herbívoro [EB09, JC07]. Herbívoros [Fon09, FMLL06, RL06, SCM^{+}07]. Herbivory [FMDK09, FEKH09, HBA09, RTN01]. Herbs [MB01]. Herd [TRSR00]. Here [Bek04]. Herencia [Kel01]. Heritability [Kel01]. Heritage [CRGG01, LSN07]. hero [CWNW01]. Herpetofauna [GRJB^{+}07].

Herpetofaunal [SOADMR08]. Herpetofaunas [Bur04]. Herpetofaunísticos [SOADMR08]. Herping [Lim06]. Herramienta [COW03, DDMMZD06, HLB^{+}03, LOM03, PCC03, RA04a, SBB^{+}08, Tho05b, Wal09]. Herramientas [BMRB04, BFMR07, CSB07, Ehr06, MC05a]. Heterogeneidad [CPD^{+}03, DJOW07, SHM04, WD06, HIS03]. Heterogeneity [CPD^{+}03, DJOW07, HIS03, SHM04]. Heterosis [VSO04]. Heuristic [BBA^{+}02]. Hibernación [BCP^{+}02, LCS04]. Híbrida [HMF^{+}04]. Hibridación [FH06, PFL01, WTR01]. Híbridas [ALH^{+}04]. Hibridización [RLC^{+}00]. Híbrido [KSB01, RNL^{+}04]. Híbridos [HMF^{+}04, RNL^{+}04]. Hidrastia [Rob00]. Hidrología [JSWS00]. Hidrológica [BRV06, SH02]. Hidrológicas [DDBA^{+}07]. Hidroperíodo [SKBB00]. Hierarchical [CJ08, ZR00]. Hierba [BK02, HH06b, LM05, LS01, LE02, PRL03, TO02, YL05]. Hierbas [HH03, VS09]. High [GSB04, Kel01, PJYO04, WMP05]. Higher [MDCK07, VIMMO05]. Higher-Taxon [MDCK07]. Highlands [LO06, RVNM08, RVVNH09]. Highly [ARPEBM09, FPL^{+}04]. Highway [FD00]. Hihi [AP00]. Hiking [SS03b]. Hill [Lam04]. Himalaya [Pan09, SH09, XGS^{+}09]. Himalayas [Kal05, XGS^{+}09]. Hindú [SH09]. Hindús [Kal05]. Hipótesis [BFH^{+}06, CT07, DSJ02, FBC^{+}06, LME08, ME00, RHA^{+}05, RA05]. Historia [Bra03a, BV01b, CVS^{+}09, Fea05, FDBM04, KD05, MCR^{+}04, OGH04, OHC08, SFD09, Wed02, Wil02c, FHL^{+}04, KPJ06, LM05, Slo02]. Historias [Pyn04]. Historic [DW05a, FFHW03, KF01a, LCC^{+}07, LW02b, PMBN01]. Histórica [FC05, LCC^{+}07, OMP^{+}05, WSLW00]. Historical [Bor05, FC05, JG04, Jon04, LE02, MBM04, Mcc09, MK02a, OMP^{+}05, PJK03, WSLW00]. Históricas [DW05a, MBM04, Mcc09]. Histórico [Bor05, FFHW03, KF01a, LSN07, LW02b, MK02a, PJK03]. Históricos [LE02, PMBN01]. Histories [Wil09b]. History [Bra03a, Bru05, BV01b, CVS^{+}09, DM04a, Fea05, FHL^{+}04, FDBM04, KF01b, KD05, KPJ06, LM05, LSF04, MCR^{+}04, MDRB05, OGH04, OHC08, SFD09, She02, Slo02, Van04, Wed02, Wil02c]. HIV [Awo08]. HIV/AIDS [Awo08]. Hoddle [Pro04]. Hogar [ME00, SS02a]. Hojarasca [ARP05, HVR08]. Hojas [BMBM06, EGPS04, FMLL06, NBM^{+}04]. Holanda [KBS^{+}04]. Holarctic [Ber07]. Holárticas [Ber07]. Holocarpha [HH06b]. Holocene [DDC^{+}00]. Holoceno [DDC^{+}00]. Holothurians [UWB04]. Holotúreos [UWB04]. Homage [Lee03]. Home [GG06, Hun07a, HN09, Lei00, ME00, WK09]. Home-Site [ME00]. Homing [SS02a]. Homogecene [Pur03]. Homogeneización [DJC^{+}07]. Homogenización [RGW09, RWRW04]. Homogenization [DJC^{+}07, ODD05, RGW09, RWRW04]. Homogocene [Ric05]. Honest [PA06b]. Honeybees [EBAW01]. Hong [CMA^{+}06]. Hongo [BRS^{+}05, MSSM09]. Hongos [BJ05, CDD^{+}05, SLS05]. Hood [AAO^{+}07]. Hope [Ano07m, Ehr09, Fle04a, Kni07, MCG^{+}09, Nor08, Orr04b, Orr04d, Orr07c, Sod06, Tuc04, Web05]. Hopeful [MCG^{+}09]. Horizon [CW08, Gun02, Tuc04]. Horizonte [Gun02]. Hormiga [HHSC00]. Hormigas [ARP05, DBL06, DOGN08, PV02, PAA^{+}08, RCP00, SMM09, TS03]. Hormone [Ano01h]. Hornbill [SH09, Smi09]. Hornbill-Dispersed [SH09]. Horned [FSC02]. Horses [LEPH02]. Horticultural [DSTPW07]. Hospedero [ØDEA00]. Host [MKK^{+}09, ØDEA00]. Hot [BL03]. Hotspot [Dav07, DMDFV08, HKC^{+}08, PDLS08]. Hotspots [AW01, Ara02, BMM^{+}02, CGR^{+}05, Fat06, Fon09, GSR^{+}04, GBC^{+}04, HBD^{+}09, JB06, Kom03, Kop01, MLN^{+}06, MN03, MSBOF^{+}09, MM03b, Ova03, RLXL01, Vee00]. Hotter [Orr07c]. House [RSJE06]. Household [PCL08]. Housing [LKG06]. How-Much-Is-Enough [Wil08a]. Huérfanos [TAP^{+}01]. Huertos [SD03]. Huésped [MKK^{+}09]. Huevos [PLB09, TSP^{+}08]. Human [Ada06, Ano01k, ANW^{+}08, BS08, Bor08, BISS06, CPR^{+}07, CY08, CNM01, DeS03, DDBA^{+}07, Fie00b, Gil08, Har05, HJ05, JB06, KGM^{+}02, KFNB09, KG02, LLS08, LFR^{+}08, LAN05, LF00, MNRD^{+}08, Mas03, MC09, Mcd04, NWRD03, ODD05, PH02, PM07, PP03, Pyk00, RAL^{+}04, Rol06a, Rud06, SRSP01, SBM06, SSK^{+}05, SBK^{+}06, SRHS09, TB01, TK03a, Tre04, TNTH^{+}04, TWW09, VVMV09, Van04, WBW06, WMD^{+}04, WMD^{+}06, WGK07]. Human-Carnivore [TK03a, TNTH^{+}04]. Human-Caused [SRHS09]. Human-Dominated [LLS08, WMD^{+}04]. Human-Modified [PP03]. Humana [BS08, CY08, CNM01, HJ05, KGM^{+}02, LAN05, LF00, MNRD^{+}08, PH02, PM07, WGK07]. Humanas [Bor08]. Humano [CPR^{+}07, JB06, KFNB09, MC09, TNTH^{+}04, VVMV09]. Humanos [ANW^{+}08, DDBA^{+}07, LLS08, LFR^{+}08, NWRD03, PP03, Pyk00, RAL^{+}04, Rud06, RIBGG08, SRHS09, TB01, TK03a, TWW09, WBW06, WMD^{+}04]. Humanos-Vida [RAL^{+}04]. Humans [Ber01, RS00, RIBGG08]. Humedal [GWS^{+}06, SH02]. Humedales [Elp00, FB00, GWS^{+}06, Gib00, HTM^{+}05, HF04, JGM03a, JMH01, Mar05, MJ03, PM05, RH08, SB03, SMC00, SKBB00]. Húmedas [MEA00]. Húmedo [RCP00]. Húmedos [ARA^{+}09b]. Humid [ARA^{+}09b]. Humo [DWWF04]. Hundred [Orr07b, SPM05, SAA^{+}09]. Hungary [SF05]. Hungría [SF05]. Hunt [JML00]. Hunted [SFD09, SH09]. Hunter [Paq07, BST^{+}05, CCM00a]. Hunter-Collected [BST^{+}05]. Hunters [PL03]. Hunting [BBB^{+}07, CWC00, ESSGL00, FYC00, FND06, FJ07, HMF03, LCT^{+}06, LCG^{+}08, LFA^{+}07, LMG06, MNA07, NTMT^{+}03, NRB00, OSSK^{+}07, PBP09, Per00, Per01, RS01, SK02a, SHM04]. Hunting-Related [HMF03]. Husbandry [OWOF03, Pyk00]. Hybognathus [AT05]. Hybrid [HMF^{+}04, KSB01, RNL^{+}04]. Hybridization [CK05, FH06, PFL01, RLC^{+}00, WTR01]. Hybridized [ALH^{+}04]. Hybrids [HMF^{+}04, RNL^{+}04]. Hydrology [JSWS00]. Hydroperiod [SKBB00]. Hylocichla [DD04b, SFS00]. Hypericum [QAMW03]. Hypotheses [BFH^{+}06, DSJ02, RHA^{+}05]. Hypothesis [CT07, FBC^{+}06, ME00, Pea01, RA05, Roo00, Sta09, LME08].

Iberian [FMT^{+}04, ALM07, BRBMS06, LMP02, PDF^{+}00, TRSR00]. Ibérica [BRBMS06, FMT^{+}04]. Ibéricas [ALM07]. Ibérico [PDF^{+}00, TRSR00]. Ibéricos [LMP02]. Iberoamerican [Mor00a]. Ibex [SS00]. Íbices [SS00]. ICO [Ano04d]. Iconography [Cor04b]. Icons [DeN07]. Idea [Bru02, JMB01, Kum08]. Ideal [CLŽ^{+}07, Chi09a]. Ideasclerosis [Orr00c, Orr00b]. Identidad [HMF^{+}04]. Identificación [CWCG02, CA05, FP02, LBR07, LSL^{+}06, MBL02, SBB^{+}08, Vee00, WHA^{+}05, CMA^{+}06, HMF^{+}04, HCZ00, MB00, MCG^{+}09, MPR^{+}01, PMC01, SCP^{+}02, WM05]. Identificar [CHW^{+}01, Dun02, MBB^{+}03, PBSR06, ZG05]. Identification [CMA^{+}06, Des06, FP02, HMF^{+}04, LBR07, LSL^{+}06, MB00, SCP^{+}02, SBB^{+}08, Vee00]. Identify [CHW^{+}01, Dun02, PBSR06, PS01a]. Identifying [CWCG02, CA05, HCZ00, MBL02, MCG^{+}09, MPR^{+}01, MBB^{+}03, Mul07, PMC01, WM05, WHA^{+}05, ZG05]. Identity [HMF^{+}04]. Ideology [Mcd04]. If [Cze06]. Ignoradas [ESB05]. II [GB05b, Lov06]. III [RKR03b]. Ilegal [Mad05, MBL02, TSP^{+}08, WCD^{+}08]. Ilegales [PBSR06]. Ilícitos [Álv02]. Illegal [PBSR06, TSP^{+}08, WCD^{+}08]. Illicit [Álv02]. Illinois [Sch03b]. Illustrated [Cre01a]. Illustration [Tem02]. Ilustrado [Cre01a]. Imágenes [SAA^{+}01]. Imagery [SAA^{+}01]. Imaginary [Got06]. Immediacy [Spe07]. Immediate [JLRBB^{+}07]. Immigrant [TTB03]. Immigration [War05]. Impact [ASM05, BS09, BMMR03, BTS09, Cal06, FBC^{+}06, FCG^{+}05, GNEE07, LBG03, LCS03, LSC07, Mal03, MDF05, MSSM09, MM03b, NTMT^{+}03, PGP04, PBS03, SRW06, SAPRB07, Sch10, Spe04, ST05]. Impacto [ASM05, BTS09, Cal06, FCG^{+}05, LCS03, MDF05, NTGT03, NTMT^{+}03, PGP04, PBS03, SRW06, SAPRB07, ST05, BS09, BMMR03, CPSS03, FBC^{+}06, GNEE07, LBG03, LSC07, MSSM09]. Impactos [Bey04, GWJ^{+}07, HY03, LFR^{+}08, LO06, Lus03, MP08, PM05, BJM04, BHB08, BC08, DDBA^{+}07, FMLL06, GMW08, HH03, Kee06, LCT^{+}06, LCG^{+}08, MM07, SMT07, WBBA03]. Impacts [BJM04, BHB08, Bey04, BC08, BRR^{+}02, CPSS03, DDBA^{+}07, Don04, FMLL06, GMW08, GWJ^{+}07, HH03, HY03, Kee06, LCT^{+}06, LCG^{+}08, LFR^{+}08, LO06, Lus03, MM07, MP08, PM05, SMT07, WBBA03]. Impenetrable [HCBK00, OBM09, RIBGG08]. Imperative [Ang00, Pov01]. Imperativo [Ang00]. Imperfections [Ehr05b]. Imperiled [Log03, PPRB06, UKC^{+}09, Wen08]. Imperilment [BRTW07, BL06]. Implementación [BRYK01, EK08, FDL^{+}05, JMS^{+}06a, KDC^{+}06, KCC06, SAPRB07, TFGJ06, KCR^{+}08, KB08]. Implementadas [CLŽ^{+}07]. Implementation [BRYK01, EK08, FDL^{+}05, JMS^{+}06a, KDC^{+}06, KCR^{+}08, LGB05, SAPRB07, TFGJ06]. Implemented [CLŽ^{+}07]. Implementing [KCC06, KB08]. Implicaciones [AMF^{+}01, BZD03, DDC^{+}00, EGPS04, FKU05, GRB02, Gav04, GWS07, GCV03, GWS^{+}06, GS05, GBH^{+}05, HMF^{+}04, HMF^{+}01, HLG03, JSW08, KFSR^{+}06, LCZ^{+}05, LARK01, LP01, MSB^{+}00, ND03, ØDEA00, PR03, PCL08, PPW09, RTN01, SMM09, SMG00, Sie04, SKBB00, SLCW08, SS06, TG04, TTB03, UMT^{+}01, VSH^{+}07, Wed02, WS03, WD03, CNC04, CBH01, MBC03, PFL01]. Implicancias [KSO^{+}03].

Implications [AMF^{+}01, ABE01, BZD03, CNC04, CBH01, DDC^{+}00, EGPS04, FKU05, GRB02, Gav04, GWS07, GCV03, GWS^{+}06, GS05, GBH^{+}05, HMF^{+}04, HMF^{+}01, HLG03, JSW08, KFSR^{+}06, KSO^{+}03, LCZ^{+}05, LARK01, LP01, Mas03, MCM^{+}00, MSB^{+}00, MBC03, ND03, ØDEA00, PFL01, PR03, PCL08, PPW09, Pur03, RTN01, SMM09, SRL^{+}07, SMG00, STWH00, Sie04, SKBB00, SLCW08, SS06, TG04, TTB03, UMT^{+}01, VSH^{+}07, Wed02, WS03, WD03]. Imponer [WG08]. Importance [APCS05, CJ08, Cor03, DM04b, FBK02, GEJK05, GRG^{+}05, Hei01, JAH08, Kai03, Kér04, LNP^{+}09, LS07, LGK02, MKH07, Nad04, Paq07, PKB^{+}08, PP03, PR08, SD03, SB08, SD01, SAA^{+}09, Tem02, TSKD08, VP06, YKIO08, YL05]. Importancia [CJ08, CGR^{+}05, DMDFV08, Fat06, Fon09, HBD^{+}09, JB06, JAH08, Kai03, KSP00, Kér04, Kom03, MLN^{+}06, MSBOF^{+}09, PKB^{+}08, PP03, SD03, SAA^{+}09, TKB^{+}06, APCS05, DM04b, FBK02, GEJK05, LGK02, MKH07, SB08, SD01, YKIO08, YL05]. Important [Ano03d, Bec07, BGD09, Don06, Fle08, Joh09a, LBR07, Man02, TKB^{+}06, WM05]. Importante [Dav07]. Importantes [AW01, BV01b, LBR07, WM05]. Importation [Har07]. Impossible [Lov04]. Impoundment [HAG^{+}09]. Impoverishment [RWRW04, SLKP06]. Impresiones [UWB04]. Improve [Bre03, GBE^{+}06, KORB02, KMGB07, Pér05b, RWM^{+}06]. Improved [GBC^{+}05]. Improving [CHBK02, CHF00, HP01, KRM^{+}00, KRW^{+}00, LBR07, Mur00, NFD^{+}09, PP08a, RANM08, SMRR02, SL00b, TG05, VO03]. Impuestos [Rom03]. Impulsa [SB04]. Inadecuado [HRS08]. Inadequate [HRS08, KDKS08]. Inbred [JRL03]. Inbreeding [AHR^{+}09, CGR01, DH00a, EC01, ENN^{+}08, FG00, Fra01, JRL03, JWB06, KHM00, KLH08, LRS^{+}09b, MRP07, VSH^{+}07, VSAG06]. Incendio [BRK^{+}04, Bey04]. Incendios [Coc01, HIS03, KF01a, KCS04, MGPM04, NSB^{+}06, Pan02, QAMW03, SRHS09, Bur04, ND03]. Incentive [MAM^{+}03]. Incentives [BMV09, CS05, Hop07, LLS00, WL04]. Incentivos [BMV09, MAM^{+}03, WL04, CS05, LLS00]. Incertidumbre [AFB^{+}00, BRV06, Cor02, For00b, MWBP07, MC05b, MWEB06, Pra05, SFT00, SLCW08, TWDB00, WMVK03]. Incidental [LC07a, RDN06, UCAGG08, WELCH09, ŽWGW09]. Incidentales [CLŽ^{+}07]. Incierta [HSR05]. Incierto [PCC03]. Inciertos [HRS00]. Incipient [HHSC00]. Incipiente [RHA^{+}05]. Incipientes [HHSC00]. Included [ALH^{+}04]. Including [Bru02, SHM04]. Incluídas [Gus02]. Inclusion [PJK03, ALH^{+}04, PJK03, SHM04]. Income [WG01]. Incompatible [GPMM08, LDG^{+}00, MRP07]. Incomplete [NL05]. Incompletos [NL05]. Inconclusiveness [Mcc05]. Incongruencias [HPF^{+}06]. Inconsistency [MM00]. Incorporación [BNH02, Cor02, JGL^{+}02, LMMF00, Lyn06, MWBP07, RM06, TWDB00]. Incorporando [LSOBA04]. Incorporating [BNH02, Cor02, JGL^{+}02, LSOBA04, LMMF00, MWBP07, RM06, SRSP01, TWDB00]. Incorporation [Lyn06]. Increased [GGI^{+}05, MS04, RTN01, SRW06]. Increases [FJ07, HAJ^{+}04, SRHS09]. Increasing [AM05, GLMN00, Gra09, KCS04, LD03, Vee06]. Incrementa [FJ07, HAJ^{+}04]. Incrementada [MS04]. Incremental [SBP07]. Incrementan [Ric04b]. Incremento [FCG^{+}05, KCS04, LD03, SRW06, AM05, FV05, GGI^{+}05, RTN01, ZMS^{+}08]. Incrementos [SRHS09]. Incumplimiento [Kri04]. Independent [PR08]. Independently [Pea01]. Index [Ano03b, Ano03g, Ano04a, Ano04l, Ano05a, Ano05l, Ano06a, Ano06k, Ano07a, Ano08b, Ano08r, Ano09a, Ano09o, BNP^{+}07, BAKHT06, LD07, NS09, QRF07, Roy04, WPS08, CLW^{+}09]. India [Ram01, BKV^{+}05, BKWB05, KCS04, KK09, Mad05, Ram01]. Indian [TRCC^{+}05, Kal05, Mad05, SH09]. Indians [Gal01b]. Indicaciones [EG07]. Indicada [PB04]. Indicador [FM09, HMK^{+}09, MCK06, QR05]. Indicadora [HMK^{+}09]. Indicadoras [BMBM06, FTN^{+}05, NF04]. Indicadores [BSS^{+}05, BMB^{+}08, CMP^{+}00, CKL^{+}00, KDD^{+}04, KSP00, KU02, LME08, LBR07, LWSM03, MB07, MB01, MEA00, MGC01, MBB^{+}03, PK01, VIMMO05, LMB00]. Indicated [PB04]. Indications [EG07]. Indicator [BMBM06, FTN^{+}05, FM09, HMK^{+}09, KSP00, LBR07, LWSM03, MCK06, MBB^{+}03, NF04, QR05, RGGS02]. Indicators [BSS^{+}05, BMB^{+}08, BAKHT06, CMP^{+}00, CKL^{+}00, KDD^{+}04, LME08, LMB00, MB07, MB01, MEA00, MGC01, QR06]. Índice [CLW^{+}09, BNP^{+}07, NS09, QRF07, WPS08]. Indices [MNH01, MNH01, McK02, Roy04]. Índico [TRCC^{+}05]. Indígena [Mcs05]. Indígenas [NSB^{+}06, SP09, SZ05, SMT07]. Indigenous [Bro04, Cho06, Col00, EBDM06, GGC03, Kal05, Mcs05, Mcs06, NSB^{+}06, SP09, SZ05, SMT07]. Indigo [Wel06]. Indirect [LH04, LEPH02, RBP01]. Indirectos [LH04, LEPH02, RBP01, VIMMO05]. Individual [BMMW07, FTC04a, KF03, MMF^{+}05, Nor03, PLB09, Tow03]. Individual- [MMF^{+}05]. Individuales [BMMW07, CKL^{+}00]. Individuals [KF02, LTC^{+}08, MBL02, RSC03]. Individuos [FTC04a, KF02, LTC^{+}08, MBL02, PLB09, Tow03]. Indo [BMP07]. Indo-0Pacific [BMP07]. Indonesia [FJS04, GFCP03, JLF09, JW02, Koh09, PPLS01, VSK01, WGW01, WMM04]. Industria [Ber06, TB03]. Industrial [Ewi09]. Industrias [NSA06]. Industries [NSA06]. Industry [Ber06, NARL^{+}04, TB03]. Inequality [HPG09]. Inequidad [HPG09]. Inesperados [LO02]. Inestabilidad [HCBK00]. Infección [GC06, KPH07]. Infecciones [Urb06b]. Infecciosas [GGI^{+}05, PJNA07, RHA^{+}05, SSL06]. Infection [GC06, KPH07, OMP^{+}05]. Infections [Urb06b]. Infectious [GGI^{+}05, PJNA07, RHA^{+}05, SSL06]. Inference [NS09]. Inferences [PD03a]. Inferencia [MRDC06, NS09, DKS^{+}09]. Inferencias [PD03a]. Inferior [TTB03]. Inferring [DKS^{+}09, MRDC06, TPW03]. Infiriendo [TPW03]. Influence [Bei08, BD06, CMM^{+}05b, CY08, FPSD09, Gun02, HMB^{+}05, JC07, JFS08, LHPH06, MR00a, NFM04, PA04, SP00, WHH00]. Influences [KL06, Wal03]. Influencia [CMM^{+}05b, CY08, LHPH06, SP00, Bei08, BD06, FPSD09, Gun02, HMB^{+}05, JC07, JFS08, MR00a, NFM04, PA04, WHH00]. Influencían [CW00]. Influencias [Wal03, KL06]. Influencing [CW00, Gra01, MMM03]. Influye [PM05]. Influyen [Gra01, MMM03]. Información [FNS^{+}03, GS01, HC09, LRO^{+}08, MP08, PBC^{+}09, RPT00, Sim03, WBSG07]. Informal [DC06b]. Informático [WM05]. Information [Bre03, FNS^{+}03, GS01, HC09, LRO^{+}08, MP08, Sim03, WM05, WBSG07]. Information-Theoretic [WM05]. Informed [PBC^{+}09]. Infraestructura [VCHBP08, RC05]. Infrastructure [RC05, VCHBP08]. Ingeniería [APSB02]. Ingeniero [BRA08]. Inglaterra [BCHS08, FO06, HBA09, SB04]. Inglés [VDHD01]. Inhabiting [Olu00, SS01]. Inherent [MH00]. Inherente [MH00]. Inhibición [NSB^{+}06]. Inhibition [NSB^{+}06]. Inicial [MM04]. Iniciales [KHHK03]. Iniciativas [HTM^{+}05, TRCC^{+}05]. Initial [KHHK03, MM04]. Initiatives [HTM^{+}05, RA04b, TRCC^{+}05]. Inland [ATG05, Lei00, LCZX09]. Inmigración [War05]. Inmigrantes [TTB03]. Input [KHP04, SDL07]. Insect [Dun05, Fon09, HCRM09, HDE00, Kom03, RL06, Sam06, Tal04, VSC06, Vic02, ZLM03]. Insect-Plant [VSC06]. Insecticida [BS01b]. Insecticide [BS01b]. Insectívoras [GRB02]. Insectivoría [JSJ05]. Insectívoros [JP09]. Insectivorous [GRB02, JP09]. Insectivory [JSJ05]. Insecto [ØDEA00, RL06, VSC06]. Insecto-Planta [VSC06]. Insectos [Dun05, Fon09, HCRM09, HDE00, Kom03, LS09, LARR03, NBM^{+}04, Pan02, SH03, Vic02, WHJJ04, ZLM03]. Insects [LS09, LARR03, NBM^{+}04, ØDEA00, Pan02, SH03, WHJJ04]. Insights [BFMM02, MF08, Mcs05, SC09, SS08a]. Inspiration [Dor06, Fie06]. Inspire [Pim06]. Instability [HCBK00]. Instalaciones [VTRG00, PA08]. Institucionales [PMC^{+}07]. Instituting [PR08]. Institutional [PMC^{+}07, Pro04]. Institutions [GOL^{+}07, VL01]. Instream [MO04]. Instrument [GLMN00]. Instrumentación [Rig01]. Instrumento [GLMN00]. Insuficientes [GZK06]. Insular [APCS05, RSJE06]. Insulares [APCS05, CSBS01, DOGN08]. insularis [DW05a, GPMM08]. Insult [SSG01]. Insumo [WG01]. Insumos [KHP04]. Integra [GSG08]. Integración [BPV09, LG05b, AAO^{+}09, Arm05, Cur02a, HLSZ02]. Integradora [KB06]. Integrados [LRO^{+}08, SGSA07]. Integrando [DWW04, ARMH04, HCM03]. Integrar [BTK08]. Integrate [BTK08]. Integrated [LRO^{+}08, Min01, SAKQ^{+}05, SGSA07]. Integrating [AAO^{+}09, ARMH04, Arm05, BGRJ00, BPV09, Cur07, DWW04, FRH04, FCN^{+}06, GSG08, HKC^{+}08, HLSZ02, HCM03, HMH03, LG05b, SSK^{+}05]. Integration [Cur02a, RANM08]. Integrative [BH01, KB06]. Integridad [IF02, RAW07, WD01]. Integrity [IF02, RAW07, WD01]. Intención [SAPRB07]. Intenciones [AJF03, JSJV03]. Intended [FVS03]. Intense [VVMV09]. Intensidad [KSDBT02, KPH07, KT02, LSC07]. Intensidades [Kes01]. Intensificación [BMV09, JP09, WHJJ04]. Intensification [JP09, MV06, WHJJ04, BMV09]. Intensified [Koh07]. Intensities [Kes01]. Intensity [KSDBT02, KPH07, KT02, LSC07]. Intensive [VP06]. Intenso [VVMV09]. Intent [PD00]. Intention [EBDM06, SAPRB07]. Intentions [AJF03, JSJV03, Par02]. Interacción [ZSMW06, BHMC01, Car07, Say05, WG06]. Interacciones [BBB08, MKK^{+}09, PP08a, RSCL04, Tho05b, BS01b, Coc01, RS01]. Interacting [CVS^{+}09, Car07, Say05]. Interaction [WG06]. Interactions [BBB08, BS01b, Coc01, MKK^{+}09, PP08a, RSCL04, RS01, Tho05b]. Interactivas [SEBM03]. Interactive [FMDK09, SEBM03]. Interactivos [FMDK09]. Interactuantes [CVS^{+}09]. Interagenciales [HSDS01]. Interagency [HSDS01]. Intercambios [CW08]. Intercharcas [Pyk05]. Intercontinental [FPM02]. Intercrosses [ALH^{+}04]. Interdisciplinarias [CMBC00, SMLB00]. Interdisciplinary [BPP05, Cam05, CMBC00, CCM00b, Kaw04, Mul07, NL03a, SMLB00]. Interés [BPL^{+}01, KB04a, RBT08, RAL^{+}04]. Interesados [BF03]. Intereses [CAR05, DWW04]. Interespecífica [BRS^{+}05]. Interespecíficos [ES08, LW02a]. Interest [AFK^{+}09, Cla01, PJB^{+}06]. Interests [CAR05]. Interface [BG04]. Interior [CP01, HWH^{+}02, LCZX09, MO04, WRW^{+}02, WWR^{+}02]. Interiores [ATG05]. Interiors [CP01]. Intermareales [LAN05]. Intermedias [Kes01]. Intermediate [Kes01]. Internacional [BM05b, FV05, LD03, ZJ04]. Internacionales [Cle06b, Mad05]. International [Aco00, Ano08l, BGRJ00, BM05b, Cle06b, Dav02, FV05, FF08, GCM^{+}08, LD03, Mad05, Mae02, Mef02, RA04b, VS05a, Wes03, ZJ04]. Internationalization [Mef03b]. Interns [Whi03]. Interplay [BHMC01, ZSMW06]. Interpond [BBH06, Pyk05]. Interpretation [LNP^{+}09]. Interpretativo [DD01]. Interpretive [DD01]. Interrupción [TR05]. Intersección [LG02]. Intersecciones [Spe02]. Intersection [LG02]. Interspecific [BRS^{+}05, ES08, LW02a]. Intertidal [LAN05]. Intervención [BV01b, KFL08, PCVC06]. Intervenciones [Fer01, TWW09]. Intervention [Awo08, BV01b, KFL08, PCVC06]. Interventions [Fer01, TWW09]. Intimate [Alt05]. Intracruza [KHM00]. Intraespecífica [BNP^{+}07, SL08]. Intraspecific [BNP^{+}07, SL08, You01]. Introducción [BEB00, FDBM04, HH06b, KDG^{+}06, MTB00, VDHD01, VSO04]. Introducciones [CTNR02, DDK05, KS05, Rue05, WSCP05, KM00]. Introduced [APSB02, Ber05, BRBMS06, Côt05, EBAW01, FDBM04, FTC^{+}04b, FR07, KGS^{+}07, PFL01, Pop08, Sim03, TS09, TPW03, TPT^{+}07, VDHD01]. Introducidas [APSB02, EBAW01, FTC^{+}04b, FR07, KGS^{+}07, Sim03, TS09, TPT^{+}07]. introducido [BIK^{+}06]. Introducidos [Ber05, BRBMS06, FDBM04, PFL01, Pop08, TPW03]. Introducing [HH06b, WSCP05]. Introduction [AL03, Ano08k, BEB00, CCM00b, FDBM04, Hou00, KDG^{+}06, LC01, LW09, MTB00, Rod09, TMKL07, TK03b, TSJ01, VSO04, WD04]. Introductions [CTNR02, DDK05, KS05, KM00, Rue05]. Introgresión [FH06]. Introgression [FH06]. Inundable [SSY06]. Inundables [GRB02].

Inundación [RR00, SK02b]. Inundaciones [SMG00]. Invader [Tow03]. Invariantes [NL05]. Invariants [NL05]. Invasion [CMC03, CCH03, DOGN08, EB09, FRD^{+}09, Gay02, GJD05, GS08, HFR07, KPJ06, MCK06, MH08, NMB09, PJ00, Par01, PAHM09, SB04, CMC03, DOGN08, DSTPW07, EB09, GJD05, GS08, HFR07, KPJ06, MCK06, MH08, NMB09, PJ00, SB04, CCH03, FRD^{+}09, NMB09, PAHM09]. Invasiones [BCMA^{+}01, CM06, DGC05, FRR07, FBL^{+}09b, GB03, Hob01, JC07, KZA09, LD03, LM07, MNB09, NHHG^{+}05, PRL03, PA04, PR03, Ric07, VK07]. Invasions [Ano01k, BCMA^{+}01, CM06, DSTPW07, DGC05, FRR07, FBL^{+}09b, GB03, Hob01, JC07, KZA09, KG02, LD03, LM07, LSF04, MNB09, PRL03, PGD^{+}07, PA04, Ped07, PR03, Ric07, VK07, WWH03]. Invasiva [SMG00]. Invasive [APCS05, AL03, ACMA04, BSW^{+}08, BS09, BTK08, Car04, CTNR02, CAFL03, DDAK04, DTCC03, DM04a, HHSC00, HBBD08, HH08, HH09, Hod04a, Hod04b, HF04, HDG^{+}07, JTZ^{+}08, Kee06, KFL08, KFK^{+}09, LRO^{+}08, MSOM^{+}06, MS08, NHHG^{+}05, NFM04, PTP^{+}08, RO08, RBT08, RBK^{+}05, SMG00, Tho05b, TS03, VJ07, Wit02]. Invasiveness [LARR03]. Invasividad [LARR03]. Invasivo [CTNR02, FHL^{+}04, FPW01]. Invasivos [PV06]. Invasor [CAFL03, Tow03]. Invasora [HHSC00, MSOM^{+}06]. Invasoras [APCS05, ACMA04, BTK08, DDAK04, HBBD08, HH08, Hod04a, HF04, JTZ^{+}08, Kee06, KFL08, LRO^{+}08, MS08, NFM04, PTP^{+}08, RO08, RBT08, RBK^{+}05, Tho05b, TS03, VJ07, Wit02, KFK^{+}09]. Invasores [BSW^{+}08, BS09, HDG^{+}07]. Inventario [HR04c, HR04c]. Inventarios [NS09]. Inventories [NS09]. Inventory [HR04c]. Invernadero [Jan07]. Invernantes [JSHM06, LF09]. Inversión [DPTT^{+}05, MWBP07]. Inversiones [Nel09]. Invertebrados [ASM05, LFP05, MD04, RMK07]. Invertebrates [ASM05, Law06, LFP05, MD04, RMK07, Sch00]. Investigación [BR01, BRR^{+}02, CMK^{+}07, FBK00, LLV^{+}02, MS09, MCM07, PPW^{+}07, RS02c, SGG07, TG05, Bau05, BFMM02, KCR^{+}08, LB05]. Investigador [Lat00]. Investigation [BR01, LLV^{+}02, RS02c]. Investment [ADD^{+}07, BW08a, DPTT^{+}05, MWBP07]. Investments [Nel09]. Investor [Jac09]. Invisible [KTN^{+}09]. Invisibles [KTN^{+}09]. Involucrando [GMB^{+}08]. Involucrar [Dal05]. Involvement [PD03b]. Involving [Dal05]. Irish [FMPB07]. Irlandesas [FMPB07]. Irregular [SKBY03]. Irregulares [SKBY03]. irregularis [PV09b]. Irremplazabilidad [NCVBW02]. Irreplaceability [NCVBW02]. Isla [VTA02, ABE01, CMM05a, DGSS03, FYC00, PLA00, RPM09]. Islam [CFO^{+}09]. Island [ARB^{+}07, CMM05a, DOGN08, DW05a, DGSS03, EC01, Fra01, MTB00, RPM09, VTA02, FYC00, PLA00]. Islands [Ano01g, BCD^{+}08, CD05, CRGG01, CE00, CSBS01, DDAK04, HDG^{+}07, NMT^{+}04, SS02a, TPT^{+}07, TSKD07, WD06, Dav05, GB05a, HLJV07, Rut01, SMI^{+}07, TR05]. Islas [CD05, DDAK04, GB05a, HLJV07, HDG^{+}07, MTB00, NMT^{+}04, SS02a, TPT^{+}07, TSKD07, WD06, Dav05, Rut01, TR05]. Isolated [Bos07, FBL09a, GGGS00, GWS^{+}06, NT01, Pan02, SKBB00, SLKP06, War05].

Isolates [RTN01]. Isolation [CHFR09, FRD^{+}09, FRSG^{+}02, GGW09, NR03, SSY06, TM04, WH01]. Isoodon [SLCW08]. Isoso [FND06]. Israel [BDSD^{+}05, SS00]. Issues [App03, Dud02, KWL^{+}09, PBC^{+}09, RMD^{+}02]. Italia [BBBM02, BMBM06, GFTB04, MFGB07, RB07, VTA02]. Italian [Lai05, RLC^{+}00]. Italianos [Lai05, RLC^{+}00]. Italy [BBBM02, BMBM06, BG04, GFTB04, MFGB07, RB07, VTA02]. Ithaca [GB01]. Ithomiinae [MK09]. Ithomiine [MK09]. Itself [Meh05]. IUCN [AFB^{+}00, AOGMS09, dLBBG^{+}08, But03, BAKHT06, EGN^{+}05, FMPB07, GHTMR01, Hay09b, KZSV05, KG03, MCG^{+}08, MRAF^{+}07, QRF07]. Ivory [COW03, WCD^{+}08].

J [PJ00]. Jacobson [Tol07]. Jaguar [SRC^{+}02]. Jaguares [SLH00]. Jamaica [KJSH08]. Jamaican [KJSH08]. January [Ano05i, Ano08p, Ano09l, Ano09n]. Japan [KA03, PKM00, RLK^{+}01]. Japanese [SFD09]. Japón [KA03, RLK^{+}01]. Japonesa [SFD09]. Jardín [Fre04]. Jardines [HN09, MCR^{+}04, GFCP03]. Jarrah [Cra02]. Jaulas [Fra09]. javanicus [BIK^{+}06]. Jay [DW05a, MSWF00]. Jeopardy [DBB^{+}09a]. Jepson [WDL^{+}02]. Jerarquías [ZR00]. Jerárquicos [CJ08]. Job [Pre04]. Join [Tho03]. Joining [KORB02]. Joint [Ben01a, LRS^{+}09b]. Jonsson [Wil09a]. Joseph [Sha01]. Journal [Ano03d, Ano05g, Ano06f]. Journalism [Ano02e]. Juan [DGSS03]. Judging [WSPL09]. Judgment [Ano02e]. Judgments [SS06]. Judicious [RA05]. Juego [MC05a]. Juicios [SS06]. Juicioso [RA05]. Julian [Cze02a]. July [Ano04g, Ano05b, Ano01a, Ano02a, Ano03c, Ano03f, Ano04i, Ano04h, Ano04g, Ano05b]. June [Ano05h, Ano06g, Ano07n, Ano08m, Ano09j, Ano02a]. Junonia [Had00]. Jurisdicción [LB01]. Jurisdiction [LB01]. Just [Lim06]. Justice [Ful02, Ros08]. Justicia [Ros08]. Justificaciones [KZSV05]. Justifications [KZSV05]. Juvenile [LARK01, RBP01, RS02c]. Juveniles [LARK01, RSCL04, RBP01, RS02c, SL00a].

Kairomona [LS09]. Kairomone [LS09]. Kakadu [RSW05]. Kalimantan [PPLS01, VSK01, FJS04]. Kalinowski [Tem02]. Karst [SLELA05]. Katavi [MCM07]. Katavi-Rukwa [MCM07]. Kaua’i [FTC^{+}04b]. Keel [Gra01]. Keel-Billed [Gra01]. Keeping [Rom09]. Kelm [Gor08]. Kemp [LCS03]. Kendall [RSC03]. Kenia [Bor08]. Kenya [Bor08]. Kernel [CMCG07]. Kestrel [ENN^{+}08]. Key [CHBK02, DMDFV08, PLL^{+}05]. Keys [Mcc09]. Keystone [Kot00, MRH^{+}00]. Keystone-Species [Kot00]. Kibale [CBG^{+}00, CBG^{+}00]. Kind [Joh08, Kun03]. Kingbird [Mur01]. Kingdom [EGN^{+}05, PK05]. Kit [MC05a]. Kite [MKH07]. Klamath [OFS^{+}04, Owe08, SD01]. Kleijn [VW05]. Knew [Bro05]. Knot [Cre01b]. Know [Bra05b, Bro05, Smi04a, Wad02]. Knowing [AFK^{+}09, GMCP09, KCR^{+}08, RFB09]. Knowledge [AJF03, AGBP09, BTBC^{+}03, BPP05, BBW^{+}07, Dre05, HP06, LW06, RNP^{+}07, VP06, You02].

Knowledge-Intensive [VP06]. Known [Hei09, HW02, SS06, Swe08]. Koala [CMBC00, CMC^{+}00, CHF00, CCM00b, LMMF00, MCM^{+}00, PSG^{+}00, Phi00, RWM^{+}06, STWH00, SMLB00]. Komodo [WGW01, WGW01]. Kona [TH03]. Kong [CMA^{+}06]. Kremen [Gha08]. Kruger [Ant08, Ant08]. Kuznets [LME08, LME08]. Kyoto [BS03, Nos01a, SS02b].

Laband [BRTW07]. Laboratories [Bre02a]. Labors [Orr02b]. Labranza [SMC00]. Lacertid [DPTT^{+}05]. Lacértida [DPTT^{+}05]. Lack [UWB04]. Lag [GWJ^{+}07]. Lagartija [DPTT^{+}05, TR05]. Lagartijas [BD06, KGS^{+}07, SG01]. Lagged [Ste05]. Lago [CCS^{+}03, FPF00, MMF^{+}05]. Lagoon [KTN^{+}09]. Lagopo [TRNH00]. Lagos [DDC^{+}00, DN00, MSOM^{+}06, BGD09, DDC^{+}00]. Laguna [KTN^{+}09]. Lake [CCS^{+}03, FPF00, MMF^{+}05]. Lakes [BGD09, DDC^{+}00, DN00, MSOM^{+}06]. Lakeshores [DDC^{+}00]. Lampiño [TRSR00]. Lanchas [Lus03]. Land [ABE^{+}02b, BT07a, Cua00, CAR05, Cze02b, DBL06, DN01, FB08, Gav04, GM01, GGW09, HR02, HNG04, JMS^{+}06a, Kau04, KB05, KSDBT02, KB08, KFNB09, LBPBG07, LLWCW09, MKG03, MNRD^{+}08, MV06, MNL03, MPR^{+}01, MHGF04, NRBM05, Nor00, Nug08, Orr08a, PA04, RJW09, Say05, SLELA05, SP09, Sei04, SVBR07, SF05, SCC^{+}09, VW04, VW05, WPS08, WTEK08, WMM04, WM01]. Land-Cover [Cua00, GGW09]. Land-Use [DBL06, KSDBT02, LBPBG07, MKG03, MPR^{+}01, NRBM05, SCC^{+}09]. Landbird [AM05]. Landowner [BZD03]. Landowners [Jam02]. Landownership [Say05]. Lands [BMRB04, BRK^{+}04, DBB^{+}09a, DWW04, FB00, HM03, LMMF00, MYFF^{+}07, MH01b, MN07, NSB^{+}06, PP03, Sul09, WE07]. Landsat [DDN01, WM01]. Landscape [APSB02, ARMH04, Ano03e, BNH02, Ber05, BFD07, BKWB05, BJB06, BV01a, BBH06, Bur04, CPSS03, Car07, CPD^{+}03, CPO03, Cur07, DeS03, DM04b, DD04b, FJ00, FPSD09, FLF04, FTR01, FCL00, GB03, GM00a, GM00b, Gra01, GBC^{+}04, GGW09, GH02, HR04a, HLB^{+}03, HCZ00, Hol01, JMH01, KSO^{+}03, KFNB09, LMS^{+}02, LB02, ML04, MU08, Nas06, NS01, PDF^{+}00, PA04, PGP04, PV02, PDS^{+}07, PMC01, PBL00, RCS^{+}03, RDEF01, RAR^{+}04, RP07, RS02c, Rot01, RK05, SFX^{+}08, SP00, Sod06, SD03, SB08, SBSL06, VVMV09, VDP00, WM01, WCPB03, WP03, Wie06, WMD^{+}04, WMS^{+}01, WBV^{+}05, Wit02, YL05]. Landscape-Scale [BV01a, LB02]. Landscapes [BW09, BMPNC^{+}03, BTS09, CNPS04, Cur02b, DCP^{+}03, DWW04, Don06, FEKH09, GLF^{+}08, HMB^{+}05, HKC^{+}08, HSB01, JTS06, KBS^{+}04, LLS08, LZD^{+}03, MM02b, MPE^{+}07, NA06, PJB^{+}06, PJYO04, PD03a, PFR^{+}09, PAA^{+}08, RL06, RMN00, RWM^{+}06, RBG^{+}04, SHYH06, TNFH07, Vee06, WFP01]. Lane [McC00, MSWF00]. Langebaan [KTN^{+}09]. Langostinos [PFL01]. Language [Arm02, Bag09, Pri01a]. Larga [Ber04a, LBG03, VK07]. Large [Ano04e, BVL^{+}03, Bus02, Côt05, FDL^{+}05, FN04, FMSL08, HBV02, HAG^{+}09, JC07, KSO^{+}03, LSZ^{+}08, Mcc09, PA08, PMC01, Pol06, ROB^{+}04, RB07, RCL^{+}06, SRW^{+}08, SCS06, SCM^{+}07, Swe08, UWB04, VGB^{+}04].

Large-Mammal [Bus02]. Large-Scale [FDL^{+}05, HBV02, RCL^{+}06, SRW^{+}08, UWB04]. Larger [ZGR02]. Largest [LCZX09, RJW09]. Largo [BSW^{+}06, CWL^{+}00, CWH07, CBG^{+}00, DBML07, FTC^{+}04b, GMW08, GWJ^{+}07, Gut05, HKN01, IMD01, LLH08, PJ00, PPW^{+}07, RR00, SRW^{+}08, Sch03a, SRH^{+}07, SMH^{+}07, SLKP06, SBP07, SBSL06, VEST02, WBDP02, WGGS07, BMA^{+}08, GFTB04, MKG03]. Largos [WE07]. Larinus [LO02]. Larkspur [KSB01]. Larval [BS01b]. Last [Ber04a, PDS^{+}07, Rol00]. Late [LW02b, RWS02, SK02b]. Late-Season [SK02b]. Late-Seral [RWS02]. Lateral [MBC03]. Laticauda [SS02a]. Latin [Ank04, GHL08, MGB^{+}07, PAA^{+}08, Rod03, YLR^{+}01]. Latinoamérica [GHL08, PAA^{+}08, Rod03, YLR^{+}01]. Latitudinal [KPH07]. Laúd [TSP^{+}08]. Laurentian [MSOM^{+}06]. Law [Fre06, HH06a, Led06, RD05]. Laws [HBLW09, KK09, MA05, TAJ04, Orr04c]. Leadership [Cla01, DAB^{+}04, MCF^{+}09]. Leading [BC09, GMCP09, GVS09]. Leading-Edge [GVS09]. Leaf [ARP05, EGPS04, FMLL06, GB05b, HVR08, NBM^{+}04]. Leaf-Chewing [FMLL06, NBM^{+}04]. Leaf-Litter [ARP05, HVR08]. Leap [Edw04, Lat00]. Learn [Jac09, Ran00]. Learned [WL04]. Learning [CW08, Kni06, LW06, Mae02, MSL06]. Leasing [BMRB04]. Leatherback [TSP^{+}08]. Leave [Hun07a]. Lecciones [CSS00, FC05, KB08, Lat00, MB09b, PW01, Rig01, SMH^{+}07, TRCC^{+}05, Wal03, WGW01, WL04, TG05]. Lectures [Mae02]. Leg [JGL^{+}02]. Leg-Banding [JGL^{+}02]. Legacy [Chr08, Pre09, Yul06]. Legal [DZM08, HMF^{+}04, MPFBH07, RD05, VNP06, WABT08]. Legales [HMF^{+}04]. Legally [PBSR06]. Legalmente [PBSR06]. Legged [DR03, KM00]. Legislation [MPFBH07]. Legitimacy [Pro04]. Length [Had00]. Leñosas [KPJ06, NBM^{+}04]. Leñoso [YE07]. Leñosos [LGK02]. Lens [Bru05, Hel08]. Lento [KHP04, UWB04]. Leones [GSG08, WSQP07]. Leopard [MAM^{+}03, SWR00]. Leopardo [SWR00]. Leopardos [KD00]. Lepidoptera [DCD^{+}02, SBVC03, DCD^{+}02]. Lepidopteran [TS09]. Lepidópteros [SBVC03, TS09]. leptorrhynchoides [Mor00b]. Leroux [LMR^{+}07]. Lesser [Hei09]. Lesson [Mef09a, Ova03]. Lessons [Bus02, CSS00, FC05, Gal01b, Hei09, KB08, Lat00, MB09b, PW01, Rig01, SMH^{+}07, TG05, TRCC^{+}05, Wal03, WGW01, WL04]. Let [Ber03, ML00]. Letal [BPB^{+}05]. Letales [PRL^{+}02, STC03, VT09]. Letanía [BADF^{+}05]. Lethal [PRL^{+}02, VT09]. Letter [Ano04f, Dal00a, Smi04b, SLTG04]. Letters [Bal04, Car02a, CMC05, CHBK03, DP03, Leo03, RN04, RA04c, SS03a, Ter05, Cle06a, Cze07b, KWV08b]. Levantamiento [LMMF00]. Level [BB06, FCG^{+}05, Kra01, MMF^{+}05, Rua00]. Levels [Alt05, BM05b, CVS^{+}09, CY08, GHTMR01, GAF00, PZG^{+}05, RM03, TG04]. Leverage [Kra01, Orr01a]. Lewis [MS02]. lewisi [ALH^{+}04]. Lexicon [SSS^{+}08, SSS^{+}08]. Ley [HH06a]. Leyendo [Pyn04]. Leyes [TAJ04]. Liana [GV02]. Lianas [GV02]. Liberables [Bau05]. Liberación [KLH08, LRS09a, SL00a, TAP^{+}01, RSCL04]. Liberaciones [MS04]. Libre [PIH00]. Libres [FPW01]. Lice [Fra09]. Lichen [BSS^{+}05, HW02]. Lichens [BJ05, MJ03]. Liderazgo [Cla01, MCF^{+}09]. Life [Ano01h, Bra03a, BMMW07, CVS^{+}09, FDBM04, KD05, Kre07, Mig00, MSWF00, Nos01b, OHC08, Orr06b, Orr09c, SFD09, Wed02]. Life-History [Bra03a, CVS^{+}09, FDBM04, KD05, SFD09, Wed02]. Lifting [DD04a]. Ligando [KLH08]. Light [Fie05, PJ00, LM07]. Lighting [BBM04]. Lightness [Orr09b]. Like [CGR^{+}05, Joh07, Per02]. Likely [PC08]. Limb [JLZW03]. Limestone [BKK03, MBQAM06]. Limit [FV05]. Limitación [FJ07]. Limitaciones [VO03, HLJV07, SWK05]. Limitada [DSE^{+}08, LCS04, WBSG07]. Limitadas [AP00, NWRD03]. Limitado [MLB08]. Limitados [SGSA07]. Limitando [OWOF03]. Limitantes [Fer01]. Limitar [BM05a]. Limitation [FJ07]. Limitations [Fer01, HLJV07, MF08, SWK05]. Límite [FV05]. Limited [AP00, DSE^{+}08, LCS04, MLB08, SGSA07, WBSG07]. Límites [Ehr05a, GDAV00, Gui07, NTMT^{+}03, RCMV06, TNRJ02]. Limiting [BM05a, Dal09, OWOF03]. Limits [Ehr05a, Lam06, TNRJ02]. Linaje [Kel01]. Lince [Car07, PDF^{+}00]. Lindenmayer [Lam02]. Lindero [MW07]. Linderos [PW01]. Lineage [Kel01]. Líneas [AYKM04, KLH08]. Lines [AYKM04, KLH08]. Link [Ben02]. Linkage [PCHN04]. Linkages [BMS08, CWCG02, Mad05, PB04, Pyk05]. Linked [HCZ00]. Linking [KLH08, SGG07, SVBR07, TFG08, TTSP04]. Links [BGHM06, BGHMb07, DHZM09, SMW^{+}07]. Linnaeus [SGLB^{+}07]. Lion [UCAGG08]. Lions [GSG08, HY03, MB09b, RBG06, WSQP07]. Líquen [HW02]. Líquenes [BSS^{+}05, BJ05]. List [dLBBG^{+}08, But03, BAKHT06, GHTMR01, Hay09b, KG03, QRF07]. Lista [dLBBG^{+}08, Gus02, Hay09b, HW08, HKM06, KG03, Nos03, QRF07, FEGB09]. Listado [BZD03, HP01, EGN^{+}05]. Listados [DP01]. Listas [DP01, FKU05, FMPB07, GM01, KB04a, KZSV05, GLMN00]. Listed [Gus02, HW08, HKM06]. Listing [BZD03, DP01, EGN^{+}05, FKU05, Fal07, FEGB09, HBC^{+}06, MRAF^{+}07, MPFBH07, QR05, RFB09]. Listings [HP01]. Lists [Bal02, FMPB07, GLMN00, GA02, GM01, KB04a, KZSV05]. Litany [BADF^{+}05]. Literacy [Bre01, TOR^{+}04]. Literatura [BADF^{+}05]. Literature [BADF^{+}05, New08, Orr01b, Sal02, Whe03]. Lithophaga [GFTB04]. Litoral [GFTB04]. Litorales [CSM^{+}06]. Litter [ARP05, HVR08]. Little [NB00, SS06]. Little-Known [SS06]. Littoral [CSM^{+}06, GFTB04]. littoralis [ARB^{+}07]. littorea [VSAG06]. Live [CMC03, DDMMZD06, MH02, RGW09, SVV07, WSCP05]. Live-Bearing [DDMMZD06]. Live-Vertebrate [RGW09]. Lived [KHP04]. Livelihood [SBK^{+}06]. Livelihoods [Ano02j, XGS^{+}09]. Liverworts [MJ03]. Lives [Whi06]. Livestock [Ber06, Bor05, BPB^{+}05, Cur02b, FFHW03, GM00b, Hob01, Mad05, MM07, MMB^{+}03, OWOF03, RIBGG08, TNTH^{+}04, VMBC02]. Living [BOJ09, Ehr05b, Ray06b, CLW^{+}09]. Lizard [DPTT^{+}05, FSC02, KGS^{+}07, TR05]. Lizards [BD06, SG01]. Llamados [Roy04]. Llamas [Pyn04]. Llanura [SK02b]. Llanuras [GRB02, GSB04, HBV02, Say05]. Llegada [PHW05]. Llevada [WPS08]. Llevado [GPZ04]. Llevando [DLCV00]. Lluviosa [MO04, NBM^{+}04, WA05]. Lluvioso [ARPEBM09, EAA^{+}09, FBL^{+}09b, HN09, JHHL04, LCT^{+}06, LLH08, LP07, Ram01, SWD00, SBSL06, WLCS00]. Lluviosos [FR07, LCG^{+}08]. loa [FBP^{+}09]. Load [JGP09, LRS^{+}09b]. Loafing [RS02b]. Loblolly [LGK02]. Lobo [HLSZ02, HY03]. Lobos [BPB^{+}05, CPSS03, CFH^{+}02, FH06, MMB^{+}03, NTGT03, RLC^{+}00, TNTH^{+}04, UCAGG08]. Local [AGBP09, Baw06, CPO03, DBB^{+}09b, DMG^{+}00, GM00b, HGH^{+}03, KDD^{+}04, KFBG08, Kra01, LP01, LRS^{+}09b, LB02, MDCK07, MGH^{+}09, NBM^{+}04, ØDEA00, PRL03, PH02, PN05, Pow06, RA04b, VL01, VSO04, VBOV05, WSKH08]. Locales [DMG^{+}00, GM00b, LP01, LB02, NBM^{+}04, Pow06, WSKH08]. Localidades [BGD09, LCS04]. Location [PCL08]. Locations [BGD09, LCS04, WBW05]. Logged [DMSP00, EAA^{+}09, GAF00, Olu00, SH09, TK07]. Loggerhead [LCS03, PLB09]. Logging [Ano01k, CBG^{+}00, CPME09, CPR05, GV02, GJLL08, Hut06, KG02, Lew01, LN06, LO06, NL06, PMZ06, PCME09, PDH00, PPPV08, RBRB06, SSSK06, SRH^{+}07, VSK01, WLCS00]. Logic [O’C00]. Logical [Ros02]. Logistic [CB01b]. Logística [CB01b]. Logrando [DPB^{+}09, KB06]. Lombardia [BBBM02, BMBM06]. Lombardy [BBBM02, BMBM06]. Lomborg [BADF^{+}05, Moo02]. Lombrices [Gun02, GJD05, HFR07, MNB09, NMB09, SZBR03]. Long [BSW^{+}06, Ber04a, Bru05, CWL^{+}00, CWH07, CBG^{+}00, Doa01, DBML07, FTC^{+}04b, GMW08, Gut05, IMD01, KHP04, LLH08, LBG03, PPW^{+}07, SRW^{+}08, Sch03a, SMH^{+}07, SLKP06, SBP07, SBSL06, VEST02, VK07, WLS^{+}08, WBDP02, WE07, WGGS07]. Long-Awaited [Doa01]. Long-Distance [Ber04a, LBG03, VK07]. Long-Lived [KHP04]. Long-Term [BSW^{+}06, CWL^{+}00, CWH07, CBG^{+}00, DBML07, FTC^{+}04b, GMW08, Gut05, IMD01, LLH08, PPW^{+}07, SRW^{+}08, Sch03a, SMH^{+}07, SLKP06, SBP07, SBSL06, VEST02, WLS^{+}08, WBDP02, WGGS07]. Longevas [KHP04]. Longhorn [BRA08]. Longleaf [aARLG03, MDJP04]. Longline [LC07a]. Loop [PR08]. Loros [WTEH^{+}01]. Losers [WCX09]. Losing [BT07a, PJB^{+}06]. Loss [BGHMb07, BJTD03, BMM^{+}02, BV00, CM03, CVS^{+}09, DBL06, DAR08, ENN^{+}08, Fai08, FMSL08, FJS04, GRH05, Gib00, HRS08, HC09, HPG09, Kel01, LM07, Mcc09, NNM00, PLL^{+}05, PAA^{+}08, RLK^{+}01, Ros06, Ros08, SAA^{+}01, SB04, SMW^{+}07, WRM05, Wie06]. Losses [BCS04, ESB05, FKU05, Koh07]. Lost [GWM^{+}07]. Lot [Ran00, Sha04]. Lotus [ME00]. Love [Bat04, Pim06]. Low [CH01, GPMM08]. Lower [FKU05]. Lowland [BBBM02, EBDM06, GAF00, Lam04, Mcs05, NBM^{+}04, SSY06, WG01]. Luck [BV01b, Tut03]. Lugar [FN02b, MS09]. Lugares [CML02]. Lughadha [MM03b]. Lumas [SS01]. Lure [Fre04]. Lurking [MM00]. luteiventris [WM05]. Luz [PJ00]. Lying [RS02a]. Lyman [MS02]. Lyme [AKO03, OK00]. Lynch [Ste08a]. Lynx [Car07, PDF^{+}00, RSF^{+}03].

Macaco [CMC^{+}01]. Macaques [CMC^{+}01]. Machos [Olu00]. macradenia [HH06b]. Macroecología [GM01]. Macroecology [GM01]. Macroeconomic [Law08]. Macroeconómica [Law08]. Macrofauna [MRM^{+}05]. Macroinvertebrados [Has00]. Macroinvertebrate [Has00]. Macrolichens [BSS^{+}05, Pyk04]. Macrolíquenes [BSS^{+}05, Pyk04]. macropomum [RW02]. Maculinea [Wal04]. Madagascar [ACC^{+}05, CSM^{+}06, GZK06, JAH^{+}05, JAH08]. Madera [CWH07, Cra02, FMLL06, HKM06, JMS06b, Log03, MR00a, PPLS01, Pow06, RBRB06, SLS05, SBB^{+}08, ZSVL07]. Maderable [FJ07, PPLS01]. Maderables [PL03, TNRJ02]. Maderera [PCME09]. Madereras [CPME09, PDH00]. Madereros [SRH^{+}07]. Maduro [WSLW00]. Maduros [STL02, SDJ06]. Magallanes [WBW06]. Magellanic [WBW05, WBW06]. Magnetic [Ber04b]. Magnitud [ØDEA00]. Magnitude [ØDEA00]. Mahogany [GB05b]. Maine [JMH01]. Maintained [FTC04a]. Maintaining [FPL^{+}04]. Maintenance [FPF00, MA08]. Maíz [PHW05]. Maize [PHW05]. Major [EYZ09, KWL^{+}09, MKVZ^{+}09, PBC^{+}09, WGGS07]. Majority [Dun05]. Make [Pim06]. Making [AFB^{+}00, Blu07, BBW^{+}07, Coc04, Dix07, Gro05a, Kar01, Lat00, MDW01, Pra05, RM03, SDL07]. Mala [BV01b]. Malasia [Lai00, SLELA05]. Malay [SBK^{+}06]. Malaysia [Lai00]. Malaysian [SLELA05]. Male [Olu00, Wil02d]. Malformed [Joh09b]. Mallee [Cun00]. Malos [Bat04]. Malthusian [PS01b]. Mamífero [BV04, GM08a]. Mamíferos [Ber04a, Bra03a, BVL^{+}03, CB01a, CWC00, CKL^{+}00, Chi00a, CT07, CLMD05, Cro02, Cua00, DCP^{+}03, FDBM04, GRH05, HM04, JG08, JP09, KSO^{+}03, Lai00, LCT^{+}06, LCG^{+}08, LLH08, LMP00, LP08, LP07, LF00, MBR^{+}02, MR00a, MP07b, NTMT^{+}03, OK01a, PH02, PJNA07, RW00, RDN06, RB06, SCS06, TWDB00, Wal03]. Mammal [BV04, BCS00, Ber04b, Bra03a, Bus02, CWC00, CLMD05, Dud02, GA02, GM08a, JG08, LLH08, MR00a, MP07b, OK01a, Ray06a, RS01, WZD^{+}00]. Mammalian [Cro02, GRH05, HM04, PH02]. Mammalogy [Cua02]. Mammals [Ano00b, Ber04a, BVL^{+}03, CB01a, CKL^{+}00, Chi00a, CT07, Cua00, DCP^{+}03, FDBM04, GNR01, JP09, KSO^{+}03, Lai00, LCT^{+}06, LCG^{+}08, LMP00, LP08, Lis04, LP07, LF00, MBR^{+}02, Mar00a, NTMT^{+}03, PJNA07, RW00, RDN06, RB06, SCS06, Swe08, TWDB00, Wal03]. Mamushi [SFD09]. Man [Por08]. Manada [TRSR00]. Manage [MML06, Sim03]. Managed [BMPNC^{+}03, FPL^{+}04, FTR01, TK07, WCPB03]. Management [Ano01j, BMP^{+}06, Bei03, BS01a, BRK^{+}04, BTK08, Blu07, BMMW07, Cha06, Che02b, Cho06, CMM05a, CSB07, Cor06, CV08, DN01, DW06, DM04a, EGPS04, EK08, FBK00, Ful00, GFCP03, Gav04, GEJK05, GWS07, GPC^{+}03, GB05a, GMB^{+}08, HM08, HM09, HH09, HCM03, Hut06, IF02, JMS^{+}06a, JWB06, JML00, KEJ06, Kar09, KF01b, Kee06, KB06, KT06, KB08, Kri04, LOM03, LSL^{+}06, LMB00, Lin03b, LRS09a, LGB05, Mag01, MDH^{+}07, MCA^{+}00, MOO09, MP07a, ND03, Nos01a, NR03, Ped03, PSF05, PLA00, PB09, PPBW08, PR08, PS04, PK05, PS06, RJW09, RT00, RGF^{+}04, Rig01, Rob00, RSW05, Roo00, Rud03, SC09, SMM09, SP09, SSSK06, SD02, SLCW08, SS06, TG05, TWDB00, TRNH00, TAJ04, TK03a, VS09, VS05a, VGB^{+}04, VBOV05, Wan04, WELCH09, WDA^{+}06]. Management [Wed02, WS03, Wil02a, WD03, WSPL09, WBV^{+}05, ZSVL07, ZSG01, KB08, PK05]. Management-Plan [PK05]. Managers [Hay03, Jon04, MC05a]. Managing [FRD^{+}09, FTC04a, HLSZ02, HH08, HH06b, LRO^{+}08, LB00, MR00b, NARL^{+}04, RBT08, RMN00, STC03, SMLB00, Whe02]. Manatees [AJF03]. Manatíes [AJF03]. Manchado [Fau01]. Manchados [SS03b]. Manchones [PMC01]. Manejado [FTR01]. Manejados [BMPNC^{+}03]. Manejar [Sim03]. Manejo [BTK08, BMMW07, CSB07, CV08, DN01, EGPS04, EK08, FRD^{+}09, FPL^{+}04, FBK00, GWS07, GPC^{+}03, GMB^{+}08, HM08, HLSZ02, IF02, JML00, LOM03, LRO^{+}08, LMB00, Lin03b, LRS09a, MCA^{+}00, MOO09, ND03, Nos01a, Ped03, PLA00, PB09, PPBW08, RJW09, RMN00, Rig01, Rob00, RSW05, SC09, SMM09, SP09, SLCW08, SMLB00, TWDB00, TRNH00, WCPB03, WELCH09, Wed02, WS03, WD03, WBV^{+}05, ZSVL07, ZSG01, FTC04a, HH08, MR00b, MP07a, RBT08, Wil02a]. Mangabey [Olu00]. Manglares [RLK^{+}01]. Mangrove [RLK^{+}01]. maniculatus [PLA00]. Manipulación [MR00b, RG04a]. Manipulaciones [Pop08]. Manipulation [MR00b, RG04a]. Manipulations [Pop08]. Mantener [FPL^{+}04, RR00]. Mantenida [FTC04a, GLMN00]. Mantenimiento [FPF00, MA08]. Manu [Ano01i, OSSK^{+}07, OSSK^{+}07]. Manual [Car06, Hay03]. Manufactura [BBW^{+}07]. Manuscripts [Ano03d]. Many [BCK^{+}09, LP05]. Maori [RNP^{+}07]. Map [CMS09, VS05b]. Mapaches [Log03]. Mapas [LMMF00]. Mapeador [DDN01]. Mapeo [BKV^{+}05, BTBC^{+}03, MD04]. Mapper [DDN01]. Mapping [Ano03e, BKV^{+}05, BTBC^{+}03, LMMF00, MD04, Moi08, RCS^{+}03]. Mar [FCG^{+}05, FV05, Fra09, GPZ04, FMSL08, Wol00]. Marbled [CLBC03, GGW09, MM02b, PBN^{+}04, Rap06, TSG06]. Marca [LG05a, PV06]. Marcadores [LCS04]. Marcaje [MP08, MP08]. March [Ano04h, Ano08p, Ano04j]. Marchetti [LM07]. Marco [Dal05, EG07, GWS07, GBC^{+}05, KSC^{+}06, MU08, SMRR02, WABT07a, Rua00, TAP^{+}01]. Marcos [BFMR07]. Mareas [RLK^{+}01]. Marfil [COW03, WCD^{+}08]. Margarita [Mor00b]. Margen [Car07]. Margin [Car07, Orr02a]. Marginal [AW01]. Marginales [AW01]. Marimono [PS00a]. Marina [AJ05, BMRB04, BB06, CY08, DLCV00, DPB^{+}09, Dre05, FC05, GPZ04, GB05a, GM08a, KTN^{+}09, LB05, Les05, LV04, LG05b, SS02a, Slo02, TFG08, VTA02, ZR00, ŽWGW09]. Marinas [BPV09, BM05a, BSKH^{+}06, CNM01, Dal05, Dav05, Dav07, DMS03, FVS03, FDL^{+}05, FCG^{+}05, Fra09, FNS^{+}03, GMMPP08, GWS07, GC09, GHL08, Gui07, JTZ^{+}08, KCK^{+}08, Kri04, LB05, LC07a, LB01, LARK01, Lyn06, MC09, MA08, RKEJP06, RBP01, RBHKK02, Rut01, VTA02, WSCP05, ZG05]. Marine [AJ05, BPV09, Bea07, BMRB04, BSKH^{+}06, CY08, Cor04b, CNM01, CV08, Dal05, Dav05, Dav07, DMS03, DPB^{+}09, Dre05, ESB05, EK08, FVS03, FDL^{+}05, FC05, FNS^{+}03, GMMPP08, GEJK05, GWS07, GPZ04, GB05a, GM08a, GHL08, Gui07, HSMK07, KTN^{+}09, KCK^{+}08, Kri04, LB05, Les05, LK04, LB01, LARK01, LV04, LGB05, LG05b, Lyn06, Lyn08, MBR^{+}02, Mar00a, Mas03, MA08, PIH00, RT00, Ray06a, Ray06b, RW00, RDN06, RKEJP06, RBP01, RBHKK02, Rut01, SC09, SRSP01, SSB^{+}08, SB06, Slo02, Ste08a, TFG08, TWDB00, TSP^{+}08, VTA02, Wal03, WSCP05, ZR00, ZR01, ZG05, ŽWGW09, GC09, MC09]. Marine-Reserve [RKEJP06]. Marino [ESB05, GM08a, Kai03, LARK01, MHJ^{+}09, PIH00, SC09, SSB^{+}08, TSP^{+}08]. Marinos [CY08, CV08, EK08, GEJK05, GSG08, HSMK07, HY03, MBR^{+}02, RW00, RLK^{+}01, TWDB00, UCAGG08, Wal03, ZG05]. marinus [BS09]. Mariposa [BTBC^{+}03, BCP^{+}02, CWNW01, Gut05, Had00, HJ05, SH03, Wal04]. Mariposas [BE00, BKK03, BTS09, CPO03, FNFM01, FN02a, FNF03, FM09, HHD^{+}08, HDDEB03, KSB04, KT02, Lew01, MKK^{+}06, MKK^{+}09, MM03a, MK09, NFM04, RDW01, SC05, VVMV09, WLCS00, VMV07]. Mariscos [CMC03, ES04]. Marisma [BCHS08]. Marismas [BCS04, BS08, HBA09, SB04]. Mark [LG05a, PV06]. Markers [LCS04]. Market [CMA^{+}06, Hal02, HAG^{+}09, Pow06, RH08]. Marketeers [Orr05b]. Markets [FYC00, Mad05]. Marking [MP08]. Markov [Lus03]. Marmoleado [MM02b]. marmoratus [CLBC03, PBN^{+}04, Rap06]. Marmota [GVTM07]. Marmots [GVTM07]. Marsh [BCS04, BCHS08]. Marshes [BS08, HBA09, SB04]. Marta [Car07, PBL00]. Marten [Car07, PBL00]. Martin [HH09, LSF04]. Maryland [IB03]. Más [Gui07, LCZX09, HJ05, Nel09, RJW09, RH01, WMLS09, Dal05, DWWF04, RH01, SDJK04, ZSVL07]. Masa [Olu00, TTB03]. Mascotas [RSW05]. Masiva [Fon09, PMKDS01]. Masivo [GWJ^{+}07]. Mass [BEH06, Fon09, GWJ^{+}07, Olu00, PMKDS01, TTB03]. Massachusetts [FD00]. Master [Mai04]. Masterful [Col06]. Masticadores [NBM^{+}04]. Mata [Ibi08]. Mate [GPMM08, RG04a]. Mating [EBAW01, FLQ03, LM05]. Mating-System [EBAW01]. Matorral [BRA^{+}03b, CKL^{+}00, RSW05, Rub01]. Matorrales [DNZK02, RAW07]. Matrices [Wil03]. Matricial [Wil03]. Matriciales [BMP^{+}06, Tho05b]. Matrix [BMP^{+}06, BTGS05, CS06, JMLG01, MK09, NAC^{+}06, PV02, Tho05b, VP07]. Matriz [BTGS05, CS06, JMLG01, MK09, NAC^{+}06, PV02]. Matsigenka [OSSK^{+}07]. Matter [Bal02, BV01b, DA02, Ter00, Dal00b]. Maturation [Che02b]. Mature [SDJ06]. Mauritius [ENN^{+}08]. Maximizar [CDD^{+}05, FTC04a, FPL^{+}04, TTSP04]. Maximize [CDD^{+}05, FTC04a, TTSP04]. Maximizing [FPL^{+}04]. Máximo [Kes01]. Maximum [Kes01]. May [Ano04j, DP01, For02, KG02, Ste08b]. Mayor [JC07, NCVBW02]. Mayores [CSS00, FN04, MKVZ^{+}09, PBC^{+}09, RB07, SCS06, SCM^{+}07, ZGR02]. Mayoría [Dun05]. Mbaracayu [HMF03, HMF03]. McCoy [Orr05b]. McMurdo [ANW^{+}08, ANW^{+}08]. McPherson [Dal00b]. McSweeney [EBDM06]. Me [Ebe04]. Meadows [Cas03, HBSD04, SD03]. Meandering [RR00]. Meaning [Bre06, Joh09b, MLMB08, NVP07, VNP06]. Meaningful [Hut06]. Meaningfulness [Wol06]. Means [Meh05, PBSR06]. Measure [BNP^{+}07, FYC00, GS08, Jac08, SDJK04, WPS08]. Measurement [Wol06]. Measures [Ant08, CLŽ^{+}07, DJOW07, FPW01, PEC01, RM06]. Measuring [BKV^{+}05, GM09, JGM03b]. Meat [Ano02j, Ben02]. Mecanicista [OIK^{+}06]. Mecanismos [You05]. Mechanismos [PJ00]. Mechanisms [GRG^{+}05, Hal02, Rob00, You05]. Mechanistic [OIK^{+}06]. Medellín [CD07]. Medianate [ŽWGW09]. Mediante [BOJ09, BRV06, BMB^{+}08, Cze08, DKS^{+}09, FEKH09, HW08, JGL^{+}02, LG05a, MR00b, Pol06, Pyk00, VTA02, Rap06]. Mediated [MP07b]. Mediates [PM05]. Medicina [LG02, PK01]. Medicinal [Kal05, OBM09, Rob00, Sie00]. Medicinales [Kal05, OBM09, Rob00]. Medicine [Jes03, LG02, Laf03, Oso00, PK01, Spe00]. Medición [BKV^{+}05, Wol06]. Medida [FYC00, GS08, SDJK04, WPS08]. Medidas [Ant08, BMMW07, CLŽ^{+}07, DJOW07, FPW01, PEC01, FBL09a]. Medio [AAO^{+}07, ESB05, Fau01, KLH08, MLB08, PBSR06, RHS05, Gav04]. Medio-Occidente [Fau01]. Medios [XGS^{+}09]. Medir [BNP^{+}07, GM09]. Mediterránea [ZGR02]. Mediterranean [DMDFV08, Fat06, FMSL08, KS05, MTB00, STP09, SRHS09, UKC^{+}09, ZGR02]. Mediterranean-Climate [SRHS09]. Mediterráneo [DMDFV08, Fat06, FMSL08, MTB00, SRHS09, UKC^{+}09]. Mediterráneos [KS05]. Medium [OK01a]. Medium-Sized [OK01a]. Meet [Arc09, Cob03, O’N07]. Meeting [Ano03a, Büs08]. Meets [Chi00b, Lei05]. Megadiverse [Rod05, VIMMO05]. Megadiverso [Rod05, VIMMO05]. Megafauna [ŽWGW09]. Megareservas [Per05a]. Megareserves [Per05a]. Mejillones [BJPL07b, PHW05]. Mejor [KDC^{+}06]. Mejora [VW04]. Mejoramiento [CHF00, NFD^{+}09, TG05, KRM^{+}00, LBR07, SMRR02]. Mejorando [CHBK02, PP08a, VO03]. Mejorar [GBC^{+}05, GBE^{+}06, KMGB07, RWM^{+}06, WMVK03]. Mejoras [HP01]. Mejores [GGT08]. Melanesia [CMM^{+}05b]. melanogaster [DH00a, KLH08]. Meltdown [GK09]. Melting [Pan09, XGS^{+}09]. Membership [CNHT09, Gra09, SHB08]. Memoriam [Ano07l]. Memories [SF00]. Mercado [CMA^{+}06, HAG^{+}09, Pow06, Rob00, RH08]. Mercados [Mad05, FYC00]. Mercurio [UMT^{+}01]. Mercury [UMT^{+}01]. Meretsky [BS01a]. Mérgulo [MM02b]. Meridional [MNL03, RBHKK02]. Meritorias [MB00]. Mesa [DDBA^{+}07]. Meseta [DDBA^{+}07]. Mesic [HH03]. Mésicos [HH03]. Meso [Sch03b]. Meso-depredadores [Sch03b]. Mesoamerican [HKC^{+}08]. Mesofilter [Hun05]. Mesofiltro [Hun05]. mesopotamica [ZSSM09]. Mesopredator [Sch03b]. Messages [Ibi08, Smi05]. Meta [BBB08, CTR02, VW05, VMM07]. Meta-Análisis [BBB08]. Meta-Analysis [BBB08, CTR02, VW05, VMM07]. Metamorfosis [BSM08]. Metamorphosis [BSM08]. Metapoblación [Arm05, Bos07, Esl00]. Metapoblacional [FRSG^{+}02, KKHS05, SLH00]. Metapoblacionales [LMP00, MT01, WBV^{+}05]. Metapoblaciones [ARMH04, FCL00, GC06, HSB01, JGM03a, LP01]. Metapopulation [ARMH04, Arm05, Bos07, Esl00, FRSG^{+}02, HSB01, KKHS05, LMP00, LP01, MT01, SLH00, WBV^{+}05]. Metapopulations [FCL00, GC06, JGM03a]. Metas [DS08, ERB01, SEBM03, SBP07, PKB^{+}08]. Metasequoia [LCZ^{+}05]. Method [AGBP09, BMBM06, BNP^{+}07, BPB^{+}05, BPL^{+}01, MP09, MW07, PBK09, QRF07, SD02, STL02]. Methodological [SGB^{+}08]. Methods [AOGMS09, BST^{+}05, BOJ09, CWM08, CB01b, LG05a, Pol06, Wad00]. Método [AGBP09, AKM07, BMBM06, BNP^{+}07, BPB^{+}05, GBC^{+}04, HBKF05, JLF09, JAH^{+}05, LTC^{+}08, MP09, MW07, OIK^{+}06, PBK09, QRF07, RMSG^{+}03, TMR^{+}07, Wal04, WM05, JHHL04, STL02, BPL^{+}01]. Metodología [ZG05]. Metodológica [SGB^{+}08]. Métodos [ARMH04, AOGMS09, BPV09, BST^{+}05, BOJ09, CWM08, CB01b, DBB^{+}09b, LG05a, Pol06, SMSB05, Wad00]. Metric [LRG^{+}02]. Métrico [LRG^{+}02]. Metrics [BMMW07, FBL09a]. Metrioptera [BBK02, Ber05]. Metropolitan [HSR05]. Metropolitanos [HSR05]. Mexican [AOGMS09, GM01, Hie05, PPRB06, SS03b]. Mexicanas [GM01, Hie05]. Mexicanos [AOGMS09, SS03b]. Mexico [ACMA04, BMPNC^{+}03, BCP^{+}02, Cur02a, DLCV00, DDMMZD06, DDBA^{+}07, ESSGL00, FFHW03, GMA07, Gra01, LCS03, NHHG^{+}05, PV02, PMEH05, RCMV06, SMFPCA05, TNRJ02, ACMA04, BMPNC^{+}03, BCP^{+}02, Cur02a, DLCV00, DDMMZD06, DDBA^{+}07, ESSGL00, FFHW03, GMA07, Gra01, LCS03, PV02, PMEH05, RCMV06, SMFPCA05, TVD03, TNRJ02]. Mexico/U.S. [Cur02a]. Mezcla [KGS^{+}07]. Mice [BBM04, PLA00, RG04a]. Microhabitat [KL06, KL06]. Micromys [RG04a]. Micronesia [BBW^{+}07]. Microsatelitales [TVB^{+}06]. Microsatélite [JGL^{+}02, KGM01]. Microsatellite [JGL^{+}02, KGM01, TVB^{+}06]. Midcontinental [HRW^{+}03]. Midst [Álv05]. Midwest [RHS05]. Midwestern [Fau01]. Migración [Ber04a, IML^{+}05, MH08]. Migraciones [PB04]. Migrant [BV01a, LCS04, PSBG05, Wan04]. Migrante [Wan04]. Migrante-por [Wan04]. Migrantes [BGD09, LZD^{+}03]. Migrants [LBG03, LZD^{+}03]. Migrating [BGD09]. Migration [Ber04a, IML^{+}05, MHS07, MH08]. Migrations [PB04]. Migratorias [BNH02, BV01a, EM03, Esl00, LBG03, LCS04, PSBG05, War05]. Migratorios [KTN^{+}09]. Migratory [BNH02, Esl00, JSHM06, KTN^{+}09, Pau08, War05]. Mild [WRL02]. Mile [Ber04a]. Milenio [Nor08]. Milestone [New08]. Milla [Ber04a]. Millennium [Bel02, Nor08]. Mills [EVD01]. Mimósido [PS00a]. Mimosoid [PS00a]. Minding [You02]. Minds [JMM07]. Minería [MO04]. Mínima [FV05, WHA^{+}05]. Minimización [ŽWGW09]. Minimizar [CWM08]. Minimization [ŽWGW09]. Minimizing [CWM08]. Minimum [Ano04e, FV05, ROB^{+}04, Spe04, WHA^{+}05]. Minimum-Dispersal [WHA^{+}05]. Minimum-Impact [Spe04]. Mining [GGC03, MO04, PDLS08]. Minnesota [PMC01]. Minnow [AT05, RNL^{+}04]. Minshan [SFX^{+}08, SFX^{+}08]. Minting [EN07]. minutus [RG04a]. Mirrors [DWWF04]. Misbehavin [Woo00]. Miscellany [Haa09]. Misdirections [Asq01]. Misión [MSN06, MCR^{+}04]. Mismatch [HBLW09]. Mismatches [HPF^{+}06]. Missing [BFMM02, GCBF09, RRCOM^{+}06]. Mission [Bri06, MSN06, MCR^{+}04]. Mission-Driven [MSN06]. Missouri [GPC^{+}03, RGF^{+}04]. Misunderstanding [Tro00]. Mitigación [BCS04, CWCG02, RCMV06, SS08a, Pyk00]. Mitigar [TWW09]. Mitigate [TWW09]. Mitigating [Pyk00, RCMV06]. Mitigation [BCS04, CWCG02, McK09, Orr09a, PAB^{+}08, SS08a]. Mito [PPP05]. Mitochondrial [COW03, RLC^{+}00, Rub06b]. Mitocondrial [COW03, RLC^{+}00, Rub06b]. Mitten [HRS^{+}07]. Mixed [ASD01, DS01, PV09b, Smi05]. Mixed-Oak [ASD01]. Mixed-Species [DS01]. Mixtas [ASD01, DS01]. Mixto [PV09b]. Mizoram [Ram01]. Mobile [GM08a]. Mobilizing [Mai04]. Mode [AHR^{+}09, FH09]. Model [Awo08, BR05, BMPNC^{+}03, Cha06, CMCG07, DR03, DW06, DFS05, FLF04, FMPB07, GSG08, GBH^{+}09, HSDS01, KCC06, Lin03a, LMP02, Lus03, Min01, Moo02, NF04, PV09b, SRC^{+}02, SSSK06, SHM04, Swe08, TNTH^{+}04, TUE06, VSH^{+}07, WMVK03]. Modelación [WSQP07]. Modelado [CP00, CPD^{+}03, Cor02, HLB^{+}03, WRL02, BM05a, BM04, Cal06, Cho06, FNFM01, FBL^{+}09b, LARK01, NA06, PBK09, WDA^{+}06]. Modelados [WRW^{+}02]. Modelaje [Cua00, ET03, Old03, RWM^{+}06, Roy04]. Modelando [PAHM09, GMA07]. Modelar [Lus03]. Modeled [SHM00, WRW^{+}02]. Modeling [ARMH04, BM05a, BS01a, BM04, Cal06, CP00, Cho06, CPD^{+}03, Cor02, Cri00, Cua00, DDMMZD06, ET03, Edw04, FNFM01, FBL^{+}09b, GMA07, HLB^{+}03, KMGB07, LARK01, MPR^{+}01, NA06, Old03, PAHM09, PBK09, PR03, RAL^{+}04, RWM^{+}06, Roy04, SGSA07, WDA^{+}06, WSQP07, WRL02]. Modelo [BR05, BMPNC^{+}03, CMCG07, FLF04, FMPB07, GSG08, GBH^{+}09, HSDS01, KCC06, Lin03a, LMP02, MPR^{+}01, NF04, PV09b, SRC^{+}02, SSSK06, SHM04, TNTH^{+}04, TUE06, VSH^{+}07, WMVK03]. Modelos [ARMH04, ADF04, BNH02, BMP^{+}06, CPSS03, CJ08, Cha06, CB01b, CMS09, DFS05, DDMMZD06, DKS^{+}09, DAB06, FNF03, GBH^{+}09, GGW09, GBE^{+}06, HCK^{+}02, HSB01, KMGB07, LB05, LMG06, LHG^{+}03, LP01, MF00, NLCAR09, PD03a, PR03, RAL^{+}04, SGSA07, SHM00, SBDD05, TWDB00, Tho05b, VMV07, VO03, WHS02, WMVK03, WBV^{+}05, WWR^{+}02, BDSD^{+}05, CWCG02, CHF00, For00b, RSB05, WSBP09]. Models [ADF04, BNH02, BDSD^{+}05, BMP^{+}06, CPSS03, CJ08, CWCG02, CHF00, CB01b, CMS09, DKS^{+}09, DAB06, FNF03, Foi00, For00b, GBH^{+}09, GGW09, GBE^{+}06, HCK^{+}02, HSB01, LB05, LMG06, LHG^{+}03, LP01, MF00, NLCAR09, PD03a, RSB05, SBDD05, TWDB00, Tho05b, VMV07, VO03, WHS02, WSBP09, WMVK03, WBV^{+}05, WWR^{+}02]. Modera [DD04b]. Moderadas [WRL02]. Moderates [DD04b]. Modern [Dun05, PJW^{+}02, Ric07]. Moderna [PJW^{+}02]. Modernas [Dun05, Ric07]. Modificación [KZSV05]. Modificadas [Ant08, PP03]. Modificados [HS06].

Modifican [Cre01a]. Modification [KZSV05]. Modified [Ant08, HS06, PP03]. Modifying [Cre01a, Rud03]. Modo [AHR^{+}09, FH09]. Mohéli [GB05a]. Mohua [Cho06]. Moist [RCP00]. Molecular [CMA^{+}06, EBAW01]. Mollusc [ES04, RFD01]. Mollusca [GFTB04]. Molluscs [KDL07]. Mollusk [RFB09]. Moluscos [ES04, KDL07, RFB09, RFD01]. Monarca [BTBC^{+}03, BCP^{+}02]. Monarch [BTBC^{+}03, BCP^{+}02]. Money [HTU^{+}06]. Mongolia [IML^{+}05]. Mongolian [IML^{+}05]. Monitorear [COW03]. Monitoreo [DBB^{+}09b, FHL^{+}04, GPZ04, GBC^{+}05, HHD^{+}08, JFWP06, KMGB07, LRO^{+}08, MT08, MKH07, MML06, PBK09, RLK^{+}01, RMK07, SHJ^{+}09, SCP^{+}02, STS05, Wal09, WD01, CWC00, CLW^{+}09, LS09, SMSB05, TMR^{+}07]. Monitoreos [CMP^{+}00, Rob00]. Monitoring [Ano08l, CMP^{+}00, CWC00, CLW^{+}09, COW03, DBB^{+}09b, FHL^{+}04, FF08, GPZ04, GBC^{+}05, HHD^{+}08, JFWP06, KMGB07, LS09, LRO^{+}08, MT08, MKH07, PBK09, Rob00, RLK^{+}01, RMK07, SHJ^{+}09, She01, SCP^{+}02, STS05, SMSB05, TMR^{+}07, Wal09, WD01]. Monje [KGM01]. Monk [KGM01]. Monkey [OALM^{+}00, PS00a]. Monkeys [GJLL08, Olu00]. Mono [OALM^{+}00]. Monofilament [AYKM04]. Monofilamento [AYKM04]. monogyna [GC07]. Monos [GJLL08, Olu00]. Monster [Haa09, Joh03a]. Montaña [KM00, MSBGP03, MY02, ODE01, RIBGG08, BRBMS06, LH04, LDG^{+}00]. Montañas [GJLL08, Has00, KU02, LZD^{+}03, OFS^{+}04, SFX^{+}08]. Montane [BRBMS06, HN09, Kes01, LO06, MSBGP03, PV02, PB09]. Montano [HN09, PB09]. Montanos [LO06]. Montañoso [PV02]. Montañosos [Kes01]. Monte [TB03]. Monumental [Edw04]. Moore [WSPL09]. Moquillo [HLSZ02]. Moraine [WE07]. mordax [MSOM^{+}06]. Morfoespecies [DCD^{+}02]. Morfología [GK06]. Morfológica [Ber05, Pod01]. Morfológicamente [CTNR02]. mormo [Gun02]. Morning [WH01]. Morphological [Ber05, Pod01]. Morphologically [CTNR02]. Morphology [GK06]. Morphospecies [DCD^{+}02]. Morrena [WE07]. Mortalidad [GS02, GS05, GSB^{+}06, MSWF00, SFT00]. Mortality [GS02, GS05, GSB^{+}06, MSWF00, PMKDS01, SFT00, WLO^{+}00]. Mortandad [PMKDS01, WLO^{+}00]. Mosaic [ESSGL00, KWvH08, MJ03, PA06a, SOADMR08]. Mosaico [ESSGL00, KWvH08, MJ03, PA06a, SOADMR08]. Mosaicos [Bur04]. Mosaics [Bur04]. Most [Bec07, Joh09a]. Moth [GBC^{+}04]. Mother [Pim06]. Moths [BEB00, MKK^{+}06, MKK^{+}09, MM03a, RDEF01]. Motivación [DS08]. Motivaciones [CA06]. Motivation [DS08]. Motivations [CA06]. Motores [RS02b]. Motoristas [AJF03]. Mountain [EMW^{+}04, KM00, MB09b, RBG06, RIBGG08, Swe08, KJSH08]. Mountains [Bee09, Cor03, GJLL08, Has00, KJSH08, KU02, LZD^{+}03, OFS^{+}04, SFX^{+}08, SFS00, VBRNP08, SFS00]. Mouse [PN05]. Movement [BBK02, dLBBG^{+}08, Chi09a, FRD^{+}09, Gra01, LSL04, MMGB05, SZS01]. Movements [BBH06, DS01, NS01]. Móvil [GM08a]. Movilidad [CF01].

Movimiento [BBK02, dLBBG^{+}08, FRD^{+}09, Gra01, HL07b, LSL04, MMGB05]. Movimientos [BBH06, DS01, NS01, SZS01]. Moving [Caf00b, Fle01, HL07b, Hun07b]. mtDNA [AMF^{+}01, LCS04, TCSD00]. Much [DCT^{+}07, Sim03, Wil08a]. Mud [KC08]. Muerta [HKM06]. Muerte [Kau04, MSWF00, Rom03]. Muestra [Kai03, Wal00]. Muestreados [RTCR03]. Muestreo [GBE^{+}06, HW08, MM04, NFM04, Pol06]. Muestreos [Roy04, SCM^{+}07, UWB04]. Mulinia [RFD01]. Multi [Lam04]. Multi-Sitio [Lam04]. Multicriteria [NCVBW02]. Multidisciplinario [JAH^{+}05]. Multidisciplinarios [TRCC^{+}05]. Multidisciplinary [JAH^{+}05, TRCC^{+}05]. Multiescala [BMBM06, GBC^{+}04]. Multiespecífica [RAG03]. Multiespecífico [CA05]. Multifaceted [LMR^{+}07]. Multilocus [MBL02, Tem02]. Multiple [CTNR02, FTN^{+}05, GCF^{+}07, GCV03, GALW09, KGS^{+}07, Lam04, NP06, RA05, VTA02, WR01, YKIO08, GCF^{+}07]. Multiple-Benefit [GALW09]. Multiple-Criteria [VTA02]. Multiple-Site [Lam04]. Multiple-Species [NP06]. Multiple-Use [GCF^{+}07]. Múltiples [CTNR02, FTN^{+}05, GCV03, KGS^{+}07, NP06, NCVBW02, RA05, VTA02, WR01, YKIO08]. Multiscale [BMBM06, GBC^{+}04]. Multispecies [CA05, RAG03]. Multivariadas [SDL07]. Multivariate [SDL07]. Mulvey [BB00]. Mundial [GBC^{+}05, TK03a]. Mundo [BCMA^{+}01, HJ05, JWB06, PCC03, Rob06b]. Murciélagos [AW02, CPR05, GGGS00, KWvH08, MEA00, PMZ06, PMEH05, SK04, TRCC^{+}05, WHJJ04]. Murray [Sch10]. Murrelet [CLBC03, PBN^{+}04, Rap06]. Murrelets [MM02b]. Museo [JLZW03, PCFC01, REC^{+}00]. Museos [MCR^{+}04]. Museum [JLZW03, PCFC01, REC^{+}00]. Museums [MCR^{+}04]. Musgos [MJ03]. Mussels [BB00, BJPL07b, PHW05]. Must [Hay03, Sal05]. Must-Have [Hay03]. mustelina [DD04b, SFS00]. Mutaciones [JGP09]. Mutation [JGP09]. Mutual [CAR05]. Mutualismos [PZG^{+}05]. Mutualisms [PZG^{+}05]. Mutuos [CAR05]. Myanmar [RRK02]. mykiss [AAO^{+}07]. Mynamar [RRK02]. mysticetus [SDRO01]. Myth [Lei05, PPP05, Tas09, Wil08a].

Nabugabo [CCS^{+}03, CCS^{+}03]. Nacidos [TAP^{+}01]. Nacional [Ant08, Bei08, Cas03, CBG^{+}00, CW00, Cze05, FFHW03, HDK^{+}00, McK02, OSSK^{+}07, OBM09, PPLS01, PPW^{+}07, RPD01, Rig01, RSW05, Rua00, RIBGG08, SF03, SFS00, TSP^{+}08, TKB^{+}06, VGB^{+}04, VTA02, WGW01, ZSG01]. Nacionales [AOGMS09, DDC^{+}00, DKS^{+}09, KZSV05, MRAF^{+}07, OGH04, SGB^{+}08, GBH^{+}05, RPP^{+}00]. Nada [FO06]. Names [CML02]. Namibia [MMM03]. Namibian [MMM03]. Naped [HPK^{+}04]. Narrowing [Ber07]. National [Ant08, BT07a, Bei08, Ber03, Cas03, CBG^{+}00, CW00, Cze05, DDC^{+}00, DKS^{+}09, FFHW03, HPG09, HDK^{+}00, KZSV05, McK02, MRAF^{+}07, OGH04, OSSK^{+}07, OBM09, OS04, PPLS01, PPW^{+}07, RPD01, Rig01, RPP^{+}00, RSW05, Rua00, RIBGG08, SGB^{+}08, SF03, SFS00, TSP^{+}08, VGB^{+}04, VTA02, WGW01, WWH03, ZSG01, AOGMS09, GBH^{+}05].

Nationwide [TKB^{+}06]. Nativa [Cho06, DGSS03, SMG00]. Nativas [Bey04, BEB00, BTS09, DGC05, FTC^{+}04b, KS03, Koe03, LSF03, MNB09, NHHG^{+}05, OSSK^{+}07, PJK03, SZBR03, TS09, Tho05b, WSKH08]. Native [APSB02, ASF^{+}09, Bey04, BEB00, BIK^{+}06, BTS09, CBH01, DGC05, DGSS03, FKU05, FRD^{+}09, FTC^{+}04b, HDA01, KS03, KM00, Koe03, MNB09, NHHG^{+}05, OSSK^{+}07, PJK03, RL06, RDEF01, SZBR03, SMG00, SWB^{+}08, TS09, Tho05b, WSKH08]. Nativo [HDA01]. Nativos [APSB02, ASF^{+}09, CBH01, FKU05, GGC03, KDL07, RL06, RDEF01, SWB^{+}08, KM00]. Natterjack [BR01, SB08]. Natura [MFGB07]. Natural [ASM05, Ang00, App03, AAO^{+}07, BJM04, BB01, BEM^{+}04, DBB^{+}09b, DHZM09, Far08, FG00, FHB^{+}08, GGC03, HLG03, JSHM06, KT06, KLH08, LB00, MCR^{+}04, Mor00b, NWRD03, Nel09, NL06, OGH04, PJB^{+}06, PCVC06, PMC01, Pov01, SSSK06, SP00, Van04, VTR^{+}09, VTRG00, WM01, Whi03]. Natural-Disturbance [SSSK06]. Natural-Resource [Whi03]. Naturales [BJM04, BEM^{+}04, DBB^{+}09b, DHZM09, FG00, FHB^{+}08, GGC03, HR02, HDK^{+}00, JSHM06, Mor00b, NWRD03, Nel09, PJK03, TW06, WM01, WU05, SP00]. Naturaleza [GLMN00, SFD09]. Naturalist [Mai04]. Naturalista [Hay09a]. Naturalistic [Hay09a]. Naturalists [Bag09, Tol07]. Naturalizadas [MS08]. Naturalized [MS08]. Naturally [Bos07, EBAW01, VTR^{+}09]. Naturalmente [Bos07, EBAW01]. Nature [CKF^{+}04, DGSS06, Fie03a, HR02, HDK^{+}00, Jac08, JGM03b, LMR^{+}07, MV06, NF05a, Orr03a, PJK03, RS00, RA09, Sar07, Sch05, TW06, WU05, Caf00c]. Nauru [GK09]. Navigating [Hol01]. Near [FN04, GB01]. Nearby [Ric04b]. Nearly [Mef09c]. Necesidad [Hay09b, Joh03b]. Necesidades [Dun02]. Necesita [Sim03]. Necesitamos [Per05a]. Necessity [Fie03a, Jac08, Joh03b]. Need [DTCC03, Hay09b, HL07a, Per05a, PBC^{+}09, SAKQ^{+}05, WD09]. Needed [Sim03]. Needs [Dun02, HTU^{+}06, PPP06, Rei05b, Wil09a]. Negative [BC08]. Negativos [BC08]. Negev [SS00]. Neglected [Dun05, SL08]. Negras [MTB00, WBDP02]. Negro [DOW^{+}06, TRSR00]. Neighborhood [EAR^{+}05]. Neo [Dal00b, Tro00]. Neo-Darwinism [Dal00b, Tro00]. Neoliberalism [Büs08]. Neoliberalization [Ehr08]. Neotropical [BV01a, CVB^{+}09, Cua00, Cua02, DCP^{+}03, GRJB^{+}07, KWvH08, Lis04, LCS04, MBMLJ09, MEA00, MO04, PA08, PP03, PJW^{+}02, RAR^{+}04, RSPM05, RVVNH09, WZD^{+}00, WTEH^{+}01]. Neotropicales [BV01a, DCP^{+}03, LCS04, Mcs05, MEA00, PA08, PP03, WTEH^{+}01]. Neotrópico [CPR05, NRB00]. Neotropics [CPR05, EBDM06, Mcs05, NRB00]. Nest [BB04, BEM^{+}04, CTR02, DWN01, Faa04, HRW^{+}03, HAJ^{+}04, PLB09, Sch03b, TB04, Wel06, WTEH^{+}01]. Nest-Site [PLB09]. Nestedness [FN02a, KSP00]. Nesting [BMM^{+}01, DD04b, FCG^{+}05, GPC^{+}03, GNEE07, KSDBT02, Koe03, KT02, MTB00, MM02b, PV09a, Rem03, RR01]. Nests [BM05a, BEM^{+}04, TB04, VP04, WMS^{+}01]. Nestwatch [EAR^{+}05]. Net [Dav07]. Netherlands [KBS^{+}04, LDG^{+}00, Wol00]. Network [BBBM02, BVL^{+}03, HCZ00, MFGB07, Nos03, RWJRR06, TKB^{+}06, UKC^{+}09]. Networks [BFMR07, JSWS00, JFS08, Kri04, SJF^{+}05]. Neuter [LRS09a]. Neutrality [Ano02e]. Newly [ABE01]. News [Ghi06, SMI^{+}07]. Newts [DDK05, JMLG01]. Next [Ada08, Nug08]. Nexus [Har07]. NGOs [Da 03, Swa09]. Niagara [ML02, MNL03]. Nicaragua [Gil01b, SMT07]. Niche [DDMMZD06, GBE^{+}06, PR03]. Niche-Based [GBE^{+}06]. Nicho [DDMMZD06, PR03]. Nichos [GBE^{+}06]. Nidación [BMM^{+}01, MTB00, PLB09, RR01]. Nido [GNEE07]. Nidos [BB04, BM05a, BEM^{+}04, CTR02, DWN01, HRW^{+}03, HAJ^{+}04, PV09a, Sch03b, TB04, WMS^{+}01, WTEH^{+}01]. Niebla [MK09, PMEH05]. Nieve [CFH^{+}02]. Nile [KDG^{+}06, Mal03]. Nilo [KDG^{+}06]. Nine [Tut03]. Niño [FJS04, WLO^{+}00]. Nivel [AHR^{+}09, BB06, FCG^{+}05, GDAV00, Gui07, HLB^{+}03, HDE00, MMF^{+}05, Rua00]. Niveles [BM05b, CVS^{+}09, CY08, GAF00, PZG^{+}05, TG04]. No [BJB^{+}02, BPB^{+}05, DGC05, DWW04, FDL^{+}05, FHL^{+}04, FPW01, FJ07, GK08, Gos07, Gru04, KDL07, KM00, KCR^{+}08, LB05, LSF03, MNB09, MM00, MBQ05, MS09, NHHG^{+}05, PPLS01, PV06, RFB09, RWRW04, TNRJ02, WBW05, WZW^{+}06, BEB00, Fau01, HDA01, Olu00, PL03, RTCR03, STC03, RFB09]. No-nativos [KM00]. No-Take [FDL^{+}05, LB05]. Noah [HRP^{+}04]. Nocivas [DDAK04]. Nocivos [Cou02]. Noctuid [MKK^{+}09]. Nocturnal [LCG^{+}08, WHJJ04]. Nocturnas [MKK^{+}06, MKK^{+}09, MM03a]. Nocturnos [LCG^{+}08, WHJJ04]. nodiflorus [KFA^{+}03]. Noise [BHB08]. Nomads [LWL09]. Nombres [CML02]. Non [BIK^{+}06, DGC05, HDA01, KM00, MNB09]. Non-Native [BIK^{+}06, DGC05, HDA01, KM00, MNB09]. Nonacademic [MS09]. Noncompliance [Kri04]. Nondestructive [KB03]. Nonindigenous [BRR^{+}02, KL05, KDL07, Lei00, LSF03]. Noninvasive [Cla09, FHL^{+}04, FPW01, PV06]. Nonlethal [BPB^{+}05, STC03]. Nonlinearity [Ova03]. Nonorganic [BJB^{+}02]. Nonrandom [MBQ05]. Nontarget [BEB00]. Nontimber [FJ07, PPLS01, PL03, TNRJ02]. Nontraditional [Nad04]. Nonviolence [Ano08a]. Nonviolent [Kum08, Mer08]. Noreste [BD06, LCTL05, NMB09, RW02]. Normas [JAH08]. Normativa [VNP06]. Normative [CCM00a, Hun00, VNP06, Wil00a]. Normativity [NVP07, WABT07b]. Normativos [WL04]. Norms [JAH08]. Noroccidentales [Bur04]. Noroeste [BPB^{+}05, BVL^{+}03, Cha06, aARLG03, MML06, NB06, Pow06, Rap06, RWBG06, TFGJ06, Dav07, LW02b, SDJ06]. Nororiental [Ram01]. Norte [BJ05, GFCP03, HR04a, HM04, HL07b, HFR07, KU02, LSN07, MB00, NWRD03, PMBN01, PCME09, RSJE06, REC^{+}00, TVB^{+}06, Urb06b, LCC^{+}07, BHMC01, Bor08, BIK^{+}06, Car07, EM03, GK06, HL07b, Hut06, JSJ05, MMB^{+}03, NMB09, OMP^{+}05, Ros08, SRW^{+}08, SBVC03, VJ07, VMM07]. Norte-Centro [JSJ05]. Norteamérica [BMMW07, MSOM^{+}06, WD01]. North [BMMW07, EM03, GK06, GNR01, HL07b, KU02, MMB^{+}03, NMB09, Orr03c, RR01, Ros08, BHMC01, Fie00b, GFCP03, GCBF09, HL07b, Hut06, JSJ05, MSOM^{+}06, NFD^{+}09, OMP^{+}05, REC^{+}00, Rot01, SRW^{+}08, SBVC03, VJ07, VMM07, VS05b, WD01]. North-Central [JSJ05]. Northeast [Ram01]. Northeastern [BD06, CT01, Ibi08, LCTL05, NMB09, RW02]. Northern [CJ08, GSR^{+}04, GJD05, HODB09, HFR07, NB06, Vic02, BJ05, Bor08, BIK^{+}06, Car07, HR04a, HM04, LSN07, LCC^{+}07, MB00, NWRD03, PMBN01, PCME09, RSJE06, TVB^{+}06, Urb06b]. Northwest [Cha06, DW06, Hoo07, aARLG03, MML06, NB06, Pow06, Rap06, RWBG06, TFGJ06, Dav07, SDJ06]. Northwestern [Bur04, BPB^{+}05, BVL^{+}03, LW02b]. Noruega [RSJE06]. Noruego [Rua00]. norvegicus [APCS05]. Norway [APCS05, RSJE06]. Norwegian [Rua00]. Nosed [GJLL08]. Noss [HRN00]. Note [Ano05e, Ano05f, BG04, EC01]. Nothing [FO06]. Novel [Lin03a, RHA^{+}05]. November [Ano04k, Ano05j, Ano07o, Ano07r, Ano08q]. Nubia [SS00]. Nubian [SS00]. Nublado [CFC02]. Nuca [HPK^{+}04]. Nuclear [COW03]. Núcleo [CMCG07, CS07b, RPD01, SS00]. Nueva [CMM05a, GB01, Lin03a, MCF^{+}09, NKBD02, Wal09, AE02, BCHS08, Cho06, CAFL03, DA07, FO06, HBA09, JRL03, JWB06, LB05, LOM03, LLW02, RNP^{+}07, SSB^{+}08, SB04, TPW03, Tow03]. Nuevas [Joh03b, Pyk04, RDAL09, BMRB04]. Nuevo [BNP^{+}07, FFHW03, GLMN00, Kot00, LTC^{+}08, MC05a]. Nuevos [JLZW03, WL04]. Nula [CT07, FBC^{+}06]. Null [CT07, FBC^{+}06]. Null-Hypothesis [CT07, FBC^{+}06]. Number [AHR^{+}09, BM05a, GCM^{+}08, MC05b, SL08]. Numbers [DA02, MLMB08, MS04, NA01]. Numerical [BBA^{+}02]. Número [AHR^{+}09, BM05a, MS04, MC05b, SL08]. Números [MLMB08, NA01]. Numerosas [RFB09]. Numerous [RFB09]. Nursery [Du 01]. Nutrias [FVS03]. Nutrient [WCP01]. Nutrientes [WCP01].

o [BFMR07, GMMPP08, GLMN00, HW08, MCR^{+}04, RAW07, WABT07a]. Oak [ASD01, Bra06, FB08, FMLL06, GMA07, GBC^{+}04, MBM04, WE07, WE07]. Oaks [BRA08, PAHM04]. obesulus [SLCW08]. Objectives [JML00, NP06, ZSVL07]. Objetivos [GV05, JML00, JFS08, LSL^{+}06, NC03a, WSS^{+}04, WG08, ZSVL07, NP06]. Obligada [CMCH00]. Obligate [CMCH00]. Observación [MF00, WMPC08]. Observation [BNM^{+}09, MF00]. Obstacles [BPP05, Cam05, EK08, Kra01]. Obstáculos [EK08]. Obstruction [RVNM08]. Obtained [PBSR06, SGSA07]. Obtenida [SGSA07]. Obtenidos [PBSR06]. Ocasiona [CW02]. Ocasionada [GFTB04]. Occidental [BJB06, CWNW01, DMDFV08, GM00b, HR04c, MMB^{+}03, OALM^{+}00, PPLS01, RCP00, BVK02, Bra03a, Cra02, TRCC^{+}05]. Occidentales [BKWB05, Kau04, KCS04, OFS^{+}04]. occidentalis [HMF^{+}04, HODB09, NB06, TG04, CJ08, PR03, TG04]. Occidente [Fau01, LCS03]. Occupancy [GM09, JMLG01, LMP00, PBK09, RB06]. Occupied [Vee06]. Occur [Bre02b]. Occurrence [BFD07, Edw04, FNFM01, FNF03, MM03a, NFB05, PSBG05, VCHBP08].

Occurrences [MC05b]. Ocean [AFK^{+}09, SW07, BPB07, CRGG01, LHPH06, TRCC^{+}05, WELCH09]. Oceania [KWL^{+}09, KWL^{+}09]. Oceanic [BCD^{+}08, TSKD07, TSKD08]. Oceánicas [TSKD07]. Océano [LHPH06, TRCC^{+}05, WELCH09]. Océanos [BCMA^{+}01, GBH^{+}05]. Oceans [BCMA^{+}01, GBH^{+}05]. Ocho [GV02]. ochrocephala [Cho06]. October [Ano09k, Ano05j, Ano06j, Ano09m]. Ocupación [GM09, JMLG01, PBK09, RB06, LMP00]. Ocupados [Vee06]. Ocurrencia [BFD07, FNF03, NFB05, PSBG05, VCHBP08]. Ocurrencias [MC05b]. Odocoileus [PLL^{+}05]. Oeste [DDC^{+}00, HVR08, RWS02, VDHD01, BRK^{+}04, Kee06, LS01, RHS05]. Off [BC08, Ehr03, HBA09]. Off-Site [BC08]. Offs [FRD^{+}09, WELCH09]. Offsets [WD09]. Offspring [EKL^{+}00, FNP09, FPL^{+}04]. Ohio [ASD01]. Oigenes [TFGJ06]. Oil [BW08b, BW09, Har07, PFR^{+}09, TB03]. Oil-Palm [BW08b, BW09, PFR^{+}09]. Old [BRA08, Cur03, MML06, Moo07, RS02c, STL02, SHYH06, SDJ06, WSLW00]. Old-Field [RS02c]. Old-Forest [STL02]. Old-Growth [MML06, Moo07, SHYH06, SDJ06]. Olfactory [RG04a]. Olfativa [RG04a]. Oligosoma [TPW03]. Olympic [GVTM07]. olympus [GVTM07]. Once [PN05]. Once-Common [PN05]. Oncorhynchus [ALH^{+}04, AAO^{+}07, BPB07, NR03, ZSMW06]. One [Caf04, Foi00, LFA^{+}07, NBM^{+}04, Orr00c, Orr07b, SPM05, Spe07, SAA^{+}09, Tut03, Wan04]. One-Migrant-per-Generation [Wan04]. Online [Edi03]. Only [Fre06]. Ontario [ML02, MNL03, WE07, ML02, MNL03, ZSG01]. opacum [GGW09, TSG06]. Opciones [BMS08, CV08, PPBW08]. Open [BVH^{+}02, GNEE07, HSR05, Mer04, PV09a]. Open-Nesting [GNEE07, PV09a]. Open-Space [BVH^{+}02, HSR05]. Operacional [KCC06]. Operacionales [BFMR07, FP02]. Operation [CMR05]. Operational [BFMR07, FP02, KCC06]. Opinion [SBDD05, SBDD05]. Oportunidad [CWM08, NA06]. Oportunidades [DCP^{+}03, NSA06, PTP^{+}08, RDAL09, TPS^{+}05]. Oposición [Bei08]. Opportunism [KC07, PB08]. Opportunistic [Smi05]. Opportunities [BDRC05, DCP^{+}03, NSA06, PTP^{+}08, RDAL09, TPS^{+}05]. Opportunity [CWM08, GB05b, KC08, NA06, SS02b]. Opposing [Cor06]. Opposition [Bei08]. Óptima [JMP09, SDL07, TTSP04]. Optimal [CPD^{+}03, HHD^{+}08, JMP09, MP08, SDL07, TTSP04, ZLM03]. Óptimas [CPD^{+}03]. Optimism [Ano07m, Bee00, Fie03c, Kni07, Orr07c]. Optimistic [Bro00]. Optimización [HCK^{+}02, MDH^{+}07]. Optimizing [HCK^{+}02, MDH^{+}07]. Óptimo [HHD^{+}08, MP08]. Óptimos [ZLM03]. Options [CV08, PPBW08, YBWZ06]. Opuestas [RHA^{+}05]. Orchard [SD03]. Orchestrating [Rig01]. Orchid [CNC04, DKS^{+}09, GG06, PCVC06, PC07]. Orchids [SMFPCA05]. Orconectes [PFL01]. Oregon [RWS02, FB08, MM02b, PJ00, PJYO04, RWS02, VB06, WSLW00]. Organic [BJB^{+}02, Gun02, JSJV03, WHJJ04]. Orgánicas [BJB^{+}02, WHJJ04].

Orgánico [Gun02]. Orgánicos [JSJV03]. Organismos [HS06]. Organisms [HS06]. Organización [CMR05]. Organizaciones [Cle06b]. Organizations [Cle06b, RA04b]. Orientación [Cla01, The03]. Oriental [CGR^{+}05, GV02, KPH07, LF00, MRM^{+}05, RLC^{+}00, SH09, Sta01, CCS^{+}03, ES04, EM03, GCF^{+}07]. Orientales [GK06, SBVC03]. Oriented [Cla01]. Origen [KSB01, MBL02, RHA^{+}05, Ols07]. Origin [KSB01, MBL02, Ols07, RHA^{+}05]. Origins [TFGJ06, Van04]. Orlog [BR05]. Ornamental [DSTPW07]. Ornamentales [DSTPW07]. Oro [MO04]. Orphaned [TAP^{+}01]. Orquídea [CNC04]. Orquídeas [DKS^{+}09, PCVC06, PC07, SMFPCA05]. Orr [Ano05c, Ehr01, Joh05b, MA05, Orr04c]. Osmerus [MSOM^{+}06]. Oso [DN01, DOW^{+}06, NWRD03]. Osos [BST^{+}05, MM02a]. Ostras [Rom03]. Other [DDAK04, FM09, Joh05a, MHM09, OBM09, Orr04d, Pan09, PDLS08, UMT^{+}01]. Otras [DDAK04, OBM09, UMT^{+}01]. Otros [FM09, MHM09]. Otters [FVS03]. Otto [SF00]. Our [Ano08a, Fie03a, Har03, KORB02, NL03a, Tas09, NL03b]. Outboard [RS02b]. Outboard-Powered [RS02b]. Outbreeding [ET03, FG00, GGI^{+}05]. Outcome [Rue05]. Outcomes [BMC^{+}09, Bei08, EAR^{+}05, MM04]. Outcrops [GN00]. Output [GMW08]. Outreach [Bre02b, JMM07, WMCB02]. Outsiders [Foo04]. Ovadia [MM03b]. Ovenbird [FTR01]. Over-the-Horizon [CW08]. Overabundant [ASM05]. Overall [JMS06b]. Overcoming [BPP05, Cam05, Kra01]. Overestimating [JSW08]. Overestimation [HODB09]. Overexploitation [RM02]. Overfished [UWB04]. Overharvested [NL05]. Overlap [LSN07]. Overlooked [ESB05]. Overpasses [CHFR09]. Overview [Ano01j, DH00b, MCM^{+}00]. Overwintering [BCP^{+}02, LF09, LCS04]. Ovina [Ber06]. Owl [CJ08, Dav04, HODB09, NB06, PR03]. Owls [HMF^{+}04, SS03b, TG04]. Owner [Car06]. Ownership [BMRB04, Bro08a]. Ownerships [TAJ04]. Oyster [Rom03]. Ozark [GPC^{+}03, RGF^{+}04]. Ozymandian [Rai00].

Pêsca [FMP^{+}05]. Pachira [FLQ03]. Pacific [WELCH09, CE00, DDAK04, GWM^{+}07, Hoo07, LHPH06, PSGA09, RC04, SDJ06, TPT^{+}07]. Pacífico [BMP07, DDAK04, GWM^{+}07, LHPH06, PSGA09, SDJ06, TPT^{+}07, WELCH09]. Pacú [RW02]. Paedomorphic [DDK05]. Pagan [FN04]. Pago [SAPRB07, RRH^{+}09]. Pagophilus [JML00]. Pagos [DJOW07, Fer01, JLF09, NTGT03, Wun07]. Painting [Mul07]. País [Rod05, VIMMO05]. Paisaje [APSB02, ARMH04, BNH02, Ber05, BFD07, BKWB05, BJB06, BV01a, BBH06, CPSS03, Car07, CPD^{+}03, CPO03, DM04b, DD04b, FJ00, FPSD09, FLF04, FTR01, FCL00, GB03, GM00a, GM00b, Gra01, GBC^{+}04, GGW09, GH02, HR04a, HLB^{+}03, HCZ00, HDDEB03, JMH01, KSO^{+}03, KFNB09, LMS^{+}02, LB02, ML04, MU08, NS01, PDF^{+}00, PA04, PGP04, PV02, PDS^{+}07, PMC01, PBL00, RWM^{+}06, RDEF01, RAR^{+}04, RS02c, RK05, SLB^{+}07, SFX^{+}08, SP00, SD03, SB08, SBSL06, VVMV09, VDP00, WM01, WCPB03, WMD^{+}04, WMS^{+}01, Wit02, YL05]. Paisajes [BMPNC^{+}03, BTS09, CNPS04, DCP^{+}03, DWW04, FEKH09, GLF^{+}08, HMB^{+}05, HSB01, JTS06, KBS^{+}04, LLS08, LZD^{+}03, MM02b, MPE^{+}07, NA06, PJYO04, PD03a, PAA^{+}08, RL06, RMN00, RBG^{+}04, SHYH06, TNFH07, Vee06, WFP01, WBV^{+}05]. Paisajísticos [Bur04]. Países [LDG^{+}00, Wol00]. Pájaros [Lam04, MGC01]. Palangres [LC07a]. Palm [BW08b, BW09, EGPS04, PFR^{+}09]. Palma [EGPS04]. Palomas [EM03]. Palung [PPLS01]. palustris [Bos07]. Panama [RCP00]. Panamanian [GGC03]. Panameños [GGC03]. Panda [LJMR^{+}01, WLS^{+}08]. Pandas [LZD^{+}03, SFX^{+}08]. Pantanal [HTM^{+}05, BF03]. Pantano [CTNR02, FTC^{+}04b]. Pantanosa [ND03]. Pantheon [Bye04]. Papas [HGH^{+}00]. Papel [ASF^{+}09, BRA08, CM06, Far08, Fer01, FRSG^{+}02, FCL00, KEJ06, KFA^{+}03, MDF05, MAM^{+}03, MCM07, OWOF03, RNL^{+}04, SZS01, SSL06, War05, DMDFV08, HBA09, TR01, WSA^{+}05]. Paper [Bea07]. Papers [LK04]. Papua [CMM05a]. Para [SMRR02, YLR^{+}01, AGBP09, AMF^{+}01, AW02, Ang00, Ant08, AP00, Arm05, AKM07, ARPEBM09, AOGMS09, ASD01, BV04, BPV09, BTM^{+}05, BBBM02, BMBM06, BCS04, Bau05, BMRB04, BR01, BMS08, BST^{+}05, BKK03, BM04, Bey04, BTK08, BNP^{+}07, BGD09, BSM08, BR05, BPB^{+}05, BZD03, BMM^{+}02, BFH^{+}06, BWS09, BVL^{+}03, BV01b, BPL^{+}01, BRR^{+}02, CWM08, CHW^{+}01, Cas03, CSS00, CVB^{+}09, CMC03, Cha06, Che09, CS05, CDD^{+}05, CT07, CNC04, CPD^{+}03, CSB07, CDM09, CMBC00, Cla01, CMK^{+}07, CW00, CWCG02, CA06, CMCG07, COW03, CA05, CHFR09, CLŽ^{+}07, Cra02, CS07b, CFH^{+}02, CGR^{+}05, CB01b, CV08, CMS09, Cze02b, Cze05, Cze08, DDAK04, Dal05, DBB^{+}09a, DSJ02, DDC^{+}00, Dav05, DN01, DK06, DW05a, DD01, DC05, DA07, DOW^{+}06, DWW04, DM04b, DP01, DJOW07, DAB06, Dun02, EAA^{+}09, Ehr06]. para [EG07, EM08, EGPS04, EK08, FKU05, FVS03, Fat06, FP02, FTC04a, Fer01, FMT^{+}04, FBL09a, FMPB07, FPL^{+}04, Fje07, FTN^{+}05, Fon09, FBP^{+}09, FBK02, GRJB^{+}07, GGA01, Gav04, GB03, GWS07, GCV03, GB09, GWS^{+}06, GLMN00, GN00, GM01, GBC^{+}04, GBH^{+}05, GB05a, GBH^{+}09, GM08a, Gre05, GBC^{+}05, GGW09, GBE00, GGT08, GV05, Gui07, GBE^{+}06, GS08, HHD^{+}08, HBC^{+}06, HM08, HMF^{+}01, HSDS01, HBD^{+}09, HRS08, Hay09b, HP01, HBBD08, HLG03, HLB^{+}03, HBKF05, HC09, Hod04a, HH06b, HY03, HW02, Hun05, Hut06, HN09, IF02, JLF09, JMS^{+}06a, Jan07, JGM03a, JSW08, JB06, JHHL04, JSHM06, JAH^{+}05, JAH08, JFWP06, KSC^{+}06, KHM00, KFSR^{+}06, KMGB07, KC01, KB04a, KZSV05, KDL07, KWvH08, KTN^{+}09, KSO^{+}03, KCC06, KCR^{+}08, Kom03, Kri04, LG02, LLS00, LBR07, LS09, Lat00, LF09]. para [LWSM03, LTC^{+}08, LOM03, LRO^{+}08, LP08, LW02a, LCZ^{+}05, LMB00, LARK01, LVBS^{+}04, LCC^{+}07, LMP02, LRG^{+}02, LGK02, LP01, LZD^{+}03, LO02, LV04, LG05b, Lus03, MB07, MCG^{+}08, MLN^{+}06, MDF05, MDCK07, MNRD^{+}08, MMGB05, MDH^{+}07, MKH07, MBQAM06, MOO09, McK02, MWBP07, MP07a, MP08, MCG^{+}09, MCK06, MS04, Mcs05, MD04, MN03, MCM^{+}00, MHGF04, MSB^{+}00, MM02b, MU08, MC05a, MWEB06, MW07, MBC03, MBB^{+}03, MH01b, MS09, MSBOF^{+}09, MMB^{+}03, NLCAR09, NSA06, NP06, NL05, NFB05, ND03, Nos03, NFD^{+}09, NR03, ØDEA00, Old03, OFN^{+}06, OIK^{+}06, PBSR06, PCL^{+}03, PRL03, PA06a, PBP09, PC02, Ped03, PSF05, PMC^{+}07, PA08, PBK09, PPBW08, PD03a, PCC03, PR03, PCL08, PP03, PLB09, PMKDS01, PCVC06, PSGA09, PJW^{+}02, PIH00, PEC01, PAHM04, PPW09, PS06, Pyk05, PTP^{+}08]. para [QAMW03, RBT08, Ram01, RL06, RRK02, RAL^{+}04, RMN00, RWM^{+}06, RR00, RDW01, RLC^{+}07, Rob00, RG04a, RH01, RS02b, Rod05, Rod03, RBP01, RBHKK02, RAG03, Ros08, REC^{+}00, RBG^{+}04, RCL^{+}06, Rua00, Rub01, RLXL01, SCD^{+}08, SCB^{+}01, SSS^{+}08, SRC^{+}02, SMM09, SC00, SJF^{+}05, SMV02, SGSA07, SGG07, SWK05, SPM05, SH03, SC05, SS08a, Sem02, SB03, SSB^{+}08, SFX^{+}08, SMG00, STC03, SCP^{+}02, Sie04, SWD00, Sim03, SFS00, SMH^{+}07, SBB^{+}08, SWR00, SDL07, SKBB00, SEBM03, SLCW08, Spe02, SS06, SD03, SHT^{+}05, SCS06, SB08, SD01, SDJK04, SAA^{+}09, TPS^{+}05, TG05, TVB^{+}06, TWDB00, TG04, TTSP04, TWGH00, TTB03, Tho05b, TS06, TK07, TWW09, TRCC^{+}05, TW06, Tur06, TKB^{+}06, TAP^{+}01, TUE06, UCAGG08, UKC^{+}09, UMT^{+}01, Van06, VSH^{+}07, VT09, VMV07]. para [VO03, VTR^{+}09, Wal09, Wal04, WGW01, Wal00, WM01, WABT07a, WFP01, WDA^{+}06, WM05, Wen08, WS03, WE07, WMD^{+}04, WL04, WHA^{+}05, WD03, Wil08b, WSBP09, WMVK03, WBV^{+}05, WWR^{+}02, WG08, YKIO08, YL05, You05, ZR00, ZG05, ZSVL07, ZJ04, ZLM03]. Paradigm [MA05, PA06a, Pas00, PBN^{+}04, VP07]. Paradigma [PA06a, PBN^{+}04]. Paradigmas [Arm05]. Paradigms [Arm05, BSR04]. Paradise [McB00, Whi06]. Paradise-Flycatcher [Whi06]. Paradoja [KIAH08]. Paradox [DGSS06, GL03, KIAH08, WA00]. Paraguas [RA04a]. Paraguay [AW02, HMF03]. Paraguayan [AW02]. Parameters [MCG^{+}09]. Parámetros [MCG^{+}09, RM06]. Parasitaria [CGR01, GC06]. Parasite [CGR01, GC06]. Parasites [Log03, Sal07]. Parasitic [JLZW03]. Parasítica [JLZW03]. Parasitism [GM00b, LAN05, VSH^{+}07]. Parasitismo [GM00b, LAN05, VSH^{+}07]. Parasitoid [FBP^{+}09]. Parasitoides [FBP^{+}09]. Parásitos [Log03]. Paratya [HGH^{+}03]. Parcelas [NFB05]. Parche [SZS01]. Parches [ARPEBM09, FE07, FRSG^{+}02, FCL00, FES01, Had00, JSWS00, LMP00, PA06a, WH01]. Parcial [VMM07]. Parciales [MMGB05]. Parda [WBBA03]. Pardo [NWRD03]. Pardos [BST^{+}05, MM02a]. Pareja [GPMM08, RG04a]. Pares [BADF^{+}05]. Park [Ant08, Ber03, CBG^{+}00, CW00, FFHW03, NTMT^{+}03, OSSK^{+}07, OBM09, PPLS01, PCVC06, PIH00, RIBGG08, Ste08a, TSP^{+}08, VGB^{+}04, WGW01, Cas03, PPW^{+}07, RPD01, RSW05, SFS00]. Parkinsoniana [MBC03]. Parkinsonism [MBC03]. Parks [And05, Bro04, DDC^{+}00, Fue06, Get05, Gro05a, MS02, NSB^{+}06, OGH04, OS04, PZ01, RRA06, RG04b, SNM00, SMN00, Ter04, WMD^{+}06, RPP^{+}00, ZSG01]. Parnassia [Bos07]. Parque [NTMT^{+}03, PCVC06, PIH00, TSP^{+}08, VGB^{+}04, Ant08, Cas03, CBG^{+}00, CW00, FFHW03, OSSK^{+}07, OBM09, PPLS01, PPW^{+}07, RPD01, RSW05, RIBGG08, SF03, SFS00, WGW01, ZSG01]. Parques [NSB^{+}06, OGH04, SMN00, DDC^{+}00, RPP^{+}00]. Parrots [WTEH^{+}01]. Part [GS08, Orr00c, Orr00b, Tut03, Mor00a]. Parte [TM04]. Partes [GS08, SCP^{+}02]. Partial [MMGB05, VMM07]. Participación [DBB^{+}09b]. Participant [EAR^{+}05]. Participation [DBB^{+}09b]. Participativa [TWW09]. Participativo [JHHL04]. Participatory [JHHL04, TWW09]. partir [Roy04]. Partitioning [GCV03]. Partnering [Put03]. Partnership [Bre02b]. Partnerships [Bre02a, Far04, Sey04]. Parts [MMK04, SCP^{+}02, Tut03]. Parus [WGGS07]. Parvadas [DS01, EM03]. Pasada [VTR^{+}09]. Pasadizos [CW00]. Pasajes [CWCG02]. Paserina [Cho06, CW02]. Paserinas [Cra02]. Paso [PA08, Rod03]. Pasos [CHFR09]. Passage [PA08]. Passages [CWCG02]. Passenger [EM03]. Passerina [Wel06]. Passerine [Cho06, CW02, Cra02]. Passerines [Sal07]. Passion [CKF^{+}04, Mul09]. Past [VTR^{+}09]. Pastiche [Ste06b]. Pastizal [BNH02, CPO03, CCH03, HIS03, JC07, LSC07, MBM04, Mor00b, RJW09, SM09, Vic02]. Pastizales [BVH^{+}02, CGR^{+}05, DBL06, GGGS00, HH03, LH04, MMLM01, MNH01, ND04, SZBR03, SP00, SWB^{+}08, VMBC02, VSGS02, CW08]. Pasto [MHJ^{+}09, VMBC02]. Pastoralists [CW08]. Pastoreo [Bor05, FFHW03, FEKH09, GM00b, HIS03, HH03, Hob01, KHRS01, KS03, KT02, LSC07, Mad05, MM07, SSM03, SK02b]. Pastores [CW08]. Pastos [BE00, RLK^{+}01]. Pastura [KWvH08]. Pasture [KWvH08, NH09]. Pasture-Remnant [NH09]. Pastures [GGGS00, SZBR03, SP00]. Pat [Gos07]. Patagonia [BPB07, BPB07]. Patagonian [GMCP09]. Patas [KM00]. Patch [FRSG^{+}02, FES01, Had00, LMP00, MK09, PA06a, PBK09, SC05, SZS01, WH01]. Patch-Occupancy [PBK09]. Patches [ARPEBM09, DMS03, FE07, FCL00, JSWS00, MJ03, PMC01, YKIO08]. Patchiness [WD06, WHH00]. Patchy [WSBP09]. Pathogen [FND06, KBM01, KDG^{+}06, KFK^{+}09, PC08, RHA^{+}05]. Pathogenic [BRS^{+}05]. Pathway [Had07]. Patient [VJ09]. Patógeno [BRS^{+}05, KBM01, RHA^{+}05, KFK^{+}09]. Patógenos [FND06, KDG^{+}06, PC08]. Patrimonio [LSN07]. Patrón [HR04a, ZGR02]. Patronatos [MHGF04]. Patrones [BS00, BL06, Fje07, FLQ03, Gra01, HDQSL06, LG05a, Lai05, LSL04, LRG^{+}02, MB01, NFM04, Rem03, RCL^{+}06, SRR00, SKBY03, SHT^{+}05, WBBA03, BVH^{+}02, BEM^{+}04, ES08, FSC02, GRH05, LJMR^{+}01, OFS^{+}04, SLH00, WMS^{+}01]. Pattern [HR04a, RCL^{+}06, SKBY03, ZGR02]. Patterns [BVH^{+}02, BS00, BL06, BEM^{+}04, ES08, FSC02, Fje07, FLQ03, GRH05, Gra01, HDQSL06, LG05a, Lai05, LSL04, LRG^{+}02, LJMR^{+}01, MB01, NFM04, OFS^{+}04, PA04, Rem03, SRR00, Sta09, SHT^{+}05, SLH00, WBBA03, WMS^{+}01]. Pay [FN04]. Paying [NTGT03, RRH^{+}09]. Payment [RA09, SAPRB07]. Payments [DJOW07, Fer01, JLF09, Wun07, ZHM08]. Paz [Ehr00]. Peace [Ehr00]. Pearl [Rom03]. Peat [ND03]. Peces [AAO^{+}07, BSKH^{+}06, CA05, CV08, DDK05, DDMMZD06, DN00, FKU05, FMT^{+}04, Fra09, GWJ^{+}07, GFTB04, KTN^{+}09, KBM01, KM00, LW02a, LHPH06, LM07, LARK01, MHJ^{+}09, Mcc09, MO04, PCL^{+}03, PA08, Pop08, Rue05, SCD^{+}08, TH03, UMT^{+}01, WBSG07]. Pecos [RNL^{+}04]. pecosensis [RNL^{+}04]. Pedomórficos [DDK05]. Pedro [KBR03]. Peer [BADF^{+}05, HP01]. Peer-Reviewed [BADF^{+}05]. Pelagic [Rob09, SCP^{+}02]. Pelágica [Rob09]. Pelágico [SCP^{+}02]. Peligro [AS01, ALH^{+}04, AMF^{+}01, BB06, BTGS05, BL06, BV00, BRA08, CGR01, CHBK02, CMGSM00, Cze05, DBB^{+}09a, DK06, DDBA^{+}07, DP01, ERB01, Fal07, FSS^{+}00, FH09, Gar03a, GS01, Gre05, HBC^{+}06, HH06b, JWB06, KSP00, KSO^{+}03, KFA^{+}03, KSB01, KDD02, LTC^{+}08, LVM02, Log03, MKH07, MBQAM06, Mor00b, PZG^{+}05, RM01, RW03, RLXL01, Say05, SH03, SDRO01, SSL06, SB08, TS06, TRCC^{+}05, UKC^{+}09, VT09, VNP06, Wal04, WABT07a, Wen08, WL04, WMPC08, BZD03, HP01]. Peligros [SMN00]. Penalty [MW07]. Penguin [WBW05]. Penguins [WBW06]. Península [FMT^{+}04, FMT^{+}04]. Peninsular [ODE01, Lai00]. People [CW08, Dor04, GGC03, KK08, MH02, RG04b, SNM00, WMD^{+}06]. Peoples [Bro04, Col00, SZ05]. Pequeña [DLCV00, HJ05, MBR^{+}02, RSJE06]. Pequeñas [BCS00, FTC04a, FNP09, KFBG08, Log03, Pan02, VSAG06]. Pequeño [BV04, MO04, RLC^{+}00]. Pequeños [ARPEBM09, CKL^{+}00, MR00a, OK01a]. Perceived [FCN^{+}06]. Percepción [BGD09, CHW^{+}01, MNH01, PMKDS01]. Percepciones [Kau04, MMM03, Mcs05, BF03]. Perceptions [And01c, BF03, Kau04, MMM03]. Perdedores [WCX09]. Perdida [GWM^{+}07, BV00, CM03, CVS^{+}09, CW02, DBL06, DAR08, ENN^{+}08, Fai08, GRH05, HRS08, HC09, HPG09, LM07, Mcc09, RLK^{+}01, Ros06, Ros08, SAA^{+}01, SB04, WRM05, BJTD03, BMM^{+}02, FMSL08, FJS04, Gib00, Kel01, NNM00, PLL^{+}05, PAA^{+}08]. Pérdidas [BCS04, ESB05, FKU05]. Peregrine [TR01]. Peregrinos [TR01]. Perenne [BJE^{+}05, LM05, LE02, LDG^{+}00, SM09, VMBC02, Vic02]. Perennes [Coc01]. Perennial [BJE^{+}05, LM05, LE02, LDG^{+}00, SM09, VMBC02, Vic02]. Perennifolia [GGGS00]. Perfect [Pim06]. Performance [BSS^{+}05, FVS03, Fer01, LBR07, MBB^{+}03, WWR^{+}02, ZHM08]. Periféricas [YE07]. Peril [PZ01]. Perils [PZ01, SMN00]. Periodically [BMC^{+}09]. Periódicamente [BMC^{+}09]. Peripheral [GVS09, YE07]. Perliferas [Rom03]. Perlita [Rub01, ZBABR00]. Permanence [Jan07]. Permanencia [Jan07]. Permanentemente [BMC^{+}09]. Permanently [BMC^{+}09]. Permeabilidad [CS06, Cas03]. Permeability [Cas03, CS06]. Permisos [Ros06]. Permits [AR09, Ros06]. pero [KCR^{+}08]. Peromyscus [PLA00]. Perros [Cre01a, CB01b, FND06, Kot00, LADM05, Shi06]. Persia [BDSD^{+}05]. Persian [BDSD^{+}05, ZSSM09]. Persist [PDH00]. Persisten [PDH00]. Persistence [EKL^{+}00, EMW^{+}04, GM09, JF04, LLS08, LWV04, MP07b, MY02, Sch03a, SHM00, SLB^{+}07, SLS05, SB08, TSG06, TTSP04, WBBA03, YL05]. Persistencia [EMW^{+}04, GM09, JF04, LLS08, LWV04, MY02, Sch03a, SLS05, SB08, TSG06, TTSP04, WBBA03, YL05, EKL^{+}00, MP07b, SHM00, SLB^{+}07]. Personal [Car04, DD01, RS02b]. Personales [RS02b]. Personnel [DD01]. Perspectiva [FC05, OB00]. Perspectivas [DK06, DP01, LC07a, MBM04, MRAF^{+}07, PMC^{+}07, SC09, SS08a, SMLB00, TK03a, Cze08, Ped03]. Perspective [Ano00b, Car04, Cha06, Fai02, FC05, Fol06, Gil01a, KK08, KT06, LC07a, LSF04, Oso00, OB00, Pim07a, VS05a, WB06, Wie06]. Perspectives [Ank04, DP01, KR00, MBM04, MRAF^{+}07, PMC^{+}07, SRL^{+}07, SMLB00, Tod00, TK03a]. Perturbación [BMA^{+}08, BS08, GN00, GBMH07, HH00, HH06b, JSWS00, KKHS05, KGM^{+}02, KFBG08, KL06, Lai00, MNRD^{+}08, MP07b, New06, PAHM09, RS02b, SSSK06, SBB^{+}08, SH02, SHYH06, WSQP07, WD03, WGK07, FES01]. Perturbaciones [BSW^{+}06, Kes01, MEA00, SMG00, GGC03]. Perturbadas [MP09]. Perturbados [WBW05]. Perturbation [BHW01]. Peru [BL01, MBR^{+}02, Mic08, NTMT^{+}03, OSSK^{+}07, SVV07, BL01, MBR^{+}02, NTMT^{+}03, OSSK^{+}07, SVV07]. Peruana [Gav04]. Peruvian [Gav04, PPPV08]. Pesca [AYKM04, HR04b, KSH00, Kai03, KB03, LARK01, Lyn06, Rut01]. Pescadores [CDM09, GEJK05, GMB^{+}08]. Peso [CT07]. Pesquería [CDM09, DMG^{+}00, GFTB04, MBR^{+}02, WELCH09]. Pesquerías [BB06, CSB07, DLCV00, ES04, FVS03, GWJ^{+}07, HAG^{+}09, KEJ06, Kri04, LCS03, RDN06, WELCH09, ŽWGW09]. Pessimism [Ano01i]. Pessimistic [Bro00]. Pest [Cho06, DDAK04]. peste [KFK^{+}09]. Pesticidas [BS00, DSJ02]. Pesticide [BS00, DSJ02]. Pesticides [LA00]. Pests [APSB02, Hod04b]. Petén [FVMG03]. Petirrojo [AE02]. Petirrojos [DA07]. Petrolera [TB03]. Pets [RSW05]. Pez [DDBA^{+}07, KB03, LG05a, Tow03]. pH [BBB08, PRL^{+}02]. phasianellus [ARMH04]. Phenological [HDQSL06]. Phenology [Bee02, BRK^{+}02, FLQ03, GB01, PC02]. Phenomenon [Paq07]. Phenotypic [ADF04, OB00]. Pheromone [LS09]. Phifer [PS01b]. philia [Fle04b]. Philopatry [SS02a]. Philosophical [BK09]. Philosophies [SC09]. Philosophy [Kaw04]. Philpott [RKR03b]. Phlox [HK00]. Photographs [Mcc09, RLK^{+}01]. Phragmites [SB04]. Phrase [Gre09]. Phylogenetic [ADF04, Fai02, FRH04, Fai08, HMF^{+}01, PEC01]. Phylogeography [DSE^{+}08]. Physical [KU02]. Physiological [WBW05]. Physiology [GMW08]. Phytophagous [ØDEA00]. Picture [Mul07]. Pieces [Dix07]. Piecing [She02]. Pigeon [DGSS06, EM03]. Pigs [RPM09, TRSR00]. Pilar [Kot00]. pilosa [HK00]. Pine [BMM^{+}01, GMA07, GBC^{+}04, aARLG03, LGK02, MDJP04]. Pingüinos [WBW05, WBW06]. Pino [GMA07, GBC^{+}04]. Pinos [BMM^{+}01, aARLG03]. Pinus [BM04, LGK02]. Piojos [Fra09]. Pioneros [Sha01]. Pipeline [Sey04]. Pipelines [RG04b]. Pírica [FEKH09]. Pirodiversidad [PA06a]. Piromancia [Pyn04]. Pisoteo [ANW^{+}08]. Pitch [GBC^{+}04]. Pitfalls [LHG^{+}03, WDA^{+}06]. Pitkin [OK01a, OK01a]. Place [BAF04, Che02a, CML02, FN02b]. Plagas [APSB02, Cho06]. Plague [KFK^{+}09, Mae04]. Plains [GSB04, HBV02, SWR00]. Plan [Cha06, CHBK02, DW06, HODB09, MML06, NB06, Pow06, PK05, Rap06, RWBG06, TFGJ06]. Planeación [CPSS03, CWCG02, FN02a, LWSM03, LHG^{+}03, RG02, MH02, HDK^{+}00, SRC^{+}02]. Planeador [Lat00]. Planes [BTK08, Cla01, CHBK02, GS01, HCE^{+}01, HP01, Lat00, PK05, Sem02, Wil02a, ZLM03]. Planet [Bin02, CLW^{+}09, Hoo07, WMLS09]. Planeta [CLW^{+}09, WMLS09]. Planicies [SWR00]. Planificación [BRV06, CWM08, Dal05, Fje07, HBKF05, Les05, MGH^{+}09, NLCAR09, NP06, RRK02, RSB05, SSB^{+}08, SHT^{+}05, TFG08, MDF05, MPE^{+}07, PCC03, RB07, TWW09, WE07, WHA^{+}05]. Planificar [MU08]. Planner [Lat00]. Planning [Ara09a, BRV06, BDR04a, CWM08, CPSS03, CWCG02, Dal05, Fje07, FN02a, GALW09, HBKF05, HDK^{+}00, Lat00, LWSM03, Les05, LHG^{+}03, Lyn08, MDF05, Mer04, MPE^{+}07, MGH^{+}09, MU08, Moi08, MN07, Nas06, NLCAR09, NP06, PCC03, Pre04, PB08, RRK02, RG02, RSB05, RB07, RP07, Rya06, SRC^{+}02, SSB^{+}08, SHT^{+}05, TFG08, Tho00, TWW09, WP03, WE07, WHA^{+}05]. Plans [Bei03, BTK08, CHBK02, GS01, HCE^{+}01, HP01, Sem02, Wil02a]. Plant [ARPEBM09, BTM^{+}05, BMMR03, Bey04, Bos07, BPL^{+}01, CM06, CSM^{+}06, CCH03, Cun00, DSTPW07, DGC05, DNZK02, EB09, FRR07, FBK02, Gar03a, GB03, GSB^{+}06, Gus02, HFR07, HJ07, HLJV07, HF04, JHHL04, JC07, Kér04, Kes01, KFA^{+}03, KD05, KDD02, LLS08, LH04, LB00, MNB09, MT08, MKK^{+}09, MCA^{+}00, MSBOF^{+}09, NMB09, PJ00, PAHM09, Pat00, PA04, PM07, Qia07, RTN01, Rob00, RAW07, RS01, SM09, SB04, Tho05b, TS06, TR05, VSC06, VTR^{+}09, VIMMO05, VK07, WMP05, WHH00]. Plant-Community [Kes01]. Plant-Mammal [RS01]. Plant-Species [ARPEBM09]. Planta [Bos07, Gar03a, KDD02, MKK^{+}09, MCA^{+}00, ØDEA00, RS01, SM09, TS06, TR05, VSC06, WHH00]. Planta-Animal [RS01]. Plantaciones [ARP05, DBB^{+}09a, GRJB^{+}07, Hie05, HN09, PV02, SMFPCA05]. Plantas [ARPEBM09, BTM^{+}05, Bey04, BTS09, BPL^{+}01, CDD^{+}05, CM06, CSM^{+}06, Cun00, DDAK04, DSTPW07, DGC05, DNZK02, EB09, Fox07, FRR07, FBK02, GB03, GSB^{+}06, GHP^{+}05, Gus02, HMH00, HCRM09, HH08, HFR07, HJ07, HLJV07, HF04, HN09, JHHL04, JC07, Kal05, Kee06, Kér04, Kes01, KS03, KFA^{+}03, KD05, LLS08, LH04, LLW02, MNB09, MT08, MS08, MSBOF^{+}09, NHHG^{+}05, NFM04, NBM^{+}04, NMB09, OBM09, PJ00, PAHM09, PA04, PM07, Qia07, RTN01, Rob00, RAW07, SM09, SL08, SB04, TS09, Tho05b, VJ07, VEST02, VTR^{+}09, VIMMO05, VK07, WCPB03, WMP05, WRL02, RAW07]. Plantation [GRJB^{+}07]. Plantations [ARP05, DBB^{+}09a, Hie05, HN09, PV02, SMFPCA05]. Planteamiento [LMS^{+}02]. Planting [CLC^{+}08, KPJ06]. Plants [BTS09, CDD^{+}05, DDAK04, Fie00b, Fox07, GHP^{+}05, HMH00, HCRM09, HH08, HN09, Kal05, Kee06, KS03, LA00, LLW02, MS08, NFM04, NBM^{+}04, OBM09, RAW07, SL08, Sie00, Tal04, TS09, VJ07, VK07, WCPB03, WRL02]. planus [LO02]. Plateado [AMF^{+}01]. Playa [BBM04, SH02]. Playero [SH02]. Playing [Fox08]. Plaza [CBG^{+}00]. Plazo [BSW^{+}06, CWL^{+}00, CWH07, DBML07, FTC^{+}04b, GMW08, GWJ^{+}07, Gut05, IMD01, LLH08, PPW^{+}07, SRW^{+}08, Sch03a, SMH^{+}07, SLKP06, SBP07, SBSL06, VEST02, WBDP02, WGGS07]. Plea [BEH06]. Pleistocene [GD05, Wil02b]. Pleistoceno [GD05]. Plethodon [GK06]. Plethodontid [WD01]. Plethodóntidas [WD01]. Pletscher [PR01a]. pleuriticus [NR03]. Pliny [Par03]. plumbeus [MKH07]. Pluralism [Sar07]. pneumonanthe [VBOV05]. Poaceae [LLW02]. Poachers [WZD^{+}00]. Poaching [GMW08, MBL02, WTEH^{+}01]. Población [AP00, AE02, BRBMS06, CLBC03, DLCV00, DW05a, EMW^{+}04, ENN^{+}08, FHL^{+}04, FHB^{+}08, Gar03a, HJ05, HGH^{+}03, JF04, JRL03, KD00, KDD02, LM05, MBQ05, MM04, Mcs05, Mur01, NKBD02, PM07, PBN^{+}04, PPBW08, RSJE06, SFS00, VBOV05, War05, WSQP07]. Poblacional [APCS05, AT05, BST^{+}05, BJTD03, BNP^{+}07, BJE^{+}05, Cal06, CWNW01, Chi00a, CNC04, Cre01a, CV08, EKL^{+}00, ED07, FMDK09, FPF00, FBCG00, HHD^{+}08, HRS00, HCK^{+}02, HLSZ02, HK00, JB06, Kal05, Kel01, KGM01, LG05a, LDG^{+}00, ML04, MPDT01, MAP03, MM04, NLCAR09, PV06, Pod01, Pop08, QAMW03, RLC^{+}00, RMD^{+}02, Ree05, SRR00, SGSA07, SM07a, Sch03b, SFS05, SH03, SHM00, Sim03, SLCW08, TWGH00, TS06, TSP^{+}08, TS03, VWQI00, WRM05, YE07, ZSMW06]. Poblacionales [BMP^{+}06, CW09, For00b, FH09, FTC^{+}04b, LRW03, Pol06, Phi00]. Poblaciones [APCS05, ALH^{+}04, AM05, AW01, ASD01, ANW^{+}08, Bat04, BNP^{+}07, Bra06, CWC00, Car07, CCC04, Cor02, Cou02, DMS03, DA07, DOW^{+}06, Dun02, FP02, FG00, FTC04a, FBL09a, FPF00, GRH05, GS02, GSB^{+}06, GSG08, GGW09, HMH00, HLG03, Her09, HY03, JTS06, JGP09, KW02, KMGB07, Kér04, KGS^{+}07, KLH08, LME08, LLS08, LCZ^{+}05, Log03, LP01, LWV04, LJMR^{+}01, MT08, MR00b, MMBP08, Mor00b, MY02, NT01, NL05, OMP^{+}05, PAHM09, RW02, RAW07, SFD09, SWK05, ST05, STS05, SG04, TTSP04, TTB03, Tow03, UCAGG08, Vee06, VSAG06, WBDP02, Wed02, WBSG07, WSBP09, Fau01, FNP09, KB03, YE07, YLR^{+}01]. Pobres [GR03]. Poco [HW02, SS06]. Poder [HMK^{+}09, SM01, TWGH00]. Podrán [DP01]. Pohnpei [BBW^{+}07]. Point [O’K01b]. Pointers [PK01]. Points [Gil04]. Policies [FP07, GGT08, Rom09, VGB^{+}04]. Policy [AAO^{+}09, BT07a, BBB^{+}07, BL07, CMBC00, CMC^{+}00, Cla01, Cur03, DWWF04, DWW04, FB08, GC09, HH06a, Hei03, KWL^{+}09, Kra01, Lac07, LW06, Law08, LSF03, MB07, Mas03, MKVZ^{+}09, Mef07a, NFD^{+}09, PBC^{+}09, PTP^{+}08, RT00, RPT00, RC05, RANM08, Rud06, SRL^{+}07, TG05, TSH06, Tur06, WGW01, Win04]. Policy-Oriented [Cla01]. Polillas [BEB00, GBC^{+}04, RDEF01]. Polinización [GC07]. Polinizado [EBAW01]. Polinizadores [DNZK02, Ric04b]. Política [Cla01, DC06b, HH06a, HCBK00, LSF03, MB07, SM01, Law08, WGW01]. Political [AO06, Bei08, CS05, HCBK00, LG05b, SM01]. Políticas [CMBC00, CMC^{+}00, DWWF04, FB08, GGT08, KWL^{+}09, PBC^{+}09, PTP^{+}08, RPT00, RC05, Rud06, TG05, Tur06, VGB^{+}04, DWW04]. Político [Bei08, GC09]. Políticos [CS05, LG05b, MKVZ^{+}09]. Politics [And01c, Ano02h, Che02a, DC06b, Joh07, Led06, Mef07b, Nos01b, Orr01c]. Pollinated [EBAW01]. Pollination [Cor00, GC07, KR00, Pat00, Rou00]. Pollinator [CE00, DNZK02, Ric04b]. Pollos [WBW05]. Pollutant [WCP01]. Pollution [PC08, SGG07]. Pond [JMLG01, PC02, PKM00]. Pond-Breeding [PC02]. Ponderación [AHTM05]. Ponds [BSM08]. Poniendo [FN02b]. Pool [Che02b, HRS08]. Pool-Breeding [HRS08]. Pools [ASF^{+}09, CMCG07]. Poor [Dal02, GR03, Mcs06, PBN^{+}04, PDH00]. Poorly [HW02]. Popular [Cua02, Ful00, JHHL04]. Popularity [GM08b]. Population [APCS05, AL03, AT05, AM05, Ano04e, AP00, AE02, BMP^{+}06, BST^{+}05, BJTD03, BNP^{+}07, BRBMS06, Bro00, BJE^{+}05, Cal06, CLBC03, CWNW01, Chi00a, CNC04, Côt05, CW09, Cre01a, CV08, DLCV00, DW05a, EKL^{+}00, EFLW02, ERB01, EBDM06, EMW^{+}04, ED07, ENN^{+}08, FMDK09, FPF00, FHL^{+}04, FHB^{+}08, For00b, FH09, FTC^{+}04b, FBCG00, Gar03a, GSG08, GNEE07, GRG^{+}05, HHD^{+}08, HRS00, HCK^{+}02, HLSZ02, HK00, HJ05, HGH^{+}03, JF04, JRL03, JB06, Kal05, KD00, Kel01, KDD02, KGM01, LG05a, LLS08, LM05, LRW03, LP01, LWV04, LDG^{+}00, ML04, MT08, MBQAM06, MPDT01, MAP03, MBQ05, MM04, Mcs05, Men03, Mur01, NLCAR09, NKBD02, PM07, PBN^{+}04, PD00, PPBW08, PV06, PR01a, Phi00, PS00b, Pod01, Pol06, Pop08, Put07, QAMW03, RLC^{+}00, RMD^{+}02, ROB^{+}04, Ree05, RSJE06, SRR00, SPFF^{+}08]. Population [SGSA07, SM07a, Sch03b, SFS05, SH03, SHM00, SSK^{+}05, Sim03, SFS00, SLCW08, STS05, TWGH00, TS06, TSP^{+}08, Tow03, TS03, VBOV05, VWQI00, War05, WSQP07, WRM05, Wie06, Wil03, YE07, YLR^{+}01, ZSMW06]. Populations [APCS05, ALH^{+}04, AW01, ASD01, ANW^{+}08, Bat04, BNP^{+}07, Bra06, But03, CWC00, Car07, CCC04, Cor02, Cou02, DMS03, DA07, DOW^{+}06, Dun02, EC01, Fau01, FP02, FG00, FTC04a, FNP09, FBL09a, FPF00, Fra01, GRH05, GS02, GVS09, GSB^{+}06, GGW09, HMH00, HLG03, Her09, HY03, JTS06, JGP09, KW02, KMGB07, Kér04, KGS^{+}07, KB03, KLH08, LME08, LCZ^{+}05, Log03, LJMR^{+}01, MR00b, MMBP08, Mor00b, MY02, NT01, NL05, OMP^{+}05, PAHM09, RW02, RAW07, Rya06, SFD09, Sch02, SWK05, ST05, SG04, TTSP04, TTB03, UCAGG08, Vee06, VSAG06, WBDP02, Wed02, WBSG07, WSBP09, YE07]. Populus [SMG00]. por [AP00, AYKM04, AHR^{+}09, BGC^{+}04, BPB07, BST^{+}05, BS08, BADF^{+}05, BV01a, BGWE00, CS06, CMM05a, Cle06b, CH01, Cre01a, DH00a, DS08, DMS03, DWN01, EBAW01, EB09, FYC00, FG00, FJ00, Fer01, For00a, GRB02, GGGS00, Gav04, GFTB04, GGC03, Had00, HPK^{+}04, HM04, JLF09, JMLG01, KHM00, KFL08, KJSH08, KFNB09, KPH07, LLS08, Mad05, MMLM01, MBQ05, MSN06, MF00, MHM09, MM02b, MLB08, MBC03, MH08, MSSM09, MMB^{+}03, NTMT^{+}03, NWRD03, NV09, NSB^{+}06, OBK^{+}04, PCL^{+}03, PL03, PLL^{+}05, PP03, PC08, PBL00, PB04, PPW09, RL06, RCP00, RLK^{+}01, RS02b, RNL^{+}04, SH09, SMG00, SS00, SRHS09, TS09, TNTH^{+}04, UMT^{+}01, Vee06, VK07, WBW05, Wan04, WMD^{+}04, WTR01, Wun07, ZR00]. Porción [WABT07a]. Porque [BV01b, NV09, Per05a, PDH00, DP01, GK08].

Portage [DGC05]. Portion [Gre09, WABT07a]. Posed [HLSZ02]. Posesión [CMM^{+}05b]. Posgrado [KSC^{+}06]. Posible [BRA08, MBC03]. Posibles [Sta01]. Positivas [LW01]. Positive [LW01]. Possibility [Jac08]. Possible [Awo08, BRA08, MBC03, Sta01]. Possum [GLV05]. Post [BRK^{+}04, Bey04, FVMG03, RBRB06]. Post-Agricultural [FVMG03]. Post-Incendio [BRK^{+}04, Bey04]. Postdispersal [PDF^{+}00]. Posterior [VMBC02]. Posterity [Led06]. Postfire [BRK^{+}04, Bey04, Hut06, RBRB06]. Postscript [Ehr01]. Potatoes [HGH^{+}00]. Potencial [ACMA04, CPD^{+}09, Fai08, FM09, Gui07, HWH^{+}02, LF09, LBG03, LHPH06, Old03, PMC^{+}07, PBS03, RL06, TAJ04, WDA^{+}06, LSN07]. Potenciales [EM03, HBBD08, LMP00, LO06, WCP01, Web01]. Potencialmente [PZG^{+}05]. Potential [ACMA04, CPD^{+}09, EM03, Fai08, FM09, Gui07, HBBD08, HWH^{+}02, Koh07, LMP00, LBG03, LSN07, LHPH06, LO06, LFA^{+}07, Mal03, Old03, PMC^{+}07, PBS03, RL06, TAJ04, WDA^{+}06, WCP01, Web01]. Potentially [PZG^{+}05]. Potreros [NH09]. Pouched [Lei05]. Poverty [BISS06, Put07]. Povilitis [Ang01]. Power [Flo05, HMK^{+}09, Put07, SM01, TWGH00]. Powered [RS02b]. Práctica [DKW01, DBML07, KDC^{+}06, LMMF00, PK01, SMRR02]. Practical [LMMF00]. Practically [ID00]. Practicantes [GBH^{+}05]. Prácticas [BE00, FBC^{+}06, FMLL06, GNEE07, KHHK03, PDH00, Pyk04, SMC00, PTP^{+}08]. Practice [Chi00b, DKW01, DBML07, KDC^{+}06, KBMN00, PK01, SMRR02, SBS^{+}09]. Practices [BE00, FBC^{+}06, FMLL06, GNEE07, KHHK03, PDH00, Pyk04, PTP^{+}08, SMC00]. Practicing [Bea09a, Ker09]. Practitioners [GBH^{+}05]. Pradera [Kot00, PPW09, Shi06, Vee06, WMP05]. Praderas [BMMW07, Cas03, HRW^{+}03, JGM03a, MMLM01, Pan02, RDW01, SMC00, SD03]. Pragmatic [Mag04, Sar07]. Pragmatism [SBS^{+}09]. Prairie [BJTD03, BMMW07, CCH03, HRW^{+}03, JGM03a, Kot00, MMLM01, MRH^{+}00, Pan02, PPW09, RJW09, RDW01, Shi06, SMC00, WMP05]. Praise [PS05]. Preaching [Nad04]. Prealpes [PCVC06]. Prealps [PCVC06]. Precautionary [DAB06, TWGH00]. Precedente [Ric07]. Precio [FN04]. Precios [Far08, WG01, WSA^{+}05]. Precision [EFLW02, AM05, MPDT01]. Precisos [SS08a]. Predation [BS01b, BEM^{+}04, GC07, HAJ^{+}04, Sch03b, STC03, TNTH^{+}04, Vic02, WMS^{+}01, WZD^{+}00]. Predator [Ber06, HDA01, LS09]. Predators [BSW^{+}08, CTR02, FN04, GBE00, HM04, RL06, TB04, TPW03]. Predatory [FMSL08]. Predecir [EM08, FTN^{+}05, GBC^{+}04, MBQAM06, NFB05, PR03, Wen08]. Predicción [FNFM01, GGA01, DGSS03, DR02, FCG^{+}05, FRSG^{+}02, GC06, GGW09, GM09, HRS08, HRS^{+}07, HC09, JGP09, KDG^{+}06, LD03, PSG^{+}00, SHT^{+}05, SWB^{+}08, TNFH07, TNTH^{+}04, WMP05, WTR01]. Predicciones [KFL08, LWV04, MAP03]. Predice [HPG09]. Predicen [GBH^{+}09, LRW03]. Predict [EM08, FTN^{+}05, GBH^{+}09, MBQAM06, NFB05, PR03, Wen08]. Predicted [MBMLJ09, PN05]. Predicting [DGSS03, DR02, FCG^{+}05, FNFM01, FRSG^{+}02, GGA01, GBC^{+}04, GGW09, GM09, HRS08, HRS^{+}07, JGP09, KDG^{+}06, LRW03, SHT^{+}05, SWB^{+}08, TNFH07, TNTH^{+}04, WMP05, WTR01]. Prediction [GC06, HC09]. Predictions [KFL08, LWV04, MAP03]. Predictiva [SBDD05]. Predictive [FNF03, LB05, LMP02, MN03, NF04, SBDD05]. Predictivo [LMP02, MN03, NF04]. Predictivos [FNF03, LB05]. Predictores [KPJ06, aARLG03, SBDD05, LB02]. Predictors [KPJ06, aARLG03, LB02, OHC08]. Predicts [HPG09]. Predispersal [PDF^{+}00]. Preemptive [FO06]. Preference [JLF09, MHM09]. Preferences [Bal00, BCS04, GRB02, War00]. Preferencia [JLF09, GRB02, MHM09]. Preferencias [BCS04]. Preguntas [SAA^{+}09]. Prehistoric [LW02b, PHW05]. Prehistórica [PHW05]. Prehistórico [LW02b]. Prehistory [Leo04]. Preliminar [IML^{+}05]. Preliminary [IML^{+}05]. Premontane [RCP00]. Premontano [RCP00]. Preparing [MSL06]. Presa [PCL^{+}03]. Presas [Ber07, ESSGL00, LS09, MY02, WSQP07]. Prescribed [ASD01, BM04, HLB^{+}03, Pan02, RO04, SF03, Vic02]. Prescribing [RR00]. Prescripción [RR00]. Prescrita [ASD01]. Prescritas [HLB^{+}03, SF03]. Prescrito [BM04, Vic02]. Presence [Gus02, JFWP06, NHHG^{+}05, Pol06, ZLM03]. Presence-Absence [Pol06]. Presencia [FNFM01, JFWP06, MM03a, Pol06, ZLM03, Gus02]. Presencia-Ausencia [JFWP06, Pol06]. Present [VTR^{+}09]. Presente [VTR^{+}09]. Preservación [DN01, SPFF^{+}08]. Preservation [CWL^{+}00, DN01]. Preserves [ZGR02]. Preserving [SPFF^{+}08]. Presettlement [EM03]. President [Bec07]. Presión [FYC00, HJ05, KPJ06]. Presiones [FPW01, RS01]. Pressey [KC08]. Pressure [FYC00, HJ05, KPJ06, Put07, SSK^{+}05]. Pressures [FPW01, RS01]. Prevalence [KFK^{+}09, KPH07, SRL^{+}07]. Prevalencia [KPH07, KFK^{+}09]. Prevención [KFL08]. Prevenir [CHFR09]. Prevent [CHFR09]. Preventing [CAFL03, KFL08]. Preventiva [FO06]. Preventivas [DAB06, TWGH00]. Previamente [BJE^{+}05]. Previenen [CAFL03]. Previously [BJE^{+}05]. Previstos [MBMLJ09]. Prey [Ber07, LS09]. Price [FN04, WG01]. Prices [Far08, WSA^{+}05]. Pricing [WGW01]. Primaria [BCHS08]. Primario [SAA^{+}01]. Primarios [GRJB^{+}07, STC03]. Primary [BCHS08, GRJB^{+}07, SAA^{+}01, STC03, Sie00]. Primate [CBG^{+}00, GC06, MN03]. Primates [CBG^{+}00, GC06, MN03]. Primera [ZSVL07]. Primeval [Wes05]. Primula [LSL^{+}06, VEST02]. Principal [BCP^{+}02]. Principales [KWL^{+}09]. Principios [BAF04, DS08, GGT08]. Principles [BAF04, DS08, GGT08, TOR^{+}04]. Prioridad [FMPB07, HK03]. Prioridades [ADF04, Álv02, Che09, HPF^{+}06, HJP05, HCM03, LV04, MWBP07, PEC01, RM06, RM01, TVB^{+}06, YLR^{+}01, NFD^{+}09]. Prioritarias [BMBM06, FMT^{+}04, KCR^{+}08, MN03, MPR^{+}01, MBB^{+}03, PMC01, SBB^{+}08, Spe02]. Priorities [ADF04, Álv02, Che09, Dud02, HPF^{+}06, HJP05, JP08, LV04, NFD^{+}09, Pem01, PEC01, RM01, She01, YLR^{+}01]. Prioritization [FBL09a, JMP09, RM06, SF05, TVB^{+}06]. Prioritize [BNP^{+}07, Hay09b]. Prioritizing [BPV09, BMB^{+}08, HLB^{+}03, NRBM05, Pyk05, Rua00]. Priority [BGRJ00, BMBM06, FMT^{+}04, FMPB07, HK03, HCM03, KC07, KCR^{+}08, MWBP07, MPR^{+}01, MBB^{+}03, PMC01, RRCOM^{+}06, SSK^{+}05, SBB^{+}08, Spe02]. Priorización [BPV09, FBL09a, JMP09, NRBM05, Pyk05, BMB^{+}08, SF05]. Priorizar [BNP^{+}07, Hay09b, HLB^{+}03, Rua00]. Privadas [HM03, LMMF00, MH01b, NRBM05, ZMS^{+}08]. Privado [PS00a, RLC^{+}07]. Privados [LLS00, WTEK08, WE07]. Private [HM03, Jam02, LLS00, LMMF00, MH01b, MN07, NRBM05, Nor00, RLC^{+}07, Tho00, WTEK08, WE07, ZMS^{+}08]. Privatization [Fue06]. Pro [Nos01b, PP08a]. Pro-Life [Nos01b]. Probabilidades [Fai08]. Probabilistic [DKS^{+}09]. Probabilísticos [DKS^{+}09]. Probabilities [Fai08]. Probability [KL05]. Probable [PC08]. Probando [SGH02, BFH^{+}06, KDD^{+}04, LME08, MPDT01]. Probar [PAHM04]. Problem [BK09, KB04b, LW06, Lin03a, Mul07, SMLB00]. Problem-Based [LW06]. Problema [Lin03a, SMLB00]. Problemas [WRL02]. Problems [Bei02b, LFA^{+}07, RO04, WRL02]. Procedimiento [SDL07]. Procedimientos [BMS08]. Procedure [SDL07]. Procedures [BMS08]. procera [FJ07]. Proceso [EGN^{+}05, Van06, ZSG01, CMC^{+}00]. Procesos [Fje07, Hay09b, LN06, RCL^{+}06, VSC06, SFS05]. Process [CMC^{+}00, EGN^{+}05, Hei03, LJI^{+}08, RCL^{+}06, Van06, ZSG01]. Processes [Fje07, Hay09b, LN06, SFS05, Sta09, VSC06]. Procnias [PB04]. Prodicción [PV09b]. Producción [Cra02, EGPS04, FPL^{+}04, GGT08, Pyk00, TO02, ZSVL07]. Produce [FMPB07]. Produced [MA05]. Producir [FMPB07]. Product [FJ07, MKDT01, PPLS01]. Production [Cra02, Don04, EGPS04, FPL^{+}04, GGT08, Koh07, TO02, ZSVL07]. Productive [Bin02]. Productividad [AM05, BCHS08, GMW08, GNEE07, HMH00]. Productivity [AM05, BCHS08, Cam08, CLH^{+}04, GOL^{+}07, GNEE07]. Producto [MKDT01, PPLS01]. Productos [PBSR06]. Products [PBSR06]. Profesional [Cla01]. Professional [Cla01, RRCOM^{+}06]. Professionals [Gra09, Whi03, Ano05g, Ano06f]. Professions [App03]. Profunda [Rob09]. Program [CPSS03, CMGSM00, DD01, EAR^{+}05, IB03, JMS^{+}06a, KHM00, LLS00, Mai04, MML06, Pro04, RMK07, SAPRB07, Sil05, Tem02, ZSG01, Bei02b, Ris02, BRYK01]. Programa [BRYK01, CMGSM00, DD01, JMS^{+}06a, KHM00, LLS00, MML06, RMK07, SMRR02, SAPRB07, ZSG01]. Programas [AAO^{+}07, CPSS03, Her09, KRM^{+}00, LRS09a, MKH07, MS04, MAM^{+}03, PBRG07, SPFF^{+}08, TRCC^{+}05, TWS06, Wal03, WL04, ZKNL02, FPL^{+}04]. Programs [AAO^{+}07, Bre02b, FPL^{+}04, Her09, KRM^{+}00, KRW^{+}00, MKH07, MS04, MGB^{+}07, MAM^{+}03, Mur00, PBRG07, SPFF^{+}08, TM02, TRCC^{+}05, TWS06, VS05a, Wal03, WL04, ZKNL02]. Progresando [SC09]. Progreso [Cze08, MK02a]. Progress [Cze03, Cze08, Har03, MK02a]. Progressive [Cob05]. Project [EAR^{+}05, JMP09, MSL06, Sey04, CHBK02, DBML07]. Project-Based [MSL06]. Projection [Wil03]. Projections [SHM00]. Projects [GALW09]. Promedios [WMVK03]. Promote [KDC^{+}06, Mae02, PPBW08]. Promotes [PKM00]. Promoting [JMM07, LF09]. Promover [BZD03, LF09, PPBW08]. Promueven [KDC^{+}06]. Prone [SHYH06]. Proneness [KSB04]. Pronóstico [MBMLJ09, PN05, BJPL07b, MSOM^{+}06]. Propagación [BGWE00, KFL08]. Propagados [LW02a]. Propagated [LW02a]. Propagation [BGWE00]. Propagation-Assisted [BGWE00]. Propagule [Ber05, HHSC00, KPJ06]. Propágulos [Ber05, HHSC00, KPJ06]. Propensión [KSB04]. Propensos [SHYH06]. Property [MC09, Tho00, Whe02]. Prophet [Caf00a]. Propiedad [BMRB04, Say05, MC09]. Propiedades [TAJ04]. Propó [PRL03]. Propó-sito [PRL03]. Proporción [CMA^{+}06, FNP09]. Proporcionados [KJSH08]. Proporcionar [CHFR09]. Proporciones [GS05]. Proportion [CMA^{+}06]. Proposal [IF02, Orr00a]. Proposals [Bow09]. Proposed [CMGSM00, SFX^{+}08]. Propósitos [Rua00]. Propuesta [IF02, SFX^{+}08]. Propuesto [CMGSM00]. Pros [FRD^{+}09]. Prosobranch [SLELA05]. Prosobranquios [SLELA05]. Prospección [HW02]. Prospect [Orr09b]. Prospectos [WDA^{+}06]. Prospects [Cze08, Dal02, DK06, Ped03, WDA^{+}06]. Prosperidad [NA01]. Prosperity [NA01]. Protagonists [OS07]. Proteaceae [WHA^{+}05]. Protección [CS07b, GLMN00, GAF00, Gut05, HCK^{+}02, Pra05, RLC^{+}07, The03, TK07, TWS06, WG08, DC06b, Fer01, HSR05, JG08, MA08, MML06]. Protect [BV01b, GMMPP08, MMB^{+}03, RS02b, TK07]. Protected [ALM07, BSKH^{+}06, BDR04b, Bro04, CWC00, CNPS04, Dal05, DK06, DHZM09, FVS03, FDL^{+}05, FRR07, FNS^{+}03, GMMPP08, GC09, GM08a, GHL08, HDK^{+}00, Inf01, IIG05, Kal05, KTN^{+}09, KCK^{+}08, KM00, Kri04, Lai00, LSN07, LB01, LARK01, LSZ^{+}08, LGB05, Lyn06, Lyn08, Mas03, MC09, McK02, MYFF^{+}07, MK02a, MA08, PH02, PA04, RRK02, RBP01, RBHKK02, RCMV06, RB05, SMV02, Sil05, TVD03, TB01, UKC^{+}09, VTR^{+}09, VTA02, Wal09, WGW01, WB06, WMD^{+}06, XM07]. Protected-Area [Kri04, MK02a, RRK02, Wal09]. Protecting [Led06, MML06, Raf03, Sar07, TWS06]. Protection [BPV09, CS07b, DC06b, EYZ09, Fer01, GAF00, Gut05, HCK^{+}02, HSR05, Hop07, JG08, MA08, Pra05, RLC^{+}07, SSK^{+}05, SCM^{+}07, The03, Whe02, WG08]. Proteger [BV01b, GMMPP08, MMB^{+}03, RS02b]. Protegida [GM08a, KTN^{+}09, LSZ^{+}08, TB01]. Protegidas [ALM07, BPV09, BSKH^{+}06, CWC00, CNPS04, DK06, DHZM09, FVS03, FDL^{+}05, FRR07, FNS^{+}03, GMMPP08, GHL08, HDK^{+}00, Kal05, KCK^{+}08, KM00, Kri04, Lai00, LSN07, LB01, LARK01, Lyn06, MC09, McK02, MYFF^{+}07, MK02a, MA08, PH02, RRK02, RBP01, RBHKK02, RCMV06, RB05, SMV02, UKC^{+}09, VTR^{+}09, VTA02, Wal09, WGW01, XM07, Dal05, GC09, TVD03]. Protegiendo [HTM^{+}05]. Protestants [Cob05]. Protocol [BS03, JMP09, SS02b, HS06]. Protocolo [JMP09, HS06]. Protocolos [RBM^{+}05]. Protocols [RBM^{+}05].

Proven [SD02]. Proveniente [ZSSM09]. Provide [CHFR09]. Provided [KJSH08]. Province [RTCR03]. Provincia [RTCR03]. Provocar [Gui07]. Provoking [Gro05a]. Proxies [KU02]. Proyecciones [SHM00]. Proyecto [CHBK02, DBML07]. Proyectos [JMP09]. Prueba [CS06, CT07, ME00, VSO04, AP00, SCB^{+}01]. Pruebas [BR01, FBC^{+}06, KSB01, MBL02, RA05, DSJ02, FNF03, PBRG07]. Psicológicos [DS08]. Psychological [DS08, KK08, MF08]. Psychology [Orr08b, SBM06]. Public [BCS04, BRK^{+}04, Bye04, Cla01, Dal05, DWW04, KB08, LSF03, MHM09, Nor03, Pow06, PK01, RH01, Sul09, Tho00, XM07]. Pública [Pow06, LSF03, PK01, RH01]. Públicas [DWW04, KB08, XM07]. Publication [LK04]. Público [BCS04, Cla01, Dal05, MHM09]. Públicos [BRK^{+}04]. Publish [Pri01a]. Pudrición [SLS05]. Puede [For02]. Pueden [CPME09, Fje07]. Puerto [Sta01, DAR08, Sta01]. Puesta [SSL06]. Puestas [VSO04]. Puget [MK02b, MK02b]. Pulmonados [SLELA05]. Pulmonate [SLELA05]. Puma [Mae04]. Pumas [MB09b]. punctatus [ENN^{+}08, TPT^{+}07]. Punta [MBR^{+}02]. Pupfish [RNL^{+}04]. Purpose [PRL03, You00]. Purposes [Asc02, Rua00]. Purshia [MBQAM06]. Put [VSO04, VP04]. Putativas [NFB05]. Putative [NFB05]. Putatively [DAKB08]. Putativos [NFB05]. Putting [Blu07, DKW01, FN02b, KBMN00, LC07a, Mul09, PV09a]. PVA [BBA^{+}02, CB01b, KG03]. Pyric [FEKH09]. Pyrodiversity [PA06a]. Pyromancy [Pyn04].

Qinling [GJLL08, LZD^{+}03, GJLL08, LZD^{+}03]. Qualified [Fie03c]. Quality [BCP^{+}02, CMC^{+}01, EKL^{+}00, FRSG^{+}02, Ful02, JSHM06, LP08, MSB^{+}01, MW07, PV02, PV09a, Ste08a, SB08, VO03]. Quantification [MCG^{+}08, Rau02, SBB^{+}08]. Quantifying [BOJ09, FKU05, Fai02, Fie03a, GWM^{+}07, HMK^{+}09, LLS08, ND04, OB00, PCVC06, RPM09]. Quantitative [BCP^{+}02, GV05, HCE^{+}01, MW07, Wal04]. Quantity [MSB^{+}01]. Quarries [BKK03]. que [BADF^{+}05, BEB00, CSB07, CDM09, CW00, CMCG07, CAFL03, CW02, CS07b, Cre01a, DGC05, ET03, FBCG00, GSG08, Gra01, GBH^{+}09, HRS08, HLSZ02, KSDBT02, KDC^{+}06, Koe03, LS01, MMM03, MY02, MS09, NHHG^{+}05, PC02, PBC^{+}09, Rue05, SP00, MHGF04, NV09, Kot00]. Quebrantahuesos [MH01b]. Queensland [WA05]. Quema [NTMT^{+}03, Ram01, VS09, ASD01, PA06a, TK07]. Quemas [HLB^{+}03, SF03]. Question [MF08]. Questionable [Cle06b]. Questions [BBA^{+}02, Joh09a, SAA^{+}09]. Quien [Har06a, Har06a, MHGF04, MHGF04]. quinata [FLQ03]. Quitridio [MSSM09]. Quitridiomicosis [Mcc05, OMP^{+}05, WA05]. Quitridios [KPH07]. quitzeoensis [DDBA^{+}07].

R [Apl01]. Rabbit [Cal06]. Rabbits [DMDFV08]. Rabia [HLSZ02]. Rabies [HLSZ02]. Raccoon [Log03]. Racionalizar [Hay09b]. Radar [BGD09, SAA^{+}01]. Radiación [BBB08, BOC00, GZK06, MHC^{+}01, SKP00].

Radiaciones [WG06]. Radiation [BBB08, BOC00, GZK06, MHC^{+}01, SKP00, WG06]. Radical [Nel06]. radicalis [EGPS04]. Raigones [Hut06]. Railroad [IML^{+}05]. Rain [Gor08, Har06b]. Rainbow [MSOM^{+}06]. Rainforest [Ano00b, ARPEBM09, EAA^{+}09, FBL^{+}09b, GGGS00, HN09, JHHL04, LSL04, LCT^{+}06, LLH08, LP07, MO04, NBM^{+}04, Ram01, SWD00, SSY06, SBSL06, WLCS00, WA05]. Rainforests [FR07, LCG^{+}08, MEA00]. Rains [GGGS00]. ramosissim [SMG00]. ramosissima [SMG00]. Rana [BRBMS06, KM00, PRL^{+}02, PGP04, SWR00, SM07a, WM05, WG06]. Ranas [BD06, BSM08, BOC00, CFC02, DAR08, FH09, GB01, HVR08, Mcc05, PMEH05, SG01, WSKH08, WA05]. Ranch [Whi08]. Ranchos [JSJV03]. Random [HW08]. Range [Car07, CT07, HP08, HRS^{+}07, PSGA09, Sch03a, SSFL08, SLCJ08, VJ07, WABT07a, Wil09a, Gre09, WSLW00]. Range-Wide [PSGA09]. Rangelands [Say05]. Ranges [JSW08, Web01]. Ranging [FPW01]. Rango [CT07, HP08, SLCJ08, VJ07]. Rangos [JSW08, Sch03a, SSFL08, Web01]. Ranid [WSKH08]. Ranidae [BS00]. Ranked [GOL^{+}07]. Ranking [Col01, HSMK07, Ova03, TS09]. Rankings [DHW01, PE01]. Ranunculus [KFA^{+}03]. rapa [WRL02]. Rapaces [BGC^{+}04, TRNH00]. Rapaz [RAL^{+}04]. Rapid [GM08a, HR04c, HC09, JHHL04, KSP00, KDD02, ML04, MP09, Nos01a, Ric05, SFD09, SZS01, SA04]. Rápida [GM08a, HC09, JHHL04, KSP00, KDD02, LDG^{+}00, MP09, SFD09, SZS01]. Rapidly [LDG^{+}00]. Rápido [HR04c, ML04]. Raptor [RAL^{+}04]. Raptors [BGC^{+}04, TRNH00]. Rara [CWNW01, CNC04, EKL^{+}00, GPMM08, Gut05, KHP04, Vic02, YL05]. Raras [DKS^{+}09, Fox07, GBE^{+}06, HHD^{+}08, HW08, HJ07, LWSM03, LTC^{+}08, LLW02, LBTS05, MML06, SL08, SS06, TWS06, VTR^{+}09]. Rare [CWNW01, CNC04, DKS^{+}09, EKL^{+}00, Fox07, GPMM08, GBE^{+}06, Gun02, Gut05, HHD^{+}08, HW08, HJ07, HKM06, KHP04, LG05a, LS09, LWSM03, LTC^{+}08, LA00, LLW02, LB00, LBTS05, MML06, Pyk04, SL08, SS06, TWS06, VTR^{+}09, Vic02, WM05, YL05]. Rareza [FKU05, GSR^{+}04, GBC^{+}04, HK03, MDCK07, MH00, RG02]. Raris [Kir08]. Rarity [FKU05, GSR^{+}04, GBC^{+}04, HK03, MH00, MDCK07, RG02]. Raro [Gun02, LG05a]. Raros [HKM06, LS09]. Rasa [KHHK03]. Rastreo [HPK^{+}04, WM01]. Rat [APCS05, APS07, WBDP02]. Rata [Log03]. Ratán [Sie04]. Ratas [JTZ^{+}08, TPT^{+}07, MTB00]. Rate [Kér04, MRDC06, PJW^{+}02]. Rates [BM05a, BBK02, dLBBG^{+}08, DWN01, Fea05, HMF03, McK02, MM04, SS01, UCAGG08, ZGR02]. Ratio [FNP09]. Rationale [Gha07, Gha08, KDKS08]. Rationality [Ano08a]. Rationalize [Hay09b]. Ratios [GS05]. Ratón [PLA00, PN05]. Ratoneras [WBDP02]. Ratones [BBM04, GM00a]. Rats [JTZ^{+}08, MTB00, TPT^{+}07]. Rattan [Sie04]. Rattus [APS07, APCS05]. Rayas [DMG^{+}00, DR02]. Raymond [Jar03]. Razas [RMSG^{+}03, Rua00, TVB^{+}06]. Re [Wal06]. Re-Visioning [Wal06]. Rea [TB01]. Reaching [Joh05a]. Reaction [Tro00]. Read [Sal05].

Reader [Fra04]. Readiness [CDM09]. Reading [Pyn04]. Ready [Kni09]. Real [All09, CLŽ^{+}07, Got06, HSB01, Mul07, PC01, Rob06b, TB04, VP04, Wil02c]. Real-World [Rob06b]. Reales [HSB01, TB04]. Realidad [BDSD^{+}05]. Realism [YBWZ06]. Realista [FM09]. Realistic [FM09]. Realities [RG06, She01]. Reality [BDSD^{+}05, Car02b, Lei05, Tas09, ZZ05]. Really [DA02, Kar09]. Realm [Mar00a]. Reanálisis [HMF^{+}01]. Reanalizar [PMKDS01]. Reanalysis [HMF^{+}01]. Reaping [ZSVL07]. Rearranging [PS01b]. Reassess [PMKDS01]. Recaptura [CLBC03, LG05a, PV06]. Recapture [CLBC03, LG05a, PV06]. Received [Ano01a, Ano02a, Ano03c, Ano03f, Ano04i, Ano04k, Ano04h, Ano04j, Ano04g, Ano05b, Ano05j, Ano05i, Ano05h, Ano06b, Ano06j, Ano06h, Ano06i, Ano06g, Ano07d, Ano07o, Ano07r, Ano07p, Ano07q, Ano07n, Ano08c, Ano08n, Ano08q, Ano08p, Ano08o, Ano08m, Ano09b, Ano09k, Ano09m, Ano09l, Ano09j, Ano09n]. Recently [Ano03f, Ano04i, Ano04k, Ano04h, Ano04j, Ano04g, Ano05j, Ano05i, Ano05h, Ano06j, Ano06h, Ano06i, Ano06g, Ano07o, Ano07r, Ano07p, Ano07q, Ano07n, Ano08n, Ano08q, Ano08p, Ano08o, Ano08m, Ano09k, Ano09m, Ano09l, Ano09j, Ano09n, NT01]. Recesión [Sch03a]. Recession [Sch03a]. Recetas [CMC03]. Reciente [CWL^{+}00, JRL03]. Recientemente [NT01]. Recientes [McC01]. Recipes [CMC03]. Reclutameinto [NAC^{+}06]. Reclutamiento [CH01, HW02, SH09]. Recocido [CA05]. Recolección [Sie04, VO03]. Recolectando [KMGB07]. Recolectar [AGBP09]. Recomendaciones [CMC^{+}00, CHBK02, Ped03, QR05, RRK02, RPT00]. Recommendations [CMC^{+}00, CHBK02, Ped03, QR05, RRK02, RPT00]. Recommended [TOR^{+}04]. Reconciliación [Cze08, HH06a, SSSK06]. Reconciling [BSR04, BBB^{+}07, Cze08, HH06a, SSSK06, PZ01]. Reconsideración [SMN00]. Reconsidering [OS04]. Reconstructed [TNFH07]. Reconstruidos [TNFH07]. Recorded [Par03]. Recording [RFB09]. Records [CMA^{+}06, LE02]. Recoupling [FEKH09]. Recovered [AS01]. Recovering [ALR05]. Recovery [Bei03, Bei02b, BRYK01, CHBK02, CAFL03, Dun04, ERB01, GS01, HODB09, HP01, LCZ^{+}05, LCC^{+}07, MRM^{+}05, MB01, PPBW08, RM01, Ris02, SZBR03, SH03, Sem02, UWB04, VMBC02, Wal03, Wil03, YH01, VNP06]. Recreational [DGC05, GPZ04, GMB^{+}08, Lyn06, Lyn08]. Recreativa [Lyn06]. Recreativos [DGC05, GPZ04, GMB^{+}08, NHHG^{+}05]. Recruitment [CH01, NAC^{+}06, SH09]. Recuperación [ALR05, CHBK02, CAFL03, ERB01, GS01, HODB09, HP01, LCZ^{+}05, LCC^{+}07, PPBW08, RM01, SH03, Sem02, UWB04, VMBC02, VNP06, Wal03, WS03, Wil03, YH01, BRYK01, Dun04, MRM^{+}05, MB01, SZBR03]. Recuperadas [AS01]. Recurso [BF03]. Recursos [CMM^{+}05b, DBB^{+}09b, GK08, GEJK05, HSDS01, JMP09, LVBS^{+}04, MDH^{+}07, Nel09, PL03, SC00, TB01]. Red [BV04, dLBBG^{+}08, BVL^{+}03, But03, BAKHT06, CVS^{+}09, EGN^{+}05, FVS03, FMPB07, FH06, GHTMR01, GK06, Gus02, HDA01, Hay09b, HW08, HCRM09, HCZ00, HKM06, KB04a, KZSV05, KG03, MFGB07, OALM^{+}00, Put05, QRF07, TM04, TRNH00, TKB^{+}06, UKC^{+}09]. Red-Backed [GK06, TM04]. Red-Listed [Gus02, HW08, HKM06]. Redemption [Orr01b, Orr04a]. Redes [BBBM02, DLCV00, Dav07, JSWS00, JFS08, Kri04, LAN05, Nos03, RWJRR06, SJF^{+}05, SFT00, VSC06, BFMR07]. Redesigning [BTBC^{+}03]. Redford [Kni06]. Rediseño [BTBC^{+}03]. Reducción [Bur04, For00b, Ant08, ARP05, Ber07, CLŽ^{+}07, GBH^{+}09, MKK^{+}06, SS01]. Reducciones [KL06]. Reduce [BPB^{+}05, BRR^{+}02, For02, RAL^{+}04, Tal04, VT09, Wel06]. Reduced [ARP05, KF02, PZG^{+}05, SS01]. Reducen [Wel06]. Reducida [KF02, WSBP09]. Reducing [For00b]. Reducir [BPB^{+}05, BRR^{+}02, For02, RAL^{+}04, VT09]. Reduction [Ant08, BISS06, Bur04, CLŽ^{+}07, GBH^{+}09]. Redundancy [Ros02]. Redux [Cze02a, WABT07b]. Redwoods.edu [Stu01]. Reed [FCL00]. Reef [BMC^{+}09, BMP07, CMM05a, DAKB08, FMP^{+}05, GWJ^{+}07, KB03, Mas03, MA08, MBB^{+}08, O’C00, SCD^{+}08, TH03, WS03, WBSG07, FDL^{+}05, UWB04]. Reefs [DDN01, GMHS00, GGC03, HR04b, MCG^{+}09]. Reemplazo [RNL^{+}04]. Reencauchando [DPB^{+}09]. Reestablecida [FHL^{+}04]. Reestablished [FHL^{+}04, Mor00b]. Reestablishing [SMG00]. Reestablishment [MSB^{+}00]. Reference [JW02]. Referencia [EG07, GWS07, JW02, MU08, SCD^{+}08]. Refinación [HJP05]. Refinamiento [AKM07]. Refinando [JGL^{+}02]. Refinement [AKM07]. Refining [HJP05, Hun00, JGL^{+}02]. Reflected [BBH06]. Reflection [Büs08]. Reflections [Ter04]. Reflejada [BBH06]. Reforestation [BS03]. Reform [CAR05, Nel09, WPS08]. Reforma [WPS08, CAR05]. Reformas [Nel09]. Refreshing [SLC^{+}04]. Refuge [BHW01, Cze05, Gri00, KTN^{+}09, LLS00]. Refuges [ASF^{+}09, BKK03, SMFPCA05]. Refugia [OFN^{+}06]. Refugio [KTN^{+}09, SMFPCA05]. Refugios [ASF^{+}09, BKK03, Cze05, KWvH08, LLS00, OFN^{+}06]. Refuting [Moo02]. Regal [Wil02c]. Regan [Tod00]. Regarding [LRS09a]. Regeneración [AE02, Dun04, MB01]. Regenerating [AE02]. Regeneration [Dun04, MB01, NHHG^{+}05]. Regime [JMB01, KF01a, Mar02a]. Régimen [KF01a]. Regimenes [BM04, VGB^{+}04, HH06b, RR00, Rob00]. Regimes [Bea09a, BM04, HH06b, Ker09, Min01, RR00, Rob00, VGB^{+}04, Whe02]. Region [BMB^{+}08, Cua00, HCBK00, KPJ06, LSN07, ML04, MSOM^{+}06, RSD03, BMB^{+}08, KPJ06, MB00, RSD03, VDP00, Cua00, HCBK00, LSN07, MSOM^{+}06]. Regional [Ank04, CMP^{+}00, CGR^{+}05, CMS09, DOW^{+}06, DAKB08, DKS^{+}09, EGN^{+}05, ESB05, FMPB07, GHTMR01, GCV03, HMH00, HCZ00, KFSR^{+}06, KZSV05, LFR^{+}08, MRAF^{+}07, MWEB06, Qia07, SLS05, SHT^{+}05, TK06, VBOV05]. Regional-Scale [MWEB06]. Regionales [CMP^{+}00, CGR^{+}05, CMS09, DKS^{+}09, FMPB07, KZSV05, MRAF^{+}07, SHT^{+}05, VBOV05]. Regionally [Gut05, VMV07]. Regionalmente [Gut05, VMV07]. Regiones [AE01, ACMA04, GR03]. Regions [AE01, ACMA04, GR03]. Registers [PBSR06]. Registrar [RFB09]. Registro [JGL^{+}02]. Registros [CMA^{+}06, LE02, PBSR06]. Regla [Wan04]. Reglamentados [Pan02]. Reglas [TW06]. Regresando [PV09a]. Regresión [CB01b]. Regresiva [MPE^{+}07]. Regression [CB01b]. Regulación [LARR03, SKBB00]. Regulado [IF02]. Regular [PR08]. Regulated [IF02]. Regulation [LARR03]. Regulations [NPT05, SKBB00]. Regulatory [WL04]. Rehabilitación [EAA^{+}09, FMP^{+}05, LCZX09]. Rehabilitating [EAA^{+}09, LCZX09]. Rehabilitation [FMP^{+}05]. Reindeer [GD05]. Reino [EGN^{+}05, PK05]. Reintegrar [SMG00]. Reinterpreting [Gre09]. Reintroducción [CPSS03, GP08, JC07, MS04, ODE01, SAM07, VSH^{+}07, ZSSM09]. Reintroducciones [BDSD^{+}05, RSCL04, SL00a]. Reintroduced [AP00, AE02, ENN^{+}08, Ols07, SRR00, TR01]. Reintroducida [AP00, AE02, ENN^{+}08]. Reintroducidas [Ols07]. Reintroducidos [SRR00, TR01]. Reintroduction [Ano07e, CPSS03, GP08, JC07, MS04, MSB^{+}01, ODE01, SAM07, VSH^{+}07, ZSSM09]. Reintroductions [BDSD^{+}05, RSCL04, SL00a]. Reinvigorating [CNHT09]. Relación [BFD07, BVH^{+}02, Che09, EKL^{+}00, FE07, GMW08, KG03, NNM00, ØDEA00, Pod01, RHS05, SMC00, SH02, VBOV05, CGR01, KW02, Ree05, SSM03]. Relacionada [OBM09, ZGR02]. Relacionadas [LRS09a]. Relacionado [BZD03]. Relacionados [CWNW01, HMF03, MHM09, TG04]. Relacionando [DBML07]. Relaciones [CJ08, Fat06, MP07b, NFM04, ALR05, JP09, MNRD^{+}08, Qia07, SKBB00]. Related [CWNW01, HMF03, MHM09, OBM09, TG04, ZGR02]. Relatedness [GMW08]. Relating [DBML07, SSM03]. Relation [BFD07, BVH^{+}02, EKL^{+}00, KG03, Pod01, RHS05, SMC00, SH02, VBOV05]. Relationship [BRTW07, CGR01, Che09, KW02, NNM00, Ova03, Ree05]. Relationships [ALR05, CJ08, Fat06, JP09, MNRD^{+}08, MP07b, NFM04, Qia07, SKBB00]. Relativa [BSW^{+}06, Cro02, PKB^{+}08, VSAG06, YKIO08, YL05]. Relativas [LWV04, MAP03, MM04]. Relative [BSW^{+}06, Cro02, FE07, GM00b, LWV04, MAP03, MM04, PKB^{+}08, PCVC06, SCJ^{+}04, VSAG06, YKIO08, YL05]. Relativo [GM00b]. Relativos [PCVC06]. Releasable [Bau05]. Release [KLH08, MS04, TAP^{+}01]. Releasing [RSCL04, SL00a]. Relevance [CMK^{+}07]. Relevancia [CMK^{+}07]. Reliability [EVD01, LWV04, MDCK07, MAP03]. Reliable [MD04]. Relict [CCM00c, Gar03a]. Relicta [Gar03a]. Relicto [CCM00c]. Religion [Awo08]. Religious [RF00]. Relocation [PLB09]. Remain [Bra05a]. Remaining [GWM^{+}07, LJMR^{+}01, RJW09]. Remanente [CLC^{+}08, DW05b, GWM^{+}07, RJW09]. Remanentes [Chi00a, DBL06, LP08, LJMR^{+}01, MMLM01, NH09]. Remarkable [GWS^{+}06]. Remnant [CLC^{+}08, DW05b, LP08, NH09]. Remnants [Chi00a, DBL06, MMLM01]. Remoción [KFBG08, KBR03, RPM09, TPT^{+}07, VMBC02, GAF00]. Remota [BGD09, CHW^{+}01, HPK^{+}04, MNH01, PMKDS01]. Remote [BGD09, BNM^{+}09, CHW^{+}01, HPK^{+}04, MNH01, OFN^{+}06, PMKDS01]. Remote-Sensing [CHW^{+}01]. Remotos [OFN^{+}06]. Removal [GAF00, KFBG08, KBR03, RPM09, STT05, TPT^{+}07, VMBC02]. Removed [BSM08]. Removidas [BSM08]. Renacuajos [BRS^{+}05, BS00, SWR00]. Renewal [Ano02h]. Renos [GD05]. Renosterveld [DNZK02]. Rentabilidad [JMS06b, LADM05]. Rentable [BSW^{+}08]. Rentables [DJOW07, HM08]. Repatriación [MD04, Ber07]. Repatriating [MD04]. Repatriation [Ber07]. Repeated [UWB04]. Repelentes [STC03]. Repellents [STC03]. Repensando [Ber04c, XM07]. Repercusiones [Rue05]. Repetidos [UWB04]. Replacement [RNL^{+}04]. Replanteamiento [BE00]. Replicadas [KLH08]. Replicated [KLH08]. Reply [Ang01, Ano05c, ADD^{+}07, BS01a, BRK^{+}03, CCM00a, Col01, CD07, EVD01, Fra01, Hod04b, KC08, KDKS08, LMR^{+}07, MS02, Mcs06, MSB^{+}01, MDW01, ML00, OS04, PD00, PR01a, PS01b, PMEC04, RKR03b, SBS^{+}09, SF04, SRDG03, SBH00, Tod00, TM02, TSKD08, VW05, War00, WDL^{+}02, Wil09a, You01, KO01]. Repoblados [DN00]. Reportados [BM05b]. Reporte [FBC^{+}06]. Reported [BM05b]. Reportes [GBH^{+}05]. Reporting [FBC^{+}06]. Reports [GBH^{+}05]. Repositories [BB01]. Represas [HAG^{+}09]. Represent [WMVK03]. Representación [AHTM05, JFS08, KU02, TK06, WMVK03, BMP^{+}06]. Representan [HLSZ02]. Representation [AHTM05, JFS08, KU02]. Representativas [FDL^{+}05]. Representative [FDL^{+}05]. Representativeness [HGH^{+}00]. Representatividad [HGH^{+}00]. Representing [TK06]. Reprieve [STT05]. Reproducción [For02, HK00, KHM00, KBR03, LCS04, ODE01, PBN^{+}04, RG04a, RK05, SC00, SPM05, VTRG00, WR01, Wed02, YE07, BBO^{+}01, GP08, MBQ05]. Reproducen [CMCG07, CS07b, HRS08, PC02]. Reproducidas [FPF00]. Reproduction [HK00, PBN^{+}04, SC00, VTRG00, YE07]. Reproductiva [Cun00, DPTT^{+}05, HL07b, Kel01, YL05]. Reproductivas [ASD01, FPW01]. Reproductive [AAO^{+}07, BV01a, CQLF02, Cun00, DOGN08, DPTT^{+}05, DNZK02, EBDM06, FPW01, FLQ03, GMW08, HDQSL06, Kel01, LDG^{+}00, Mor00b, TSG06, WH01, YL05]. Reproductivo [AAO^{+}07, BV01a, CQLF02, DOGN08, DNZK02, FLQ03, HDQSL06, LDG^{+}00, Mor00b, SM07a, TSG06, WH01]. Reproductivos [FPL^{+}04]. Reproductoras [MR00b, MKDT01]. Reproductores [SS03b]. Reptile [TA08, VS09]. Reptiles [GB09, RGF^{+}04, Rod05, SB03, Sha04, TA08, VS09]. Republic [DC06b, IIG05, KFK^{+}09, Lat00]. República [DC06b, KFK^{+}09, Lat00]. Requerimientos [BOJ09, SLCW08]. Requested [Edi03]. Requiems [DTCC03]. Requieren [PBC^{+}09]. Required [GCM^{+}08]. Requirements [BOJ09, SLCW08]. Requires [CKF^{+}04]. Rescatar [SWK05]. Rescate [Lin03a, Bos07]. Rescue [Bos07, Lin03a, SWK05]. Research [AR09, Bau05, BFMM02, BPP05, Bre02b, BL07, BRR^{+}02, Cam08, Cam05, CMK^{+}07, DTCC03, EAR^{+}05, Far04, FBK00, GMCP09, GLV05, GOL^{+}07, HL07a, KNH07, KCR^{+}08, KK09, LB05, Mas03, MBH07, MS09, MCM07, OS01, PGD^{+}07, PPW^{+}07, Ray06a, SMRR02, SAKQ^{+}05, SGG07, SL00b, Spe04, TG05]. Researcher [Lat00]. Researchers [Hay03]. Resentment [CS07a]. Reserva [AKM07, BTBC^{+}03, DM04b, FNS^{+}03, Mad05, PH02, SBP07, VW04, HMF03, SMT07, TVD03, TAP^{+}01, VTA02]. Reservado [CJ08]. Reservados [Her09]. Reservas [AW02, AHTM05, BMC^{+}09, BMP07, CDD^{+}05, Dav05, FBP^{+}09, GR03, GWS07, Gui07, Gut05, HR02, HCZ00, HDK^{+}00, KFSR^{+}06, LB05, MC05b, MWEB06, MW07, Pan02, PKB^{+}08, PJK03, RPD01, RKEJP06, RBG^{+}04, STL02, SBP07, TW06, TWS06, WSS^{+}04, WU05, ZGR02]. Reservation [SBP07]. Reserve [AHTM05, AKM07, BMP07, BTBC^{+}03, CDD^{+}05, Dav05, DM04b, FBP^{+}09, FNS^{+}03, GR03, HMF03, HCZ00, KFSR^{+}06, Mad05, MC05b, MWEB06, MW07, PH02, PKB^{+}08, PC01, RKEJP06, SBP07, TVD03, TWS06, TAP^{+}01, VTA02, WSS^{+}04, WU05, SMT07]. Reserved [CJ08]. Reserves [AW02, BMC^{+}09, GWS07, Gui07, GNR01, Gut05, HR02, HDK^{+}00, LB05, Pan02, PZ01, PJK03, RPD01, RBG^{+}04, STL02, TW06]. Reservoirs [RAW07]. Reservorios [RAW07]. Resident [LF09, PFL01]. Residentes [LF09, Olu00, PFL01, QRF07]. Resilience [GMMPP08, PRMV06, SCC^{+}09, WS03]. Resiliencia [GMMPP08, SCC^{+}09, PRMV06]. Resistance [DH00a, RS02c, WS03]. Resistant [CMCG07]. Resistant-Kernel [CMCG07]. Resistencia [RS02c, DH00a, WS03]. Resistente [CMCG07]. Resolución [RKEJP06, HH06a]. Resolution [RKEJP06]. Resolving [HH06a]. Resource [App03, BF03, CMM^{+}05b, DBB^{+}09b, HSDS01, KT06, Nel09, PL03, Rai03, Roo00, Whi03]. Resources [Che02b, GL03, GK08, GEJK05, JMP09, LVBS^{+}04, MDH^{+}07, Pan03, SC00, TB01]. Respecting [Sch05]. Respond [MCG^{+}09]. Responder [MCG^{+}09]. Responsabilidades [SGB^{+}08]. Response [BRTW07, Bei02b, BM06, BC09, CNHT09, CPD^{+}03, Dal00b, Des06, DGG^{+}01, Dun03, EBDM06, FB00, FF08, GCM^{+}08, Gha08, GBMH07, GV05, GJLL08, HRN00, HM09, Her09, HH09, HC09, KF01b, KBR03, Lam02, LSF04, Lyn08, MBFT09, MP09, McD03b, MA05, MDJP04, MRH^{+}00, MC06, MM03b, NVP07, NB00, Orr05b, Orr05c, PMBN01, Pri01b, Ris02, RSC03, Say09, SSY06, Ste08a, Tay03b, Tem02, TA08, Tol07, VGB^{+}04, VDP00, WABT08, Wen08, Whe02, Wil00a, WSPL09, NL03b]. Responses [BZD03, CFH^{+}02, DW05b, GBC^{+}05, KD05, LMD05, LARK01, MHJ^{+}09, MW07, OHC08, Orr04d, RJW09, SJP^{+}07, SK02b, TPT^{+}07, VMM07, WD06]. Responsibilities [SGB^{+}08]. Responsibility [Cob05, Rol06a]. Respuesta [FB00, GV05, HC09, LARK01, MP09, TA08, Wen08, CPD^{+}03, GBMH07, GJLL08, Her09, KBR03, PMBN01, SSY06, TPT^{+}07, VGB^{+}04, VDP00]. Respuestas [CFH^{+}02, DW05b, GBC^{+}05, KD05, MW07, OHC08, VMM07, WD06, BZD03, LMD05, MHJ^{+}09, RJW09, SK02b]. Rest [Cur02b, JWB06]. Restablecidas [Mor00b]. Restablecimiento [FEKH09, MSB^{+}00, BRA^{+}03b]. Restauración [ASF^{+}09, Bra06, BGWE00, BC08, CWL^{+}00, CPSS03, CA06, DN00, GRB02, HWH^{+}02, HLG03, Kau04, LMS^{+}02, MH00, MK02b, RDW01, SWK05, SC05, TUE06, WRW^{+}02, BAF04]. Restaurados [ALR05, FVS03]. Restaurando [Hod04a]. Restaurar [ASD01, WL04]. Resto [JWB06]. Restoration [Ano07e, ASF^{+}09, BGWE00, BAF04, BC08, CWL^{+}00, CPSS03, CA06, Cur02b, DN00, GRB02, GALW09, HWH^{+}02, HLG03, Hol08, Kau04, KWV08a, Lam02, LMS^{+}02, LS04, MH00, McB02, PKM00, RDW01, Sad05, SWK05, SC05, SS02b, TUE06, WRW^{+}02]. Restore [ASD01, WL04]. Restored [ALR05, FVS03]. Restoring [Bra06, BRA^{+}03b, Hod04a, MK02b]. Restricciones [CWC00, KSH00]. Restricted [Cou02]. Restrictions [CWC00, KSH00]. Restringido [Cou02]. Result [HL07b]. Resultado [HL07b, KB03]. Resultados [BMC^{+}09, Bei08, KG03, MM04]. Resultantes [LHG^{+}03]. Resulting [GGI^{+}05]. Results [KG03]. Resurgence [CCS^{+}03]. Resurgimiento [CCS^{+}03]. Resurrection [Yul06]. Resurvey [JLZW03]. Retención [HNG04, HKM06]. Retention [HNG04, HKM06]. Rethinking [BE00, Ber04c, SMN00, XM07]. Retiring [RS03]. Reto [MPE^{+}07]. Retos [LCC^{+}07, MB07, MKVZ^{+}09, Rod05, CVB^{+}09, HH06b, TPS^{+}05]. Retreading [DPB^{+}09]. Retroalimentaciones [LW01]. Retrospect [Orr09b]. Retrospective [Joh01]. Return [Bow09, LRS09a]. Reubicación [PLB09]. Reveal [GN00]. Revealed [JLF09]. Revelada [JLF09]. Revelan [GN00]. Reverence [Ano02h]. Reversal [CW09]. Reversión [CW09]. Review [ACC^{+}05, CTR02, Fle01, HP01, JTZ^{+}08, MOO09, NMT^{+}04, PAA^{+}08, PR08, PS06, RRK02, SGB^{+}08, SMSB05, TF00]. Reviewed [BADF^{+}05]. Reviews [Ano00a, Fie03b]. Revisada [BADF^{+}05]. Revisado [PM07]. Revisión [CTR02, DH00b, MOO09, PAA^{+}08, PS06, SGB^{+}08, SMSB05, ACC^{+}05, JTZ^{+}08, NMT^{+}04, RRK02, TF00]. Revisiones [HP01, MD04]. Revisions [MD04]. Revisita [Kér04]. Revisitation [Kér04]. Revisited [BCD^{+}08, Nas06, PM07, Por08, TM04]. Revisiting [BGHMb07, SMW^{+}07]. Revival [Mei01]. Revolution [Rol06b, Vit04]. Revolutionize [BBC^{+}04]. Rewilding [FEKH09]. Rewriting [Orr01c]. Rhetoric [Rol06a]. Rhinoceros [Lin03b]. Rhinopithecus [GJLL08]. Ribereña [TS06]. Ribereñas [CS07b, FPSD09, FRR07, RBRB06]. Ribereños [JTS06, MM07]. Rica [BD06, CWC00, DCP^{+}03, LLS00, SAPRB07, SLB^{+}07, SSY06, TSP^{+}08]. Rican [CWC00, SLB^{+}07]. Rice [Elp00]. Richness [BMMR03, BSS^{+}05, CKL^{+}00, CDD^{+}05, DBL06, FTN^{+}05, HH00, HFR07, JMS06b, LBG03, aARLG03, LMP02, MDCK07, NF04, NBM^{+}04, ØDEA00, PM07, Qia07, RPP^{+}00, SKBB00, SD03, SDJK04, VSGS02, WMM04, WHJJ04, WMP05]. Ricketts [SC09]. Rico [DAR08, Sta01]. Rides [Kau04]. Ridges [WE07]. Ridley [LCS03]. Riesgo [AKO03, ACC^{+}05, ACMA04, AOGMS09, BJE^{+}05, CB01a, Cho06, DDAK04, ES08, Fai08, FEGB09, FRR07, GZK06, HODB09, HP08, HK03, HLSZ02, KDL07, LW02a, LE02, LRG^{+}02, LARR03, MCG^{+}08, MBQAM06, MKK^{+}06, MH08, NLCAR09, OK00, PJNA07, PV09b, QR05, RBM^{+}05, SK04, SM09, SWB^{+}08, VT09, VSAG06, VWQI00, WPS08, WSCP05, WMS^{+}01, WSBP09, WTR01, WG08, GMMPP08, KKHS05]. Riesgos [FKU05, GM08a, HRS08, OBK^{+}04, PCVC06, SFT00, STS05, TTB03]. Right [Arm02, KMGB07, Orr06b, Rei05a, REC^{+}00]. Rights [BISS06, Hut08, MC09, PP08b, PP08a]. Rigueur [Lei09]. Rinocerontes [Lin03b]. Río [FKU05, LCZX09, MO04, VTRG00, KFK^{+}09, AT05, AAO^{+}07, BRYK01, FRD^{+}09, FMT^{+}04, HM08, IF02, JAH^{+}05, KBR03, RFD01, TS06, Wil03]. Ríos [Cas03, HAG^{+}09, PA08, PW01, RR00]. Riparian [CS07b, FPSD09, FRR07, GG02, HM04, HNG04, JTS06, LP08, MM07, RBRB06, RR00, SSM03, SB03, TS06]. Riparias [GG02]. Riparios [HM04, HNG04, LP08, RR00, SSM03, SB03]. Riqueza [BMMR03, BSS^{+}05, CKL^{+}00, CDD^{+}05, DBL06, FTN^{+}05, HH00, HFR07, JMS06b, LBG03, aARLG03, LMP02, MDCK07, NF04, ØDEA00, PM07, Qia07, RPP^{+}00, SKBB00, SD03, SDJK04, VSGS02, WMM04, WHJJ04, WSA^{+}05, WMP05, NBM^{+}04]. Rise [FCG^{+}05, Lei05]. Risebrough [Bei02b]. Rising [Bye04]. Risk [AKO03, ACMA04, AOGMS09, CB01a, Cho06, DDAK04, ES08, Fai08, FEGB09, FRR07, GZK06, HODB09, HP08, HK03, HLSZ02, KKHS05, KDL07, LW02a, LE02, LRG^{+}02, LARR03, MCG^{+}08, MBQAM06, MKK^{+}06, MH08, NLCAR09, OK00, PJNA07, PV09b, QR05, RBM^{+}05, SK04, SM09, SFT00, STS05, SWB^{+}08, VT09, VSAG06, VWQI00, WPS08, WSCP05, WMS^{+}01, WSBP09, WTR01, WG08, GMMPP08]. Risk-Assessment [DDAK04]. Risk-Based [STS05]. Risks [ACC^{+}05, BJE^{+}05, FKU05, GM08a, HRS08, OBK^{+}04, PCVC06, SAS^{+}09, TTB03]. River [BRV06, BRYK01, FKU05, FMT^{+}04, HAG^{+}09, IF02, KBR03, LCZX09, NHHG^{+}05, TS06, VTRG00, AAO^{+}07, GALW09, HM08, HM09, RFD01, Wil03, WSPL09]. Rivers [Cas03, HAG^{+}09, KG02, PA08, PW01, RR00]. Road [BGC^{+}04, BMS08, CF01, FB00, For00a, FD00, GS02, GS05, HR04a, JF04, LSL04, MSWF00, SRH^{+}07, Urb06b, VS05b]. Road-Effect [FD00]. Roadless [SD01, Tur06]. Roads [Cas03, DS01, GB03, Has00, HR04a, JSWS00, KGM^{+}02, LCT^{+}06, LCG^{+}08, MMGB05, PJ00, Sei04, SRH^{+}07, SG04, TF00, WCPB03, WSR^{+}00]. Roadside [RDW01]. Roadsides [CM06, PA04]. Rob [FB08]. Robin [AE02]. Robins [DA07]. Robinson [ABE01, DGSS03, DGSS03]. Roble [Bra06, FMLL06, MBM04]. Robles [BRA08, PAHM04]. Robust [Pov01]. Roca [GN00, KL06]. Rocas [ML02]. Rock [GN00, KL06, ML02, MNL03]. Rock-Climbing [KL06]. Rocky [GFTB04]. Rocosas [GFTB04]. Rodent [BVH^{+}02, BJB06, HDG^{+}07]. Roebuck [PS01b]. Roedores [BVH^{+}02, BJB06, HDG^{+}07]. Roesel [BBK02, BBK02, Ber05]. roeseli [BBK02, Ber05]. Roff [SRSP01, ZR01]. Roja [BV04, dLBBG^{+}08, Gus02, Hay09b, HW08, HKM06, KG03, QRF07]. Rojas [FMPB07, KB04a, KZSV05]. Rojo [EGN^{+}05, FVS03, HDA01, OALM^{+}00, TRNH00]. Rojos [FH06]. Role [ASF^{+}09, Awo08, BRA08, CM06, DMDFV08, DM04a, Far08, Fer01, FCL00, HHSC00, HBA09, KEJ06, KFA^{+}03, MDF05, Man02, Mer04, MRH^{+}00, MAM^{+}03, MCM07, MN07, OWOF03, PJ00, Rai03, RNL^{+}04, Rya06, SZS01, SSL06, TR01, War05, Wes03, WSA^{+}05]. Roles [Bee00, FRSG^{+}02, HN09]. Roll [Mei06]. Rome [Cze06]. Romero [QAMW03]. Rominger [MC06]. Room [McD03a]. Roosevelt [Mei01]. Roosts [Hol08, KWvH08, KWV08a]. Roots [Mai04]. Rootworm [MMLM01]. Rosaceae [LLW02]. Rosemary [QAMW03]. Rostrhamus [MKH07]. Rotación [Ram01]. Rotting [SLS05]. Roundworm [Log03]. Routes [MR00a]. roxellana [GJLL08]. Roza [NTMT^{+}03]. Rubinoff [Des06]. Ruderales [RDW01]. Ruedas [DPB^{+}09]. rufus [RSF^{+}03]. Ruido [BHB08]. Ruin [MF00]. Rukwa [MCM07]. Rule [Sad05, Wan04]. Rules [TW06]. Rural [BL01, MKG03, Mcs06, NTGT03, RHS05, RDEF01]. Rusia [HPK^{+}04]. Russia [HPK^{+}04]. Rutas [MR00a]. Rutidosis [Mor00b].

S. [HMF^{+}04]. Sabah [SLELA05]. Sabana [FBL^{+}09b, MBMLJ09, VGB^{+}04]. Sabanas [Bra06, VS09]. Sabiendo [KCR^{+}08]. Sacramento [GALW09]. Sacred [BKWB05, DHZM09, Pea01]. Safe [Ano00d]. Safe-Fail [Ano00d]. Safeguarding [BWS09, HTM^{+}05]. Sage [BRA^{+}03b, CKL^{+}00, Rub01, WRW^{+}02, WWR^{+}02]. Sage-Grouse [WRW^{+}02, WWR^{+}02]. Sagebrush [HWH^{+}02, WRW^{+}02]. Sagebrush-Steppe [HWH^{+}02, WRW^{+}02]. Sagrados [BKWB05, DHZM09]. Sahara [MBB^{+}03]. Saharan [MBB^{+}03]. Saint [VDHD01, VDHD01]. Saint-Brieuc [VDHD01, VDHD01]. Sake [Vit04]. Salafsky [BCK^{+}09]. Salamander [GK06, GGW09, KHHK03, LB02, MNB09, MMGB05, Pyk05, SS08a, SRH^{+}07, TSG06]. Salamanders [CS07b, WD01, WD03]. Salamandra [SS08a, TSG06, SFS05]. Salamandras [CS07b, KHHK03, LB02, MNB09, MMGB05, SRH^{+}07, WD01, WD03]. Salmon [BPB07, FHB^{+}08, GL03, GWM^{+}07, HM08, HM09, HLG03, LW02a, PPBW08, PSGA09, Ril00, VTRG00, Wil03, WSPL09, ZSMW06, BPB07, FHB^{+}08, GWM^{+}07, HM08, HLG03, LW02a, PPBW08, PSGA09, Wil03]. Salmones [VTRG00]. Salmónidos [BRBMS06, FRD^{+}09]. Salmonids [BRBMS06, FRD^{+}09]. Salomon [ZR01]. Salt [BCS04, BS08, BCHS08, HBA09, SB04]. Saltcedar [NHHG^{+}05]. Salud [DKW01, MK02b, PK01]. Salvage [FO06, Hut06, LN06, LO06, NL06, SSSK06]. Salvaging [Tho09]. Salvaguardar [BWS09]. Salvajes [KB03, LADM05]. Salvamento [FO06, Hut06, LN06, LO06, SSSK06]. Salvar [SRC^{+}02, UKC^{+}09]. Salvelinus [PD03a]. Salvia [BRA^{+}03b, CKL^{+}00, Rub01]. saman [CQLF02]. Samanea [CQLF02]. Same [Kri05]. Sample [Kai03, Wal00]. Sampling [Cla09, GBE^{+}06, HW08, MM04, NFM04]. Samuel [BM03]. San [KBR03, MBR^{+}02, KBR03]. Sanción [MW07]. Sanctuary [BKV^{+}05]. Sandhills [aARLG03]. Sangihe [SMI^{+}07]. Santa [RPM09, STE^{+}01, RPM09]. Santuario [BKV^{+}05]. Sapo [BR01]. Sapos [PBS03]. Saproxílicos [HW08]. Saproxylic [HW08]. Saqueo [WTEH^{+}01]. Saskatchewan [HBV02, SMC00]. Satélite [HPK^{+}04, MHC^{+}01]. Satellite [HPK^{+}04, MHC^{+}01]. Saturniid [BEB00]. Saturniidae [BEB00]. Sauces [SK02b]. Saudi [HMF^{+}01]. Savanna [FBL^{+}09b, MBMLJ09, VGB^{+}04]. Savannas [Bra06, VS09]. Save [Kni07, SRC^{+}02, SBM06, UKC^{+}09]. Saving [HRP^{+}04, Hoo07, JLRBB^{+}07, PJB^{+}06, RA09]. Says [BADF^{+}05]. SBC [CHBK02]. Scaffolds [McB02]. Scalar [DAB06]. Scale [AW02, Arm05, BV01a, Cha06, FDL^{+}05, Fje07, FNFM01, GC07, GSR^{+}04, GM08a, HH00, HK03, HLB^{+}03, HBV02, JP08, Kai03, KDD^{+}04, LB02, MBR^{+}02, MDCK07, MWEB06, MO04, NFM04, Qia07, RCL^{+}06, SRW^{+}08, SP00, UWB04, WSS^{+}04]. Scale-Dependent [GC07, HH00]. Scales [CPD^{+}03, GCV03]. Scanning [Tuc04]. scapigera [KDD02]. Scarce [CWNW01]. Scattered [GLF^{+}08]. SCB [Bei03, Büs08, CNHT09, CHBK02, Gra09, VBRNP08]. Scenario [BRV06, Ibi08, PCC03]. Scenarios [Fai08]. Schemes [KBS^{+}04, SSB^{+}08, ZLM03]. Schmiegelow [LMR^{+}07]. Schoolyard [Bre02a]. Schurbon [MDJP04, RO04]. Schwartz [CNHT09, PR01a]. Science [AO06, And05, Ano02e, ACD^{+}07, AFK^{+}09, BPV09, BHW01, BV00, Büs08, Cle06b, Cor04a, Cri00, Cur02a, Cur07, Cze07a, Da 03, Dav05, DN01, DWWF04, EAR^{+}05, FCN^{+}06, FN02b, GT03, GC09, HTU^{+}06, KBMN00, Lac07, LW06, Lin03b, MB07, MCF^{+}09, Mas03, Mcd04, Mcs05, Moo02, ML00, Nel06, NPT05, Nos00, Orr04a, Pan03, PBC^{+}09, QR06, Rei05b, RVNM08, Sag07, SMRR02, SRL^{+}07, SAM07, SM01, Tur06, Tut03, Wes03, Win04, Wol06, Wri07, ZMS^{+}08, Mef07a]. Science-Based [BPV09, Cur07]. Sciences [MBD^{+}03, Sar06, EYZ09]. Scientific [DWW04, Fle01, HCE^{+}01, KK09, LG05b, Moo02, NFD^{+}09, Pro04, RPT00, RSPM05, SHB08, Ste05, Van06]. Scientism [Har03]. Scientist [Ter04]. Scientists [Lac07, LJI^{+}08, MN07, NV09, Sch06, Tho03]. Sclerosis [MBC03]. scoparius [ME00]. Scratching [KO01]. Screening [DDAK04]. Scrub [BRA^{+}03b, CKL^{+}00, DW05a, GBC^{+}04, MSWF00, QAMW03, Rub01]. Scrub-Jay [DW05a, MSWF00]. Scuba [GPZ04]. Se [Sim03, XGS^{+}09, CMCG07, CS07b, HRS08, PC02, MF00]. Sea [FVS03, FCG^{+}05, Fra09, GSG08, HY03, LC07a, MBR^{+}02, Por08, SS02a, SVBR07, UCAGG08, VBRNP08, FMSL08, Wol00]. Sea-Cage [Fra09]. Sea-Level [FCG^{+}05]. Seabed [Kai03]. Seabird [BM05a, BB06]. Seabirds [JTZ^{+}08]. Seafloor [LARK01]. Seaflower [FNS^{+}03, FNS^{+}03]. Seafood [CMC03, KL05]. Seagrass [MHJ^{+}09, RLK^{+}01]. Seahorses [FV05, GPZ04]. Seal [JML00, KGM01, BM03]. Searching [Kaw04, LMP02]. Seas [Bee09, Tay03b]. Season [SK02b]. Seasonal [Lai05, PB04]. Sección [CVB^{+}09]. Seco [CWL^{+}00, CQLF02, FLQ03, HDQSL06, SOADMR08]. Secondary [CPD^{+}09, GRJB^{+}07, HP04, PBP09, STC03]. Secos [ANW^{+}08]. Secreción [WBW06]. Secretion [WBW06]. Section [BDRC05, CVB^{+}09, DW06, LGB05, Mef06a, NL06, RP07]. Sections [CNHT09]. Sector [BHB08, RANM08]. Secundario [HP04]. Secundarios [CPD^{+}09, GRJB^{+}07, MNA07, PBP09, STC03]. Security [Ano02j, Ben02, Har07]. Sedentarization [LWL09]. Sedimentación [MMF^{+}05]. Sedimentation [MMF^{+}05]. Seed [GGGS00, GC07, Gor08, GAF00, HLJV07, KWvH08, KFBG08, NH09, PS00a, Sch07, TO02, TR05, WZD^{+}00]. Seed-Plant [HLJV07]. Seed-Rain [Gor08]. Seeded [KFBG08]. Seeding [Bey04]. Seeps [HMH00]. Segment [IF02]. Segmento [IF02]. Segments [PD00]. Segunda [TM04]. Seigel [TM02]. Seis [KDD^{+}04, VSGS02]. Seldom [Ste08b]. Selección [AHTM05, AKM07, BV01a, CM03, DFS05, For02, FHB^{+}08, FBP^{+}09, GMMPP08, GSG08, HLG03, KCR^{+}08, MC05b, MWEB06, MW07, PKB^{+}08, PMC01, RG04a, SJF^{+}05, STL02, SDL07, Vee00, WSS^{+}04, WSQP07, FMT^{+}04, GR03, PLB09, PSGA09, PBL00, Wed02]. Seleccionados [PDF^{+}00]. Seleccionar [BMBM06, DDAK04, RBG^{+}04]. Selecciones [PCL08]. Select [RBG^{+}04]. Selecting [BMBM06, KCR^{+}08]. Selection [AHTM05, AKM07, BV01a, CM03, CWH07, DFS05, FMT^{+}04, For02, FHB^{+}08, FBP^{+}09, GMMPP08, GR03, GSG08, HLG03, KC07, MC05b, MWEB06, MW07, PKB^{+}08, PLB09, PSGA09, PMC01, PBL00, PC01, SJF^{+}05, STL02, SDL07, WSS^{+}04, Wed02, WSQP07]. Selectiva [CWH07, Lew01, MNA07, PMZ06]. Selectivamente [GAF00, RR01]. Selective [Lew01, MNA07, PMZ06, WLCS00]. Selectively [GAF00, RR01]. Selectivo [WLCS00]. Self [AFK^{+}09, Col00, GPMM08, LDG^{+}00, MRP07, PIH00, SP09, VJ07]. Self-Compatibility [VJ07]. Self-Determination [Col00, SP09]. Self-Guided [PIH00]. Self-Incompatible [GPMM08, LDG^{+}00, MRP07]. Self-Interest [AFK^{+}09]. Selfless [Rom07]. Selva [BMMR03, ESSGL00, GGGS00, Lim06, MP07b, MO04, NBM^{+}04, SSY06, WA05]. Selvas [MEA00]. Semana [BGC^{+}04]. Semi [SP00]. Semi-Natural [SP00]. Semi-naturales [SP00]. Semiacuáticos [FPSD09]. Semiaquatic [FPSD09]. Semiarid [CCM00c, GB03]. Semiárido [CCM00c, GB03]. Semillas [GGGS00, GC07, GAF00, HLJV07, KWvH08, KFBG08, NH09, PS00a, TO02, TR05, WZD^{+}00]. Seminatural [DBL06, Elp00, VSGS02]. Seminaturales [DBL06, Elp00, VSGS02]. Semiurban [RWM^{+}06]. Semiurbanos [RWM^{+}06]. Señalización [RG04a]. Senderos [DGC05, PIH00, NHHG^{+}05]. Sensationalizing [BBM07]. Sense [Roo00]. Sensibilidad [BMP^{+}06, BFMM02, CB01b, ED07, JP09, NLCAR09, SLCW08, Cro02, RKEJP06, WSS^{+}04, ZG05]. Sensing [BGD09, BNM^{+}09, CHW^{+}01, MNH01, PMKDS01]. Sensitividad [SKP00]. Sensitivities [Cro02]. Sensitivity [BMP^{+}06, BFMM02, CB01b, ED07, JP09, NLCAR09, RKEJP06, SLCW08, SKP00, WSS^{+}04, ZG05]. Separating [Bal00]. September [Ano04i, Ano04k]. Septentrional [REC^{+}00]. Sequel [VP04]. Sequía [GSB^{+}06, LW01, LSC07, WLO^{+}00]. Sequoia [Rig01]. Ser [CJ08, DP01]. Seral [RWS02]. Serengeti [CMK^{+}07, DBML07, KD00, Kel01, SMH^{+}07]. Serengueti [Kel01]. Serial [GCM^{+}08]. Series [LWV04]. Serpentina [WH01]. Serpentine [HMH00, WHH00, WH01]. Serpentinos [HMH00]. Serpiente [GN00, SFD09, SS02a]. Serpientes [FBL^{+}09b, PBS03, RS02a, ZJ04]. Serve [Fje07]. Service [ACD^{+}07, Gha07, Gha08, KDKS08, NPT05, Tur06, Yul06, KB08]. Services [BRV06, HH08, HH09, HAG^{+}09, JLF09, KJSH08, MB09a, MBB^{+}08, RA09, Sal05, SAPRB07, SMM09, Wun07]. Servicio [Tur06, KB08]. Servicios [HH08, HAG^{+}09, JLF09, MHGF04, MBB^{+}08, SAPRB07, SMM09, Wun07, BRV06, KJSH08]. Servidumbre [BR05]. Servir [Fje07]. Sesgada [MBQ05]. Sesgo [AW01]. Sesgos [MF00, TB04, SGG07]. Set [Che09, Her09, JLF09, KB05, LV04, PEC01, VW04, VW05]. Set-Aside [Her09, KB05, VW04, VW05]. Sets [RC00, Tod00]. Setting [BGRJ00, BPL^{+}01, HCM03, MWBP07, RRCOM^{+}06, SSK^{+}05]. Settlement [Rem03]. Severidad [ET03, JTZ^{+}08, OFS^{+}04]. Severity [ET03, JTZ^{+}08, OFS^{+}04]. Sex [CW09, FNP09, GS05, JRL03]. Sex-Ratio [FNP09]. Sex-Specific [JRL03]. Sexo [CW09, JRL03]. Sexos [FNP09, GS05]. Sexual [Wed02, YE07]. Shad [BGWE00]. Shade [RCP00, SMFPCA05]. Shade-Coffee [SMFPCA05]. Shadow [Moo07]. Shakes [GX09]. Shape [ED07, YKIO08]. Shared [Juk01]. Shark [CMA^{+}06, Cor02, SCP^{+}02, WELCH09]. Sharks [FMSL08]. Sharp [ARMH04]. Sharp-Tailed [ARMH04]. Shearing [SVV07]. Sheep [Ber06, EMW^{+}04, ODE01, RBG06, SZS01]. Sheepshead [RNL^{+}04]. Shelden [Tay03b]. Shelf [O’C05, Orr09c]. Shell [VDHD01]. Shellfish [ES04]. Shibboleths [RRA06]. Shifting [Cor04a, NSW08, Nor00, Ram01]. Shifts [DMG^{+}00, DDN01, LHPH06, SSFL08]. Ship [APS07, PS01b]. Shore [VSAG06]. Shoreline [SB04]. Shortcomings [CD07, VO03]. Shortfall [RSB05]. Should [ALH^{+}04, AR09, GMMPP08, Law06, LA00, TSH06]. Shrimp [HGH^{+}03]. Shrinking [Rud06]. Shrub [EBAW01, MBQAM06, SK02b, YE07, YL05]. Shrub-Willow [SK02b]. Shrublands [DNZK02, KF01a]. Shrubs [CBH01]. Shrubsteppe [VDP00]. si [AP00, GM01, WABT07a, NV09]. Sichuan [GJLL08]. Side [MNA07]. Siembra [CLC^{+}08, KPJ06, Bey04]. Siempre [DP01]. SIG [AOGMS09, WGGS07]. Sighting [MRDC06]. Siglo [Rom03]. Signaling [RG04a]. Significado [MLMB08, VNP06, Wol06, MRDC06, MD04]. Significance [CT07, FBC^{+}06, MRDC06, MD04]. Significancia [CT07, FBC^{+}06]. Significant [DGG^{+}01, WABT07a, Gre09]. Significativa [WABT07a]. Significativas [Hut06]. Silence [Bee09]. Silenciosa [Fon09]. Silene [JTS06, VSAG06]. Silent [Fon09]. Silky [WELCH09]. Silphium [Par03]. Silvery [AT05, AMF^{+}01]. Silvestre [BKV^{+}05, BM05b, BL01, BD05, CW00, CWCG02, CHFR09, CMR05, Cze05, DC06b, DKW01, DGP^{+}02, FYC00, FPM02, LLS00, Mad05, MBL02, MNRD^{+}08, MDH^{+}07, MOO09, NTMT^{+}03, NKBD02, OFN^{+}06, PBSR06, PLL^{+}05, PCME09, RHA^{+}05, RAL^{+}04, SRW^{+}08, Slo02, TWW09, UMT^{+}01, WSA^{+}05, WSBP09]. Silvestres [BMS08, Cre01a, CB01b, For02, Fra09, FEKH09, HGH^{+}00, LEPH02, LCZ^{+}05, MCA^{+}00, Nos03, OBM09, PJNA07, RSF^{+}03, SRR00, SMFPCA05, TTSP04, TAP^{+}01]. Silvícolas [KHHK03, Pyk04]. Silvicultural [BMM^{+}01, DWN01, KHHK03]. Silviculturales [BMM^{+}01, DWN01]. Similares [CGR^{+}05]. Similarity [SHT^{+}05, SDJK04]. Similitud [SHT^{+}05, SDJK04]. Simon [Cze02a]. Simpátricas [BOC00, Fox07]. Simple [Fla00, HLB^{+}03, MF00, WZW^{+}06]. Simples [MF00]. Simulación [TCSD00, WHS02, WSLW00]. Simulada [WSQP07]. Simulado [CA05]. Simulated [CA05, WSQP07]. Simulating [WSLW00]. Simulation [TCSD00, WHS02]. sin [DN00, FDL^{+}05, LB05, Ric07, SD01, TPW03, Tur06, WG08]. Since [FJS04]. Sincrónica [WBDP02]. sinensis [HRS^{+}07]. Sinergéticas [BBB08]. Sinérgicas [Coc01]. Sinergismo [PRL^{+}02]. Sinergismos [LU09, WCP01, Hob01]. Sinergistas [Per01]. Singapore [CSS00]. Singapur [CSS00]. Single [CKL^{+}00, FV05]. Singular [Hei09]. Sink [FCL00, LW02b, Mur01, NRB00]. Sinking [BH04]. Sinks [Fau01, NFB05]. Síntesis [DWWF04, HM08, WD06, Les05]. Sinuosos [RR00]. Sir [SF00]. Siskiyou [SD01, SD01]. Sistema [Cze05, DDAK04, For00a, HH08, HCZ00, HDK^{+}00, JHHL04, KFSR^{+}06, LS09, LM05, LVM02, LRG^{+}02, MCG^{+}08, MK02a, RRK02, TR05, WU05]. Sistemas [BMP07, CPR05, EBAW01, KSDBT02, NRB00, OBK^{+}04, SM01, SBP07, LRO^{+}08]. Sistemática [PS06, RB07, SCB^{+}01, WSS^{+}04]. Sistemáticas [KDC^{+}06]. Sisyphus [Orr02b]. Site [BC08, CM03, HSR05, Lam04, ME00, PLB09]. Site-Selection [CM03]. Sites [AKM07, DHZM09, LSN07, LH04, MHC^{+}01, NCVBW02, TK07]. Sitio [BAF04, Dav07, Lam04]. Sitios [AKM07, CM03, CGR^{+}05, DHZM09, Fat06, Fon09, HSR05, HBD^{+}09, JB06, LSN07, MLN^{+}06, MCA^{+}00, MHC^{+}01, NCVBW02, PLB09, PSGA09, TK07, WBW05, MSBOF^{+}09]. sito [PRL03]. Situ [CMGSM00, GG06, SPFF^{+}08]. Six [KDD^{+}04, VSGS02]. Sixteenth [Rom03]. Sixteenth-Century [Rom03]. Size [AT05, BST^{+}05, Ber05, BPL^{+}01, CB01a, CWNW01, Chi00a, CNC04, DAR08, DM04b, DW05b, EKL^{+}00, ED07, FPF00, FV05, HRS08, HP08, HHSC00, HK00, Kai03, KW02, Kel01, LP07, LDG^{+}00, MBQ05, MM04, MK09, PH02, PM07, PV06, Pod01, RLC^{+}00, Ree05, SRR00, SFD09, SM07a, SC05, SLCJ08, SZS01, SKBB00, Vic02, Wal00, WH01, YE07]. Size- [MBQ05]. Sized [OK01a]. Sizes [Ano04e, Ova03, ROB^{+}04]. Skates [DMG^{+}00, DR02]. Skeptical [Moo02]. Skink [TPW03]. Skinks [BTGS05]. Slash [NTMT^{+}03, Ram01]. Slash-and-Burn [NTMT^{+}03, Ram01]. Slaying [Joh03a]. Slipper [VDHD01]. Slow [KHP04, McC00, UWB04]. Slow-Growing [KHP04]. Small [ARPEBM09, BV04, BCS00, CKL^{+}00, DLCV00, FTC04a, FNP09, GNR01, JP08, KFBG08, Log03, MBR^{+}02, MR00a, MO04, Pan02, RLC^{+}00, RSJE06, Rya06, VSAG06, WSBP09]. Small-Mammal [MR00a]. Small-Scale [JP08, MBR^{+}02, MO04]. Small-Seeded [KFBG08]. Smallest [HJ05]. Smallholders [CN06]. Smelt [MSOM^{+}06]. Smoke [DWWF04]. Smoky [SFS00]. Snag [Hut06]. Snag-Management [Hut06]. Snail [MKH07, MNL03, SLELA05]. Snails [HNG04, MMF^{+}05, SF05, Urb06b]. Snake [FBL^{+}09b, GN00, PV09b, SFD09, WBBA03, ZJ04, HM08, HM09, Wil03, WSPL09]. Snakes [PBS03, RS02a, SS02a, WBDP02]. Snorkeling [PIH00]. Snow [Cor03, MAM^{+}03]. Snowmobile [CFH^{+}02]. Snub [GJLL08]. Snub-Nosed [GJLL08]. Soberón [CD07]. Sobre [BB04, BBK02, BTGS05, BEB00, CQLF02, DGP^{+}02, FCG^{+}05, GS02, Gra01, HH00, KD05, LBG03, ML02, MO04, PSBG05, Pra05, RGF^{+}04, Say05, SP00, TNRJ02, VSC06, WD01, WHH00, AKO03, ASM05, AT05, AGBP09, ASD01, ANW^{+}08, BJM04, BBB08, BGC^{+}04, BHB08, BM05a, BS09, Bei08, BD06, BMMR03, Ber06, Ber05, Bey04, BFMM02, BBM04, Bor08, BRBMS06, Bra06, BBW^{+}07, BJE^{+}05, BTS09, Cal06, CWH07, CF01, CPSS03, CCC04, Cho06, CMM^{+}05b, CY08, CPO03, CNM01, Cua00, CLC^{+}08, DBL06, DOGN08, Dav07, DDK05, DS01, DJC^{+}07, DDBA^{+}07, DNZK02, DP01, EM03, EGPS04, EMW^{+}04, ED07, FMDK09, FH05, FDL^{+}05, FPSD09, FBK00, FFHW03, FMLL06, FBL^{+}09b, FLQ03, GC07, GK06, GMW08, GMA07, GWJ^{+}07, GPC^{+}03, GSB04, GNEE07, GVTM07, GFTB04, HMB^{+}05, HP04, HIS03, Has00]. sobre [HR04b, HH03, HCRM09, HLG03, HRW^{+}03, HVR08, HAG^{+}09, HFR07, HF04, HNG04, IMD01, IML^{+}05, JF04, JC07, JTZ^{+}08, JFS08, Kai03, KHRS01, Kee06, KWvH08, KGM^{+}02, KS03, KFBG08, KSDBT02, KHHK03, Koe03, Kri04, KL06, KA03, Lai00, LSL04, LCT^{+}06, LCG^{+}08, LBPBG07, LLWCW09, LFR^{+}08, LH04, LCS03, LO06, LAN05, LRS^{+}09b, Lus03, MR00a, ML04, MM07, Mar05, MBQAM06, MMLM01, MYFF^{+}07, MMF^{+}05, MJ03, MY02, MK09, NA01, NFM04, NAC^{+}06, NTMT^{+}03, PA04, PGP04, Per00, Per01, PMZ06, Phi00, PBS03, PC07, PDS^{+}07, Pyk00, Pyk04, PM05, PTP^{+}08, RO08, Rem03, RSF^{+}03, RBK^{+}05, RS02c, SRW06, SMM09, SLELA05, SG01, SF03, SLCJ08, Shi06, SLS05, Sim03, SLCW08, ST05, SG04, SBP07, SMT07, SOADMR08, SS03b, TG05, TS06, TH03, TSP^{+}08, Tow03, TR05]. sobre [VCHBP08, VS09, VDHD01, VJ07, Vic02, WMM04, WSS^{+}04, WCPB03, Wed02, WG06, WHJJ04, WRM05, WGGS07, WLCS00, WD03, WGK07, WRW^{+}02, XGS^{+}09, NV09, Sag07]. Sobreabundancia [ASM05]. Sobreestimación [HODB09]. Sobreexplotación [NL05]. Sobreexplotados [UWB04]. Sobrestimación [JSW08]. Sobrestimadas [RH09]. Sobrevivencia [AP00, NTMT^{+}03]. sociabilis [MKH07]. Social [Büs08, Cha06, CY08, DS08, DWW04, FPW01, FCN^{+}06, Ful02, JAH08, MBD^{+}03, Mas03, McB02, MCG^{+}09, Mcs05, PS04, Sch06, SS06, Tho03, Van06, Wal03, WTEK08]. Social-Psychological [DS08]. Sociales [DS08, DWW04, FPW01, JAH08, MCG^{+}09, Mcs05, Wal03, WTEK08]. Sociedad [Dav05]. Sociedades [WG01]. Societal [DZM08, WABT08]. Societies [SHB08, WG01]. Society [And05, Ano03a, Ano05g, Ano05k, Ano06f, Bei05, Dav05, TOR^{+}04, Wil09a, Ano01l]. Sociocultural [CMM^{+}05b]. Socioculturales [CMM^{+}05b]. Socioecological [Hol01]. Socioeconomic [CSB07, CDM09, KCK^{+}08, LG05b, MWBP07, RKEJP06]. Socioeconómica [MWBP07]. Socioeconómico [RKEJP06]. Socioeconómicos [CSB07, CDM09, LG05b]. Sociológica [FNS^{+}03].

Sociological [FNS^{+}03]. Software [Fie00a, PS01a]. Soil [ANW^{+}08, BMB^{+}08, CPD^{+}03, Gun02, Has00, JLF09, MRM^{+}05]. Soils [Ste06a]. Solar [SKP00]. Solidago [Web01, Web01]. Solidaridad [DS08]. Soluciones [TRNH00]. Solution [BW08b, Joh07, LFA^{+}07, WZW^{+}06]. Solutions [GCBF09, TRNH00]. Solving [CNHT09]. Sombra [RCP00, SMFPCA05]. Sombrilla [OIK^{+}06, PMC01, RB06, Rub01, SGH02]. Some [Don04, EG07, Har05, LARR03, MB09b, WB06]. son [AP00, BSM08, BV01b, Fau01, GM01, HRS00, Ric07]. Sondeo [PK05, DH00b]. Songbird [BFD07, DWN01, JSHM06, NS01, OK01a, TB04]. Songbirds [BHB08, Cas03, PSBG05, Sch03b, WS04]. Soporte [Shi06, TUE06]. Sorrow [Sod06]. Sostener [Ber04a]. Sostenibilidad [SVV07]. Sostenible [LMB00, Sie04]. Sostenibles [BMPNC^{+}03, TB01]. Sotobosque [BMMR03, BK02, CS06, DMSP00, DS01, KHP04, LSL04, MB01, New06, RWRW04, TO02, WCPB03]. Sotobosques [RRH^{+}09]. Sound [MK02b, MK02b]. Soup [Ano01h]. Source [CMS09, DMS03, DA07, FCL00, Mur01, NFB05, NRB00, PC08, SFS00]. Source-Destination [CMS09]. Source-Sink [FCL00, Mur01, NRB00]. Sourcebook [Doa01]. Sources [HR04a, KGS^{+}07]. South [Ant08, BHMC01, MB00, PLBBJC09, PEC01, SWD00, Büs08, DNZK02, KTN^{+}09, KDC^{+}06, LMP00, LADM05, Mad05, MHC^{+}01, ND04, PA04, Ros08]. South-Central [PA04]. South-Temperate [SWD00]. Southbound [Orr03c]. Southeast [CSS00, SL00b, SCS06]. Southeastern [RSD03, BMM^{+}01, CTNR02, GFTB04, PBK09, PMZ06, SF03, WFP01]. Southern [GSB04, SLCW08, ASD01, BKV^{+}05, BPB07, BRV06, Cam08, CRGG01, Gra01, Has00, JMH01, KF01a, ML02, MNL03, MM02b, MB09b, Pat00, PLBBJC09, PV02, PEC01, RSF^{+}03, RBK^{+}05, RBHKK02, RRH^{+}09, Rub01, SAA^{+}01, SKP00, WSS^{+}04]. Southwest [DeN07]. Southwestern [BJB06, CMR05, Cra02, Hob01, PRMV06, Say05, Sul09, TB01, XM07]. Soy [NSA06]. Soya [NSA06]. Space [BGC^{+}04, BVH^{+}02, HSR05, Mer04, SRR00]. Space-Use [SRR00]. Spain [NWRD03]. Spanish [Gra03, Rau02, Rod03]. Spared [MV06]. Sparrows [RSJE06]. Spatial [AW02, BRTW07, BL06, CP00, CPSS03, CJ08, DSJ02, FP02, FSC02, GCV03, HLB^{+}03, KFSR^{+}06, Lai05, LSN07, LMP02, Lyn08, NLCAR09, RKEJP06, SK02a, SHM04, SBVC03, TNTH^{+}04, VTA02, WHS02, WHH00, ZGR02]. Spatial-Accumulation [ZGR02]. Spatially [KKHS05, RBHKK02, TUE06]. Spatiotemporal [DJOW07, RLXL01]. Spawning [SCD^{+}08]. Speak [Flo05]. Special [BDRC05, JW02, Mef06a, DW06, LGB05, NL06, RP07]. Specialist [PZG^{+}05]. Specialists [FN04]. Specialization [SK04]. Speciation [Rab09]. Specie [ED07, Gre09, HP04, RWM^{+}06]. Species [AS01, APCS05, AL03, ALH^{+}04, AMF^{+}01, ABE01, ACC^{+}05, Ano08l, Arm02, Arm05, AHTM05, ACMA04, ARPEBM09, BRTW07, Bal02, BBBM02, BMBM06, BFB^{+}09, Bei03, BMMR03, BSS^{+}05, Ber03, BOJ09, BTK08, BSKH^{+}06, BEB00, Bos07, BV01a, Bow09, Bre02b, BS00, BZD03, dLBBG^{+}08, BDR04b, BDR04a, BOC00, BL06, BMMW07, BV00, BV01b, But03, BRR^{+}02, CWL^{+}00, CP01, CVS^{+}09, CK05, Car00, CES05, CF01, Car04, CMC03, CKL^{+}00, CPD^{+}09, Che09, CDD^{+}05, CHBK02, Cla09, CMA^{+}06, Col01, CCM00c, CKF^{+}04, Cra02, CD07, CCH03, Cun00, Cze05, DZM08, Dal02, DBL06, DC06a, DC06b, DK06, DCD^{+}02, Des06, DMS03, DS01, DFS05, DDMMZD06, DTCC03, DM04a, DP01, DKS^{+}09, DAB06, DHW01, ESB05, Edw04, Ehr06, EM08, Fai08, Fal07, Fat06, FEGB09, FMPB07, FPL^{+}04, FNFM01]. Species [FTN^{+}05, FNM05, FF08, FBK02, GRB02, GGA01, GS01, Ger03, GWS07, GCV03, GV02, GLMN00, GPMM08, GDAV00, GMA07, GM01, GPC^{+}03, GBH^{+}09, Gre05, Gre09, GNEE07, GG02, GBE^{+}06, Gus02, HBC^{+}06, HRS00, HMK^{+}09, HH00, HR02, HP01, HW08, HBBD08, HH09, Hew04, HRP^{+}04, HL07b, HC09, Hod04a, HFR07, HJ07, HBSD04, Hop07, HF04, Hun07b, JTS06, JWB06, JSW08, JFWP06, JMP09, JMH01, JMS06b, KB04a, KL05, KFL08, KTN^{+}09, KU02, KB04b, KD05, Kot00, KPJ06, KHP04, Lam02, LBR07, LU09, LWSM03, LTC^{+}08, LRO^{+}08, LS01, LBG03, LBPBG07, LLWCW09, LP05, LCZ^{+}05, LMS^{+}02, aARLG03, LMP02, LSF03, LHG^{+}03, LB00, LBTS05, MCG^{+}08, MLN^{+}06, MKH07, MP09, McK02, MP08, MBH07, MD04, Men03, MHM09, MRH^{+}00, MRAF^{+}07, MW07, MML06, MPFBH07, MBB^{+}08, MSSM09, NA01]. Species [NF04, NLCAR09, NWRD03, NNM00, NP06, ND03, NBM^{+}04, OBK^{+}04, ØDEA00, Old03, OS04, Ova03, OIK^{+}06, Par03, PCL^{+}03, PE01, PM07, Ped07, PS01a, PMC01, Pra05, PTP^{+}08, PJK03, Qia07, QR05, RO08, RBT08, RNP^{+}07, RC00, RH09, RM01, RBK^{+}05, RPP^{+}00, RA04a, RW03, RGGS02, RM02, Rub01, Rub06b, RLXL01, SRC^{+}02, Say05, SGB^{+}08, Sch10, SPM05, SDRO01, Sim03, SSL06, SKBB00, SM07b, SEBM03, SS06, SD03, SHT^{+}05, SCC^{+}09, SDJK04, SBVC03, SGH02, Tay03a, TTSP04, Tho00, TTB03, TNFH07, TNRJ02, TK06, TWS06, VMV07, Vee00, VTR^{+}09, VSGS02, VNP06, WMM04, WABT07a, WFP01, WSCP05, Wen08, WHJJ04, WL04, WBBA03, WMP05, WMLS09, ZJ04, ZLM03, ZGR02, FEGB09]. Species-Accumulation [DFS05]. Species-Area [Fat06, Ova03, Vee00, WMP05]. Species-Area-Habitat [Che09]. Species-Occurrence [Edw04]. Species-Richness [VSGS02]. Species-Specific [Old03]. Specific [JRL03, LME08, Old03]. Specificity [ØDEA00]. Specimen [MD04]. Specimens [JLZW03, REC^{+}00]. Speke [KHM00, Tem02]. Spend [HTU^{+}06]. Spending [Cle06b, HPF^{+}06, McK02]. Spent [Cze03]. Sphenodon [NKBD02, TPT^{+}07]. Spider [PS00a]. Spikes [Rol00]. Spillover [RL06]. Spiritual [Ano02h, Moo07]. Spirituality [Ano02g, Ano02i, Gar03b]. Split [Caf07]. SPOIR [WABT08]. Spontaneous [SKBY03]. Spotlight [Lov05]. Spotted [CJ08, HMF^{+}04, HODB09, NB06, PR03, SS03b, TG04, WM05]. Sprawl [BRTW07, RHS05, Ste06b]. Spread [FDBM04, KFL08, KDG^{+}06, MSOM^{+}06, Wit02]. Spreading [WMCB02]. Spring [Wil03]. Squirrel [BV04]. Squirrels [RMN00]. Stability [DMG^{+}00, FBL09a, MSBGP03, Rou00, San00]. Stabilized [GLMN00]. Stable [WRL02]. Stakeholder [BF03]. Stakeholders [RAL^{+}04]. Stand [Bor05]. Standard [SSS^{+}08]. Stapp [MRH^{+}00]. Star [Caf00c, Vic02]. Starfield [MSB^{+}01]. Starlings [Koe03]. Stars [Sch02]. Start [Fre06]. State [BV01b, FP07, Whe03]. State-of-the-Art [Whe03]. States [BRK^{+}04, BJPL07b, BPB^{+}05, BL06, CTNR02, CML02, CMCH00, Cze05, DC05, EG07, Fau01, For00a, FNM05, GS05, GOL^{+}07, HRW^{+}03, Kee06, Lei00, LB01, LW02b, MM02a, PM07, Ped03, RW00, RLXL01, SD01]. Static [CPSS03]. Station [SSY06]. Statistical [DFS05, FBC^{+}06, Mar04]. statistics [Fle04b]. Status [AM05, BAKHT06, CRGG01, FBD^{+}00, FNM05, Gre05, GHL08, Hay09b, LNP^{+}09, MRAF^{+}07, NKW09, QR05, QR06, QRF07, RRK02, RW00, RS02a, SRDG03, SSK^{+}05, SDJ06, VCHBP08, Wil02c, ZJ04]. Stay [Bek04]. Ste [APS07]. Ste. [APS07]. Steady [FP07]. Steelhead [AAO^{+}07, MBQ05]. Steffe [Lyn08]. Steller [HY03]. Step [JWB06, Rod03]. Steppe [CGR^{+}05, HWH^{+}02, WRW^{+}02]. Steppe-Like [CGR^{+}05]. Stewardship [CPDO01, Ter00]. Still [Pim07b]. Stoat [Cho06]. Stochasticity [Fox07, KF02, KF03, RSC03, SK02a, SRW06]. Stocked [DN00]. Stool [DR03]. Stop [Jol04]. Stopover [BGD09]. Stories [Pyn04]. Storks [Ols07]. Story [BK09, Caf06, Hen04, Kri05]. Strandings [BEH06, LCS03]. Strange [Wil09b]. Strategic [HH08, KC01]. Strategies [BFH^{+}06, HRS00, KU02, LG05b, MBM04, MOO09, MCM07, PCL^{+}03, PC02, RBG^{+}04, SCJ^{+}04, SMV02, SCM^{+}07, TTSP04, VSO04, WWH03, Wil03]. Strategy [Flo05, Hun05, Inf01, MM03b, NR03, RWBG06, SCB^{+}01]. Stream [CS07b, HLG03, JSWS00, MO04, MY02, WD03, WA05]. Stream-Breeding [CS07b]. Stream-Dwelling [HLG03, MY02]. Streams [LB02, MK02b, NR03, PJ00, PW01, RBK^{+}05, Tow03]. Streets [FJ00]. Strength [Hod04b, Jac02]. Strengths [BAKHT06]. Stress [CWNW01, CFH^{+}02, DH00a, FPW01, GMW08, O’C00, TG04, ZSSM09]. Stressed [Koz07]. Stressor [Wen08]. Strife [Dud00, DGP^{+}02]. Strikes [VT09]. Striking [Kle03, KCK^{+}08]. Strip [HNG04]. Stripping [CMC^{+}01]. Strix [CJ08, HMF^{+}04, HODB09, NB06, PR03, TG04]. strobus [BM04]. Strong [GMMPP08, Whi02]. Structure [APCS05, BD06, Ber05, BFD07, BK02, CP01, CQLF02, CNC04, CY08, CNM01, CMR05, DMG^{+}00, EM03, HMB^{+}05, HY03, KGM01, MR00a, ML04, MRP07, New06, PS00a, Per00, PJYO04, RTN01, SG04, TS03, VS05a, WHS02, WHH00]. Structured [KMGB07]. Structures [BPP05]. Struggling [Ano02e]. Studbook [JGL^{+}02]. Student [MS09]. Students [ID00, MSL06, Pér05b, Pim06, TSH06, WMCB02]. Studies [DBML07, Faa04, Fje07, GMB^{+}08, Kér04, LF09, TB04]. Study [APS07, ACMA04, BMBM06, BR01, BTGS05, Che09, DCD^{+}02, DA07, FVS03, Fat06, Gay02, GPZ04, GM01, GZK06, HRN00, HDA01, HW08, HY03, HLJV07, HW02, IB03, JHHL04, JMH01, KZSV05, KU02, LB05, Lin03b, LZD^{+}03, MBM04, MKH07, MCA^{+}00, MS09, New06, ODE01, PBK09, PV09a, PJK03, RAL^{+}04, RWJRR06, RWM^{+}06, RG04a, RMK07, RCMV06, Rua00, SH03, SDRO01, SFT00, SD01, TVD03, TPW03, VEST02, WSS^{+}04, WMPC08]. Studying [HM03, Swe08]. Sub [MBB^{+}03]. Sub-Sahara [MBB^{+}03].

Sub-Saharan [MBB^{+}03]. Subastas [JLF09]. Subdivision [GM00a, GM00a]. Subespecies [HBC^{+}06]. subintegra [MBQAM06]. Subject [Ano03g, Ano08r, Ano09o, Whi06, Ano04l, Ano05l, Ano06k, LD07]. Subjective [HW08]. Subjetivo [HW08]. Subletales [PRL^{+}02]. Sublethal [PRL^{+}02]. Submareas [VDHD01]. Submerged [BMRB04]. Submittal [Ano03d]. Subpoblacional [MMBP08]. Subpopulation [MMBP08]. Subshrub [ME00]. Subsidiado [Ber06]. Subsidiados [RL06]. Subsidized [Ber06, RL06]. Subsistence [Bor08, OSSK^{+}07, Per00, Per01]. Subsistencia [Bor08, OSSK^{+}07, Per00, Per01]. Subspecies [HBC^{+}06]. Substitute [CES05]. Substitutos [OIK^{+}06]. Subtidal [VDHD01]. Suburban [BVH^{+}02, BTS09, FD00, RHS05]. Suburbana [FD00, RHS05]. Suburbanos [BVH^{+}02, BTS09]. Subyacente [TB03]. Subyacentes [SFS05, Sta01]. Success [Ant08, AAO^{+}07, BMM^{+}01, BB04, BV01a, BFH^{+}06, BWS09, CPSS03, CQLF02, DOGN08, DNZK02, DD04b, FLQ03, GPC^{+}03, HHSC00, Hen04, HRW^{+}03, HDQSL06, KS05, Kri04, LDG^{+}00, Mef06b, Mor00b, Ols07, RPM09, Rem03, RR01, Shi06, TR01, VSH^{+}07, Wal09, Wel06, WH01]. Successful [LG05b, NF04]. Succession [BE00, FVMG03, Gav04]. Sucesión [BE00, Gav04, FVMG03]. Sudáfrica [LMP00, MB00, ND04]. Sudafricana [BRV06]. Sudamérica [PLBBJC09]. Sudoriental [CSS00]. Suecia [BJ05, HW08, VSGS02]. Suecos [Gus02]. Suelero [FTR01]. Suelo [ABE^{+}02b, ANW^{+}08, BMB^{+}08, CPD^{+}03, Cua00, DBL06, GGW09, Gun02, HR02, Has00, Kau04, KSDBT02, KFNB09, LBPBG07, LLWCW09, MKG03, MNRD^{+}08, MRM^{+}05, MPR^{+}01, NRBM05, PA04, Say05, SCC^{+}09, WMM04, WM01]. Suelos [HMH00, JLF09]. Suerte [BV01b]. Suffer [KB03]. Sufren [KB03]. Sugerencias [TFGJ06]. Suggestions [TFGJ06]. Suid [CMGSM00]. Suitability [BSM08, FDBM04, GB09, Pyk05, RSB05]. Suitable [PD03a]. Suiza [GLMN00, HBSD04]. Sulawesi [WMM04]. Sullivan [Say09]. Sum [MMK04]. Sumamente [ARPEBM09]. Sumatra [GFCP03, GFCP03, LSZ^{+}08]. Sumatran [LSZ^{+}08]. Sumergidas [BMRB04]. Sumidero [FCL00, LW02b, NRB00]. Suministro [AP00, JLF09]. Summary [Ano00c, KF01b, Lis04, Mar00a]. Summer [Wil03]. Summit [GBC^{+}05]. Sunflowers [MMLM01]. Super [FN04, FN04]. Superficial [BM04]. Superficie [PM07, YKIO08]. Superfluous [Wil02d]. Superior [MDCK07]. Superiores [VIMMO05]. Superman [Pat08]. Supervivencia [AM05, BBB08, DWN01, TSG06, VEST02, WMS^{+}01, ZSSM09]. Suplementarios [FPL^{+}04]. Suposiciones [Pow06, FB08]. Supplemental [FPL^{+}04]. Supply [AP00, JLF09, PCME09, Pow06]. Support [Inf01, RANM08, Sey04, Shi06, TUE06]. Supportive [DS08, For02, WR01, Wed02]. Suppression [BJM04, BSW^{+}08, JMB01]. Suprabenthic [VDHD01]. Suprabéntica [VDHD01]. Supresión [BJM04, BSW^{+}08]. Supuestamente [DAKB08]. Suráfrica [DNZK02]. Suramérica [BHMC01]. Sureste [BMM^{+}01, CTNR02, PBK09, PMZ06, RSD03, SF03, WFP01, GFTB04, SCS06]. Surface [BM04]. Surface-Fire [BM04]. Suroeste [CMR05, Cra02, Hob01, PRMV06, Say05, TB01, XM07].

Surrogate [Arm02, BMP07, CDD^{+}05, JMS06b]. Surrogates [FBP^{+}09, GGA01, MBB^{+}08, OIK^{+}06, SJF^{+}05, VIMMO05, Wen08]. Surrounding [NAC^{+}06]. Surroundings [MYFF^{+}07]. Survey [DH00b, Edi03, Gre06, HW02, MML06, Ped03, Pol06, PK05, Ste08a, SCM^{+}07]. Surveys [Roy04, UWB04]. Survival [AM05, AP00, BBB08, DWN01, MF08, NTMT^{+}03, Orr08b, Sam06, TSG06, VEST02, WMS^{+}01, ZSSM09]. Surviving [McK09]. Sus [BZD03, DDBA^{+}07, DGP^{+}02, Gav04, KSO^{+}03, LFP05, LO06, PR03, RCMV06, RNL^{+}04, SGG07, SS06, BBB08, Gui07, HMF^{+}04, HR02, HP04, LARK01, MYFF^{+}07, ØDEA00, PLB09, PC07, Sie04, TR05, UMT^{+}01]. Susceptibilidad [BRS^{+}05, GGI^{+}05, PJYO04, GJD05, HJ07]. Susceptibility [BRS^{+}05, GGI^{+}05, GJD05, HJ07, PJYO04]. Sustain [Ber04a, RR00]. Sustainability [Ano02f, Ano02g, Ano02i, Ano02h, Coc04, EN07, Ehr05b, FVS03, FV05, Gar03b, JAH^{+}05, Jor01, Kaw04, LMG06, NF05b, OSSK^{+}07, Orr06a, PS05, Pae05, PBP09, RW02, SVV07]. Sustainability-Biodiversity-Agriculture [Jor01]. Sustainable [BL01, BMPNC^{+}03, DR03, EGPS04, GBC^{+}05, IB03, Kra01, LMB00, SK02a, SAKQ^{+}05, Sch00, Sey04, Sie04, TB01, UCAGG08, WBV^{+}05, You02]. Sustaining [Chr08]. Sustentabilidad [FV05, JAH^{+}05, LMG06, NF05b, OSSK^{+}07, PBP09, RW02]. Sustentable [EGPS04, SK02a, WBV^{+}05, GBC^{+}05]. Sustentables [FVS03, UCAGG08]. Sustitutas [CES05]. Sustituto [CDD^{+}05, JMS06b]. Sustitutos [BMP07, FBP^{+}09, MBB^{+}08, SJF^{+}05, Wen08]. Suter [TPW03]. suteri [TPW03]. Swamp [CTNR02, FTC^{+}04b]. Swan [Gib07]. Swayne [FSS^{+}00]. Sweden [HW08, ZHM08]. Swedish [BJ05, Gus02, VSGS02]. Swiss [HBSD04]. Sword [LS04]. Swordfish [Bea09b]. sylvatica [WG06]. Sylvia [Rem03]. Symmetry [Sei07]. Sympatric [BOC00, Fox07]. Synchronous [WBDP02]. Synergism [PRL^{+}02]. Synergisms [Hob01, LU09, WCP01]. Synergistic [BBB08, Coc01, Per01]. Synthesis [DWWF04, HM08, Les05, Lis04, Mar00a, WD06]. System [BNM^{+}09, Cze05, DDAK04, EBAW01, For00a, HH08, HH09, HCZ00, HDK^{+}00, JHHL04, KFSR^{+}06, LS09, LM05, LVM02, MCG^{+}08, MK02a, RRK02, Rou00, SM01, TR05, WU05]. Systematic [Ara09a, KDC^{+}06, Moi08, PB08, PS06, RB07, RP07, SCB^{+}01, WSS^{+}04]. Systems [BMP07, CPR05, CD07, DC06a, Don04, KSDBT02, LRO^{+}08, Nas06, NRB00, O’C00, OBK^{+}04, Pat00, SM07b, SBP07].

Taber [Kni06]. Taboos [JAH08]. Tabús [JAH08]. Tackling [RB07]. Tadpoles [BRS^{+}05, BS00, SWR00]. taeda [LGK02]. Taiga [Sch03a]. Tailed [ARMH04, Côt05, Shi06]. Taita [LVM02]. Takahe [JRL03]. Take [FDL^{+}05, LB05]. Takes [PBSR06]. Taking [RM03]. Tala [CBG^{+}00, GV02, HNG04, KHHK03, Lew01, MM07, MRM^{+}05, PMZ06, SH09, STE^{+}01]. Talada [HKN01]. Talado [DMSP00, PBL00, WLCS00]. Talados [GAF00, Olu00, TK07]. Tale [RW03]. Tales [Sal07]. Talla [FV05].

Tallapoosa [IF02, IF02]. Tallgrass [CCH03, MMLM01, RJW09]. Tamaño [AT05, BST^{+}05, Ber05, BPL^{+}01, CWNW01, Chi00a, CNC04, DAR08, DM04b, DW05b, EKL^{+}00, ED07, HRS08, HHSC00, Kai03, KW02, Kel01, LP07, MBQ05, MM04, MK09, PH02, PM07, PV06, Pod01, RLC^{+}00, Ree05, SRR00, SFD09, SM07a, SZS01, SKBB00, Vic02, WH01, YE07, CB01a, FPF00, HK00, LDG^{+}00, SC05, Wal00]. Tamarix [SMG00]. Tambopata [NTMT^{+}03]. Taming [Caf01]. Tampa [AJF03]. tan [Kot00]. Tanganyika [MMF^{+}05]. Tanzania [KNH07, MCM07, New06, PB09, SCM^{+}07]. Tapajós [UMT^{+}01, UMT^{+}01]. Tapires [SK02a]. Tapirs [SK02a]. Tardío [LW02b, RWS02, SK02b]. Tardío/Histórico [LW02b]. Target [BPL^{+}01, Cla09]. Target-Species [Cla09]. Targeting [PCVC06, The03]. Targets [GV05, JFS08, LSL^{+}06, NC03a, WSS^{+}04, WG08]. Tarifaria [WGW01]. Tasa [MRDC06]. Tasas [BM05a, BBK02, DWN01, Fea05, HMF03, Kér04, MM04, PJW^{+}02, SS01, UCAGG08, ZGR02, dLBBG^{+}08]. Tasmania [ES04, ESB05, Lei05, MK02a]. tatarica [JTS06]. Taxa [BMP07, GA02, HW02, KSP00, Qia07, VSGS02, VIMMO05, YKIO08, VIMMO05]. Taxes [Rom03]. taxifolia [SHM00]. Taxon [CDD^{+}05, CTNR02, MDCK07, SDJK04, CDD^{+}05, CTNR02, MDCK07]. Taxones [FPL^{+}04, HW02, SDJK04, VSGS02, KSP00]. Taxonomía [MN03, ZBABR00]. Taxonomic [ALR05, DCD^{+}02, FTN^{+}05, FM09, HBC^{+}06, KDD^{+}04, LRG^{+}02, MD04, PC07, RGW09, San00, Wil02c, GCM^{+}08]. Taxonómica [PC07, RGW09]. Taxonómicas [DCD^{+}02, HBC^{+}06, MD04]. Taxonómico [Wil02c]. Taxonómicos [ALR05, FTN^{+}05, FM09, KDD^{+}04, LRG^{+}02]. Taxonomists [GT03, HW02]. Taxónomos [HW02]. Taxonomy [GT03, MN03, Yah04, ZBABR00]. Taxus [KHP04]. Taylor [SRDG03]. Teaching [Fie00a, NL03a]. Teachings [RF00]. Teams [HSDS01]. Technique [SD02]. Techniques [HR04c, Mar04, MBH07, STC03, WGGS07]. Technological [Cze03, Cze08]. Technologies [SC00]. Technology [Bal01, Bre03, Cor04a, HPK^{+}04, HMH03, McN03, Ril00]. Técnicas [HR04c, WGGS07, STC03]. Tecnología [HPK^{+}04]. Tecnologías [SC00]. Tecnológico [Cze08]. Tehuacán [TVD03, TVD03]. Tejón [RPD01]. Tell [Kri05, Sal07]. Telling [Caf06]. Temas [PBC^{+}09, RMD^{+}02]. Temático [DDN01]. Temperate [BHMC01, BSKH^{+}06, CCM00c, GMCP09, GG02, HF04, KPJ06, LP07, MRP07, ND03, PJK03, RR01, SK04, SWD00, SWB^{+}08]. Temperate-Deciduous [GG02]. Temperate-Zone [SK04]. Temperaturas [KLH08]. Temperatures [KLH08]. Templada [KPJ06, ND03]. Templadas [BSKH^{+}06, PJK03, SK04]. Templado [CCM00c, GG02, LP07, RR01, SWD00]. Templados [BHMC01, HF04, MRP07, SWB^{+}08]. Temporal [BGD09, CP00, KG03, Lyn08, NFM04, VMBC02]. temporaria [SM07a]. Temporización [KZA09]. Temporospatial [PCD^{+}03]. Temprana [LVM02]. Temprano [LW02b]. Tempranos [MRP07]. Ten [Orr01c]. Tendencia [QRF07]. Tendencias [For00b, GS05, MT08, PCHN04, Phi00, QR05, SMSB05, WBDP02, KSO^{+}03, MSBGP03, MMB^{+}03]. Tenerife [HLJV07]. Tenses [Rob06a]. Tenure [CMM^{+}05b]. Teoría [DKW01, Esl00, FDL^{+}05, LG02]. Teorías [Gil01b]. Teórico [MU08, WM05]. Term [BSW^{+}06, CWL^{+}00, CWH07, CBG^{+}00, DBML07, FTC^{+}04b, GMW08, Gut05, IMD01, LLH08, PPW^{+}07, SRW^{+}08, Sch03a, SMH^{+}07, SLKP06, SBP07, SBSL06, VEST02, WLS^{+}08, WBDP02, WGGS07]. Terms [Ano02i]. Terratenientes [BZD03]. Terreno [SLKP06]. Terrenos [BRK^{+}04, DBB^{+}09a, Her09, NSB^{+}06, VW04, WPS08, WTEK08, WE07]. Terrestre [ALM07, CNC04, CS07b, HRS08, TFG08, TSG06]. Terrestres [AM05, BHMC01, GM01, GJD05, HNG04, KHHK03, LFP05, LB01, MMGB05, MNL03, SLELA05, SF05, TF00]. Terrestrial [ALM07, BHMC01, BNM^{+}09, CNC04, CS07b, HRS08, KHHK03, LFP05, LB01, MMGB05, TFG08, TSG06, TF00]. Territorial [Rem03, WS04]. Territoriales [WS04]. Territories [O’N07]. Territorio [MM02a]. Territory [Rem03]. Test [Ano01l, CS06, PAHM04, RO04, SCB^{+}01, VSO04]. Testable [Pea01]. Testing [AP00, BR01, BFH^{+}06, CT07, FBC^{+}06, KDD^{+}04, LME08, MPDT01, ME00, PBRG07, Roo00, SGH02]. Tests [DSJ02, FNF03, KSB01, MBL02, PMEC04, RA05]. Tether [Ano08a]. Tetrao [SGH02]. Texas [KHRS01]. Text [JG04]. Textbooks [Pri03]. Textural [Sta09]. thapsus [PRL03]. Their [CHF00, Cor02, Gav04, Gui07, HMF^{+}04, Koe03, LFP05, LARR03, MB01, MYFF^{+}07, RCMV06, SGG07, Sie04, SS06, Van04, KZSV05]. Them [GBB^{+}08, MA05]. Thematic [DDN01]. Theology [Ano05c]. Theoretic [WM05]. Theories [Gil01b, Lea06]. Theory [Chi00b, DKW01, Esl00, FDL^{+}05, HMH03, LG02, Man02, MU08, Ped07, Roo00, WP03]. There [Ben02, Hol08, KWV08a, LP05]. Thing [Lei05, Nos07]. Things [Gal03, JGM03b, NB00, Pre06]. Thinking [Ada06, Nas06]. Thistle [LO02]. Thompson [Sha01]. Thoreau [Chi09a]. Thought [Gro05a]. Thought-Provoking [Gro05a]. Thoughts [Flo05]. Thread [Chi09a, Pas00]. Threat [Ant08, BCK^{+}09, CGR^{+}05, Hay09b, KCS04, MP09, McK02, MRDC06, ND04, NKW09, PAB^{+}08, RL06, SL08, VCHBP08]. Threatened [dLBBG^{+}08, Col01, CD07, CV08, Cze05, DC06a, DAB06, Fai08, GLMN00, GBC^{+}04, GBH^{+}09, HY03, JFWP06, JMP09, Kal05, LS09, LLWCW09, MCG^{+}08, MP08, MRAF^{+}07, Mor00b, MSSM09, NA01, ND03, OBK^{+}04, PCL^{+}03, PGP04, PD03a, RC00, Sch10, SM07b, Tay03a, TR05, TWS06, VMV07, WABT07a, YE07, ZBABR00]. Threatening [Hay09b]. Threats [BKV^{+}05, BCD^{+}08, DGSS03, HSMK07, HTM^{+}05, PDLS08, PB08, RAW07, RSW05, SSS^{+}08, SBS^{+}09, SRHS09, TFG08, The03, TSKD07, TSKD08]. Three [Apl01, AAO^{+}07, BEB00, CVS^{+}09, Dal09, DR03, LMP00, LCS04, PCL^{+}03, PB04, Rob06a, Ste08b, Web01, WL04, WCX09, PCL^{+}03]. Three-Legged-Stool [DR03]. Three-Wattled [PB04]. Threshold [HO02, Rol06b]. Thresholds [BFD07, CSB07, GV05, HBLW09, SC05]. Throwing [BBA^{+}02]. Thrush [Fau01, LVM02, SFS00]. Thrushes [DD04b]. Tibetan [LWL09]. Tiburón [CMA^{+}06, SCP^{+}02]. Tiburones [Cor02, FMSL08]. Tidal [RLK^{+}01]. Tidal-Flat [RLK^{+}01]. Tiempo [LWV04, MBQ05, NH09, Pol06, FB00]. Tiempos [Nos01a, Ros06]. Tierra [HFR07, NMB09, RJW09, SZBR03]. Tierras [BMRB04, BMS08, CLC^{+}08, Cur02a, Cze02b, DN01, DWW04, FB00, FB08, Gav04, GAF00, HM03, JMS^{+}06a, Jan07, JSJ05, Lam04, LMMF00, Mcs05, MH01b, NRBM05, Nos03, PP03, Say05, SP09, VW04, WG01, KB08, STE^{+}01]. Tiger [Ber01, Pyk05, SS08a]. Tigers [Din03a, GBB^{+}08, KGM^{+}02, WMD^{+}04]. Tigre [SS08a]. Tigres [KGM^{+}02, WMD^{+}04]. Timber [Ben00, Ben01a, Ben01b, CWH07, Cra02, FMLL06, MR00a, NARL^{+}04, Pow06, PR01b, PPPV08, ZSVL07]. Time [Fie05, FB00, LWV04, NH09, Nos01a, Orr07c, Pol06, Ste05, Whi03, Woo03, YBWZ06]. Time-Lagged [Ste05]. Time-to-Encounter [Pol06]. Timelessly [Gre06]. Times [FP07, Orr04b, DW06]. Timescale [VMBC02]. Timing [KZA09, MBQ05]. Timing-Biased [MBQ05]. Tipo [MKK^{+}09, MK09]. Tipos [HSB01]. Tirano [Mur01]. Tit [WGGS07]. Toads [BS09, BR01, PBS03, SB08]. Toda [PSGA09]. Todas [Fau01]. Today [Het04]. Todd [Col01]. Together [She02]. Tolerance [BS00, HP04, MMM03, NTGT03]. Tolerancia [BS00, MMM03, HP04]. Toma [Pra05]. Tomadas [RLK^{+}01]. Tongass [Bei08, Bei08]. Too [DCT^{+}07, GBB^{+}08, GNR01, LK04]. Tool [BPP05, COW03, DDMMZD06, HLB^{+}03, Hol00, LOM03, LB00, Mag04, MC05a, PGD^{+}07, PS01a, PCC03, RA04a, SBB^{+}08, Tho05b, Wal09]. Toolbox [Tho05a]. Tools [BMRB04, BFMR07, CSB07, Ehr06, Men03, Ray06b, Tho05a]. Top [AAO^{+}09, DAB^{+}04, FVS03]. Top-Down [AAO^{+}09]. Topográficos [FN02a]. Topographic [FN02a]. Tordo [LVM02]. Tordos [GM00b]. Toromiro [MCA^{+}00]. Torreya [SHM00]. Tortoise [KHRS01]. Tortoises [FBCG00]. Tortuga [BBH06, KHRS01, ML04]. Tortugas [FCG^{+}05, FBCG00, GS02, GS05, JMH01, LCS03, LC07a, PLB09, ST05, SG04, TSP^{+}08]. Total [JFS08, JMS06b]. Toucans [Gra01]. Tough [Fox08]. Tour [Lus03]. Tourism [WGW01]. Tourist [WBW05]. Tourist-Visited [WBW05]. Tourists [GVTM07]. Toxic [SSG01]. Trabaja [MH02]. Trabajan [MS09]. Trabajo [MS09, Rua00, TAP^{+}01]. Tracción [DPB^{+}09]. Track [Rei05a]. Tracking [HPK^{+}04, WM01]. Traction [DPB^{+}09]. Tradable [Ros06]. Trade [Ano08l, BM05b, BM06, CMM^{+}05b, CMA^{+}06, COW03, CV08, DC06b, DSTPW07, FRD^{+}09, FV05, FF08, GCM^{+}08, KB03, LD03, PPPV08, Rob00, RGW09, SCP^{+}02, TA06, WCD^{+}08, WELCH09, WSCP05, ZJ04]. Trade-Management [CV08]. Trade-Offs [FRD^{+}09, WELCH09]. Tradición [CMM^{+}05b]. Tradicional [CMM05a, Dre05, Pyk00, RNP^{+}07]. Tradicionales [RPD01, TWGH00]. Tradition [CMM^{+}05b, Gru04]. Traditional [CMM05a, Dre05, Pyk00, RNP^{+}07, RPD01, TWGH00]. Traffic [BGC^{+}04, CF01]. Tráfico [BGC^{+}04, CF01]. Trailblazers [Sha01]. Trails [DGC05, PIH00]. Train [Ehr03, Jol04, Orr03c]. Trained [ID00]. Training [Bau05, BRD03, GBE00, MS09, VO03]. Trait [LLF08]. Trait-Dependent [LLF08]. Traits [CVS^{+}09, DW05b, FDBM04, JP09, KD05, LRW03, MKK^{+}09, SFD09]. Trampas [Bat04, PA08]. Trampling [ANW^{+}08]. Trans [SDJK04]. Trans-Taxones [SDJK04]. Transdisciplinario [Cze02b]. Transdisciplinary [Cze02b]. Transfer [KBM01, You02]. Transferability [GBH^{+}09, VMV07]. Transferencia [KBM01]. Transferibilidad [GBH^{+}09, VMV07]. Transformación [ND04]. Transformado [KFNB09]. Transformation [ND04]. Transformed [KFNB09]. Transforming [Par02]. Transgenética [Ehr06]. Transgenics [Ehr06]. Transición [MS08, NA06]. Transition [Ara02]. Transitional [NA06]. Transitioning [MS08]. Translate [TSH06]. Translating [Bre06]. Translocación [BPB^{+}05, DA07, GB09, HGH^{+}03, TTSP04]. Translocada [KDD02]. Translocados [Shi06]. Translocated [GBE00, KDD02, Shi06]. Translocation [BPB^{+}05, DA07, GB09, HRS00, HGH^{+}03, TTSP04, TM02]. Translocations [SD02]. Transmisión [RIBGG08]. Transmission [RIBGG08]. Transmitidos [Log03]. Transmitted [Log03]. Transnacional [HPG09]. Transplantation [ME00]. Transplante [ME00]. Transponiendo [KTN^{+}09]. Transport [PV09b]. Transport-Establishment [PV09b]. Transporte [PV09b]. Transvalued [RH09]. Trap [KSDBT02, KT02, LRS09a]. Trap-Nesting [KSDBT02, KT02]. Traps [Bat04, PA08]. Traslado [HRS00]. Traslape [LSN07]. Trasquilado [SVV07]. Tratamientos [BMM^{+}01, DWN01]. través [LB05, TAJ04, WBW06, WSCP05, RG02, The03]. Treasure [Ste07]. Treatment [VJ09]. Treatments [BMM^{+}01, DWN01]. Treaty [Pau08]. Tree [BFB^{+}09, BOJ09, CQLF02, CLC^{+}08, FJ07, FLQ03, GRB02, GFCP03, GGA01, HDQSL06, HN09, HKM06, KFBG08, Kre07, KHP04, MR00a, MM07, MRP07, MCA^{+}00, NH09, NAC^{+}06, PS00a, PGP04, PS01a, PV09b, SHM00, WBBA03, WLO^{+}00]. Tree-Living [BOJ09]. Tree-Species [GRB02]. Treecreeper [HAJ^{+}04]. Trees [APSB02, CH01, CBH01, GGGS00, GLF^{+}08, LVBS^{+}04, NHHG^{+}05, PAA^{+}08, RWJRR06, SH09]. Trematode [Urb06b]. Trematodos [Urb06b]. Trend [QR05, QRF07]. Trends [BAKHT06, For00b, GS05, KSO^{+}03, MT08, MSBGP03, MMB^{+}03, PCHN04, Phi00, QR06, SMSB05, WBDP02]. Tres [AAO^{+}07, BEB00, CVS^{+}09, LMP00, LCS04, Web01, WL04, WCX09]. Triage [MMBP08, MMBP08]. Trial [Got06, Orr06c]. Triangle [DSE^{+}08]. Triángulo [DSE^{+}08]. Tribute [DeN07, Est04]. Tributes [SF00]. tricarunculata [PB04]. Trickle [Bre01]. Trickle-Down [Bre01]. Tridacna [DSE^{+}08]. Trillium [TO02]. Tritones [DDK05, JMLG01]. Trivers [Lin03a]. Trofeo [Mcc09]. Trófica [CVS^{+}09]. Tróficamente [Log03]. Tróficas [LAN05]. Trófico [BB06]. Tróficos [CVS^{+}09, CY08, VSC06]. Trombulak [Dal00b]. Trophic [BB06, CVS^{+}09, CY08, VSC06]. Trophically [Log03]. Trophy [LFA^{+}07, Mcc09, WSQP07]. Tropical [And01a, Ano00b, Ano01j, Ano01k, Ano04d, ARA^{+}09b, BRD03, BS03, BWS09, CWL^{+}00, CQLF02, CPD^{+}09, CPME09, Coc01, Col00, Dun04, EAA^{+}09, ESSGL00, FPM02, FLQ03, GGGS00, Gra01, GBMH07, Gri00, HH00, Har06b, Hel08, HDQSL06, HR04c, HAG^{+}09, Hol08, HDDEB03, JHHL04, KSC^{+}06, KWV08a, KSO^{+}03, KFBG08, KSDBT02, KSB04, LU09, Lew01, PBP09, PL03, PS01a, PV02, PIH00, PCME09, PB04, Ric04b, RS01, RA04b, Rou00, Rud06, SAKQ^{+}05, SZBR03, SG01, SNM00, SMN00, SSY06, SBB^{+}08, SLKP06, SMFPCA05, STE^{+}01, SCC^{+}09, SOADMR08, Ter00, VS09, WK09, WLCS00, WA05, WMLS09, Wun07]. Tropicales [ARA^{+}09b, BWS09, CPD^{+}09, CPME09, Coc01, FPM02, GBMH07, HH00, HAG^{+}09, HDDEB03, KFBG08, KSDBT02, KSB04, LU09, Lew01, PBP09, PL03, RS01, Rud06, SZBR03, SMN00, SBB^{+}08, SCC^{+}09, VS09, WMLS09]. Trópico [HDK^{+}00]. Trópicos [Cle06b, MPR^{+}01, PDH00, Wun07]. Tropics [Cle06b, HDK^{+}00, MPR^{+}01, PDH00, She01]. Trout [ALH^{+}04, ASF^{+}09, AAO^{+}07, CK05, DN00, MBQ05, NR03, PD03a, YH01]. Trove [Ste07]. Trucha [ASF^{+}09, MBQ05, MY02, NR03, YH01]. Truchas [AAO^{+}07, DN00]. True [Wil09b, Win04]. Truly [Faa04]. Trusts [MHGF04]. Truth [Flo05]. tshawytscha [BPB07, ZSMW06]. TSN [GCM^{+}08]. Tsunami [LT05]. Tuatara [NKBD02, RNP^{+}07, TPT^{+}07]. Tucanes [Gra01]. Tucker [RSW05]. Tumba [Ram01]. Tuna [SK08, WELCH09]. Turba [ND03]. Turismo [GVTM07, Lus03, WGW01]. Turistas [WBW05]. Turning [WLS^{+}08, Wes05]. Turnover [HMH00]. Tursiops [Lus03]. Turtle [BBH06, CFO^{+}09, FCG^{+}05, GS02, JMH01, LCS03, ML04, ST05, SG04]. Turtles [GS05, HPLTH07, LC07a, PLB09, TSP^{+}08]. Twenty [MEN06, Par01]. Twenty-First [Par01]. Two [BV01a, BOC00, CF01, Chi09b, DAR08, DOW^{+}06, FN04, FBB^{+}09, GBH^{+}05, GAF00, GS08, HSB01, JMH01, Kri05, Min01, Orr00b, RW03, RB06, Sei04, SS02a, Ste08b, SS01, VMV07, WBDP02, ZSSM09]. Two-Part [GS08]. Tympanuchus [ARMH04, BJTD03]. Type [MKK^{+}09, MK09]. Types [BT07b, HSB01]. Tyranny [PS00b].

U.S [SF03]. U.S. [ALH^{+}04, BMM^{+}01, CK05, CHBK02, Cur02a, DZM08, Dal05, DD01, DP01, Fal07, Gre09, HBC^{+}06, Har07, KB08, Lav03, NPT05, PH02, Pau08, PCHN04, RHS05, RDN06, Say05, SDRO01, Sta01, VNP06, Wal03, ZMS^{+}08]. U.S.A [KBR03]. U.S.A. [BJTD03, BJB06, DGC05, FD00, HWH^{+}02, IF02, KU02, RWS02, Sul09, VDP00, WM05, WRW^{+}02, WWR^{+}02]. Ubicación [ME00, PCL08]. Uganda [CBG^{+}00, DMSP00, Ghi06, HCBK00, HDK^{+}00, OBM09, RIBGG08, TKB^{+}06]. UICN [Hay09b, KG03, MCG^{+}08]. Última [Ber04a]. Ultimate [Ano01l]. Últimos [PDS^{+}07]. Ultraviolet [BBB08, BOC00, MHC^{+}01]. Ultraviolet-B [BOC00]. Ultravioleta [BBB08, BOC00, MHC^{+}01]. Ultravioleta-B [BOC00]. Ulysses [BM03]. Umbral [HO02]. Umbrales [SC05, BFD07, CSB07, GV05]. Umbrella [OIK^{+}06, PMC01, RA04a, RB06, Rub01, SGH02]. Umbrella-Species [SGH02]. Unbearable [Mef09c, Orr09b]. Uncertain [HRS00, HSR05, PCC03]. Uncertainty [AFB^{+}00, BRV06, Cor02, For00b, MWBP07, MC05b, MWEB06, Pra05, RT00, SFT00, SLCW08, TWDB00, WMVK03]. Uncia [MAM^{+}03]. Undergoing [ML04]. Underlying [SFS05, Sta01, TB03]. Underpasses [CW00]. Understanding [Alt05, Arm05, BC08, Cla09, DWW04, EHJ^{+}04, FCL00, Hew04, JMS^{+}06a, LLF08, LF09, MC09, PRL03, Tho05b]. Understories [RRH^{+}09]. Understory [BMMR03, BK02, CS06, DMSP00, DS01, KHP04, LSL04, MB01, New06, RWRW04, TO02, WCPB03]. Underview [Col06]. Undisturbed [WBW05]. Uneasy [Sei04]. Unexpected [Dor06, Fie06, LO02, WB06]. Unfragmented [RWRW04]. Ungulados [CWL^{+}00, CGR01, MNA07]. Ungulate [CWL^{+}00]. Ungulates [MNA07]. Único [FV05]. Unidades [FP02, RG02, Wal00, DW05a, Gre05, WG08]. Unido [EGN^{+}05, PK05]. Unidos [BMM^{+}01, BJPL07b, BPB^{+}05, BL06, CTNR02, CML02, Cur02a, Cze05, DC05, DP01, Fau01, For00a, GS05, HRW^{+}03, Kee06, LB01, LW02b, MM02a, PH02, PM07, Ped03, RW00, RLXL01, Sta01, Wal03]. Unificadas [SSS^{+}08]. Unified [SSS^{+}08]. Uniform [VMM07]. Uniforme [VMM07]. Unifying [Chi09a]. Union [FBD^{+}00, Wes05]. Unique [RVVNH09]. Uniqueness [Rol06a]. Unit [DGG^{+}01]. United [BRK^{+}04, BPB^{+}05, CTNR02, Fau01, Kee06, LW02b, Ped03, BJPL07b, BL06, CML02, CMCH00, Cze05, DC05, EGN^{+}05, EG07, For00a, FNM05, GS05, GOL^{+}07, HRW^{+}03, Lei00, LB01, MM02a, PM07, PK05, RW00, RLXL01, SD01]. Units [DW05a, FP02, Gal01b, Gre05, Mag01, Pea01, RG02, Wal00, WG08, You01]. Unity [Mef01c]. University [IB03]. Unlogged [Olu00]. Unnoticed [RFB09]. Unprecedented [Ric07]. Unresolved [Bei02b]. Unsampled [RTCR03]. Unstructured [KF03]. Unveiled [Lee09]. Update [Dav04]. Updated [LDC06]. Uplands [SMC00]. Upper [BRYK01]. Upstream [FRD^{+}09, RFD01]. Uptake [Gha07]. Urban [BRTW07, BG04, DM04b, DGSS06, EYZ09, FJ00, MNRD^{+}08, Mer04, MK02b, RWJRR06, RAW07, SMM09, WM01]. Urban-Countryside [BG04]. Urbana [RHS05]. Urbanas [DM04b, RAW07]. Urbanización [DJC^{+}07, PGP04, RSF^{+}03, RBK^{+}05]. Urbanizado [RK05]. Urbanization [DJC^{+}07, PGP04, RSF^{+}03, RBK^{+}05]. Urbanized [RK05]. Urbano [FJ00, MNRD^{+}08, SMM09, WM01]. Urbanos [MK02b, RWJRR06]. Urgent [HL07a]. Urocyon [ARB^{+}07]. Urogallos [PPW09, WRW^{+}02, WWR^{+}02]. urogallus [SGH02]. Urraca [MSWF00]. US$ [Koh09]. Usados [Cra02]. Usando [FNFM01, HCZ00, TCSD00, SBDD05]. Use [ABE^{+}02b, Ano02j, Ant08, BBBM02, BGC^{+}04, BL01, BMB^{+}08, CHW^{+}01, CES05, CS06, CMM^{+}05b, CSB07, CA05, CV08, CMS09, DBL06, Dre05, EM08, FJ00, GCF^{+}07, Gav04, GS01, GCM^{+}08, HR02, HM04, HC09, JLF09, Kau04, KSDBT02, KFNB09, LWSM03, LOM03, LBPBG07, LLWCW09, LA00, MKG03, MNRD^{+}08, MB01, MF00, MPR^{+}01, MM02b, MMB^{+}03, NS09, NRBM05, Nug08, OHC08, Orr08a, PA04, PZ01, PCFC01, RLC^{+}07, RA05, RCP00, RGGS02, SRR00, Say05, SGSA07, SS00, SCC^{+}09, TS09, VVMV09, WMM04, Wen08, Whi03, WMVK03]. Used [Cra02, DC06a, Hay09b]. Useful [OBM09]. Usefulness [RA04a]. Uses [Kal05]. Using [BW08a, Ber04b, But03, CDD^{+}05, CB01b, DDMMZD06, DAB06, Dun02, FNFM01, FTN^{+}05, GBE^{+}06, HCK^{+}02, HPK^{+}04, HCZ00, Hod04a, HY03, KMGB07, LLW02, LHG^{+}03, LV04, Lus03, MBQAM06, NL05, NFB05, PC02, PR03, PMKDS01, PEC01, PAHM04, RAL^{+}04, RBHKK02, SBM06, Sch05, Sch00, SH03, TCSD00, TS06, WD01, WL04, Wil03]. Uso [BGC^{+}04, BMB^{+}08, CHW^{+}01, CES05, CS06, CMM^{+}05b, CSB07, DBL06, EM08, GCF^{+}07, Gav04, GS01, HR02, JLF09, Kau04, KSDBT02, KFNB09, LWSM03, LBPBG07, LLWCW09, LHG^{+}03, MKG03, MNRD^{+}08, MB01, MF00, MPR^{+}01, MMB^{+}03, NRBM05, PA04, PCFC01, RLC^{+}07, SRR00, Say05, SS00, SCC^{+}09, TS09, VVMV09, WMM04, WD01, WMVK03, ABE^{+}02b, Ant08, BBBM02, CA05, Dre05, FJ00, HC09, LV04, MM02b, NL05, OHC08, PC02, PR03, RMD^{+}02, RA05]. Usos [Kal05, RLK^{+}01]. Utah [WM05]. Útiles [OBM09]. Utilidad [CHF00, GM09, LCS04, RA04a, REC^{+}00, Rub06b]. Utility [AFK^{+}09, CHF00, GM09, LCS04, REC^{+}00, Rub06b, WBV^{+}05]. Utilización [BL01, Hod04a, LOM03, Lus03, CDD^{+}05, CB01b, CV08, CMS09, DDMMZD06, DAB06, Dun02, FMPB07, FTN^{+}05, GBE^{+}06, HPK^{+}04, HM04, HY03, KMGB07, MBQAM06, NS09, NFB05, PMKDS01, PEC01, PAHM04, RAL^{+}04, RCP00, RBHKK02, SGSA07, SH03, Wen08, WL04, WBV^{+}05]. Utilizadas [RCS^{+}03]. Utilizados [Hay09b]. Utilizando [CPSS03, HCK^{+}02, HR04c, LLW02, NL05, SAA^{+}01]. Utlización [TS06]. UV [DSJ02, PRL^{+}02, SWR00, SKP00, WG06]. UV-B [DSJ02, SWR00, SKP00, WG06, PRL^{+}02]. UV-B/pH [PRL^{+}02]. UV/pH [PRL^{+}02].

Vadeadoras [BF03]. Vagility [CF01]. Validación [FNF03, Sch03b]. Validation [FNF03]. Valle [FB08]. Valles [ANW^{+}08]. Valley [FB08]. Valleys [ANW^{+}08]. Valor [AHTM05, ARPEBM09, EAA^{+}09, GCF^{+}07, GRJB^{+}07, Gav04, HAG^{+}09, KDD^{+}04, LP08, MN03, Mur01, TKB^{+}06, BIK^{+}06, LKG06, MH01b, RDW01]. Valoración [RAG03, MLMB08]. Valorando [FLF04]. Valuable [Gre06]. Valuation [MLMB08, MB09a, RAG03]. Value [Ada06, AHTM05, ARPEBM09, BIK^{+}06, Cha08, DHW01, EAA^{+}09, GCF^{+}07, GRJB^{+}07, Gav04, Gra09, HAG^{+}09, KDD^{+}04, KRW^{+}00, LP08, LKG06, MN03, MH01b, Mur00, Mur01, PE01, Pea01, PD03b, RDW01, TKB^{+}06, ZR01]. Values [DZM08, Inf01, Moo07, Nos07, Pem01, Van04, WABT08]. Vanishing [Wes05]. Vaquita [DLCV00, JLRBB^{+}07]. Varamientos [LCS03]. varia [HMF^{+}04, PR03]. Variabilidad [GGA01, WSLW00, ZSMW06, AE01, AMF^{+}01, FSS^{+}00, RLC^{+}00, VWQI00]. Variability [AE01, AMF^{+}01, FSS^{+}00, GGA01, RLC^{+}00, VWQI00, WSLW00, ZSMW06]. Variable [LP08]. Variables [FNFM01, FNF03, LMP02, SDL07, VDP00, WM05]. Variación [BFB^{+}09, BJTD03, BSKH^{+}06, EBAW01, ENN^{+}08, KGS^{+}07, LM05, LDG^{+}00, PZG^{+}05, Pod01, TNRJ02, WRL02, YE07, BRS^{+}05, BS00, KF02, KPH07, LP07, MS04, MC05b, OBM09, Pod01, SL08, SBVC03, WBDP02]. Variaciones [LARK01]. Varían [LH04]. Variance [Kel01, MC05b]. Varianza [Kel01]. Variation [BFB^{+}09, BJTD03, BRS^{+}05, BSKH^{+}06, BS00, EBAW01, ENN^{+}08, KF02, KF03, KGS^{+}07, KPH07, LM05, LP07, LDG^{+}00, Lyn08, MS04, OBM09, PZG^{+}05, Pod01, RSC03, SL08, SBVC03, TNRJ02, WBDP02, WRL02, YE07]. Variations [LARK01]. Variegados [WFP01]. Variegated [WFP01]. variegatus [RNL^{+}04]. Vary [LH04]. Varying [LP08, Ova03]. Vascular [CDD^{+}05, Hie05, RVVNH09]. Vasculares [CDD^{+}05, Hie05]. Vast [BHW01]. Vaticinadores [OHC08]. Vectores [KZA09]. Vectors [KZA09]. Vegetación [BTGS05, CLC^{+}08, DGSS03, FJ00, FFHW03, FBP^{+}09, HWH^{+}02, KBR03, ML02, ND03, PMC01, RLK^{+}01]. Vegetal [HH03]. Vegetales [KL06]. Vegetation [BTGS05, CLC^{+}08, DGSS03, FFHW03, FBP^{+}09, HH03, HWH^{+}02, KBR03, KL06, ML02, ND03, PMC01, RLK^{+}01]. Vehicles [VK07]. Vehículos [CFH^{+}02, VK07]. Venado [CAFL03]. Venados [ASM05, FMDK09, MR00b]. Venezolanos [ABE^{+}02b]. Venezuela [Rom03]. Venezuelan [ABE^{+}02b]. Ventajas [SHJ^{+}09, WELCH09]. Veracruz [SMFPCA05, PMEH05, SMFPCA05]. Verbascum [PRL03]. Verde [YH01]. Verdes [HKM06]. Verificación [BJ05]. Verifying [BJ05]. veris [LSL^{+}06]. Vernal [CMCG07]. Vernales [CMCG07]. Versus [Orr05a, TS09, Ano07m, BRYK01, DCD^{+}02, Des06, FRD^{+}09, Fol06, FP07, Ful02, GM09, JFWP06, Kni07, LE02, MSB^{+}01, Mor00b, PZG^{+}05, RM01, RSCL04, RS02c, San06, SL00a, SNM00, VSO04]. Vertebrados [BHMC01, CLW^{+}09, Ehr06, MBB^{+}03, Per00, Per01, PL03, RGW09, RSB05]. Vertebrate [BHMC01, CLW^{+}09, Ehr06, Per00, RGW09]. Vertebrates [MBB^{+}03, Per01, PL03, RSB05]. Vertedero [Fau01, Mur01]. Vertederos [NFB05]. Vessel [VT09]. Veterinary [Oso00]. via [WSCP05]. Viabilidad [ARMH04, AE02, CV08, HCK^{+}02, HR02, HLSZ02, HJ05, KCK^{+}08, MPDT01, MAP03, NLCAR09, QAMW03, RMD^{+}02, SH03, SLCW08, TWGH00, WBSG07, ZSMW06, Gar03a, KD00]. Viability [ARMH04, AE02, Bro00, CV08, DZM08, EFLW02, Gar03a, GRG^{+}05, HCK^{+}02, HR02, HLSZ02, HJ05, KD00, KCK^{+}08, MBQAM06, MPDT01, MAP03, MM04, Men03, NLCAR09, QAMW03, RMD^{+}02, SH03, SLCW08, TWGH00, WABT08, WBSG07, ZSMW06]. Viable [Ano04e, FO06, ROB^{+}04, STS05]. Viables [STS05]. Vicosa [RSD03]. Victoria [FPF00, LO06]. Vicuña [SVV07]. Vicuñas [SVV07]. Vid [BCS00]. Vida [BKV^{+}05, BM05b, BL01, Bra03a, BMMW07, CVS^{+}09, CWCG02, CHFR09, Cze05, DKW01, DGP^{+}02, FDBM04, KD05, LLS00, Mad05, MBL02, MNRD^{+}08, MDH^{+}07, MOO09, OHC08, OFN^{+}06, PBSR06, PLL^{+}05, RHA^{+}05, RAL^{+}04, SRW^{+}08, SFD09, TWW09, UMT^{+}01, Wed02, WSBP09, XGS^{+}09, CW00, MSWF00]. Viejos [BRA08, MML06, RS02c, SHYH06, SDJ06]. View [DAB^{+}04, Foo04, Gil04, O’K01b, Orr02a, Spe00]. Views [Cor06, Sar06]. Vigor [RNL^{+}04]. Village [Mad05]. Villard [Wil09a]. Vinculación [PCHN04]. Vínculos [BGHM06]. Vine [BCS00]. Viñedos [HM04]. Vineyards [HM04]. Viral [CMK^{+}07]. Virales [CMK^{+}07]. virginianus [PLL^{+}05]. Virtual [All09, HAG^{+}09, PSF05, Wes05]. Virtuales [PSF05]. Virtud [Van06]. Virtue [Pal05, Van06]. Virus [KDG^{+}06, Mal03]. Visión [MCM^{+}00]. Visioning [Wal06]. Visita [WBW06]. Visitados [WBW05]. Visitan [KSDBT02]. Visitantes [BM05a]. Visitation [WBW06]. Visited [WBW05]. Visiting [KSDBT02]. Visitor [Bal00, War00]. Visitors [BM05a]. Visualization [Par08]. Vital [MM04]. Vitales [MM04]. Vivas [CMC03, SVV07, WSCP05]. Vive [MH02]. Vivienda [LKG06]. Viviendas [PCL08]. Viviente [CLW^{+}09]. Vivíparos [DDMMZD06]. Vivos [RGW09]. Vocalizaciones [GB01]. Voice [Ben00, Ben01b]. Void [Kel08]. Voladoras [RMN00]. Voladores [MBC03]. Volcanic [PJYO04]. Volcánicos [PJYO04]. Voles [GM00a, TM04]. Volume [Ano04a, Ano04l, Ano05a, Ano05l, Ano06a, Ano06k, Ano07a, Ano08b, Ano08r, Ano09a, Ano09o, LD07]. Voluntaria [MYH09, VT09]. Voluntarios [SHJ^{+}09, WMPC08, GPZ04]. Voluntary [MYH09, VT09, WMPC08]. Volunteer [SHJ^{+}09]. Volunteer-Based [SHJ^{+}09]. Volunteers [GPZ04, Whi03]. vs [ZR01]. Vulnerabilidad [DR02, HSMK07, ND04, NCVBW02, SCC^{+}09, TPW03, WCX09, ZG05]. Vulnerability [DR02, HSMK07, ND04, NCVBW02, SCC^{+}09, TPW03, WCX09, ZG05]. Vulnerable [RAR^{+}04, ZG05]. Vulnerables [PLB09, ZG05]. Vultures [PCD^{+}03].

Wadden [Wol00]. Wading [BF03]. Wait [RS02a, Tro06]. Walk [Ehr03]. Walking [Orr03c]. Want [Dor04]. Wanted [Bro05, LJI^{+}08, Smi04a, Wad02]. Waples [NVP07, PD00]. War [DGP^{+}02, Ehr00, Fre03, McN03, Owe08]. Warbler [FCL00]. Warfare [HBD^{+}09]. Warmer [WMLS09]. Warming [CBH01, GB01, GK06, GD05, HJ05, Kop01, MLN^{+}06, RVVNH09]. Warning [Bra05a, LVM02]. Wars [GCBF09]. Washington [VDP00, VDP00]. Wasps [FBP^{+}09, KSDBT02, KT02, PAHM04, SD03]. Watching [Cor04b, HL07a, WMPC08]. Water [All09, BBA^{+}02, CMC^{+}01, Owe08, Per02, RFD01, XGS^{+}09]. Waterbirds [RS02b]. Watercraft [RS02b]. Waters [ATG05, Bar03, DGC05, Lei00, NHHG^{+}05]. Watershed [Kni07, SH02, WD03]. Watersheds [ABE^{+}02b]. Watson [WP03]. Wattled [PB04]. Way [ACD^{+}07]. Ways [AFK^{+}09, Dor06, Fie06]. We-Know-Best [Bra05b]. Weak [GMMPP08, Whi02]. Wealth [Dal02, WMCB02, WSA^{+}05]. Web [CVS^{+}09, HCRM09, NC03b]. Webs [Cor00, LAN05, STP09, VSC06]. Wedekind [CW09]. Weed [PRL03, Sch07, VS09]. Weed-Management [VS09]. Weekend [BGC^{+}04]. Weevil [LO02]. Weight [CT07]. Weighting [AHTM05]. Weitzman [RMSG^{+}03]. Welcome [Sal02]. Welfare [CPR^{+}07, WMD^{+}06]. Well [ADD^{+}07, Kot00, MKH07, SBM06]. Well-Being [ADD^{+}07, SBM06]. Well-Designed [MKH07]. Were [Pre06]. West [PPLS01, BVK02, BBB^{+}07, Bra03a, Cur03, HVR08, Joh01, KDG^{+}06, LS01, Mal03, OALM^{+}00, Sey04]. West-East [Sey04]. Western [Cra02, GM00b, Kau04, RF00, BRK^{+}04, BKWB05, CWNW01, DDC^{+}00, DMDFV08, HR04c, Kee06, KCS04, LCS03, MMB^{+}03, OFS^{+}04, RWS02, RCP00, SK08, TRCC^{+}05]. Westslope [ALH^{+}04, CK05]. Wetland [Elp00, FB00, GWS^{+}06, Gib00, HF04, JGM03a, JMH01, MJ03, RH08, SH02, SKBB00]. Wetlands [Bee09, HTM^{+}05, Mar05, PM05, SB03, SMC00]. Whale [Cor04b, HL07a, REC^{+}00, SRDG03, Tay03b, WMPC08]. Whales [SDRO01, VT09, WMPC08]. Whaling [HL07a]. Wheels [DPB^{+}09]. Where [Ben00, Ben01b, Bre02b, Cob03, Gal03, Jac08, MH02, Pre06, VBRNP08, BL03, VBB^{+}08]. Whether [NV09, WABT07a, AP00]. While [FPL^{+}04]. White [Côt05, HPK^{+}04, Ols07, Wie07]. White-Naped [HPK^{+}04]. White-Tailed [Côt05]. Whither [OS01]. Whittaker [WDL^{+}02]. Whitten [KO01]. Who [Har06a, HTU^{+}06, MHGF04, MS09, Rei05b, ZR01, KO01]. Whole [Dix07, HP06, Sha04]. Whom [Get05, KO01, Mei06, MHGF04]. Whooping [JGL^{+}02]. Wide [BMA^{+}08, Gui07, NC03b, PSGA09, WE07]. Widespread [DDK05, Mcc05, OMP^{+}05, VVMV09]. Wild [Ano02j, AAO^{+}07, Ben02, Cre01a, CB01b, Din03b, Fle03, For02, Fra09, Gal03, GA02, GBB^{+}08, HGH^{+}00, JGM03b, KB03, KLH08, LCZ^{+}05, LADM05, MCA^{+}00, NKBD02, OBM09, PJNA07, Pre06, SRR00, SRW06, SFD09, SMFPCA05, TTSP04, TAP^{+}01]. Wild-Born [TAP^{+}01]. Wild-Meat [Ano02j]. Wilderness [DGC05, NHHG^{+}05, Slo02, WM01]. Wildfire [Bur04, DWW04, JMB01, Mar02a, MGPM04]. Wildland [BMS08]. Wildlands [Nos03]. Wildlife [Ano08l, BKV^{+}05, BM05b, BM06, BL01, BD05, CW00, CWCG02, CHFR09, Cze05, DKW01, DGP^{+}02, ESSGL00, FF08, Ful00, GCM^{+}08, Gru08, GNR01, KNH07, LLS00, Mad05, MBL02, MNRD^{+}08, MDH^{+}07, MOO09, NTMT^{+}03, OFN^{+}06, PBSR06, PLL^{+}05, RHA^{+}05, RAL^{+}04, SRW^{+}08, TA06, TWW09, UMT^{+}01, VS05a, WSBP09]. Wildness [Lee09]. Will [MV06, MF00]. Willamette [FB08, FB08]. Willard [Lin03a]. Willers [CCM00a]. Willi [KB05]. Willow [SK02b]. Wilson [HM09]. Wind [PPW09]. Winners [WCX09]. Wintering [JSHM06]. Wisconsin [BJTD03, HR04a, RRH^{+}09]. Wisdom [Alt07, Ful00]. Wit [Alt07]. within [Ant08, BFB^{+}09, Bea09a, BS00, DS08, Gal01b, GSB^{+}06, Ker09, KM00, MA08, RCMV06, SRH^{+}07, VW04, VW05]. Without [Hun07a, Ano02i, Moo02, PZ01, PJB^{+}06, TPW03, WG08]. Witness [Caf00c]. Wolf [BPB^{+}05, CPSS03, HLSZ02, MMB^{+}03, NTGT03, PPRB06, TNTH^{+}04]. Wolverine [FHL^{+}04]. Wolverton [MS02]. Wolves [CFH^{+}02, Dor04, FH06, RLC^{+}00]. Won [LS01]. Wood [DD04b, Fau01, JMS06b, PPLS01, SLS05, SFS00, SBB^{+}08, WG06]. Wood-Rotting [SLS05]. Wooded [FJ00]. Woodland [Cun00, GG02, MNB09, MM07]. Woodlot [SLKP06]. Woodlots [NFB05]. Woodpecker [Lam04, Sta03]. Woodpeckers [MGC01]. Woodrat [Log03].

Woody [KPJ06, LGK02, NBM^{+}04, YE07]. Words [PMEC04]. Work [Gro05a, Kot00, MH02, MS09, Mul09, Pan09, Tut03, Yah04]. Worked [GK08]. Workplace [MS09]. World [All09, BCMA^{+}01, Bro04, Caf04, CRGG01, GBB^{+}08, GX09, HJ05, JWB06, KORB02, NC03b, PCC03, PC01, Rob06b, Ter04, Woo03, And05, GBC^{+}05]. Worldwide [TK03a]. Worse [JGM03b]. Worst [Ibi08]. Worst-Case [Ibi08]. Worth [Tro06]. Worthy [MB00]. Wright [Sha01]. Writing [Ano00d, Kni06]. Written [PDS^{+}07]. Wrong [BBA^{+}02]. Wrongs [Ano00d, Kni06].

Xerófilas [BKK03]. Xerófitos [RAW07]. Xerophilous [BKK03]. Xishuangbanna [JHHL04]. Xu [Pan09]. XVI [Rom03].

Year [CBG^{+}00, LLV^{+}02, New06]. Years [GV02, Nug08, PDS^{+}07, SPM05, SSY06, TNTH^{+}04]. Yellow [KM00, KSB01]. Yellow-Legged [KM00]. Yellowstone [NCVBW02, NCVBW02]. York [Ano03a, DD04b, DD04b, GB01]. Young [Gra09, RSCL04, SL00a, DGG^{+}01]. Yukon [WU05].

Zacharias [SRSP01, ZR01]. Zealand [AE02, Cho06, CAFL03, DA07, GLV05, JRL03, JWB06, LB05, LOM03, LLW02, RNP^{+}07, SSB^{+}08, TPW03, Tow03]. Zebra [BJPL07b]. Zelanda [AE02, Cho06, CAFL03, DA07, JRL03, JWB06, LB05, LOM03, LLW02, RNP^{+}07, SSB^{+}08, TPW03, Tow03]. Zhoushan [Che09, Che09]. Zingel [LG05a]. Zona [HKN01, LF00, SS00, FD00]. Zonas [FPSD09, GG02, HRS08, KF01a, RLK^{+}01, SK04, SB03, Vee00, WD03]. Zone [FD00, RS02b, SK04]. Zones [Ara02, HRS08, SB03, WD03]. Zonificación [VTA02]. Zoning [Ste08a, VTA02]. Zoo [Bal00, War00]. Zoogoneticus [DDBA^{+}07]. Zoológicos [MCR^{+}04]. Zoos [MCR^{+}04, Swa09]. Zorro [HDA01]. Zorros [MBC03, ST05]. Zorzalito [Fau01, SFS00]. Zululand [MB00, MB00].

References

Araki:2007:RSC

[AAO^{+}07] Hitoshi Araki, William R. Ardren, Erik Olsen, Becky Cooper, and Michael S. Blouin. Reproductive success of captive-bred steelhead trout in the wild: Evaluation of three hatchery programs in the Hood River. Éxito reproductivo en el medio natural de truchas Oncorhynchus mykiss criadas en cautiverio: Evaluación de tres programas de criaderos de peces en el Río Hood. Conservation Biology, 21(1):181–190, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abrams:2009:ITB

[AAO^{+}09] R. W. Abrams, Eno D. Anwana, Alison Ormsby, Delali B. K. Dovie, Ademola Ajagbe, and Amber Abrams. Integrating topdown with bottom-up conservation policy in Africa. Integración de la conservación arriba-abajo con la conservación abajo-arriba en África. Conservation Biology, 23(4):799–804, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Litt:2003:PSR

[aARLG03] and Andrea R. Litt and Doria R. Gordon. Predictors of species richness in northwest Florida longleaf pine sandhills predictores de riqueza de especies en colinas de pinos del noroeste de Florida. Conservation Biology, 17(6):1660–1671, December 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:2001:CIN

[ABE01] Gregory J. Anderson, Gabriel Bernardello, and Michael S. Engel. Conservation implications of a newly discovered bee species on Isla Robinson Crusoe, Chile. Conservation Biology, 15(3):803– 805, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abell:2002:CBB

[Abe02a] Robin Abell. Conservation biology for the biodiversity crisis: a freshwater follow-up. Conservation Biology, 16(5):1435–1437, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allan:2002:LUW

[ABE^{+}02b] J. D. Allan, A. J. Brenner, J. Erazo, L. Fernandez, A. S. Flecker, D. L. Karwan, Samuel Segnini, and D. C. Taphorn. Land use in watersheds of the Venezuelan Andes: a comparative analysis. Uso del suelo en cuencas de los Andes Venezolanos: un análisis comparativo. Conservation Biology, 16(2):527–538, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andreone:2005:SRA

[ACC^{+}05] Franco Andreone, John E. Cadle, Neil Cox, Frank Glaw, Ronald A. Nussbaum, Christopher J. Raxworthy, Simon N. Stuart, Denis Vallan, and Miguel Vences. Species review of amphibian extinction risks in Madagascar: Conclusions from the global amphibian assessment. Revisión del riesgo de extinción de anfibios en Madagascar: Conclusiones de evaluación global de anfibios. Conservation Biology, 19(6):1790–1802, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armsworth:2007:ESS

[ACD^{+}07] P. R. Armsworth, K. M. A. Chan, G. C. Daily, P. R. Ehrlich, C. Kremen, T. H. Ricketts, and M. A. Sanjayan. Ecosystemservice science and the way forward for conservation. Conservation Biology, 21(6):1383–1384, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arriaga:2004:PED

[ACMA04] Laura Arriaga, Alejandro E. Castellanos, Elizabeth Moreno, and Jesús Alarcón. Potential ecological distribution of alien invasive species and risk assessment: a case study of buffel grass in arid regions of Mexico. Distribución ecológica potencial de especies exóticas invasoras y evaluación de riesgo: un estudio de caso de Cenchrus ciliaris en regiones Áridas de México. Conservation Biology, 18(6):1504–1514, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Acosta:2000:ICE

[Aco00] Charles A. Acosta. International conservation education. Conservation Biology, 14(4):924, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adams:2006:TBH

[Ada06] Rod D. Adams. Thinking beyond human value. Conservation Biology, 20(5):1557–1558, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adams:2008:ANC

[Ada08] Jonathan Adams. Anticipating the next challenge. Conservation Biology, 22(2):501, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arrow:2007:CIF

[ADD^{+}07] Kenneth J. Arrow, Gretchen Daily, Partha Dasgupta, Paul Ehrlich, Lawrence Goulder, Geoffrey Heal, Simon Levin, Karl-Göran Mäler, Stephen Schneider, David Starrett, and Brian Walker. Consumption, investment, and future well-being: Reply to Daly et al. Conservation Biology, 21(5):1363–1365, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alexandre:2004:PDC

[ADF04] José Alexandre and Felizola Diniz-Filho. Phylogenetic diversity and conservation priorities under distinct models of phenotypic evolution. Diversidad filogenética y prioridades de conservación bajo diferentes modelos de evolución fenotípica. Conservation Biology, 18(3):698–704, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alexander:2001:CVR

[AE01] Michael A. Alexander and Jon K. Eischeid. Climate variability in regions of amphibian declines. Variabilidad climática en regiones de declinaciones de anfibios. Conservation Biology, 15 (4):930–942, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2002:DVN

[AE02] Doug P. Armstrong and John G. Ewen. Dynamics and viability of a New Zealand robin population reintroduced to regenerating fragmented habitat. Dinámica y viabilidad de una población de petirrojo de Nueva Zelanda reintroducida en un fragmento de hábitat en regeneración. Conservation Biology, 16(4):1074–1085, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Akcakaya:2000:MCI

[AFB^{+}00] H. Reşit Akçakaya, Scott Ferson, Mark A. Burgman, David A. Keith, Georgina M. Mace, and Charles R. Todd. Making consistent IUCN classifications under uncertainty establecimiento de clasificaciones consistentes de IUCN bajo incertidumbre. Conservation Biology, 14(4):1001–1013, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Auster:2009:DOE

[AFK^{+}09] Peter J. Auster, Rod Fujita, Stephen R. Kellert, John Avise, Claudio Campagna, Benjamin Cuker, Paul Dayton, Burr Heneman, Richard Kenchington, Greg Stone, Giuseppe Notarbartolo Di Sciara, and Polita Glynn. Developing an ocean ethic: Science, utility, aesthetics, self-interest, and different ways of knowing.

Conservation Biology, 23(1):233–235, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anadon:2009:ELE

[AGBP09] José Daniel Anadón, AndréS Giménez, Rubén Ballestar, and Irene Pérez. Evaluation of local ecological knowledge as a method for collecting extensive data on animal abundance. Evaluación del conocimiento ecológico local como método para recolectar datos extensivos sobre la abundancia animal. Conservation Biology, 23(3):617–625, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ayroles:2009:GAI

[AHR^{+}09] Julien F. Ayroles, Kimberly A. Hughes, Kevin C. Rowe, Melissa M. Reedy, Sandra L. Rodriguez-Zas, Jenny M. Drnevich, Carla E. Cáceres, and Ken N. Paige. A genomewide assessment of inbreeding depression: Gene number, function, and mode of action. Una evaluación a nivel de genoma de la depresión por endogamia: Número, función y modo de acción de genes. Conservation Biology, 23(4):920–930, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arponen:2005:VBR

[AHTM05] Anni Arponen, Risto K. Heikkinen, Chris D. Thomas, and Atte Moilanen. The value of biodiversity in reserve selection: Representation, species weighting, and benefit functions. El valor de la biodiversidad en la selección de reservas: Representación, Ponderación de especies y funciones de beneficio. Conservation Biology, 19(6):2009–2014, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Amaral:2005:CMC

[AJ05] Antônia Cecília Z. Amaral and Sílvio Jablonski. Conservation of marine and coastal biodiversity in Brazil. Conservación de la biodiversidad marina y costera en Brasil. Conservation Biology, 19(3):625–631, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aipanjiguly:2003:CMK

[AJF03] Sampreethi Aipanjiguly, Susan K. Jacobson, and Richard Flamm. Conserving manatees: Knowledge, attitudes, and intentions of boaters in Tampa Bay, Florida. Conservación de manatíes: Conocimiento, actitudes e intenciones de motoristas en Tampa Bay, Florida. Conservation Biology, 17(4):1098–1105, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arponen:2007:ABR

[AKM07] Anni Arponen, Hanna Kondelin, and Atte Moilanen. Area-based refinement for selection of reserve sites with the benefit-function approach. Refinamiento basado en Área para la selección de sitios de reserva con el método de función de beneficio. Conservation Biology, 21(2):527–533, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allan:2003:EFF

[AKO03] Brian F. Allan, Felicia Keesing, and Richard S. Ostfeld. Effect of forest fragmentation on Lyme disease risk. Efecto de la fragmentación de bosques sobre el riesgo de la enfermedad de Lyme. Conservation Biology, 17(1):267–272, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allendorf:2003:IPB

[AL03] Fred W. Allendorf and Laura L. Lundquist. Introduction: Population biology, evolution, and control of invasive species. Conservation Biology, 17(1):24–30, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alagona:2008:C

[Ala08] Peter S. Alagona. Credibility. Conservation Biology, 22(6): 1365–1367, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allendorf:2004:IUE

[ALH^{+}04] Fred W. Allendorf, Robb F. Leary, Nathaniel P. Hitt, Kathy L. Knudsen, Laura L. Lundquist, and Paul Spruell. Intercrosses and the U.S. Endangered Species Act: Should hybridized populations be included as westslope cutthroat trout? Entrecruzas y el Acta de Especies en Peligro de E.U.A.: Inclusión de poblaciones híbridas de Oncorhynchus clarki lewisi. Conservation Biology, 18(5):1203–1213, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allan:2009:VWR

[All09] J. A. Tony Allan. Virtual water in the real world. Conservation Biology, 23(5):1331–1332, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Araujo:2007:EIP

[ALM07] Miguel B. Araújo, Jorge M. Lobo, and Juan C. Moreno. The effectiveness of Iberian protected areas in conserving terrestrial biodiversity. La efectividad de las Áreas protegidas Ibéricas en la conservación de la biodiversidad terrestre. Conservation Biology, 21(6):1423–1432, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anand:2005:DRA

[ALR05] Madhur Anand, Sophie Laurence, and Bronwyn Rayfield. Diversity relationships among taxonomic groups in recovering and restored forests. Relaciones de diversidad entre grupos taxonómicos en bosques en recuperación y restaurados. Conservation Biology, 19(3):955–962, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Altieri:2005:ILU

[Alt05] Miguel A. Altieri. Intimate levels of understanding. Conservation Biology, 19(4):1337–1338, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alter:2007:WW

[Alt07] S. Elizabeth Alter. Wit and wisdom. Conservation Biology, 21(4):1129, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alvarez:2002:ICB

[Álv02] Maria D. Álvarez. Illicit crops and bird conservation priorities in Colombia. Cultivos ilícitos y prioridades de conservación de aves en Colombia. Conservation Biology, 16(4):1086–1096, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alvarez:2005:GM

[Álv05] María D. Álvarez. Guerrillas in the midst. Conservation Biology, 19(6):2040–2043, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anders:2005:IAP

[AM05] Angela D. Anders and Matthew R. Marshall. Increasing the accuracy of productivity and survival estimates in assessing landbird population status. Incremento en la precisión de estimaciones de productividad y supervivencia en la evaluación del estatus de poblaciones de aves terrestres. Conservation Biology, 19(1):66–74, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andayani:2001:GVM

[AMF^{+}01] Noviar Andayani, Juan Carlos Morales, Michael R. J. Forstner, Jatna Supriatna, and Don J. Melnick. Genetic variability in mtDNA of the silvery gibbon: Implications for the conservation of a critically endangered species. Variabilidad genética en el ADNm del gibón plateado: Implicaciones para la conservación de una especie críticamente en peligro. Conservation Biology, 15(3):770–775, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:2001:TFG

[And01a] E. N. Anderson. Tropical forest game conservation. Conservation Biology, 15(3):791–792, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andrade:2001:CTC

[And01b] Germán I. Andrade. Conservation through confiscation. Conservation Biology, 15(5):1467–1469, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andrews:2001:PEP

[And01c] Richard N. L. Andrews. Politics, economics, and perceptions. Conservation Biology, 15(6):1485–1487, December 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andrade:2005:SSW

[And05] Germán I. Andrade. Science and society at the World Parks Congress. Conservation Biology, 19(1):4–5, February 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Angermeier:2000:NIB

[Ang00] Paul L. Angermeier. The natural imperative for biological conservation. El imperativo natural para la conservación biológica.

Conservation Biology, 14(2):373–381, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Angermeier:2001:RPA

[Ang01] Paul L. Angermeier. Reply to Povilitis: Another conservation conundrum. Conservation Biology, 15(2):536–538, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ankersen:2004:RPL

[Ank04] Thomas T. Ankersen. Regional perspectives on Latin American conservation. Conservation Biology, 18(5):1444–1446, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2000:BR

[Ano00a] Anonymous. Book reviews. Conservation Biology, 14(5):1555– 1564, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2000:GPH

[Ano00b] Anonymous. A global perspective on habitat disturbance and tropical rainforest mammals. Conservation Biology, 14(6):1574– 1579, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2000:S

[Ano00c] Anonymous. Summary. Conservation Biology, 14(3):702–704, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2000:WWD

[Ano00d] Anonymous. Writing the wrongs: Developing a safe-fail culture in conservation. Conservation Biology, 14(6):1567–1568, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:BRJ

[Ano01a] Anonymous. Books received July 2000 to July 15, 2001. Conservation Biology, 15(6):1829–1833, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:EAa

[Ano01b] Anonymous. Editor’s acknowledgments. Conservation Biology, 15(1):6, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:EAb

[Ano01c] Anonymous. Editor’s acknowledgments. Conservation Biology, 15(6):1479, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:Ec

[Ano01d] Anonymous. Errata. Conservation Biology, 15(4):1199, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:Ed

[Ano01e] Anonymous. Errata. Conservation Biology, 15(6):1833, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:Eb

[Ano01f] Anonymous. Erratum. Conservation Biology, 15(1):292, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:IC

[Ano01g] Anonymous. Islands at the crossroads. Conservation Biology, 15 (1):300–301, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:LHS

[Ano01h] Anonymous. Life in a hormone soup. Conservation Biology, 15 (1):298–300, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:PM

[Ano01i] Anonymous. Pessimism from manu. Conservation Biology, 15 (1):296–298, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:TFM

[Ano01j] Anonymous. Tropical forest management and conservation of biodiversity: an overview. Conservation Biology, 15(1):7–20, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:TLH

[Ano01k] Anonymous. Tropical logging and human invasions. Conservation Biology, 15(1):4–5, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2001:UTS

[Ano01l] Anonymous. The ultimate test of Society’s commitment. Conservation Biology, 15(1):293–294, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:BRJ

[Ano02a] Anonymous. Books received July 2001 to June 30, 2002. Conservation Biology, 16(6):1670–1674, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:CAF

[Ano02b] Anonymous. Coming of age in forest conservation. Conservation Biology, 16(1):272–274, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:EA

[Ano02c] Anonymous. Editor’s acknowledgments. Conservation Biology, 16(6):1456, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:E

[Ano02d] Anonymous. Erratum. Conservation Biology, 16(5):1448–1449, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:EJS

[Ano02e] Anonymous. The error of judgment: Struggling for neutrality in science and journalism. Conservation Biology, 16(6):1451–1453, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:FCS

[Ano02f] Anonymous. Four challenges of sustainability. Conservation Biology, 16(6):1457–1460, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:SS

[Ano02g] Anonymous. Spirituality and sustainability. Conservation Biology, 16(6):1461–1464, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:SSR

[Ano02h] Anonymous. Sustainability and spiritual renewal: the challenge of creating a politics of reverence. Conservation Biology, 16(6): 1466–1468, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:SSC

[Ano02i] Anonymous. Sustainability without spirituality: a contradiction in terms? Conservation Biology, 16(6):1465, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2002:WMU

[Ano02j] Anonymous. Wild-meat use, food security, livelihoods, and conservation. Conservation Biology, 16(3):580–583, June 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2003:SCB

[Ano03a] Anonymous. 2004 Society for Conservation Biology Annual Meeting, New York City. Conservation Biology, 17(5):vi, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2003:AI

[Ano03b] Anonymous. Author index. Conservation Biology, 17(6):1879– 1884, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2003:BRJ

[Ano03c] Anonymous. Books received to July 15, 2003. Conservation Biology, 17(6):1861–1864, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2003:IJA

[Ano03d] Anonymous. Important journal announcement: Electronic submittal of manuscripts. Conservation Biology, 17(6):1477, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2003:MCL

[Ano03e] Anonymous. Mapping the conservation landscape. Conservation Biology, 17(2):647, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2003:RRB

[Ano03f] Anonymous. Recently received books: (July 2002–February 2003). Conservation Biology, 17(4):1195, August 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2003:SI

[Ano03g] Anonymous. Subject index. Conservation Biology, 17(6):1865– 1878, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:AIV

[Ano04a] Anonymous. Author Index, Volume 18, 2004. Conservation Biology, 18(6):1712–1717, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:E

[Ano04b] Anonymous. Erratum. Conservation Biology, 18(1):289, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:EA

[Ano04c] Anonymous. Editor’s acknowledgments. Conservation Biology, 18(6):1456, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:IAT

[Ano04d] Anonymous. ICO approaches to tropical forest conservation. Conservation Biology, 18(6):1452, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:LEM

[Ano04e] Anonymous. Large estimates of minimum viable population sizes. Conservation Biology, 18(5):1178–1179, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:L

[Ano04f] Anonymous. Letter. Conservation Biology, 18(1):5–6, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:RRBe

[Ano04g] Anonymous. Recently received books (July 2003 –July 2004). Conservation Biology, 18(6):1701–1703, December 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:RRBc

[Ano04h] Anonymous. Recently received books (July 2003–March 2004). Conservation Biology, 18(4):1173, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:RRBa

[Ano04i] Anonymous. Recently received books July to September 2003. Conservation Biology, 18(1):288, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:RRBd

[Ano04j] Anonymous. Recently received books (March 2003–May 2004). Conservation Biology, 18(5):1448, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:RRBb

[Ano04k] Anonymous. Recently received books: (September–November 2003). Conservation Biology, 18(2):595, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2004:SIV

[Ano04l] Anonymous. Subject Index, Volume 18, 2004. Conservation Biology, 18(6):1704–1711, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:AIV

[Ano05a] Anonymous. Author Index, Volume 19, 2005. Conservation Biology, 19(6):2058–2065, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:BRJ

[Ano05b] Anonymous. Books received (July 2004 –July 2005). Conservation Biology, 19(6):2047–2050, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:CTC

[Ano05c] Anonymous. Conservation theology for conservation biologists — a reply to David Orr. Conservation Biology, 19(6):1689–1692, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:EA

[Ano05d] Anonymous. Editor’s acknowledgments. Conservation Biology, 19(6):1684, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:ENa

[Ano05e] Anonymous. Editor’s note. Conservation Biology, 19(6):1685, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:ENb

[Ano05f] Anonymous. Editor’s note. Conservation Biology, 19(6):2039, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:JSC

[Ano05g] Anonymous. The Journal of the Society for Conservation Biology: a global community of conservation professionals. Conservation Biology, 19(6):i, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:RRBc

[Ano05h] Anonymous. Recently received books (April 2004–June 2005). Conservation Biology, 19(5):1679, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:RRBb

[Ano05i] Anonymous. Recently received books (December 2004–January 2005). Conservation Biology, 19(3):985, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:RRBa

[Ano05j] Anonymous. Recently received books (October–November 2004). Conservation Biology, 19(2):586, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:SCB

[Ano05k] Anonymous. Society for conservation biology code of ethics. Conservation Biology, 19(1):3, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2005:SIV

[Ano05l] Anonymous. Subject Index, Volume 19, 2005. Conservation Biology, 19(6):2051–2057, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:AIV

[Ano06a] Anonymous. Author Index, Volume 20, 2006. Conservation Biology, 21(2):1844–1850, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:BRA

[Ano06b] Anonymous. Books received (August 2005 –August 2006). Conservation Biology, 21(2):1832–1835, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:CC

[Ano06c] Anonymous. Cover caption. Conservation Biology, 21(3):i, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:EA

[Ano06d] Anonymous. Editor’s acknowledgments. Conservation Biology, 20(6):1569, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:E

[Ano06e] Anonymous. Erratum. Conservation Biology, 33(5):1335, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:JSC

[Ano06f] Anonymous. The Journal of the Society for Conservation Biology: a global community of conservation professionals. Conservation Biology, 20(2):i, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:RRBd

[Ano06g] Anonymous. Recently received books (April 2006 –June 2006). Conservation Biology, 20(5):1562, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:RRBb

[Ano06h] Anonymous. Recently received books December 2005 to February 2006. Conservation Biology, 21(3):929, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:RRBc

[Ano06i] Anonymous. Recently received books (February 2006 –April 2006). Conservation Biology, 33(5):1334, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:RRBa

[Ano06j] Anonymous. Recently received books (October 2005 –December 2005). Conservation Biology, 20(2):594, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2006:SIV

[Ano06k] Anonymous. Subject Index, Volume 20, 2006. Conservation Biology, 21(2):1836–1843, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:AIV

[Ano07a] Anonymous. Author Index, Volume 21, 2007. Conservation Biology, 17(1):1674–1680, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:Eb

[Ano07b] Anonymous. Erratum. Conservation Biology, 21(2):895, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:Ec

[Ano07c] Anonymous. Erratum. Conservation Biology, 21(4):1131, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:BRA

[Ano07d] Anonymous. Books received (August 2006 — August 2007). Conservation Biology, 22(1):1663–1665, December 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:CFR

[Ano07e] Anonymous. Combining the fields of reintroduction biology and restoration ecology. Conservation Biology, 21(6):1387–1390, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:CCa

[Ano07f] Anonymous. Cover caption. Conservation Biology, 21(1):i, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:CCb

[Ano07g] Anonymous. Cover caption. Conservation Biology, 21(4):i, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:CCc

[Ano07h] Anonymous. Cover caption. Conservation Biology, 21(5):i, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:EAa

[Ano07i] Anonymous. Editor’s acknowledgments. Conservation Biology, 21(6):1391, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:Ea

[Ano07j] Anonymous. Errata. Conservation Biology, 21(2):566, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:Ed

[Ano07k] Anonymous. Errata. Conservation Biology, 22(1):1666, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:M

[Ano07l] Anonymous. In memoriam. Conservation Biology, 21(1):278– 280, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:OVH

[Ano07m] Anonymous. Optimism versus hope. Conservation Biology, 21 (6):1396, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:RRBe

[Ano07n] Anonymous. Recently received books (April 2007 –June 2007). Conservation Biology, 21(5):1382, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:RRBa

[Ano07o] Anonymous. Recently received books (August 2006 –November 2006). Conservation Biology, 21(1):285, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:RRBc

[Ano07p] Anonymous. Recently received books (December 2006 – February 2007). Conservation Biology, 21(2):895, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:RRBd

[Ano07q] Anonymous. Recently received books (February 2007 –April 2007). Conservation Biology, 21(4):1131, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2007:RRBb

[Ano07r] Anonymous. Recently received books (November 2006 – December 2006). Conservation Biology, 21(2):565, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:EOT

[Ano08a] Anonymous. At the end of our tether: the rationality of nonviolence. Conservation Biology, 22(2):235–238, April 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:AIV

[Ano08b] Anonymous. Author index, volume 22, 2008. Conservation Biology, 22(6):1680–1686, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:BRA

[Ano08c] Anonymous. Books received August 2007 –August 2008. Conservation Biology, 22(6):1668–1671, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:CCa

[Ano08d] Anonymous. Cover caption. Conservation Biology, 22(2):i, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:CCb

[Ano08e] Anonymous. Cover caption. Conservation Biology, 22(3):i, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:CCc

[Ano08f] Anonymous. Cover caption. Conservation Biology, 22(4):i, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:CCd

[Ano08g] Anonymous. Cover caption. Conservation Biology, 22(5):i, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:CCe

[Ano08h] Anonymous. Cover caption. Conservation Biology, 22(6):i, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:EA

[Ano08i] Anonymous. Editor’s acknowledgments. Conservation Biology, 22(6):1371, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:E

[Ano08j] Anonymous. Erratum. Conservation Biology, 22(1):228, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:I

[Ano08k] Anonymous. Introduction. Conservation Biology, 22(3):518– 520, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:MIW

[Ano08l] Anonymous. Monitoring international wildlife trade with coded species data. Conservation Biology, 22(1):4–7, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:RRBe

[Ano08m] Anonymous. Recently received books April–June 2008. Conservation Biology, 22(5):1363, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:RRBa

[Ano08n] Anonymous. Recently received books (August 2007 –December 2007). Conservation Biology, 22(1):227, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:RRBd

[Ano08o] Anonymous. Recently received books (February 2008–April 2008). Conservation Biology, 22(4):1085, August 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:RRBc

[Ano08p] Anonymous. Recently received books (January 2008 –March 2008). Conservation Biology, 22(3):809, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:RRBb

[Ano08q] Anonymous. Recently received books (November 2007 – December 2007). Conservation Biology, 22(2):503, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2008:SIV

[Ano08r] Anonymous. Subject index, volume 22, 2008. Conservation Biology, 22(6):1672–1679, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:AIV

[Ano09a] Anonymous. Author index, volume 23, 2009. Conservation Biology, 23(6):1664–1671, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:BRA

[Ano09b] Anonymous. Books received (August 2008 –August 2009). Conservation Biology, 23(6):1654–1655, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:CCa

[Ano09c] Anonymous. Cover caption. Conservation Biology, 23(1):i, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:CCb

[Ano09d] Anonymous. Cover caption. Conservation Biology, 23(2):i, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:CCc

[Ano09e] Anonymous. Cover caption. Conservation Biology, 23(3):i, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:CCd

[Ano09f] Anonymous. Cover caption. Conservation Biology, 23(4):i, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:CCe

[Ano09g] Anonymous. Cover caption. Conservation Biology, 23(5):i, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:CCf

[Ano09h] Anonymous. Cover caption. Conservation Biology, 23(6):i, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:EA

[Ano09i] Anonymous. Editor’s acknowledgments. Conservation Biology, 23(6):1348, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:RRBd

[Ano09j] Anonymous. Recently received books (April 2009 –June 2009). Conservation Biology, 23(5):1341, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:RRBa

[Ano09k] Anonymous. Recently received books August–October 2008. Conservation Biology, 23(1):239, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:RRBc

[Ano09l] Anonymous. Recently received books (January–April 2009). Conservation Biology, 23(4):1064, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:RRBb

[Ano09m] Anonymous. Recently received books (October–December 2008). Conservation Biology, 23(2):512, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:RRJ

[Ano09n] Anonymous. Recently received (January–February 2009). Conservation Biology, 23(3):784, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2009:SIV

[Ano09o] Anonymous. Subject index, volume 23, 2009. Conservation Biology, 23(6):1656–1663, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anthony:2008:UMT

[Ant08] Brandon P. Anthony. Use of modified threat reduction assessments to estimate success of conservation measures within and adjacent to Kruger National Park, South Africa. Uso de evaluaciones de Reducción de amenaza modificadas para estimar el Éxito de medidas de conservación dentro de y alrededor del Parque Nacional Kruger, África del Sur. Conservation Biology, 22 (6):1497–1505, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ayres:2008:EHT

[ANW^{+}08] Edward Ayres, Johnson N. Nkem, Diana H. Wall, Byron J. Adams, J. E. Barrett, Emma J. Broos, Andrew N. Parsons, Laura E. Powers, Breana L. Simmons, and Ross A. Virginia. Effects of human trampling on populations of soil fauna in the McMurdo Dry Valleys, Antarctica. Efectos del pisoteo de humanos sobre poblaciones de fauna del suelo en los Valles Secos McMurdo, Antártida. Conservation Biology, 22(6):1544–1551, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Agrawal:2006:PSC

[AO06] Arun Agrawal and Elinor Ostrom. Political science and conservation biology: a dialog of the deaf. Conservation Biology, 20(3):681–682, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arroyo:2009:GBC

[AOGMS09] Teresa P. Feria Arroyo, Mark E. Olson, Abisaí García-Mendoza, and Eloy Solano. A GIS-Based comparison of the Mexican National and IUCN methods for determining extinction risk. Una comparación basada en SIG de los métodos nacionales mexicanos y de la IUCN para la determinación del riesgo de extinción. Conservation Biology, 23(5):1156–1166, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2000:ETW

[AP00] Doug P. Armstrong and John K. Perrott. An experiment testing whether condition and survival are limited by food supply in a reintroduced hihi population. Prueba experimental para determinari si la condición Física y la sobrevivencia son limitadas por el suministro de alimento en una población reintroducida de hihi. Conservation Biology, 14(4):1171–1181, August 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abdelkrim:2005:IAP

[APCS05] Jawad Abdelkrim, Michel Pascal, Claire Calmet, and Sarah Samadi. Importance of assessing population genetic structure before eradication of invasive species: Examples from insular Norway rat populations. Importancia de la Estimación de la estructura genética poblacional antes de la erradicación de especies invasoras: Ejemplos con poblaciones insulares de Rattus norvegicus. Conservation Biology, 19(5):1509–1518, October 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aplet:2001:TR

[Apl01] Gregory H. Aplet. The three R’s. Conservation Biology, 15(3): 810–811, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Applegate:2003:CIN

[App03] Roger D. Applegate. Compensation issues in natural resource professions. Conservation Biology, 17(6):1475, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abdelkrim:2007:ECE

[APS07] Jawad Abdelkrim, Michel Pascal, and Sarah Samadi. Establishing causes of eradication failure based on genetics: Case study of ship rat eradication in Ste. Anne Archipelago. Establecimiento de las causas de la erradicación fallida con base en genética: Estudio de caso de la erradicación de rattus rattus en el Archipiélago Ste. Anne. Conservation Biology, 21(3):

719–730, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adams:2002:CGE

[APSB02] Jonathan M. Adams, Gianluca Piovesan, Steve Strauss, and Sandra Brown. The case for genetic engineering of native and landscape trees against introduced pests and diseases. El caso de la Ingeniería genética de Árboles nativos y de paisaje contra plagas y enfermedades introducidas. Conservation Biology, 16 (4):874–879, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [Hod04b].

Antonelli:2009:BSF

[AR09] Alexandre Antonelli and Victor Rodriguez. Brazil should facilitate research permits. Conservation Biology, 23(5):1068–1069, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Araujo:2002:BHZ

[Ara02] Miguel B. Araújo. Biodiversity hotspots and zones of ecological transition. Conservation Biology, 16(6):1662–1663, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Araujo:2009:SCP

[Ara09a] Miguel B. Araújo. Systematic conservation planning comes of age. Conservation Biology, 23(5):1332–1333, October 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Asner:2009:CAC

[ARA^{+}09b] Gregory P. Asner, Thomas K. Rudel, T. Mitchell Aide, Ruth Defries, and Ruth Emerson. A contemporary assessment of change in humid tropical forests. Una evaluación contemporánea del cambio en bosques tropicales Húmedos. Conservation Biology, 23(6):1386–1395, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Angulo:2007:DAE

[ARB^{+}07] Elena Angulo, Gary W. Roemer, Luděk Berec, Joanna Gascoigne, and Franck Courchamp. Double Allee effects and extinction in the island fox. Efectos Allee dobles y extinción en Urocyon littoralis. Conservation Biology, 21(4):1082–1091, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Archer:2009:MD

[Arc09] David Archer. Meet Dave. Conservation Biology, 23(6):1647, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2002:FSS

[Arm02] Doug Armstrong. Focal and surrogate species: Getting the language right. Conservation Biology, 16(2):285–286, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2005:IMH

[Arm05] Doug P. Armstrong. Integrating the metapopulation and habitat paradigms for understanding broad-scale declines of species. Integración de los paradigmas de metapoblación y hábitat para entender la declinación de especies a gran escala. Conservation Biology, 19(5):1402–1410, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Akcakaya:2004:ILM

[ARMH04] H. Reşit Akçakaya, Volker C. Radeloff, David J. Mladenoff, and Hong S. He. Integrating landscape and metapopulation modeling approaches: Viability of the sharp-tailed grouse in a dynamic landscape. Integrando métodos basados en modelos de paisaje y metapoblaciones: Viabilidad de Tympanuchus phasianellus en un paisaje dinámico. Conservation Biology, 18(2):526–537, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armbrecht:2005:RDC

[ARP05] Inge Armbrecht, Leonardo Rivera, and Ivette Perfecto. Reduced diversity and complexity in the leaf-litter ant assemblage of Colombian coffee plantations. Reducción en la diversidad y complejidad del ensamble de hormigas de la hojarasca en plantaciones de café colombianas. Conservation Biology, 19(3):897– 907, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arroyo-Rodriguez:2009:VSP

[ARPEBM09] Víctor Arroyo-Rodríguez, Eduardo Pineda, Federico Escobar, and Julieta Benítez-Malvido. Value of small patches in the conservation of plant-species diversity in highly fragmented rainforest. El valor de parches Pequeños para la conservación de la diversidad de especies de plantas en un bosque lluvioso sumamente fragmentado. Conservation Biology, 23(3):729–739, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abbitt:2001:EDB

[AS01] Robbyn J. F. Abbitt and J. Michael Scott. Examining differences between recovered and declining endangered species. Examinando las diferencias entre especies en peligro recuperadas y en declinación. Conservation Biology, 15(5):1274–1284, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ascencio:2002:ADC

[Asc02] Pedro F. Quintana Ascencio. Assessing demographic change for conservation purposes. Conservation Biology, 16(2):566–568, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Artman:2001:PBR

[ASD01] Vanessa L. Artman, Elaine K. Sutherland, and Jerry F. Downhower. Prescribed burning to restore mixed-oak communities in Southern Ohio: Effects on breeding- bird populations. Quema prescrita para restaurar comunidades mixtas de encino en el sur de Ohio: Efectos sobre poblaciones de aves reproductivas. Conservation Biology, 15(5):1423–1434, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Araguas:2009:RGR

[ASF^{+}09] Rosa M. Araguas, Núria Sanz, Raquel Fernández, Fred M. Utter, Carles Pla, and José-Luis García-Marín. Role of genetic refuges in the restoration of native gene pools of brown trout. El papel de los refugios genéticos en la restauración de acervos genéticos nativos de trucha común. Conservation Biology, 23(4):871–878, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allombert:2005:NEI

[ASM05] Sylvain Allombert, Steve Stockton, and Jean-Louis Martin. A natural experiment on the impact of overabundant deer on forest invertebrates un experimento natural del impacto de la sobreabundancia de venados sobre invertebrados de bosque. Conservation Biology, 19(6):1917–1929, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Asquith:2001:MCB

[Asq01] Nigel M. Asquith. Misdirections in conservation biology. Dirección errónea en conservación biológica. Conservation Biology, 15(2):345–352, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alo:2005:EHF

[AT05] Dominique Alò and Thomas F. Turner. Effects of habitat fragmentation on effective population size in the endangered Rio Grande silvery minnow. Efectos de la fragmentación del hábitat sobre el tamaño poblacional efectivo de Hybognathus amarus. Conservation Biology, 19(4):1138–1148, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Agostinho:2005:CBB

[ATG05] Angelo A. Agostinho, Sidinei M. Thomaz, and Luiz C. Gomes. Conservation of the biodiversity of Brazil’s inland waters. Conservación de la biodiversidad de las aguas interiores de Brasil. Conservation Biology, 19(3):646–652, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Araujo:2001:BCH

[AW01] Miguel B. Araújo and Paul H. Williams. The bias of complementarity hotspots toward marginal populations. El sesgo de la complementaridad de Áreas importantes par la biodiversidad (hotspots) hacia poblaciones marginales. Conservation Biology, 15(6):1710–1720, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andelman:2002:ACC

[AW02] Sandy J. Andelman and Michael R. Willig. Alternative configurations of conservation reserves for Paraguayan bats: Considerations of spatial scale. Configuraciones alternativas de reservas de conservación para murciélagos del Paraguay: Consideraciones de escala espacial. Conservation Biology, 16(5):1352–1363, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Awoyemi:2008:RRH

[Awo08] Stephen Mufutau Awoyemi. The role of religion in the HIV/ AIDS intervention in Africa: a possible model for conservation biology. Conservation Biology, 22(4):811–813, August 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Asoh:2004:DCC

[AYKM04] Kazue Asoh, Tomoko Yoshikawa, Randall Kosaki, and Elizabeth A. Marschall. Damage to cauliflower coral by monofilament fishing lines in Hawaii. Daño a coral coliflor por líneas de pesca de monofilamento en Hawai. Conservation Biology, 18(6): 1645–1650, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bini:2005:LLB

[BADF^{+}05] Luis Mauricio Bini, José Alexandre, Felizola Diniz-Filho, Priscilla Carvalho, Miriam Plaza Pinto, and Thiago Fernando L. V. B. Rangel. Lomborg and the litany of biodiversity crisis: What the peer-reviewed literature says. Lomborg y la letanía de la crisis de la biodiversidad: Lo que dice la literatura revisada por pares. Conservation Biology, 19(4):1301–1305, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2004:FRF

[BAF04] Richard T. Brown, James K. Agee, and Jerry F. Franklin. Forest restoration and fire: Principles in the context of place. Restauración de bosques y fuego — principios en el contexto del sitio. Conservation Biology, 18(4):903–912, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bagdassarian:2009:NAL

[Bag09] Carey K. Bagdassarian. Naturalists, artists, and language. Conservation Biology, 23(6):1639–1640, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Butchart:2006:BIB

[BAKHT06] Stuart H. M. Butchart, H. Resit Akcakaya, Elizabeth Kennedy, and Craig Hilton-Taylor. Biodiversity indicators based on trends in conservation status: Strengths of the IUCN red list index. Conservation Biology, 20(2):579–581, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balmford:2000:SFA

[Bal00] Andrew Balmford. Separating fact from artifact in analyses of zoo visitor preferences. Conservation Biology, 14(4):1193–1195, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [War00].

Baldwin:2001:BTC

[Bal01] Rob Baldwin. Biodiversity technology in the classroom. Conservation Biology, 15(6):1826–1827, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balmer:2002:SLE

[Bal02] Oliver Balmer. Species lists in ecology and conservation: Abundances matter. Conservation Biology, 16(4):1160–1161, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baldwin:2004:L

[Bal04] Robert F. Baldwin. Letters. Conservation Biology, 18(3):601, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barlow:2003:CFC

[Bar03] Michele Barlow. “Cool” fish in cold waters. Conservation Biology, 17(5):1464–1465, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Battin:2004:WGA

[Bat04] James Battin. When good animals love bad habitats: Ecological traps and the conservation of animal populations. Cuando animales buenos aman a hábitats malos: Trampas ecológicas y la conservación de poblaciones animales. Conservation Biology, 18(6):1482–1491, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bauer:2005:RTR

[Bau05] Gordon B. Bauer. Research training for releasable animals. Investigación para el entrenamiento de animales liberables. Conservation Biology, 19(6):1779–1789, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:2005:MTC

[Baw05] Kamaljit S. Bawa. More than a conservation assessment. Conservation Biology, 19(1):281–282, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:2006:GDL

[Baw06] Kamaljit S. Bawa. Globally dispersed local challenges in conservation biology. Conservation Biology, 20(3):696–699, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berg:2000:CGF

[BB00] David J. Berg and Paul H. Berg. Conservation genetics of freshwater mussels: Comments on Mulvey et al. Conservation Biology, 14(6):1920–1923, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [ML00].

Barrett:2001:CRN

[BB01] Gary W. Barrett and Terry L. Barrett. Cemeteries as repositories of natural and cultural diversity. Conservation Biology, 15 (6):1820–1824, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Batary:2004:EEE

[BB04] Péter Batáry and András Báldi. Evidence of an edge effect on avian nest success. Evidencia de un efecto de borde sobre el Éxito de nidos de aves. Conservation Biology, 18(2):389–400, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Becker:2006:CDT

[BB06] Benjamin H. Becker and Steven R. Beissinger. Centennial decline in the trophic level of an endangered seabird after fisheries decline. Declinación centenaria en el nivel trófico de una especie de ave marina en peligro después del colapso de las pesquerías. Conservation Biology, 20(2):470–479, April 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brook:2002:CPA

[BBA^{+}02] Barry W. Brook, Mark A. Burgman, H. Resit Akçakaya, Julian J. O’Grady, and Richard Frankham. Critiques of PVA ask the wrong questions: Throwing the heuristic baby out with the numerical bath water. Conservation Biology, 16(1):262–263, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2007:HCR

[BBB^{+}07] Elizabeth L. Bennett, Eric Blencowe, Katrina Brandon, David Brown, Robert W. Burn, Guy Cowlishaw, Glyn Davies, Holly Dublin, Julia E. Fa, E. J. Milner-Gulland, John G. Robinson, J. Marcus Rowcliffe, Fiona M. Underwood, and David S. Wilkie. Hunting for consensus: Reconciling bushmeat harvest, conservation, and development policy in West and Central Africa. Conservation Biology, 21(2):884–887, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bancroft:2008:MAE

[BBB08] Betsy A. Bancroft, Nick J. Baker, and Andrew R. Blaustein. A meta-analysis of the effects of ultraviolet B radiation and its synergistic interactions with pH, contaminants, and disease on amphibian survival. Un meta-análisis de los efectos de la radiación ultravioleta B y sus interacciones Sinergéticas con el pH, contaminantes y enfermedades sobre la supervivencia de anfibios. Conservation Biology, 22(4):987–996, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bani:2002:UFS

[BBBM02] Luciano Bani, Marco Baietto, Luciana Bottoni, and Renato Massa. The use of focal species in designing a habitat network for a lowland area of Lombardy, Italy. Uso de especies focales en el diseño de redes de hábitat para un Área baja de Lombardia, Italia. Conservation Biology, 16(3):826–831, June 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bertorelle:2004:NFB

[BBC^{+}04] Giorgio Bertorelle, Mike Bruford, Claudio Chemini, Cristiano Vernesi, and Heidi C. Hauffe. New, flexible Bayesian approaches to revolutionize conservation genetics. Conservation Biology, 18 (2):584, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowne:2006:CAL

[BBH06] David R. Bowne, Michael A. Bowers, and James E. Hines. Connectivity in an agricultural landscape as reflected by interpond movements of a freshwater turtle. Conectividad en un paisaje agrícola reflejada en movimientos entre charcas de una tortuga dulceacuícola. Conservation Biology, 20(3):780–791, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berggren:2002:ECH

[BBK02] Åsa Berggren, Bengt Birath, and Oskar Kindvall. Effect of corridors and habitat edges on dispersal behavior, movement rates, and movement angles in Roesel’s bush-Cricket (Metrioptera roeseli). Efecto de los corredores y los bordes del hábitat sobre la conducta de dispersión, las tasas de movimiento y los Ángulos de movimiento del grillo de arbusto de roesel (Metrioptera roeseli). Conservation Biology, 16(6):1562–1569, December 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bird:2004:ECL

[BBM04] Brittany L. Bird, Lyn C. Branch, and Deborah L. Miller. Effects of coastal lighting on foraging behavior of beach mice. Efectos del alumbrado costero sobre el comportamiento de forrajeo de ratones de playa. Conservation Biology, 18(5):1435–1439, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradshaw:2007:DSC

[BBM07] Corey J. A. Bradshaw, Barry W. Brook, and Clive R. McMahon. Dangers of sensationalizing conservation biology. Conservation Biology, 21(3):570–571, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blaustein:2001:ABC

[BBO^{+}01] Andrew R. Blaustein, Lisa K. Belden, Deanna H. Olson, David M. Green, Terry L. Root, and Joseph M. Kiesecker. Amphibian breeding and climate change. Reproducción de anfibios y cambio climático. Conservation Biology, 15(6):1804–1809, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brosi:2007:CEB

[BBW^{+}07] Berry J. Brosi, Michael J. Balick, Robert Wolkow, Roberta Lee, Mark Kostka, William Raynor, Robert Gallen, Ally Raynor, Pelihter Raynor, and Dana Lee Ling. Cultural erosion and biodiversity: Canoe-making knowledge in Pohnpei, Micronesia. Erosión cultural y biodiversidad: Conocimiento sobre la manufactura de canoas en Pohnpei, Micronesia. Conservation Biology, 21(2):875–879, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balmford:2006:FFF

[BC06] Andrew Balmford and Richard M. Cowling. Fusion or failure? The future of conservation biology. Conservation Biology, 20(3): 692–695, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buckley:2008:NSI

[BC08] Mark C. Buckley and Elizabeth E. Crone. Negative off-site impacts of ecological restoration: Understanding and addressing the conflict. Impactos negativos de la restauración ecológica: Comprendiendo y atendiendo el conflicto. Conservation Biology, 22(5):1118–1124, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buckley:2009:LER

[BC09] Mark C. Buckley and Elizabeth E. Crone. Leading by example: Response to Golet et al. Conservation Biology, 23(6):1638, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blackburn:2008:TAO

[BCD^{+}08] Tim M. Blackburn, Phillip Cassey, Richard P. Duncan, Karl L. Evans, and Kevin J. Gaston. Threats to avifauna on oceanic islands revisited. Conservation Biology, 22(2):492–494, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bertness:2008:ECC

[BCHS08] Mark D. Bertness, Caitlin Crain, Christine Holdredge, and Nicholas Sala. Eutrophication and consumer control of New England salt marsh primary productivity. Eutrofización y control de consumidores de la productividad primaria de una marisma en Nueva Inglaterra. Conservation Biology, 22(1):131– 139, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balmford:2009:CMD

[BCK^{+}09] Andrew Balmford, Pete Carey, Valerie Kapos, Andrea Manica, Ana S. L. Rodrigues, Jörn P. W. Scharlemann, and Rhys E. Green. Capturing the many dimensions of threat: Comment on Salafsky et al. Conservation Biology, 23(2):482–487, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SSS^{+}08].

Bax:2001:CBI

[BCMA^{+}01] N. Bax, J. T. Carlton, A. Mathews-Amos, R. L. Haedrich, F. G. Howarth, J. E. Purcell, A. Rieser, and A. Gray. The control of biological invasions in the world’s oceans. El control de invasiones biológicas en los océanos del mundo. Conservation Biology, 15(5):1234–1246, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brower:2002:QCF

[BCP^{+}02] Lincoln P. Brower, Guillermo Castilleja, Armando Peralta, Jose Lopez-Garcia, Luis Bojorquez-Tapia, Salomon Diaz, Daniela Melgarejo, and Monica Missrie. Quantitative changes in forest quality in a principal overwintering area of the Monarch butterfly in Mexico, 1971–1999. Cambios cuantitativos en la calidad del bosque en un Área principal de hibernación de la mariposa monarca en México, 1971 a 1999. Conservation Biology, 16(2): 346–359, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bentley:2000:EFA

[BCS00] Jocelyn M. Bentley, Carla P. Catterall, and Geoffrey C. Smith. Effects of fragmentation of araucarian vine forest on small mammal communities. Efectos de la fragmentación de un bosque de vid araucario en comunidades Pequeñas. Conservation Biology, 14(4):1075–1087, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bauer:2004:PPC

[BCS04] Dana Marie Bauer, Nicole E. Cyr, and Stephen K. Swallow. Public preferences for compensatory mitigation of salt marsh losses: a contingent choice of alternatives. Preferencias del público para mitigación compensatoria de pérdidas de marismas: una elección contingente de alternativas. Conservation Biology, 18(2):401– 411, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bulte:2005:EAW

[BD05] Erwin H. Bulte and Richard Damania. An economic assessment of wildlife farming and conservation. Una evaluación económica de la crianza y conservación de fauna silvestre. Conservation Biology, 19(4):1222–1233, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bell:2006:IFF

[BD06] Kristen E. Bell and Maureen A. Donnelly. Influence of forest fragmentation on community structure of frogs and lizards in northeastern Costa Rica. Influencia de la Fragmemtación de bosques sobre la estructura de la comunidad de ranas y lagartijas en el noreste de Costa Rica. Conservation Biology, 20(6): 1750–1760, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:2004:SDC

[BDR04a] Thomas Brooks, Gustavo A. B. Da Fonseca, and Ana S. L. Rodrigues. Species, data, and conservation planning. Conservation Biology, 18(6):1682–1688, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:2004:PAS

[BDR04b] Thomas M. Brooks, Gustavo A. B. Da Fonseca, and Ana S. L. Rodrigues. Protected areas and species. Conservation Biology, 18(3):616–618, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brandon:2005:SSB

[BDRC05] Katrina Brandon, Gustavo A. B. Da Fonseca, Anthony B. Rylands, and José Maria Cardoso Da Silva. Special section: Brazilian conservation: Challenges and opportunities. Conservation Biology, 19(3):595–600, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bar-David:2005:DMR

[BDSD^{+}05] Shirli Bar-David, David Saltz, Tamar Dayan, Amir Perelberg, and Amit Dolev. Demographic models and reality in reintroductions: Persian fallow deer in Israel. Modelos demográficos y la realidad en reintroducciones: Ciervo dama de Persia en Israel. Conservation Biology, 19(1):131–138, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balmer:2000:CSE

[BE00] Oliver Balmer and Andreas Erhardt. Consequences of succession on extensively grazed grasslands for Central European butterfly communities: Rethinking conservation practices. Consecuencias de la sucesión en los pastos de cultivo extensivo en las comunidades de mariposas en Europa Central replanteamiento de prácticas de conservación. Conservation Biology, 14(3):746– 757, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bearzi:2007:MCP

[Bea07] Giovanni Bearzi. Marine conservation on paper. Conservation Biology, 21(1):1–3, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bearzi:2009:PFC

[Bea09a] Giovanni Bearzi. Practicing fisheries conservation biology within harvesting regimes. Conservation Biology, 23(5):1073–1074, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bearzi:2009:WSC

[Bea09b] Giovanni Bearzi. When swordfish conservation biologists eat swordfish. Conservation Biology, 23(1):1–2, February 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boettner:2000:EBC

[BEB00] George H. Boettner, Joseph S. Elkinton, and Cynthia J. Boettner. Effects of a biological control introduction on three nontarget native species of saturniid moths. Efectos de la introducción de control biológico sobre tres especies de polillas saturniidae nativas que no eran el blanco del control. Conservation Biology, 14(6):1798–1806, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Becker:2007:MIP

[Bec07] William S. Becker. The most important President and us. Conservation Biology, 21(4):912–913, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Becker:2009:CAF

[Bec09] William Becker. Climate and the age of Faustian choices. Conservation Biology, 23(4):794–795, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beever:2000:ROC

[Bee00] Erik Beever. The roles of optimism in conservation biology. Conservation Biology, 14(3):907–909, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beebee:2002:APC

[Bee02] Trevor J. C. Beebee. Amphibian phenology and climate change. Conservation Biology, 16(6):1454, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beever:2009:ESG

[Bee09] Erik A. Beever. Ecological silence of the grasslands, forests, wetlands, mountains, and seas. Conservation Biology, 23(5): 1320–1322, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradshaw:2006:MCS

[BEH06] Corey J. A. Bradshaw, Karen Evans, and Mark A. Hindell. Mass cetacean strandings — a plea for empiricism. Conservation Biology, 20(2):584–586, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beissinger:2001:CCF

[Bei01] Steven R. Beissinger. The California Condor: a flagship adrift. Conservation Biology, 15(4):1197–1199, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [Ris02].

Beier:2002:DDC

[Bei02a] Paul Beier. Details of a desert carnivore. Conservation Biology, 16(5):1441–1443, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beissinger:2002:UPC

[Bei02b] Steven R. Beissinger. Unresolved problems in the Condor Recovery Program: Response to Risebrough. Conservation Biology, 16(4):1158–1159, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2003:AMS

[Bei03] Paul Beier. Adaptive management and SCB’s evaluation of species recovery plans. Conservation Biology, 17(3):653–655, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2005:BEC

[Bei05] Paul Beier. Being ethical as conservation biologists and as a society. Conservation Biology, 19(1):1–2, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2008:IPO

[Bei08] Colin M. Beier. Influence of political opposition and compromise on conservation outcomes in the Tongass National Forest, Alaska. Influencia de la Oposición y compromiso Político sobre los resultados de conservación en el Bosque Nacional Tongass, Alaska. Conservation Biology, 22(6):1485–1496, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bekoff:2004:CBH

[Bek04] Marc Bekoff. Conservation behavior is here to stay. Conservation Biology, 18(2):591–593, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Belfiore:2002:HMB

[Bel02] Natalia M. Belfiore. Half a Millennium of bad habits. Conservation Biology, 16(5):1439–1441, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burke:2004:PNP

[BEM^{+}04] Dawn M. Burke, Ken Elliott, Levi Moore, Wendy Dunford, Erica Nol, Judith Phillips, Stephen Holmes, and Kathryn Freemark. Patterns of nest predation on artificial and natural nests in forests. Patrones de depredación de nidos artificiales y naturales en bosques. Conservation Biology, 18(2):381–388, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2000:TCW

[Ben00] Elizabeth L. Bennett. Timber certification: Where is the voice of the biologist? Conservation Biology, 14(4):921–923, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2001:JET

[Ben01a] Elizabeth L. Bennett. The joint effort of timber certification. Conservation Biology, 15(2):318–319, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2001:TCW

[Ben01b] Elizabeth L. Bennett. Timber certification: Where is the voice of the biologist? Conservation Biology, 15(2):308–310, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2002:TLB

[Ben02] Elizabeth L. Bennett. Is there a link between wild meat and food security? Conservation Biology, 16(3):590–592, June 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2001:THF

[Ber01] Joel Berger. Tiger, humans, and the face of conservation. Conservation Biology, 15(1):292–293, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2003:IAL

[Ber03] Joel Berger. Is it acceptable to let a species go extinct in a national park? Conservation Biology, 17(5):1451–1454, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2004:LMH

[Ber04a] Joel Berger. The last mile: How to sustain long-distance migration in mammals. La Última milla: Como sostener la migración de larga distancia en mamíferos. Conservation Biology, 18(2): 320–331, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2004:UMM

[Ber04b] Joel Berger. Using a magnetic mammal to conserve an ecosystem. Conservation Biology, 18(3):853–854, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berkes:2004:RCB

[Ber04c] Fikret Berkes. Rethinking community-based conservation. Repensando la conservación basada en comunidades. Conservation Biology, 18(3):621–630, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berggren:2005:EPS

[Ber05] ÅSa Berggren. Effect of propagule size and landscape structure on morphological differentiation and asymmetry in experimentally introduced Roesel’s bush-crickets. Efecto del tamaño de los propágulos y de la estructura del paisaje sobre la Diferenciación morfológica y asimetría en grillos (Metrioptera roeseli) introducidos experimentalmente. Conservation Biology, 19(4): 1095–1102, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2006:CLC

[Ber06] Kim Murray Berger. Carnivore-livestock conflicts: Effects of subsidized predator control and economic correlates on the sheep industry. Conflictos carnívoros–ganado: Efectos del control subsidiado de depredadores y correlaciones económicas sobre la industria ovina. Conservation Biology, 20(3):751–761, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2007:CRH

[Ber07] Joel Berger. Carnivore repatriation and holarctic prey: Narrowing the deficit in ecological effectiveness. Repatriación de carnívoros y presas Holárticas: Reducción del Déficit de eficacia ecológica. Conservation Biology, 21(4):1105–1116, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beyers:2004:PSE

[Bey04] Jan L. Beyers. Postfire seeding for erosion control: Effectiveness and impacts on native plant communities. Siembra postincendio para control de erosión: Efectividad e impactos sobre comunidades de plantas nativas. Conservation Biology, 18 (4):947–956, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bouton:2003:SPW

[BF03] Shannon N. Bouton and Peter C. Frederick. Stakeholders’ perceptions of a wading bird colony as a community resource in the Brazilian Pantanal. Percepciones de interesados de una colonia de aves vadeadoras como recurso comunitario en El Pantanal brasileño. Conservation Biology, 17(1):297–306, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barbour:2009:BCG

[BFB^{+}09] Robert C. Barbour, Lynne G. Forster, Susan C. Baker, Dorothy A. Steane, and Bradley M. Potts. Biodiversity consequences of genetic variation in bark characteristics within a foundation tree species. Consecuencias de la variación genética en las características de la corteza en una especie de Árbol fundadora. Conservation Biology, 23(5):1146–1155, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Betts:2007:TSO

[BFD07] Matthew G. Betts, Graham J. Forbes, and Antony W. Diamond. Thresholds in songbird occurrence in relation to landscape structure. Umbrales en la ocurrencia de aves canoras en relación con la estructura del paisaje. Conservation Biology, 21(4):1046–1058, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:2006:THS

[BFH^{+}06] Jeremy S. Brooks, Margaret A. Franzen, Christopher M. Holmes, Mark N. Grote, and Monique Borgerhoff Mulder. Testing hypotheses for the success of different conservation strategies. Probando hipótesis para el Éxito de diferentes estrategias de conservación. Conservation Biology, 20(5):1528–1538, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blumstein:2004:ECB

[BFJ04] Daniel T. Blumstein and Esteban Fernández-Juricic. The emergence of conservation behavior. Conservation Biology, 18(5):

1175–1177, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Biek:2002:WMA

[BFMM02] Roman Biek, W. Chris Funk, Bryce A. Maxell, and L. Scott Mills. What is missing in amphibian decline research: Insights from ecological sensitivity analysis. Investigación sobre la disminución de anfibios: Comprensión del análisis de sensibilidad ecológica. Conservation Biology, 16(3):728–734, June 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boitani:2007:ENC

[BFMR07] Luigi Boitani, Alessandra Falcucci, Luigi Maiorano, and Carlo Rondinini. Ecological networks as conceptual frameworks or operational tools in conservation. Redes ecológicas como marcos conceptuales o herramientas operacionales de la conservación. Conservation Biology, 21(6):1414–1422, December 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Battisti:2004:CUC

[BG04] Corrado Battisti and Spartaco Gippoliti. Conservation in the urban-countryside interface: a cautionary note from Italy. Conservation Biology, 18(2):581–583, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bautista:2004:EWR

[BGC^{+}04] Luis M. Bautista, Jesús T. García, Ricardo G. Calmaestra, Carlos Palacín, Carlos A. Martín, Manuel B. Morales, Raúl Bonal, and Javier Viñuela. Effect of weekend road traffic on the use of space by raptors. Efecto del tráfico de fin de semana sobre el uso de espacio por rapaces. Conservation Biology, 18(3):726– 732, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bonter:2009:CIS

[BGD09] David N. Bonter, Sidney A. Gauthreaux, Jr., and Therese M. Donovan. Characteristics of important stopover locations for migrating birds: Remote sensing with radar in the Great Lakes Basin. Características de localidades de escala temporal para aves migrantes: Percepción remota con radar en la Cuenca de los Grandes Lagos. Conservation Biology, 23(2):440–448, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barrett:2006:CLB

[BGHM06] Christopher B. Barrett, Clark C. Gibson, Barak Hoffman, and Mathew D. Mccubbins. The complex links between governance and biodiversity. Los complejos vínculos entre la autoridad y la biodiversidad. Conservation Biology, 20(5):1358–1366, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barrett:2007:RLB

[BGHMb07] Christopher B. Barrett, Clark C. Gibson, Barak Hoffman, and Mathew D. McCub-bins. Revisiting the links between governance and biodiversity loss. Conservation Biology, 21(4):900– 901, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balmford:2000:ICC

[BGRJ00] Andrew Balmford, Kevin J. Gaston, Ana S. L. Rodrigues, and Alexander James. Integrating costs of conservation into international priority setting. Conservation Biology, 14(3):597–605, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2000:EGD

[BGWE00] B. L. Brown, T. P. Gunter, J. M. Waters, and J. M. Epifanio. Evaluating genetic diversity associated with propagationassisted restoration of American shad. Evaluación de la diversidad genética asociada con la restauración por propagación de la alosa americana. Conservation Biology, 14(1):294–303, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berkson:2001:ICC

[BH01] Jim Berkson and Autumn-Lynn Harrison. An integrative capstone course for the conservation biology curriculum. Conservation Biology, 15(5):1461–1463, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blake:2004:SFC

[BH04] Stephen Blake and Simon Hedges. Sinking the flagship: the case of forest elephants in Asia and Africa. Conservation Biology, 18 (5):1191–1202, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bayne:2008:ICA

[BHB08] Erin M. Bayne, Lucas Habib, and Stan Boutin. Impacts of chronic anthropogenic noise from energy-sector activity on abundance of songbirds in the boreal forest. Impactos de ruido antropogénico Crónico de la actividad del sector Energético sobre la abundancia de aves canoras en el bosque boreal. Conservation Biology, 22(5):1186–1193, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barnosky:2001:TTV

[BHMC01] Anthony D. Barnosky, Elizabeth A. Hadly, Brian A. Maurer, and Michael I. Christie. Temperate terrestrial vertebrate faunas in North and South America: Interplay of ecology, evolution, and geography with biodiversity. Fauna de vertebrados terrestres templados en Norte y Suramérica: Interacción entre la ecología, la evolución y la Geografía con la biodiversidad. Conservation Biology, 15(3):658–674, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:2001:PVE

[BHW01] Joel Berger, Anne Hoylman, and William Weber. Perturbation of vast ecosystems in the absence of adequate science: Alaska’s Arctic refuge. Conservation Biology, 15(2):539–541, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradshaw:2006:CVN

[BIK^{+}06] Corey J. A. Bradshaw, Yuji Isagi, Shingo Kaneko, David M. J. S. Bowman, and Barry W. Brook. Conservation value of nonnative banteng in Northern Australia. Valor de conservación de Bos javanicus introducido en el Norte de Australia. Conservation Biology, 33(5):1306–1311, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Binkley:2002:HPY

[Bin02] Dan Binkley. How productive is your planet? Conservation Biology, 16(6):1664–1665, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brockington:2006:CHR

[BISS06] Dan Brockington, Jim Igoe, and Kai Schmidt-Soltau. Conservation, human rights, and poverty reduction. Conservation Biology, 20(1):250–252, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berglund:2005:VED

[BJ05] HÅKan Berglund and Bengt Gunnar Jonsson. Verifying an extinction debt among lichens and fungi in Northern Swedish boreal forests. Verificación de una deuda de extinción entre líquenes y hongos en bosques boreales del norte de Suecia. Conservation Biology, 19(2):338–348, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beecher:2002:ABO

[BJB^{+}02] Nancy A. Beecher, Ron J. Johnson, James R. Brandle, Ronald M. Case, and Linda J. Young. Agroecology of birds in organic and nonorganic farmland. Agroecología de aves en granjas Orgánicas y no Orgánicas. Conservation Biology, 16(6): 1620–1631, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bock:2006:RCE

[BJB06] Carl E. Bock, Zach F. Jones, and Jane H. Bock. Rodent communities in an exurbanizing southwestern landscape (U.S.A.) comunidades de roedores en un paisaje exurbano sur occidental (E.U.A.). Conservation Biology, 33(5):1242–1250, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brys:2005:EHD

[BJE^{+}05] Rein Brys, Hans Jacquemyn, Patrick Endels, Geert De Blust, and Martin Hermy. Effect of habitat deterioration on population dynamics and extinction risks in a previously common perennial. Efecto del deterioro del hábitat sobre la dinámica poblacional y el riesgo de extinción de una perenne previamente común. Conservation Biology, 19(5):1633–1643, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Backer:2004:IFS

[BJM04] Dana M. Backer, Sara E. Jensen, and Guy R. Mcpherson. Impacts of fire-suppression activities on natural communities. Impactos de actividades de supresión de fuego sobre comunidades naturales. Conservation Biology, 18(4):937–946, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bossenbroek:2007:CC

[BJPL07a] Jonathan M. Bossenbroek, Ladd E. Johnson, Brett Peters, and David M. Lodge. Cover caption. Conservation Biology, 21(3): i, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bossenbroek:2007:FEZ

[BJPL07b] Jonathan M. Bossenbroek, Ladd E. Johnson, Brett Peters, and David M. Lodge. Forecasting the expansion of zebra mussels in the United States. Pronóstico de la expansión de mejillones cebra en los Estados Unidos. Conservation Biology, 21(3):800– 810, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bellinger:2003:LGV

[BJTD03] M. Renee Bellinger, Jeff A. Johnson, John Toepfer, and Peter Dunn. Loss of genetic variation in greater prairie chickens following a population bottleneck in Wisconsin, U.S.A. Pérdida de variación genética en Tympanuchus cupido después de un cuello de botella poblacional en Wisconsin, EE.UU. Conservation Biology, 17(3):717–724, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bruna:2002:HFD

[BK02] Emilio M. Bruna and W. John Kress. Habitat fragmentation and the demographic structure of an Amazonian understory herb (Heliconia acuminata). Fragmentación del hábitat y la estructura demográfica de una hierba (Heliconia acuminata) del sotobosque amazónico. Conservation Biology, 16(5):1256–1266, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bendik-Keymer:2009:SPP

[BK09] Jeremy Bendik-Keymer. The story of a philosophical problem. Conservation Biology, 23(3):778–781, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Benes:2003:LQR

[BKK03] Jiří Beneš, Pavel Kepka, and Martin Konvička. Limestone quarries as refuges for European xerophilous butterflies. Canteras de caliza como refugios para mariposas europeas xerófilas.

Conservation Biology, 17(4):1058–1069, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barve:2005:MMT

[BKV^{+}05] N. Barve, M. C. Kiran, G. Vanaraj, N. A. Aravind, D. Rao, R. Uma Shaanker, K. N. Ganeshaiah, and J. G. Poulsen. Measuring and mapping threats to a wildlife sanctuary in Southern India. Medición y mapeo de amenazas a un santuario de vida silvestre en el Sur de India. Conservation Biology, 19(1):122–130, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bhagwat:2005:LAB

[BKWB05] Shonil A. Bhagwat, Cheppudira G. Kushalappa, Paul H. Williams, and Nick D. Brown. A landscape approach to biodiversity conservation of sacred groves in the Western Ghats of India. Un enfoque de paisaje la conservación de bosques sagrados en los Ghats Occidentales de India. Conservation Biology, 19(6): 1853–1862, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bodmer:2001:RDS

[BL01] Richard E. Bodmer and Etersit Pezo Lozano. Rural development and sustainable wildlife use in Peru. Desarrollo rural y utilización de vida silvestre en Perú. Conservation Biology, 15 (4):1163–1170, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brummitt:2003:BWH

[BL03] Neil Brummitt and Eimear Nic Lughadha. Biodiversity: Where’s hot and Where’s not. Conservation Biology, 17(5): 1442–1448, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2006:SIS

[BL06] Roger M. Brown and David N. Laband. Species imperilment and spatial patterns of development in the United States especies en peligro y patrones de desarrollo en los Estados Unidos. Conservation Biology, 20(1):239–244, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2007:ARF

[BL07] Roger M. Brown and David N. Laband. Applicability of research findings to conservation policy. Conservation Biology, 21(2): 549–551, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blumstein:2007:DDM

[Blu07] Daniel T. Blumstein. Darwinian decision making: Putting the adaptive into adaptive management. Conservation Biology, 21 (2):552–553, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Byers:2003:USS

[BM03] Onnie Byers and Michael Maunder. Ulysses Samuel Seal, 1929– 2003. Conservation Biology, 17(5):1461–1462, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beverly:2004:MPS

[BM04] Jennifer L. Beverly and David L. Martell. Modeling prescribed surface-fire regimes for Pinus strobus conservation. Modelado de regimenes de fuego superficial prescrito para la conservación de Pinus strobus. Conservation Biology, 18(6):1541–1552, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beale:2005:MEL

[BM05a] Colin M. Beale and P. Monaghan. Modeling the effects of limiting the number of visitors on failure rates of seabird nests. Modelado de los efectos de limitar el número de visitantes sobre las tasas de fracaso de nidos de aves marinas. Conservation Biology, 19(6):2015–2019, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blundell:2005:DRL

[BM05b] Arthur G. Blundell and Michael B. Mascia. Discrepancies in reported levels of international wildlife trade discrepancias en los niveles reportados de comercio internacional de vida silvestre. Conservation Biology, 19(6):2020–2025, December 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blundell:2006:RDW

[BM06] Arthur G. Blundell and Michael B. Mascia. Response to data on wildlife trade. Conservation Biology, 20(3):598–599, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Basset:2008:CAA

[BMA^{+}08] Yves Basset, Olivier Missa, Alfonso Alonso, Scott E. Miller, Gianfranco Curletti, Marc De Meyer, Connal Eardley, Owen T. Lewis, Mervyn W. Mansell, Vojtech Novotny, and Thomas Wagner. Changes in arthropod assemblages along a wide gradient of disturbance in Gabon. Cambios en ensambles de artrópodos a lo largo de un gradiente de perturbación amplio en Gabón. Conservation Biology, 22(6):1552–1563, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowker:2008:PCE

[BMB^{+}08] Matthew A. Bowker, Mark E. Miller, Jayne Belnap, Thomas D. Sisk, and Nancy C. Johnson. Prioritizing conservation effort through the use of biological soil crusts as ecosystem function indicators in an arid region. Priorización del esfuerzo de conservación mediante el uso de capas de suelo biológico como indicadores del funcionamiento del ecosistema en una región Árida. Conservation Biology, 22(6):1533–1543, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bani:2006:MMS

[BMBM06] Luciano Bani, Dario Massimino, Luciana Bottoni, and Renato Massa. A multiscale method for selecting indicator species and priority conservation areas: a case study for broadleaved forests in Lombardy, Italy. Un método multiescala para seleccionar especies indicadoras y Áreas prioritarias para la conservación: un estudio de caso para bosques de hojas anchas en Lombardía, Italia. Conservation Biology, 20(2):512–526, April 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bartlett:2009:COP

[BMC^{+}09] C. Y. Bartlett, C. Manua, J. Cinner, S. Sutton, R. Jimmy, R. South, J. Nilsson, and J. Raina. Comparison of outcomes of permanently closed and periodically harvested coral reef reserves. Comparación de resultados de reservas de arrecifes de coral cerradas permanentemente y cosechadas periódicamente.

Conservation Biology, 23(6):1475–1484, December 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barber:2001:NSB

[BMM^{+}01] David R. Barber, Thomas E. Martin, M. Anthony Melchiors, Ronald E. Thill, and T. Bently Wigley. Nesting success of birds in different silvicultural treatments in Southeastern U.S. pine forests. Éxito de nidación de aves en diferentes tratamientos silviculturales en bosques de pinos del sureste de los Estados Unidos. Conservation Biology, 15(1):196–207, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:2002:HLE

[BMM^{+}02] Thomas M. Brooks, Russell A. Mittermeier, Cristina G. Mittermeier, Gustavo A. B. Da Fonseca, Anthony B. Rylands, William R. Konstant, Penny Flick, John Pilgrim, Sara Oldfield, Georgina Magin, and Craig Hilton-Taylor. Habitat loss and extinction in the hotspots of biodiversity. Pérdida de hábitat y extinciones en Áreas críticas para la biodiversidad. Conservation Biology, 16(4):909–923, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Benitez-Malvido:2003:IFF

[BMMR03] Julieta Benítez-Malvido and Miguel Martínez-Ramos. Impact of forest fragmentation on understory plant species richness in Amazonia. Impacto de la fragmentación de la selva sobre la riqueza de especies del sotobosque en el Amazonas. Conservation Biology, 17(2):389–400, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brudvig:2007:ECN

[BMMW07] Lars A. Brudvig, Catherine M. Mabry, James R. Miller, and Tracy A. Walker. Evaluation of Central North American prairie management based on species diversity, life form, and individual species metrics. Evaluación del manejo de praderas en Norteamérica Central con base en la diversidad de especies, formas de vida y medidas de especies individuales. Conservation Biology, 21(2):864–874, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baxter:2006:AMC

[BMP^{+}06] Peter W. J. Baxter, Michael A. Mccarthy, Hugh P. Possingham, Peter W. Menkhorst, and Natasha Mclean. Accounting for management costs in sensitivity analyses of matrix population models. Representación de gastos de administración en el análisis de sensibilidad de modelos poblacionales matriciales. Conservation Biology, 21(3):893–905, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beger:2007:EST

[BMP07] Maria Beger, Sheila A. McKenna, and Hugh P. Possingham. Effectiveness of surrogate taxa in the design of coral reef reserve systems in the Indo-0Pacific. Efectividad de taxa sustitutos en el diseño de sistemas de reservas de arrecifes de coral en el Indo–Pacífico. Conservation Biology, 22(1):1584–1593, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bray:2003:MCM

[BMPNC^{+}03] David Barton Bray, Leticia Merino-Pérez, Patricia Negreros-Castillo, Gerardo Segura-Warnholtz, Juan Manuel Torres-Rojo, and Henricus F. M. Vester. Mexico’s community-managed forests as a global model for sustainable landscapes. Bosques manejados comunalmente en México como un modelo global de paisajes sostenibles. Conservation Biology, 17(3):672–677, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beck:2004:NTM

[BMRB04] Michael W. Beck, Theresa D. Marsh, Shauna E. Reisewitz, and Marci L. Bortman. New tools for marine conservation: the leasing and ownership of submerged lands. Nuevas herramientas para la conservación marina: El arrendamiento y propiedad de tierras sumergidas. Conservation Biology, 18(5):1214–1223, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2008:FRC

[BMS08] Paul Beier, Daniel R. Majka, and Wayne D. Spencer. Forks in the road: Choices in procedures for designing wildland linkages bifurcaciones en el camino: Opciones de procedimientos para el diseño de enlaces de tierras silvestres. Conservation Biology, 22 (4):836–851, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bisseleua:2009:BCE

[BMV09] D. H. B. Bisseleua, A. D. Missoup, and S. Vidal. Biodiversity conservation, ecosystem functioning, and economic incentives under cocoa agroforestry. Intensification conservación de biodiversidad, funcionamiento del ecosistema e incentivos económicos en la intensificación de la agroforestería de cacao. Conservation Biology, 23(5):1176–1184, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bakker:2002:ILA

[BNH02] Kristel K. Bakker, David E. Naugle, and Kenneth F. Higgins. Incorporating landscape attributes into models for migratory grassland bird conservation. Incorporación de atributos del paisaje en modelos de conservación de aves migratorias de pastizal. Conservation Biology, 16(6):1638–1646, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buchanan:2009:DGT

[BNM^{+}09] Graeme M. Buchanan, Andrew Nelson, Philippe Mayaux, Andrew Hartley, and Paul F. Donald. Delivering a global, terrestrial, biodiversity observation system through remote sensing. Conservation Biology, 23(2):499–502, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bonin:2007:PAI

[BNP^{+}07] Aurelie Bonin, Florence Nicole, François Pompanon, Claude Miaud, and Pierre Taberlet. Population adaptive index: a new method to help measure intraspecific genetic diversity and prioritize populations for conservation. Índice de la adaptación poblacional: un nuevo método para ayudar a medir la diversidad genética Intraespecífica y priorizar poblaciones para conservación. Conservation Biology, 21(3):697–708, June 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Broomhall:2000:CEA

[BOC00] Sara D. Broomhall, William S. Osborne, and Ross B. Cunningham. Comparative effects of ambient ultraviolet-B radiation on two sympatric species of Australian frogs. Efectos comparativos de la radiación ultravioleta-B ambiental en dos especies simpátricas de ranas Australianas. Conservation Biology, 14(2): 420–427, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berglund:2009:QHR

[BOJ09] HÅKan Berglund, Robert B. O’Hara, and Bengt Gunnar Jonsson. Quantifying habitat requirements of tree-living species in fragmented boreal forests with Bayesian methods. Cuantificación de los requerimientos de hábitat de especies arborícolas en bosques boreales fragmentados mediante métodos Bayesianos. Conservation Biology, 23(5):1127–1137, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Borman:2005:FSD

[Bor05] Michael M. Borman. Forest stand dynamics and livestock grazing in historical context. Dinámica de bosques y pastoreo de ganado en un contexto histórico. Conservation Biology, 19(5): 1658–1662, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Borghesio:2008:EHS

[Bor08] Luca Borghesio. Effects of human subsistence activities on forest birds in Northern Kenya. Efecto de las actividades humanas de subsistencia sobre aves de bosque en el Norte de Kenia. Conservation Biology, 22(2):384–394, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bossuyt:2007:GRI

[Bos07] Beatrijs Bossuyt. Genetic rescue in an isolated metapopulation of a naturally fragmented plant species, Parnassia palustris. Rescate genético en una metapoblación aislada de una especie de planta fragmentada naturalmente, Parnassia palustris. Conservation Biology, 21(2):832–841, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowkett:2009:RCB

[Bow09] Andrew E. Bowkett. Recent captive-breeding proposals and the return of the ark concept to global species conservation. Conservation Biology, 23(3):773–776, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradley:2005:EWT

[BPB^{+}05] Elizabeth H. Bradley, Daniel H. Pletscher, Edward E. Bangs, Kyran E. Kunkel, Douglas W. Smith, Curt M. Mack, Thomas J. Meier, Joseph A. Fontaine, Carter C. Niemeyer, and Michael D. Jimenez. Evaluating wolf translocation as a nonlethal method to reduce livestock conflicts in the Northwestern United States. Evaluación de la translocación de lobos como un método no letal para reducir conflictos con el ganado en el noroeste de Estados Unidos. Conservation Biology, 19(5):1498–1508, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Becker:2007:CSP

[BPB07] Leandro A. Becker, Miguel A. Pascual, and Néstor G. Basso. Colonization of the Southern Patagonia Ocean by exotic Chinook salmon. Colonización del sur de la Patagonia por el salmón exótico Oncorhynchus tshawytscha. Conservation Biology, 21 (5):1347–1352, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burgman:2001:MSS

[BPL^{+}01] Mark A. Burgman, Hugh P. Possingham, A. Jasmyn J. Lynch, David A. Keith, Michael A. McCarthy, Stephen D. Hopper, Wendy L. Drury, Jason A. Passioura, and Robert J. Devries. A method for setting the size of plant conservation target areas. Método para determinar el tamaño de las Áreas de interés para la conservación de plantas. Conservation Biology, 15(3):603– 616, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boulton:2005:EKS

[BPP05] Andrew J. Boulton, Debra Panizzon, and Julian Prior. Explicit knowledge structures as a tool for overcoming obstacles to interdisciplinary research. Conservation Biology, 19(6):2026–2029, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ban:2009:CIC

[BPV09] Natalie C. Ban, Chris R. Picard, and Amanda C. J. Vincent. Comparing and integrating community-based and science-based approaches to prioritizing marine areas for protection. Comparación e integración de métodos basados en evaluaciones comunitarias y cientificas para la priorización de Áreas marinas protegidas. Conservation Biology, 23(4):899–910, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beebee:2001:AGB

[BR01] Trevor Beebee and Graham Rowe. Application of genetic bottleneck testing to the investigation of amphibian declines: a case study with natterjack toads. Aplicaciones de pruebas de cuello de botella genético para la investigación de disminuciones de anfibios: un caso de estudio con el sapo (Bufo calamita). Conservation Biology, 15(1):266–270, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowen:2005:GHO

[BR05] Brian W. Bowen and Joe Roman. Gaia’s handmaidens: the orlog model for conservation biology. La servidumbre de Gaia: El modelo de orlog para la biología de la conservación. Conservation Biology, 19(4):1037–1043, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brashares:2003:EBL

[Bra03a] Justin S. Brashares. Ecological, behavioral, and life-history correlates of mammal extinctions in West Africa. Correlaciones ecológicas, conductuales y de historia de vida de extinciones de mamíferos en África Occidental. Conservation Biology, 17 (3):733–743, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burger:2003:RAC

[BRA^{+}03b] J. C. Burger, R. A. Redak, E. B. Allen, J. T. Rotenberry, and M. F. Allen. Restoring arthropod communities in coastal sage scrub. Restablecimiento de comunidades de artrópodos en el matorral costero de salvia. Conservation Biology, 17(2):460–467, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Braden:2005:WCR

[Bra05a] Kathleen Braden. A warning of challenges that remain. Conservation Biology, 19(2):581–582, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bray:2005:CWK

[Bra05b] David Barton Bray. Challenging the we-know-best approach. Conservation Biology, 19(1):278–279, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brawn:2006:ERO

[Bra06] Jeffrey D. Brawn. Effects of restoring oak savannas on bird communities and populations. Efectos de la restauración de sabanas de roble sobre comunidades y poblaciones de aves. Conservation Biology, 20(2):460–469, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buse:2008:ELB

[BRA08] J. Buse, T. Ranius, and T. Assmann. An endangered longhorn beetle associated with old oaks and its possible role as an ecosystem engineer un escarabajo en peligro asociados con robles viejos y su posible papel como ingeniero del ecosistema. Conservation Biology, 22(2):329–337, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bosch:2006:EIS

[BRBMS06] Jaime Bosch, Pedro A. Rincón, Luz Boyero, and Iñigo Martínez-Solano. Effects of introduced salmonids on a montane population of Iberian frogs. Efectos de Salmónidos introducidos sobre una población Montana de rana Ibérica. Conservation Biology, 20 (1):180–189, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bonine:2003:TET

[BRD03] Kimberly Bonine, John Reid, and Robyn Dalzen. Training and education for tropical conservation. Conservation Biology, 17(5): 1209–1218, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brewer:2001:CCL

[Bre01] Carol Brewer. Cultivating conservation literacy: “Trickle-Down” education is not enough. Conservation Biology, 15(5): 1203–1205, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brewer:2002:CEP

[Bre02a] Carol Brewer. Conservation education partnerships in schoolyard laboratories: a call back to action. Conservation Biology, 16(3):577–579, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brewer:2002:OPP

[Bre02b] Carol Brewer. Outreach and partnership programs for conservation education where endangered species conservation and research occur. Conservation Biology, 16(1):4–6, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brewer:2003:CCH

[Bre03] Carol Brewer. Computers in the classroom: How information technology can improve conservation education. Conservation Biology, 17(3):657–660, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brewer:2006:TDM

[Bre06] Carol Brewer. Translating data into meaning: Education in conservation biology. Conservation Biology, 20(3):689–691, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bride:2006:CCE

[Bri06] Ian Bride. The conundrum of conservation education and the conservation mission. Conservation Biology, 20(5):1337–1339, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blaustein:2002:APC

[BRK^{+}02] Andrew R. Blaustein, Terry L. Root, Joseph M. Kiesecker, Lisa K. Belden, Deanna H. Olson, and David M. Green. Amphibian phenology and climate change. Conservation Biology, 16 (6):1454–1455, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blaustein:2003:ABC

[BRK^{+}03] Andrew R. Blaustein, Terry L. Root, Joseph M. Kiesecker, Lisa K. Belden, Deanna H. Olson, and David M. Green. Amphibian breeding and climate change: Reply to Corn. Conservation Biology, 17(2):626–627, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beschta:2004:PMF

[BRK^{+}04] Robert L. Beschta, Jonathan J. Rhodes, J. Boone Kauffman, Robert E. Gresswell, G. Wayne Minshall, James R. Karr, David A. Perry, F. Richard Hauer, and Christopher A. Frissell. Postfire management on forested public lands of the Western United States. Gestión post-incendio en terrenos boscosos Públicos en el Oeste de E.U.A. Conservation Biology, 18(4):957– 967, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brook:2000:POB

[Bro00] Barry W. Brook. Pessimistic and optimistic bias in population viability analysis. Conservation Biology, 14(2):564–566, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brosius:2004:IPP

[Bro04] J. Peter Brosius. Indigenous peoples and protected areas at the world parks Congress. Conservation Biology, 18(3):609–612, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:2005:EYE

[Bro05] Daniel M. Brooks. Everything you ever wanted to know about cracids but already knew. Conservation Biology, 19(3):982–983, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brosius:2006:CGB

[Bro06] J. Peter Brosius. Common ground between anthropology and conservation biology. Conservation Biology, 20(3):683–685, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bromley:2008:OE

[Bro08a] Daniel W. Bromley. Ownership explored. Conservation Biology, 22(4):1079–1081, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooke:2008:CAC

[Bro08b] Cassandra Brooke. Conservation and adaptation to climate change. Conservación y adaptación al cambio climático. Conservation Biology, 22(6):1471–1476, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Byers:2002:DRR

[BRR^{+}02] James E. Byers, Sarah Reichard, John M. Randall, Ingrid M. Parker, Carey S. Smith, W. M. Lonsdale, I. A. E. Atkinson, T. R. Seastedt, Mark Williamson, E. Chornesky, and D. Hayes. Directing research to reduce the impacts of nonindigenous species. Dirección de la investigación para reducir los efectos de especies exóticas. Conservation Biology, 16(3):630–640, June 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blaustein:2005:IVS

[BRS^{+}05] Andrew R. Blaustein, John M. Romansic, Erin A. Scheessele, Barbara A. Han, Allan P. Pessier, and Joyce E. Longcore. Interspecific variation in susceptibility of frog tadpoles to the pathogenic fungus Batrachochytrium dendrobatidis. Variación interespecífica en la susceptibilidad de renacuajos al hongo patógeno Batrachochytrium dendrobatidis. Conservation Biology, 19(5):1460–1468, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baldwin:2007:RBS

[BRTW07] Robert F. Baldwin, Justina C. Ray, Stephen C. Trombulak, and Gillian Woolmer. Relationship between spatial distribution of urban sprawl and species imperilment: Response to Brown and Laband. Conservation Biology, 21(2):546–548, April 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bruna:2002:BGF

[Bru02] Emilio M. Bruna. A Biologist’s guide to fragmenting the Amazon (including why That’s a bad idea). Conservation Biology, 16(5):1446–1448, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brush:2005:LLF

[Bru05] Grace S. Brush. The long lens of forest history. Conservation Biology, 19(3):979–981, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bohensky:2006:FES

[BRV06] Erin L. Bohensky, Belinda Reyers, and Albert S. Van Jaarsveld. Future ecosystem services in a Southern African river basin: a scenario planning approach to uncertainty. Servicios ecosistémicos futuro en una cuenca hidrológica sudafricana: un acercamiento a la incertidumbre mediante la planificación de escenarios. Conservation Biology, 33(5):1051–1061, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brower:2001:CVC

[BRYK01] Ann Brower, Chanel Reedy, and Jennifer Yelin-Kefer. Consensus versus conservation in the Upper Colorado River Basin Recovery Implementation Program. Consenso versus conservación en el programa de implementación de la Recuperación de la Cuenca Alta del Río Colorado. Conservation Biology, 15(4):1001–1007, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bridges:2000:VPT

[BS00] Christine M. Bridges and Raymond D. Semlitsch. Variation in pesticide tolerance of tadpoles among and within species of Ranidae and patterns of amphibian decline. Variación de la tolerancia a pesticidas en renacuajos entre y dentro de especies de la familia Ranidae y patrones de disminución de anfibios. Conservation Biology, 14(5):1490–1499, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beres:2001:DCC

[BS01a] Diane L. Beres and Anthony M. Starfield. “Demography of the California condor” and management decisions based on modeling: a reply to Meretsky et al. Conservation Biology, 15(5): 1445–1448, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boone:2001:IIL

[BS01b] Michelle D. Boone and Raymond D. Semlitsch. Interactions of an insecticide with larval density and predation in experimental amphibian communities. Interacciones de un insecticida con la densidad larval y la depredación en comunidades experimentales de anfibios. Conservation Biology, 15(1):228–238, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bonnie:2003:TRD

[BS03] Robert Bonnie and Stephan Schwartzman. Tropical reforestation and deforestation and the Kyoto protocol. Conservation Biology, 17(1):4–5, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bertness:2008:CCS

[BS08] Mark D. Bertness and Brian R. Silliman. Consumer control of salt marshes driven by human disturbance. Control por consumidores en marismas conducido por perturbación humana. Conservation Biology, 22(3):618–623, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beckmann:2009:IIC

[BS09] Christa Beckmann and Richard Shine. Impact of invasive cane toads on Australian birds. Impacto de Bufo marinus invasores sobre aves Australianas. Conservation Biology, 23(6):1544–1549, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blyth-Skyrme:2006:CBT

[BSKH^{+}06] Robert E. Blyth-Skyrme, Michel J. Kaiser, Jan G. Hiddink, Gareth Edwards-Jones, and Paul J. B. Hart. Conservation benefits of temperate marine protected areas: Variation among fish species. Beneficios de conservación de Áreas marinas templadas protegidas: Variación entre especies de peces. Conservation Biology, 20(3):811–820, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blyth-Skyrme:2007:CC

[BSKH^{+}07] Robert E. Blyth-Skyrme, Michel J. Kaiser, Jan G. Hiddink, Gareth Edwards-Jones, and Paul J. B. Hart. Cover caption. Conservation Biology, 20(3):i, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boone:2008:SGC

[BSM08] Michelle D. Boone, Raymond D. Semlitsch, and Cory Mosby. Suitability of golf course ponds for amphibian metamorphosis when bullfrogs are removed beneficio de la charcas en campos de golf para la metamorfosis de anfibios cuando son removidas las ranas. Conservation Biology, 22(1):172–179, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:2004:RCP

[BSR04] Kamaljit S. Bawa, Reinmar Seidler, and Peter H. Raven. Reconciling conservation paradigms. Conservation Biology, 18(4):859– 860, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bergamini:2005:PML

[BSS^{+}05] Ariel Bergamini, Christoph Scheidegger, Silvia Stofer, Palmira Carvalho, Simon Davey, Michael Dietrich, Florence Dubs, Edit Farkas, Urs Groner, Kati Kärkkäinen, Christine Keller, László Lökös, Sampsa Lommi, Cristina Máguas, Ruth Mitchell, Pedro Pinho, Víctor J. Rico, Gregorio Aragón, Anne-Marie Truscott, Pat Wolseley, and Alan Watt. Performance of macrolichens and lichen genera as indicators of lichen species richness and composition. El funcionamiento de Géneros de macrolíquenes y líquenes como indicadores de la riqueza y composición de especies de líquenes. Conservation Biology, 19(4):1051–1062, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bellemain:2005:EPS

[BST^{+}05] Eva Bellemain, Jon E. Swenson, David Tallmon, Sven Brunberg, and Pierre Taberlet. Estimating population size of elusive animals with DNA from Hunter-collected feces: Four methods for brown bears. Estimación del tamaño poblacional de animales elusivos con ADN de heces colectadas por cazadores: Cuatro métodos para osos pardos. Conservation Biology, 19(1):150– 161, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bejder:2006:DRA

[BSW^{+}06] Lars Bejder, Amy Samuels, Hal Whitehead, Nick Gales, Janet Mann, Richard Connor, Mike Heithaus, Jana Watson-Capps, Cindy Flaherty, and Michael Krützen. Decline in relative abundance of bottlenose dolphins exposed to long-term disturbance. Declinación de la abundancia relativa de delfines expuestos a perturbaciones de largo plazo. Conservation Biology, 21(2): 1791–1798, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baxter:2008:CES

[BSW^{+}08] Peter W. J. Baxter, John L. Sabo, Chris Wilcox, Michael A. Mccarthy, and Hugh P. Possingham. Cost-effective suppression and eradication of invasive predators. Supresión y erradicación rentable de depredadores invasores. Conservation Biology, 22 (1):89–98, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baldwin:2007:LDC

[BT07a] Robert F. Baldwin and Stephen C. Trombulak. Losing the dark: a case for a national policy on land conservation. Conservation Biology, 21(5):1133–1134, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brussard:2007:CBF

[BT07b] Peter F. Brussard and John C. Tull. Conservation biology and four types of advocacy. Conservation Biology, 21(1):21– 24, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bojorquez-Tapia:2003:MEK

[BTBC^{+}03] Luis A. Bojórquez-Tapia, Lincoln P. Brower, Guillermo Castilleja, Salvador Sánchez-Colón, Mariano Hernández, William Calvert, Salomón Díaz, Paola Gómez-Priego, Georgina Alcantar, Erika Daniela Melgarejo, María José Solares, Liliana Gutiérrez, and María Del Lourdes Juárez. Mapping expert knowledge: Redesigning the Monarch butterfly Biosphere reserve mapeo del conocimiento experto: Rediseño de la reserva de la biosfera mariposa monarca. Conservation Biology, 17(2): 367–379, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berry:2005:EVM

[BTGS05] Oliver Berry, Mandy D. Tocher, Dianne M. Gleeson, and Stephen D. Sarre. Effect of vegetation matrix on animal dispersal: Genetic evidence from a study of endangered skinks. Efecto de la matriz de vegetación sobre la dispersión animal: Evidencia genética de un estudio de esquincos en peligro. Conservation Biology, 19(3):855–864, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bierwagen:2008:CMP

[BTK08] Britta G. Bierwagen, Roxanne Thomas, and Austin Kane. Capacity of management plans for aquatic invasive species to integrate climate change. Capacidad de los planes de manejo de especies acuáticas invasoras para integrar el cambio climático. Conservation Biology, 22(3):568–574, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bangert:2005:BCP

[BTM^{+}05] Randy K. Bangert, Richard J. Turek, Gregory D. Martinsen, Gina M. Wimp, Joseph K. Bailey, and Thomas G. Whitham. Benefits of conservation of plant genetic diversity to arthropod diversity. Beneficios de la conservación de diversidad genética de plantas para la diversidad de artrópodos. Conservation Biology, 19(2):379–390, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burghardt:2009:INP

[BTS09] Karin T. Burghardt, Douglas W. Tallamy, and W. Gregory Shriver. Impact of native plants on bird and butterfly biodiversity in suburban landscapes impacto de plantas nativas sobre la biodiversidad de aves y mariposas en paisajes suburbanos. Conservation Biology, 23(1):219–224, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bury:2004:WFR

[Bur04] R. Bruce Bury. Wildfire, fuel reduction, and herpetofaunas across diverse landscape mosaics in northwestern forests. Incendios, reducción de combustibles y herpetofaunas en mosaicos Paisajísticos diversos en bosques noroccidentales. Conservation Biology, 18(4):968–975, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buskirk:2002:LLM

[Bus02] Steven Buskirk. Lessons for large-mammal conservationists. Conservation Biology, 16(5):1443–1444, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buscher:2008:CNS

[Büs08] Bram E. Büscher. Conservation, neoliberalism, and social science: a critical reflection on the SCB 2007 annual meeting in South Africa. Conservation Biology, 22(2):229–231, April 2008.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Butchart:2003:UIR

[But03] Stuart Butchart. Using the IUCN red list criteria to assess species with declining populations. Conservation Biology, 17(5): 1200–1201, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bulte:2000:ESE

[BV00] Erwin Bulte and G. C. Van Kooten. Economic science, endangered species, and biodiversity loss. Ciencias económicas, especies en peligro y pérdida de biodiversidad. Conservation Biology, 14(1):113–119, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bourque:2001:ESC

[BV01a] Julie Bourque and Marc-André Villard. Effects of selection cutting and landscape-scale harvesting on the reproductive success of two neotropical migrant bird species. Efectos de corte por selección y cosecha a escala de paisaje en el exito reproductivo de dos especies de aves migratorias neotropicales. Conservation Biology, 15(1):184–195, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bulte:2001:SIP

[BV01b] Erwin H. Bulte and G. Cornelis Van Kooten. State intervention to protect endangered species: Why history and bad luck matter. Intervención estatal para proteger especies amenazadas: Porque la historia y la mala suerte son importantes. Conservation Biology, 15(6):1799–1803, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bakker:2004:GCD

[BV04] Victoria J. Bakker and Dirk H. Van Vuren. Gap-crossing decisions by the red squirrel, a forest-dependent small mammal. Decisiones para atravesar claros de la ardilla roja, un pequeño mamífero dependiente de bosque. Conservation Biology, 18(3): 689–697, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bock:2002:PRA

[BVH^{+}02] Carl E. Bock, Kerri T. Vierling, Sandra L. Haire, John D. Boone, and William W. Merkle. Patterns of rodent abundance on open-space grasslands in relation to suburban edges. Patrones de abundancia de roedores en pastizales a espacio abierto en relación con los bordes suburbanos. Conservation Biology, 16 (6):1653–1658, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:2002:ACW

[BVK02] Paul Beier, Maryann Van Drielen, and Bright O. Kankam. Avifaunal collapse in West African forest fragments. Colapso de la avifauna en fragmentos de bosque en África Occidental. Conservation Biology, 16(4):1097–1111, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bruinderink:2003:DCE

[BVL^{+}03] Geert Groot Bruinderink, Theo Van Der Sluis, Dennis Lammertsma, Paul Opdam, and Rogier Pouwels. Designing a coherent ecological network for large mammals in Northwestern Europe. Diseño de una red ecológica coherente para mamíferos en el noroeste de Europa. Conservation Biology, 17(2):549–557, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bekessy:2008:UCI

[BW08a] Sarah A. Bekessy and Brendan A. Wintle. Using carbon investment to grow the biodiversity bank. Conservation Biology, 22(3):510–513, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bhagwat:2008:ASO

[BW08b] Shonil A. Bhagwat and Katherine J. Willis. Agroforestry as a solution to the oil-palm debate. Conservation Biology, 22(6): 1368–1369, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bhagwat:2009:COP

[BW09] Shonil A. Bhagwat and Katherine J. Willis. Conservation in oilpalm landscapes. Conservation Biology, 23(2):245–246, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:2009:ESC

[BWS09] Thomas M. Brooks, S. Joseph Wright, and Douglas Sheil. Evaluating the success of conservation actions in safeguarding tropical forest biodiversity. Evaluación del Éxito de acciones de conservación para salvaguardar la biodiversidad de bosques tropicales. Conservation Biology, 23(6):1448–1457, December 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Byers:2004:RPP

[Bye04] John A. Byers. Rising into the Pantheon of public adoration. Conservation Biology, 18(4):1165–1166, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brook:2003:LRE

[BZD03] Amara Brook, Michaela Zint, and Raymond De Young. Landowners’ responses to an Endangered Species Act listing and implications for encouraging conservation. Respuestas de terratenientes a un listado relacionado al Acta de Especies en Peligro y sus implicaciones para promover la conservación. Conservation Biology, 17(6):1638–1649, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cook:2005:USA

[CA05] Rosamonde R. Cook and Peter J. Auster. Use of simulated annealing for identifying essential fish habitat in a multispecies context. Uso de recocido simulado para la identificación de hábitat esencial para peces en un contexto Multiespecífico. Conservation Biology, 19(3):876–886, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clewell:2006:MRE

[CA06] Andre F. Clewell and James Aronson. Motivations for the restoration of ecosystems. Motivaciones para la restauración de ecosistemas. Conservation Biology, 20(2):420–428, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:2000:CP

[Caf00a] Philip Cafaro. Conservation prophet. Conservation Biology, 14 (4):1217–1219, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:2000:MBE

[Caf00b] Philip Cafaro. Moving beyond eco-dread. Conservation Biology, 14(1):328, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:2000:NSW

[Caf00c] Philip Cafaro. Nature’s star witness. Conservation Biology, 14 (3):911–912, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:2001:TH

[Caf01] Philip Cafaro. Taming Henry. Conservation Biology, 15(5): 1470–1472, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:2004:OWE

[Caf04] Philip Cafaro. One world: the ethics of globalization. Conservation Biology, 18(2):585–586, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:2006:TBS

[Caf06] Philip Cafaro. Telling the better story. Conservation Biology, 20 (1):255–257, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:2007:SD

[Caf07] Phil Cafaro. Split decision. Conservation Biology, 21(2):888– 890, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coomes:2003:FPR

[CAFL03] David A. Coomes, Robert B. Allen, David M. Forsyth, and William G. Lee. Factors preventing the recovery of New Zealand forests following control of invasive deer. Factores que previenen la recuperación de bosques de Nueva Zelanda después del control del venado invasor. Conservation Biology, 17(2):450–459, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Calder:2002:DD

[Cal02] William A. Calder. Diversity in the desert. Conservation Biology, 16(1):280–281, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Calvete:2006:MEP

[Cal06] Carlos Calvete. Modeling the effect of population dynamics on the impact of rabbit hemorrhagic disease. Modelado del efecto de la dinámica poblacional sobre el impacto de enfermedad Hemorrágica de conejos. Conservation Biology, 33(5): 1232–1241, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Campbell:2005:OOI

[Cam05] Lisa M. Campbell. Overcoming obstacles to interdisciplinary research. Conservation Biology, 19(2):574–577, April 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cameron:2008:PCR

[Cam08] Elissa Z. Cameron. Productivity in conservation research in the Southern Hemisphere. Conservation Biology, 22(2):232–233, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caro:2000:FS

[Car00] Tim Caro. Focal species. Conservation Biology, 14(6):1569– 1570, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caro:2002:L

[Car02a] Tim Caro. Letters. Conservation Biology, 16(2):286–287, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2002:BRC

[Car02b] C. Ronald Carroll. A biophysical reality check for economists. Conservation Biology, 16(2):568–569, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carruthers:2004:BCI

[Car04] Raymond I. Carruthers. Biological control of invasive species, a personal perspective. Conservation Biology, 18(1):54–57, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cullen:2005:LRB

[CAR05] Laury Cullen, Jr., Keith Alger, and Denise M. Rambaldi. Land reform and biodiversity conservation in Brazil in the 1990s: Conflict and the articulation of mutual interests. Reforma agraria y conservación de biodiversidad en Brasil en los 1990s: Conflictos y la articulación de intereses mutuos. Conservation Biology, 19(3):747–755, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carthy:2006:OMC

[Car06] Raymond R. Carthy. An Owner’s manual for the chelonioidae. Conservation Biology, 20(1):259–260, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2007:IEC

[Car07] Carlos Carroll. Interacting effects of climate change, landscape conversion, and harvest on carnivore populations at the range margin: Marten and lynx in the Northern Appalachians. Interacción de los efectos del cambio climático, Conversión del paisaje y cosecha de poblaciones de carnívoros en el margen de su distribución: Marta y lince en el Norte de las Apalaches. Conservation Biology, 21(4):1092–1104, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

CassadyStClair:2003:CPR

[Cas03] Colleen Cassady St. Clair. Comparative permeability of roads, rivers, and meadows to songbirds in Banff National Park. Permeabilidad comparada de caminos, ríos y praderas para aves canoras en el Parque Nacional Banff. Conservation Biology, 17 (4):1151–1160, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cardillo:2001:BSR

[CB01a] Marcel Cardillo and Lindell Bromham. Body size and risk of extinction in Australian mammals. Tamaño del cuerpo y riesgo de extinción en mamíferos australianos. Conservation Biology, 15(5):1435–1440, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cross:2001:ULR

[CB01b] Paul C. Cross and Steven R. Beissinger. Using logistic regression to analyze the sensitivity of PVA models: a comparison of methods based on African wild dog models. Utilización de la Regresión Logística para analizar la sensibilidad de modelos de AVP: una comparación de métodos basados en modelos de perros silvestres africanos. Conservation Biology, 15(5):1335–1346, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:2000:LTE

[CBG^{+}00] Colin A. Chapman, Sophia R. Balcomb, Thomas R. Gillespie, Joseph P. Skorupa, and Thomas T. Struhsaker. Long-term effects of logging on African primate communities: a 28-year comparison from Kibale National Park, Uganda. Efectos de largo plaza de la tala en comunidades de primates africanos: Una comparación de 28 años en el Parque Nacional Kibale, Uganda. Conservation Biology, 14(1):207–217, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crumpacker:2001:ICW

[CBH01] David W. Crumpacker, Elgene O. Box, and E. Dennis Hardin. Implications of climatic warming for conservation of native trees and shrubs in Florida. Implicaciones del calentamiento global en la conservación de arboles y arbustos nativos de Florida. Conservation Biology, 15(4):1008–1020, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cederbaum:2004:EAC

[CCC04] Sandra B. Cederbaum, John P. Carroll, and Robert J. Cooper. Effects of alternative cotton agriculture on avian and arthropod populations. Efectos de agricultura alternativa de algodón sobre poblaciones de aves y artrópodos. Conservation Biology, 18(5): 1272–1282, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cully:2003:IEP

[CCH03] Anne C. Cully, Jack F. Cully, Jr., and Ronald D. Hiebert. Invasion of exotic plant species in tallgrass prairie fragments. Invasión de especies exóticas en fragmentos de pastizal alto. Conservation Biology, 17(4):990–998, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Callicott:2000:NCC

[CCM00a] J. Baird Callicott, Larry B. Crowder, and Karen Mumford. Normative concepts in conservation biology: Reply to Willers and Hunter. Conservation Biology, 14(2):575–578, April 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cork:2000:IIE

[CCM00b] Steven J. Cork, Tim W. Clark, and Nicole Mazur. Introduction: an interdisciplinary effort for koala conservation. Conservation Biology, 14(3):606–609, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cornelius:2000:EHF

[CCM00c] Cintia Cornelius, Hernán Cofré, and Pablo A. Marquet. Effects of habitat fragmentation on bird species in a relict temperate forest in semiarid Chile. Efectos de la fragmentación del hábitat en especies de aves en un bosque relicto templado semiárido de Chile. Conservation Biology, 14(2):534–543, April 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:2003:FFR

[CCS^{+}03] Lauren J. Chapman, Colin A. Chapman, Pamela J. Schofield, John P. Olowo, Les Kaufman, Ole Seehausen, and Richard Ogutu-Ohwayo. Fish faunal resurgence in Lake Nabugabo, East Africa. Resurgimiento de la faunaíctica en el Lago Nabugabo, África Oriental. Conservation Biology, 17(2):500–511, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Campbell:2005:FGE

[CD05] Karl Campbell and C. Josh Donlan. Feral goat eradications on islands. Erradicaciones de cabras ferales en islas. Conservation Biology, 19(5):1362–1374, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cuaron:2007:STS

[CD07] Alfredo D. Cuarón and Paloma C. De Grammont. Shortcomings of threatened species categorization systems: Reply to Soberón and Medellín. Conservation Biology, 21(5):1368–1370, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chiarucci:2005:UVP

[CDD^{+}05] Alessandro Chiarucci, Francesca D’Auria, Vincenzo De Dominicis, Angela Laganà, Claudia Perini, and Elena Salerni. Using vascular plants as a surrogate taxon to maximize fungal species richness in reserve design. Utilización de plantas vasculares como un taxón sustituto para maximizar la riqueza de especies de hongos en el diseño de reservas. Conservation Biology, 19(5):1644– 1652, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cinner:2009:SFA

[CDM09] J. E. Cinner, T. Daw, and T. R. Mcclanahan. Socioeconomic factors that affect artisanal fishers’ readiness to exit a declining fishery. Factores socioeconómicos que afectan la disponibilidad de pescadores artesanales para abandonar una pesquería en declinación. Conservation Biology, 23(1):124–130, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cox:2000:PEP

[CE00] Paul Alan Cox and Thomas Elmqvist. Pollinator extinction in the Pacific islands. Conservation Biology, 14(5):1237–1239, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caro:2005:USS

[CES05] Tim Caro, John Eadie, and Andrew Sih. Use of substitute species in conservation biology. El uso de especies sustitutas en biología de la conservación. Conservation Biology, 19(6): 1821–1826, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carr:2001:ERT

[CF01] Laurie W. Carr and Lenore Fahrig. Effect of road traffic on two amphibian species of differing vagility. Efectos del tráfico carretero sobre dos especies de anfibios de diferente movilidad. Conservation Biology, 15(4):1071–1078, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

C:2002:FCF

[CFC02] Eduardo Toral C., Peter Feinsinger, and Martha L. Crump. Frogs and a cloud-forest edge in Ecuador ranas y el borde de un bosque nublado ecuatonano. Conservation Biology, 16(3): 735–744, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Creel:2002:SAG

[CFH^{+}02] Scott Creel, Jennifer E. Fox, Amanda Hardy, Jennifer Sands, Bob Garrott, and Rolf O. Peterson. Snowmobile activity and glucocorticoid stress responses in wolves and elk. Actividad de vehículos para nieve y respuestas de stress glucocorticoide en lobos y alces. Conservation Biology, 16(3):809–814, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clements:2009:ITC

[CFO^{+}09] Reuben Clements, River Foo, Syahaneem Othman, Umi Rahman, Sharifah Ruqaiyah Syed Mustafa, and Rahayu Zulkifli. Islam, turtle conservation, and coastal communities. Conservation Biology, 23(3):516–517, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cassinello:2001:RBC

[CGR01] Jorge Cassinello, Montserrat Gomendio, and Eduardo R. S. Roldan. Relationship between coefficient of inbreeding and parasite burden in endangered gazelles. Relación entre el coeficiente de consanguinidad y la carga parasitaria en ungulados en peligro de extinción. Conservation Biology, 15(4):1171–1174, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cremene:2005:ASL

[CGR^{+}05] Cristina Cremene, Gheorghe Groza, Laszlo Rakosy, Anatoli A. Schileyko, Anette Baur, Andreas Erhardt, and Bruno Baur. Alterations of steppe-like grasslands in Eastern Europe: a threat to regional biodiversity hotspots alteraciones de pastizales similares a estepas en Europa oriental: una amenaza para sitios de importancia para la biodiversidad regionales. Conservation Biology, 19(5):1606–1618, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cordeiro:2001:LRT

[CH01] N. J. Cordeiro and H. F. Howe. Low recruitment of trees dispersed by animals in African forest fragments. Reclutamiento bajo de Árboles dispersados por animales en fragmentos de bosque africanos. Conservation Biology, 15(6):1733–1741, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Charnley:2006:NFP

[Cha06] Susan Charnley. The northwest forest plan as a model for broadscale ecosystem management: a social perspective. El plan forestal del noroeste como un modelos para la gestión de ecosistemas a gran escala. Conservation Biology, 20(2):330–340, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chan:2008:VAC

[Cha08] Kai M. A. Chan. Value and advocacy in conservation biology: Crisis discipline or discipline in crisis? Conservation Biology, 22(1):1–3, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2002:IUE

[CHBK02] J. Alan Clark, Jonathan M. Hoekstra, P. Dee Boersma, and Peter Kareiva. Improving U.S. Endangered Species Act recovery plans: Key findings and recommendations of the SCB Recovery Plan Project. Mejorando los planes de recuperación del Acta de Especies en Peligro de E.U.A.: Hallazgos clave y recomendaciones del Proyecto Plan de Recuperación de la SBC. Conservation Biology, 16(6):1510–1519, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2003:L

[CHBK03] J. Alan Clark, Jonathan M. Hoekstra, P. Dee Boersma, and Peter Kareiva. Letters. Conservation Biology, 17(3):655, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cheng:2002:FDP

[Che02a] Antony S. Cheng. Further debate on the politics of place. Conservation Biology, 16(4):1165–1168, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cheng:2002:MCP

[Che02b] Antony S. Cheng. The maturation of common-pool resources management. Conservation Biology, 16(3):847–850, June 2002.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chen:2009:CSA

[Che09] You-Hua Chen. Combining the species-area-habitat relationship and environmental cluster analysis to set conservation priorities: a study in the Zhoushan Archipelago, China. Combinación de la relación de Especies-Área–Hábitat y el análisis clúster ambiental para definir prioridades de conservación: un estudio en el Archipiélago Zhoushan, China. Conservation Biology, 23(3): 537–545, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cork:2000:IHM

[CHF00] Steven J. Cork, Ian D. Hume, and William J. Foley. Improving habitat models and their utility in koala conservation. Modelos de mejoramiento del hábitat y su utilidad en la conservación del koala. Conservation Biology, 14(3):660–668, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Corlatti:2009:AWO

[CHFR09] Luca Corlatti, Klaus Hackländer, and Fredy Frey-Roos. Ability of wildlife overpasses to provide connectivity and prevent genetic isolation. Capacidad de los pasos elevados para vida silvestre para proporcionar conectividad y prevenir el aislamiento genético. Conservation Biology, 23(3):548–556, June 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chiarello:2000:DPS

[Chi00a] Adriano G. Chiarello. Density and population size of mammals in remnants of Brazilian Atlantic forest. Densidad y tamaño poblacional de mamíferos en remanentes del bosque del Atlántico brasileño. Conservation Biology, 14(6):1649–1657, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chicchon:2000:CTM

[Chi00b] Avecita Chicchón. Conservation theory meets practice. Conservation Biology, 14(5):1368–1369, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Child:2009:TIU

[Chi09a] Matthew F. Child. The Thoreau Ideal as a unifying thread in the conservation movement. Conservation Biology, 23(2):241– 243, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Child:2009:TCC

[Chi09b] Matthew F. Child. Two cultures of conservation. Conservation Biology, 23(5):1071–1072, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cabeza:2005:EBA

[CHMP05] Mar Cabeza, Jenni Hottola, Evgeniy Meyke, and Marcelo A. H. Penna. Extending the benefits of attending a conference abroad. Conservation Biology, 19(6):1683, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Choquenot:2006:BMC

[Cho06] David Choquenot. Bioeconomic modeling in conservation pest management: Effect of stoat control on extinction risk of an indigenous New Zealand passerine, Mohua ochrocephala. Modelado bioeconómico en la gestión de plagas de conservación: Efecto del control de armiños sobre el riesgo de extinción de una paserina nativa de Nueva Zelanda, Mohua ochrocephala. Conservation Biology, 20(2):480–489, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Christensen:2008:SAF

[Chr08] Norman L. Christensen, Jr. Sustaining America’s forest legacy. Conservation Biology, 22(6):1378–1379, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carey:2001:ADE

[CHW^{+}01] Cynthia Carey, W. Ronald Heyer, John Wilkinson, Ross A. Alford, J. W. Arntzen, Tim Halliday, Laura Hungerford, Karen R. Lips, Elizabeth M. Middleton, Stan A. Orchard, and A. Stanley Rand. Amphibian declines and environmental change: Use of remote-sensing data to identify environmental correlates. Declinaciones de anfibios y cambio ambiental: Uso de datos de percepción remota para identificar correlaciones ambientales. Conservation Biology, 15(4):903–913, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2008:IBS

[CJ08] Carlos Carroll and Devin S. Johnson. The importance of being spatial (and reserved): Assessing northern spotted owl habitat relationships with hierarchical Bayesian models. La importancia de ser espacial (y reservado): Evaluación de las relaciones del hábitat del búho Strix occidentalis con modelos Bayesianos Jerárquicos. Conservation Biology, 22(4):1026–1036, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Campton:2005:WCT

[CK05] Donald E. Campton and Lynn R. Kaeding. Westslope cutthroat trout, hybridization, and the U.S. Endangered Species Act. Conservation Biology, 19(4):1323–1325, August 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cowling:2004:NCR

[CKF^{+}04] Richard M. Cowling, Andrew T. Knight, Daniel P. Faith, Simon Ferrier, Amanda T. Lombard, Amanda Driver, Mathieu Rouget, Kristal Maze, and Philip G. Desmet. Nature conservation requires more than a passion for species. Conservation Biology, 18(6):1674–1676, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chase:2000:SSI

[CKL^{+}00] Mary K. Chase, William B. Kristan III, Anthony J. Lynam, Mary V. Price, and John T. Rotenberry. Single species as indicators of species richness and composition in California coastal sage scrub birds and small mammals. Especies individuales como indicadores de la riqueza y la composición de especies de aves y mamíferos Pequeños en un matorral costero de salvia en California. Conservation Biology, 14(2):474–487, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2001:DPO

[Cla01] Tim W. Clark. Developing policy-oriented curricula for conservation biology: Professional and leadership education in the public interest. Desarrollo de planes de estudio con orientación política para la biología de la conservación: Educación profesional y de liderazgo para beneficio del público. Conservation Biology, 15(1):31–39, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2009:ANS

[Cla09] Howard O. Clark, Jr. Advances in noninvasive sampling through a better understanding of target-species ecology. Conservation Biology, 23(5):1336–1338, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cam:2003:DAM

[CLBC03] Emmanuelle Cam, Lynn Lougheed, Russell Bradley, and Fred Cooke. Demographic assessment of a marbled murrelet population from capture–recapture data. Evaluación demográfica de una población de Brachyrampus marmoratus con datos de captura–recaptura. Conservation Biology, 17(4):1118–1126, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cunningham:2008:CER

[CLC^{+}08] Ross B. Cunningham, David B. Lindenmayer, Mason Crane, Damian Michael, Christopher Macgregor, Rebecca Montague-Drake, and Joern Fischer. The combined effects of remnant vegetation and tree planting on farmland birds. Efectos combinados de la vegetación remanente y la siembra de Árboles sobre aves de tierras agrícolas. Conservation Biology, 22(3):742–752, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cleary:2006:L

[Cle06a] David Cleary. Letters. Conservation Biology, 20(6):1567–1568, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cleary:2006:QES

[Cle06b] David Cleary. The questionable effectiveness of science spending by international conservation organizations in the tropics. La cuestionable efectividad del gasto en ciencia por organizaciones internacionales de conservación en los trópicos. Conservation Biology, 20(3):733–738, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cloyd:2004:DPG

[CLH^{+}04] Emily T. Cloyd, Melissa S. Lucash, Nicole E. Muir Hotaling, J. Hope Hornbeck, Heather N. Jensen, Byron J. Norelius, Stephen H. Reynolds, and Jacob D. Wickham. Diversity and productivity in grassland ecosystems. Conservation Biology, 18 (4):1171–1173, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Costa:2005:MCB

[CLMD05] Leonora Pires Costa, Yuri Luiz Reis Leite, Sérgio Lucena Mendes, and Albert David Ditchfield. Mammal conservation in Brazil. Conservación de mamíferos en Brasil. Conservation Biology, 19(3):672–679, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Collen:2009:MCV

[CLW^{+}09] Ben Collen, Jonathan Loh, Sarah Whitmee, Louise Mcrae, Rajan Amin, and Jonathan E. M. Baillie. Monitoring change in vertebrate abundance: the Living Planet Index. Monitoreo del cambio en la abundancia de vertebrados: El Índice del Planeta Viviente. Conservation Biology, 23(2):317–327, April 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cox:2007:CEE

[CLŽ^{+}07] T. M. Cox, R. L. Lewison, R. Žydelis, L. B. Crowder, C. Safina, and A. J. Read. Comparing effectiveness of experimental and implemented bycatch reduction measures: the ideal and the real. Comparación de la efectividad de medidas experimentales e implementadas para la Reducción de capturas incidentales: lo ideal y lo real. Conservation Biology, 21(5):1155–1164, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cabeza:2003:SSA

[CM03] Mar Cabeza and Atte Moilanen. Site-selection algorithms and habitat loss. Algoritmos de selección de sitios y pérdida de hábitat. Conservation Biology, 17(5):1402–1413, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Christen:2006:RRP

[CM06] Douglas Christen and Glenn Matlack. The role of roadsides in plant invasions: a demographic approach. El papel de bordes carreteras en invasiones de plantas: un enfoque Demográfico. Conservation Biology, 20(2):385–391, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clarke:2006:ISS

[CMA^{+}06] Shelley C. Clarke, Jennifer E. Magnussen, Debra L. Abercrombie, Murdoch K. Mcallister, and Mahmood S. Shivji. Identification of shark species composition and proportion in the Hong Kong shark fin market based on molecular genetics and trade records. Identificación de la composición y Proporción de especies de Tiburón en el mercado de aletas de Tiburón en Hong Kong con base en genética molecular y registros comerciales. Conservation Biology, 20(1):201–211, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2000:IGD

[CMBC00] Tim W. Clark, Nicole Mazur, Robert J. Begg, and Steven J. Cork. Interdisciplinary guidelines for developing effective koala conservation policy. Guías interdisciplinarias para el desarrollo efectivo de políticas de conservación del koala. Conservation Biology, 14(3):691–701, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2000:KCP

[CMC^{+}00] Tim W. Clark, Nicole Mazur, Steven J. Cork, Steve Dovers, and Ronnie Harding. Koala conservation policy process: Appraisal and recommendations. Proceso de las políticas de conservación: Evaluación y recomendaciones. Conservation Biology, 14(3):681–690, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ciani:2001:EWA

[CMC^{+}01] Andrea Camperio Ciani, Loredana Martinoli, Claudio Capiluppi, Mohamed Arahou, and Mohamed Mouna. Effects of water availability and habitat quality on bark-stripping behavior in Barbary macaques. Efectos de la disponibilidad de agua y la calidad del hábitat en la conducta descortezadora del macaco de Barbary. Conservation Biology, 15(1):259–265, February 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:2003:LSS

[CMC03] John W. Chapman, Todd W. Miller, and Eugene V. Coan. Live seafood species as recipes for invasion. Especies de mariscos vivas como recetas para invasión. Conservation Biology, 17(5):1386– 1395, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:2005:L

[CMC05] John W. Chapman, Todd W. Miller, and Eugene V. Coan. Letters: 2. Conservation Biology, 19(2):288–289, April 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Compton:2007:RKM

[CMCG07] Bradley W. Compton, Kevin Mcgarigal, Samuel A. Cushman, and Lloyd R. Gamble. A resistant-kernel model of connectivity for amphibians that breed in vernal pools. Un modelo de núcleo resistente de la conectividad para anfibios que se reproducen en charcos vernales. Conservation Biology, 21(3):788–799, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Culver:2000:OCF

[CMCH00] David C. Culver, Lawrence L. Master, Mary C. Christman, and Horton H. Hobbs III. Obligate cave fauna of the 48 contiguous United States fauna. Cavernícola obligada de los 48 estados contiguos de E.U.A. Conservation Biology, 14(2):386–401, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clayton:2000:EPE

[CMGSM00] L. M. Clayton, E. J. Milner-Gulland, D. W. Sinaga, and A. H. Mustari. Effects of a proposed ex situ conservation program on in situ conservation of the babirusa, an endangered suid. Efectos de un programa de conservación propuesto ex situ en la conservación in situ de la babirusa, una especie en peligro de extinción. Conservation Biology, 14(2):382–385, April 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cleaveland:2007:CRE

[CMK^{+}07] Sarah Cleaveland, Titus Mlengeya, Magai Kaare, Dan Haydon, Tiziana Lembo, M. Karen Laurenson, and Craig Packer. The conservation relevance of epidemiological research into carnivore viral diseases in the Serengeti. La relevancia para la conservación de la investigación Epidemiológica de enfermedades virales de carnívoros en el Serengeti. Conservation Biology, 21(3):612– 622, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cox:2002:BFP

[CML02] John J. Cox, David S. Maehr, and Jeffery L. Larkin. The biogeography of faunal place names in the United States. Biogeografía de nombres de lugares asociados a la fauna en los Estados Unidos. Conservation Biology, 16(4):1143–1150, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cinner:2005:CCB

[CMM05a] Joshua E. Cinner, Michael J. Marnane, and Tim R. Mcclanahan. Conservation and community benefits from traditional coral reef management at Ahus Island, Papua New Guinea. Beneficios de conservación y de la comunidad por la gestión tradicional de arrecifes de coral en la Isla Ahus, Papua Nueva Guinea. Conservation Biology, 19(6):1714–1723, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cinner:2005:TTT

[CMM^{+}05b] Joshua E. Cinner, Michael J. Marnane, Timothy R. Mcclanahan, Tracy H. Clark, and John Ben. Trade, tenure, and tradition: Influence of sociocultural factors on resource use in Melanesia. Comercio, posesión y tradición: Influencia de factores socioculturales sobre el uso de recursos en Melanesia. Conservation Biology, 19(5):1469–1477, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Canterbury:2000:BCH

[CMP^{+}00] Grant E. Canterbury, Thomas E. Martin, Daniel R. Petit, Lisa J. Petit, and David F. Bradford. Bird communities and habitat as ecological indicators of forest condition in regional monitoring. Comunidades de aves y hábitat como indicadores ecológicos de la condición del bosque en monitoreos regionales. Conservation Biology, 14(2):544–558, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cowlishaw:2005:SOB

[CMR05] Guy Cowlishaw, Samantha Mendelson, and J. Marcus Rowcliffe. Structure and operation of a bushmeat commodity chain in Southwestern Ghana. Estructura y Organización de una cadena de comercialización de carne silvestre en el suroeste de Ghana. Conservation Biology, 19(1):139–149, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cushman:2009:UED

[CMS09] Samuel A. Cushman, Kevin S. Mckelvey, and Michael K. Schwartz. Use of empirically derived source-destination models to map regional conservation corridors. Utilización de modelos fuente-destino Empíricamente derivados para cartografiar corredores de conservación regionales. Conservation Biology, 23 (2):368–376, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Campos:2006:SAN

[CN06] Marina T. Campos and Daniel C. Nepstad. Smallholders, the Amazon’s new conservationists. Conservation Biology, 20(5): 1553–1556, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chung:2004:ICS

[CNC04] Mi Yoon Chung, John D. Nason, and Myong Gi Chung. Implications of clonal structure for effective population size and genetic drift in a rare terrestrial orchid, Cremastra appendiculata. Implicaciones de la estructura clonal para el tamaño poblacional efectivo y la deriva génica en una orquídea terrestre rara, Cremastra appendiculata. Conservation Biology, 18(6):1515–1524, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2009:SSM

[CNHT09] Carlos Carroll, Reed F. Noss, Jodi Hilty, and Stephen C. Trombulak. Solving SCB’s membership crisis by reinvigorating the sections: Response to Schwartz et al. Conservation Biology, 23 (1):5–6, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cornelius:2001:EHA

[CNM01] Cintia Cornelius, Sergio A. Navarrete, and Pablo A. Marquet. Effects of human activity on the structure of coastal marine bird assemblages in Central Chile. Efectos de la actividad humana sobre la estructura de ensamblajes de aves marinas costeras en Chile Central. Conservation Biology, 15(5):1396–1404, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2004:EDP

[CNPS04] Carlos Carroll, Reed F. Noss, Paul C. Paquet, and Nathan H. Schumaker. Extinction debt of protected areas in developing landscapes. Deuda de extinción de Áreas protegidas en paisajes en desarrollo. Conservation Biology, 18(4):1110–1120, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coblentz:2003:WCB

[Cob03] Bruce Coblentz. Where conservation biology and conservation ballistics meet. Conservation Biology, 17(6):1858–1859, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cobb:2005:RPP

[Cob05] John B. Cobb, Jr. The responsibility of progressive protestants. Conservation Biology, 19(2):294, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cochrane:2001:SIB

[Coc01] Mark A. Cochrane. Synergistic interactions between habitat fragmentation and fire in evergreen tropical forests. Interacciones sinérgicas entre la fragmentación del hábitat y los incendios en bosques tropicales perennes. Conservation Biology, 15 (6):1515–1521, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cochrane:2004:MSA

[Coc04] Mark A. Cochrane. Making sustainability accessible and achievable. Conservation Biology, 18(5):1446–1447, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Colchester:2000:SDE

[Col00] Marcus Colchester. Self-determination or environmental determinism for indigenous peoples in tropical forest conservation. Conservation Biology, 14(5):1365–1367, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Colyvan:2001:RTS

[Col01] Mark Colyvan. Ranking in threatened species classification: Reply to Todd. Conservation Biology, 15(6):1814–1816, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coleman:2006:MU

[Col06] David C. Coleman. A masterful underview. Conservation Biology, 33(5):1328–1330, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Colwell:2009:AB

[Col09] Rita R. Colwell. Along the banks. Conservation Biology, 23(3): 783–784, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Corbet:2000:CCP

[Cor00] Sarah A. Corbet. Conserving compartments in pollination webs. Conservation Biology, 14(5):1229–1231, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cortes:2002:IUD

[Cor02] Enric Cortés. Incorporating uncertainty into demographic modeling: Application to shark populations and their conservation. Incorporación de la incertidumbre en el modelado Demográfico: Aplicación en poblaciones de tiburones y su conservación. Conservation Biology, 16(4):1048–1062, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Corn:2003:ABC

[Cor03] Paul Stephen Corn. Amphibian breeding and climate change: Importance of snow in the mountains. Conservation Biology, 17(2):622–625, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cordes:2004:SAS

[Cor04a] Colleen Cordes. Shifting the agenda of science and technology. Conservation Biology, 18(4):867–868, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Corkeron:2004:WWI

[Cor04b] Peter J. Corkeron. Whale watching, iconography, and marine conservation. Conservation Biology, 18(3):847–849, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Corkeron:2006:OVE

[Cor06] Peter J. Corkeron. Opposing views of the “Ecosystem approach” to fisheries management. Conservation Biology, 20(3):617–619, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cote:2005:ELB

[Côt05] Steeve D. Côté. Extirpation of a large black bear population by introduced white-tailed deer. Conservation Biology, 19(5): 1668–1671, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Couvet:2002:DER

[Cou02] Denis Couvet. Deleterious effects of restricted gene flow in fragmented populations. Efectos nocivos del flujo restringido de genes en poblaciones fragmentadas. Conservation Biology, 16 (2):369–376, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Comstock:2003:ANM

[COW03] Kenine E. Comstock, Elaine A. Ostrander, and Samuel K. Wasser. Amplifying nuclear and mitochondrial DNA from African elephant ivory: a tool for monitoring the ivory trade. Amplificación de ADN nuclear y mitocondrial de marfil de elefante africano: Una herramienta para monitorear el comercio de marfil. Conservation Biology, 17(6):1840–1843, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2000:DMS

[CP00] Steven S. Carroll and David L. Pearson. Detecting and modeling spatial and temporal dependence in conservation biology. Detección y modelado de la dependencia espacial y temporal en la biología de la conservación. Conservation Biology, 14(6): 1893–1897, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cadenasso:2001:EES

[CP01] M. L. Cadenasso and S. T. A. Pickett. Effect of edge structure on the flux of species into forest interiors. Efecto de la estructura de bordes en el flujo de especies hacia el interior de bosques.

Conservation Biology, 15(1):91–97, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chust:2003:RSF

[CPD^{+}03] G. Chust, J. L. Pretus, D. Ducrot, A. Bedòs, and L. Deharveng. Response of soil fauna to landscape heterogeneity: Determining optimal scales for biodiversity modeling. Respuesta de la fauna del suelo a la heterogeneidad del paisaje: Determinación de escalas Óptimas para el modelado de biodiversidad. Conservation Biology, 17(6):1712–1723, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chazdon:2009:PSC

[CPD^{+}09] Robin L. Chazdon, Carlos A. Peres, Daisy Dent, Douglas Sheil, Ariel E. Lugo, David Lamb, Nigel E. Stork, and Scott E. Miller. The potential for species conservation in tropical secondary forests. El potencial de conservación de especies en bosques tropicales secundarios. Conservation Biology, 23(6):1406–1417, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cauley:2001:FSC

[CPDO01] Henry A. Cauley, Charles M. Peters, Richard Z. Donovan, and Jennifer M. O’Connor. Forest Stewardship Council forest certification. Conservation Biology, 15(2):311–312, April 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2009:LCC

[CPME09] C. J. Clark, J. R. Poulsen, R. Malonga, and P. W. Elkan, Jr. Logging concessions can extend the conservation estate for Central African tropical forests. Las concesiones madereras pueden extender el estado de conservación de los bosques tropicales de África Central. Conservation Biology, 23(5):1281–1293, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Collinge:2003:ELH

[CPO03] Sharon K. Collinge, Kathleen L. Prudic, and Jeffrey C. Oliver. Effects of local habitat characteristics and landscape context on grassland butterfly diversity. Efectos de las características del hábitat local y el contexto del paisaje sobre la diversidad de mariposas de pastizal. Conservation Biology, 17(1):178–187, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clarke:2005:CLS

[CPR05] F. M. Clarke, D. V. Pio, and P. A. Racey. A comparison of logging systems and bat diversity in the neotropics. Una comparación de sistemas de explotación forestal y diversidad de murciélagos en el Neotrópico. Conservation Biology, 19(4):1194– 1204, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chan:2007:WAC

[CPR^{+}07] Kai M. A. Chan, Robert M. Pringle, Jai Ranganathan, Carol L. Boggs, Yvonne L. Chan, Paul R. Ehrlich, Peter K. Haff, Nicole E. Heller, Karim Al-Khafaji, and Dena P. Macmynowski. When agendas collide: Human welfare and biological conservation. Cuando las agendas chocan: Bienestar humano y conservación biológica. Conservation Biology, 21(1):59–68, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2003:ILC

[CPSS03] Carlos Carroll, Michael K. Phillips, Nathan H. Schumaker, and Douglas W. Smith. Impacts of landscape change on wolf restoration success: Planning a reintroduction program based on static and dynamic spatial models. Impacto del cambio de paisaje sobre el exito de restauración de lobos: Planeación de programas de reintroducción utilizando modelos espaciales estáticos y dinámicos. Conservation Biology, 17(2):536–548, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cascante:2002:EDT

[CQLF02] Alfredo Cascante, Mauricio Quesada, Jorge J. Lobo, and Eric A. Fuchs. Effects of dry tropical forest fragmentation on the reproductive success and genetic structure of the tree Samanea saman. Efectos de la fragmentación del bosque tropical seco sobre el Éxito reproductivo y la estructura genética del arbol Samanea saman. Conservation Biology, 16(1):137–147, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Craig:2002:CEF

[Cra02] Michael D. Craig. Comparative ecology of four passerine species in jarrah forests used for timber production in Southwestern Western Australia. Ecología comparativa de cuatro especies paserinas en bosques de jarrah usados para la producción de madera en el suroeste de Australia Occidental. Conservation Biology, 16(6):1609–1619, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

CardosoDaSilva:2005:FAA

[CRD05] José Maria Cardoso Da Silva, Anthony B. Rylands, and Gustavo A. B. Da Fonseca. The fate of the Amazonian areas of endemism. El destino de las Áreas de endemismo amazónicas. Conservation Biology, 19(3):689–694, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Creel:2001:FFM

[Cre01a] Scott Creel. Four factors modifying the effect of competition on carnivore population dynamics as illustrated by African wild dogs. Cuatro factores que modifican el efecto de la competencia en la dinámica poblacional de carnívoros ilustrado por perros silvestres africanos. Conservation Biology, 15(1):271–274, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crews:2001:CGK

[Cre01b] David Crews. Cutting the Gordian knot. Conservation Biology, 15(3):808–809, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chown:2001:WHS

[CRGG01] Steven L. Chown, Ana S. L. Rodrigues, Niek J. M. Gremmen, and Kevin J. Gaston. World heritage status and conservation of Southern Ocean islands. Conservation Biology, 15(3):550–557, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crist:2000:ADM

[Cri00] Patrick J. Crist. Animal-distribution modeling in gap analysis: an evolving science. Conservation Biology, 14(5):1224, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [SBH00].

Crooks:2002:RSM

[Cro02] Kevin R. Crooks. Relative sensitivities of mammalian carnivores to habitat fragmentation. Sensibilidad relativa a la fragmentación del hábitat de mamíferos carnívoros. Conservation Biology, 16(2):488–502, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chhatre:2005:PIB

[CS05] Ashwini Chhatre and Vasant Saberwal. Political incentives for biodiversity conservation. Incentivos políticos para la conservación de biodiversidad. Conservation Biology, 19(2):310–317, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Castellon:2006:ETM

[CS06] Traci D. Castellón and Kathryn E. Sieving. An experimental test of matrix permeability and corridor use by an endemic understory bird. Una prueba experimental de la permeabilidad de la matriz y el uso de corredores por una especie de sotobosque endémica. Conservation Biology, 20(1):135–145, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chan:2007:CRD

[CS07a] Kai M. A. Chan and Terre Satterfield. Conservation resentment dissected. Conservation Biology, 21(5):1380–1382, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crawford:2007:ECT

[CS07b] John A. Crawford and Raymond D. Semlitsch. Estimation of core terrestrial habitat for stream-breeding salamanders and delineation of riparian buffers for protection of biodiversity. Estimación del hábitat terrestre núcleo para salamandras que se reproducen en arroyos y definición de Áreas de amortiguamiento ribereñas para la protección de biodiversidad. Conservation Biology, 21(1):152–158, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cinner:2007:STA

[CSB07] Joshua E. Cinner, Stephen G. Sutton, and Trevor G. Bond. Socioeconomic thresholds that affect use of customary fisheries management tools. Umbrales socioeconómicos que afectan el uso de herramientas convencionales para el manejo de pesquerías. Conservation Biology, 22(1):1603–1611, December 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crooks:2001:ECB

[CSBS01] Kevin R. Crooks, Andrew V. Suarez, Douglas T. Bolger, and Michael E. Soulé. Extinction and colonization of birds on habitat islands. Extinciones y colonizaciones de aves en hábitats insulares. Conservation Biology, 15(1):159–172, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Consiglio:2006:DPD

[CSM^{+}06] Trisha Consiglio, George E. Schatz, Gordon Mcpherson, Porter P. Lowry II, Johny Rabenantoandro, Zachary S. Rogers, Raymond Rabevohitra, and David Rabehevitra. Deforestation and plant diversity of Madagascar’s littoral forests. Deforestación y la diversidad de plantas en los bosques litorales de Madagascar. Conservation Biology, 21(2):1799–1803, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Castelletta:2000:HEF

[CSS00] Marjorie Castelletta, Navjot S. Sodhi, and R. Subaraj. Heavy extinctions of forest avifauna in Singapore: Lessons for biodiversity conservation in Southeast Asia. Extinciones mayores de avifauna de bosque en singapur: Lecciones para la conservación en Asia sudoriental. Conservation Biology, 14(6):1870–1880, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

CardosoDaSilva:2001:FAF

[CT01] Jose Maria Cardoso Da Silva and Marcelo Tabarelli. The future of the Atlantic forest in northeastern Brazil. Conservation Biology, 15(4):819–820, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chisholm:2007:NHS

[CT07] Ryan Chisholm and Robert Taylor. Null-hypothesis significance testing and the critical weight range for Australian mammals. La prueba de significancia de la hipótesis nula y el rango de peso crítico para mamíferos australianos. Conservation Biology, 22 (1):1641–1645, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Collins:2002:GDM

[CTNR02] Timothy M. Collins, Joel C. Trexler, Leo G. Nico, and Timothy A. Rawlings. Genetic diversity in a morphologically conservative invasive taxon: Multiple introductions of swamp eels to the Southeastern United States. Diversidad genética en un taxón invasivo Morfológicamente conservativo: Introducciones múltiples de anguila de pantano en el sureste de Estados Unidos. Conservation Biology, 16(4):1024–1035, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chalfoun:2002:NPF

[CTR02] Anna D. Chalfoun, Frank R. Thompson III, and Mary J. Ratnaswamy. Nest predators and fragmentation: a review and metaanalysis. Depredadores de nidos y fragmentación: una revisión y meta–análisis. Conservation Biology, 16(2):306–318, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cuaron:2000:ELC

[Cua00] Alfredo D. Cuarón. Effects of land-cover changes on mammals in a neotropical region: a modeling approach. Efectos de los cambios en la cobertura del suelo sobre los mamíferos de una Región neotropical: una estrategia de modelaje. Conservation Biology, 14(6):1676–1692, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cuaron:2002:NMP

[Cua02] Alfredo D. Cuarón. Neotropical mammalogy by popular demand. Conservation Biology, 16(3):844–845, June 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cunningham:2000:EHF

[Cun00] Saul A. Cunningham. Effects of habitat fragmentation on the reproductive ecology of four plant species in mallee woodland. Efectos de la fragmentación del hábitat en la ecología reproductiva de cuatro especies de plantas en bosques arbolados. Conservation Biology, 14(3):758–768, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtin:2002:ISC

[Cur02a] Charles G. Curtin. Integration of science and community-based conservation in the Mexico/U.S. borderlands. Integración de la ciencia y la conservación basada en la comunidad en las tierras de la frontera entre México y Estados Unidos. Conservation Biology, 16(4):880–886, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtin:2002:LGR

[Cur02b] Charles G. Curtin. Livestock grazing, rest, and restoration in arid landscapes. Conservation Biology, 16(3):840–842, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtin:2003:CEP

[Cur03] Charles Curtin. Culture, ecology, and policy in the old and New West. Conservation Biology, 17(4):1188–1190, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtin:2007:ILE

[Cur07] Charles Curtin. Integrating landscape and ecosystems approaches through science-based collaborative conservation. Conservation Biology, 21(4):1117–1119, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtis:2008:UPV

[CV08] Janelle M. R. Curtis and Amanda C. J. Vincent. Use of population viability analysis to evaluate CITES trade-management options for threatened marine fishes. Utilización del análisis de viabilidad poblacional para evaluar las opciones de manejo de comercio de CITES para peces marinos amenazados. Conservation Biology, 22(5):1225–1232, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ceballos:2009:CCA

[CVB^{+}09] Gerardo Ceballos, Mariana M. Vale, Cristian Bonacic, Julio Calvo-Alvarado, Rurik List, Nora Bynum, Rodrigo A. Medellín, Javier A. Simonetti, and Jon Paul Rodríguez. Conservation challenges for the Austral and neotropical America section. Retos de conservación para la sección América Austral y neotropical. Conservation Biology, 23(4):811–817, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cagnolo:2009:HFS

[CVS^{+}09] Luciano Cagnolo, Graciela Valladares, Adriana Salvo, Marcelo Cabido, and Marcelo Zak. Habitat fragmentation and species loss across three interacting trophic levels: Effects of life-history and food-web traits. Fragmentación de hábitat y pérdida de especies en tres niveles tróficos interactuantes: Efectos de la historia de vida y los atributos de la red Trófica. Conservation Biology, 23(5):1167–1175, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clevenger:2000:FIE

[CW00] Anthony P. Clevenger and Nigel Waltho. Factors influencing the effectiveness of wildlife underpasses in Banff National Park, Alberta, Canada. Factores que influencían la efectivadad de pasadizos para vida silvestre en el Parque Nacional Banff, Alberta, Canada. Conservation Biology, 14(1):47–56, February 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cooper:2002:EED

[CW02] Caren B. Cooper and Jeffrey R. Walters. Experimental evidence of disrupted dispersal causing decline of an Australian passerine in fragmented habitat. Evidencia experimental de la disrupción de la dispersión que ocasiona la pérdida de una especie paserina Australiana en un hábitat fragmentado. Conservation Biology, 16(2):471–478, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtin:2008:GPC

[CW08] Charles Curtin and David Western. Grasslands, people, and conservation: Over-the-horizon learning exchanges between African and American pastoralists. Pastizales, gente y conservación: Intercambios de aprendizaje entre pastores africanos y americanos. Conservation Biology, 22(4):870–877, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cotton:2009:PCE

[CW09] Samuel Cotton and Claus Wedekind. Population consequences of environmental sex reversal. Consecuencias poblacionales de la Reversión de sexo ambiental Cotton & Wedekind. Conservation Biology, 23(1):196–206, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carrillo:2000:MMP

[CWC00] Eduardo Carrillo, Grace Wong, and Alfredo D. Cuarón. Monitoring mammal populations in Costa Rican protected areas under different hunting restrictions. Monitoreo de poblaciones de mamíferos en Áreas protegidas de Costa Rica con diferentes restricciones de cacería. Conservation Biology, 14(6):1580–1591, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clevenger:2002:GGE

[CWCG02] Anthony P. Clevenger, Jack Wierzchowski, Bryan Chruszcz, and Kari Gunson. GIS-Generated, expert-based models for identifying wildlife habitat linkages and planning mitigation passages. Modelos generados con GIS y basados en expertos para la identificación de conexiones del hábitat para vida silvestre y la planeación de pasajes de mitigación. Conservation Biology, 16(2):503–514, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Campbell:2007:LTE

[CWH07] Steven P. Campbell, Jack W. Witham, and Malcolm L. Hunter, Jr. Long-term effects of group-selection timber harvesting on abundance of forest birds. Efectos de largo plazo de la cosecha selectiva de madera sobre la abundancia de aves forestales. Conservation Biology, 21(5):1218–1229, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cabin:2000:ELT

[CWL^{+}00] Robert J. Cabin, Stephen G. Weller, David H. Lorence, Tim W. Flynn, Ann K. Sakai, Darren Sandquist, and Lisa J. Hadway. Effects of long-term ungulate exclusion and recent alien species control on the preservation and restoration of a Hawaiian tropical dry forest. Efectos a largo plazo de la exclusión de ungulados y el reciente control de especies exóticas en la conservación y restauración de un bosque tropical seco hawaiano. Conservation Biology, 14(2):439–453, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cameron:2008:ECA

[CWM08] Susan E. Cameron, Kristen J. Williams, and David K. Mitchell. Efficiency and concordance of alternative methods for minimizing opportunity costs in conservation planning. Eficiencia y concordancia de métodos alternativos para minimizar los costos de oportunidad en la planificación de la conservación. Conservation Biology, 22(4):886–896, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cassel:2001:EPS

[CWNW01] Anna Cassel, Jack Windig, Sören Nylin, and Christer Wiklund. Effects of population size and food stress on fitness-related characters in the scarce Heath, a rare butterfly in Western Europe. Efectos del tamaño poblacional y el estrés alimenticio en los caracteres relacionados con la adaptación de la mariposa Coenonympha hero, una mariposa rara de Europa occidental. Conservation Biology, 15(6):1667–1673, December 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clausen:2008:EGM

[CY08] Rebecca Clausen and Richard York. Economic growth and marine biodiversity: Influence of human social structure on decline of marine trophic levels. Crecimiento económico y biodiversidad marina: Influencia de la estructura social humana sobre la declinación de niveles tróficos marinos. Conservation Biology, 22 (2):458–466, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:2002:JSR

[Cze02a] Brian Czech. Julian Simon redux. Conservation Biology, 16 (2):570–572, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:2002:TAC

[Cze02b] Brian Czech. A transdisciplinary approach to conservation land acquisition. Un enfogue transdisciplinario para la adquisición de tierras para conservación. Conservation Biology, 16(6):1488– 1497, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:2003:TPB

[Cze03] Brian Czech. Technological progress and biodiversity conservation: a dollar spent, a dollar burned. Conservation Biology, 17 (5):1455–1457, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:2005:CNW

[Cze05] Brian Czech. The capacity of the national wildlife refuge system to conserve threatened and endangered animal species in the United States la capacidad del sistema nacional de refugios de vida silvestre para conservar especies amenazadas y en peligro en Estados Unidos. Conservation Biology, 19(4):1246–1253, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:2006:IRB

[Cze06] Brian Czech. If Rome is burning, why are we fiddling? Conservation Biology, 20(6):1563–1565, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:2007:ASD

[Cze07a] Brian Czech. The advocacy and science divide. Conservation Biology, 21(4):902–903, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:2007:L

[Cze07b] Brian Czech. Letters. Conservation Biology, 21(5):1137–1138, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:2008:PRC

[Cze08] Brian Czech. Prospects for reconciling the conflict between economic growth and biodiversity conservation with technological progress. Perspectivas para la reconciliación del conflicto entre crecimiento económico y conservación de la biodiversidad mediante progreso tecnológico. Conservation Biology, 22(6): 1389–1398, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DaFonseca:2003:CSN

[Da 03] Gustavo A. B. Da Fonseca. Conservation science and NGOs. Conservation Biology, 17(2):345–347, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daugherty:2002:NRM

[DA02] Charles H. Daugherty and Fred W. Allendorf. The numbers that really matter. Conservation Biology, 16(2):283–284, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dimond:2007:AHS

[DA07] Wendy J. Dimond and Doug P. Armstrong. Adaptive harvesting of source populations for translocation: a case study with New Zealand robins. Cosecha adaptativa de poblaciones fuente para translocación: un caso de estudio con petirrojos de Nueva Zelanda. Conservation Biology, 21(1):114–124, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dietz:2004:DLC

[DAB^{+}04] James M. Dietz, Rina Aviram, Sophia Bickford, Karen Douthwaite, Amy Goodstine, Jose-Luis Izursa, Stephanie Kavanaugh, Katie Maccarthy, Michelle O’Herron, and Keri Parker. Defining leadership in conservation: a view from the top. Conservation Biology, 18(1):274–278, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunham:2006:USM

[DAB06] Amy E. Dunham, H. Resit Akçakaya, and Todd S. Bridges. Using scalar models for precautionary assessments of threatened species. Utilización de modelos escalares para evaluaciones preventivas de especies amenazadas. Conservation Biology, 20(5): 1499–1506, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drew:2008:ERC

[DAKB08] Joshua Drew, Gerald R. Allen, Les Kaufman, and Paul H. Barber. Endemism and regional color and genetic differences in five putatively cosmopolitan reef fishes. Endemismo y color regional y diferencias genéticas en cinco especies de arrecife supuestamente cosmopolitas. Conservation Biology, 22(4):965–975, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daly:2000:L

[Dal00a] Herman E. Daly. Letter. Conservation Biology, 14(1):4, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [Tro00].

Daly:2000:WND

[Dal00b] Herman E. Daly. What neo-Darwinism does not explain: Response to Trombulak and Matter and McPherson. Conservation Biology, 14(4):1206–1207, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daly:2002:WES

[Dal02] Douglas C. Daly. A wealth of endemic species with poor prospects. Conservation Biology, 16(4):1168–1169, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dalton:2005:BBF

[Dal05] Tracey Morin Dalton. Beyond biogeography: a framework for involving the public in planning of U.S. Marine Protected Areas. Más allá de la biogeografía: un marco para involucrar al público en la planificación de Áreas Marinas Protegidas en E.U.A. Conservation Biology, 19(5):1392–1401, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daly:2009:TAL

[Dal09] Herman Daly. Three anathemas on limiting economic growth. Conservation Biology, 23(2):252–253, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delgado-Acevedo:2008:CHL

[DAR08] Johanna Delgado-Acevedo and Carla Restrepo. The contribution of habitat loss to changes in body size, allometry, and bilateral asymmetry in two Eleutherodactylus frogs from Puerto Rico. La Contribución de la pérdida de hábitat a cambios en el tamaño corporal, alometría y asimetría bilateral en dos ranas Eleutherodactylus de Puerto Rico. Conservation Biology, 22(3): 773–782, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davies:2002:BID

[Dav02] Glyn Davies. Bushmeat and international development. Conservation Biology, 16(3):587–589, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davis:2004:OU

[Dav04] Jeff N. Davis. Owl update. Conservation Biology, 18(6):1696– 1697, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davis:2005:SSM

[Dav05] Gary E. Davis. Science and society: Marine reserve design for the California Channel Islands. Ciencia y sociedad: Diseño de reservas marinas para las Islas Channel de California. Conservation Biology, 19(6):1745–1751, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davoren:2007:EGN

[Dav07] Gail K. Davoren. Effects of gill-net fishing on marine birds in a biological hotspot in the Northwest Atlantic. Efectos de las redes agalleras sobre aves marinas en un sitio biológicamente importante en el Noroeste del Atlántico. Conservation Biology, 21(4):1032–1045, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Danielsen:2009:BPF

[DBB^{+}09a] Finn Danielsen, Hendrien Beukema, Neil D. Burgess, Faizal Parish, Carsten A. Brühl, Paul F. Donald, Daniel Murdiyarso, Ben Phalan, Lucas Reijnders, Matthew Struebig, and Emily B. Fitzherbert. Biofuel plantations on forested lands: Double jeopardy for biodiversity and climate plantaciones de biocombustible en terrenos boscosos: Doble peligro para la biodiversidad y el clima. Conservation Biology, 23(2):348–358, April 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Danielsen:2009:LPN

[DBB^{+}09b] Finn Danielsen, Neil D. Burgess, Andrew Balmford, Paul F. Donald, Mikkel Funder, Julia P. G. Jones, Philip Alviola, Danilo S. Balete, Tom Blomley, Justin Brashares, Brian Child, Martin Enghoff, Jon Fjeldså, Sune Holt, Hanne Hübertz, Arne E. Jensen, Per M. Jensen, John Massao, Marlynn M. Mendoza, Yonika Ngaga, Michael K. Poulsen, Ricardo Rueda, Moses Sam, Thomas Skielboe, Greg Stuart-Hill, Elmer Topp-JØRgensen, and Deki Yonten. Local participation in natural resource monitoring: a characterization of approaches. Participación local en el monitoreo de recursos naturales: una caracterización de métodos. Conservation Biology, 23(1):31–42, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dauber:2006:EEH

[DBL06] Jens Dauber, Jan Bengtsson, and Lisette Lenoir. Evaluating effects of habitat loss and land-use continuity on ant species richness in seminatural grassland remnants. Evaluación de los efectos de la pérdida de hábitat y de continuidad en el uso de suelo sobre la riqueza de especies de hormigas en remanentes de pastizales seminaturales. Conservation Biology, 33(5):1150– 1160, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Durant:2007:RLT

[DBML07] Sarah M. Durant, Sultana Bashir, Thomas Maddox, and M. Karen Laurenson. Relating long-term studies to conservation practice: the case of the Serengeti Cheetah Project. Relacionando estudios de largo plazo con la conservación en la Práctica: El caso del Proyecto Guepardo del Serengeti Durant et al. Conservation Biology, 21(3):602–611, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dietz:2005:CDC

[DC05] Robert W. Dietz and Brian Czech. Conservation deficits for the continental United States: an ecosystem gap analysis. Déficit de conservación para los Estados Unidos: Un análisis de congruencias de ecosistemas. Conservation Biology, 19(5):1478–1487, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DeGrammont:2006:ETS

[DC06a] Paloma C. De Grammont and Alfredo D. Cuarón. An evaluation of threatened species categorization systems used on the American continent. Conservation Biology, 20(1):14–27, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DeMerode:2006:SPC

[DC06b] Emmanuel De Merode and Guy Cowlishaw. Species protection, the changing informal economy, and the politics of access to the bushmeat trade in the Democratic Republic of Congo. Protección de especies, la economía informal cambiante y la política de acceso al comercio de carne silvestre en la República Democrática de Congo. Conservation Biology, 33(5):1262–1271, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Derraik:2002:AMV

[DCD^{+}02] José G. B. Derraik, Gerard P. Closs, Katharine J. M. Dickinson, Philip Sirvid, Barbara I. P. Barratt, and Brian H. Patrick. Arthropod morphospecies versus taxonomic species: a case study with Araneae, Coleoptera, and Lepidoptera. Morfoespecies de artrópodos versus especies taxonómicas: Un estudio de caso con aranae, Coleóptera y Lepidóptera. Conservation Biology, 16(4):1015–1023, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daily:2003:CBN

[DCP^{+}03] Gretchen C. Daily, Gerardo Ceballos, Jesús Pacheco, Gerardo Suzán, and Arturo Sánchez-Azofeifa. Countryside biogeography of neotropical mammals: Conservation opportunities in agricultural landscapes of Costa Rica. Biogeografía del campo de mamíferos neotropicales: Oportunidades de conservación en paisajes agrícolas de Costa Rica. Conservation Biology, 17(6): 1814–1826, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daly:2007:WCT

[DCT^{+}07] Herman E. Daly, Brian Czech, David L. Trauger, William E. Rees, Mansi Grover, Tracy Dobson, and Stephen C. Trombulak. Are we consuming too much — for what? Conservation Biology, 21(5):1359–1362, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DiMauro:2001:AIP

[DD01] Desiree Di Mauro and Thomas Dietz. Assessment of an interpretive program for U.S. Coast guard personnel evaluación de un programa interpretativo para el personal de la guardia costera de U.S. Conservation Biology, 15(3):776–779, June 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diffendorfer:2004:LCC

[DD04a] James E. Diffendorfer and Paul F. Doherty, Jr. Lifting Cassandra’s curse. Conservation Biology, 18(3):600, June 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Driscoll:2004:LCM

[DD04b] Melanie J. L. Driscoll and Therese M. Donovan. Landscape context moderates edge effects: Nesting success of wood thrushes in Central New York. El contexto del paisaje modera los efectos de borde: Éxito de anidación de Hylocichla mustelina en New York Central. Conservation Biology, 18(5):1330–1338, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daehler:2004:RAS

[DDAK04] Curtis C. Daehler, Julie S. Denslow, Shahin Ansari, and Huang-Chi Kuo. A risk-assessment system for screening out invasive pest plants from Hawaii and other Pacific islands. Un sistema de evaluación de riesgo para seleccionar plantas invasoras nocivas de Hawai’i y otras islas del Pacífico. Conservation Biology, 18 (2):360–368, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dominguez-Dominguez:2007:HID

[DDBA^{+}07] Omar Domínguez-Domínguez, Luis Boto, Fernando Alda, Gerardo Pérez-Ponce De León, and Ignacio Doadrio. Human impacts on drainages of the Mesa Central, Mexico, and its genetic effects on an endangered fish, Zoogoneticus quitzeoensis. Impactos humanos en cuencas hidrológicas de la Meseta Central, México, y sus efectos genéticos sobre una especie de pez en peligro de extinción, Zoogoneticus quitzeoensis. Conservation Biology, 21(1):168–180, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davis:2000:HCW

[DDC^{+}00] Margaret Davis, Christine Douglas, Randy Calcote, Kenneth L. Cole, Marjorie Green Winkler, and Robyn Flakne. Holocene climate in the Western Great Lakes National Parks and Lakeshores: Implications for future climate change. Clima en el Holoceno en los Parques Nacionales de los Grandes Lagos del Oeste y en Costas de Lagos: Implicaciones para cambios climáticos futuros. Conservation Biology, 14(4):968–983, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Denoel:2005:EWF

[DDK05] Mathieu Denoel, Georg Dzukic, and Milos L. Kalezic. Effects of widespread fish introductions on paedomorphic newts in Europe. Efectos de introducciones de peces sobre tritones Pedomórficos en Europa. Conservation Biology, 19(1):162–170, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dominguez-Dominguez:2006:UEN

[DDMMZD06] Omar Domínguez-Domínguez, Enrique Martínez-Meyer, Luis Zambrano, and Gerardo Pérez-Ponce De León. Using ecologicalniche modeling as a conservation tool for freshwater species: Live-bearing fishes in Central Mexico. Utilización de modelos de nicho ecológico como herramienta de conservación de especies dulceacuícolas: Peces Vivíparos en el Centro de México. Conservation Biology, 20(6):1730–1739, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dustan:2001:LTM

[DDN01] Phillip Dustan, Eric Dobson, and George Nelson. Landsat Thematic Mapper: Detection of shifts in community composition of coral reefs. Mapeador Temático Landsat: Detección de desplazamientos en la composición de comunidades en arrecifes de coral. Conservation Biology, 15(4):892–902, August 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DeNardo:2007:TIA

[DeN07] Dale DeNardo. Tribute to icons of the American southwest. Conservation Biology, 21(2):564–565, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DeStefano:2003:BHL

[DeS03] Stephen DeStefano. Birds in the human landscape. Conservation Biology, 17(5):1467–1468, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Desalle:2006:SDV

[Des06] Rob Desalle. Species discovery versus species identification in DNA barcoding efforts: Response to Rubinoff. Conservation Biology, 20(5):1545–1547, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diaz-Frances:2005:SEM

[DFS05] Eloísa Díaz-Francés and Jorge Soberón. Statistical estimation and model selection of species-accumulation functions. Estimación Estadística y selección de modelos de funciones de acumulación de especies. Conservation Biology, 19(2):569–573, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dickens:2005:RPT

[DGC05] Sara Jo M. Dickens, Fritz Gerhardt, and Sharon K. Collinge. Recreational portage trails as corridors facilitating non-native plant invasions of the boundary waters canoe area wilderness (U.S.A.) senderos recreativos como corredores que facilitan invasiones de plantas no nativas en el boundary waters canoe area wilderness (E.U.A.). Conservation Biology, 19(5):1653–1657, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dimmick:2001:ESU

[DGG^{+}01] Walter Wheaton Dimmick, Michael J. Ghedotti, Michael J. Grose, Anne M. Maglia, Daniel J. Meinhardt, and David S. Pennock. The evolutionarily significant unit and adaptive criteria: a response to Young. Conservation Biology, 15(3):788–790, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dudley:2002:EWC

[DGP^{+}02] Joseph P. Dudley, Joshua R. Ginsberg, Andrew J. Plumptre, John A. Hart, and Liliana C. Campos. Effects of war and civil strife on wildlife and wildlife habitats. Efectos de la guerra y conflictos civiles sobre la vida silvestre y sus hábitats. Conservation Biology, 16(2):319–329, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dirnbock:2003:PFT

[DGSS03] Thomas Dirnböck, Josef Greimler, Patricio Lopez S., and Tod F. Stuessy. Predicting future threats to the native vegetation of Robinson Crusoe Island, Juan Fernandez Archipelago, Chile. Predicción de futuras amenazas a la vegetación nativa de la Isla Robinson Crusoe, Archipiélago Juan Fernández, Chile. Conservation Biology, 17(6):1650–1659, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:2006:PPD

[DGSS06] Robert R. Dunn, Michael C. Gavin, Monica C. Sanchez, and Jennifer N. Solomon. The pigeon paradox: Dependence of global conservation on urban nature. Conservation Biology, 21(2): 1814–1816, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dahlgaard:2000:SRE

[DH00a] Jesper Dahlgaard and Ary A. Hoffmann. Stress resistance and environmental dependency of inbreeding depression in Drosophila melanogaster. Resistencia al estrés y dependencia ambiental de la depresión por endogamia en Drosophila melanogaster. Conservation Biology, 14(4):1187–1192, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Debinski:2000:SOH

[DH00b] Diane M. Debinski and Robert D. Holt. A survey and overview of habitat fragmentation experiments. Sondeo y revisión de experimentos de fragmentación de hábitat. Conservation Biology, 14(2):342–355, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:2001:CVR

[DHW01] Erica H. Dunn, David J. T. Hussell, and Daniel A. Welsh. Conservation value and rankings of exotic species. Conservation Biology, 15(4):818, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dudley:2009:LBP

[DHZM09] Nigel Dudley, Liza Higgins-Zogib, and Stephanie Mansourian. The links between protected areas, faiths, and sacred natural sites. Enlaces entre Áreas protegidas, creencias y sitios naturales sagrados. Conservation Biology, 23(3):568–577, June 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dinerstein:2003:CT

[Din03a] Eric Dinerstein. Counting tigers. Conservation Biology, 17(6): 1852–1853, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dinerstein:2003:WC

[Din03b] Eric Dinerstein. I ♡ wild cats. Conservation Biology, 17(6): 1853–1855, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dixon:2007:MBP

[Dix07] Jeremy Dixon. Making broken pieces whole again. Conservation Biology, 21(1):283–284, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Devictor:2007:FHE

[DJC^{+}07] Vincent Devictor, Romain Julliard, Denis Couvet, Alexandre Lee, and Frédéric Jiguet. Functional homogenization effect of urbanization on bird communities. Efecto de la Homogeneización funcional de la urbanización sobre comunidades de aves. Conservation Biology, 21(3):741–751, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drechsler:2007:DCE

[DJOW07] Martin Drechsler, Karin Johst, Cornelia Ohl, and Frank Wätzold. Designing cost-effective payments for conservation measures to generate spatiotemporal habitat heterogeneity. Diseño de pagos rentables a medidas de conservación para generar heterogeneidad espaciotemporal de hábitat. Conservation Biology, 21(6):1475–1486, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Deguise:2006:PAP

[DK06] Isabelle E. Deguise and Jeremy T. Kerr. Protected areas and prospects for endangered species conservation in Canada. Áreas protegidas y perspectivas para la conservación de especies en peligro en Canadá. Conservation Biology, 20(1):48–55, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffy:2009:INR

[DKS^{+}09] Karl J. Duffy, Naomi E. Kingston, Brendan A. Sayers, David L. Roberts, and Jane C. Stout. Inferring national and regional declines of rare orchid species with probabilistic models. Inferencia de declinaciones nacionales y regionales de especies de orquídeas raras mediante modelos Probabilísticos. Conservation Biology, 23(1):184–195, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Deem:2001:PTP

[DKW01] Sharon L. Deem, William B. Karesh, and Wendy Weisman. Putting theory into practice: Wildlife health in conservation. De la teoría a la Práctica: la salud de la vida silvestre en la conservación. Conservation Biology, 15(5):1224–1233, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooke:2008:RMT

[dLBBG^{+}08] M. de L. Brooke, S. H. M. Butchart, S. T. Garnett, G. M. Crowley, N. B. Mantilla-Beniers, and A. J. Stattersfield. Rates of movement of threatened bird species between IUCN red list categories and toward extinction. Tasas de movimiento de especies de aves amenazadas entre categorías de la lista roja y hacia la extinción. Conservation Biology, 22(2):417–427, April 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dagrosa:2000:VBM

[DLCV00] Caterina D’agrosa, Cleridy E. Lennert-Cody, and Omar Vidal. Vaquita bycatch in Mexico’s artisanal gillnet fisheries: Driving a small population to extinction. Captura accidental de la vaquita marina en redes agalleras de pesquerías artesanales de México: Llevando una población pequeña hacia la extinción. Conservation Biology, 14(4):1110–1119, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donlan:2004:REH

[DM04a] C. Josh Donlan and Paul S. Martin. Role of ecological history in invasive species management and conservation. Conservation Biology, 18(1):267–269, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donnelly:2004:IRS

[DM04b] Roarke Donnelly and John M. Marzluff. Importance of reserve size and landscape context to urban bird conservation. Importancia del tamaño de la reserva y el contexto del paisaje para la conservación de aves urbanas. Conservation Biology, 18(3): 733–745, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delibes-Mateos:2008:KRE

[DMDFV08] Miguel Delibes-Mateos, Miguel Delibes, Pablo Ferreras, and Rafael Villafuerte. Key role of European rabbits in the conservation of the Western Mediterranean Basin hotspot. Papel clave de conejos Europeos en la conservación del Área de importancia del Mediterráneo occidental. Conservation Biology, 22(5): 1106–1117, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dulvy:2000:FSL

[DMG^{+}00] Nicholas K. Dulvy, Julian D. Metcalfe, Jamie Glanville, Mike G. Pawson, and John D. Reynolds. Fishery stability, local extinctions, and shifts in community structure in skates. Estabilidad de la pesquería, extinciones locales y cambios en la estructura de comunidades de rayas. Conservation Biology, 14(1):283–293, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dethier:2003:CCH

[DMS03] Megan N. Dethier, Kathryn McDonald, and Richard R. Strathmann. Colonization and connectivity of habitat patches for coastal marine species distant from source populations. Conectividad de fragmentos de hábitat y su colonización por especies marinas costeras distantes de poblaciones fuente. Conservation Biology, 17(4):1024–1035, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dale:2000:EEU

[DMSP00] Svein Dale, Kjetil Mork, Rune Solvang, and Andrew J. Plumptre. Edge effects on the understory bird community in a logged forest in Uganda. Efectos de bordes en comunidades de aves del sotobosque en un bosque talado de Uganda. Conservation Biology, 14(1):265–276, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drake:2000:ERE

[DN00] Deanne C. Drake and Robert J. Naiman. An evaluation of restoration efforts in fishless lakes stocked with exotic trout. Evaluación de los esfuerzos de restauración en lagos sin peces repoblados con truchas exóticas. Conservation Biology, 14(6): 1807–1820, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davradou:2001:SEA

[DN01] Maria Davradou and Gene Namkoong. Science, ethical arguments, and management in the preservation of land for grizzly bear conservation. Ciencia, argumentos Éticos y manejo en la preservación de tierras para la conservación del oso gris. Conservation Biology, 15(3):570–577, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donaldson:2002:EHF

[DNZK02] John Donaldson, Ingrid Nänni, Costas Zachariades, and Jessica Kemper. Effects of habitat fragmentation on pollinator diversity and plant reproductive success in renosterveld shrublands of South Africa. Efectos de la fragmentación del hábitat sobre la diversidad de polinizadores y el Éxito reproductivo de plantas en matorrales renosterveld (Suráfrica). Conservation Biology, 16(5):1267–1276, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doak:2001:LAS

[Doa01] Daniel F. Doak. The long-awaited sourcebook arrives. Conservation Biology, 15(6):1825–1826, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davis:2008:EAA

[DOGN08] Naomi E. Davis, Dennis J. O’Dowd, Peter T. Green, and Ralph Mac Nally. Effects of an alien ant invasion on abundance, behavior, and reproductive success of endemic island birds. Efectos de la invasión de hormigas exóticas sobre la abundancia y el Éxito reproductivo de aves insulares endémicas. Conservation Biology, 22(5):1165–1176, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donald:2004:BIS

[Don04] Paul F. Donald. Biodiversity impacts of some agricultural commodity production systems. Conservation Biology, 18(1):17–38, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donner:2006:DWI

[Don06] Deahn Donner. Determining What’s important about landscapes. Conservation Biology, 20(2):592–594, April 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doremus:2004:WPW

[Dor04] Holly Doremus. What people want from wolves. Conservation Biology, 18(4):1163–1164, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doremus:2006:IUW

[Dor06] Holly Doremus. Inspiration in unexpected ways. Conservation Biology, 21(2):1830–1831, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dixon:2006:ERC

[DOW^{+}06] Jeremy D. Dixon, Madan K. Oli, Michael C. Wooten, Thomas H. Eason, J. Walter Mccown, and David Paetkau. Effectiveness of a regional corridor in connecting two Florida black bear populations. Efectividad de un corredor regional para la conectividad de dos poblaciones de oso negro de Florida. Conservation Biology, 20(1):155–162, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doremus:2001:WLM

[DP01] Holly Doremus and Joel E. Pagel. Why listing may be forever: Perspectives on delisting under the U.S. Endangered Species Act. Porqué los listados Podrán Ser para siempre: Perspectivas sobre eliminar especies de las listas bajo el Acta de Especies en Peligro de los Estados Unidos. Conservation Biology, 15(5): 1258–1268, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doremus:2003:L

[DP03] Holly Doremus and Joel E. Pagel. Letters. Conservation Biology, 17(3):652–653, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doukakis:2009:GTR

[DPB^{+}09] Phaedra Doukakis, E. C. M. Parsons, William C. G. Burns, Anne K. Salomon, Ellen Hines, and John A. Cigliano. Gaining traction: Retreading the wheels of marine conservation. Logrando tracción: Reencauchando las ruedas de la conservación marina. Conservation Biology, 23(4):841–846, August 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diaz:2005:RIL

[DPTT^{+}05] José A. Díaz, Javier Pérez-Tris, José L. Tellería, Roberto Carbonell, and Tomás Santos. Reproductive investment of a lacertid lizard in fragmented habitat. Inversión reproductiva de una lagartija lacértida en hábitat fragmentado. Conservation Biology, 19(5):1578–1585, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dulvy:2002:PEV

[DR02] Nicholas K. Dulvy and John D. Reynolds. Predicting extinction vulnerability in skates. Predicción de la vulnerabilidad de extinción en rayas. Conservation Biology, 16(2):440–450, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dawe:2003:FTL

[DR03] Neil K. Dawe and Kenneth L. Ryan. The faulty three-legged-stool model of sustainable development. Conservation Biology, 17(5):1458–1460, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drew:2005:UTE

[Dre05] Joshua A. Drew. Use of traditional ecological knowledge in marine conservation. Uso del conocimiento ecológico tradicional en la conservación marina. Conservation Biology, 19(4):1286–1293, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Develey:2001:ERM

[DS01] Pedro F. Develey and Philip C. Stouffer. Effects of roads on movements by understory birds in mixed-species flocks in Central Amazonian Brazil. Efectos de los caminos sobre los movimientos de aves de sotobosque en parvadas mixtas en la Amazonía Central Brasil. Conservation Biology, 15(5):1416– 1422, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Decaro:2008:SPP

[DS08] Daniel Decaro and Michael Stokes. Social-psychological principles of community-based conservation and conservancy motivation: Attaining goals within an autonomy-supportive environment principios. Psicológicos sociales de la conservación basada en comunidades y la Motivación por la conservación: Alcanzando metas en un ambiente de autonomía y solidaridad. Conservation Biology, 22(6):1443–1451, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Deboer:2008:PLG

[DSE^{+}08] Timery S. Deboer, Matthew D. Subia, Mark V. Erdmann, Katie Kovitvongsa, and Paul H. Barber. Phylogeography and limited genetic connectivity in the endangered boring giant clam across the coral triangle. Filogeografía y conectividad genética limitada en la almeja gigante Tridacna crocea en el Triángulo de coral. Conservation Biology, 22(5):1255–1266, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davidson:2002:STP

[DSJ02] Carlos Davidson, H. Bradley Shaffer, and Mark R. Jennings. Spatial tests of the pesticide drift, habitat destruction, UV-B, and climate-change hypotheses for California amphibian declines. Pruebas espaciales de la deriva de pesticidas, destrucción de hábitat, UV-B e hipótesis de cambio climático para la declinación de anfibios de California. Conservation Biology, 16(6): 1588–1601, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dehnen-Schmutz:2007:HTO

[DSTPW07] Katharina Dehnen-Schmutz, Julia Touza, Charles Perrings, and Mark Williamson. The horticultural trade and ornamental plant invasions in Britain. El comercio horticultural y la invasión de plantas ornamentales en Gran Bretaña. Conservation Biology, 21(1):224–231, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donlan:2003:RRN

[DTCC03] C. Josh Donlan, Bernie R. Tershy, Karl Campbell, and Felipe Cruz. Research for requiems: the need for more collaborative action in eradication of invasive species. Conservation Biology, 17 (6):1850–1851, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DuPlessis:2001:ANG

[Du 01] Morné A. Du Plessis. Academia as a nursery ground for conservation biology. Conservation Biology, 15(6):1477, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dudley:2000:ECS

[Dud00] Joseph P. Dudley. Environment and civil strife. Conservation Biology, 14(4):1215–1216, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dudley:2002:IPM

[Dud02] Joseph P. Dudley. Issues and priorities for mammal conservation. Conservation Biology, 16(4):1169–1171, August 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:2002:UDB

[Dun02] Erica H. Dunn. Using decline in bird populations to identify needs for conservation action. Utilización de la declinación de poblaciones de aves para identificar necesidades de acciones de conservación. Conservation Biology, 16(6):1632–1637, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:2003:R

[Dun03] Erica Dunn. Response. Conservation Biology, 17(5):1201–1202, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:2004:RFC

[Dun04] Robert R. Dunn. Recovery of faunal communities during tropical forest regeneration. Recuperación de comunidades de fauna durante la regeneración del bosque tropical. Conservation Biology, 18(2):302–309, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:2005:MIE

[Dun05] Robert R. Dunn. Modern insect extinctions, the neglected majority. Extinciones modernas de insectos, la mayoría desatendida. Conservation Biology, 19(4):1030–1036, August 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delaney:2005:AUC

[DW05a] Kathleen S. Delaney and Robert K. Wayne. Adaptive units for conservation: Population distinction and historic extinctions in the island scrub-jay. Unidades adaptativas para conservación: Diferenciación de la población y extinciones históricas de Aphelocoma insularis. Conservation Biology, 19(2):523–533, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Driscoll:2005:BRH

[DW05b] Don A. Driscoll and Tom Weir. Beetle responses to habitat fragmentation depend on ecological traits, habitat condition, and remnant size. Las respuestas de escarbajos a la fragmentación de hábitat dependen de los atributos ecológicos, condición del hábitat y tamaño remanente. Conservation Biology, 19(1):182– 194, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DellaSala:2006:SSN

[DW06] Dominick A. DellaSala and Jack E. Williams. Special Section: The Northwest Forest Plan: a Global Model of Forest Management in Contentious Times. Conservation Biology, 20(2):274– 276, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duguay:2001:SAA

[DWN01] Jeffrey P. Duguay, Petra Bohall Wood, and Jeffrey V. Nichols. Songbird abundance and avian nest survival rates in forests fragmented by different silvicultural treatments abundancia de aves canoras y tasas de supervivencia de nidos de aves en bosques fragmentados por diferentes tratamientos silviculturales. Conservation Biology, 15(5):1405–1415, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dombeck:2004:WPP

[DWW04] Michael P. Dombeck, Jack E. Williams, and Christopher A. Wood. Wildfire policy and public lands: Integrating scientific understanding with social concerns across landscapes. Políticas para fuegos no controlados y tierras públicas: Integrando el conocimiento científico con los intereses sociales en paisajes. Conservation Biology, 18(4):883–889, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dellasala:2004:BSM

[DWWF04] Dominick A. Dellasala, Jack E. Williams, Cindy Deacon Williams, and Jerry F. Franklin. Beyond smoke and mirrors: a synthesis of fire policy and science. Más allá de humo y espejos: una síntesis de políticas y ciencia del fuego. Conservation Biology, 18(4):976–986, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DElia:2008:CLV

[DZM08] Jesse D’Elia, Michele Zwartjes, and Scott Mccarthy. Considering legal viability and societal values when deciding what to conserve under the U.S. Endangered Species Act. Conservation Biology, 22(4):1072–1074, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [WABT08].

Edwards:2009:VRL

[EAA^{+}09] David P. Edwards, Felicity A. Ansell, Abdul H. Ahmad, Reuben Nilus, and Keith C. Hamer. The value of rehabilitating logged rainforest for birds. El valor de la rehabilitación de un bosque tropical lluvioso para aves. Conservation Biology, 23(6):1628– 1633, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Evans:2005:NNP

[EAR^{+}05] Celia Evans, Eleanor Abrams, Robert Reitsma, Karin Roux, Laura Salmonsen, and Peter P. Marra. The neighborhood nestwatch program: Participant outcomes of a citizen-science ecological research project. Conservation Biology, 19(3):589–594, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eschtruth:2009:AEP

[EB09] Anne K. Eschtruth and John J. Battles. Acceleration of exotic plant invasion in a forested ecosystem by a generalist herbivore. Aceleración de la invasión de plantas exóticas en un ecosistema boscoso por un herbívoro generalista. Conservation Biology, 23 (2):388–399, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

England:2001:MGA

[EBAW01] Phillip R. England, Fiona Beynon, David J. Ayre, and Robert J. Whelan. A molecular genetic assessment of mating-system variation in a naturally bird-pollinated shrub: Contributions from birds and introduced honeybees. Evaluación genética molecular de la variación de sistemas de apareamiento en un arbusto naturalmente polinizado por aves: Contribuciones de aves y de abejas introducidas. Conservation Biology, 15(6):1645–1655, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Engelman:2006:IPF

[EBDM06] Robert Engelman, Jason Bremner, Roger-Mark De Souza, and Kathleen Mogelgaard. Indigenous population, fertility, and reproductive intention in the lowland neotropics: Response to Mc-Sweeney. Conservation Biology, 33(5):1315–1317, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ebersole:2004:DYB

[Ebe04] John P. Ebersole. So, do you believe me now? Conservation Biology, 18(1):280–281, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elgar:2001:IEI

[EC01] Mark A. Elgar and Danielle Clode. Inbreeding and extinction in island populations: a cautionary note. Conservation Biology, 15 (1):284–286, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [Fra01].

Ewers:2007:EFS

[ED07] Robert M. Ewers and Raphael K. Didham. The effect of fragment shape and species’ sensitivity to habitat edges on animal population size. El efecto de la forma del fragmento y la sensibilidad de las especies a los bordes de hábitat sobre el tamaño poblacional de animales Ewers & Didham. Conservation Biology, 21(4):926–936, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

EOCB:2003:FRO

[Edi03] The Editorial Office of Conservation Biology. Feedback requested — online survey for Conservation Biology. Conservation Biology, 17(4):1195, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edwards:2004:MLF

[Edw04] Thomas C. Edwards, Jr. A monumental leap forward in speciesoccurrence modeling. Conservation Biology, 18(3):855–857, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ellner:2002:PPV

[EFLW02] Stephen P. Ellner, John Fieberg, Donald Ludwig, and Chris Wilcox. Precision of population viability analysis. Conservation Biology, 16(1):258–261, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrlich:2007:CCE

[EG07] Paul R. Ehrlich and Lawrence H. Goulder. Is current consumption excessive? A general framework and some indications for the United States `?Es excesivo el consumo actual? Un marco de referencia general y algunas indicaciones para E.U.A. Conservation Biology, 21(5):1145–1154, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eaton:2005:RIR

[EGN^{+}05] M. A. Eaton, R. D. Gregory, D. G. Noble, J. A. Robinson, J. Hughes, D. Procter, A. F. Brown, and D. W. Gibbons. Regional IUCN red listing: the process as applied to birds in the United Kingdom. Listado rojo IUCN regional: el proceso aplicado a aves en el Reino Unido. Conservation Biology, 19(5): 1557–1570, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Endress:2004:HPC

[EGPS04] Bryan A. Endress, David L. Gorchov, Maren B. Peterson, and Eduardo Padrón Serrano. Harvest of the palm Chamaedorea radicalis, its effects on leaf production, and implications for sustainable management. Efectos de la cosecha de hojas de palma (Chamaedorea radicalis) sobre la producción de hojas e implicaciones para un manejo sustentable. Conservation Biology, 18(3): 822–830, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edgerton:2004:UCD

[EHJ^{+}04] Brett F. Edgerton, Paula Henttonen, Japo Jussila, Ari Mannonen, Pietari Paasonen, Trond TaugbØL, Lennart Edsman, and Catherine Souty-Grosset. Understanding the causes of disease in European freshwater crayfish entendiendo las causas de enfermedad en cangrejos europeos de agua dulce. Conservation Biology, 18(6):1466–1474, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:2000:WPC

[Ehr00] David Ehrenfeld. War and peace and conservation biology. La guerra y la paz y la biología de la conservación. Conservation Biology, 14(1):105–112, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:2001:POC

[Ehr01] David Ehrenfeld. A postscript to Orr’s commandments. Conservation Biology, 15(4):825–826, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrlich:2003:GTW

[Ehr03] Paul R. Ehrlich. Get off the train and walk. Conservation Biology, 17(2):352–353, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:2005:ELG

[Ehr05a] David Ehrenfeld. The environmental limits to globalization. Los límites ambientales de la globalización. Conservation Biology, 19(2):318–326, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:2005:SLI

[Ehr05b] David Ehrenfeld. Sustainability: Living with the imperfections. Conservation Biology, 19(1):33–35, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:2006:TVC

[Ehr06] David Ehrenfeld. Transgenics and vertebrate cloning as tools for species conservation. La transgenética y la clonación de vertebrados como herramientas para la conservación de especies. Conservation Biology, 20(3):723–732, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:2008:NC

[Ehr08] David Ehrenfeld. Neoliberalization of conservation. Conservation Biology, 22(5):1091–1092, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:2009:AH

[Ehr09] David Ehrenfeld. Authentic hope. Conservation Biology, 23(6): 1352–1353, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Evans:2008:OBI

[EK08] Kirsten E. Evans and Terrie Klinger. Obstacles to bottom-up implementation of marine ecosystem management. Obstáculos para la implementación de abajo hacia arriba del manejo de ecosistemas marinos. Conservation Biology, 22(5):1135–1143, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eisto:2000:PPO

[EKL^{+}00] Anna-Kaija Eisto, Markku Kuitunen, Antti Lammi, Veli Saari, Jukka Suhonen, Susanna Syrjäsuo, and Päivi M. Tikka. Population persistence and offspring fitness in the rare bellflower Campanula cervicaria in relation to population size and habitat quality. Persistencia poblacional y adaptabilidad de crías en la especie rara Campanula cervicaria con relación al tamaño poblacional y la calidad del hábitat. Conservation Biology, 14(5): 1413–1421, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elphick:2000:FEB

[Elp00] Chris S. Elphick. Functional equivalency between rice fields and seminatural wetland habitats. Equivalencia funcional entre hábitats de arrozales y humedales seminaturales. Conservation Biology, 14(1):181–191, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ellsworth:2003:PEP

[EM03] Joshua W. Ellsworth and Brenda C. Mccomb. Potential effects of passenger pigeon flocks on the structure and composition of presettlement forests of Eastern North America. Efectos potenciales de parvadas de palomas migratorias sobre la estructura y composición de bosques en Norte América Oriental antes de la colonización. Conservation Biology, 17(6):1548–1558, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elmendorf:2008:UCC

[EM08] Sarah C. Elmendorf and Kara A. Moore. Use of communitycomposition data to predict the fecundity and abundance of species. El uso de datos de composición de la comunidad para predecir la fecundidad y abundancia de especies. Conservation Biology, 22(6):1523–1532, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Epps:2004:ECC

[EMW^{+}04] Clinton W. Epps, Dale R. Mccullough, John D. Wehausen, Vernon C. Bleich, and Jennifer L. Rechel. Effects of climate change on population persistence of desert-dwelling mountain sheep in California. Efectos del cambio climático sobre la persistencia de la población de borrego cimarrón en California. Conservation Biology, 18(1):102–113, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edghill:2007:MCC

[EN07] Janet Edghill and Peter Newman. Minting common currency in contemporary sustainability discourse. Conservation Biology, 21(2):563–564, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ewing:2008:ILG

[ENN^{+}08] Steven R. Ewing, Ruedi G. Nager, Malcolm A. C. Nicoll, Aurelien Aumjaud, Carl G. Jones, and Lukas F. Keller. Inbreeding and loss of genetic variation in a reintroduced population of Mauritius kestrel. Endogamia y pérdida de variación genética en una población reintroducida de Falco punctatus. Conservation Biology, 22(2):395–404, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elphick:2001:CPR

[ERB01] Chris S. Elphick, J. Michael Reed, and J. Marcelo Bonta. Correlates of population recovery goals in endangered birds. Correlaciones de metas de recuperación en aves en peligro. Conservation Biology, 15(5):1285–1291, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edgar:2004:CDM

[ES04] G. J. Edgar and C. R. Samson. Catastrophic decline in mollusc diversity in Eastern Tasmania and its concurrence with shellfish fisheries. Declinación catastrófica de la diversidad de moluscos en Tasmania Oriental y su concurrencia con pesquerías de mariscos. Conservation Biology, 18(6):1579–1588, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Evans:2008:IPG

[ES08] Simon R. Evans and Ben C. Sheldon. Interspecific patterns of genetic diversity in birds: Correlations with extinction risk. Patrones interespecíficos de diversidad genética en aves: Correlaciones con el riesgo de extinción. Conservation Biology, 22(4): 1016–1025, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edgar:2005:SEM

[ESB05] Graham J. Edgar, Cath R. Samson, and Neville S. Barrett. Species extinction in the marine environment: Tasmania as a regional example of overlooked losses in biodiversity. Extinción de especies en el medio marino: Tasmania como ejemplo regional de pérdidas de biodiversidad ignoradas. Conservation Biology, 19(4):1294–1300, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Esler:2000:AMT

[Esl00] Daniel Esler. Applying metapopulation theory to conservation of migratory birds. Aplicación de la teoría de metapoblación en la conservación de aves migratorias. Conservation Biology, 14 (2):366–372, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Escamilla:2000:HMW

[ESSGL00] Alfredo Escamilla, Mauro Sanvicente, Miguel Sosa, and Carlos Galindo-Leal. Habitat mosaic, wildlife availability, and hunting in the tropical forest of Calakmul, Mexico. El mosaico de hábitats, la disponibilidad de presas y la cacería en la selva tropical de Calakmul, México. Conservation Biology, 14(6): 1592–1601, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Estes:2004:FT

[Est04] J. A. Estes. A fitting tribute. Conservation Biology, 18(1):279– 280, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edmands:2003:MFA

[ET03] Suzanne Edmands and Charles C. Timmerman. Modeling factors affecting the severity of outbreeding depression. Factores de modelaje que afectan la severidad de la depresión exogámica. Conservation Biology, 17(3):883–892, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrlen:2001:REA

[EVD01] Johan Ehrlén, Jan Van Groenendael, and Hans De Kroon. Reliability of elasticity analysis: Reply to Mills et al. Conservation Biology, 15(1):278–280, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ewing:2009:DEI

[Ewi09] Kern Ewing. Disappearing into the environmental–industrial complex. Conservation Biology, 23(3):777–778, June 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:2009:UCE

[EYZ09] David Ehrenfeld, Shen Yu, and Yongguan Zhu. Urban conservation and environmental protection in China: a major effort by the Chinese Academy of Sciences. Conservation Biology, 23(3): 546–547, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Faaborg:2004:TAN

[Faa04] John Faaborg. Truly artificial nest studies. Conservation Biology, 18(2):369–370, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Faith:2002:QBP

[Fai02] Daniel P. Faith. Quantifying biodiversity: a phylogenetic perspective. Conservation Biology, 16(1):248–252, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Faith:2008:TSP

[Fai08] Daniel P. Faith. Threatened species and the potential loss of phylogenetic diversity: Conservation scenarios based on estimated extinction probabilities and phylogenetic risk analysis. Especies amenazadas y la pérdida potencial de diversidad filogenética: Escenarios de conservación derivados de probabilidades de extinción estimadas y de análisis de riesgo filogenético. Conservation Biology, 22(6):1461–1470, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fallon:2007:GDL

[Fal07] Sylvia M. Fallon. Genetic data and the listing of species under the U.S. Endangered Species Act. Datos genéticos y el enlistado de especies en el Acta de Especies en Peligro de E.U.A. Conservation Biology, 21(5):1186–1195, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farnsworth:2004:FRP

[Far04] Elizabeth J. Farnsworth. Forging research partnerships across the academic–agency divide. Conservation Biology, 18(2):291– 293, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farley:2008:RPC

[Far08] Joshua Farley. The role of prices in conserving critical natural capital. El papel de los precios en la conservación del capital natural crítico. Conservation Biology, 22(6):1399–1408, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fattorini:2006:DBH

[Fat06] Simone Fattorini. Detecting biodiversity hotspots by speciesarea relationships: a case study of Mediterranean beetles. Detección de sitios de importancia para la biodiversidad con relaciones especies-Área: Un estudio de caso de escarabajos del Mediterráneo. Conservation Biology, 33(5):1169–1180, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fauth:2001:WTP

[Fau01] Peter T. Fauth. Wood thrush populations are not all sinks in the agricultural Midwestern United States. Poblaciones de zorzalito manchado: no todas son vertedero en el medio-occidente agrícola de los Estados Unidos. Conservation Biology, 15(2):523–527, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Findlay:2000:RTW

[FB00] C. Scot Findlay and Josée Bourdages. Response time of wetland biodiversity to road construction on adjacent lands. Tiempo de respuesta de la biodiversidad de humedales a la construcción de caminos en tierras adyacentes. Conservation Biology, 14(1):86– 94, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fischer:2008:BAC

[FB08] A. Paige Fischer and John C. Bliss. Behavioral assumptions of conservation policy: Conserving oak habitat on family-forest land in the Willamette Valley, Oregon. Suposiciones conductuales de las políticas de conservación: la conservación de hábitat de rob le en tierras forestales familiares en el Valle Willamette, Oregon. Conservation Biology, 22(2):275–283, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fisher:2009:TCC

[FBB^{+}09] Brendan Fisher, Katharine Bolt, Richard B. Bradbury, Toby A. Gardner, Jonathan M. H. Green, David G. Hole, and Robin Naidoo. Two cultures of conservation. Conservation Biology, 23 (5):1069–1071, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fidler:2006:ICN

[FBC^{+}06] Fiona Fidler, Mark A. Burgman, Geoff Cumming, Robert Buttrose, and Neil Thomason. Impact of criticism of null-hypothesis significance testing on statistical reporting practices in conservation biology. Impacto de la crítica a las pruebas de significancia de hipótesis nula en las prácticas de reporte Estadístico en biología de la conservación. Conservation Biology, 20(5):1539– 1544, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freilich:2000:FAP

[FBCG00] Jerome E. Freilich, Kenneth P. Burnham, Christopher M. Collins, and C. Ann Garry. Factors affecting population assessments of desert tortoises. Factores que afectan la evaluación poblacional de tortugas del desierto. Conservation Biology, 14 (5):1479–1489, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2000:SU

[FBD^{+}00] Erica Fleishman, Carol L. Boggs, M. Claire Devine, Salit Kark, and Taylor H. Ricketts. Status of the union. Conservation Biology, 14(6):1926–1927, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flaspohler:2000:ACB

[FBK00] David J. Flaspohler, Brian R. Bub, and Beth A. Kaplin. Application of conservation biology research to management. Aplicación de la investigación sobre conservación biológica en el manejo. Conservation Biology, 14(6):1898–1902, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freemark:2002:IFH

[FBK02] Kathryn E. Freemark, Céline Boutin, and Cathy J. Keddy. Importance of farmland habitats for conservation of plant species.

Importancia de hábitats agropecuarios para la conservación de especies de plantas. Conservation Biology, 16(2):399–412, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Finn:2009:DSM

[FBL09a] Debra S. Finn, Michael T. Bogan, and David A. Lytle. Demographic stability metrics for conservation prioritization of isolated populations. Medidas de estabilidad demográfica para la priorización de la conservación de poblaciones aisladas. Conservation Biology, 23(5):1185–1194, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freedman:2009:MEA

[FBL^{+}09b] Adam H. Freedman, Wolfgang Buermann, Matthew Lebreton, Laurent Chirio, and Thomas B. Smith. Modeling the effects of anthropogenic habitat change on savanna snake invasions into African rainforest. Modelado de los efectos del cambio antropogénico de hábitat sobre las invasiones de serpientes de la sabana al bosque lluvioso africano. Conservation Biology, 23 (1):81–92, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fraser:2009:EVS

[FBP^{+}09] Sally E. M. Fraser, Alison E. Beresford, Jennifer Peters, John W. Redhead, Alastair J. Welch, Peter J. Mayhew, and Calvin Dytham. Effectiveness of vegetation surrogates for parasitoid wasps in reserve selection. Efectividad de loa sustitutos de vegetación para avispas parasitoides en la selección de reservas. Conservation Biology, 23(1):142–150, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fernandez:2005:MCC

[FC05] Miriam Fernández and Juan Carlos Castilla. Marine conservation in Chile: Historical perspective, lessons, and challenges. Conservación marina en Chile: Perspectiva histórica, lecciones y desafíos. Conservation Biology, 19(6):1752–1762, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fish:2005:PIS

[FCG^{+}05] Marianne R. Fish, Isabelle M. Côté, Jennifer A. Gill, Andrew P. Jones, Saskia Renshoff, and Andrew R. Watkinson. Predicting the impact of sea-level rise on Caribbean sea turtle nesting habitat. Predicción del impacto del incremento del nivel del mar sobre el hábitat de anidación de tortugas marinas del Caribe. Conservation Biology, 19(2):482–491, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foppen:2000:URS

[FCL00] Ruud P. B. Foppen, J. Paul Chardon, and Wendy Liefveld. Understanding the role of sink patches in source-sink metapopulations: Reed warbler in an agricultural landscape. Comprensión del papel de parches sumidero en metapoblaciones fuente-sumidero: el carricero común en un paisaje agrícola. Conservation Biology, 14(6):1881–1892, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fox:2006:PBI

[FCN^{+}06] Helen E. Fox, Caroline Christian, J. Cully Nordby, Oliver R. W. Pergams, Garry D. Peterson, and Christopher R. Pyke. Perceived barriers to integrating social science and conservation. Conservation Biology, 21(2):1817–1820, December 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forman:2000:ERE

[FD00] Richard T. T. Forman and Robert D. Deblinger. The ecological road-effect zone of a Massachusetts (U.S.A.) suburban highway. Zona ecológica de efecto carretero en una autopista suburbana de Massachusetts (U.S.A.). Conservation Biology, 14(1):36–46, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forsyth:2004:CSL

[FDBM04] David M. Forsyth, Richard P. Duncan, Mary Bomford, and Geoff Moore. Climatic suitability, life-history traits, introduction effort, and the establishment and spread of introduced mammals in Australia. Adaptabilidad climática, características de historia de vida, esfuerzo de introducción y el establecimiento y Extensión de mamíferos introducidos en Australia. Conservation Biology, 18(2):557–569, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fernandes:2005:ERN

[FDL^{+}05] Leanne Fernandes, Jon Day, Adam Lewis, Suzanne Slegers, Brigid Kerrigan, Dan Breen, Darren Cameron, Belinda Jago, James Hall, Dave Lowe, James Innes, John Tanzer, Virginia Chadwick, Leanne Thompson, Kerrie Gorman, Mark Simmons, Bryony Barnett, Kirsti Sampson, Glenn De’Ath, Bruce Mapstone, Helene Marsh, Hugh Possingham, Ian Ball, Trevor Ward, Kirstin Dobbs, James Aumend, Deb Slater, and Kate Stapleton. Establishing representative no-take areas in the Great Barrier Reef: Large-scale implementation of theory on marine protected areas. Establecimiento de Áreas sin captura representativas en la Gran Barrera Arrecifal: Implementación a gran escala de la teoría sobre Áreas marinas protegidas. Conservation Biology, 19 (6):1733–1744, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Falcy:2007:ECR

[FE07] Matthew R. Falcy and Cristián F. Estades. Effectiveness of corridors relative to enlargement of habitat patches. La efectividad de los corredores en relación con la ampliación de parches de hábitat. Conservation Biology, 21(5):1341–1346, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fearnside:2005:DBA

[Fea05] Philip M. Fearnside. Deforestation in Brazilian Amazonia: History, rates, and consequences. Deforestación en la Amazonía brasileña: Historia, tasas y consecuencias. Conservation Biology, 19(3):680–688, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Findlay:2009:SLU

[FEGB09] C. Scott Findlay, Stewart Elgie, Brian Giles, and Linda Burr. Species listing under Canada’s Species at Risk Act. Lista de especies en el contexto del Acta de Especies en Riesgo de Canadá. Conservation Biology, 23(6):1609–1617, December 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuhlendorf:2009:PHR

[FEKH09] Samuel D. Fuhlendorf, David M. Engle, Jay Kerby, and Robert Hamilton. Pyric herbivory: Rewilding landscapes through the recoupling of fire and grazing. Herbivoría pírica: Restablecimiento de paisajes silvestres mediante la combinación de fuego y pastoreo. Conservation Biology, 23(3):588–598, June 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferraro:2001:GHP

[Fer01] Paul J. Ferraro. Global habitat protection: Limitations of development interventions and a role for conservation performance payments. Protección global del hábitat: Limitantes para el desarrollo de intervenciones y el papel de los pagos por la Ejecución de actividades de conservación. Conservation Biology, 15 (4):990–1000, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forbes:2001:ADP

[FES01] Bruce C. Forbes, James J. Ebersole, and Beate Strandberg. Anthropogenic disturbance and patch dynamics in circumpolar Arctic ecosystems. Perturbación antropogénica y dinámica de parches en ecosistemas del Ártico circumpolar. Conservation Biology, 15(4):954–969, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fragoso:2008:MIW

[FF08] Gerardo Fragoso and Sarah Ferriss. Monitoring international wildlife trade with coded species data: Response to Gerson et al. Conservation Biology, 22(6):1648–1650, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Floyd:2003:EHL

[FFHW03] M. Lisa Floyd, Thomas L. Fleischner, David Hanna, and Paul Whitefield. Effects of historic livestock grazing on vegetation at Chaco Culture National Historic Park, New Mexico. Efectos del pastoreo histórico de ganado sobre la vegetación en el Parque Histórico Nacional Cultura Chaco, Nuevo México. Conservation Biology, 17(6):1703–1711, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fenster:2000:IOD

[FG00] Charles B. Fenster and Laura F. Galloway. Inbreeding and outbreeding depression in natural populations of Chamaecrista fasciculata (Fabaceae). Depresión por endogamia y exogamia en poblaciones naturales de Chamaecrista fasciculata (Fabaceae). Conservation Biology, 14(5):1406–1412, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Feer:2005:EFF

[FH05] François Feer and Yves Hingrat. Effects of forest fragmentation on a dung beetle community in French Guiana. Efectos de la fragmentación de bosques sobre una comunidad de escarabajos coprófagos en La Guyana Francesa. Conservation Biology, 19 (4):1103–1112, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fredrickson:2006:DHI

[FH06] Richard J. Fredrickson and Philip W. Hedrick. Dynamics of hybridization and introgression in red wolves and coyotes. Dinámica de la hibridación e introgresión en lobos rojos y coyotes. Conservation Biology, 33(5):1272–1283, August 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forsman:2009:ACM

[FH09] Anders Forsman and Mattias Hagman. Association of coloration mode with population declines and endangerment in Australian frogs. Asociación del modo de coloración con las declinaciones poblacionales y el peligro de extinción de ranas Australianas. Conservation Biology, 23(6):1535–1543, December 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ford:2008:ENS

[FHB^{+}08] Michael J. Ford, Jeffrey J. Hard, Brant Boelts, Eric Lahood, and Jason Miller. Estimates of natural selection in a salmon population in captive and natural environments. Estimaciones de selección natural en una población de salmón en ambientes cautivos y naturales. Conservation Biology, 22(3):783–794, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flagstad:2004:CHN

[FHL^{+}04] Øystein Flagstad, Eva Hedmark, Arild Landa, Henrik Brøseth, Jens Persson, Roy Andersen, Peter Segerström, and Hans Ellegren. Colonization history and noninvasive monitoring of a reestablished wolverine population. Historia de colonización y monitoreo no invasivo de una población reestablecida de Gulo gulo. Conservation Biology, 18(3):676–688, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2000:STC

[Fie00a] Joseph R. Fiedler. Software for teaching conservation biologists. Conservation Biology, 14(2):585–586, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2000:PHA

[Fie00b] Peggy L. Fiedler. Plants and human affairs in North America. Conservation Biology, 14(2):589–590, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2003:QOC

[Fie03a] Diane Fiedler. Quantifying our Children’s necessity for nature. Conservation Biology, 17(2):643–644, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2003:BR

[Fie03b] Peggy L. Fiedler. Book reviews. Conservation Biology, 17(4): 1179, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2003:QO

[Fie03c] Winifred M. Fiedler. Qualified optimism. Conservation Biology, 17(5):1465–1467, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2004:EA

[Fie04] Peggie L. Fiedler. Elegant arguments. Conservation Biology, 18 (2):585, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2005:LT

[Fie05] Peggy L. Fiedler. Light and time. Conservation Biology, 19 (2):585–586, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2006:IUW

[Fie06] Diane C. Fiedler. Inspiration in unexpected ways. Conservation Biology, 21(2):1821–1823, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiedler:2009:EA

[Fie09] Peggy L. Fiedler. Editor’s acknowledgments. Conservation Biology, 23(6):1645–1647, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fernandez-Juricic:2000:AUW

[FJ00] Esteban Fernández-Juricic. Avifaunal use of wooded streets in an urban landscape. Uso de calles con vegetación por aves en un paisaje urbano. Conservation Biology, 14(2):513–521, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forget:2007:HID

[FJ07] Pierre-Michel Forget and Patrick A. Jansen. Hunting increases dispersal limitation in the tree Carapa procera, a nontimber forest product. La cacería incrementa la Limitación de la dispersión de Carapa procera, una especie de Árbol no maderable. Conservation Biology, 21(1):106–113, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fjeldsaa:2007:HBS

[Fje07] Jon Fjeldså. How broad-scale studies of patterns and processes can serve to guide conservation planning in Africa. Cómo pueden servir los estudios de patrones y procesos a gran escala para guiar la planificación de la conservación en África. Conservation Biology, 21(3):659–667, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuller:2004:LFC

[FJS04] D. O. Fuller, T. C. Jessup, and A. Salim. Loss of forest cover in Kalimantan, Indonesia, since the 1997–1998 El Niño. Pérdida de cobertura forestal en Kalimantan, Indonesia, desde El Niño de 1997–1998. Conservation Biology, 18(1):249–254, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fagan:2005:QRL

[FKU05] William F. Fagan, Christina M. Kennedy, and Peter J. Unmack. Quantifying rarity, losses, and risks for native fishes of the Lower Colorado River Basin: Implications for conservation listing. Cuantificación de rareza, pérdidas y riesgos para peces nativos de la Cuenca Baja del Río Colorado: Implicaciones para las listas de conservación. Conservation Biology, 19(6):1872–1882, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flaspohler:2000:SCE

[Fla00] David J. Flaspohler. Simple concepts, elusive clarity. Conservation Biology, 14(4):1216–1217, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2001:MSR

[Fle01] Erica Fleishman. Moving scientific review beyond academia. Conservation Biology, 15(3):547–549, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleischner:2003:DDW

[Fle03] Thomas L. Fleischner. Diversity deep and wild. Conservation Biology, 17(4):952–953, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleischner:2004:BH

[Fle04a] Thomas L. Fleischner. The book of hope. Conservation Biology, 18(6):1694–1695, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2004:B

[Fle04b] Erica Fleishman. Bio[statistics]philia. Conservation Biology, 18 (1):286–288, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flessa:2008:IB

[Fle08] Karl W. Flessa. An important beginning. Conservation Biology, 22(5):1361–1362, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fischer:2004:AEC

[FLF04] Joern Fischer, David B. Lindenmayer, and Ioan Fazey. Appreciating ecological complexity: Habitat contours as a conceptual landscape model. Valorando la complejidad ecológica: Contornos de hábitat como un modelo conceptual del paisaje. Conservation Biology, 18(5):1245–1253, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flood:2005:STP

[Flo05] Rick Flood. Speak truth to power: Thoughts on a bold strategy for conservation biologists. Conservation Biology, 19(2):293, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuchs:2003:EFF

[FLQ03] Eric J. Fuchs, Jorge A. Lobo, and Mauricio Quesada. Effects of forest fragmentation and flowering phenology on the reproductive success and mating patterns of the tropical dry forest tree Pachira quinata. Efectos de la fragmentación de bosques y la fenología de Floración sobre el Éxito reproductivo y los patrones de apareamiento de Pachira quinata, un Árbol de bosque tropical seco. Conservation Biology, 17(1):149–157, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2009:RAI

[FM09] Erica Fleishman and Dennis D. Murphy. A realistic assessment of the indicator potential of butterflies and other charismatic taxonomic groups. Una evaluación realista del potencial indicador de mariposas y otros grupos taxonómicos carismáticos. Conservation Biology, 23(5):1109–1116, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farrington:2009:IEH

[FMDK09] Susan J. Farrington, Rose-Marie Muzika, Dan Drees, and Tiffany M. Knight. Interactive effects of harvest and deer herbivory on the population dynamics of American ginseng. Efectos interactivos de la cosecha y la herbivoría de venados sobre la dinámica poblacional del ginseng americano. Conservation Biology, 23(3):719–728, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forkner:2006:IAT

[FMLL06] Rebecca E. Forkner, Robert J. Marquis, John T. Lill, and Josiane Le Corff. Impacts of alternative timber harvest practices on leaf-chewing herbivores of oak. Impactos de prácticas alternativas del corte de madera sobre herbívoros de las hojas de roble. Conservation Biology, 20(2):429–440, April 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fox:2005:EAC

[FMP^{+}05] Helen E. Fox, Peter J. Mous, Jos S. Pet, Andreas H. Muljadi, and Roy L. Caldwell. Experimental assessment of coral reef rehabilitation following blast fishing. Evaluación experimental de la Rehabilitación de arrecifes de coral después de pêsca con explosivos. Conservation Biology, 19(1):98–107, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fitzpatrick:2007:BIR

[FMPB07] ÚNa Fitzpatrick, Tomás E. Murray, Robert J. Paxton, and Mark J. F. Brown. Building on IUCN regional red lists to produce lists of species of conservation priority: a model with Irish bees. Utilización de listas rojas IUCN regionales para producir listas de especies de prioridad para la conservación: un modelo con abejas irlandesas. Conservation Biology, 21(5):1324–1332, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferretti:2008:LLP

[FMSL08] Francesco Ferretti, Ransom A. Myers, Fabrizio Serena, and Heike K. Lotze. Loss of large predatory sharks from the Mediterranean Sea. Pérdida de tiburones depredadores grandes en el Mar Mediterráneo. Conservation Biology, 22(4):952–964, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Filipe:2004:SPA

[FMT^{+}04] A. F. Filipe, T. A. Marques, P. Tiago, F. Ribeiro, L. Moreira Da Costa, I. G. Cowx, and M. J. Collares-Pereira. Selection of priority areas for fish conservation in Guadiana River Basin, Iberian Peninsula. Selección de Áreas prioritarias para la conservación de peces en la cuenca del Río Guadiana, Península Ibérica. Conservation Biology, 18(1):189–200, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2002:TDF

[FN02a] Erica Fleishman and Ralph Mac Nally. Topographic determinants of faunal nestedness in Great Basin butterfly assemblages: Applications to conservation planning. Determinantes topográficos de la anidación faunística de ensamblajes de mariposas de la gran cuenca: Aplicaciones en la planeación de la conservación. Conservation Biology, 16(2):422–429, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freyfogle:2002:PSP

[FN02b] Eric T. Freyfogle and Julianne Lutz Newton. Putting science in its place poniendo a la ciencia en su lugar. Conservation Biology, 16(4):863–873, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferrer:2004:NET

[FN04] Miguel Ferrer and Juan José Negro. The near extinction of two large European predators: Super specialists pay a price. La cuasi extinción de dos grandes depredadores europeos mayores: los Súper especialistas pagan el precio. Conservation Biology, 18(2):344–349, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiorello:2006:DHE

[FND06] Christine V. Fiorello, Andrew J. Noss, and Sharon L. Deem. Demography, hunting ecology, and pathogen exposure of domestic dogs in the Isoso of Bolivia. Demografía, ecología cinegética y exposición a patógenos de perros domésticos en el Isoso de Bolivia. Conservation Biology, 20(3):762–771, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2003:VTP

[FNF03] Erica Fleishman, Ralph Mac Nally, and John P. Fay. Validation tests of predictive models of butterfly occurrence based on environmental variables. Pruebas de validación de modelos predictivos de ocurrencia de mariposas basados en variables ambientales. Conservation Biology, 17(3):806–817, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2001:MPS

[FNFM01] Erica Fleishman, Ralph Mac Nally, John P. Fay, and Dennis D. Murphy. Modeling and predicting species occurrence using broad-scale environmental variables: an example with butterflies of the Great Basin. Modelado y predicción de la presencia de especies usando variables ambientales de amplia escala: un ejemplo con mariposas de la gran depresión. Conservation Biology, 15(6):1674–1685, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fox:2005:SSC

[FNM05] Jessica Fox and Anamaria Nino-Murcia. Status of species conservation banking in the United States. Conservation Biology, 19 (4):996–1007, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferrer:2009:SPO

[FNP09] Miguel Ferrer, Ian Newton, and Massimo Pandolfi. Small populations and offspring sex-ratio deviations in eagles. Poblaciones Pequeñas y desviaciones en la Proporción de sexos de crías en Águilas. Conservation Biology, 23(4):1017–1025, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Friedlander:2003:DEM

[FNS^{+}03] A. Friedlander, J. Sladek Nowlis, J. A. Sanchez, R. Appeldoorn, P. Usseglio, C. Mccormick, S. Bejarano, and A. Mitchell-Chui. Designing effective marine protected areas in Seaflower Biosphere Reserve, Colombia, based on biological and sociological information. Diseño de Áreas marinas protegidas efectivas en la reserva de la Biosfera Seaflower, Colombia, en base a información biológica y sociológica. Conservation Biology, 17(6): 1769–1784, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foster:2006:PSH

[FO06] David R. Foster and David A. Orwig. Preemptive and salvage harvesting of New England forests: When doing nothing is a viable alternative. Cosecha preventiva y de salvamento en bosques de Nueva Inglaterra: Cuando hacer nada es una alternativa viable. Conservation Biology, 20(4):959–970, August 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foin:2000:OMM

[Foi00] Ted Foin. One for models, and models for all. Conservation Biology, 14(4):1214–1215, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Folke:2006:EPC

[Fol06] Carl Folke. The economic perspective: Conservation against development versus conservation for development. Conservation Biology, 20(3):686–688, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fonseca:2009:SME

[Fon09] Carlos Roberto Fonseca. The silent mass extinction of insect herbivores in biodiversity hotspots. La extinción masiva silenciosa de insectos herbívoros en sitios de importancia para la biodiversidad. Conservation Biology, 23(6):1507–1515, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foote:2004:IVO

[Foo04] Michael Foote. The inside view from outsiders. Conservation Biology, 18(3):857–858, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forman:2000:EAA

[For00a] Richard T. T. Forman. Estimate of the area affected ecologically by the road system in the United States. Estimación del area ecológicamente afectada por el sistema carretero de los Estados Unidos. Conservation Biology, 14(1):31–35, February 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forney:2000:EMC

[For00b] Karin A. Forney. Environmental models of cetacean abundance: Reducing uncertainty in population trends. Modelos ambientales de abundancia de cetáceos: Reducción de la incertidumbre en tendencias poblacionales. Conservation Biology, 14(5):1271– 1286, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ford:2002:SCD

[For02] Michael J. Ford. Selection in captivity during supportive breeding may reduce fitness in the wild. La selección en cautiverio durante la reproducción de apoyo puede reducir la adaptabilidad en condiciones silvestres. Conservation Biology, 16(3):815– 825, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fox:2007:CBS

[Fox07] Laurel R. Fox. Climatic and biotic stochasticity: Disparate causes of convergent demographies in rare, sympatric plants. Estocasticidad climática y biótica: Causas dispares de Demografías convergentes en plantas Simpátricas, raras. Conservation Biology, 21(6):1556–1561, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fox:2008:PTC

[Fox08] Gordon A. Fox. Playing to a tough crowd. Conservation Biology, 22(1):225–226, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

FelizolaDiniz-Filho:2002:SAA

[FP02] José Alexandre Felizola Diniz-Filho and Mariana Pires De Campos Telles. Spatial autocorrelation analysis and the identification of operational units for conservation in continuous populations. Análisis de autocorrelación espacial e identificación de unidades operacionales para la conservación en poblaciones continuas. Conservation Biology, 16(4):924–935, August 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Folmer:2007:EEP

[FP07] Henk Folmer and Theunis Piersma. Ecological and environmental policies versus a steady state economy in times of crisis. Conservation Biology, 21(5):1136–1137, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiumera:2000:EPS

[FPF00] A. C. Fiumera, P. G. Parker, and P. A. Fuerst. Effective population size and maintenance of genetic diversity in captive-bred populations of a Lake Victoria cichlid. Tamaño poblacional efectivo y mantenimiento de la diversidad genética en poblaciones de un Cíclido del Lago Victoria reproducidas en cautiverio. Conservation Biology, 14(3):886–892, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fiumera:2004:MOP

[FPL^{+}04] Anthony C. Fiumera, Brady A. Porter, Greg Looney, Marjorie A. Asmussen, and John C. Avise. Maximizing offspring production while maintaining genetic diversity in supplemental breeding programs of highly fecund managed species. Programas reproductivos suplementarios para especies bajo manejo: Estrategias para maximizar la producción de crías y mantener la diversidad genética en taxones altamente fecundos. Conservation Biology, 18(1):94–101, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fa:2002:BET

[FPM02] Julia E. Fa, Carlos A. Peres, and Jessica Meeuwig. Bushmeat exploitation in tropical forests: an intercontinental comparison. Explotación de carne silvestre en bosques tropicales: una comparación intercontinental. Conservation Biology, 16(1):232–237, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ficetola:2009:ILE

[FPSD09] Gentile Francesco Ficetola, Emilio Padoa-Schioppa, and Fiorenza De Bernardi. Influence of landscape elements in riparian buffers on the conservation of semiaquatic amphibians. Influencia de elementos del paisaje en zonas de amortiguamiento ribereñas sobre la conservación de anfibios semiacuáticos. Conservation Biology, 23(1):114–123, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foley:2001:NSR

[FPW01] C. A. H. Foley, S. Papageorge, and S. K. Wasser. Noninvasive stress and reproductive measures of social and ecological pressures in free-ranging African elephants. Estrés no invasivo y medidas reproductivas de presiones sociales y ecológicas en elefantes africanos libres. Conservation Biology, 15(4):1134–1142, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foster:2007:IBF

[FR07] Jeffrey T. Foster and Scott K. Robinson. Introduced birds and the fate of Hawaiian rainforests. Aves introducidas y el destino de bosques lluviosos de Hawai. Conservation Biology, 21(5): 1248–1257, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frankham:2001:IEI

[Fra01] Richard Frankham. Inbreeding and extinction in island populations: Reply to Elgar and Clode. Conservation Biology, 15(1):

287–289, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [EC01].

Frasz:2004:EEF

[Fra04] Geoffrey Frasz. Environmental ethics and forestry: a reader. Conservation Biology, 18(2):586–587, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frazer:2009:SCA

[Fra09] L. Neil Frazer. Sea-cage aquaculture, sea lice, and declines of wild fish acuacultura en jaulas marinas, piojos de mar y declinaciones de peces silvestres. Conservation Biology, 23(3):599–607, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fausch:2009:IVI

[FRD^{+}09] Kurt D. Fausch, Bruce E. Rieman, Jason B. Dunham, Michael K. Young, and Douglas P. Peterson. Invasion versus isolation: Trade-offs in managing native salmonids with barriers to upstream movement. Invasión versus aislamiento: Pros y contras del manejo de Salmónidos con barreras al movimiento Río arriba. Conservation Biology, 23(4):859–870, August 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freyfogle:2003:CCW

[Fre03] Eric T. Freyfogle. Conservation and the culture war. Conservation Biology, 17(2):354–355, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freyfogle:2004:CLG

[Fre04] Eric T. Freyfogle. Conservation and the lure of the garden. Conservación y el encanto del jardín. Conservation Biology, 18 (4):995–1003, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freyfogle:2006:CBL

[Fre06] Eric T. Freyfogle. Conservation biology and law: Only a start. Conservation Biology, 20(3):679–680, June 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Faith:2004:IPD

[FRH04] Daniel P. Faith, C. A. M. Reid, and James Hunter. Integrating phylogenetic diversity, complementarity, and endemism for conservation assessment. Conservation Biology, 18(1):255–261, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foxcroft:2007:RAR

[FRR07] Llewellyn C. Foxcroft, Mathieu Rouget, and David M. Richardson. Risk assessment of riparian plant invasions into protected areas. Evaluación del riesgo de invasiones de plantas ribereñas Áreas protegidas. Conservation Biology, 21(2):412–421, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2002:ARP

[FRSG^{+}02] Erica Fleishman, Chris Ray, Per Sjögren-Gulve, Carol L. Boggs, and Dennis D. Murphy. Assessing the roles of patch quality, area, and isolation in predicting metapopulation dynamics. Evaluación del papel de la calidad, Área y aislamiento de parches en la Predicción de la dinámica metapoblacional. Conservation Biology, 16(3):706–716, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fisher:2002:SPA

[FSC02] Robert N. Fisher, Andrew V. Suarez, and Ted J. Case. Spatial patterns in the abundance of the coastal horned lizard. Patrones espaciales en la abundancia del falso camaleón costeño. Conservation Biology, 16(1):205–215, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flagstad:2000:GVS

[FSS^{+}00] Øystein Flagstad, Per Ole Syvertsen, Nils ChR. Stenseth, John E. Stacy, Ingrid Olsaker, Knut H. Røed, and Kjetill S. Jakobsen. Genetic variability in Swayne’s hartebeest, an endangered antelope of Ethiopia. Variabilidad genética en el alcelafo de Swayne, un antílope de Etiopía en peligro. Conservation Biology, 14(1):254–264, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fernandez:2004:MIC

[FTC04a] J. Fernández, M. A. Toro, and A. Caballero. Managing individuals’ contributions to maximize the allelic diversity maintained in small, conserved populations. Manejo de contribuciones de individuos para maximizar la diversidad alélica mantenida en poblaciones conservadas Pequeñas. Conservation Biology, 18(5):

1358–1367, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foster:2004:LTP

[FTC^{+}04b] Jeffrey T. Foster, Erik J. Tweed, Richard J. Camp, Bethany L. Woodworth, Corey D. Adler, and Tom Telfer. Long-term population changes of native and introduced birds in the Alaka’i Swamp, Kaua’i. Cambios poblacionales a largo plazo de aves nativas e introducidas en el Pantano Alaka’i, Kaua’i. Conservation Biology, 18(3):716–725, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:2005:UIS

[FTN^{+}05] Erica Fleishman, James R. Thomson, Ralph Mac Nally, Dennis D. Murphy, and John P. Fay. Using indicator species to predict species richness of multiple taxonomic groups. Utilización de especies indicadoras para predecir la riqueza de especies de múltiples grupos taxonómicos. Conservation Biology, 19(4): 1125–1137, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flaspohler:2001:EFE

[FTR01] David J. Flaspohler, Stanley A. Temple, and Robert N. Rosenfield. Effects of forest edges on ovenbird demography in a managed forest landscape. Efecto de los bordes del bosque en la demografía del chipe suelero en un paisaje de bosque manejado. Conservation Biology, 15(1):173–183, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuentes:2006:PP

[Fue06] Marcelino Fuentes. Privatization of parks. Conservation Biology, 20(4):934, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuller:2000:FPW

[Ful00] D. O. Fuller. The fallacy of popular wisdom in African wildlife management. Conservation Biology, 14(2):586–587, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuller:2002:EQV

[Ful02] Steven G. Fuller. Environmental quality versus social justice. Conservation Biology, 16(2):569–570, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foster:2005:ESI

[FV05] S. J. Foster and A. C. J. Vincent. Enhancing sustainability of the international trade in seahorses with a single minimum size limit. Incremento de la sustentabilidad del comercio internacional de caballitos de mar con un límite de talla mínima Único. Conservation Biology, 19(4):1044–1050, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferguson:2003:PAS

[FVMG03] Bruce G. Ferguson, John Vandermeer, Helda Morales, and Daniel M. Griffith. Post-agricultural succession in El Petén, Guatemala. Sucesión post–agrícola en El Petén, Guatemala. Conservation Biology, 17(3):818–828, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fanshawe:2003:RTC

[FVS03] Samantha Fanshawe, Glenn R. Vanblaricom, and Alice A. Shelly. Restored top carnivores as detriments to the performance of marine protected areas intended for fishery sustainability: a case study with red abalones and sea otters. Carnívoros restaurados como detrimento del funcionamiento de Áreas marinas protegidas para pesquerías sustentables: un estudio de caso con Abulón rojo y nutrias marinas. Conservation Biology, 17(1):273–283, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fa:2000:BMB

[FYC00] Julia E. Fa, Juan E. Garcia Yuste, and Ramon Castelo. Bushmeat markets on Bioko Island as a measure of hunting pressure. Mercados de carne silvestre en la Isla Bioko como medida de la presión por caza. Conservation Biology, 14(6):1602–1613, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gippoliti:2002:AWM

[GA02] Spartaco Gippoliti and Giovanni Amori. Anthropochorous wild mammal taxa and conservation lists. Conservation Biology, 16 (4):1162–1164, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guariguata:2000:SRF

[GAF00] Manuel R. Guariguata, Jesús J. Rosales Adame, and Bryan Finegan. Seed removal and fate in two selectively logged lowland forests with constrasting protection levels. Remoción y destino de semillas en dos bosques de tierras bajas selectivamente talados con niveles contrastantes de protección. Conservation Biology, 14(4):1046–1054, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galetti:2001:FAF

[Gal01a] Mauro Galetti. The future of the Atlantic forest. Conservation Biology, 15(1):4, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galetti:2001:IWC

[Gal01b] Mauro Galetti. Indians within conservation units: Lessons from the Atlantic forest. Conservation Biology, 15(3):798–799, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galvin:2003:WWT

[Gal03] Kathleen A. Galvin. Where the wild things are. Conservation Biology, 17(4):1192–1193, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Golet:2009:CPF

[GALW09] Gregory H. Golet, Beverley Anderson, Ryan A. Luster, and Gregg Werner. Collaborative planning fosters multiple-benefit restoration projects on the Sacramento River. Conservation Biology, 23(6):1634–1637, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia:2003:DVR

[Gar03a] María B. García. Demographic viability of a relict population of the critically endangered plant Borderea chouardii. Viabilidad demográfica de una población relicta de la planta en peligro crítico Borderea chouardii. Conservation Biology, 17(6): 1672–1680, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garner:2003:SS

[Gar03b] Alisse Garner. Spirituality and sustainability. Conservation Biology, 17(4):946–947, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gavin:2004:CFU

[Gav04] Michael C. Gavin. Changes in forest use value through ecological succession and their implications for land management in the Peruvian Amazon. Cambios en el valor de uso del bosque por medio de la sucesión ecológica y sus implicaciones para la gestión de tierras en la Amazonía peruana. Conservation Biology, 18(6): 1562–1570, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaydos:2002:BIC

[Gay02] Joseph K. Gaydos. Biological invasion: a case study. Conservation Biology, 16(6):1665–1666, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbs:2001:CWC

[GB01] James P. Gibbs and Alvin R. Breisch. Climate warming and calling phenology of frogs near Ithaca, New York, 1900–1999. Calentamiento del clima y fenología de vocalizaciones de ranas cerca de Ithaca, Nueva York, 1900–1999. Conservation Biology, 15(4):1175–1178, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gelbard:2003:RCE

[GB03] Jonathan L. Gelbard and Jayne Belnap. Roads as conduits for exotic plant invasions in a semiarid landscape. Caminos como conductos para invasiones de plantas en un paisaje semiárido. Conservation Biology, 17(2):420–432, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Granek:2005:CMA

[GB05a] Elise F. Granek and Mark A. Brown. Co-management approach to marine conservation in Mohéli, Comoros Islands. Estrategia de cogestión para la conservación marina en Mohéli, Islas Comoros. Conservation Biology, 19(6):1724–1732, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grogan:2005:BLM

[GB05b] James Grogan and Paulo Barreto. Big-leaf mahogany on CITES appendix II: Big challenge, big opportunity. Conservation Biology, 19(3):973–976, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Germano:2009:SAR

[GB09] Jennifer M. Germano and Phillip J. Bishop. Suitability of amphibians and reptiles for translocation. Aptitud de anfibios y reptiles para la translocación. Conservation Biology, 23(1):7– 15, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gratwicke:2008:WCW

[GBB^{+}08] Brian Gratwicke, Elizabeth L. Bennett, Steven Broad, Sarah Christie, Adam Dutton, Grace Gabriel, Craig Kirkpatrick, and Kristin Nowell. The world can’t have wild tigers and eat them, too. Conservation Biology, 22(1):222–223, February 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grand:2004:MLA

[GBC^{+}04] Joanna Grand, John Buonaccorsi, Samuel A. Cushman, Curtice R. Griffin, and Maile C. Neel. A multiscale landscape approach to predicting bird and moth rarity hotspots in a threatened pitch pine–scrub oak community. Un método de paisaje multiescala para predecir Áreas de rareza crítica de aves y polillas en una comunidad amenazada de pino–encino. Conservation Biology, 18(4):1063–1077, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Green:2005:FIM

[GBC^{+}05] Rhys E. Green, Andrew Balmford, Peter R. Crane, Georgina M. Mace, John D. Reynolds, and R. Kerry Turner. A framework for improved monitoring of biodiversity: Responses to the World Summit on Sustainable Development. Un marco para mejorar el monitoreo de biodiversidad: Respuestas a la Cumbre Mundial de Desarrollo Sustentable. Conservation Biology, 19(1):56–65, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffin:2000:TCB

[GBE00] Andrea S. Griffin, Daniel T. Blumstein, and Christopher S. Evans. Training captive-bred or translocated animals to avoid predators. Entrenamiento de animales para evitar depredadores. Conservation Biology, 14(5):1317–1326, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guisan:2006:UNB

[GBE^{+}06] Antoine Guisan, Olivier Broennimann, Robin Engler, Mathias Vust, Nigel G. Yoccoz, Anthony Lehmann, and Niklaus E. Zimmermann. Using niche-based models to improve the sampling of rare species. Utilización de modelos basados en nichos para mejorar el muestreo de especies raras. Conservation Biology, 20 (2):501–511, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Granek:2005:BOI

[GBH^{+}05] Elise F. Granek, Daniel R. Brumbaugh, Scott A. Heppell, Selina S. Heppell, and David Secord. A blueprint for the oceans: Implications of two National Commission reports for conservation practitioners. Un anteproyecto para los océanos: Implicaciones de dos reportes de Comisiones Nacionales para practicantes de la conservación. Conservation Biology, 19(4):1008– 1018, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gray:2009:GMP

[GBH^{+}09] Thomas N. E. Gray, Ro Borey, Seng Kim Hout, Hong Chamnan, Nigel, J. Collar, and Paul M. Dolman. Generality of models that predict the distribution of species: Conservation activity and reduction of model transferability for a threatened bustard. Generalidad de los modelos que predicen la distribución de especies: Actividades de conservación y Reducción de la transferibilidad del modelo para una avutarda amenazada. Conservation Biology, 23(2):433–439, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gray:2007:RAF

[GBMH07] Michael A. Gray, Sandra L. Baldauf, Peter J. Mayhew, and Jane K. Hill. The response of avian feeding guilds to tropical forest disturbance. Respuesta de gremios alimentarios de aves a la perturbación de bosques tropicales. Conservation Biology, 21 (1):133–141, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gillespie:2006:PPI

[GC06] Thomas R. Gillespie and Colin A. Chapman. Prediction of parasite infection dynamics in primate metapopulations based on attributes of forest fragmentation. Predicción de la dinámica de la infección parasitaria en metapoblaciones de primates con base en atributos de la fragmentación de bosques. Conservation Biology, 20(2):441–448, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia:2007:SDE

[GC07] Daniel García and Natacha P. Chacoff. Scale-dependent effects of habitat fragmentation on hawthorn pollination, frugivory, and seed predation. Efectos de la fragmentación dependientes de la escala sobre la polinización, frugivoría y depredación de semillas de Crataegus monogyna. Conservation Biology, 21(2):400–411, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gray:2009:SPA

[GC09] Noella J. Gray and Lisa M. Campbell. Science, policy advocacy, and Marine Protected Areas. Ciencia, cabildeo Político y Áreas Marinas Protegidas. Conservation Biology, 23(2):460–468, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gleason:2009:WMB

[GCBF09] Mary Gleason, Chuck Cook, Michael Bell, and Erika Feller. Are we missing the boat? Collaborative solutions for North American fish wars. Conservation Biology, 23(5):1065–1067, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gardner:2007:CVM

[GCF^{+}07] Toby A. Gardner, Tim Caro, Emily B. Fitzherbert, Tasila Banda, and Punit Lalbhai. Conservation value of multipleuse areas in East Africa. El valor de conservación de Áreas de uso Múltiple en África Oriental. Conservation Biology, 21(6): 1516–1525, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerson:2008:UTS

[GCM^{+}08] Helen Gerson, Becky Cudmore, Nicholas E. Mandrak, Lonny D. Coote, Ken Farr, and Guy Baillargeon. Use of the Taxonomic Serial Number (TSN) as a required data element in international wildlife trade: Response to Fragoso and Ferriss. Conservation Biology, 22(6):1651–1654, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gering:2003:APS

[GCV03] Jon C. Gering, Thomas O. Crist, and Joseph A. Veech. Additive partitioning of species diversity across multiple spatial scales: Implications for regional conservation of biodiversity. División aditiva de la diversidad de especies en escalas espaciales múltiples: Implicaciones para la conservación de biodiversidad regional. Conservation Biology, 17(2):488–499, April 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grayson:2005:PRG

[GD05] Donald K. Grayson and Françoise Delpech. Pleistocene reindeer and global warming. Renos del Pleistoceno y calentamiento global. Conservation Biology, 19(2):557–562, April 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goldstein:2000:CGS

[GDAV00] Paul Z. Goldstein, Robert Desalle, George Amato, and Alfried P. Vogler. Conservation genetics at the species boundary. Genética de la conservación en los límites del nivel de especie. Conservation Biology, 14(1):120–131, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gelcich:2005:IAD

[GEJK05] Stefan Gelcich, Gareth Edwards-Jones, and Michel J. Kaiser. Importance of attitudinal differences among artisanal fishers toward co-management and conservation of marine resources. Importancia de las diferencias de actitudes entre pescadores artesanales en la cogestión y conservación de recursos marinos. Conservation Biology, 19(3):865–875, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerber:2003:DSU

[Ger03] Leah Gerber. Delisting of species under the ESA. Conservation Biology, 17(3):651–652, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Getz:2005:WP

[Get05] Wayne M. Getz. Whom are parks for? Conservation Biology, 19 (4):1333–1334, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia-Fernandez:2003:BGN

[GFCP03] Carmen García-Fernández, Miguel A. Casado, and Manuel Ruiz Pérez. Benzoin gardens in North Sumatra, Indonesia: Effects of management on tree diversity. Jardines de benjuí en Sumatra Norte, Indonesia: Consecuencias de la gestión en la diversidad de especies arbóreas. Conservation Biology, 17(3):829–836, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guidetti:2004:EDC

[GFTB04] Paolo Guidetti, Simonetta Fraschetti, Antonio Terlizzi, and Ferdinando Boero. Effects of desertification caused by Lithophaga lithophaga (Mollusca) fishery on littoral fish assemblages along rocky coasts of Southeastern Italy. Efectos de la Desertificación ocasionada por la pesquería de Lithophaga lithophaga (Mollusca) sobre ensambles de peces de litoral a lo largo de Costas rocosas del Sureste de Italia. Conservation Biology, 18(5):1417–1423, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Groom:2002:BSA

[GG02] Jeremiah D. Groom and Thomas C. Grubb, Jr. Bird species associated with riparian woodland in fragmented, temperatedeciduous forest especies de aves asociadas con zonas riparias arboladas fragmentadas de bosque templado caducifolio. Conservation Biology, 16(3):832–836, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goldman:2006:BES

[GG06] Douglas H. Goldman and Barbara Gravendeel. Bringing ex situ orchid conservation to the home garden. Conservation Biology, 21(2):1829–1830, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garnier-Gere:2001:ESP

[GGA01] Pauline Hélène Garnier-Géré and Peter Kevin Ades. Environmental surrogates for predicting and conserving adaptive genetic variability in tree species. Caracteres ambientales para la predicción y conservación de la variabilidad genética adaptativa en especies de Árboles. Conservation Biology, 15(6):1632–1644, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guzman:2003:NDM

[GGC03] Héctor M. Guzmán, Carlos Guevara, and Arcadio Castillo. Natural disturbances and mining of Panamanian coral reefs by indigenous people. Perturbaciones naturales y explotación de arrecifes coralinos panameños por habitantes nativos. Conservation Biology, 17(5):1396–1401, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galindo-Gonzalez:2000:BBG

[GGGS00] Jorge Galindo-González, Sergio Guevara, and Vinicio J. Sosa. Bat- and bird-generated seed rains at isolated trees in pastures in a tropical rainforest. Dispersión de semillas generada por murciélagos y aves bajo arboles aislados en pastizales de una selva alta perennifolia. Conservation Biology, 14(6):1693–1703, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goldberg:2005:IID

[GGI^{+}05] Tony L. Goldberg, Emily C. Grant, Kate R. Inendino, Todd W. Kassler, Julie E. Claussen, and David P. Philipp. Increased infectious disease susceptibility resulting from outbreeding depression. Incremento de la susceptibilidad a enfermedades infecciosas debido a depresión exogámica. Conservation Biology, 19 (2):455–462, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Groom:2008:BBP

[GGT08] Martha J. Groom, Elizabeth M. Gray, and Patricia A. Townsend. Biofuels and biodiversity: Principles for creating better policies for biofuel production. Biocombustibles y biodiversidad: Principios para la Creación de mejores políticas para la producción de biocombustible. Conservation Biology, 22(3): 602–609, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Greenwald:2009:ELC

[GGW09] Katherine R. Greenwald, H. Lisle Gibbs, and Thomas A. Waite. Efficacy of land-cover models in predicting isolation of marbled salamander populations in a fragmented landscape. Eficacia de los modelos de cobertura de suelo para la Predicción del aislamiento de poblaciones de Ambystoma opacum en un paisaje fragmentado. Conservation Biology, 23(5):1232–1241, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guerry:2002:ADL

[GH02] Anne D. Guerry and Malcolm L. Hunter, Jr. Amphibian distributions in a landscape of forests and agriculture: an examination of landscape composition and configuration. Distribución de anfibios en un paisaje de bosques y agricultura: un análisis de la composición y configuración del paisaje. Conservation Biology, 16(3):745–754, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ghazoul:2001:BBC

[Gha01] Jaboury Ghazoul. Barriers to biodiversity conservation in forest certification. Conservation Biology, 15(2):315–317, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ghazoul:2007:CUE

[Gha07] Jaboury Ghazoul. Challenges to the uptake of the ecosystem service rationale for conservation. Conservation Biology, 22(1): 1651–1652, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ghazoul:2008:DES

[Gha08] Jaboury Ghazoul. Debating the ecosystem service rationale for conservation: Response to Kremen et al. Conservation Biology, 22(3):799–801, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ghiglieri:2006:GNB

[Ghi06] Michael P. Ghiglieri. Good news, bad news in Uganda. Conservation Biology, 20(5):1561–1562, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guarderas:2008:CSM

[GHL08] A. Paulina Guarderas, Sally D. Hacker, and Jane Lubchenco. Current status of marine protected areas in Latin America and the Caribbean. Estado actual de las Áreas marinas protegidas en Latinoamérica y el Caribe. Conservation Biology, 22(6): 1630–1640, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Giulietti:2005:BCP

[GHP^{+}05] Ana Maria Giulietti, Raymond M. Harley, Luciano Paganucci De Queiroz, Maria Das Graças Lapa Wanderley, and Cassio Van Den Berg. Biodiversity and conservation of plants in Brazil. Biodiversidad y conservación de plantas en Brasil. Conservation Biology, 19(3):632–639, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gardenfors:2001:AIR

[GHTMR01] Ulf Gärdenfors, Craig Hilton-Taylor, Georgina M. Mace, and Jon Paul Rodríguez. The application of IUCN red list criteria at regional levels. Conservation Biology, 15(5):1206–1212, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbs:2000:WLB

[Gib00] James P. Gibbs. Wetland loss and biodiversity conservation. Pérdida de humedales y conservación de la biodiversidad. Conservation Biology, 14(1):314–317, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbons:2007:SAG

[Gib07] Michael Gibbons. A swan among geese. Conservation Biology, 21(5):1377–1378, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gill:2001:EDF

[Gil01a] A. Malcolm Gill. Economically destructive fires and biodiversity conservation: an Australian perspective. Conservation Biology, 15(6):1558–1560, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gillespie:2001:AEC

[Gil01b] Thomas W. Gillespie. Application of extinction and conservation theories for forest birds in Nicaragua. Aplicación de Teorías de extinción y conservación en aves de bosque de Nicaragua. Conservation Biology, 15(3):699–709, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilbert:2004:PV

[Gil04] Barrie K. Gilbert. Points of view. Conservation Biology, 18(5): 1443–1444, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gillespie:2008:HE

[Gil08] Thomas R. Gillespie. 98% human: 91% extinct. Conservation Biology, 22(1):224–225, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ginsberg:2002:C

[Gin02] Joshua Ginsberg. CITES at 30, or 40. Conservation Biology, 16(5):1184–1191, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gundale:2005:SNH

[GJD05] Michael J. Gundale, William M. Jolly, and Thomas H. Deluca. Susceptibility of a northern hardwood forest to exotic earthworm invasion. Susceptibilidad de un bosque boreal a la invasión de lombrices terrestres exóticas. Conservation Biology, 19(4):1075– 1083, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guo:2008:RGS

[GJLL08] Songtao Guo, Weihong Ji, Baoguo Li, and Ming Li. Response of a group of Sichuan snub-nosed monkeys to commercial logging in the Qinling Mountains, China. Respuesta a la explotación forestal comercial de un grupo de monos Rhinopithecus roxellana en las Montañas Qinling, China. Conservation Biology, 22(4): 1055–1064, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbs:2006:EWC

[GK06] James P. Gibbs and Nancy E. Karraker. Effects of warming conditions in Eastern North American forests on red-backed salamander morphology. Efectos de las condiciones de calentamiento en bosques orientales de Norte América sobre la morfología de Plethodon cinereus. Conservation Biology, 21(3):913–917, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Geburek:2008:WCF

[GK08] Thomas Geburek and H. Konrad. Why the conservation of forest genetic resources has not worked. Porqué no ha funcionado la conservación de recursos genéticos forestales. Conservation Biology, 22(2):267–274, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gowdy:2009:FNG

[GK09] John Gowdy and Lisi Krall. The fate of nauru and the global financial meltdown. Conservation Biology, 23(2):257–258, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gajardo:2003:CAB

[GL03] Gonzalo Gajardo and Linda Laikre. Chilean aquaculture boom is based on exotic salmon resources: a conservation paradox. Conservation Biology, 17(4):1173–1174, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbons:2008:FST

[GLF^{+}08] P. Gibbons, D. B. Lindenmayer, J. Fischer, A. D. Manning, A. Weinberg, J. Seddon, P. Ryan, and G. Barrett. The future of scattered trees in agricultural landscapes. El futuro de Árboles dispersos en paisajes agrícolas. Conservation Biology, 22(5):1309–1319, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gigon:2000:BLT

[GLMN00] Andreas Gigon, Regula Langenauer, Claude Meier, and Bernhard Nievergelt. Blue lists of threatened species with stabilized or increasing abundance: a new instrument for conservation. Listas azules de especies amenazadas con abundancia mantenida o aumentada — un nuevo instrumento para la protección de la naturaleza en Suiza. Conservation Biology, 14(2):402–413, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilna:2005:DNZ

[GLV05] B. Gilna, D. B. Lindenmayer, and K. L. Viggers. Dangers of New Zealand possum biocontrol research to endogenous Australian fauna. Conservation Biology, 19(6):2030–2032, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerlach:2000:FLC

[GM00a] Gabriele Gerlach and Kerstin Musolf. Fragmentation of landscape as a cause for genetic subdivision in bank voles. Fragmentación del paisaje como causa de la subdivisión genética en ratones de campo. Conservation Biology, 14(4):1066–1074, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goguen:2000:LGC

[GM00b] Christopher B. Goguen and Nancy E. Mathews. Local gradients of cowbird abundance and parasitism relative to livestock grazing in a western landscape gradientes de abundancia locales de tordos y parasitismo relativo al pastoreo de ganado en un paisaje occidental. Conservation Biology, 14(6):1862–1869, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

GomezDeSilva:2001:ECS

[GM01] Héctor Gómez De Silva and Rodrigo A. Medellín. Evaluating completeness of species lists for conservation and macroecology: a case study of Mexican land birds. Evaluación para determinar si las listas de especies para conservación y Macroecología son completas: Caso de estudio de aves terrestres mexicanas. Conservation Biology, 15(5):1384–1395, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grech:2008:RAR

[GM08a] A. Grech and H. Marsh. Rapid assessment of risks to a mobile marine mammal in an ecosystem-scale marine protected area. Evaluación rápida de riesgos para un mamífero marino móvil en una Área marina protegida a escala de ecosistema. Conservation Biology, 22(3):711–720, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guadarrama-Maillot:2008:PB

[GM08b] Veronica Guadarrama-Maillot. The popularity of biodiversity. Conservation Biology, 22(2):233–234, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grouios:2009:UMA

[GM09] Christopher P. Grouios and Lisa L. Manne. Utility of measuring abundance versus consistent occupancy in predicting biodiversity persistence. La utilidad de medir la abundancia versus la ocupación consistente en la Predicción de la persistencia de la biodiversidad. Conservation Biology, 23(5):1260–1269, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gomez-Mendoza:2007:MEC

[GMA07] Leticia Gómez-Mendoza and Laura Arriaga. Modeling the effect of climate change on the distribution of oak and pine species of Mexico. Modelando el efecto del cambio climático sobre la distribución de especies de encino y pino de México. Conservation Biology, 21(6):1545–1555, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Granek:2008:ERF

[GMB^{+}08] E. F. Granek, E. M. P. Madin, M. A. Brown, W. Figueira, D. S. Cameron, Z. Hogan, G. Kristianson, P. De Villiers, J. E. Williams, J. Post, S. Zahn, and R. Arlinghaus. Engaging recreational fishers in management and conservation: Global case studies. Involucrando a pescadores recreativos en el manejo y la conservación: Estudios de caso globales. Conservation Biology, 22(5):1125–1134, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gallo:2009:KDR

[GMCP09] Leonardo A. Gallo, Paula Marchelli, Luis Chauchard, and Marcelo Gonzalez Peñalba. Knowing and doing: Research leading to action in the conservation of forest genetic diversity of Patagonian temperate forests. Conservation Biology, 23(4):895– 898, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goreau:2000:CCR

[GMHS00] Tom Goreau, Tim McClanahan, Ray Hayes, and Al Strong. Conservation of coral reefs after the 1998 global bleaching event. Conservation Biology, 14(1):5–15, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Game:2008:SWP

[GMMPP08] Edward T. Game, Eve Mcdonald-Madden, Marji L. Puotinen, and Hugh P. Possingham. Should we protect the strong or the weak? Risk, resilience, and the selection of marine protected areas. `?Debemos proteger al fuerte o al débil? Riesgo, resiliencia y la selección de Áreas marinas protegidas. Conservation Biology, 22(6):1619–1629, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gobush:2008:LTI

[GMW08] K. S. Gobush, B. M. Mutayoba, and S. K. Wasser. Long-term impacts of poaching on relatedness, stress physiology, and reproductive output of adult female African elephants. Impactos a largo plazo de la cacería furtiva sobre el grado de relación, la fisiología del estrés y la productividad de hembras adultas de elefantes africanos. Conservation Biology, 22(6):1590–1599, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goldingay:2000:ERO

[GN00] Ross L. Goldingay and David A. Newell. Experimental rock outcrops reveal continuing habitat disturbance for an endangered Australian snake. Farallones de roca experimentales revelan la perturbación continua del hábitat para una serpiente Australiana amenazada. Conservation Biology, 14(6):1908–1912, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griesser:2007:IFP

[GNEE07] Michael Griesser, Magdalena Nystrand, Sönke Eggers, and Jan Ekman. Impact of forestry practices on fitness correlates and population productivity in an open-nesting bird species. Impacto de las prácticas forestales sobre la adaptabilidad y la productividad de una especie de ave de nido abierto. Conservation Biology, 21(3):767–774, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gurd:2001:CME

[GNR01] D. Brent Gurd, Thomas D. Nudds, and Donald H. Rivard. Conservation of mammals in Eastern North American wildlife reserves: How small is too small? Conservation Biology, 15(5): 1355–1363, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grant:2007:AIU

[GOL^{+}07] Jacqualine B. Grant, Julian D. Olden, Joshua J. Lawler, Cara R. Nelson, and Brian R. Silliman. Academic institutions in the United States and Canada ranked according to research productivity in the field of conservation biology. Conservation Biology, 21(5):1139–1144, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gorchov:2008:CSR

[Gor08] David L. Gorchov. Clarification on seed-rain data in Kelm et al. Conservation Biology, 22(6):1369–1370, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gossow:2007:NPA

[Gos07] Hartmut Gossow. No “Pat” answers. Conservation Biology, 21 (1):282–283, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gottlieb:2006:RCI

[Got06] Roger S. Gottlieb. Real coda to an imaginary trial. Conservation Biology, 20(6):1574–1575, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffiths:2008:CBR

[GP08] Richard A. Griffiths and Lissette Pavajeau. Captive breeding, reintroduction, and the conservation of amphibians. Reproducción en cautiverio, reintroducción y la conservación de anfibios. Conservation Biology, 22(4):852–861, August 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gram:2003:EEF

[GPC^{+}03] Wendy K. Gram, Paul A. Porneluzi, Richard L. Clawson, John Faaborg, and Stephen C. Richter. Effects of experimental forest management on density and nesting success of bird species in Missouri Ozark forests. Efectos del manejo experimental de bosques sobre la densidad y el Éxito de anidación de especies de aves en los bosques Ozark de Missouri. Conservation Biology, 17(5):1324–1337, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Glemin:2008:CLM

[GPMM08] Sylvain Glémin, Christophe Petit, Sandrine Maurice, and Agnès Mignot. Consequences of low mate availability in the rare self-incompatible species Brassica insularis. Consecuencias de la baja disponibilidad de pareja en Brassica insularis, especie autoincompatible rara. Conservation Biology, 22(1):216–221, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goffredo:2004:VMC

[GPZ04] Stefano Goffredo, Corrado Piccinetti, and Francesco Zaccanti. Volunteers in marine conservation monitoring: a study of the distribution of seahorses carried out in collaboration with recreational scuba divers. Voluntarios en el monitoreo de conservación marina: un estudio de distribución de caballitos de mar llevado a cabo con buzos scuba recreativos. Conservation Biology, 18(6): 1492–1503, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaston:2003:RSR

[GR03] Kevin J. Gaston and Ana S. L. Rodrigues. Reserve selection in regions with poor biological data. Selección de reservas en regiones con datos biológicos pobres. Conservation Biology, 17 (1):188–195, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Graham:2001:FIM

[Gra01] Catherine H. Graham. Factors influencing movement patterns of keel-billed toucans in a fragmented tropical landscape in Southern Mexico. Factores que influyen sobre los patrones de movimiento de tucanes en un paisaje tropical fragmentado en el sur de México. Conservation Biology, 15(6):1789–1798, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grauel:2003:CEF

[Gra03] William T. Grauel. Un compendio de ecología forestal. (Spanish) [A compendium of forest ecology]. Conservation Biology, 21(6):

341–342, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grajal:2009:IVS

[Gra09] Alejandro Grajal. Increasing the value of SCB membership for young professionals. Conservation Biology, 23(2):246–247, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gabbe:2002:TSP

[GRB02] Aaron P. Gabbe, Scott K. Robinson, and Jeffrey D. Brawn. Treespecies preferences of foraging insectivorous birds: Implications for floodplain forest restoration. Preferencia de aves forrajeras insectívoras por especies de Árboles: Implicaciones en la restauración de bosques de llanuras inundables. Conservation Biology, 16(2):462–470, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Green:2005:DUS

[Gre05] David M. Green. Designatable units for status assessment of endangered species. Unidades designatables para la evaluación del estatus de especies en peligro. Conservation Biology, 19(6): 1813–1820, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Greene:2006:TVS

[Gre06] Harry W. Greene. A timelessly valuable survey. Conservation Biology, 21(3):921–923, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Greenwald:2009:ESC

[Gre09] D. Noah Greenwald. Effects on species’ conservation of reinterpreting the phrase “Significant Portion of its Range” in the U.S. Endangered Species Act. Conservation Biology, 23(6): 1374–1377, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grimm:2005:IBM

[GRG^{+}05] Volker Grimm, Eloy Revilla, Jürgen Groeneveld, Stephanie Kramer-Schadt, Monika Schwager, Jörg Tews, Matthias C. Wichmann, and Florian Jeltsch. Importance of buffer mechanisms for population viability analysis. Conservation Biology, 19(2):578–580, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garner:2005:PGD

[GRH05] Alisse Garner, Janet L. Rachlow, and Jason F. Hicks. Patterns of genetic diversity and its loss in mammalian populations. Patrones de diversidad genética y su pérdida en poblaciones de mamíferos. Conservation Biology, 19(4):1215–1221, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffith:2000:ART

[Gri00] Daniel M. Griffith. Agroforestry: a refuge for tropical biodiversity after fire. Conservation Biology, 14(1):325–326, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gardner:2007:VPS

[GRJB^{+}07] Toby A. Gardner, Marco Antônio Ribeiro-Júnior, Jos Barlow, Teresa Cristina Sauer Ávila-Pires, Marinus S. Hoogmoed, and Carlos A. Peres. The value of primary, secondary, and plantation forests for a neotropical herpetofauna. El valor de la biodiversidad de bosques primarios, secundarios y plantaciones para una herpetofauna neotropical. Conservation Biology, 21(3):775–787, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Group:2005:MPW

[Gro05a] Centread Working Group. Making parks work: a thoughtprovoking argument, but not a guide. Conservation Biology, 19 (1):279–281, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Group:2005:FFC

[Gro05b] The Centread Working Group. Farmers and the forest: Can agroforestry actually conserve biodiversity? Conservation Biology, 19(6):2043–2044, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grumbine:2004:NGB

[Gru04] Ed Grumbine. No going back to tradition. Conservation Biology, 18(2):594–595, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grumbine:2008:CCW

[Gru08] R. Edward Grumbine. Conserving China’s wildlife. Conservation Biology, 22(4):1083–1084, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerber:2001:AUB

[GS01] Leah R. Gerber and Cheryl B. Schultz. Authorship and the use of biological information in endangered species recovery plans. Autoría y el uso de la información biológica en los planes de recuperación de especies en peligro. Conservation Biology, 15 (5):1308–1314, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbs:2002:EER

[GS02] James P. Gibbs and W. Gregory Shriver. Estimating the effects of road mortality on turtle populations. Estimación de los efectos de la mortalidad en carreteras sobre poblaciones de tortugas. Conservation Biology, 16(6):1647–1652, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbs:2005:TSR

[GS05] James P. Gibbs and David A. Steen. Trends in sex ratios of turtles in the United States: Implications of road mortality tendencias en las proporciones de sexos de tortugas en los Estados Unidos: Implicaciones de la mortalidad en caminos. Conservation Biology, 19(2):552–556, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guo:2008:TPM

[GS08] Qinfeng Guo and Amy Symstad. A two-part measure of degree of invasion for cross-community comparisons. Una medida de dos partes del grado de invasión para comparaciones entre comunidades. Conservation Biology, 22(3):666–672, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gray:2004:EAC

[GSB04] Matthew J. Gray, Loren M. Smith, and Roberto Brenes. Effects of agricultural cultivation on demographics of southern high plains amphibians. Efectos de cultivos agrícolas sobre la demografía de anfibios de llanuras del sur (E.U.A.). Conservation Biology, 18(5):1368–1377, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gitlin:2006:MGW

[GSB^{+}06] Alicyn R. Gitlin, Christopher M. Sthultz, Matthew A. Bowker, Stacy Stumpf, Kristina L. Paxton, Karla Kennedy, Axhel Muñoz, Joseph K. Bailey, and Thomas G. Whitham. Mortality gradients within and among dominant plant populations as barometers of ecosystem change during extreme drought. Gradientes de mortalidad dentro y entre poblaciones de plantas dominantes como barómetros de cambios en el ecosistema durante sequía extrema. Conservation Biology, 20(5):1477–1486, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gonzalez-Suarez:2008:BED

[GSG08] Manuela González-Suárez and Leah R. Gerber. A behaviorally explicit demographic model integrating habitat selection and population dynamics in California sea lions. Un modelo Demográfico conductualmente explícito que integra selección de hábitat y dinámica de poblaciones de leones marinos de California. Conservation Biology, 22(6):1608–1618, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gjerde:2004:FSD

[GSR^{+}04] Ivar Gjerde, Magne SæTersdal, Jørund Rolstad, Hans H. Blom, and Ken Olaf Storaunet. Fine-scale diversity and rarity hotspots in northern forests. Diversidad a escala fina y Áreas de rareza crítica en bosques boreales. Conservation Biology, 18(4):1032– 1042, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Golding:2003:HTC

[GT03] Janice S. Golding and Jonathan Timberlake. How taxonomists can bridge the gap between taxonomy and conservation science. Conservation Biology, 17(4):1177–1178, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guidetti:2007:PMR

[Gui07] Paolo Guidetti. Potential of marine reserves to cause community-wide changes beyond their boundaries. El potencial de reservas marinas para provocar cambios a nivel de comunidad más allá de sus límites. Conservation Biology, 21(2):540–545, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gundale:2002:IEE

[Gun02] Michael J. Gundale. Influence of exotic earthworms on the soil organic horizon and the rare fern Botrychium mormo. Influencia de lombrices exóticas en el horizonte orgánico del suelo y el helecho raro Botrychium mormo. Conservation Biology, 16(6): 1555–1561, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gustafsson:2002:PAR

[Gus02] Lena Gustafsson. Presence and abundance of red-listed plant species in Swedish forests. Presencia y abundancia de especies de plantas en bosques suecos incluídas en la lista roja. Conservation Biology, 16(2):377–388, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gutierrez:2005:EER

[Gut05] David Gutiérrez. Effectiveness of existing reserves in the longterm protection of a regionally rare butterfly. Efectividad de las reservas actuales en la protección a largo plazo de una mariposa regionalmente rara. Conservation Biology, 19(5):1586–1597, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerwing:2002:CLA

[GV02] Jeffrey J. Gerwing and Edson Vidal. Changes in liana abundance and species diversity eight years after liana cutting and logging in an Eastern Amazonian forest. Cambios en la abundancia de lianas y la diversidad de especies después de ocho años del corte de lianas y de tala en un bosque de la Amazonia oriental. Conservation Biology, 16(2):544–548, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guenette:2005:TFB

[GV05] Jean-Sébastien Guénette and Marc-André Villard. Thresholds in forest bird response to habitat alteration as quantitative targets for conservation. Umbrales en la respuesta de aves de bosque a la alteración del hábitat como objetivos cuantitativos para la conservación. Conservation Biology, 19(4):1168–1180, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibson:2009:CCC

[GVS09] Sachiko Y. Gibson, Ryan C. Van Der Marel, and Brian M. Starzomski. Climate change and conservation of leading-edge peripheral populations. Conservation Biology, 23(6):1369–1373, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffin:2007:ETB

[GVTM07] Suzanne C. Griffin, Tanguy Valois, Mark L. Taper, and L. Scott Mills. Effects of tourists on behavior and demography of Olympic marmots. Efectos del turismo sobre el comportamiento de Marmota olympus. Conservation Biology, 21(4):1070–1081, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gower:2005:CBC

[GW05] David J. Gower and Mark Wilkinson. Conservation biology of caecilian amphibians. Conservación de anfibios cecilianos. Conservation Biology, 19(1):45–55, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Graham:2007:LEI

[GWJ^{+}07] Nicholas A. J. Graham, Shaun K. Wilson, Simon Jennings, Nicholas V. C. Polunin, Jan Robinson, Jude P. Bijoux, and Tim M. Daw. Lag effects in the impacts of mass coral bleaching on coral reef fish, fisheries, and ecosystems. Efectos a largo plazo de los impactos del blanqueado masivo de corales sobre peces de arrecifes coralinos, pesquerías y ecosistemas. Conservation Biology, 21(5):1291–1300, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gustafson:2007:PSE

[GWM^{+}07] Richard G. Gustafson, Robin S. Waples, James M. Myers, Laurie A. Weitkamp, Gregory J. Bryant, Orlay W. Johnson, and Jeffrey J. Hard. Pacific salmon extinctions: Quantifying lost and remaining diversity. Extinciones de salmón del Pacífico: Cuantificación de la diversidad perdida y la remanente. Conservation Biology, 21(4):1009–1020, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbons:2006:RAB

[GWS^{+}06] J. Whitfield Gibbons, Christopher T. Winne, David E. Scott, John D. Willson, Xavier Glaudas, Kimberly M. Andrews, Brian D. Todd, Luke A. Fedewa, Lucas Wilkinson, Ria N. Tsaliagos, Steven J. Harper, Judith L. Greene, Tracey D. Tuberville, Brian S. Metts, Michael E. Dorcas, John P. Nestor, Cameron A. Young, Tom Akre, Robert N. Reed, Kurt A. Buhlmann, Jason Norman, Dean A. Croshaw, Cris Hagen, and Betsie B. Rothermel. Remarkable amphibian biomass and abundance in an isolated wetland: Implications for wetland conservation. Biomasa y abundancia de anfibios extraordinaria en un humedal aislado: Implicaciones para la conservación de humedales. Conservation Biology, 20(5):1457–1465, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerber:2007:DFA

[GWS07] Leah R. Gerber, Jeffrey Wielgus, and Enric Sala. A decision framework for the adaptive management of an exploited species with implications for marine reserves un marco de referencia para decisiones para el manejo adaptativo de una especie explotada con implicaciones para reservas marinas. Conservation Biology, 22(1):1594–1602, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grumbine:2009:CSW

[GX09] R. Edward Grumbine and Jianchu Xu. China shakes the world — and then what? Conservation Biology, 23(3):513–515, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Good:2006:ADD

[GZK06] Tatjana C. Good, Michelle L. Zjhra, and Claire Kremen. Addressing data deficiency in classifying extinction risk: a case study of a radiation of bignoniaceae from Madagascar. Enfrentando datos insuficientes en la clasificación de riesgo de extinción: un estudio de caso de una radiación de bignoniaceae de Madagascar. Conservation Biology, 33(5):1099–1110, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haag:2009:HBM

[Haa09] Wendell R. Haag. How to build a “Monster” from miscellany. Conservation Biology, 23(2):504–506, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haddad:2000:CLP

[Had00] Nick Haddad. Corridor length and patch colonization by a butterfly, Junonia coenia. Extensión de corredores y colonización de parches por una mariposa, Junonia coenia. Conservation Biology, 14(3):738–745, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haddad:2007:HCP

[Had07] Nick Haddad. Helping to create a pathway to better conservation. Conservation Biology, 21(2):890–891, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hafernik:2008:DCB

[Haf08] John E. Hafernik. Downsizing conservation biology. Conservation Biology, 22(6):1666–1667, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoeinghaus:2009:ERI

[HAG^{+}09] David J. Hoeinghaus, Angelo A. Agostinho, Luiz C. Gomes, Fernando M. Pelicice, Edson K. Okada, João D. Latini, Elaine A. L. Kashiwaqui, and Kirk O. Winemiller. Effects of river impoundment on ecosystem services of large tropical rivers: Embodied energy and market value of artisanal fisheries. Efectos de las represas en ríos tropicales sobre los servicios del ecosistema: Energía virtual y valor de mercado de las pesquerías artesanales. Conservation Biology, 23(5):1222–1231, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huhta:2004:FFI

[HAJ^{+}04] Esa Huhta, Teija Aho, Ari Jäntti, Petri Suorsa, Markku Kuitunen, Ari Nikula, and Harri Hakkarainen. Forest fragmentation increases nest predation in the Eurasian treecreeper. La fragmentación del bosque incrementa la depredación de nidos de Certhia familiaris. Conservation Biology, 18(1):148–155, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:2002:AMM

[Hal02] Charles A. S. Hall. Applying market mechanisms to ecosystem conservation. Conservation Biology, 16(3):850–852, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harding:2003:SBO

[Har03] Elaine K. Harding. Scientism as a barrier to our progress. Conservation Biology, 17(3):651, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2005:HDL

[Har05] Larry D. Harris. Human dimensions … and then some. Conservation Biology, 19(1):282–284, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:2006:WWC

[Har06a] Autumn-Lynn Harrison. Who’s who in Conservation Biology — an authorship analysis. Quien es quien en Conservation Biology — un análisis de autorías. Conservation Biology, 20(3):652–657, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartshorn:2006:TRF

[Har06b] Gary S. Hartshorn. A tropical rain forest gem. Conservation Biology, 33(5):1332–1333, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hart:2007:USN

[Har07] Gary Hart. U.S. security: the nexus of oil importation and climate change. Conservation Biology, 21(4):915, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haskell:2000:EFR

[Has00] David G. Haskell. Effects of forest roads on macroinvertebrate soil fauna of the Southern Appalachian mountains. Efectos de carreteras en bosques sobre la fauna de macroinvertebrados del suelo en las Montañas del Sur de los Apalaches. Conservation Biology, 14(1):57–63, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayes:2003:MMC

[Hay03] Grey Hayes. A must-have manual for conservation managers and researchers. Conservation Biology, 17(2):642–643, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayes:2004:DEB

[Hay04] Denis Hayes. Deflating the ecological bubble. Conservation Biology, 18(6):1461–1462, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayes:2009:FNE

[Hay09a] Mark A. Hayes. Into the field: Naturalistic education and the future of conservation. Hacia el campo: Educación naturalista y el futuro de la conservación. Conservation Biology, 23(5):1075– 1079, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayward:2009:NRP

[Hay09b] Matt W. Hayward. The need to rationalize and prioritize threatening processes used to determine threat status in the IUCN red list la necesidad de racionalizar y priorizar los procesos amenazantes utilizados para determinar el estatus de amenaza en la lista roja de la UICN. Conservation Biology, 23(6):1568–1576, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holdredge:2009:RCH

[HBA09] Christine Holdredge, Mark D. Bertness, and Andrew H. Altieri. Role of crab herbivory in die-off of New England salt marshes. Papel de la herbivoría de cangrejos en la declinación de marismas de Nueva Inglaterra. Conservation Biology, 23(3):672–679, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hellmann:2008:FPC

[HBBD08] Jessica J. Hellmann, James E. Byers, Britta G. Bierwagen, and Jeffrey S. Dukes. Five potential consequences of climate change for invasive species. Cinco consecuencias potenciales del cambio climático para especies invasoras. Conservation Biology, 22(3): 534–543, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haig:2006:TCL

[HBC^{+}06] Susan M. Haig, Erik A. Beever, Steven M. Chambers, Hope M. Draheim, Bruce D. Dugger, Susie Dunham, Elise Elliott-Smith, Joseph B. Fontaine, Dylan C. Kesler, Brian J. Knaus, Iara F. Lopes, Pete Loschl, Thomas D. Mullins, and Lisa M. Sheffield. Taxonomic considerations in listing subspecies under the U.S. Endangered Species Act. Consideraciones taxonómicas para enlistar subespecies en el Acta de Especies en Peligro de E.U.A. Conservation Biology, 20(6):1584–1594, December 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanson:2009:WBH

[HBD^{+}09] Thor Hanson, Thomas M. Brooks, Gustavo A. B. Da Fonseca, Michael Hoffmann, John F. Lamoreux, Gary Machlis, Cristina G. Mittermeier, Russell A. Mittermeier, and John D. Pilgrim. Warfare in biodiversity hotspots guerra en sitios de importancia para la biodiversidad. Conservation Biology, 23(3): 578–587, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higgins:2005:FCA

[HBKF05] Jonathan V. Higgins, Mark T. Bryer, Mary L. Khoury, and Thomas W. Fitzhugh. A freshwater classification approach for biodiversity conservation planning. Un método de clasificación de agua dulce para planificación de conservación de biodiversidad. Conservation Biology, 19(2):432–445, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2009:TMB

[HBLW09] Malcolm L. Hunter, Jr., Michael J. Bean, David B. Lindenmayer, and David S. Wilcove. Thresholds and the mismatch between environmental laws and ecosystems. Conservation Biology, 23 (4):1053–1055, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hooftman:2004:GEH

[HBSD04] Danny A. P. Hooftman, Regula C. Billeter, Bernhard Schmid, and Matthias Diemer. Genetic effects of habitat fragmentation on common species of Swiss fen meadows. Efectos genéticos de la fragmentación de hábitat en especies comunes en ciénegas de Suiza. Conservation Biology, 18(4):1043–1051, August 2004.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hobson:2002:LSC

[HBV02] Keith A. Hobson, Erin M. Bayne, and Steve L. Van Wilgenburg. Large-scale conversion of forest to agriculture in the boreal plains of Saskatchewan. Conversión a gran escala del bosque a agricultura en las llanuras boreales de Saskatchewan. Conservation Biology, 16(6):1530–1541, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hockey:2009:UBB

[HC09] Philip A. R. Hockey and Odette E. Curtis. Use of basic biological information for rapid prediction of the response of species to habitat loss. Uso de información biológica básica para la Predicción rápida de la respuesta de especies a la pérdida de hábitat. Conservation Biology, 23(1):64–71, February 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamilton:2000:CRP

[HCBK00] A. Hamilton, A. Cunningham, D. Byarugaba, and F. Kayanja. Conservation in a region of political instability: Bwindi Impenetrable Forest, Uganda. Conservación en una Región de inestabilidad política: El bosque impenetrable bwindi en Uganda. Conservation Biology, 14(6):1722–1725, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harding:2001:SFH

[HCE^{+}01] Elaine K. Harding, Elizabeth E. Crone, Bret D. Elderd, Jonathan M. Hoekstra, Alexa J. McKerrow, John D. Perrine, Jim Regetz, Leslie J. Rissler, Amanda G. Stanley, Eric L. Walters, and NCEAS Habitat Conservation Plan Working Group. The scientific foundations of habitat conservation plans: a quantitative assessment. Las bases científicas de los planes de conservación del hábitat: una evaluación cuantitativa. Conservation Biology, 15(2):488–500, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haight:2002:OHP

[HCK^{+}02] Robert G. Haight, Brian Cypher, Patrick A. Kelly, Scott Phillips, Hugh P. Possingham, Katherine Ralls, Anthony M. Starfield, P. J. White, and Daniel Williams. Optimizing habitat protection using demographic models of population viability.

Optimización de la protección de hábitat utilizando modelos de viabilidad poblacional. Conservation Biology, 16(5):1386–1397, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hughey:2003:IEP

[HCM03] Kenneth F. D. Hughey, Ross Cullen, and Emma Moran. Integrating economics into priority setting and evaluation in conservation management. Integrando la economía a la definición de prioridades y la evaluación de la gestión de la conservación. Conservation Biology, 17(1):93–103, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heleno:2009:EAP

[HCRM09] Rúben H. Heleno, Ricardo S. Ceia, Jaime A. Ramos, and Jane Memmott. Effects of alien plants on insect abundance and biomass: a food-Web approach. Efectos de plantas exóticas sobre la abundancia y biomasa de insectos: un enfoque de red alimenticia. Conservation Biology, 23(2):410–419, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoctor:2000:ILR

[HCZ00] Thomas S. Hoctor, Margaret H. Carr, and Paul D. Zwick. Identifying a linked reserve system using a regional landscape approach: the Florida ecological network. Identificación de un sistema de reservas conectado usando una estrategia de paisaje: la red ecológica de Florida. Conservation Biology, 14(4):984– 1000, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harding:2001:EEP

[HDA01] Elaine K. Harding, Daniel F. Doak, and Joy D. Albertson. Evaluating the effectiveness of predator control: the non-native red fox as a case study. Evaluación de la efectividad del control de depredadores: el zorro rojo no nativo como estudio de caso. Conservation Biology, 15(4):1114–1122, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Horner-Devine:2003:CBT

[HDDEB03] M. Claire Horner-Devine, Gretchen C. Daily, Paul R. Ehrlich, and Carol L. Boggs. Countryside biogeography of tropical butterflies. Biogeografía de paisaje de mariposas tropicales. Conservation Biology, 17(1):168–177, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hughes:2000:CID

[HDE00] Jennifer B. Hughes, Gretchen C. Daily, and Paul R. Ehrlich. Conservation of insect diversity: a habitat approach. Conservación de la diversidad de insectos: una estrategia a nivel de hábitat. Conservation Biology, 14(6):1788–1797, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Howald:2007:IRE

[HDG^{+}07] Gregg Howald, C. Josh Donlan, Juan Pablo Galván, James C. Russell, John Parkes, Araceli Samaniego, Yiwei Wang, Dick Veitch, Piero Genovesi, Michel Pascal, Alan Saunders, and Bernie Tershy. Invasive rodent eradication on islands. Erradicación de roedores invasores de islas. Conservation Biology, 21 (5):1258–1268, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Howard:2000:PAP

[HDK^{+}00] P. C. Howard, T. R. B. Davenport, F. W. Kigenyi, P. Viskanic, M. C. Baltzer, C. J. Dickinson, J. Lwanga, R. A. Matthews, and E. Mupada. Protected area planning in the tropics: Uganda’s national system of forest nature reserves. Planeación de areas protegidas en el Trópico: Sistema nacional de reservas forestales naturales de Uganda. Conservation Biology, 14(3):858– 875, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Herrerias-Diego:2006:EFF

[HDQSL06] Yvonne Herrerías-Diego, Mauricio Quesada, Kathryn E. Stoner, and Jorge A. Lobo. Effects of forest fragmentation on phenological patterns and reproductive success of the tropical dry forest tree Ceiba aesculifolia. Efectos de la fragmentación de patrones fenológicos y Éxito reproductivo del Árbol de bosque tropical seco Ceiba aesculifolia. Conservation Biology, 33(5):1111–1120, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:2001:IC

[Hei01] Joel T. Heinen. The importance of context. Conservation Biology, 15(4):1194–1195, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:2003:CPP

[Hei03] Joel T. Heinen. Clarifying the policy process. Conservation Biology, 17(6):1857–1858, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:2009:SLL

[Hei09] Joel Heinen. Singular lessons for the lesser known. Conservation Biology, 23(1):237–238, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Helgen:2008:NGT

[Hel08] Kristofer M. Helgen. New Guinea as a tropical lens. Conservation Biology, 22(6):1667–1668, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Henny:2004:ASS

[Hen04] Charles J. Henny. An avian success story. Conservation Biology, 18(6):1697–1699, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Henderson:2005:EC

[Hen05] David Henderson. Evangelicals are conservationists. Conservation Biology, 19(6):1687–1688, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Herkert:2009:RBP

[Her09] James R. Herkert. Response of bird populations to farmland set-aside programs. Respuesta de las poblaciones de aves a los programas de terrenos agrícolas reservados. Conservation Biology, 23(4):1036–1040, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hettinger:2004:EET

[Het04] Ned Hettinger. Environmental ethics today. Conservation Biology, 18(2):587–589, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hewitt:2004:UDA

[Hew04] David G. Hewitt. Understanding the distribution and abundance of species. Conservation Biology, 18(4):1168–1171, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Houlahan:2004:EIP

[HF04] Jeff E. Houlahan and C. Scott Findlay. Effect of invasive plant species on temperate wetland plant diversity. Efecto de especies de plantas invasoras sobre la diversidad de plantas de humedales templados. Conservation Biology, 18(4):1132–1138, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holdsworth:2007:EEI

[HFR07] Andrew R. Holdsworth, Lee E. Frelich, and Peter B. Reich. Effects of earthworm invasion on plant species richness in northern hardwood forests. Efectos de la invasión de lombrices de tierra sobre la riqueza de especies de plantas en bosques deciduos del norte. Conservation Biology, 21(4):997–1008, August 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hijmans:2000:AGR

[HGH^{+}00] R. J. Hijmans, K. A. Garrett, Z. Huamán, D. P. Zhang, M. Schreuder, and M. Bonierbale. Assessing the geographic representativeness of Genebank collections: the case of Bolivian wild potatoes. Evaluación de la representatividad geográfica de las colecciones de bancos de genes: El caso de las papas silvestres bolivianas. Conservation Biology, 14(6):1755–1765, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hughes:2003:TCE

[HGH^{+}03] Jane Hughes, Katrina Goudkamp, David Hurwood, Martyn Hancock, and Stuart Bunn. Translocation causes extinction of a local population of the freshwater shrimp Paratya australiensis. La translocación causa la extinción de una población local de camarón de agua dulce Paratya australiensis. Conservation Biology, 17(4):1007–1012, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamer:2000:SDE

[HH00] K. C. Hamer and J. K. Hill. Scale-dependent effects of habitat disturbance on species richness in tropical forests. Efectos escaladependientes de la perturbación del hábitat sobre la riqueza de especies en bosques tropicales. Conservation Biology, 14(5): 1435–1440, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayes:2003:CGI

[HH03] Grey F. Hayes and Karen D. Holl. Cattle grazing impacts on annual forbs and vegetation composition of mesic grasslands in California. Impactos del pastoreo de ganado sobre hierbas anuales y la composición vegetal de pastizales Mésicos en California. Conservation Biology, 17(6):1694–1702, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hagen:2006:RCH

[HH06a] Amy N. Hagen and Karen E. Hodges. Resolving critical habitat designation failures: Reconciling law, policy, and biology. Resolución de fallas en la Designación de hábitats Críticos: Reconciliación de la ley, la política y la biología. Conservation Biology, 20(2):399–407, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holl:2006:CIM

[HH06b] Karen D. Holl and Grey F. Hayes. Challenges to introducing and managing disturbance regimes for Holocarpha macradenia, an endangered annual grassland forb. Retos de la introducción y gestión de regímenes de perturbación para Holocarpha macradenia, una hierba anual en peligro. Conservation Biology, 33(5): 1121–1131, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hershner:2008:MIA

[HH08] Carl Hershner and Kirk J. Havens. Managing invasive aquatic plants in a changing system: Strategic consideration of ecosystem services. Manejo de plantas acuáticas invasoras en un sistema cambiante: Consideración Estratégica de los servicios del ecosistema. Conservation Biology, 22(3):544–550, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hershner:2009:ESM

[HH09] Carl Hershner and Kirk Havens. Ecosystem services and management of invasive species in a changing system: Response to Martin and Blossey. Conservation Biology, 23(2):497–498, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haddad:2008:DOP

[HHD^{+}08] Nick M. Haddad, Brian Hudgens, Chris Damiani, Kevin Gross, Daniel Kuefler, and Ken Pollock. Determining optimal population monitoring for rare butterflies. Determinación del monitoreo poblacional Óptimo para mariposas raras. Conservation Biology, 22(4):929–940, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hee:2000:RPS

[HHSC00] Jason J. Hee, David A. Holway, Andrew V. Suarez, and Ted J. Case. Role of propagule size in the success of incipient colonies of the invasive Argentine ant. El tamaño de propágulos como factor en el Éxito de colonias incipientes de la hormiga invasora Argentina. Conservation Biology, 14(2):559–563, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hietz:2005:CVE

[Hie05] Peter Hietz. Conservation of vascular epiphyte diversity in Mexican coffee plantations. Conservación de la diversidad de epifitas vasculares en plantaciones mexicanas de café. Conservation Biology, 19(2):391–399, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hilborn:2003:CC

[Hil03] Ray Hilborn. A contradiction in conservation. Conservation Biology, 21(6):340–341, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:2003:EHE

[HIS03] S. Harrison, B. D. Inouye, and H. D. Safford. Ecological heterogeneity in the effects of grazing and fire on grassland diversity. Heterogeneidad ecológica en los efectos del pastoreo y los incendios sobre la diversidad del pastizal. Conservation Biology, 17(3):837–845, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hitt:2007:AEC

[Hit07] Mary Anne Hitt. America at the energy crossroads. Conservation Biology, 21(2):293–294, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoyle:2005:GWH

[HJ05] Martin Hoyle and Mike James. Global warming, human population pressure, and viability of the world’s smallest butterfly. Calentamiento global, presión de la población humana y viabilidad de la mariposa más pequeña del mundo. Conservation Biology, 19(4):1113–1124, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Honnay:2007:SCR

[HJ07] Olivier Honnay and Hans Jacquemyn. Susceptibility of common and rare plant species to the genetic consequences of habitat fragmentation. Susceptibilidad de especies de plantas comunes y raras a las consecuencias genéticas de la fragmentación del hábitat. Conservation Biology, 21(2):823–831, June 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2005:RBC

[HJP05] Grant M. Harris, Clinton N. Jenkins, and Stuart L. Pimm. Refining biodiversity conservation priorities. Refinación de las prioridades de conservación de la biodiversidad. Conservation Biology, 19(6):1957–1968, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hendrix:2000:PSR

[HK00] Stephen D. Hendrix and John F. Kyhl. Population size and reproduction in Phlox pilosa. Tamaño poblacional y reproducción de Phlox pilosa. Conservation Biology, 14(1):304–313, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartley:2003:SDR

[HK03] Stephen Hartley and William E. Kunin. Scale dependency of rarity, extinction risk, and conservation priority. Dependencia de la escala de la rareza, riesgo de extinción y prioridad de conservación. Conservation Biology, 17(6):1559–1570, December 2003.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harvey:2008:IAL

[HKC^{+}08] Celia A. Harvey, Oliver Komar, Robin Chazdon, Bruce G. Ferguson, Bryan Finegan, Daniel M. Griffith, Miguel Martínez-Ramos, Helda Morales, Ronald Nigh, Lorena Soto-Pinto, Michiel Van Breugel, and Mark Wishnie. Integrating agricultural landscapes with biodiversity conservation in the Mesoamerican hotspot. Conservation Biology, 22(1):8–15, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hyvarinen:2006:FGT

[HKM06] Esko Hyvärinen, Jari Kouki, and Petri Martikainen. Fire and green-tree retention in conservation of red-listed and rare deadwood-dependent beetles in Finnish boreal forests. Fuego y retención de Árboles verdes en la conservación de escarabajos en la lista roja y raros y dependientes de madera muerta en bosques boreales finlandeses. Conservation Biology, 20(6): 1710–1719, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heliola:2001:DCB

[HKN01] Janne Heliölä, Matti Koivula, and Jari Niemelä. Distribution of carabid beetles (Coleoptera, Carabidae) across a boreal forest–clearcut ecotone. Distribución de escarabajos carábidos (Coleoptera, Carabidae) a lo largo de un ecotono de bosque boreal–zona talada. Conservation Biology, 15(2):370–377, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higham:2007:UNE

[HL07a] J. E. S. Higham and D. Lusseau. Urgent need for empirical research into whaling and whale watching. Conservation Biology, 21(2):554–558, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hitch:2007:BDN

[HL07b] Alan T. Hitch and Paul L. Leberg. Breeding distributions of North American bird species moving north as a result of climate change. Distribución reproductiva de especies de aves Norte Americanas en movimiento hacia el norte como resultado del cambio climático. Conservation Biology, 21(2):534–539, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hiers:2003:SSM

[HLB^{+}03] J. Kevin Hiers, Stephen C. Laine, J. J. Bachant, James H. Furman, Weldon W. Greene, Jr., and Vernon Compton. Simple spatial modeling tool for prioritizing prescribed burning activities at the landscape scale herramienta simple de modelado espacial para priorizar actividades de quemas prescritas al nivel de paisaje. Conservation Biology, 17(6):1571–1578, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hendry:2003:ENS

[HLG03] Andrew P. Hendry, Benjamin H. Letcher, and Gabe Gries. Estimating natural selection acting on stream-dwelling Atlantic salmon: Implications for the restoration of extirpated populations. Estimación de la acción de la selección natural sobre el salmón del Atlántico en arroyos: Implicaciones para la restauración de poblaciones extirpadas. Conservation Biology, 17(3): 795–805, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hortal:2007:LBD

[HLJV07] Joaquín Hortal, Jorge M. Lobo, and Alberto Jiménez-Valverde. Limitations of biodiversity databases: Case study on seed-plant diversity in Tenerife, Canary Islands. Limitaciones de las bases de datos de biodiversidad: Estudio de caso de la diversidad de plantas con semillas en Tenerife, Islas Canarias. Conservation Biology, 21(2):853–863, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haydon:2002:IEP

[HLSZ02] D. T. Haydon, M. K. Laurenson, and C. Sillero-Zubiri. Integrating epidemiology into population viability analysis: Managing the risk posed by rabies and canine distemper to the Ethiopian wolf. Integración de la Epidemiología dentro del análisis de viabilidad poblacional: Manejo del riesgo que representan la rabia y el moquillo canino en el lobo Etíope. Conservation Biology, 16(5):1372–1385, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hackney:2001:EDA

[HM01] Erin E. Hackney and James B. McGraw. Experimental demonstration of an Allee effect in American ginseng. Demostración experimental de un efecto de Allee en el ginseng americano. Conservation Biology, 15(1):129–136, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hersch-Martinez:2002:CA

[HM02] Paul Hersch-Martínez. A comprehensive approach. Conservation Biology, 16(6):1668–1669, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hilty:2003:SBP

[HM03] Jodi Hilty and Adina M. Merenlender. Studying biodiversity on private lands estudio de la biodiversidad en tierras privadas. Conservation Biology, 17(1):132–137, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hilty:2004:URC

[HM04] Jodi A. Hilty and Adina M. Merenlender. Use of riparian corridors and vineyards by mammalian predators in Northern California. Utilización de corredores riparios y Viñedos por mamíferos depredadores en el norte de California. Conservation Biology, 18(1):126–135, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halsing:2008:CEM

[HM08] David L. Halsing and Michael R. Moore. Cost-effective management alternatives for Snake River Chinook salmon: a biologicaleconomic synthesis. Alternativas rentables para el manejo de salmón en el Río Snake: una síntesis biológica–económica. Conservation Biology, 22(2):338–350, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halsing:2009:CEM

[HM09] David L. Halsing and Michael R. Moore. Cost-effective management of Snake River Chinook salmon: Response to Wilson et al. Conservation Biology, 23(2):479–481, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harper:2005:EIF

[HMB^{+}05] Karen A. Harper, S. Ellen Macdonald, Philip J. Burton, Jiquan Chen, Kimberley D. Brosofske, Sari C. Saunders, Eugénie S. Euskirchen, Dar Roberts, Malanding S. Jaiteh, and Per-Anders Esseen. Edge influence on forest structure and composition in fragmented landscapes. Influencia del borde sobre la composición y estructura de bosques en paisajes fragmentados. Conservation Biology, 19(3):768–782, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hammond:2001:PRS

[HMF^{+}01] Robert L. Hammond, William Macasero, Benito Flores, Osama B. Mohammed, Tim Wacher, and Michael W. Bruford. Phylogenetic reanalysis of the Saudi gazelle and its implications for conservation. Reanálisis filogenético de la gacela Saudi y las implicaciones para su conservación. Conservation Biology, 15(4): 1123–1133, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hill:2003:HRC

[HMF03] Kim Hill, Garnett McMillan, and Rosalia Fariña. Huntingrelated changes in game encounter rates from 1994 to 2001 in the Mbaracayu Reserve, Paraguay. Cambios relacionados con la cacería en las tasas de encuentro de especies cinegéticas de 1994 a 2001 en la Reserva Mbaracayu, Paraguay. Conservation Biology, 17(5):1312–1323, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haig:2004:GIS

[HMF^{+}04] Susan M. Haig, Thomas D. Mullins, Eric D. Forsman, Pepper W. Trail, and Liv Wennerberg. Genetic identification of spotted owls, barred owls, and their hybrids: Legal implications of hybrid identity. Identificación genética de Strix occidentalis caurina y S. varia y sus Híbridos: Implicaciones legales de la identidad Híbrida. Conservation Biology, 18(5):1347–1357, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:2000:RTF

[HMH00] Susan Harrison, John Maron, and Gary Huxel. Regional turnover and fluctuation in populations of five plants confined to serpentine seeps. Productividad regional y Fluctuación en poblaciones de cinco plantas confinadas a suelos serpentinos. Conservation Biology, 14(3):769–779, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hurt:2003:EEI

[HMH03] C. R. Hurt, S. Meredith, and P. Hamilton. Evolutionary ecology: Integrating theory, technology, and ecology. Conservation Biology, 21(6):342–343, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halme:2009:QIP

[HMK^{+}09] Panu Halme, Mikko Mönkkönen, Janne S. Kotiaho, Anna-Liisa Ylisirniö, and Anni Markkanen. Quantifying the indicator power of an indicator species. Cuantificación del poder indicador de una especie indicadora. Conservation Biology, 23(4):1008–1016, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hannah:2002:CBC

[HML^{+}02] L. Hannah, G. F. Midgley, T. Lovejoy, W. J. Bond, M. Bush, J. C. Lovett, D. Scott, and F. I. Woodward. Conservation of biodiversity in a changing climate. Conservation Biology, 16(1): 264–268, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hylander:2009:CRH

[HN09] Kristoffer Hylander and Sileshi Nemomissa. Complementary roles of home gardens and exotic tree plantations as alternative habitats for plants of the Ethiopian montane rainforest roles. Complementarios de jardines domésticos y plantaciones de Árboles Exóticos como hábitats alternativos para plantas del bosque lluvioso montano Etíope. Conservation Biology, 23(2): 400–409, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hylander:2004:EBS

[HNG04] Kristoffer Hylander, Christer Nilsson, and Tove Göthner. Effects of buffer-strip retention and clearcutting on land snails in boreal riparian forests. Efectos de la retención de una franja de amortiguamiento y de la tala sobre caracoles terrestres en bosques boreales riparios. Conservation Biology, 18(4):1052–1062, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanski:2002:EDE

[HO02] Ilkka Hanski and Otso Ovaskainen. Extinction debt at extinction threshold. Deuda de extinción y umbral de extinción. Conservation Biology, 16(3):666–673, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hobbs:2001:SAH

[Hob01] Richard J. Hobbs. Synergisms among habitat fragmentation, livestock grazing, and biotic invasions in Southwestern Australia. Sinergismos entre la fragmentacion del habitat, el pastoreo de ganado y las invasiones bióticas en el suroeste de Australia. Conservation Biology, 15(6):1522–1528, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoddle:2004:RBU

[Hod04a] Mark S. Hoddle. Restoring balance: Using exotic species to control invasive exotic species. Restaurando el balance: Utilización de Exóticos para controlar exóticas invasoras. Conservation Biology, 18(1):38–49, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoddle:2004:SBC

[Hod04b] Mark S. Hoddle. The strength of biological control in the battle against invasive pests: a reply. Conservation Biology, 18(1):61– 65, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [APSB02, LS04].

Hanson:2009:OFR

[HODB09] Chad T. Hanson, Dennis C. Odion, Dominick A. Dellasala, and William L. Baker. Overestimation of fire risk in the northern spotted owl recovery plan. Sobreestimación del riesgo de fuego en el plan de recuperación de Strix occidentalis caurina. Conservation Biology, 23(5):1314–1319, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holzman:2000:TC

[Hol00] Barbara A. Holzman. A tool for change. Conservation Biology, 14(4):1213–1214, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmes:2001:NSL

[Hol01] Christopher M. Holmes. Navigating the socioecological landscape. Conservation Biology, 15(5):1466–1467, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmes:2002:CC

[Hol02] Elizabeth E. Holmes. Compute and conserve. Conservation Biology, 16(1):275–276, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holl:2008:TBB

[Hol08] Karen D. Holl. Are there benefits of bat roosts for tropical forest restoration? Conservation Biology, 22(5):1090, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hood:2007:SPP

[Hoo07] W. Gregory Hood. Saving the planet in the Pacific northwest. Conservation Biology, 21(5):1379–1380, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hopkins:2007:EIS

[Hop07] John Hopkins. Exploring incentives in species protection. Conservation Biology, 22(1):1656–1657, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hourdequin:2000:I

[Hou00] Marion Hourdequin. Introduction. Conservation Biology, 14 (1):16–17, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hecht:2001:IPR

[HP01] Anne Hecht and Mary J. Parkin. Improving peer review of listings and recovery plans under the Endangered Species Act. Mejoras a la revisiones colegiadas de los planes para el listado y recuperación del Acta de Especies en Peligro. Conservation Biology, 15(5):1269–1273, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2004:BST

[HP04] Grant M. Harris and Stuart L. Pimm. Bird species’ tolerance of secondary forest habitats and its effects on extinction. Tolerancia de especies de aves a hábitats de bosque secundario y sus efectos sobre extinción. Conservation Biology, 18(6):1607–1616, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Humphrey:2006:WK

[HP06] Stephen R. Humphrey and Ignacio Porzecanski. Whole knowledge. Conservation Biology, 21(3):926–927, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2008:RSE

[HP08] Grant Harris and Stuart L. Pimm. Range size and extinction risk in forest birds. Amplitud de rango de distribución y riesgo de extinción en aves de bosque. Conservation Biology, 22(1): 163–171, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halpern:2006:GMB

[HPF^{+}06] Benjamin S. Halpern, Christopher R. Pyke, Helen E. Fox, J. Chris Haney, Martin A. Schlaepfer, and Patricia Zaradic. Gaps and mismatches between global conservation priorities and spending. Brechas e incongruencias entre prioridades y gasto de conservación global. Conservation Biology, 20(1):56–64, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holland:2009:CNA

[HPG09] Tim G. Holland, Garry D. Peterson, and Andrew Gonzalez. A cross-national analysis of how economic inequality predicts biodiversity loss. Análisis transnacional de Cómo la inequidad económica predice la pérdida de biodiversidad. Conservation Biology, 23(5):1304–1313, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higuchi:2004:URT

[HPK^{+}04] Hiroyoshi Higuchi, Johanna P. Pierre, Vladimir Krever, Vladimir Andronov, Go Fujita, Kiyoaki Ozaki, Oleg Goroshko, Mutsuyuki Ueta, Sergei Smirensky, and Nagahisa Mita. Using a remote technology in conservation: Satellite tracking whitenaped cranes in Russia and Asia. Utilización de una Tecnología remota en conservación: Rastreo por Satélite de grullas de nuca blanca en Rusia y China. Conservation Biology, 18(1):136–147, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haitao:2007:FET

[HPLTH07] Shi Haitao, James F. Parham, Michael Lau, and Chen Tien-Hsi. Farming endangered turtles to extinction in China. Conservation Biology, 21(1):5–6, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hansen:2002:BFL

[HR02] Andrew J. Hansen and Jay J. Rotella. Biophysical factors, land use, and species viability in and around nature reserves. Factores biofísicos, uso de suelo y viabilidad de especies en reservas naturales y sus alrededores. Conservation Biology, 16(4):1112– 1122, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hawbaker:2004:RLP

[HR04a] Todd J. Hawbaker and Volker C. Radeloff. Roads and landscape pattern in Northern Wisconsin based on a comparison of four road data sources. Caminos y patrón del paisaje en el norte de Wisconsin con base en una comparación de cuatro fuentes de datos de caminos. Conservation Biology, 18(5):1233–1244, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hawkins:2004:EAF

[HR04b] Julie P. Hawkins and Callum M. Roberts. Effects of artisanal fishing on Caribbean coral reefs. Efectos de la pesca artesanal sobre arrecifes de coral en el Caribe. Conservation Biology, 18 (1):215–226, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higgins:2004:RTF

[HR04c] Mark A. Higgins and Kalle Ruokolainen. Rapid tropical forest inventory: a comparison of techniques based on inventory data from Western Amazonia. Inventario forestal tropical rápido: Comparación de técnicas utilizando datos del inventario de la Amazonía occidental. Conservation Biology, 18(3):799– 811, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haddad:2000:ESC

[HRN00] Nick M. Haddad, Daniel K. Rosenberg, and Barry R. Noon. On experimentation and the study of corridors: Response to Beier and Noss. Conservation Biology, 14(5):1543–1545, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higgins:2004:BNS

[HRP^{+}04] Jonathan V. Higgins, Taylor H. Ricketts, Jeffrey D. Parrish, Eric Dinerstein, George Powell, Suzanne Palminteri, Jonathan M. Hoekstra, John Morrison, Adam Tomasek, and Jonathan Adams. Beyond Noah: Saving species is not enough. Conservation Biology, 18(6):1672–1673, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haight:2000:DST

[HRS00] Robert G. Haight, Katherine Ralls, and Anthony M. Starfield. Designing species translocation strategies when population growth and future funding are uncertain diseño de estrategias de traslado de especies cuando el crecimiento poblacional y el financiamiento a futuro son inciertos. Conservation Biology, 14(5): 1298–1307, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Herborg:2007:PRC

[HRS^{+}07] Leif-Matthias Herborg, Deborah A. Rudnick, Yang Siliang, David M. Lodge, and Hugh J. Macisaac. Predicting the range of Chinese mitten crabs in Europe. Predicción de la distribución de Eriocheir sinensis en Europa. Conservation Biology, 21(5): 1316–1323, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harper:2008:DCT

[HRS08] Elizabeth B. Harper, Tracy A. G. Rittenhouse, and Raymond D. Semlitsch. Demographic consequences of terrestrial habitat loss for pool-breeding amphibians: Predicting extinction risks associated with inadequate size of buffer zones. Consecuencias demográficas de la pérdida de hábitat terrestre para anfibios que se reproducen en charcas: Predicción de los riesgos de extinción asociados con el tamaño inadecuado de las zonas de amortiguamiento. Conservation Biology, 22(5):1205–1215, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Herkert:2003:EPF

[HRW^{+}03] James R. Herkert, Dan L. Reinking, David A. Wiedenfeld, Maiken Winter, John L. Zimmerman, William E. Jensen, Elmer J. Finck, Rolf R. Koford, Donald H. Wolfe, Steve K. Sherrod, M. Alan Jenkins, John Faaborg, and Scott K. Robinson. Effects of prairie fragmentation on the nest success of breeding birds in the midcontinental United States. Efectos de la fragmentación de praderas sobre el Éxito de nidos de aves en el centro de los Estados Unidos. Conservation Biology, 17(2):587– 594, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hill:2006:CBG

[HS06] Ryan Hill and Cyrie Sendashonga. Conservation biology, genetically modified organisms, and the Biosafety Protocol. Biología de la conservación, organismos genéticamente modificados y el Protocolo de Bioseguridad. Conservation Biology, 20 (6):1620–1625, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hokit:2001:CTT

[HSB01] D. Grant Hokit, Bradley M. Stith, and Lyn C. Branch. Comparison of two types of metapopulation models in real and artificial landscapes. Comparación de dos tipos de modelos de metapoblaciones en paisajes reales y artificiales. Conservation Biology, 15 (4):1102–1113, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hancock:2001:IRT

[HSDS01] Christine Hancock, Annalisa Storey, James Downing, and Susan May Szewczak. Interagency resource teams: a model for collaborative approaches to environmental education. Equipos de recursos interagenciales: un modelo para aproximaciones de colaboración para la educación ambiental. Conservation Biology, 15(3):596–602, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halpern:2007:ERV

[HSMK07] Benjamin S. Halpern, Kimberly A. Selkoe, Fiorenza Micheli, and Carrie V. Kappel. Evaluating and ranking the vulnerability of global marine ecosystems to anthropogenic threats. Evaluación y clasificación de la vulnerabilidad a las amenazas antropogénicas de los ecosistemas marinos globales. Conservation Biology, 21 (5):1301–1315, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haight:2005:MOS

[HSR05] Robert G. Haight, Stephanie A. Snyder, and Charles S. Revelle. Metropolitan open-space protection with uncertain site availability. Protección de espacios abiertos metropolitanos con disponibilidad de sitios incierta. Conservation Biology, 19(2): 327–337, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:2005:SPW

[HTM^{+}05] Mônica B. Harris, Walfrido Tomas, Guilherme Mourão, Carolina J. Da Silva, Erika Guimarães, Fátima Sonoda, and Eliani Fachim. Safeguarding the pantanal wetlands: Threats and conservation initiatives. Protegiendo a los humedales del pantanal: Amenazas e iniciativas de conservación. Conservation Biology, 19(3):714–720, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higgins:2006:WNS

[HTU^{+}06] Jonathan V. Higgins, Jerome L. Touval, Robert S. Unnasch, Steffen Reichle, David C. Oren, William R. Waldman, and Jonathan M. Hoekstra. Who needs to spend money on conservation science anyway? Conservation Biology, 20(6):1566–1567, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2000:RNC

[Hun00] Malcolm L. Hunter, Jr. Refining normative concepts in conservation. Conservation Biology, 14(2):573–574, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2005:MCS

[Hun05] Malcolm L. Hunter, Jr. A mesofilter conservation strategy to complement fine and coarse filters. Una estrategia de conservación de mesofiltro para complementar filtros finos y gruesos. Conservation Biology, 19(4):1025–1029, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2007:DLH

[Hun07a] David Hunter. Don’t leave home without it. Conservation Biology, 21(2):892–893, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2007:CCM

[Hun07b] Malcolm L. Hunter, Jr. Climate change and moving species: Furthering the debate on assisted colonization. Conservation Biology, 21(5):1356–1358, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hutto:2006:TMS

[Hut06] Richard L. Hutto. Toward meaningful snag-management guidelines for postfire salvage logging in North American conifer forests. Hacia directrices significativas para la gestión de raigones en la cosecha de salvamento en bosques de Coníferas de Norte América. Conservation Biology, 20(4):984–993, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hutchins:2008:ARC

[Hut08] Michael Hutchins. Animal rights and conservation. Conservation Biology, 22(4):815–816, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hillers:2008:EFF

[HVR08] Annika Hillers, Michael Veith, and Mark-Oliver Rödel. Effects of forest fragmentation and habitat degradation on West African leaf-litter frogs. Efectos de la fragmentación de bosques y la degradación del hábitat sobre ranas de la hojarasca en el oeste de África. Conservation Biology, 22(3):762–772, June 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2002:ETS

[HW02] Malcolm L. Hunter, Jr. and Sara L. Webb. Enlisting taxonomists to survey poorly known taxa for biodiversity conservation: a lichen case study. Reclutamiento de taxónomos para la Prospección de taxones poco conocidos para la conservación de la biodiversidad: Estudio de caso de un líquen. Conservation Biology, 16(3):660–665, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hedgren:2008:DRS

[HW08] Olof Hedgren and Jan Weslien. Detecting rare species with random or subjective sampling: a case study of red-listed saproxylic beetles in boreal Sweden. Detección de especies raras mediante muestreo aleatorio o subjetivo: un estudio de caso de escarabajos saproxílicos en la lista roja en Suecia boreal. Conservation Biology, 22(1):212–215, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hemstrom:2002:SSV

[HWH^{+}02] Miles A. Hemstrom, Michael J. Wisdom, Wendel J. Hann, Mary M. Rowland, Barbara C. Wales, and Rebecca A. Gravenmier. Sagebrush-steppe vegetation dynamics and restoration potential in the interior Columbia Basin, U.S.A. dinámica de la vegetación y restauración potencial de una estepa de artemisa en la cuenca Columbia interior, E.U.A. Conservation Biology, 16(5):1243–1255, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmes:2003:UAS

[HY03] E. E. Holmes and A. E. York. Using age structure to detect impacts on threatened populations: a case study with Steller sea lions. Utilización de la estructura de edades para detectar impactos en poblaciones amenazadas: Un estudio de caso con lobo marinos de Steller. Conservation Biology, 17(6):1794–1806, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Inouye:2003:CSP

[IB03] David W. Inouye and Carol Brewer. A case study of the program in sustainable development and conservation biology at the University of Maryland. Conservation Biology, 17(5):1204– 1208, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ibisch:2008:MWC

[Ibi08] Pierre L. Ibisch. Messages from a worst-case scenario: Mata atlantica in northeastern Brazil. Conservation Biology, 22(4):

1078–1079, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Inouye:2000:CAP

[ID00] David W. Inouye and James M. Dietz. Creating academically and practically trained graduate students. Conservation Biology, 14(3):595–596, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Irwin:2002:PAM

[IF02] Elise R. Irwin and Mary C. Freeman. Proposal for adaptive management to conserve biotic integrity in a regulated segment of the Tallapoosa River, Alabama, U.S.A. Una propuesta de manejo adaptivo para conservar la integridad biótica en un segmento regulado del Río Tallapoosa, Alabama (E.U.A.). Conservation Biology, 16(5):1212–1222, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Inogwabini:2005:PAD

[IIG05] Bila-Isia Inogwabini, Omari Ilambu, and Mbayma Atalia Gbanzi. Protected areas of the Democratic Republic of Congo. Conservation Biology, 19(1):15–22, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Imbeau:2001:LTE

[IMD01] Louis Imbeau, Mikko Mönkkönen, and André Desrochers. Longterm effects of forestry on birds of the Eastern Canadian boreal forests: a comparison with Fennoscandia. Efectos a largo plazo de actividades forestales sobre aves de los bosques boreales canadienses del este: una comparación con Fennoscandia. Conservation Biology, 15(4):1151–1162, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ito:2005:PEB

[IML^{+}05] Takehiko Y. Ito, Naoko Miura, Badamjav Lhagvasuren, Dulamtseren Enkhbileg, Seiki Takatsuki, Atsushi Tsunekawa, and Zhaowen Jiang. Preliminary evidence of a barrier effect of a railroad on the migration of Mongolian gazelles. Evidencia preliminar del efecto de barrera de un ferrocarril sobre la migración de gacelas de Mongolia. Conservation Biology, 19(3):945–948, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Infield:2001:CVF

[Inf01] Mark Infield. Cultural values: a forgotten strategy for building community support for protected areas in Africa. Conservation Biology, 15(3):800–802, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jackson:2002:WST

[Jac02] Wes Jackson. When strength through exhaustion fails. Conservation Biology, 16(2):291, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jackson:2008:NPA

[Jac08] Wes Jackson. The necessity and possibility of an agriculture where nature is the measure. Conservation Biology, 22(6):1376– 1377, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacquet:2009:WCC

[Jac09] Jennifer Jacquet. What can conservationists learn from investor behavior? Conservation Biology, 23(3):518–519, June 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2005:MAA

[JAH^{+}05] Julia P. G. Jones, Fortunat B. Andriahajaina, Neal J. Hockley, Andrew Balmford, and Olga R. Ravoahangimalala. A multidisciplinary approach to assessing the sustainability of freshwater crayfish harvesting in Madagascar. Un método multidisciplinario para evaluar la sustentabilidad de la captura de cangrejos de Río en Madagascar. Conservation Biology, 19(6):1863–1871, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2008:ITS

[JAH08] Julia P. G. Jones, Mijasoa M. Andriamarovololona, and Neal Hockley. The importance of taboos and social norms to conservation in Madagascar. La importancia de los tabús y las normas sociales para la conservación en Madagascar. Conservation Biology, 22(4):976–986, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

James:2002:BGB

[Jam02] Stacy M. James. Bridging the gap between private landowners and conservationists. Conservation Biology, 16(1):269–271, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Janzen:2007:GGC

[Jan07] H. Henry Janzen. Greenhouse gases as clues to permanence of farmlands. Los gases de invernadero como claves para la permanencia de tierras de cultivo. Conservation Biology, 21(3):668– 674, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jarrell:2003:RFD

[Jar03] Randall Jarrell. Raymond F. Dasmann: 1919–2002. Conservation Biology, 17(2):636–637, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jha:2006:PGH

[JB06] S. Jha and K. S. Bawa. Population growth, human development, and deforestation in biodiversity hotspots. Crecimiento poblacional, desarrollo humano y deforestación en sitios de importancia para la biodiversidad. Conservation Biology, 21(3): 906–912, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:2007:ILH

[JC07] Brent E. Johnson and J. Hall Cushman. Influence of a large herbivore reintroduction on plant invasions and community composition in a California grassland. Influencia de la reintroducción de un herbívoro mayor sobre invasiones de plantas y la composición de la comunidad en un pastizal de California. Conservation Biology, 21(2):515–526, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jessup:2003:GMC

[Jes03] David A. Jessup. Good medicine for conservation biology: Comments, corrections, and connections. Conservation Biology, 17 (3):921–924, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [LG02].

Jaeger:2004:ERF

[JF04] Jochen A. G. Jaeger and Lenore Fahrig. Effects of road fencing on population persistence. Efectos del cercado de caminos sobre la persistencia de la población. Conservation Biology, 18(6): 1651–1657, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Justus:2008:IRT

[JFS08] James Justus, Trevon Fuller, and Sahotra Sarkar. Influence of representation targets on the total area of conservation-area networks. Influencia de los objetivos de representación sobre el Área total de redes de Áreas de conservación. Conservation Biology, 22(3):673–682, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joseph:2006:PAV

[JFWP06] Liana N. Joseph, Scott A. Field, Chris Wilcox, and Hugh P. Possingham. Presence–absence versus abundance data for monitoring threatened species datos de presencia-ausencia versus abundancia para el monitoreo de especies amenazadas. Conservation Biology, 20(6):1679–1687, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson-Groh:2004:CTH

[JG04] Cindy Johnson-Groh. Conservation text in historical context. Conservation Biology, 18(3):852–853, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jenkins:2008:PMD

[JG08] Clinton N. Jenkins and Chandra Giri. Protection of mammal diversity in Central America. Protección de la diversidad de mamíferos en América Central. Conservation Biology, 22(4): 1037–1044, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2002:RWC

[JGL^{+}02] Kenneth L. Jones, Travis C. Glenn, Robert C. Lacy, James R. Pierce, Natalie Unruh, Claire M. Mirande, and Felipe Chavez-Ramirez. Refining the whooping crane studbook by incorporating microsatellite DNA and leg-banding analyses. Refinando el registro Genealógico de la grulla americana mediante la Incorporación de ADN Microsatélite y el análisis de agrupamientos. Conservation Biology, 16(3):789–799, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jenkins:2003:CPW

[JGM03a] David G. Jenkins, Scott Grissom, and Keith Miller. Consequences of prairie wetland drainage for crustacean biodiversity and metapopulations. Consecuencias del drenado de praderas de humedales para la biodiversidad y metapoblaciones de Crustáceos. Conservation Biology, 17(1):158–167, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jenkins:2003:CMG

[JGM03b] Martin Jenkins, Rhys E. Green, and Joah Madden. The challenge of measuring global change in wild nature: Are things getting better or worse? Conservation Biology, 17(1):20–23, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jaquiery:2009:PDE

[JGP09] J. Jaquiéry, F. Guillaume, and N. Perrin. Predicting the deleterious effects of mutation load in fragmented populations. Predicción de los efectos Deletéreos de la carga de mutaciones en poblaciones fragmentadas. Conservation Biology, 23(1):207– 218, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2005:CCC

[JH05] Clara B. Jones and Robert H. Horwich. Constructive criticism of community-based conservation. Conservation Biology, 19(4): 990–991, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jinxiu:2004:PAR

[JHHL04] Wang Jinxiu, Liu Hongmao, Hu Huabin, and Gao Lei. Participatory approach for rapid assessment of plant diversity through a folk classification system in a tropical rainforest: Case study in Xishuangbanna, China. Método participativo para la evaluación rápida de diversidad de plantas con un sistema de clasificación popular en un bosque tropical lluvioso: Caso de estudio en Xishuangbanna, China. Conservation Biology, 18(4):1139– 1142, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jack:2009:RPA

[JLF09] B. Kelsey Jack, Beria Leimona, and Paul J. Ferraro. A revealed preference approach to estimating supply curves for ecosystem services: Use of auctions to set payments for soil erosion control in Indonesia. Un método de preferencia revelada para la estimación de curvas de suministro de servicios del ecosistema: Uso de subastas para fijar pagos por la erosión de suelos en Indonesia. Conservation Biology, 23(2):359–367, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jaramillo-Legorreta:2007:SVI

[JLRBB^{+}07] Armando Jaramillo-Legorreta, Lorenzo Rojas-Bracho, Robert L. Brownell, Jr., Andrew J. Read, Randall R. Reeves, Katherine Ralls, and Barbara L. Taylor. Saving the vaquita: Immediate action, not more data. Conservation Biology, 22(1):1653–1655, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:2003:LDE

[JLZW03] Pieter T. J. Johnson, Kevin B. Lunde, Derek A. Zelmer, and J. Kirwin Werner. Limb deformities as an emerging parasitic disease in amphibians: Evidence from museum specimens and resurvey data. Extremidades deformes como una enfermedad parasítica emergente en anfibios: Evidencia de especimenes de museo y datos de campo nuevos. Conservation Biology, 17(6): 1724–1737, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:2001:WRB

[JMB01] E. A. Johnson, K. Miyanishi, and S. R. J. Bridge. Wildfire regime in the boreal forest and the idea of suppression and fuel buildup. Conservation Biology, 15(6):1554–1557, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joyal:2001:LEA

[JMH01] Lisa A. Joyal, Mark McCollough, and Malcolm L. Hunter, Jr. Landscape ecology approaches to wetland species conservation: a case study of two turtle species in Southern Maine. Aproximaciones de la ecología de paisaje a la conservación de especies de humedales: Estudio de caso de dos especies de tortugas en el sur de Maine. Conservation Biology, 15(6):1755–1762, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnston:2000:EMO

[JML00] D. W. Johnston, P. Meisenheimer, and D. M. Lavigne. An evaluation of management objectives for Canada’s commercial harp seal hunt, 1996-1998. Evaluación de los objetivos de manejo de la caza de la foca Pagophilus groenlandicus de Canadá, 1996– 1998. Conservation Biology, 14(3):729–737, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joly:2001:HME

[JMLG01] Pierre Joly, Claude Miaud, Anthony Lehmann, and Odile Grolet. Habitat matrix effects on pond occupancy in newts. Efecto de la matriz del hábitat en la ocupación de estanques por tritones. Conservation Biology, 15(1):239–248, February 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:2007:PCT

[JMM07] Susan K. Jacobson, Mallory D. Mcduff, and Martha C. Monroe. Promoting conservation through the arts: Outreach for hearts and minds. Conservation Biology, 21(1):7–10, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joseph:2009:OAR

[JMP09] Liana N. Joseph, Richard F. Maloney, and Hugh P. Possingham. Optimal allocation of resources among threatened species: a project prioritization protocol. Asignación Óptima de recursos entre especies amenazadas: un protocolo de priorización de proyectos. Conservation Biology, 23(2):328–338, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:2006:UBI

[JMS^{+}06a] Susan K. Jacobson, Julie K. Morris, J. Scott Sanders, Eugene N. Wiley, Michael Brooks, Robert E. Bennetts, H. Franklin Percival, and Susan Marynowski. Understanding barriers to implementation of an adaptive land management program. Entendiendo las barreras para la implementación de un programa adaptativo de gestión de tierras. Conservation Biology, 20(5):1516– 1527, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Juutinen:2006:CED

[JMS06b] Artti Juutinen, Mikko Mönkkönen, and Anna-Liisa Sippola. Cost-efficiency of decaying wood as a surrogate for overall species richness in boreal forests. La rentabilidad de la madera en descomposición como un sustituto de la riqueza total de especies en bosques boreales. Conservation Biology, 20(1):74–84, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2001:RCW

[Joh01] David Johns. A retrospective on conservation in West Africa. Conservation Biology, 15(2):543–544, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2003:SGM

[Joh03a] David Johns. Slaying the growth monster. Conservation Biology, 17(3):937–939, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2003:GCN

[Joh03b] David M. Johns. Growth, conservation, and the necessity of new alliances. Crecimiento, conservación y la necesidad de nuevas alianzas. Conservation Biology, 17(5):1229–1237, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2005:OCR

[Joh05a] David Johns. The other connectivity: Reaching beyond the choir. Conservation Biology, 19(6):1681–1682, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2005:OAB

[Joh05b] David M. Johns. Orr and Armageddon: Building a coalition. Conservation Biology, 19(6):1685–1686, December 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2007:LIP

[Joh07] David Johns. Like it or not, politics is the solution. Conservation Biology, 21(2):287–288, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnsingh:2008:FK

[Joh08] A. J. T. Johnsingh. A first of its kind. Conservation Biology, 22 (4):1084–1085, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:2009:MIQ

[Joh09a] David Johns. 1 10 most important questions. Conservation Biology, 23(5):1069, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:2009:MBM

[Joh09b] Pieter Johnson. The meaning behind malformed frogs. Conservation Biology, 23(2):508–511, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joly:2004:ST

[Jol04] Kyle Joly. Stop the train! Conservation Biology, 18(1):4, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2004:HEC

[Jon04] Menna E. Jones. Historical ecology for conservation managers. Conservation Biology, 18(1):281–282, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jordan:2001:SBA

[Jor01] Carl F. Jordan. The sustainability-biodiversity-agriculture conundrum. Conservation Biology, 15(1):294–296, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacquet:2008:FPB

[JP08] Jennifer Jacquet and Daniel Pauly. Funding priorities: Big barriers to small-scale fisheries. Conservation Biology, 22(4):832– 835, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jennings:2009:RBS

[JP09] Nancy Jennings and Michael J. O. Pocock. Relationships between sensitivity to agricultural intensification and ecological traits of insectivorous mammals and arthropods. Relaciones entre la sensibilidad a la intensificación agrícola y características ecológicas de mamíferos insectívoros y artrópodos. Conservation Biology, 23(5):1195–1203, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jamieson:2003:SSC

[JRL03] Ian G. Jamieson, Michael S. Roy, and Marieke Lettink. Sexspecific consequences of recent inbreeding in an ancestrally inbred population of New Zealand takahe. Consecuencias sexo– específicas de endogamia reciente en una población ancestralmente endogámica de takahe de Nueva Zelanda. Conservation Biology, 17(3):708–716, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:2006:AHQ

[JSHM06] Matthew D. Johnson, Thomas W. Sherry, Richard T. Holmes, and Peter P. Marra. Assessing habitat quality for a migratory songbird wintering in natural and agricultural habitats. Evaluación de la calidad de hábitat para aves canoras invernantes en hábitats naturales y agrícolas. Conservation Biology, 20(5): 1433–1444, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2005:ADF

[JSJ05] Gregory A. Jones, Kathryn E. Sieving, and Susan K. Jacobson. Avian diversity and functional insectivory on North-Central Florida farmlands. Diversidad aviar e insectivoría funcional en tierras agrícolas del Norte-Centro de Florida. Conservation Biology, 19(4):1234–1245, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:2003:AFA

[JSJV03] Susan K. Jacobson, Kathryn E. Sieving, Gregory A. Jones, and Annamaria Van Doorn. Assessment of Farmer attitudes and behavioral intentions toward bird conservation on organic and conventional Florida farms. Evaluación de las actitudes e intenciones conductuales de agricultores hacia la conservación de aves en ranchos orgánicos y convencionales de Florida. Conservation Biology, 17(2):595–606, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jetz:2008:ECC

[JSW08] Walter Jetz, Cagan H. Sekercioglu, and James E. M. Watson. Ecological correlates and conservation implications of overestimating species geographic ranges. Correlaciones ecológicas e implicaciones para la conservación de la sobrestimación de los rangos geográficos de especies. Conservation Biology, 22(1):110– 119, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2000:ERH

[JSWS00] Julia A. Jones, Frederick J. Swanson, Beverley C. Wemple, and Kai U. Snyder. Effects of roads on hydrology, geomorphology, and disturbance patches in stream networks. Efectos de carreteras en la hidrología, geomorfología y parches de perturbación en redes de arroyos. Conservation Biology, 14(1):76–85, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jakalaniemi:2006:CSD

[JTS06] Anne Jäkäläniemi, Juha Tuomi, and Pirkko Siikamäki. Conservation of species in dynamic landscapes: Divergent fates of Silene tatarica populations in riparian habitats. Conservación de especies en paisajes dinámicos: Destinos divergentes de poblaciones de Silene tatarica en hábitats ribereños. Conservation Biology, 21(3):844–852, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2008:SEI

[JTZ^{+}08] Holly P. Jones, Bernie R. Tershy, Erika S. Zavaleta, Donald A. Croll, Bradford S. Keitt, Myra E. Finkelstein, and Gregg R. Howald. Severity of the effects of invasive rats on seabirds: A global review. Revisión de la severidad global de los efectos de ratas invasoras sobre aves marinas. Conservation Biology, 22 (1):16–26, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jukofsky:2001:SCE

[Juk01] Diane Jukofsky. Shared conservation experiences. Conservation Biology, 15(4):818–819, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jepson:2002:ECC

[JW02] Paul Jepson and Robert J. Whittaker. Ecoregions in context: a critique with special reference to Indonesia ecoregiones en contexto: una critica con especial referencia a Indonesia. Conservation Biology, 16(1):42–57, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jamieson:2006:IES

[JWB06] Ian G. Jamieson, Graham P. Wallis, and James V. Briskie. Inbreeding and endangered species management: Is New Zealand out of step with the rest of the world? Endogamia y gestión de especies en peligro: `?Está Nueva Zelanda desfasada del resto del mundo? Conservation Biology, 20(1):38–47, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kurosawa:2003:EHF

[KA03] Reiko Kurosawa and Robert A. Askins. Effects of habitat fragmentation on birds in deciduous forests in Japan. Efectos de la fragmentación del hábitat sobre aves en bosques deciduos del Japón. Conservation Biology, 17(3):695–707, June 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaiser:2003:DEF

[Kai03] Michel J. Kaiser. Detecting the effects of fishing on seabed community diversity: Importance of scale and sample size. Detección de los efectos de la pesca sobre las comunidades del fondo marino: Importancia de la escala y el tamaño de muestra. Conservation Biology, 17(2):512–520, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kala:2005:IUP

[Kal05] Chandra Prakash Kala. Indigenous uses, population density, and conservation of threatened medicinal plants in protected areas of the Indian Himalayas. Usos autóctonos, densidad poblacional y conservación de plantas medicinales amenazadas en Áreas protegidas de los Himalayas Hindús. Conservation Biology, 19(2): 368–378, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kareiva:2001:MBD

[Kar01] Peter M. Kareiva. Making a big difference. Conservation Biology, 15(5):1472–1473, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karr:2009:EBM

[Kar09] James R. Karr. Ecosystem-based management: What really happens? Conservation Biology, 23(6):1648–1651, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kauffman:2004:DRF

[Kau04] J. Boone Kauffman. Death rides the forest: Perceptions of fire, land use, and ecological restoration of western forests. La muerte cabalga en el bosque — percepciones del fuego, uso del suelo y restauración ecológica de bosques occidentales. Conservation Biology, 18(4):878–882, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kawall:2004:SSI

[Kaw04] Jason Kawall. Searching for sustainability: Interdisciplinary essays in the philosophy of conservation biology. Conservation Biology, 18(2):589–590, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kolm:2003:WPR

[KB03] Niclas Kolm and Anders Berglund. Wild populations of a reef fish suffer from the “Nondestructive” aquarium trade fishery. Poblaciones salvajes de un pez de arrecife sufren como resultado de la pesca comercial de acuario. Conservation Biology, 17(3): 910–914, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keller:2004:RLS

[KB04a] Verena Keller and Kurt Bollmann. From red lists to species of conservation concern. De listas rojas a especies interés para la conservación. Conservation Biology, 18(6):1636–1644, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kodric-Brown:2004:CES

[KB04b] Astrid Kodric-Brown. Complexities of the exotic species problem. Conservation Biology, 18(4):1171, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kleijn:2005:ESA

[KB05] David Kleijn and András Báldi. Effects of set-aside land on farmland biodiversity: Comments on Van Buskirk and Willi. Conservation Biology, 19(3):963–966, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [VW04].

Keough:2006:AIC

[KB06] Heather L. Keough and Dale J. Blahna. Achieving integrative, collaborative ecosystem management. Logrando la gestión integradora y cooperativa de ecosistemas. Conservation Biology, 20 (5):1373–1382, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koontz:2008:IEM

[KB08] Tomas M. Koontz and Jennifer Bodine. Implementing ecosystem management in public agencies: Lessons from the U.S. Bureau of Land Management and the Forest Service. Implementación de la gestión de ecosistemas en agencias públicas: Lecciones del Buró de Gestión de Tierras y del Servicio Forestal de E.U.A. Conservation Biology, 22(1):60–69, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kiesecker:2001:TPF

[KBM01] Joseph M. Kiesecker, Andrew R. Blaustein, and Cheri L. Miller. Transfer of a pathogen from fish to amphibians. Transferencia de un patógeno de peces a anfibios. Conservation Biology, 15(4): 1064–1070, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kohm:2000:PSP

[KBMN00] Kathryn Kohm, P. Dee Boersma, Gary K. Meffe, and Reed Noss. Putting the science into practice and the practice into science. Conservation Biology, 14(3):593–594, June 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Krueper:2003:RVB

[KBR03] David Krueper, Jonathan Bart, and Terrell D. Rich. Response of vegetation and breeding birds to the removal of cattle on the San Pedro River, Arizona (U.S.A.). Respuesta de la vegetación y las aves en reproducción a la remoción de ganado en el Río San Pedro, Arizona (U.S.A.). Conservation Biology, 17(2):607– 615, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kleijn:2004:EEA

[KBS^{+}04] David Kleijn, Frank Berendse, Ruben Smit, Niels Gilissen, John Smit, Bastiaan Brak, and Rolf Groeneveld. Ecological effectiveness of agri-environment schemes in different agricultural landscapes in the Netherlands. Efectividad ecológica de esquemas agri-ambientes en diferentes paisajes agrícolas en Holanda. Conservation Biology, 18(3):775–786, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kautz:2001:SHB

[KC01] Randy S. Kautz and James A. Cox. Strategic habitats for biodiversity conservation in Florida. Hábitats estratégicos para la conservación de la biodiversidad en Florida. Conservation Biology, 15(1):55–77, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2007:EOS

[KC07] Andrew T. Knight and Richard M. Cowling. Embracing opportunism in the selection of priority conservation areas. Conservation Biology, 21(4):1124–1126, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2008:CMC

[KC08] Andrew T. Knight and Richard M. Cowling. Clearing the mud from the conservation opportunity debate: Reply to Pressey and Bottrill. Conservation Biology, 22(5):1346–1348, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2006:OMI

[KCC06] Andrew T. Knight, Richard M. Cowling, and Bruce M. Campbell. An operational model for implementing conservation action. Un modelo operacional para la implementación de acciones de conservación. Conservation Biology, 20(2):408–419, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Klein:2008:SBB

[KCK^{+}08] C. J. Klein, A. Chan, L. Kircher, A. J. Cundiff, N. Gardner, Y. Hrovat, A. Scholz, B. E. Kendall, and S. Airamé. Striking a balance between biodiversity conservation and socioeconomic viability in the design of marine protected areas. Estableciendo un balance entre la conservación de la biodiversidad y la viabilidad económica en el diseño de Áreas marinas protegidas. Conservation Biology, 22(3):691–700, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2008:KDS

[KCR^{+}08] Andrew T. Knight, Richard M. Cowling, Mathieu Rouget, Andrew Balmford, Amanda T. Lombard, and Bruce M. Campbell. Knowing but not doing: Selecting priority conservation areas and the research–implementation gap. Sabiendo pero no haciendo: Selección de Áreas prioritarias para la conservación y la brecha Investigación–Implementación. Conservation Biology, 22(3):610–617, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kodandapani:2004:CTI

[KCS04] Narendran Kodandapani, Mark A. Cochrane, and R. Sukumar. Conservation threat of increasing fire frequencies in the Western Ghats, India. Amenaza del incremento de frecuencias de incendios a la conservación en las Ghats Occidentales, India. Conservation Biology, 18(6):1553–1561, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:2000:VSC

[KD00] Marcella J. Kelly and Sarah M. Durant. Viability of the Serengeti cheetah population. Viabilidad de la población de leopardos en el Serengeti. Conservation Biology, 14(3):786–797, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kolb:2005:ELH

[KD05] Annette Kolb and Martin Diekmann. Effects of life-history traits on responses of plant species to forest fragmentation. Efectos de las características de la historia de vida sobre las respuestas de especies de plantas a la fragmentación de bosques. Conservation Biology, 19(3):929–938, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2006:DSC

[KDC^{+}06] Andrew T. Knight, Amanda Driver, Richard M. Cowling, Kristal Maze, Philip G. Desmet, Amanda T. Lombard, Mathieu Rouget, Mark A. Botha, Andre F. Boshoff, J. Guy Castley, Peter S. Goodman, Kathy Mackinnon, Shirley M. Pierce, Rebecca Sims-Castley, Warrick I. Stewart, and Amrei Von Hase. Designing systematic conservation assessments that promote effective implementation: Best practice from South Africa. Diseño de evaluaciones sistemáticas de la conservación que promueven la implementación efectiva: la mejor Práctica en África del Sur. Conservation Biology, 20(3):739–750, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Krauss:2002:RGD

[KDD02] Siegfried L. Krauss, Bob Dixon, and Kingsley W. Dixon. Rapid genetic decline in a translocated population of the endangered plant Grevillea scapigera. Declinación genética rápida en una población translocada de una planta en peligro Grevillea scapigera. Conservation Biology, 16(4):986–994, August 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kati:2004:TVS

[KDD^{+}04] Vassiliki Kati, Pierre Devillers, Marc Dufrêne, Anastasios Legakis, Despina Vokou, and Philippe Lebrun. Testing the value of six taxonomic groups as biodiversity indicators at a local scale. Probando el valor de seis grupos taxonómicos como indicadores de biodiversidad en una escala local. Conservation Biology, 18 (3):667–675, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kilpatrick:2006:PPI

[KDG^{+}06] A. Marm Kilpatrick, Peter Daszak, Simon J. Goodman, Helmuth Rogg, Laura D. Kramer, Virna Cedeño, and Andrew A. Cunningham. Predicting pathogen introduction: West Nile virus spread to Galápagos. Predicción de la introducción de patógenos: Dispersión del virus del Nilo a las Galápagos. Conservation Biology, 33(5):1224–1231, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kremen:2008:IAE

[KDKS08] Claire Kremen, Gretchen C. Daily, Alexandra-Maria Klein, and Drew Scofield. Inadequate assessment of the ecosystem service rationale for conservation: Reply to Ghazoul. Conservation Biology, 22(3):795–798, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keller:2007:FBR

[KDL07] Reuben P. Keller, John M. Drake, and David M. Lodge. Fecundity as a basis for risk assessment of nonindigenous freshwater molluscs. La fecundidad como base para la evaluación de riesgo de moluscos dulceacuícolas no nativos. Conservation Biology, 21 (1):191–200, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keeley:2006:FMI

[Kee06] Jon E. Keeley. Fire management impacts on invasive plants in the Western United States. Impactos de la gestión de fuego sobre plantas invasoras en el Oeste de Estados Unidos. Conservation Biology, 20(2):375–384, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaiser:2006:REF

[KEJ06] Michel J. Kaiser and Gareth Edwards-Jones. The role of ecolabeling in fisheries management and conservation. El papel del ecorotulado en la gestión y conservación de pesquerías. Conservation Biology, 20(2):392–398, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:2001:LLS

[Kel01] Marcella J. Kelly. Lineage loss in Serengeti cheetahs: Consequences of high reproductive variance and heritability of fitness on effective population size. Pérdida de linaje en cheetas del serengueti: Consecuencias de una varianza reproductiva alta y de la herencia de la adaptabilidad en el tamaño poblacional. Conservation Biology, 15(1):137–147, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelt:2008:FV

[Kel08] Douglas A. Kelt. Filling the void. Conservation Biology, 22(4): 1081–1082, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kery:2004:ERE

[Kér04] Marc Kéry. Extinction rate estimates for plant populations in revisitation studies: Importance of detectability. Estimaciones de tasas de extinción de poblaciones de plantas en estudios de revisita: Importancia de la detectabilidad. Conservation Biology, 18(2):570–574, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kerstetter:2009:PFC

[Ker09] David W. Kerstetter. Practicing fisheries conservation biology within harvesting regimes. Conservation Biology, 23(5):1072– 1073, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kessler:2001:MPC

[Kes01] Michael Kessler. Maximum plant-community endemism at intermediate intensities of anthropogenic disturbance in Bolivian montane forests. Endemismo máximo en comunidades de plantas con perturbaciones antropogénicas de intensidades intermedias en bosques montañosos bolivianos. Conservation Biology, 15(3):634–641, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keeley:2001:HFR

[KF01a] Jon E. Keeley and C. J. Fotheringham. Historic fire regime in Southern California shrublands. Régimen histórico de incendios en zonas arbustivas del sur de California. Conservation Biology, 15(6):1536–1548, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keeley:2001:HMC

[KF01b] Jon E. Keeley and C. J. Fotheringham. History and management of crown-fire ecosystems: a summary and response. Conservation Biology, 15(6):1561–1567, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kendall:2002:VAI

[KF02] Bruce E. Kendall and Gordon A. Fox. Variation among individuals and reduced demographic stochasticity. Variación entre individuos y estocasticidad demográfica reducida. Conservation Biology, 16(1):109–116, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kendall:2003:UIV

[KF03] Bruce E. Kendall and Gordon A. Fox. Unstructured individual variation and demographic stochasticity. Conservation Biology, 17(4):1170–1172, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kirchner:2003:RCP

[KFA^{+}03] Florian Kirchner, Jean-Baptiste Ferdy, Christophe Andalo, Bruno Colas, and Jacques Moret. Role of corridors in plant dispersal: an example with the endangered Ranunculus nodiflorus. El papel de los corredores en la dispersión de plantas: el ejemplo de la especie en peligro Ranunculus nodiflorus. Conservation Biology, 17(2):401–410, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kirika:2008:ELD

[KFBG08] Jasper Mbae Kirika, Nina Farwig, and Katrin Böhning-Gaese. Effects of local disturbance of tropical forests on frugivores and seed removal of a small-seeded afrotropical tree. Efectos de la perturbación local en bosques tropicales sobre Frugívoros y la remoción de semillas de un Árbol afrotropical con semillas Pequeñas. Conservation Biology, 22(2):318–328, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kozubikova:2009:PCP

[KFK^{+}09] Eva Kozubíková, L. Filipová, P. Kozák, Z. Ďuriš, M. P. Martín, J. Diéguez-Uribeondo, B. Oidtmann, and A. Petrusek. Prevalence of the crayfish plague pathogen Aphanomyces astaci in invasive American crayfishes in the Czech Republic. Prevalencia del patógeno causante de la “peste del cangrejo,” Aphanomyces astaci, en especies invasoras de cangrejo de río americano de la República Checa. Conservation Biology, 23(5):1204–1213, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keller:2008:PSI

[KFL08] Reuben P. Keller, Kristin Frang, and David M. Lodge. Preventing the spread of invasive species: Economic benefits of intervention guided by ecological predictions. Prevención de la propagación de especies invasoras: Beneficios económicos de la Intervención guiada por predicciones ecológicas. Conservation Biology, 22(1):80–88, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koster:2009:CED

[KFNB09] Nils Köster, Karoline Friedrich, Jürgen Nieder, and Wilhelm Barthlott. Conservation of epiphyte diversity in an Andean landscape transformed by human land use. Conservación de la diversidad de epífitas en una paisaje andino transformado por uso de suelo humano. Conservation Biology, 23(4):911–919, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kattan:2006:SCB

[KFSR^{+}06] Gustavo H. Kattan, Padu Franco, Carlos A. Saavedra-Rodríguez, Carlos Valderrama, Vladimir Rojas, Daniel Osorio, and Jesús Martínez. Spatial components of bird diversity in the Andes of Colombia: Implications for designing a regional reserve system componentes espaciales de la diversidad de aves en los Andes de Colombia: Implicaciones para el diseño de un sistema regional de reservas. Conservation Biology, 33(5):1203– 1211, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kursar:2002:LRM

[KG02] Thomas A. Kursar and William Grauel. Logging from rivers may control human invasions. Conservation Biology, 16(2):285, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kindvall:2003:TEP

[KG03] Oskar Kindvall and Ulf Gärdenfors. Temporal extrapolation of PVA results in relation to the IUCN red list criterion E. Extrapolación temporal de resultados de PVA en relación con el criterio E de los criterios de la lista roja de la UICN. Conservation Biology, 17(1):316–321, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kretzmann:2001:MAP

[KGM01] Maria B. Kretzmann, Neil J. Gemmell, and Axel Meyer. Microsatellite analysis of population structure in the endangered Hawaiian monk seal. Análisis Microsatélite de la estructura poblacional de la foca monje de Hawai. Conservation Biology, 15(2):457–466, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kerley:2002:ERH

[KGM^{+}02] Linda L. Kerley, John M. Goodrich, Dale G. Miquelle, Evgeny N. Smirnov, Howard B. Quigley, and Maurice G. Hornocker. Effects of roads and human disturbance on Amur tigers. Efectos de caminos y perturbación humana sobre tigres Amur. Conservation Biology, 16(1):97–108, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kolbe:2007:MSA

[KGS^{+}07] Jason J. Kolbe, Richard E. Glor, Lourdes Rodríguez Schettino, Ada Chamizo Lara, Allan Larson, and Jonathan B. Losos. Multiple sources, admixture, and genetic variation in introduced Anolis lizard populations. Fuentes múltiples, mezcla y variación genética en poblaciones de lagartijas Anolis introducidas. Conservation Biology, 22(1):1612–1625, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knapp:2003:IEC

[KHHK03] Shannon M. Knapp, Carola A. Haas, Douglas N. Harpole, and Roy L. Kirkpatrick. Initial effects of clearcutting and alternative silvicultural practices on terrestrial salamander abundance. Efectos iniciales de la tala rasa y de prácticas silvícolas alternativas sobre la abundancia de salamandras terrestres. Conservation Biology, 17(3):752–762, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kalinowski:2000:IDS

[KHM00] Steven T. Kalinowski, Philip W. Hedrick, and Philip S. Miller. Inbreeding depression in the Speke’s gazelle captive breeding program. Depresión por intracruza en el programa de reproducción en cautiverio para la gacela de Speke. Conservation Biology, 14(5):1375–1384, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kwit:2004:CSG

[KHP04] Charles Kwit, Carol C. Horvitz, and William J. Platt. Conserving slow-growing, long-lived tree species: Input from the demography of a rare understory conifer, Taxus floridana. Conservación de especies arbóreas de crecimiento lento y longevas: Insumos de la demografía de una Conífera de sotobosque rara, Taxus floridana. Conservation Biology, 18(2):432–443, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kazmaier:2001:EGD

[KHRS01] Richard T. Kazmaier, Eric C. Hellgren, Donald C. Ruthven III, and David R. Synatzske. Effects of grazing on the demography and growth of the Texas tortoise. Efectos del pastoreo sobre la demografía y el crecimiento de la tortuga de Texas. Conservation Biology, 15(4):1091–1101, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kramer:2008:PFF

[KIAH08] Andrea T. Kramer, Jennifer L. Ison, Mary V. Ashley, and Henry F. Howe. The paradox of forest fragmentation genetics. La paradoja de la genética de la fragmentación de bosques. Conservation Biology, 22(4):878–885, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kirch:2008:AR

[Kir08] Patrick V. Kirch. Avis raris. Conservation Biology, 22(2):500– 501, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kitchell:2007:FYE

[Kit07] James F. Kitchell. Finding your ecosystem episteme. Conservation Biology, 21(2):891–892, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kellermann:2008:EES

[KJSH08] Jherime L. Kellermann, Matthew D. Johnson, Amy M. Stercho, and Steven C. Hackett. Ecological and economic services provided by birds on Jamaican Blue Mountain coffee farms. Servicios ecológicos y económicos proporcionados por aves en fincas cafetaleras en Blue Mountains, Jamaica. Conservation Biology, 22(5):1177–1185, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaplan:2008:BBP

[KK08] Rachel Kaplan and Stephen Kaplan. Bringing out the best in people: a psychological perspective. Conservation Biology, 22 (4):826–829, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kothamasi:2009:ECB

[KK09] David Kothamasi and E. Toby Kiers. Emerging conflicts between biodiversity conservation laws and scientific research: the case of the Czech entomologists in India. Conservation Biology, 23(5):1328–1330, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kallimanis:2005:MER

[KKHS05] A. S. Kallimanis, W. E. Kunin, J. M. Halley, and S. P. Sgardelis. Metapopulation extinction risk under spatially autocorrelated disturbance. Riesgo de extinción metapoblacional bajo perturbación autocorrelacionada espacialmente. Conservation Biology, 19(2):534–546, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keller:2005:PNS

[KL05] Reuben P. Keller and David M. Lodge. Probability of nonindigenous seafood species becoming established. Conservation Biology, 19(2):287–288, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kuntz:2006:IMC

[KL06] Kathryn Lynne Kuntz and Douglas W. Larson. Influences of microhabitat constraints and rock-climbing disturbance on cliffface vegetation communities. Influencias de las reducciones de microhábitat y de la perturbación del escalado de roca sobre las comunidades vegetales en acantilados. Conservation Biology, 21 (3):821–832, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kleiman:2003:SB

[Kle03] Devra G. Kleiman. Striking a balance. Conservation Biology, 17(2):628–629, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kleiman:2006:CFB

[Kle06] Devra G. Kleiman. Conservation focus: Baiji, freshwater dolphin. Conservation Biology, 20(3):620, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kristensen:2008:LIE

[KLH08] Torsten N. Kristensen, Volker Loeschcke, and Ary A. Hoffmann. Linking inbreeding effects in captive populations with fitness in the wild: Release of replicated Drosophila melanogaster lines under different temperatures. Ligando los efectos de la endogamia en poblaciones cautivas con la adecuación biológica en el medio natural: Liberación de líneas de Drosophila melanogaster replicadas bajo temperaturas diferentes. Conservation Biology, 22 (1):189–199, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knapp:2000:NNF

[KM00] Roland A. Knapp and Kathleen R. Matthews. Non-native fish introductions and the decline of the mountain yellow-legged frog from within protected areas. Introducciones de peces no-nativos y disminuciones de la rana de montaña de patas amarillas dentro de Áreas protegidas. Conservation Biology, 14(2):428–438, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Klink:2005:CBC

[KM05] Carlos A. Klink and Ricardo B. Machado. Conservation of the Brazilian cerrado. Conservación del cerrado brasileño. Conservation Biology, 19(3):707–713, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Katzner:2007:UMI

[KMGB07] Todd Katzner, E. J. Milner-Gulland, and Evgeny Bragin. Using modeling to improve monitoring of structured populations: Are we collecting the right data? Utilización de modelos para mejorar el monitoreo de poblaciones estructuradas: `?estamos recolectando los datos correctos? Conservation Biology, 21(1): 241–252, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kideghesho:2007:WRT

[KNH07] Jafari R. Kideghesho, Julius W. Nyahongo, and Tomas Holmern. Wildlife research in Tanzania: Capacity building is feasible. Conservation Biology, 21(6):1385–1386, December 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2006:FLW

[Kni06] Andrew T. Knight. Failing but learning: Writing the wrongs after Redford and Taber. Conservation Biology, 33(5):1312– 1314, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2007:SES

[Kni07] Richard L. Knight. To save the Earth or to save a watershed: Optimism versus hope. Conservation Biology, 21(6):1397–1398, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:2009:CBR

[Kni09] Andrew T. Knight. Is conservation biology ready to fail? Conservation Biology, 23(3):517, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kinnaird:2001:WSW

[KO01] Margaret F. Kinnaird and Timothy G. O’Brien. Who’s scratching whom? Reply to Whitten et al. Conservation Biology, 15(5): 1459–1460, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koenig:2003:EST

[Koe03] Walter D. Koenig. European starlings and their effect on native cavity-nesting birds. Estorninos europeos y su efecto sobre aves nativas que anidan en cavidades. Conservation Biology, 17(4): 1134–1140, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koh:2007:PHB

[Koh07] Lian Pin Koh. Potential habitat and biodiversity losses from intensified biodiesel feedstock production. Conservation Biology, 21(5):1373–1375, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koh:2009:CIU

[Koh09] Lian Pin Koh. Calling Indonesia’s US$ 13 billion bluff. Conservation Biology, 23(4):789, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Komonen:2003:HID

[Kom03] Atte Komonen. Hotspots of insect diversity in boreal forests. Áreas de importancia para la diversidad de insectos en bosques boreales. Conservation Biology, 17(4):976–981, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koptur:2001:WH

[Kop01] Suzanne Koptur. Warming up to hotspots. Conservation Biology, 15(2):542–543, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karesh:2002:JFI

[KORB02] William B. Karesh, Steven A. Osofsky, Tonie E. Rocke, and Paul L. Barrows. Joining forces to improve our world. Conservation Biology, 16(5):1432–1434, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kotliar:2000:ANK

[Kot00] Natasha B. Kotliar. Application of the new keystone-species concept to prairie dogs: How well does it work? Aplicación del nuevo concepto de especie pilar en perros de pradera: `?Qué tan bien funciona? Conservation Biology, 14(6):1715–1721, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kozloff:2007:S

[Koz07] Eugene N. Kozloff. Stressed out? Conservation Biology, 21(1): 284, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kriger:2007:LVP

[KPH07] Kerry M. Kriger, Felicia Pereoglou, and Jean-Marc Hero. Latitudinal variation in the prevalence and intensity of chytrid (Batrachochytrium dendrobatidis) infection in Eastern Australia. Variación latitudinal en la prevalencia e intensidad de infección por quitridios (Batrachochytrium dendrobatidis) en Australia oriental. Conservation Biology, 21(5):1280–1290, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Krivanek:2006:PHP

[KPJ06] Martin Křivánek, Petr Pyšek, and Vojtěch Jarošík. Planting history and propagule pressure as predictors of invasion by woody species in a temperate region. Historia de siembra y presión de propágulos como predictores de la invasión de especies leñosas en una región templada. Conservation Biology, 20(5):1487–1498, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kremen:2000:GPP

[KR00] Claire Kremen and Taylor Ricketts. Global perspectives on pollination disruptions. Conservation Biology, 14(5):1226–1228, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kraft:2001:LSC

[Kra01] Michael E. Kraft. Leverage and sustainable communities: Overcoming policy obstacles at the local level. Conservation Biology, 15(6):1483–1484, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kress:2007:CTL

[Kre07] W. John Kress. Conserving the tree of life. Conservation Biology, 21(2):561–563, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kritzer:2004:ENS

[Kri04] Jacob P. Kritzer. Effects of noncompliance on the success of alternative designs of marine protected-area networks for conservation and fisheries management. Efectos del incumplimiento sobre el Éxito de Diseños alternativos de redes de Áreas marinas protegidas para conservación y gestión de pesquerías. Conservation Biology, 18(4):1021–1031, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kricher:2005:TBT

[Kri05] John Kricher. Two books tell the same story. Conservation Biology, 19(4):1335–1337, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kleiman:2000:IEC

[KRM^{+}00] Devra G. Kleiman, Richard P. Reading, Brian J. Miller, Tim W. Clark, J. Michael Scott, John Robinson, Richard L. Wallace, Robert J. Cabin, and Fred Felleman. Improving the evaluation of conservation programs. Mejoramiento de la evaluación de programas de conservación. Conservation Biology, 14(2):356–365, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kleiman:2000:IVC

[KRW^{+}00] Devra G. Kleiman, Richard P. Reading, Richard L. Wallace, John Robinson, Robert J. Cabin, and F. Felleman. Improving the value of conservation programs. Conservation Biology, 14 (6):1569, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kimball:2003:DEC

[KS03] Sarah Kimball and Paula M. Schiffman. Differing effects of cattle grazing on native and alien plants diferentes. Efectos del pastoreo de ganado sobre plantas nativas y exóticas. Conservation Biology, 17(6):1681–1693, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kark:2005:ESA

[KS05] Salit Kark and Daniel Sol. Establishment success across convergent Mediterranean ecosystems: an analysis of bird introductions. Exitoso de establecimiento en ecosistemas Mediterráneos convergentes: un análisis de introducciones de aves. Conservation Biology, 19(5):1519–1527, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koontz:2001:GDT

[KSB01] Jason A. Koontz, Pamela S. Soltis, and Steven J. Brunsfeld. Genetic diversity and tests of the hybrid origin of the endangered yellow larkspur. Diversidad genética y pruebas del origen Híbrido del delfinio amarillo en peligro de extinción. Conservation Biology, 15(6):1608–1618, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koh:2004:ECE

[KSB04] Lian Pin Koh, Navjot S. Sodhi, and Barry W. Brook. Ecological correlates of extinction proneness in tropical butterflies. Correlativos ecológicos de la Propensión a la extinción de mariposas tropicales. Conservation Biology, 18(6):1571–1578, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kainer:2006:GEF

[KSC^{+}06] Karen A. Kainer, Marianne Schmink, Hannah Covert, John Richard Stepp, Emilio M. Bruna, Jonathan L. Dain, Santiago Espinosa, and Shoana Humphries. A graduate education framework for tropical conservation and development. Un marco de educación de posgrado para conservación y desarrollo tropical. Conservation Biology, 20(1):3–13, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Klein:2002:ELU

[KSDBT02] Alexandra-Maria Klein, Ingolf Steffan-Dewenter, Damayanti Buchori, and Teja Tscharntke. Effects of land-use intensity in tropical agroforestry systems on coffee flower-visiting and trapnesting bees and wasps. Efectos de la intensidad del uso del suelo en sistemas agroforestales tropicales sobre las abejas y avispas que visitan las flores del café. Conservation Biology, 16(4):1003– 1014, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaiser:2000:FGR

[KSH00] Michel J. Kaiser, Fiona E. Spence, and Paul J. B. Hart. Fishinggear restrictions and conservation of benthic habitat complexity. Restricciones en las artes de pesca y conservación de la complejidad del hábitat béntico. Conservation Biology, 14(5):1512–1525, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kinnaird:2003:DTT

[KSO^{+}03] Margaret F. Kinnaird, Eric W. Sanderson, Timothy G. O’Brien, Hariyo T. Wibisono, and Gillian Woolmer. Deforestation trends in a tropical landscape and implications for endangered large mammals. Tendencias de deforestación en un paisaje tropical y sus implicancias para mamíferos grandes en peligro. Conservation Biology, 17(1):245–257, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kerr:2000:ITR

[KSP00] Jeremy T. Kerr, Alissa Sugar, and Laurence Packer. Indicator taxa, rapid biodiversity assessment, and nestedness in an endangered ecosystem. Taxones indicadores, evaluación rápida de la biodiversidad y la importancia del anidamiento en un ecosistema en peligro. Conservation Biology, 14(6):1726–1734, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kruess:2002:GID

[KT02] Andreas Kruess and Teja Tscharntke. Grazing intensity and the diversity of grasshoppers, butterflies, and trap-nesting bees and wasps intensidad de pastoreo y la diversidad de chapulines, mariposas y abejas y avispas. Conservation Biology, 16 (6):1570–1580, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kessler:2006:CBP

[KT06] Winifred B. Kessler and Jack Ward Thomas. Conservation biology from the perspective of natural resource management disciplines. Conservation Biology, 20(3):670–673, June 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kerwath:2009:CIB

[KTN^{+}09] Sven E. Kerwath, Eva B. Thorstad, Tor F. Næsje, Paul D. Cowley, Finn Økland, Chris Wilke, and Colin G. Attwood. Crossing invisible boundaries: the effectiveness of the Langebaan Lagoon Marine Protected Area as a harvest refuge for a migratory fish species in South Africa. Transponiendo barreras invisibles: la efectividad del Área Marina Protegida Laguna Langebaan como un refugio de cosecha para especies de peces migratorios en África del Sur. Conservation Biology, 23(3):653–661, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kintsch:2002:FSC

[KU02] Julia A. Kintsch and Dean L. Urban. Focal species, community representation, and physical proxies as conservation strategies: a case study in the Amphibolite Mountains, North Carolina, U.S.A. Especies focales, representación comunitaria e indicadores Físicos como estrategias de conservación: un estudio de caso en las Montañas Amphibolite, Carolina del Norte, E.U.A.

Conservation Biology, 16(4):936–947, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kumar:2008:NCG

[Kum08] Satish Kumar. Nonviolent change, a good idea. Conservation Biology, 22(2):239–240, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kunz:2003:EEB

[Kun03] Thomas H. Kunz. Encyclopedic efforts behind a first-of-its-kind encyclopedia. Conservation Biology, 17(3):939–940, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kalinowski:2002:REC

[KW02] Steven T. Kalinowski and Robin S. Waples. Relationship of effective to census size in fluctuating populations. Relación entre tamaño efectivo y censal en poblaciones fluctuantes. Conservation Biology, 16(1):129–136, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kingsford:2009:MCP

[KWL^{+}09] R. T. Kingsford, J. E. M. Watson, C. J. Lundquist, O. Venter, L. Hughes, E. L. Johnston, J. Atherton, M. Gawel, D. A. Keith, B. G. Mackey, C. Morley, H. P. Possingham, B. Raynor, H. F. Recher, and K. A. Wilson. Major conservation policy issues for biodiversity in Oceania. Principales cuestiones de políticas de conservación de la biodiversidad en Oceanía. Conservation Biology, 23(4):834–840, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelm:2008:TBB

[KWV08a] Detlev H. Kelm, Kerstin R. Wiesner, and Otto Von Helversen. Are there benefits of bat roosts for tropical forest restoration? Conservation Biology, 22(5):1090–1091, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelm:2008:L

[KWV08b] Detlev H. Kelm, Kerstin R. Wiesner, and Otto Von Helversen. Letters. Conservation Biology, 22(6):1370, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelm:2008:EAR

[KWvH08] Detlev H. Kelm, Kerstin R. Wiesner, and Otto von Helversen. Effects of artificial roosts for frugivorous bats on seed dispersal in a neotropical forest pasture mosaic. Efectos de refugios artificiales para murciélagos frugívoros sobre la dispersión de semillas en un mosaico bosque-pastura neotropical. Conservation Biology, 22(3):733–741, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keller:2009:VTF

[KZA09] Reuben P. Keller, Philine S. E. Zu Ermgassen, and David C. Aldridge. Vectors and timing of freshwater invasions in Great Britain. Vectores y temporización de invasiones dulceacuícolas en Gran Bretaña. Conservation Biology, 23(6):1526–1534, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keller:2005:CSA

[KZSV05] Verena Keller, Niklaus Zbinden, Hans Schmid, and Bernard Volet. A case study in applying the IUCN regional guidelines for national red lists and justifications for their modification. Un caso de estudio de la aplicación de “Directrices regionales” de IUCN para listas rojas nacionales y justificaciones para su Modificación. Conservation Biology, 19(6):1827–1834, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lesica:2000:SWU

[LA00] Peter Lesica and Helen E. Atthowe. Should we use pesticides to conserve rare plants? Conservation Biology, 14(5):1549–1550, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lackey:2007:SSP

[Lac07] Robert T. Lackey. Science, scientists, and policy advocacy. Conservation Biology, 21(1):12–17, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindsey:2005:CEW

[LADM05] P. A. Lindsey, R. Alexander, J. T. Du Toit, and M. G. L. Mills. The cost efficiency of wild dog conservation in South Africa. La rentabilidad de la conservación de perros salvajes en África del Sur. Conservation Biology, 19(4):1205–1214, August 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lafferty:2003:CDC

[Laf03] Kevin D. Lafferty. The crisis discipline of conservation medicine. Conservation Biology, 17(6):1859–1860, December 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laidlaw:2000:EHD

[Lai00] Ruth K. Laidlaw. Effects of habitat disturbance and protected areas on mammals of Peninsular Malaysia. Efectos de la perturbación del hábitat y las Áreas protegidas sobre los mamíferos de Malasia Peninsular. Conservation Biology, 14(6):1639–1648, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laiolo:2005:SSP

[Lai05] Paola Laiolo. Spatial and seasonal patterns of bird communities in Italian agroecosystems patrones espaciales y estacionales de comunidades de aves en agroecosistemas italianos. Conservation Biology, 19(5):1547–1556, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lake:2001:GF

[Lak01] Larry M. Lake. A generous facilitator. Conservation Biology, 15(3):811–813, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lambeck:2002:FSR

[Lam02] Robert J. Lambeck. Focal species and restoration ecology: Response to Lindenmayer et al. Conservation Biology, 16(2):549– 551, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lammertink:2004:MSC

[Lam04] Martjan Lammertink. A multiple-site comparison of woodpecker communities in Bornean lowland and hill forests. Una comparación multi-sitio de comunidades de pájaros carpinteros en bosques de tierras bajas y colinas en Borneo. Conservation Biology, 18(3):746–757, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lamm:2006:CGC

[Lam06] Richard D. Lamm. The culture of growth and the culture of limits. Conservation Biology, 20(2):269–271, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loot:2005:EHE

[LAN05] Géraldine Loot, Marcela Aldana, and Sergio A. Navarrete. Effects of human exclusion on parasitism in intertidal food webs of Central Chile. Efectos de exclusión humana sobre parasitismo en redes Tróficas intermareales del Centro de Chile. Conservation Biology, 19(1):203–212, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindholm:2001:MEF

[LARK01] James B. Lindholm, Peter J. Auster, Matthias Ruth, and Les Kaufman. Modeling the effects of fishing and implications for the design of marine protected areas: Juvenile fish responses to variations in seafloor habitat. Modelado de los efectos de la pesca y sus implicaciones para el diseño de areas marinas protegidas: Respuesta de peces juveniles a variaciones en el hábitat del fondo marino. Conservation Biology, 15(2):424–437, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Louda:2003:ISB

[LARR03] S. M. Louda, A. E. Arnett, T. A. Rand, and F. L. Russell. Invasiveness of some biological control insects and adequacy of their ecological risk assessment and regulation. Invasividad de algunos insectos de control biológico y adecuación de su evaluación de riesgo ecológico y regulación. Conservation Biology, 17 (1):73–82, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Latta:2000:MLR

[Lat00] Steven C. Latta. Making the leap from researcher to planner: Lessons from avian conservation planning in the Dominican Republic. Brincando de investigador a planeador: Lecciones de planes de conservación para la conservación de aves en la República Dominicana. Conservation Biology, 14(1):132–139, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lavigne:2003:GUC

[Lav03] Peter M. Lavigne. Greening the U.S. constitution. Conservation Biology, 17(6):1485–1486, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawler:2006:FCS

[Law06] Sharon P. Lawler. The Farmer and the conservationist should be friends: Finding common ground for invertebrates. Conservation Biology, 21(3):923–925, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawn:2008:MPG

[Law08] Philip Lawn. Macroeconomic policy, growth, and biodiversity conservation. Política Macroeconómica, crecimiento y conservación de la biodiversidad. Conservation Biology, 22(6):1418– 1423, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Louda:2000:ENE

[LB00] S. M. Louda and R. L. Bevill. Exclusion of natural enemies as a tool in managing rare plant species. Conservation Biology, 14(5): 1551–1552, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindholm:2001:CMT

[LB01] James Lindholm and Brad Barr. Comparison of marine and terrestrial protected areas under federal jurisdiction in the United States. Comparaciones de areas marinas y terrestres protegidas bajo jurisdicción federal en los Estados Unidos. Conservation Biology, 15(5):1441–1444, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lowe:2002:LLS

[LB02] Winsor H. Lowe and Douglas T. Bolger. Local and landscapescale predictors of salamander abundance in New Hampshire headwater streams. Predictores locales y a escala de paisaje de la abundancia de salamandras en arroyos de New Hampshire. Conservation Biology, 16(1):183–193, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Langlois:2005:MER

[LB05] Timothy J. Langlois and William J. Ballantine. Marine ecological research in New Zealand: Developing predictive models through the study of no-take marine reserves. Investigación ecológica marina en Nueva Zelanda: Desarrollo de modelos predictivos a través del estudio de reservas marinas sin captura. Conservation Biology, 19(6):1763–1770, December 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lemoine:2003:PIG

[LBG03] Nicole Lemoine and Katrin Böhning-Gaese. Potential impact of global climate change on species richness of long-distance migrants. Impacto potencial del cambio climático global sobre la riqueza de especies migratorias de larga distancia. Conservation Biology, 17(2):577–586, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lemoine:2007:ECL

[LBPBG07] Nicole Lemoine, Hans-Günther Bauer, Markus Peintinger, and Katrin Böhning-Gaese. Effects of climate and land-use change on species abundance in a Central European bird community. Efectos del clima y del cambio de uso de suelo sobre especies en una comunidad de aves en Europa Central. Conservation Biology, 21(2):495–503, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Larsen:2007:IPI

[LBR07] Frank Wugt Larsen, Jesper Bladt, and Carsten Rahbek. Improving the performance of indicator groups for the identification of important areas for species conservation. Mejoramiento del desempeño de grupos indicadores para la identificación de Áreas importantes para la conservación de especies. Conservation Biology, 21(3):731–740, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lyons:2005:RSE

[LBTS05] K. G. Lyons, C. A. Brigham, B. H. Traut, and M. W. Schwartz. Rare species and ecosystem functioning especies raras y funcionamiento del ecosistema. Conservation Biology, 19(4):1019– 1024, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2001:I

[LC01] William F. Laurance and Mark A. Cochrane. Introduction. Conservation Biology, 15(6):1488–1489, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewison:2007:PLB

[LC07a] Rebecca L. Lewison and Larry B. Crowder. Putting longline bycatch of sea turtles into perspective perspectivas de la captura incidental de tortugas marinas en palangres. Conservation Biology, 21(1):79–86, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lundquist:2007:BBG

[LC07b] Carolyn J. Lundquist and John A. Cigliano. Big blue gets its due. Conservation Biology, 21(2):560–561, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

List:2007:HDC

[LCC^{+}07] Rurik List, Gerardo Ceballos, Charles Curtin, Peter J. P. Gogan, Jesús Pacheco, and Joe Truett. Historic distribution and challenges to bison recovery in the Northern Chihuahuan desert. Distribución histórica y retos para la recuperación del bisonte en el norte del desierto chihuahuense. Conservation Biology, 21(6): 1487–1494, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2008:IRH

[LCG^{+}08] William F. Laurance, Barbara M. Croes, Nicaise Guissouegou, Ralph Buij, Marc Dethier, and Alfonso Alonso. Impacts of roads, hunting, and habitat alteration on nocturnal mammals in African rainforests. Impactos de carreteras, cacería y alteración del hábitat sobre mamíferos nocturnos de bosques lluviosos africanos. Conservation Biology, 22(3):721–732, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewison:2003:ITE

[LCS03] Rebecca L. Lewison, Larry B. Crowder, and Donna J. Shaver. The impact of turtle excluder devices and fisheries closures on loggerhead and Kemp’s ridley strandings in the Western Gulf of Mexico. El impacto de dispositivos excluyentes de tortugas y del cierre de pesquerías sobre varamientos de tortugas de Carey y de Kemp en el occidente del Golfo de México. Conservation Biology, 17(4):1089–1097, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovette:2004:LUM

[LCS04] Irby J. Lovette, Sonya M. Clegg, and Thomas B. Smith. Limited utility of mtDNA markers for determining connectivity among breeding and overwintering locations in three neotropical migrant birds. Utilidad limitada de los marcadores de ADNmt en la determinación de conectividad entre localidades de reproducción y de Hibernación de tres aves migratorias neotropicales. Conservation Biology, 18(1):156–166, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2006:IRH

[LCT^{+}06] William F. Laurance, Barbara M. Croes, Landry Tchignoumba, Sally A. Lahm, Alfonso Alonso, Michelle E. Lee, Patrick Campbell, and Claude Ondzeano. Impacts of roads and hunting on Central African rainforest mammals. Impactos de carreteras y cacería sobre mamíferos de bosque lluvioso en África Central. Conservation Biology, 33(5):1251–1261, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leal:2005:CCB

[LCTL05] Inara R. Leal, José Maria Cardoso Da Silva, Marcelo Tabarelli, and Thomas E. Lacher, Jr. Changing the course of biodiversity conservation in the caatinga of northeastern Brazil. Cambiando el curso de la conservación de biodiversidad en la caatinga del noreste de Brasil. Conservation Biology, 19(3):701–706, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2005:GDB

[LCZ^{+}05] Yuan-Yuan Li, Xiao-Yong Chen, Xin Zhang, Tian-Yi Wu, Hui-Ping Lu, and Yue-Wei Cai. Genetic differences between wild and artificial populations of Metasequoia glyptostroboides: Implications for species recovery. Diferencias genéticas entre poblaciones silvestres y artificiales de Metasequoia glyptostroboides: Implicaciones para la recuperación de especies. Conservation Biology, 19(1):224–231, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2009:RCL

[LCZX09] Yiqing Li, Yaning Chen, Yaoqi Zhang, and Yang Xia. Rehabilitating China’s largest inland river. Rehabilitación del río interior más extenso de China. Conservation Biology, 23(3):531–536, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levine:2003:FBI

[LD03] Jonathan M. Levine and Carla M. D’Antonio. Forecasting biological invasions with increasing international trade. Predicción de las invasiones biológicas con el incremento del comercio internacional. Conservation Biology, 17(1):322–326, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levine:2007:SIV

[LD07] Jonathan M. Levine and Carla M. D’Antonio. Subject Index, Volume 21, 2007. Conservation Biology, 17(1):1667–1673, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindzey:2006:UCC

[LDC06] Summer Lindzey, John Doudna, and Edward F. Connor. Updated conservation classic. Conservation Biology, 21(3):925–926, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Luijten:2000:PSG

[LDG^{+}00] Sheila H. Luijten, Angelo Dierick, J. Gerard, B. Oostermeijer, Léon E. L. Raijmann, and Hans C. M. Den Nijs. Population size, genetic variation, and reproductive success in a rapidly declining, self-incompatible perennial (Arnica montana) in The Netherlands. Tamaño poblacional, variación genética y exito reproductivo de una especie perenne en declinación rápida y autoincompatible (Arnica montana) en los Países Bajos. Conservation Biology, 14(6):1776–1787, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindborg:2002:EER

[LE02] Regina Lindborg and Johan Ehrlén. Evaluating the extinction risk of a perennial herb: Demographic data versus historical records. Evaluación del riesgo de extinción de una hierba perenne: Datos demográficos contra registros históricos. Conservation Biology, 16(3):683–690, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leach:2006:TAC

[Lea06] William D. Leach. Theories about consensus-based conservation. Conservation Biology, 20(2):573–575, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ledewitz:2006:PPE

[Led06] Bruce Ledewitz. Protecting posterity: Economics, abortion politics, and the law. Conservation Biology, 20(4):940–941, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lee:2003:HD

[Lee03] David W. Lee. Homage due. Conservation Biology, 17(4):1193– 1195, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lee:2009:CWU

[Lee09] Jared B. Lee. China’s wildness unveiled. Conservation Biology, 23(2):503–504, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leidy:2000:NFH

[Lei00] Robert A. Leidy. Nonindigenous fishes at home in inland waters of the United States. Conservation Biology, 14(1):328–330, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leidy:2005:MMR

[Lei05] Robert A. Leidy. Myth meets reality in Tasmania: the fall and rise of the pouched thing with a dog head. Conservation Biology, 19(4):1331–1333, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leidy:2009:ERC

[Lei09] Robert A. Leidy. Eloquence De rigueur for California deserts. Conservation Biology, 23(4):1060–1061, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leonard:2003:L

[Leo03] Patrick Leonard. Letters. Conservation Biology, 17(3):655–656, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leopold:2004:DEP

[Leo04] Donald J. Leopold. Deforesting the Earth: from prehistory to global crisis. Conservation Biology, 18(6):1693–1694, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levin:2002:IEF

[LEPH02] Phillip S. Levin, Julie Ellis, Rachel Petrik, and Mark E. Hay. Indirect effects of feral horses on estuarine communities. Efectos indirectos de caballos silvestres en comunidades de estuario. Conservation Biology, 16(5):1364–1371, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leslie:2005:SMC

[Les05] Heather M. Leslie. A synthesis of marine conservation planning approaches. Síntesis de estrategias de planificación de conservación marina. Conservation Biology, 19(6):1701–1713, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewis:2001:EES

[Lew01] Owen T. Lewis. Effect of experimental selective logging on tropical butterflies. Efectos de una tala selectiva experimental en mariposas tropicales. Conservation Biology, 15(2):389–400, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lopes:2000:EHC

[LF00] M. Aparecida Lopes and Stephen F. Ferrari. Effects of human colonization on the abundance and diversity of mammals in Eastern Brazilian Amazonia. Efectos de la colonización humana en la abundancia y diversidad de mamíferos en la zona oriental del Amazonas Brasileño. Conservation Biology, 14(6): 1658–1665, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Latta:2009:BSO

[LF09] Steven C. Latta and John Faaborg. Benefits of studies of overwintering birds for understanding resident bird ecology and promoting development of conservation capacity. Beneficios de los estudios de aves invernantes para entender la ecología de aves residentes y promover el desarrollo del potencial para conservar. Conservation Biology, 23(2):286–293, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindsey:2007:THC

[LFA^{+}07] Peter A. Lindsey, L. G. Frank, R. Alexander, A. Mathieson, and S. S. Romañach. Trophy hunting and conservation in Africa: Problems and one potential solution. Conservation Biology, 21 (2):880–883, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewinsohn:2005:CTI

[LFP05] Thomas M. Lewinsohn, André Victor Lucci Freitas, and Paulo Inácio Prado. Conservation of terrestrial invertebrates and their habitats in Brazil. Conservación de invertebrados terrestres y sus hábitats en Brasil. Conservation Biology, 19(3): 640–645, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lepczyk:2008:HIR

[LFR^{+}08] Christopher A. Lepczyk, Curtis H. Flather, Volker C. Radeloff, Anna M. Pidgeon, Roger B. Hammer, and Jianguo Liu. Human impacts on regional avian diversity and abundance impactos humanos sobre la diversidad y abundancia regional de aves. Conservation Biology, 22(2):405–416, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lafferty:2002:GMC

[LG02] Kevin D. Lafferty and Leah R. Gerber. Good medicine for conservation biology: the intersection of epidemiology and conservation theory. Buena medicina para la biología de la conservación: la Intersección de la Epidemiología y la teoría de conservación. Conservation Biology, 16(3):593–604, June 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [Jes03].

Labonne:2005:EPD

[LG05a] Jacques Labonne and Philippe Gaudin. Exploring population dynamics patterns in a rare fish, Zingel asper, through capturemark-recapture methods. Exploración de los patrones de la dinámica poblacional de un pez raro, Zingel asper, mediante métodos de captura-marca-recaptura. Conservation Biology, 19 (2):463–472, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lundquist:2005:SSM

[LG05b] Carolyn J. Lundquist and Elise F. Granek. Strategies for successful marine conservation: Integrating socioeconomic, political, and scientific factors. Estrategias para la conservación marina exitosa: Integración de factores socioeconómicos, políticos y científicos. Conservation Biology, 19(6):1771–1778, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lundquist:2005:SSI

[LGB05] Carolyn J. Lundquist, Elise F. Granek, and Rodrigo H. Bustamante. Special Section: Implementation and Management of Marine Protected Areas. Conservation Biology, 19(6):1699– 1700, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lohr:2002:ICW

[LGK02] Steven M. Lohr, Sidney A. Gauthreaux, and John C. Kilgo. Importance of coarse woody debris to avian communities in loblolly pine forests. Importancia de desechos leñosos gruesos para las comunidades de aves en bosques de Pinus taeda. Conservation Biology, 16(3):767–777, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lesica:2004:IEB

[LH04] Peter Lesica and David Hanna. Indirect effects of biological control on plant diversity vary across sites in Montana grasslands. Los efectos indirectos del control biológico sobre la diversidad de plantas Varían en pastizales de Montana. Conservation Biology, 18(2):444–454, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loiselle:2003:APU

[LHG^{+}03] Bette A. Loiselle, Christine A. Howell, Catherine H. Graham, Jaqueline M. Goerck, Thomas Brooks, Kimberly G. Smith, and Paul H. Williams. Avoiding pitfalls of using species distribution models in conservation planning. Evitando dificultades resultantes del uso de modelos de distribución de especies en planeación de conservación. Conservation Biology, 17(6):1591– 1600, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levin:2006:SPO

[LHPH06] Phillip S. Levin, Elizabeth E. Holmes, Kevin R. Piner, and Chris J. Harvey. Shifts in a Pacific Ocean fish assemblage: the potential influence of exploitation. Cambios en un ensamble de peces del Océano Pacífico: la influencia potencial de la explotación. Conservation Biology, 33(5):1181–1190, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lim:2006:HSJ

[Lim06] Tammy Lim. Herping in la selva just got easier. Conservation Biology, 20(2):590, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linklater:2003:NAT

[Lin03a] Wayne L. Linklater. A novel application of the Trivers–Willard model to the problem of genetic rescue. Una nueva aplicación del modelo Trivers–Willard al problema de rescate genético. Conservation Biology, 17(3):906–909, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linklater:2003:SMC

[Lin03b] Wayne L. Linklater. Science and management in a conservation crisis: a case study with rhinoceros. Ciencia y manejo en una crisis de conservación: Un estudio de caso con rinocerontes. Conservation Biology, 17(4):968–975, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

List:2004:SSN

[Lis04] Rurik List. Summary and synthesis for neotropical mammals. Conservation Biology, 18(1):288, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laikre:2008:WSC

[LJI^{+}08] Linda Laikre, Bengt-Gunnar Jonsson, Margareta Ihse, Mark Marissink, Ann-Marie Dock Gustavsson, Torbjörn Ebenhard, Lovisa Hagberg, Pär-Olof Stål, Susanne Von Walter, and Per Wramner. Wanted: Scientists in the CBD process. Conservation Biology, 22(4):814–815, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lu:2001:PGD

[LJMR^{+}01] Zhi Lu, Warren E. Johnson, Marilyn Menotti-Raymond, Naoya Yuhki, Janice S. Martenson, Susan Mainka, Huang Shi-Qiang, Zheng Zhihe, Guanghan Li, Wenshi Pan, Xiarong Mao, and Stephen J. O’Brien. Patterns of genetic diversity in remaining giant panda populations. Patrones de diversidad genética en poblaciones remanentes de panda gigante. Conservation Biology, 15(6):1596–1607, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levin:2004:PMC

[LK04] Phillip S. Levin and Beth F. Kochin. Publication of marine conservation papers: Is conservation biology too dry? Conservation Biology, 18(4):1160–1162, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lenth:2006:CVC

[LKG06] Buffy A. Lenth, Richard L. Knight, and Wendell C. Gilgert. Conservation value of clustered housing developments. Valor de conservación de desarrollos de vivienda agrupados. Conservation Biology, 20(5):1445–1456, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Larsen:2008:UTD

[LLF08] Trond H. Larsen, Alejandro Lopera, and Adrian Forsyth. Understanding trait-dependent community disassembly: Dung beetles, density functions, and forest fragmentation. Entendiendo el desensamble dependiente de atributos de la comunidad: Escarabajos del estiércol, funciones de la densidad y fragmentación de bosques. Conservation Biology, 22(5):1288–1298, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2008:LTD

[LLH08] William F. Laurance, Susan G. Laurance, and David W. Hilbert. Long-term dynamics of a fragmented rainforest mammal assemblage. Dinámica a largo plazo de un ensamble de mamíferos de un bosque lluvioso fragmentado. Conservation Biology, 22(5): 1154–1164, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Langholz:2000:IBC

[LLS00] Jeff Langholz, James Lassoie, and John Schelhas. Incentives for biological conservation: Costa Rica’s private wildlife refuge program. Incentivos para la conservación biológica: el programa de refugios privados de vida silvestre de Costa Rica. Conservation Biology, 14(6):1735–1743, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawson:2008:QPP

[LLS08] Dawn M. Lawson, Cerina K. Lamar, and Mark W. Schwartz. Quantifying plant population persistence in human-dominated landscapes. Cuantificación de la persistencia de poblaciones de plantas en paisajes dominados por humanos. Conservation Biology, 22(4):922–928, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2002:EDA

[LLV^{+}02] William F. Laurance, Thomas E. Lovejoy, Heraldo L. Vasconcelos, Emilio M. Bruna, Raphael K. Didham, Philip C. Stouffer, Claude Gascon, Richard O. Bierregaard, Susan G. Laurance, and Erica Sampaio. Ecosystem decay of Amazonian forest fragments: a 22-year investigation. Descomposición del ecosistema en fragmentos de bosque amazónico, una investigación de 22 años. Conservation Biology, 16(3):605–618, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lloyd:2002:CAR

[LLW02] Kelvin M. Lloyd, William G. Lee, and J. Bastow Wilson. Competitive abilities of rare and common plants: Comparisons using Acaena (Rosaceae) and Chionochloa (Poaceae) from New Zealand. Habilidades competitivas de plantas raras y comunes: Comparaciones utilizando Acaena (Rosaceae) y Chionochloa (Poaceae) de Nueva Zelanda. Conservation Biology, 16(4):975– 985, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lenzen:2009:ELU

[LLWCW09] Manfred Lenzen, Amanda Lane, Asaph Widmer-Cooper, and Moira Williams. Effects of land use on threatened species. Efectos del uso de suelo sobre especies amenazadas. Conservation Biology, 23(2):294–306, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leimu:2005:PHM

[LM05] Roosa Leimu and Pia Mutikainen. Population history, mating system, and fitness variation in a perennial herb with a fragmented distribution. Historia de la población, sistema de apareamiento y variación de adaptabilidad en una hierba perenne con distribución fragmentada. Conservation Biology, 19(2):349–356, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Light:2007:DBI

[LM07] Theo Light and Michael P. Marchetti. Distinguishing between invasions and habitat changes as drivers of diversity loss among California’s freshwater fishes. Distinguiendo entre invasiones y cambios de hábitat como conductores de la pérdida de diversidad en peces dulceacuícolas de California Light & Marchetti. Conservation Biology, 21(2):434–446, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:2000:IBE

[LMB00] David B. Lindenmayer, Chris R. Margules, and Daniel B. Botkin. Indicators of biodiversity for ecologically sustainable forest management. Indicadores de biodiversidad para el manejo ecológicamente sostenible de bosques. Conservation Biology, 14 (4):941–950, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lampila:2005:DRB

[LMD05] Petri Lampila, Mikko Mönkkönen, and André Desrochers. Demographic responses by birds to forest fragmentation. Respuestas demográficas de aves a la fragmentación de bosques. Conservation Biology, 19(5):1537–1546, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lantz:2008:TEK

[LME08] Van Lantz and Roberto Martínez-Espiñeira. Testing the Environmental Kuznets Curve Hypothesis with bird populations as habitat-specific environmental indicators: Evidence from Canada. Probando la hipótesis de la Curva Ambiental Kuznets con poblaciones de AES como indicadores ambientales de hábitats específico: Evidencia de Canadá. Conservation Biology, 22(2):428–438, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ling:2006:ASB

[LMG06] S. Ling and E. J. Milner-Gulland. Assessment of the sustainability of bushmeat hunting based on dynamic bioeconomic models. Evaluación de la sustentabilidad de la cacería basada en modelos bioeconómicos dinámicos. Conservation Biology, 33(5):1294– 1299, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lunney:2000:IHM

[LMMF00] Daniel Lunney, Alison Matthews, Chris Moon, and Simon Ferrier. Incorporating habitat mapping into practical koala conservation on private lands. Incorporación del levantamiento de mapas del hábitat en la conservación Práctica del koala en tierras privadas. Conservation Biology, 14(3):669–680, June 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawes:2000:POP

[LMP00] Michael J. Lawes, Peter E. Mealin, and Steven E. Piper. Patch occupancy and potential metapopulation dynamics of three forest mammals in fragmented afromontane forest in South Africa. Ocupación de parches y Dinámicas metapoblacionales potenciales de tres mamíferos del bosque en un bosque Afromontañoso fragmentado en Sudáfrica. Conservation Biology, 14(4):1088– 1098, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lobo:2002:SPM

[LMP02] Jorge M. Lobo and Fermín Martín-Piera. Searching for a predictive model for species richness of Iberian dung beetle based on spatial and environmental variables. En busca de un modelo predictivo para la riqueza de especies de escarabajos coprófilos ibéricos basado en variables espaciales y ambientales. Conservation Biology, 16(1):158–173, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lamoreux:2007:MNB

[LMR^{+}07] John F. Lamoreux, John C. Morrison, Taylor H. Ricketts, David M. Olson, Eric Dinerstein, Meghan W. Mcknight, and Herman H. Shugart. The multifaceted nature of biodiversity conservation: Reply to Leroux and Schmiegelow. Conservation Biology, 21(1):269–270, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [LS07].

Lindenmayer:2002:FSA

[LMS^{+}02] D. B. Lindenmayer, A. D. Manning, P. L. Smith, H. P. Possingham, J. Fischer, I. Oliver, and M. A. McCarthy. The focal-species approach and landscape restoration: a critique. El planteamiento de especie focal y restauración de paisaje: una crítica. Conservation Biology, 16(2):338–345, April 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:2006:SLE

[LN06] D. B. Lindenmayer and R. F. Noss. Salvage logging, ecosystem processes, and biodiversity conservation. Cosecha de salvamento, procesos ecológicos y conservación de la biodiversidad. Conservation Biology, 20(4):949–958, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laikre:2009:IGI

[LNP^{+}09] Linda Laikre, Torbjörn Nilsson, Craig R. Primmer, Nils Ryman, and Fred W. Allendorf. Importance of genetics in the interpretation of favourable conservation status. Conservation Biology, 23 (6):1378–1381, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Louda:2002:UEE

[LO02] Sváa M. Louda and Charles W. O’Brien. Unexpected ecological effects of distributing the exotic weevil, Larinus planus (F.), for the biological control of Canada thistle. Efectos ecológicos inesperados de la distribución del gorgojo Exótico, Larinus planus (F.), para el control biológico del cardo de Canadá. Conservation Biology, 16(3):717–727, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:2006:SLM

[LO06] D. B. Lindenmayer and K. Ough. Salvage logging in the montane ash eucalypt forests of the central highlands of Victoria and its potential impacts on biodiversity cosecha de salvamento en los bosques montanos de eucalipto del altiplano central de Victoria y sus impactos potenciales sobre la biodiversidad. Conservation Biology, 20(4):1005–1015, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Logiudice:2003:TTP

[Log03] Kathleen Logiudice. Trophically transmitted parasites and the conservation of small populations: Raccoon roundworm and the imperiled allegheny woodrat. Parásitos transmitidos Tróficamente y la conservación de poblaciones Pequeñas: el ascárido de los mapaches y la rata de la madera de allegheny en peligro de extinción. Conservation Biology, 17(1):258–266, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leathwick:2003:EDC

[LOM03] J. R. Leathwick, J. Mcc Overton, and M. Mcleod. An environmental domain classification of New Zealand and its use as a tool for biodiversity management. Una clasificación ambiental de Nueva Zelanda y su utilización como una herramienta para el manejo de biodiversidad. Conservation Biology, 17(6): 1612–1623, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovins:2002:DFC

[Lov02] L. Hunter Lovins. Design failure and conservation. Conservation Biology, 16(2):292–293, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:2004:DI

[Lov04] Thomas E. Lovejoy. Dreaming the impossible. Conservation Biology, 18(6):1465, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:2005:SB

[Lov05] Thomas E. Lovejoy. Spotlight on Brazil. Conservation Biology, 19(3):587–588, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:2006:GCB

[Lov06] Thomas E. Lovejoy. Glimpses of conservation biology, act II. Conservation Biology, 20(3):711–712, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lopez:2001:LPD

[LP01] Jorge E. Lopez and Catherine A. Pfister. Local population dynamics in metapopulation models: Implications for conservation. Dinámica de poblaciones locales en modelos de metapoblaciones: Implicaciones para la conservación. Conservation Biology, 15(6):1700–1709, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewinsohn:2005:HMS

[LP05] Thomas M. Lewinsohn and Paulo Inácio Prado. How many species are there in Brazil? ‘cuantas especies hay en brasil? Conservation Biology, 19(3):619–624, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lomolino:2007:BSV

[LP07] Mark V. Lomolino and David R. Perault. Body size variation of mammals in a fragmented, temperate rainforest. Variación del tamaño del cuerpo de mamíferos en un bosque lluvioso templado, fragmentado. Conservation Biology, 21(4):1059–1069, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lees:2008:CVR

[LP08] Alexander C. Lees and Carlos A. Peres. Conservation value of remnant riparian forest corridors of varying quality for Amazonian birds and mammals. El valor de conservación de corredores forestales riparios remanentes con calidad variable para aves y mamíferos amazónicos. Conservation Biology, 22(2):439–449, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lockwood:2002:MAT

[LRG^{+}02] Julie L. Lockwood, Gareth J. Russell, John L. Gittleman, Curtis C. Daehler, Michael L. McKinney, and Andy Purvis. A metric for analyzing taxonomic patterns of extinction risk. Un sistema métrico para analizar los patrones taxonómicos de riesgo de extinción. Conservation Biology, 16(4):1137–1142, August 2002.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lee:2008:IMI

[LRO^{+}08] Henry Lee II, Deborah A. Reusser, Julian D. Olden, Scott S. Smith, Jim Graham, Virginia Burkett, Jeffrey S. Dukes, Robert J. Piorkowski, and John Mcphedran. Integrated monitoring and information systems for managing aquatic invasive species in a changing climate. Sistemas integrados de monitoreo e información para el manejo de especies acuáticas invasoras en un clima cambiante. Conservation Biology, 22(3):575–584, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Longcore:2009:CAC

[LRS09a] Travis Longcore, Catherine Rich, and Lauren M. Sullivan. Critical assessment of claims regarding management of feral cats by trap–neuter–return. Evaluación crítica de las demandas relacionadas con el manejo de gatos ferales en programas de captura–esterilización–liberación. Conservation Biology, 23 (4):887–894, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lopez:2009:JEI

[LRS^{+}09b] Sébastien Lopez, François Rousset, Frank H. Shaw, Ruth G. Shaw, and Ophélie Ronce. Joint effects of inbreeding and local adaptation on the evolution of genetic load after fragmentation. Efectos conjuntos de la endogamia y la adaptación local sobre la evolución de la carga genética después de la fragmentación. Conservation Biology, 23(6):1618–1627, December 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lips:2003:ETP

[LRW03] Karen R. Lips, John D. Reeve, and Lani R. Witters. Ecological traits predicting amphibian population declines in Central America. Atributos ecológicos predicen declinaciones poblacionales de anfibios en Centro América. Conservation Biology, 17(4):1078–1088, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lejeune:2001:CSF

[LS01] Katherine D. Lejeune and Timothy R. Seastedt. Centaurea species: the forb that won the West. Especies de Centaurea: la hierba que conquistó el Oeste. Conservation Biology, 15(6): 1568–1574, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lowry:2003:CGB

[LS03] Porter P. Lowry II and Paul P. Smith. Closing the Gulf between botanists and conservationists. Conservation Biology, 17 (4):1175–1176, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Louda:2004:DES

[LS04] Svata M. Louda and Peter Stiling. The double-edged sword of biological control in conservation and restoration. Conservation Biology, 18(1):50–53, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Hod04b].

Leroux:2007:BCI

[LS07] Shawn J. Leroux and Fiona K. A. Schmiegelow. Biodiversity concordance and the importance of endemism. Conservation Biology, 21(1):266–268, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [LMR^{+}07].

Larsson:2009:PMR

[LS09] Mattias C. Larsson and Glenn P. Svensson. Pheromone monitoring of rare and threatened insects: Exploiting a pheromone– kairomone system to estimate prey and predator abundance. Monitoreo de feromonas de insectos raros y amenazados: Explotación de un sistema feromona-kairomona para estimar la abundancia de presas y depredadores. Conservation Biology, 23 (6):1516–1525, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loeser:2007:IGI

[LSC07] Matthew R. R. Loeser, Thomas D. Sisk, and Timothy E. Crews. Impact of grazing intensity during drought in an Arizona grassland. Impacto de la intensidad de pastoreo durante la sequía en un pastizal de Arizona. Conservation Biology, 21(1):87–97, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lodge:2003:NSE

[LSF03] David M. Lodge and Kristin Shrader-Frechette. Nonindigenous species: Ecological explanation, environmental ethics, and public policy. Especies no nativas: Explicación ecológica, Ética ambiental y política Pública. Conservation Biology, 17(1):31–37, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lodge:2004:DHP

[LSF04] David M. Lodge and Kristin Shrader-Frechette. Deep-history perspective on biological invasions: Response to Donlan and Martin. Conservation Biology, 18(1):270–271, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2004:ERC

[LSL04] Susan G. W. Laurance, Philip C. Stouffer, and William F. Laurance. Effects of road clearings on movement patterns of understory rainforest birds in Central Amazonia. Efectos de los claros de caminos sobre patrones de movimiento de aves de sotobosque en la Amazonía Central. Conservation Biology, 18(4): 1099–1109, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lehtila:2006:HCD

[LSL^{+}06] Kari Lehtilä, Kimmo Syrjänen, Roosa Leimu, Maria Begoña Garcia, and Johan Ehrlén. Habitat change and demography of Primula veris: Identification of management targets. Cambio de hábitat y demografía de Primula veris: Identificación de objetivos de conservación. Conservation Biology, 21(3):833–843, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leroux:2007:PSO

[LSN07] Shawn J. Leroux, Fiona K. A. Schmiegelow, and John A. Nagy. Potential spatial overlap of heritage sites and protected areas in a boreal region of Northern Canada. Potencial traslape espacial de sitios de patrimonio histórico y Áreas protegidas en una Región boreal del norte de Canadá. Conservation Biology, 21(2):376–386, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linacre:2004:ICD

[LSOBA04] Nicholas A. Linacre, Allan Stewart-Oaten, Mark A. Burgman, and Peter K. Ades. Incorporating collateral data in conservation biology. Incorporando datos colaterales a la biología de la conservación. Conservation Biology, 18(3):768–774, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linkie:2008:EBC

[LSZ^{+}08] Matthew Linkie, Robert J. Smith, Yu Zhu, Deborah J. Martyr, Beth Suedmeyer, Joko Pramono, and Nigel Leader-Williams. Evaluating biodiversity conservation around a large Sumatran protected area. Evaluación de la conservación de la biodiversidad alrededor de una Área protegida extensa en Sumatra. Conservation Biology, 22(3):683–690, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:2005:TC

[LT05] D. B. Lindenmayer and C. R. Tambiah. 2004 tsunami cleanup. Conservation Biology, 19(4):991, August 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawrence:2008:NGA

[LTC^{+}08] Hayley A. Lawrence, Graeme A. Taylor, David E. Crockett, Craig D. Millar, and David M. Lambert. New genetic approach to detecting individuals of rare and endangered species. Un nuevo método para detectar individuos de especies raras y en peligro. Conservation Biology, 22(5):1267–1276, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2009:ESE

[LU09] William F. Laurance and Diana C. Useche. Environmental synergisms and extinctions of tropical species sinergismos ambientales y extinciones de especies tropicales. Conservation Biology, 23(6):1427–1437, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lusseau:2003:ETB

[Lus03] David Lusseau. Effects of tour boats on the behavior of bottlenose dolphins: Using Markov chains to model anthropogenic impacts. Efectos de lanchas de turismo sobre el comportamiento de delfines Tursiops: Utilización de cadenas de Markov para modelar impactos antropogénicos. Conservation Biology, 17(6): 1785–1793, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lourie:2004:UBH

[LV04] Sara A. Lourie and Amanda C. J. Vincent. Using biogeography to help set priorities in marine conservation. Uso de la biogeografía como ayuda para definir prioridades en conservación marina. Conservation Biology, 18(4):1004–1020, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lipow:2004:GAC

[LVBS^{+}04] Sara R. Lipow, Kenneth Vance-Borland, J. Bradley St. Clair, Jan Henderson, and Cindy Mccain. Gap analysis of conserved genetic resources for forest trees. Análisis de claros de recursos genéticos conservados para Árboles de bosque. Conservation Biology, 18(2):412–423, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lens:2002:FAE

[LVM02] Luc Lens, Stefan Van Dongen, and Erik Matthysen. Fluctuating asymmetry as an early warning system in the critically endangered taita thrush. Fluctuación de asimetría como un sistema de alerta temprana en el tordo taita críticamente en peligro. Conservation Biology, 16(2):479–487, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2001:PFA

[LW01] William F. Laurance and G. Bruce Williamson. Positive feedbacks among forest fragmentation, drought, and climate change in the Amazon. Retroalimentaciones positivas entre fragmentación de bosques, sequía y cambio climático en el Amazonas. Conservation Biology, 15(6):1529–1535, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levin:2002:IEA

[LW02a] Phillip S. Levin and John G. Williams. Interspecific effects of artifically propagated fish: an additional conservation risk for salmon. Efectos interespecíficos de peces propagados artificialmente: un riesgo adicional para la conservación del salmón. Conservation Biology, 16(6):1581–1587, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lyman:2002:LPE

[LW02b] R. Lee Lyman and Steve Wolverton. The late prehistoric–early historic game sink in the Northwestern United States. El sumidero de especies de caza del prehistórico tardío/histórico temprano del Noroeste de los Estados Unidos. Conservation Biology, 16(1):73–85, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Langen:2006:EPB

[LW06] Tom A. Langen and Rick Welsh. Effects of a problem-based learning approach on attitude change and science and policy content knowledge. Conservation Biology, 20(3):600–608, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:2009:I

[LW09] William F. Laurance and S. Joseph Wright. Introduction. Conservation Biology, 23(6):1382–1385, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lu:2009:STN

[LWL09] Tao Lu, Ning Wu, and Peng Luo. Sedentarization of Tibetan nomads. Conservation Biology, 23(5):1074, October 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawler:2003:RSU

[LWSM03] Joshua J. Lawler, Denis White, Jean C. Sifneos, and Lawrence L. Master. Rare species and the use of indicator groups for conservation planning. Especies raras y el uso de grupos indicadores para la planeación de la conservación. Conservation Biology, 17(3):875–882, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lotts:2004:RAR

[LWV04] Kelly C. Lotts, Thomas A. Waite, and John A. Vucetich. Reliability of absolute and relative predictions of population persistence based on time series confiabilidad de predicciones absolutas y relativas de la persistencia de poblaciones basadas en series de tiempo. Conservation Biology, 18(5):1224–1232, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lynch:2006:IRF

[Lyn06] Tim P. Lynch. Incorporation of recreational fishing effort into design of marine protected areas. Incorporación del esfuerzo de pesca recreativa en el diseño de Áreas marinas protegidas. Conservation Biology, 20(5):1466–1476, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lynch:2008:DBS

[Lyn08] Tim P. Lynch. The difference between spatial and temporal variation in recreational fisheries for planning of marine protected areas: Response to Steffe. Conservation Biology, 22(2):486–491, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loucks:2003:GPQ

[LZD^{+}03] Colby J. Loucks, Lü Zhi, Eric Dinerstein, Wang Dajun, Fu Dali, and Wang Hao. The giant pandas of the Qinling Mountains, China: a case study in designing conservation landscapes for elevational migrants. Los pandas gigantes de las Montañas Qinling, China: un estudio de caso de diseño de paisajes de conservación para migrantes altitudinales. Conservation Biology, 17 (2):558–565, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mccoy:2005:FOL

[MA05] Nicole Haynes Mccoy and Sheldon Atwood. Flaws in Orr’s laws (and the paradigm that produced them): an abbreviated response. Conservation Biology, 19(4):1318–1320, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2008:PGD

[MA08] Karen J. Miller and David J. Ayre. Protection of genetic diversity and maintenance of connectivity among reef corals within marine protected areas. Protección de la diversidad genética y mantenimiento de la conectividad entre arrecifes de coral en Áreas marinas protegidas. Conservation Biology, 22(5):1245– 1254, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Madhusudan:2005:GVL

[Mad05] M. D. Madhusudan. The Global Village: Linkages between international coffee markets and grazing by livestock in a South Indian wildlife reserve. La Aldea Global: Conexiones entre mercados internacionales de café y pastoreo ilegal por ganado en una reserva de vida silvestre del Sur de India. Conservation Biology, 19(2):411–420, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maehr:2002:IGL

[Mae02] David S. Maehr. International guest lectures promote conservation learning. Conservation Biology, 16(5):1179–1183, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maehr:2004:PP

[Mae04] David S. Maehr. A plague of puma? Conservation Biology, 18 (4):1166–1168, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Magnusson:2001:CBU

[Mag01] William E. Magnusson. Catchments as basic units of management in conservation biology courses. Conservation Biology, 15 (5):1464–1465, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Magnusson:2004:EPT

[Mag04] William E. Magnusson. Ecoregion as a pragmatic tool. Conservation Biology, 18(1):4–5, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Main:2004:MGR

[Mai04] Martin B. Main. Mobilizing grass-roots conservation education: the Florida master naturalist program. Conservation Biology, 18 (1):11–17, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Male:2003:PIW

[Mal03] Timothy Male. Potential impact of West Nile virus on American avifaunas. Conservation Biology, 17(3):928–930, June 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mishra:2003:RIP

[MAM^{+}03] Charudutt Mishra, Priscilla Allen, Tom Mccarthy, M. D. Madhusudan, Agvaantserengiin Bayarjargal, and Herbert H. T. Prins. The role of incentive programs in conserving the snow leopard. El papel de programas de incentivos en la conservación del Uncia uncia. Conservation Biology, 17(6):1512–1520, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mangel:2002:IRT

[Man02] Marc Mangel. The important role of theory in conservation biology. Conservation Biology, 16(3):843–844, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCarthy:2003:RRP

[MAP03] Michael A. McCarthy, Sandy J. Andelman, and Hugh P. Possingham. Reliability of relative predictions in population viability analysis. Confiabilidad de predicciones relativas en el análisis de viabilidad poblacional. Conservation Biology, 17(4):982–989, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Markowitz:2000:SSR

[Mar00a] Hal Markowitz. Summary and synthesis in the realm of marine mammals. Conservation Biology, 14(3):916–918, June 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marzluff:2000:DA

[Mar00b] John M. Marzluff. Documenting the apocalypse. Conservation Biology, 14(4):1219–1220, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martell:2002:WRB

[Mar02a] David L. Martell. Wildfire regime in the boreal forest. Conservation Biology, 16(5):1177, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marzluff:2002:FCC

[Mar02b] John M. Marzluff. Fringe conservation: a call to action. Conservation Biology, 16(5):1175–1176, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marvier:2004:BAS

[Mar04] Michelle A. Marvier. A bridge to advanced statistical techniques. Conservation Biology, 18(3):854–855, June 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marty:2005:ECG

[Mar05] Jaymee T. Marty. Effects of cattle grazing on diversity in ephemeral wetlands. Efectos del apacentamiento de ganado sobre la diversidad en humedales Efímeros. Conservation Biology, 19(5):1626–1632, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mascia:2003:HDC

[Mas03] Michael B. Mascia. The human dimension of coral reef marine protected areas: Recent social science research and its policy implications. Conservation Biology, 17(2):630–632, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maddock:2000:ICW

[MB00] Ant Maddock and Grant A. Benn. Identification of conservationworthy areas in Northern Zululand, South Africa. Identificación de areas meritorias de conservación en la región norte de Zululand, Sudáfrica. Conservation Biology, 14(1):155–166, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McLachlan:2001:RPU

[MB01] Stephane M. McLachlan and Dawn R. Bazely. Recovery patterns of understory herbs and their use as indicators of deciduous forest regeneration. Recuperación de patrones del sotobosque Herbáceo y su uso como indicadores de regeneración del bosque deciduo. Conservation Biology, 15(1):98–110, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mace:2007:BIC

[MB07] Georgina M. Mace and Jonathan E. M. Baillie. The 2010 biodiversity indicators: Challenges for science and policy. Los indicadores de biodiversidad al 2010: Retos para la ciencia y la política. Conservation Biology, 21(6):1406–1413, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2009:FES

[MB09a] Laura J. Martin and Bernd Blossey. A framework for ecosystem services valuation. Conservation Biology, 23(2):494–496, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morrison:2009:CCS

[MB09b] Scott A. Morrison and Walter M. Boyce. Conserving connectivity: Some lessons from mountain lions in Southern California. Conservación de la conectividad: Algunas lecciones de pumas en el Sur de California. Conservation Biology, 23(2):275–285, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moore:2003:PSS

[MBB^{+}03] Joslin L. Moore, Andrew Balmford, Thomas Brooks, Neil D. Burgess, Louis A. Hansen, Carsten Rahbek, and Paul H. Williams. Performance of sub-Saharan vertebrates as indicator groups for identifying priority areas for conservation. La efectividad de los vertebrados del sub-Sahara como grupos indicadores para identificar Áreas prioritarias para la conservación. Conservation Biology, 17(1):207–218, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mumby:2008:CRH

[MBB^{+}08] Peter J. Mumby, Kenneth Broad, Daniel R. Brumbaugh, Craig, P. Dahlgren, Alastair R. Harborne, Alan Hastings, Katherine E. Holmes, Carrie V. Kappel, Fiorenza Micheli, and James N. Sanchirico. Coral reef habitats as surrogates of species, ecological functions, and ecosystem services. Hábitats de arrecifes de coral como sustitutos de especies, funciones ecológicas y servicios de ecosistemas. Conservation Biology, 22(4):941–951, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Monson:2003:CIC

[MBC03] Clark S. Monson, Sandra Anne Banack, and Paul Alan Cox. Conservation implications of chamorro consumption of flying foxes as a possible cause of amyotrophic lateral sclerosis– Parkinsonism dementia complex in Guam. Implicaciones para la conservación del consumo de zorros voladores por chamorros como causa posible del complejo demencia Parkinsoniana– esclerosis lateral amiotrófica en Guam. Conservation Biology, 17(3):678–686, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mascia:2003:CSS

[MBD^{+}03] Michael B. Mascia, J. Peter Brosius, Tracy A. Dobson, Bruce C. Forbes, Leah Horowitz, Margaret A. McKean, and Nancy J. Turner. Conservation and the social sciences. Conservation Biology, 17(3):649–650, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mace:2009:RCA

[MBFT09] Georgina Mace, Luigi Boitani, Erica Fleishman, and Alan D. Thornhill. Response to Carroll et al. Conservation Biology, 23 (2):246, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcmahon:2007:AHR

[MBH07] Clive R. Mcmahon, Corey J. A. Bradshaw, and Graeme C. Hays. Applying the heat to research techniques for species conservation. Conservation Biology, 21(1):271–273, February 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Manel:2002:DWP

[MBL02] Stéphanie Manel, Pierre Berthier, and Gordon Luikart. Detecting wildlife poaching: Identifying the origin of individuals with Bayesian assignment tests and multilocus genotypes. Detección de la caza ilegal de vida silvestre: Identificación del origen de individuos con pruebas de asignación Bayesiana y genotipos multilocus. Conservation Biology, 16(3):650–659, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Macdougall:2004:DCS

[MBM04] Andrew S. Macdougall, Brenda R. Beckwith, and Carrina Y. Maslovat. Defining conservation strategies with historical perspectives: a case study from a degraded oak grassland ecosystem. Definición de estrategias de conservación con perspectivas históricas: un estudio de caso de un ecosistema degradado de pastizal y roble. Conservation Biology, 18(2):455–465, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marini:2009:PCD

[MBMLJ09] Miguel Ângelo Marini, Morgane Barbet-Massin, Leonardo Esteves Lopes, and Frédéric Jiguet. Predicted climate-driven bird distribution changes and forecasted conservation conflicts in a neotropical savanna. Cambios previstos en la distribución de aves debido al cambio climático y pronóstico de conflictos de conservación en una sabana neotropical. Conservation Biology, 23 (6):1558–1567, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mclean:2005:NST

[MBQ05] Jennifer E. Mclean, Paul Bentzen, and Thomas P. Quinn. Nonrandom, size- and timing-biased breeding in a hatchery population of steelhead trout. Reproducción no aleatoria, sesgada por tamaño y tiempo en una población de cultivo de la trucha arcoiris. Conservation Biology, 19(2):446–454, April 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maschinski:2006:UPV

[MBQAM06] Joyce Maschinski, Joanne E. Baggs, Pedro F. Quintana-Ascencio, and Eric S. Menges. Using population viability analysis to predict the effects of climate change on the extinction risk of an endangered limestone endemic shrub, Arizona cliffrose. Utilización de AVP para predecir los efectos del cambio climático sobre el riesgo de extinción de Purshia subintegra, un arbusto Endémico en peligro. Conservation Biology, 20(1): 218–228, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Majluf:2002:CBS

[MBR^{+}02] Patricia Majluf, Elizabeth A. Babcock, Juan Carlos Riveros, Milena Arias Schreiber, and William Alderete. Catch and bycatch of sea birds and marine mammals in the small-scale fishery of Punta San Juan, Peru. Captura accidental de aves y mamíferos marinos en la pesquería de pequeña escala en Punta San Juan, Perú. Conservation Biology, 16(5):1333–1343, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minteer:2005:EEB

[MC05a] Ben A. Minteer and James P. Collins. Ecological ethics: Building a new tool kit for ecologists and biodiversity managers. Ética ecológica: Construcción de un juego de herramientas nuevo para ecologistas y gestores de biodiversidad. Conservation Biology, 19 (6):1803–1812, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moilanen:2005:VUE

[MC05b] Atte Moilanen and Mar Cabeza. Variance and uncertainty in the expected number of occurrences in reserve selection. Variación e incertidumbre en el número esperado de ocurrencias en la selección de reservas. Conservation Biology, 19(5):1663–1667, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minteer:2006:RR

[MC06] Ben A. Minteer and James P. Collins. Response to Rominger. Conservation Biology, 20(5):1341, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mascia:2009:PRA

[MC09] Michael B. Mascia and C. Anne Claus. A property rights approach to understanding human displacement from protected areas: the case of Marine Protected Areas. Desplazamiento humano y Áreas marinas protegidas: una estrategia de Derechos de Propiedad. Conservation Biology, 23(1):16–23, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maunder:2000:CTT

[MCA^{+}00] Mike Maunder, Alastair Culham, Bjorn Alden, Georg Zizka, Cathérine Orliac, Wolfram Lobin, Alberto Bordeu, Jose M. Ramirez, and Sabine Glissmann-Gough. Conservation of the toromiro tree: Case study in the management of a plant extinct in the wild. Conservación del arbol toromiro, un caso de estudio en el manejo de una planta extinta en sitios silvestres. Conservation Biology, 14(5):1341–1350, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McBride:2000:AP

[McB00] Joe R. McBride. Americans in paradise. Conservation Biology, 14(3):910–911, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McBride:2002:ESE

[McB02] Joe R. McBride. Exploring the social and ethical scaffolds of ecological restoration. Conservation Biology, 16(3):845–847, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCosker:2000:GSL

[McC00] John E. McCosker. A glimpse of the slow lane. Conservation Biology, 14(1):331, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCarty:2001:ECR

[McC01] John P. McCarty. Ecological consequences of recent climate change. Consecuencias biológicas de cambios climáticos recientes. Conservation Biology, 15(2):320–331, April 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCullough:2003:BF

[McC03] Dale R. McCullough. Back to the future. Conservation Biology, 17(3):940–942, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mccallum:2005:ICA

[Mcc05] Hamish Mccallum. Inconclusiveness of chytridiomycosis as the agent in widespread frog declines falta de datos concluyentes de la quitridiomicosis como agente de las declinaciones globales de ranas. Conservation Biology, 19(5):1421–1430, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcclenachan:2009:DLL

[Mcc09] Loren Mcclenachan. Documenting loss of large trophy fish from the Florida Keys with historical photographs. Documentación de la pérdida de peces de trofeo en los Cayos de Florida con fotografías históricas. Conservation Biology, 23(3):636–643, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McDaniel:2003:ER

[McD03a] Carl N. McDaniel. Elephant in the room. Conservation Biology, 17(4):954–955, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McDaniel:2003:R

[McD03b] Jay McDaniel. Response. Conservation Biology, 17(4):947, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcdaniel:2004:HCI

[Mcd04] Carl N. Mcdaniel. The human cost of ideology as science. Conservation Biology, 18(4):869–871, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McDaniel:2009:WFC

[McD09] Carl N. McDaniel. When fantasy has consequences. Conservation Biology, 23(2):254–255, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Manolis:2009:LNF

[MCF^{+}09] Jim C. Manolis, Kai M. Chan, Myra E. Finkelstein, Scott Stephens, Cara R. Nelson, Jacqualine B. Grant, and Michael P. Dombeck. Leadership: a new frontier in conservation science. Liderazgo: una frontera nueva en ciencia de la conservación. Conservation Biology, 23(4):879–886, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mace:2008:QER

[MCG^{+}08] Georgina M. Mace, Nigel J. Collar, Kevin J. Gaston, Craig Hilton-Taylor, H. Resit Akçakaya, Nigel Leader-Williams, E. J. Milner-Gulland, and Simon N. Stuart. Quantification of extinction risk: IUCN’s system for classifying threatened species. Cuantificación del riesgo de extinción: Sistema de la UICN para la clasificación de especies amenazadas. Conservation Biology, 22(6):1424–1442, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcclanahan:2009:IRH

[MCG^{+}09] T. R. Mcclanahan, J. E. Cinner, N. A. J. Graham, T. M. Daw, J. Maina, S. M. Stead, A. Wamukota, K. Brown, V. Venus, and N. V. C. Polunin. Identifying reefs of hope and hopeful actions: Contextualizing environmental, ecological, and social parameters to respond effectively to climate change. Identificación de arrecifes de esperanza y acciones esperanzadoras: Contextualización de Parámetros ambientales, ecológicos y sociales para responder efectivamente al cambio climático. Conservation Biology, 23(3):662–671, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McKinney:2002:ENC

[McK02] Michael L. McKinney. Effects of national conservation spending and amount of protected area on species threat rates. Efectos del gasto nacional para la conservación y la cantidad de Áreas protegidas en los Índices de especies amenazadas. Conservation Biology, 16(2):539–543, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcgeoch:2006:GIB

[MCK06] Melodie A. Mcgeoch, Steven L. Chown, and Jesse M. Kalwij. A global indicator for biological invasion. Un indicador global para la invasión biológica. Conservation Biology, 20(6):1635–1646, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McKibben:2009:SCC

[McK09] Bill McKibben. Surviving climate change through mitigation and adaptation. Conservation Biology, 23(4):796, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Melzer:2000:OCA

[MCM^{+}00] Alistair Melzer, Frank Carrick, Peter Menkhorst, Daniel Lunney, and Barbara St. John. Overview, critical assessment, and conservation implications of koala distribution and abundance. Visión general, evaluación crítica y consecuencias para la conservación de la distribución y abundancia del koala. Conservation Biology, 14(3):619–628, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mulder:2007:RRE

[MCM07] Monique Borgerhoff Mulder, Tim Caro, and Omari Ayubu Msago. The role of research in evaluating conservation strategies in Tanzania: the case of the Katavi–Rukwa ecosystem. El papel de la investigación en la evaluación de estrategias de conservación en Tanzania: el caso del ecosistema Katavi-Rukwa. Conservation Biology, 21(3):647–658, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McNeal:2003:DCW

[McN03] Lyle G. McNeal. Domesticated casualties of war and technology. Conservation Biology, 17(2):644–645, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2004:ECM

[MCR^{+}04] Brian Miller, William Conway, Richard P. Reading, Chris Wemmer, David Wildt, Devra Kleiman, Steven Monfort, Alan Rabinowitz, Beth Armstrong, and Michael Hutchins. Evaluating the conservation mission of zoos, aquariums, botanical gardens, and natural history museums. Evaluación de la Misión de conservación de Zoológicos, acuarios, jardines botánicos o museos de historia natural. Conservation Biology, 18(1):86–93, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McShane:2003:DDB

[McS03] Thomas O. McShane. The Devil in the detail of biodiversity conservation. Conservation Biology, 17(1):1–3, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcsweeney:2005:IPG

[Mcs05] Kendra Mcsweeney. Indigenous population growth in the lowland neotropics: Social science insights for biodiversity conservation. Crecimiento de la población indígena en las tierras bajas neotropicales: Percepciones de las ciencias sociales para la conservación de biodiversidad. Conservation Biology, 19(5):1375– 1384, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcsweeney:2006:DIR

[Mcs06] Kendra Mcsweeney. Distinguishing the indigenous from the rural poor: Reply to Engelman et al. Conservation Biology, 33(5): 1318–1320, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meier:2004:SSD

[MD04] Rudolf Meier and Torsten Dikow. Significance of specimen databases from taxonomic revisions for estimating and mapping the global species diversity of invertebrates and repatriating reliable specimen data. Significado de bases de datos de especimenes de revisiones taxonómicas para la estimación y el mapeo de la diversidad global de especies de invertebrados y de la repatriación de datos confiables de especimenes. Conservation Biology, 18 (2):478–488, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mandelik:2007:RHT

[MDCK07] Yael Mandelik, Tamar Dayan, Vladimir Chikatunov, and Vasiliy Kravchenko. Reliability of a higher-taxon approach to richness, rarity, and composition assessments at the local scale. Confiabilidad de un enfoque de taxón superior para evaluaciones de riqueza, rareza y composición a escala local. Conservation Biology, 21(6):1506–1515, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mandelik:2005:PBR

[MDF05] Yael Mandelik, Tamar Dayan, and Eran Feitelson. Planning for biodiversity: the role of ecological impact assessment. Planificación para la biodiversidad: El papel de la evaluación de impacto ecológico. Conservation Biology, 19(4):1254–1261, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marsh:2007:OAM

[MDH^{+}07] Helene Marsh, Andrew Dennis, Harry Hines, Alex Kutt, Keith Mcdonald, Ellen Weber, Stephen Williams, and John Winter. Optimizing allocation of management resources for wildlife. Optimización de la asignación de recursos para la gestión de vida silvestre. Conservation Biology, 21(2):387–399, April 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Means:2004:AFL

[MDJP04] D. Bruce Means, C. Kenneth Dodd, Jr., Steve A. Johnson, and John G. Palis. Amphibians and fire in longleaf pine ecosystems: Response to Schurbon and Fauth. Conservation Biology, 18(4): 1149–1153, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SF03] and reply [SF04].

Mittermeier:2005:BHB

[MDRB05] Russell A. Mittermeier, Gustavo A. B. Da Fonseca, Anthony B. Rylands, and Katrina Brandon. A brief history of biodiversity conservation in Brazil. Conservation Biology, 19(3):601– 607, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:2001:EAC

[MDW01] L. Scott Mills, Daniel F. Doak, and Michael J. Wisdom. Elasticity analysis for conservation decision making: Reply to Ehrlén et al. Conservation Biology, 15(1):281–283, February 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Montalvo:2000:TSL

[ME00] Arlee M. Montalvo and Norman C. Ellstrand. Transplantation of the subshrub Lotus scoparius: Testing the home-site advantage hypothesis. Transplante del arbusto Lotus scoparius: una prueba de la hipótesis de ubicación del hogar. Conservation Biology, 14 (4):1034–1045, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Medellin:2000:BDA

[MEA00] Rodrigo A. Medellín, Miguel Equihua, and Miguel A. Amin. Bat diversity and abundance as indicators of disturbance in neotropical rainforests. Diversidad y abundancia de murciélagos como indicadores de perturbaciones en selvas húmedas neotropicales. Conservation Biology, 14(6):1666–1675, December 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2001:CCB

[Mef01a] Gary K. Meffe. The context of conservation biology. Conservation Biology, 15(4):815–816, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2001:CCD

[Mef01b] Gary K. Meffe. Crisis in a crisis discipline. Conservation Biology, 15(2):303–304, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2001:U

[Mef01c] Gary K. Meffe. Unity. Conservation Biology, 15(6):1475–1476, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2002:GI

[Mef02] Gary K. Meffe. Going international. Conservation Biology, 16 (3):573–574, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2003:EA

[Mef03a] Gary K. Meffe. Editor’s acknowledgments. Conservation Biology, 17(6):1477, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2003:TFI

[Mef03b] Gary K. Meffe. Toward further internationalization. Conservation Biology, 17(5):1197–1199, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2006:SSA

[Mef06a] Gary K. Meffe. Special section: 20^{th} anniversary of Conservation Biology. Conservation Biology, 20(3):629–630, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2006:SCC

[Mef06b] Gary K. Meffe. The success — and challenges — of conservation biology. Conservation Biology, 20(4):931–933, August 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2007:CFP

[Mef07a] Gary K. Meffe. Conservation Focus: Policy Advocacy and Conservation Science. Conservation Biology, 21(1):11, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2007:PC

[Mef07b] Gary K. Meffe. The politics of carbon. Conservation Biology, 21(2):295–296, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2008:CFE

[Mef08] Gary K. Meffe. Conservation Focus: Economic Growth and Biodiversity Conservation. Conservation Biology, 22(6):1388, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2009:AL

[Mef09a] Gary K. Meffe. Ancestral lesson. Conservation Biology, 23(6): 1651–1654, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2009:C

[Mef09b] Gary K. Meffe. Changes. Conservation Biology, 23(6):1343– 1345, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2009:NUE

[Mef09c] Gary K. Meffe. The nearly unbearable end. Conservation Biology, 23(6):1356, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mehrhoff:2005:MEE

[Meh05] Loyal A. Mehrhoff. A means to an end, not an end in itself. Conservation Biology, 19(2):583–585, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meine:2001:RCR

[Mei01] Curt Meine. Roosevelt, conservation, and the revival of democracy. Conservation Biology, 15(4):829–831, August 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meine:2003:BGD

[Mei03] Curt Meine. Bio(graphical) diversity. Conservation Biology, 17 (4):1179–1186, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meine:2006:WRC

[Mei06] Curt Meine. For whom the roll calls. Conservation Biology, 20 (2):272–273, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Menges:2003:PVE

[Men03] Eric Menges. Population viability: Evolving tools for declining species. Conservation Biology, 17(2):638–639, April 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:2006:CBT

[MEN06] Gary K. Meffe, David Ehrenfeld, and Reed F. Noss. Conservation biology at twenty. Conservation Biology, 20(3):595–596, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Merrill:2004:ROS

[Mer04] Samuel B. Merrill. The role of open space in urban planning. Conservation Biology, 18(2):294, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Merriman:2008:ACN

[Mer08] Hardy Merriman. Agents of change and nonviolent action. Conservation Biology, 22(2):241–242, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meir:2000:WOE

[MF00] Eli Meir and William F. Fagan. Will observation error and biases ruin the use of simple extinction models? `?Se arruinará el uso de modelos simples de extinción por errores y sesgos de observación? Conservation Biology, 14(1):148–154, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mayer:2008:FQS

[MF08] F. Stephan Mayer and Cynthia McPherson Frantz. Framing the question of survival: Psychological insights and limitations. Conservation Biology, 22(4):823–825, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maiorano:2007:CNN

[MFGB07] Luigi Maiorano, Alessandra Falcucci, Edward O. Garton, and Luigi Boitani. Contribution of the Natura 2000 network to biodiversity conservation in Italy. Contribución de la red Natura 2000 a la conservación de la biodiversidad en Italia. Conservation Biology, 21(6):1433–1444, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marini:2005:BCB

[MG05] Miguel Ângelo Marini and Frederico Innecco Garcia. Bird conservation in Brazil. Conservación de aves en Brasil. Conservation Biology, 19(3):665–671, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mendez:2007:AFA

[MGB^{+}07] Martín Mendez, Andrés Gómez, Nora Bynum, Rodrigo Medellín, Ana L. Porzecanski, and Eleanor Sterling. Availability of formal academic programs in conservation biology in Latin America. Conservation Biology, 21(6):1399–1403, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mikusinski:2001:WIF

[MGC01] Grzegorz Mikusiński, Maciej Gromadzki, and Przemyslaw Chylarecki. Woodpeckers as indicators of forest bird diversity. Pájaros carpinteros como indicadores de la diversidad de aves forestales. Conservation Biology, 15(1):208–217, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2009:BCL

[MGH^{+}09] James R. Miller, Martha Groom, George R. Hess, Toddi Steelman, David L. Stokes, Jan Thompson, Troy Bowman, Laura Fricke, Brandon King, and Ryan Marquardt. Biodiversity conservation in local planning. Conservación de la biodiversidad en la planificación local. Conservation Biology, 23(1):53–63, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mckenzie:2004:CCW

[MGPM04] Donald Mckenzie, Ze’Ev Gedalof, David L. Peterson, and Philip Mote. Climatic change, wildfire, and conservation. Cambio climático, incendios y conservación. Conservation Biology, 18 (4):890–902, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maina:2000:IRC

[MH00] Godfrey G. Maina and Henry F. Howe. Inherent rarity in community restoration. Rareza inherente en la restauración de comunidades. Conservation Biology, 14(5):1335–1340, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mech:2001:EEC

[MH01a] Stephen G. Mech and James G. Hallett. Evaluating the effectiveness of corridors: a genetic approach. Evaluación de la efectividad de corredores: Una estrategia genética. Conservation Biology, 15(2):467–474, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morrison:2001:CVP

[MH01b] Joan L. Morrison and Stephen R. Humphrey. Conservation value of private lands for crested caracaras in Florida. Valor de tierras privadas para la conservación de quebrantahuesos en Florida. Conservation Biology, 15(3):675–684, June 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2002:CWP

[MH02] James R. Miller and Richard J. Hobbs. Conservation where people live and work. Conservación donde la gente vive y trabaja da planeación eficaz de la conservación. Conservation Biology, 16(2):330–337, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mueller:2008:AIR

[MH08] Jillian M. Mueller and Jessica J. Hellmann. An assessment of invasion risk from assisted migration. Una evaluación del riesgo de invasión por migración asistida. Conservation Biology, 22(3): 562–567, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Middleton:2001:EUR

[MHC^{+}01] Elizabeth M. Middleton, Jay R. Herman, Edward A. Celarier, John W. Wilkinson, Cynthia Carey, and Robert J. Rusin. Evaluating ultraviolet radiation exposure with satellite data at sites of amphibian declines in central and South America. Evaluación de la exposición a la radiación ultravioleta con datos de Satélite en sitios de declinaciones de anfibios en centro y sur América. Conservation Biology, 15(4):914–929, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Merenlender:2004:LTC

[MHGF04] A. M. Merenlender, L. Huntsinger, G. Guthey, and S. K. Fairfax. Land trusts and conservation easements: Who is conserving what for whom? Patronatos agrarios y servicios de conservación:

`?Quién Está conservando Qué para Quién? Conservation Biology, 18(1):65–76, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Macreadie:2009:FRE

[MHJ^{+}09] Peter I. Macreadie, Jeremy S. Hindell, Gregory P. Jenkins, Rod M. Connolly, and Michael J. Keough. Fish responses to experimental fragmentation of seagrass habitat. Respuestas de peces a la fragmentación experimental de hábitat de pasto marino. Conservation Biology, 23(3):644–652, June 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meuser:2009:PPE

[MHM09] Emily Meuser, Howard W. Harshaw, and Arne Ø. Mooers. Public preference for endemism over other conservation-related species attributes. Preferencia del público por el endemismo por encima de otros atributos de las especies relacionados con la conservación. Conservation Biology, 23(4):1041–1046, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mclachlan:2007:FDA

[MHS07] Jason S. Mclachlan, Jessica J. Hellmann, and Mark W. Schwartz. A framework for debate of assisted migration in an era of climate change. Conservation Biology, 21(2):297–302, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Michelangeli:2008:CEF

[Mic08] Fabián A. Michelangeli. Cataloguing the endemic flora of Peru. Conservation Biology, 22(2):501–503, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Migliaccio:2000:LC

[Mig00] Christopher P. Migliaccio. Life in the clouds. Conservation Biology, 14(3):918–920, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minnich:2001:IMT

[Min01] Richard A. Minnich. An integrated model of two fire regimes. Conservation Biology, 15(6):1549–1553, December 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moen:2003:EEL

[MJ03] Jon Moen and Bengt Gunnar Jonsson. Edge effects on liverworts and lichens in forest patches in a mosaic of boreal forest and wetland. Efectos de borde sobre Hepáticas y musgos en fragmentos de bosque en un mosaico de bosque boreal y humedales. Conservation Biology, 17(2):380–388, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mendel:2002:HPB

[MK02a] Louise C. Mendel and J. B. Kirkpatrick. Historical progress of biodiversity conservation in the protected-area system of Tasmania, Australia. Progreso histórico de la conservación de biodiversidad en el sistema de Áreas protegidas de Tasmania, Australia. Conservation Biology, 16(6):1520–1529, December 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morley:2002:ARH

[MK02b] Sarah A. Morley and James R. Karr. Assessing and restoring the health of urban streams in the Puget Sound Basin. Evaluación y restauración de la salud de arroyos urbanos en la Cuenca Puget Sound. Conservation Biology, 16(6):1498–1509, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muriel:2009:EPS

[MK09] Sandra B. Muriel and Gustavo H. Kattan. Effects of patch size and type of coffee matrix on ithomiine butterfly diversity and dispersal in cloud-forest fragments. Efectos del tamaño del fragmento y tipo de matriz de café sobre la diversidad y dispersión de mariposas ithomiinae en fragmentos de bosque de niebla. Conservation Biology, 23(4):948–956, August 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Musters:2001:BBF

[MKDT01] C. J. M. Musters, M. Kruk, H. J. De Graaf, and W. J. Ter Keurs. Breeding birds as a farm product aves reproductoras como un producto de granja. Conservation Biology, 15(2):363–369, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maestas:2003:BAR

[MKG03] Jeremy D. Maestas, Richard L. Knight, and Wendell C. Gilgert. Biodiversity across a rural land-use gradient biodiversidad a lo largo de un gradiente de uso de suelo rural. Conservation Biology, 17(5):1425–1434, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2007:IWD

[MKH07] Julien Martin, Wiley M. Kitchens, and James E. Hines. Importance of well-designed monitoring programs for the conservation of endangered species: Case study of the snail kite. Importancia de programas de monitoreo para la conservación de especies en peligro bien diseñados: Caso de estudio de Rostrhamus sociabilis plumbeus. Conservation Biology, 21(2):472–481, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mattila:2006:EDD

[MKK^{+}06] Niina Mattila, Veijo Kaitala, Atte Komonen, Janne S. Kotiaho, and Jussi Päivinen. Ecological determinants of distribution decline and risk of extinction in moths determinantes. Ecológicos de la Reducción en la distribución y del riesgo de extinción en mariposas nocturnas. Conservation Biology, 33(5):1161–1168, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mattila:2009:IBE

[MKK^{+}09] Niina Mattila, Janne S. Kotiaho, Veijo Kaitala, Atte Komonen, and Jussi Päivinen. Interactions between ecological traits and host plant type explain distribution change in noctuid moths. Las interacciones entre atributos ecológicos y tipo de planta huésped explican el cambio de distribución de mariposas nocturnas. Conservation Biology, 23(3):703–709, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcneely:2009:CBA

[MKVZ^{+}09] Jeffrey A. Mcneely, Promila Kapoor-Vijay, Lu Zhi, Linda Olsvig-Whittaker, Kashif M. Sheikh, and Andrew T. Smith. Conservation biology in Asia: the major policy challenges. Biología de la conservación en Asia: los retos políticos mayores. Conservation Biology, 23(4):805–810, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mulvey:2000:LAS

[ML00] Margaret Mulvey and Charles Lydeard. Let’s not abandon science for advocacy: Reply to Berg and Berg. Conservation Biology, 14(6):1924–1925, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [BB00].

McMillan:2002:ERC

[ML02] Michele A. McMillan and Douglas W. Larson. Effects of rock climbing on the vegetation of the Niagara Escarpment in Southern Ontario, Canada. Efectos del escalado de rocas sobre la vegetación del Escarpado del Niagara en el sur de Ontario, Canadá. Conservation Biology, 16(2):389–398, April 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marchand:2004:EHF

[ML04] Michael N. Marchand and John A. Litvaitis. Effects of habitat features and landscape composition on the population structure of a common aquatic turtle in a region undergoing rapid development. Efectos de las características del hábitat y la composición del paisaje sobre la estructura poblacional de una tortuga acuática común en un paisaje bajo desarrollo rápido. Conservation Biology, 18(3):758–767, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milder:2008:CBE

[MLB08] Jeffrey C. Milder, James P. Lassoie, and Barbara L. Bedford. Conserving biodiversity and ecosystem function through limited development: an empirical evaluation. Conservación de la biodiversidad y el funcionamiento del ecosistema por medio del desarrollo limitado: una evaluación empírica. Conservation Biology, 22(1):70–79, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin-Lopez:2008:EVB

[MLMB08] Berta Martín-López, Carlos Montes, and Javier Benayas. Economic valuation of biodiversity conservation: the meaning of numbers. Valoración económica de la conservación de la biodiversidad: el significado de los números. Conservation Biology, 22(3):624–635, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Malcolm:2006:GWE

[MLN^{+}06] Jay R. Malcolm, Canran Liu, Ronald P. Neilson, Lara Hansen, and Lee Hannah. Global warming and extinctions of endemic species from biodiversity hotspots. Calentamiento global y extinciones de especies endémicas en sitios de importancia para la biodiversidad. Conservation Biology, 20(2):538–548, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matter:2000:NLI

[MM00] William J. Matter and Guy R. McPherson. No lurking inconsistency. Conservation Biology, 14(4):1204–1205, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mattson:2002:EGB

[MM02a] David J. Mattson and Troy Merrill. Extirpations of grizzly bears in the contiguous United States, 1850–2000 extirpaciones de osos pardos en el territorio contiguo de los Estados Unidos, 1850– 2000. Conservation Biology, 16(4):1123–1136, August 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meyer:2002:UFL

[MM02b] C. B. Meyer and S. L. Miller. Use of fragmented landscapes by marbled murrelets for nesting in Southern Oregon. Uso de paisajes fragmentados para anidación por el Mérgulo marmoleado en el sur de Oregon. Conservation Biology, 16(3):755– 766, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Monkkonen:2003:OMB

[MM03a] Mikko Mönkkönen and Marko Mutanen. Occurrence of moths in boreal forest corridors presencia de mariposas nocturnas en corredores de bosques boreales. Conservation Biology, 17(2): 468–475, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Myers:2003:IAH

[MM03b] Norman Myers and Russell A. Mittermeier. Impact and acceptance of the hotspots strategy: Response to Ovadia and to Brummitt and Lughadha. Conservation Biology, 17(5):1449– 1450, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcloughlin:2004:RCS

[MM04] Philip D. Mcloughlin and François Messier. Relative contributions of sampling error in initial population size and vital rates to outcomes of population viability analysis. Contribuciones relativas del error de muestreo en el tamaño poblacional inicial y tasas vitales a los resultados de análisis de población poblacional. Conservation Biology, 18(6):1665–1669, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2007:ILG

[MM07] Tara G. Martin and S. Mcintyre. Impacts of livestock grazing and tree clearing on birds of woodland and riparian habitats. Impactos del pastoreo de ganado y la tala de Árboles sobre aves de hábitats boscosos y ribereños. Conservation Biology, 21(2): 504–514, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Musiani:2003:WDT

[MMB^{+}03] Marco Musiani, Charles Mamo, Luigi Boitani, Carolyn Callaghan, C. Cormack Gates, Livia Mattei, Elisabetta Visalberghi, Stewart Breck, and Giulia Volpi. Wolf depredation trends and the use of fladry barriers to protect livestock in Western North America. Tendencias en la depredación por lobos y el uso de barreras para proteger ganado en Norte América occidental. Conservation Biology, 17(6):1538–1547, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcdonald-Madden:2008:STH

[MMBP08] Eve Mcdonald-Madden, Peter W. J. Baxter, and Hugh P. Possingham. Subpopulation triage: How to allocate conservation effort among populations. Triage subpoblacional: Cómo asignar esfuerzos de conservación entre poblaciones. Conservation Biology, 22(3):656–665, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcintyre:2005:IAL

[MMF^{+}05] Peter B. Mcintyre, Ellinor Michel, Kristin France, Adam Rivers, Paul Hakizimana, and Andrew S. Cohen. Individual- and assemblage-level effects of anthropogenic sedimentation on snails in Lake Tanganyika. Efectos a nivel individual y de ensamble de la Sedimentación antropogénica sobre caracoles en el Lago Tanganyika. Conservation Biology, 19(1):171–181, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marsh:2005:FRP

[MMGB05] David M. Marsh, Graham S. Milam, Nicholas P. Gorham, and Noelle G. Beckman. Forest roads as partial barriers to terrestrial salamander movement carreteras en bosques como barreras parciales para el movimiento de salamandras terrestres. Conservation Biology, 19(6):2004–2008, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Molnar:2004:SGT

[MMK04] Jennifer Molnar, Michelle Marvier, and Peter Kareiva. The sum is greater than the parts. Conservation Biology, 18(6): 1670–1671, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Molina:2006:PRO

[MML06] Randy Molina, Bruce G. Marcot, and Robin Lesher. Protecting rare, old-growth, forest-associated species under the survey and manage program guidelines of the northwest forest plan. Protección de especies raras y asociadas a bosques viejos bajos el programa de monitoreo y gestión del plan forestal del noroeste. Conservation Biology, 20(2):306–318, April 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McKone:2001:EEC

[MMLM01] Mark J. McKone, Kendra K. McLauchlan, Edward G. Lebrun, and Andrew C. McCall. An edge effect caused by adult cornrootworm beetles on sunflowers in tallgrass prairie remnants. Un efecto de borde causado por escarabajos adultos sobre girasoles en remanentes de pastizales de praderas. Conservation Biology, 15(5):1315–1324, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marker:2003:FIP

[MMM03] L. L. Marker, M. G. L. Mills, and D. W. Macdonald. Factors influencing perceptions of conflict and tolerance toward cheetahs on Namibian farmlands. Factores que influyen en las percepciones de conflicto y tolerancia de guepardos en granjas de Namibia. Conservation Biology, 17(5):1290–1298, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meijaard:2003:PHB

[MN03] E. Meijaard and V. Nijman. Primate hotspots on Borneo: Predictive value for general biodiversity and the effects of taxonomy. reas prioritarias para primates en Borneo: Valor predictivo para la biodiversidad general y los efectos de la taxonomía. Conservation Biology, 17(3):725–732, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:2007:RSC

[MN07] Dennis D. Murphy and Barry R. Noon. The role of scientists in conservation planning on private lands. Conservation Biology, 21(1):25–28, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milner:2007:DSE

[MNA07] Jos M. Milner, Erlend B. Nilsen, and Harry P. Andreassen. Demographic side effects of selective hunting in ungulates and carnivores. Efectos demográficos secundarios de la cacería selectiva en ungulados y carnívoros. Conservation Biology, 21(1):36–47, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maerz:2009:DWS

[MNB09] John C. Maerz, Victoria A. Nuzzo, and Bernd Blossey. Declines in woodland salamander abundance associated with non-native earthworm and plant invasions declinaciones en la abundancia de salamandras asociadas con invasiones de lombrices y plantas no nativas. Conservation Biology, 23(4):975–981, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muldavin:2001:IGB

[MNH01] Esteban H. Muldavin, Paul Neville, and Glenn Harper. Indices of grassland biodiversity in the Chihuahuan desert ecoregion derived from remote sensing. Indices de biodiversidad en pastizales derivados de percepción remota de la ecoregión del desierto de Chihuahua. Conservation Biology, 15(4):844–855, August 2001.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McMillan:2003:ERC

[MNL03] Michele A. McMillan, Jeffrey C. Nekola, and Douglas W. Larson. Effects of rock climbing on the land snail community of the Niagara Escarpment in Southern Ontario, Canada. Efectos del alpinismo en la comunidad de caracoles terrestres del Acantilado del Niagara en Ontario Meridional, Canadá. Conservation Biology, 17(2):616–621, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Markovchick-Nicholls:2008:RBH

[MNRD^{+}08] Lisa Markovchick-Nicholls, Helen M. Regan, Douglas H. Deutschman, Astrid Widyanata, Barry Martin, Lani Noreke, and Timothy Ann Hunt. Relationships between human disturbance and wildlife land use in urban habitat fragments. Relaciones entre perturbación humana y uso de suelo para vida silvestre en fragmentos de hábitat urbano. Conservation Biology, 22(1):99–109, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mol:2004:DEE

[MO04] Jan H. Mol and Paul E. Ouboter. Downstream effects of erosion from small-scale gold mining on the instream habitat and fish community of a small neotropical rainforest stream. Efectos río abajo de la erosión de Minería de oro a baja escala sobre el hábitat interior y la comunidad de peces en un pequeño arroyo en una selva lluviosa neotropical. Conservation Biology, 18(1): 201–214, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moilanen:2008:GCM

[Moi08] Atte Moilanen. Generalized complementarity and mapping of the concepts of systematic conservation planning. Conservation Biology, 22(6):1655–1658, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moomaw:2002:LSE

[Moo02] William R. Moomaw. Lomborg’s The Skeptical Environmentalist: Refuting a scientific model without science. Conservation Biology, 16(4):861–862, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moore:2007:SCS

[Moo07] Kathleen Dean Moore. In the shadow of the cedars: the spiritual values of old-growth forests. Conservation Biology, 21(4):1120– 1123, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mawdsley:2009:RCC

[MOO09] Jonathan R. Mawdsley, Robin O’Malley, and Dennis S. Ojima. A review of climate-change adaptation strategies for wildlife management and biodiversity conservation. Una revisión de las estrategias de adaptación al cambio climático para el manejo de vida silvestre y conservación de la biodiversidad. Conservation Biology, 23(5):1080–1089, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moreno:2000:IBP

[Mor00a] Claudia E. Moreno. Iberoamerican biodiversity,Part 2. Conservation Biology, 14(1):330–331, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morgan:2000:RSR

[Mor00b] John W. Morgan. Reproductive success in reestablished versus natural populations of a threatened grassland daisy (Rutidosis leptorrhynchoides). Exito reproductivo en poblaciones restablecidas contra poblaciones naturales de una margarita de pastizal (Rutidosis leptorrhynchoides) en peligro de extinción. Conservation Biology, 14(3):780–785, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mccarthy:2007:AAM

[MP07a] Michael A. Mccarthy and Hugh P. Possingham. Active adaptive management for conservation. Manejo adaptativo activo para conservar. Conservation Biology, 21(4):956–963, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Michalski:2007:DMM

[MP07b] Fernanda Michalski and Carlos A. Peres. Disturbance-mediated mammal persistence and abundance-area relationships in Amazonian forest fragments. Persistencia y relaciones abundancia-Área de mamíferos en condiciones de perturbación en fragmentos de la selva amazónica. Conservation Biology, 22(1):1626–1640, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mccarthy:2008:OMT

[MP08] Michael A. Mccarthy and Kirsten M. Parris. Optimal marking of threatened species to balance benefits of information with impacts of marking. Marcaje Óptimo de especies amenazadas para balancear beneficios de la información con impactos del marcaje. Conservation Biology, 22(6):1506–1512, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martorell:2009:DRA

[MP09] Carlos Martorell and Edward M. Peters. Disturbance-response analysis: a method for rapid assessment of the threat to species in disturbed areas. Análisis de la respuesta al disturbio: un método de evaluación rápida de la amenaza a especies en Áreas perturbadas. Conservation Biology, 23(2):377–387, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCarthy:2001:TAP

[MPDT01] M. A. McCarthy, H. P. Possingham, J. R. Day, and A. J. Tyre. Testing the accuracy of population viability analysis. Probando la precisión del análisis de viabilidad poblacional. Conservation Biology, 15(4):1030–1038, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mikusinski:2007:CPF

[MPE^{+}07] G. Mikusiński, R. L. Pressey, L. Edenius, H. Kujala, A. Moilanen, J. Niemelä, and T. Ranius. Conservation planning in forest landscapes of Fennoscandia and an approach to the challenge of Countdown 2010. Planificación de la conservación en paisajes forestales de Fennoscandia y un acercamiento al reto de la Cuenta Regresiva 2010. Conservation Biology, 21(6):1445– 1454, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mooers:2007:BLL

[MPFBH07] A.Ø. Mooers, L. R. Prugh, M. Festa-Bianchet, and J. A. Hutchings. Biases in legal listing under Canadian endangered species legislation. Conservation Biology, 21(3):572–575, June 2007.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Menon:2001:ICP

[MPR^{+}01] Shaily Menon, R. Gil Pontius, Jr., Joseph Rose, M. L. Khan, and Kamaljit S. Bawa. Identifying conservation-priority areas in the tropics: a land-use change modeling approach. Identificación de areas prioritarias de conservación en los trópicos: un modelo enfocado al cambio en el uso del suelo. Conservation Biology, 15(2):501–512, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Malcolm:2000:ITE

[MR00a] Jay R. Malcolm and Justina C. Ray. Influence of timber extraction routes on Central African small-mammal communities, forest structure, and tree diversity. Influencia de las rutas de extracción de madera sobre las comunidades de mamíferos Pequeños, la estructura del bosque y la diversidad de arboles en Africa Central. Conservation Biology, 14(6):1623–1638, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McShea:2000:MAD

[MR00b] William J. McShea and John H. Rappole. Managing the abundance and diversity of breeding bird populations through manipulation of deer populations. Manejo de la abundancia y diversidad de aves reproductoras mediante la manipulación de poblaciones de venados. Conservation Biology, 14(4):1161–1170, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2007:NTS

[MRAF^{+}07] Rebecca M. Miller, Jon Paul Rodríguez, Theresa Aniskowicz-Fowler, Channa Bambaradeniya, Ruben Boles, Mark A. Eaton, Ulf Gärdenfors, Verena Keller, Sanjay Molur, Sally Walker, and Caroline Pollock. National threatened species listing based on IUCN criteria and regional guidelines: Current status and future perspectives enlistado de especies nacionales amenazadas con base en criterios de IUCN y directrices regionales: Estatus actual y perspectivas futuras. Conservation Biology, 21(3):684–696, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcinerny:2006:SSR

[MRDC06] Greg J. Mcinerny, David L. Roberts, Anthony J. Davy, and Phillip J. Cribb. Significance of sighting rate in inferring extinction and threat. Significado de la tasa de avistamiento en la inferencia de extinción y amenaza. Conservation Biology, 20 (2):562–567, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:2000:RPD

[MRH^{+}00] B. Miller, R. Reading, J. Hoogland, T. Clark, G. Ceballos, R. List, S. Forrest, L. Hanebury, P. Manzano, J. Pacheco, and D. Uresk. The role of prairie dogs as a Keystone species: Response to Stapp. Conservation Biology, 14(1):318–321, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mathieu:2005:RSM

[MRM^{+}05] J. Mathieu, J.-P. Rossi, P. Mora, P. Lavelle, P. F. da S. Martins, C. Rouland, and M. Grimaldi. Recovery of soil macrofauna communities after forest clearance in Eastern Amazonia, Brazil. Recuperación de comunidades de macrofauna del suelo después de la tala de bosques en la Amazonía oriental, Brasil. Conservation Biology, 19(5):1598–1605, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mathiasen:2007:GSE

[MRP07] Paula Mathiasen, Adriana E. Rovere, and Andrea C. Premoli. Genetic structure and early effects of inbreeding in fragmented temperate forests of a self-incompatible tree, Embothrium coccineum. Estructura genética y efectos tempranos de la endogamia en bosques templados fragmentados de un Árbol auto incompatible, Embothrium coccineum. Conservation Biology, 21 (1):232–240, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2002:GPB

[MS02] Paul S. Martin and Christine R. Szuter. Game parks before and after Lewis and Clark: Reply to Lyman and Wolverton. Conservation Biology, 16(1):244–247, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcphee:2004:IBV

[MS04] M. Elsbeth Mcphee and Emily D. Silverman. Increased behavioral variation and the calculation of release numbers for reintroduction programs. Variación conductual incrementada y el Cálculo del número de liberaciones para programas de reintroducción. Conservation Biology, 18(3):705–715, June 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milbau:2008:FAA

[MS08] Ann Milbau and Jane C. Stout. Factors associated with alien plants transitioning from casual, to naturalized, to invasive. Factores asociados con la transición de plantas exóticas de casuales, a naturalizadas, a invasoras. Conservation Biology, 22(2):308– 317, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muir:2009:ART

[MS09] Matthew J. Muir and Mark W. Schwartz. Academic research training for a nonacademic workplace: a case study of graduate student alumni who work in conservation. Formación en investigación académica para un lugar de trabajo no académico: un estudio de caso graduados que trabajan en conservación. Conservation Biology, 23(6):1357–1368, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meretsky:2000:DCC

[MSB^{+}00] Vicky J. Meretsky, Noel F. R. Snyder, Steven R. Beissinger, David A. Clendenen, and James W. Wiley. Demography of the California Condor: Implications for reestablishment. Demografía del cóndor de California: Implicaciones para su restablecimiento. Conservation Biology, 14(4):957–967, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meretsky:2001:QVQ

[MSB^{+}01] Vicky Meretsky, Noel F. R. Snyder, Steven R. Beissinger, David A. Clendenen, and James W. Wiley. Quantity versus quality in California Condor reintroduction: Reply to Beres and Starfield. Conservation Biology, 15(5):1449–1451, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martinez-Solano:2003:DTC

[MSBGP03] Iñigo Martínez-Solano, Jaime Bosch, and Mario García-París. Demographic trends and community stability in a montane amphibian assemblage. Tendencias demográficas y estabilidad en una comunidad de anfibios de montaña. Conservation Biology, 17(1):238–244, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murray-Smith:2009:PDH

[MSBOF^{+}09] Charlotte Murray-Smith, Neil A. Brummitt, Ary T. Oliveira-Filho, Steven Bachman, Justin Moat, Eimear M. Nic Lughadha, and Eve J. Lucas. Plant diversity hotspots in the Atlantic coastal forests of Brazil. Sitios de importancia para la conservación de la diversidad de plantas en los bosques de la costa del Atlántico de Brasil. Conservation Biology, 23(1):151–163, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martinich:2006:PSC

[MSL06] Jeremy A. Martinich, Susan L. Solarz, and James R. Lyons. Preparing students for conservation careers through projectbased learning. Conservation Biology, 20(6):1579–1583, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meine:2006:MDD

[MSN06] Curt Meine, Michael Soulé, and Reed F. Noss. “A mission-driven discipline”: the growth of conservation biology. “Una disciplina dirigida por una misión”: el crecimiento de conservation biology. Conservation Biology, 20(3):631–651, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mercado-Silva:2006:FSI

[MSOM^{+}06] Norman Mercado-Silva, Julian D. Olden, Jeffrey T. Maxted, Thomas R. Hrabik, and M. Jake Vander Zanden. Forecasting the spread of invasive rainbow smelt in the Laurentian Great Lakes region of North America. Pronóstico de la dispersión de la especie invasora Osmerus mordax en la Región de los Grandes Lagos de Norteamérica. Conservation Biology, 20(6):1740–1749, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murray:2009:IDD

[MSSM09] Kris A. Murray, Lee F. Skerratt, Rick Speare, and Hamish Mccallum. Impact and dynamics of disease in species threatened by the amphibian chytrid fungus, Batrachochytrium dendrobatidis. Impacto y dinámica de la enfermedad en especies amenazadas por el hongo quitridio anfibio Batrachochytrium dendrobatidis. Conservation Biology, 23(5):1242–1252, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comment [Sch10].

Mumme:2000:LDF

[MSWF00] Ronald L. Mumme, Stephan J. Schoech, Glen E. Woolfenden, and John W. Fitzpatrick. Life and death in the fast lane: Demographic consequences of road mortality in the Florida scrubjay. Vida y muerte en la carretera: Consecuencias demográficas de mortalidad asociada a la carretera en la urraca azuleja de Florida. Conservation Biology, 14(2):501–512, April 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marsh:2001:MDA

[MT01] David M. Marsh and Peter C. Trenham. Metapopulation dynamics and amphibian conservation. Dinámicas metapoblacionales y conservación de anfibios. Conservation Biology, 15 (1):40–49, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marsh:2008:CTP

[MT08] David M. Marsh and Peter C. Trenham. Current trends in plant and animal population monitoring tendencias actuales en el monitoreo de poblaciones de plantas y animales. Conservation Biology, 22(3):647–655, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2000:BRI

[MTB00] Jean-Louis Martin, Jean-Claude Thibault, and Vincent Bretagnolle. Black rats, island characteristics, and colonial nesting birds in the Mediterranean: Consequences of an ancient introduction. Ratas negras, características de islas, y nidación colonial de aves en el Mediterráneo: Consecuencias de una introducción ancestral. Conservation Biology, 14(5):1452–1466, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minor:2008:GTF

[MU08] Emily S. Minor and Dean L. Urban. A graph-theory framework for evaluating landscape connectivity and conservation planning. Un marco de referencia Teórico–Gráfico para evaluar la conectividad del paisaje y planificar conservación. Conservation Biology, 22(2):297–307, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mulder:2007:ICP

[Mul07] Monique Borgerhoff Mulder. Interdisciplinary collaboration: Painting a brighter picture and identifying the real problem. Conservation Biology, 21(4):903–904, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mullen:2009:PPW

[Mul09] Renée B. Mullen. Putting passion to work. Conservation Biology, 23(3):781–782, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:2000:IVC

[Mur00] Jack Murphy. Improving the value of conservation programs. Conservation Biology, 14(6):1569, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:2001:SSD

[Mur01] Michael T. Murphy. Source-sink dynamics of a declining eastern kingbird population and the value of sink habitats. Dinámica fuente-vertedero de una población en declinación de tirano dorsinegro y el valor de hábitats vertedero. Conservation Biology, 15(3):737–748, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matson:2006:AIW

[MV06] Pamela A. Matson and Peter M. Vitousek. Agricultural intensification: Will land spared from farming be land spared for nature? Conservation Biology, 20(3):709–710, June 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moilanen:2007:BQP

[MW07] Atte Moilanen and Brendan A. Wintle. The boundary-quality penalty: a quantitative method for approximating species responses to fragmentation in reserve selection. La sanción calidad del lindero: un método cuantitativo para aproximarse a las respuestas de especies a la fragmentación en la selección de reservas. Conservation Biology, 21(2):355–364, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcbride:2007:IES

[MWBP07] Marissa F. Mcbride, Kerrie A. Wilson, Michael Bode, and Hugh P. Possingham. Incorporating the effects of socioeconomic uncertainty into priority setting for conservation investment. Incorporación de los efectos de la incertidumbre Socioeconómica en la definición de prioridades para la Inversión en conservación. Conservation Biology, 21(6):1463–1474, December 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moilanen:2006:UAR

[MWEB06] Atte Moilanen, Brendan A. Wintle, Jane Elith, and Mark Burgman. Uncertainty analysis for regional-scale reserve selection. Análisis de incertidumbre para la selección de reservas a escala regional. Conservation Biology, 20(6):1688–1697, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morita:2002:EHF

[MY02] Kentaro Morita and Shoichiro Yamamoto. Effects of habitat fragmentation by damming on the persistence of stream-dwelling charr populations. Efectos de la fragmentación del hábitat debido a la construcción de presas sobre la persistencia de poblaciones de la trucha de montaña que habita arroyos. Conservation Biology, 16(5):1318–1323, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcdonald:2007:EEP

[MYFF^{+}07] Robert I. Mcdonald, Chris Yuan-Farrell, Charles Fievet, Matthias Moeller, Peter Kareiva, David Foster, Ted Gragson, Ann Kinzig, Lauren Kuby, and Charles Redman. Estimating the effect of protected lands on the development and conservation of their surroundings. Estimación del efecto de Áreas protegidas sobre el desarrollo y la conservación de sus alrededores. Conservation Biology, 21(6):1526–1536, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Monkkonen:2009:EEV

[MYH09] Mikko Mönkkönen, Anna-Liisa Ylisirniö, and Tanja Hämäläinen. Ecological efficiency of voluntary conservation of boreal-forest biodiversity. Eficiencia ecológica de la conservación voluntaria de la biodiversidad de bosques boreales. Conservation Biology, 23(2):339–347, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naidoo:2001:EEP

[NA01] Robin Naidoo and Wiktor L. Adamowicz. Effects of economic prosperity on numbers of threatened species. Efectos de la prosperidad económica sobre los números de especies amenazadas. Conservation Biology, 15(4):1021–1029, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naidoo:2006:MOC

[NA06] Robin Naidoo and Wiktor L. Adamowicz. Modeling opportunity costs of conservation in transitional landscapes. Modelado de costos de oportunidad de la conservación en paisajes de transición. Conservation Biology, 20(2):490–500, April 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nascimento:2006:ESM

[NAC^{+}06] Henrique E. M. Nascimento, Ana C. S. Andrade, José Luís C. Camargo, William F. Laurance, Susan G. Laurance, and José Eduardo L. Ribeiro. Effects of the surrounding matrix on tree recruitment in Amazonian forest fragments. Efectos de la matriz circundante sobre el reclutameinto de Árboles en fragmentos de bosque amazónicos. Conservation Biology, 21(3):853–860, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nadkarni:2004:PCC

[Nad04] Nalini M. Nadkarni. Not preaching to the choir: Communicating the importance of forest conservation to nontraditional audiences. Conservation Biology, 18(3):602–606, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nepstad:2004:MAT

[NARL^{+}04] Daniel Nepstad, Claudia Azevedo-Ramos, Eirivelthon Lima, David McGrath, Cássio Pereira, and Frank Merry. Managing the Amazon timber industry. Conservation Biology, 18(2):575– 577, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nassauer:2006:LPC

[Nas06] Joan Iverson Nassauer. Landscape planning and conservation biology: Systems thinking revisited. Conservation Biology, 20 (3):677–678, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2000:ALT

[NB00] Reed F. Noss and Paul Beier. Arguing over little things: Response to Haddad et al. Conservation Biology, 14(5):1546–1548, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noon:2006:CNS

[NB06] Barry R. Noon and Jennifer A. Blakesley. Conservation of the northern spotted owl under the northwest forest plan. Conservación de Strix occidentalis caurina bajo el plan forestal del noroeste. Conservation Biology, 20(2):288–296, April 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Novotny:2004:LSR

[NBM^{+}04] V. Novotny, Y. Basset, S. E. Miller, R. L. Kitching, M. Laidlaw, P. Drozd, and L. Cizek. Local species richness of leaf-chewing insects feeding on woody plants from one hectare of a lowland rainforest. Riqueza de especies locales de insectos masticadores de hojas en plantas leñosas de una Hectárea de selva lluviosa. Conservation Biology, 18(1):227–237, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Neel:2003:ECT

[NC03a] Maile C. Neel and Michael P. Cummings. Effectiveness of conservation targets in capturing genetic diversity. Efectividad de los objetivos de conservación en la captura de diversidad genética. Conservation Biology, 17(1):219–229, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norse:2003:WWW

[NC03b] Elliott A. Norse and James T. Carlton. World wide Web buzz about biodiversity. Conservation Biology, 17(6):1475–1476, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2002:MAI

[NCVBW02] Reed F. Noss, Carlos Carroll, Ken Vance-Borland, and George Wuerthner. A multicriteria assessment of the irreplaceability and vulnerability of sites in the Greater Yellowstone Ecosystem. Evaluación de criterios múltiples de la irremplazabilidad y vulnerabilidad de sitios en el Ecosistema Mayor de Yellowstone. Conservation Biology, 16(4):895–908, August 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norton:2003:FVT

[ND03] David A. Norton and Peter J. De Lange. Fire and vegetation in a temperate peat bog: Implications for the management of threatened species. Incendios y vegetación en una turba pantanosa templada: Implicaciones para el manejo de especies. Conservation Biology, 17(1):138–148, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Neke:2004:TTQ

[ND04] Kirsten S. Neke and Morné A. Du Plessis. The threat of transformation: Quantifying the vulnerability of grasslands in South Africa. La amenaza de la transformación: Cuantificando la vulnerabilidad de pastizales en Sudáfrica. Conservation Biology, 18(2):466–477, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nelson:2006:RS

[Nel06] Fred Nelson. A radical science. Conservation Biology, 33(5): 1333–1334, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nelson:2009:CAD

[Nel09] Fred Nelson. Conservation and aid: Designing more effective investments in natural resource governance reform. Conservación y asistencia: Diseño de inversiones más efectivas en las reformas a la gestión de los recursos naturales. Conservation Biology, 23 (5):1102–1108, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newmark:2006:YSF

[New06] William D. Newmark. A 16-year study of forest disturbance and understory bird community structure and composition in Tanzania. Un estudio de 16 años de la perturbación de bosques y la estructura y composición de la comunidad de aves de sotobosque en Tanzania. Conservation Biology, 20(1):122–134, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newmark:2008:AMC

[New08] William Newmark. Another milestone in conservation literature. Conservation Biology, 22(1):224, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nally:2004:SPM

[NF04] Ralph Mac Nally and Erica Fleishman. A successful predictive model of species richness based on indicator species exitoso modelo predictivo de la riqueza de especies basado en especies indicadoras. Conservation Biology, 18(3):646–654, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newton:2005:AAN

[NF05a] Julianne Lutz Newton and Eric T. Freyfogle. All about nature. Conservation Biology, 19(1):42–44, February 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newton:2005:SDS

[NF05b] Julianne Lutz Newton and Eric T. Freyfogle. Sustainability: a dissent sustentabilidad: un disentimiento. Conservation Biology, 19(1):23–32, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nol:2005:UDP

[NFB05] Erica Nol, Charles M. Francis, and Dawn M. Burke. Using distance from putative source woodlots to predict occurrence of forest birds in putative sinks. Utilización de la distancia de parcelas putativas de Árboles fuente para predecir la ocurrencia de aves de bosque en vertederos putativos. Conservation Biology, 19(3): 836–844, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2009:PIS

[NFD^{+}09] Reed F. Noss, Erica Fleishman, Dominick A. Dellasala, John M. Fitzgerald, Mart R. Gross, Martin B. Main, Fiona Nagle, Stacey L. O’Malley, and Jon Rosales. Priorities for improving the scientific foundation of conservation policy in North America. Prioridades para el mejoramiento de la ciencia de la conservación. Conservation Biology, 23(4):825–833, August 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nally:2004:ITS

[NFM04] Ralph Mac Nally, Erica Fleishman, and Dennis D. Murphy. Influence of temporal scale of sampling on detection of relationships between invasive plants and the diversity patterns of plants and butterflies. Influencia de la escala temporal de muestreo sobre la detección de relaciones entre especies invasoras y patrones de diversidad de plantas y mariposas. Conservation Biology, 18 (6):1525–1532, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nadkarni:2009:CSB

[NH09] Nalini M. Nadkarni and Willam A. Haber. Canopy seed banks as time capsules of biodiversity in pasture-remnant tree crowns. Bancos de semillas en el dosel como cápsulas de tiempo de biodiversidad en las copas de Árboles remanentes en potreros. Conservation Biology, 23(5):1117–1126, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nagler:2005:RNT

[NHHG^{+}05] Pamela L. Nagler, Osvel Hinojosa-Huerta, Edward P. Glenn, Jaqueline Garcia-Hernandez, Reggie Romo, Charles Curtis, Alfredo R. Huete, and Stephen G. Nelson. Regeneration of native trees in the presence of invasive saltcedar in the Colorado River Delta, Mexico. Senderos recreativos como corredores que facilitan invasiones de plantas no nativas en el Boundary Waters Canoe Area Wilderness (E.U.A.). Conservation Biology, 19(6): 1842–1852, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nelson:2002:ENW

[NKBD02] Nicola J. Nelson, Susan N. Keall, Derek Brown, and Charles H. Daugherty. Establishing a new wild population of tuatara (Sphenodon guntheri). Establecimiento de una nueva población silvestre de tuatara (Sphenodon guntheri). Conservation Biology, 16(4):887–894, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nicholson:2009:ATS

[NKW09] Emily Nicholson, David A. Keith, and David S. Wilcove. Assessing the threat status of ecological communities. Evaluación del estatus de amenaza de comunidades ecológicas. Conservation Biology, 23(2):259–274, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niesenbaum:2003:GCB

[NL03a] Richard A. Niesenbaum and Tammy Lewis. Ghettoization in conservation biology: How interdisciplinary is our teaching? Conservation Biology, 17(1):6–10, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niesenbaum:2003:ROA

[NL03b] Richard A. Niesenbaum and Tammy Lewis. In response to our article. Conservation Biology, 17(5):1203, October 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niel:2005:UDI

[NL05] Colin Niel and Jean-Dominique Lebreton. Using demographic invariants to detect overharvested bird populations from incomplete data. Uso de invariantes demográficas para la detección de la Sobreexplotación de poblaciones de aves utilizando datos incompletos. Conservation Biology, 19(3):826–835, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2006:SSE

[NL06] Reed F. Noss and David B. Lindenmayer. Special Section: The Ecological Effects of Salvage Logging after Natural Disturbance. Conservation Biology, 20(4):946–948, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naujokaitis-Lewis:2009:SAS

[NLCAR09] Ilona R. Naujokaitis-Lewis, Janelle M. R. Curtis, Peter Arcese, and Jordan Rosenfeld. Sensitivity analyses of spatial population viability analysis models for species at risk and habitat conservation planning. Análisis de sensibilidad de modelos de análisis de viabilidad poblacional espaciales para especies en riesgo y la planificación de la conservación de hábitat. Conservation Biology, 23(1):225–229, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nuzzo:2009:EID

[NMB09] Victoria A. Nuzzo, John C. Maerz, and Bernd Blossey. Earthworm invasion as the driving force behind plant invasion and community change in Northeastern North American forests. Invasión de lombrices de tierra como la fuerza conductora detrás de la invasión de plantas y cambios en la comunidad en bosques del noreste de América del Norte. Conservation Biology, 23 (4):966–974, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nogales:2004:RFC

[NMT^{+}04] Manuel Nogales, Aurelio Martín, Bernie R. Tershy, C. Josh Donlan, Dick Veitch, Néstor Puerta, Bill Wood, and Jesús Alonso. A review of feral cat eradication on islands. Revisión de la erradicación de gatos asilvestrados en islas. Conservation Biology, 18(2):310–319, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ney-Nifle:2000:HLC

[NNM00] Muriel Ney-Nifle and Marc Mangel. Habitat loss and changes in the species–area relationship. Pérdida del hábitat y cambios en la relación especie–area. Conservation Biology, 14(3):893– 898, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norton:2000:CBP

[Nor00] David A. Norton. Conservation biology and private land: Shifting the focus. Conservation Biology, 14(5):1221–1223, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norgaard:2003:IGP

[Nor03] Richard B. Norgaard. Individual greed and public goods. Conservation Biology, 17(4):1190–1192, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norgaard:2008:FHM

[Nor08] Richard B. Norgaard. Finding hope in the Millennium ecosystem assessment. Encontrando esperanza en la evaluación de ecosistemas milenio. Conservation Biology, 22(4):862–869, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2000:SB

[Nos00] Reed F. Noss. Science on the bridge. Conservation Biology, 14 (2):333–335, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2001:BKF

[Nos01a] Reed F. Noss. Beyond Kyoto: Forest management in a time of rapid climate change. Después de Kyoto: Manejo forestal en tiempos de cambio climático acelerado. Conservation Biology, 15(3):578–590, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2001:TPL

[Nos01b] Reed F. Noss. Toward a Pro-life politics. Conservation Biology, 15(4):827–828, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2003:CWN

[Nos03] Reed F. Noss. A checklist for wildlands network designs lista de control para el diseño de redes de tierras silvestres. Conservation Biology, 17(5):1270–1275, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2007:VGT

[Nos07] Reed F. Noss. Values are a good thing in conservation biology. Conservation Biology, 21(1):18–20, February 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2009:HBC

[Nos09] Reed F. Noss. The heavy burden of conservation. Conservation Biology, 23(6):1354–1355, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nicholson:2006:OMS

[NP06] Emily Nicholson and Hugh P. Possingham. Objectives for multiple-species conservation planning. Objetivos para la planificación de la conservación de múltiples especies. Conservation Biology, 21(3):871–881, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noon:2005:CSB

[NPT05] Barry R. Noon, Patrick Parenteau, and Stephen C. Trombulak. Conservation science, biodiversity, and the 2005 U.S. forest service regulations. Conservation Biology, 19(5):1359–1361, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Novinger:2003:IMA

[NR03] Douglas C. Novinger and Frank J. Rahel. Isolation management with artificial barriers as a conservation strategy for cutthroat trout in headwater streams. El aislamiento con barreras artificiales como una estrategia de conservación para la trucha (Oncorhynchus clarki pleuriticus) en arroyos de cabecera. Conservation Biology, 17(3):772–781, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Novaro:2000:EHS

[NRB00] Andres J. Novaro, Kent H. Redford, and Richard E. Bodmer. Effect of hunting in source-sink systems in the neotropics. Efectas de la cacería en sistemas fuente-sumidero del Neotrópico. Conservation Biology, 14(3):713–721, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newburn:2005:ELU

[NRBM05] David Newburn, Sarah Reed, Peter Berck, and Adina Merenlender. Economics and land-use change in prioritizing private land conservation. Economía y cambio en el uso de suelo en la priorización de la conservación de tierras privadas. Conservation Biology, 19(5):1411–1420, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norris:2001:EMF

[NS01] D. Ryan Norris and Bridget J. M. Stutchbury. Extraterritorial movements of a forest songbird in a fragmented landscape movimientos extraterritoriales de una ave canora de bosque en un paisaje fragmentado. Conservation Biology, 15(3):729–736, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nakamura:2009:UAI

[NS09] Miguel Nakamura and Jorge Soberón. Use of approximate inference in an index of completeness of biological inventories. Utilización de la inferencia aproximada como un Índice de completitud de inventarios biológicos. Conservation Biology, 23(2): 469–474, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nepstad:2006:GAS

[NSA06] Daniel C. Nepstad, Claudia M. Stickler, and Oriana T. Almeida. Globalization of the Amazon soy and beef industries: Opportunities for conservation. Globalización de las industrias de soya y ganado del Amazonas: Oportunidades para la conservación. Conservation Biology, 20(6):1595–1603, December 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nepstad:2006:IAD

[NSB^{+}06] D. Nepstad, S. Schwartzman, B. Bamberger, M. Santilli, D. Ray, P. Schlesinger, P. Lefebvre, A. Alencar, E. Prinz, Greg Fiske, and Alicia Rolla. Inhibition of Amazon deforestation and fire by parks and indigenous lands. Inhibición de deforestación e incendios por parques y terrenos indígenas en la Amazonia. Conservation Biology, 20(1):65–73, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Namgyel:2008:SCB

[NSW08] Ugyen Namgyel, Stephen F. Siebert, and Sonam Wang. Shifting cultivation and biodiversity conservation in Bhutan. Conservation Biology, 22(5):1349–1351, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newman:2001:EEB

[NT01] Dara Newman and David A. Tallmon. Experimental evidence for beneficial fitness effects of gene flow in recently isolated populations. Evidencia experimental de los efectos benéficos del flujo de genes en la adaptabilidad de poblaciones recientemente aisladas. Conservation Biology, 15(4):1054–1063, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naughton-Treves:2003:PTR

[NTGT03] Lisa Naughton-Treves, Rebecca Grossberg, and Adrian Treves. Paying for tolerance: Rural citizens’ attitudes toward wolf depredation and compensation. El impacto de la depredación y de los pagos compensatorios en las actitudes de ciudadanos hacia lobos. Conservation Biology, 17(6):1500–1511, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naughton-Treves:2003:WSB

[NTMT^{+}03] Lisa Naughton-Treves, Jose Luis Mena, Adrian Treves, Nora Alvarez, and Volker Christian Radeloff. Wildlife survival beyond park boundaries: the impact of slash-and-burn agriculture and hunting on mammals in Tambopata, Peru. Sobrevivencia de fauna silvestre por fuera de los límites del parque: el impacto de la agricultura de roza y quema y de la cacería sobre mamíferos en Tambopata, Perú. Conservation Biology, 17(4): 1106–1117, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nugent:2008:ALU

[Nug08] Conn Nugent. American land use: the next 50 years. Conservation Biology, 22(6):1375, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nelson:2009:AES

[NV09] Michael P. Nelson and John A. Vucetich. On advocacy by environmental scientists: What, whether, why, and how. Sobre la abogacía por científicos ambientales: Qué, Sí, porque y Cómo. Conservation Biology, 23(5):1090–1101, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nelson:2007:NME

[NVP07] Michael P. Nelson, John A. Vucetich, and Michael K. Phillips. Normativity and the meaning of Endangered: Response to Waples et al. 2007. Conservation Biology, 22(1):1646–1648, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [VNP06].

Naves:2003:ESC

[NWRD03] Javier Naves, Thorsten Wiegand, Eloy Revilla, and Miguel Delibes. Endangered species constrained by natural and human factors: the case of brown bears in Northern Spain especies amenazadas limitadas por factores naturales y humanos: el caso del oso pardo en el norte de España. Conservation Biology, 17(5): 1276–1289, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oates:2000:EWA

[OALM^{+}00] John F. Oates, Michael Abedi-Lartey, W. Scott McGraw, Thomas T. Struhsaker, and George H. Whitesides. Extinction of a West African red colobus monkey. Extinción del mono colobus rojo en Africa occidental. Conservation Biology, 14(5):1526– 1532, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Owens:2000:QBP

[OB00] Ian P. F. Owens and Peter M. Bennett. Quantifying biodiversity: a phenotypic perspective. Cuantificación de la biodiversidad: una perspectiva fenotípica. Conservation Biology, 14(4): 1014–1022, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OGrady:2004:CAE

[OBK^{+}04] Julian J. O’Grady, Mark A. Burgman, David A. Keith, Lawrence L. Master, Sandy J. Andelman, Barry W. Brook, Geoffrey A. Hammerson, Tracey Regan, and Richard Frankham. Correlations among extinction risks assessed by different systems of threatened species categorization. Correlaciones entre riesgos de extinción evaluados por diferentes sistemas de categorización de especies amenazadas. Conservation Biology, 18(6): 1624–1635, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olupot:2009:ERV

[OBM09] William Olupot, Robert Barigyira, and Alastair J. Mcneilage. Edge-related variation in medicinal and other “Useful” wild plants of Bwindi Impenetrable National Park, Uganda. Variación relacionada con el borde en plantas medicinales y otras plantas silvestres “Útiles” en el Parque Nacional Bwindi Impenetrable, Uganda. Conservation Biology, 23(5):1138–1145, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OConnor:2000:ESF

[O’C00] Raymond J. O’Connor. Expert systems, fuzzy logic, and coral reef development under environmental stress. Conservation Biology, 14(3):904–906, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OConnell:2005:ACB

[O’C05] Michael O’Connell. Additions to the conservation book shelf. Conservation Biology, 19(4):1334–1335, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olden:2005:HDB

[ODD05] Julian D. Olden, Michael E. Douglas, and Marlis R. Douglas. The human dimensions of biotic homogenization. Conservation Biology, 19(6):2036–2038, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ostermann:2001:CBR

[ODE01] Stacey D. Ostermann, James R. Deforge, and W. Daniel Edge. Captive breeding and reintroduction evaluation criteria: a case study of peninsular bighorn sheep. Criterios de evaluación de la reproducción en cautiverio y la reintroducción: un caso de estudio del carnero peninsular de montaña. Conservation Biology, 15(3):749–760, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Odenbaugh:2009:DB

[Ode09] Jay Odenbaugh. The diversity of biodiversity. Conservation Biology, 23(6):1647–1648, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Odegaard:2000:MLH

[ØDEA00] Frode Ødegaard, Ola H. Diserud, Steinar Engen, and Kaare Aagaard. The magnitude of local host specificity for phytophagous insects and its implications for estimates of global species richness. Magnitud de la especificidad de la relación insecto fitófago– planta hospedero y sus implicaciones para la estimación de la riqueza específica global. Conservation Biology, 14(4):1182– 1186, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olson:2002:CBB

[ODPW02] David M. Olson, Eric Dinerstein, George V. N. Powell, and Eric D. Wikramanayake. Conservation biology for the biodiversity crisis. Conservation Biology, 16(1):1–3, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olson:2006:RFR

[OFN^{+}06] David Olson, Linda Farley, Waisea Naisilisili, Alipate Raikabula, Om Prasad, James Atherton, and Craig Morley. Remote forest refugia for Fijian wildlife. Refugios forestales remotos para vida silvestre de Fiji. Conservation Biology, 20(2):568–572, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Odion:2004:PFS

[OFS^{+}04] Dennis C. Odion, Evan J. Frost, James R. Strittholt, Hong Jiang, Dominick A. Dellasala, and Max A. Moritz. Patterns of fire severity and forest conditions in the Western Klamath Mountains, California. Patrones de severidad de fuego y condiciones del bosque en las Montañas Klamath Occidentales, California. Conservation Biology, 18(4):927–936, August 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OConnell:2004:CNH

[OGH04] Allan F. O’Connell, Jr., Andrew T. Gilbert, and Jeff S. Hatfield. Contribution of natural history collection data to biodiversity assessment in national parks. Contribución de datos de colecciones de historia natural a la evaluación de la biodiversidad en parques nacionales. Conservation Biology, 18(5):1254–1261, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Okes:2008:HUL

[OHC08] Nicola C. Okes, Philip A. R. Hockey, and Graeme S. Cumming. Habitat use and life history as predictors of bird responses to habitat change. Uso de hábitat e historia de vida como vaticinadores de las respuestas de aves al cambio de hábitat. Conservation Biology, 22(1):151–162, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ozaki:2006:MAE

[OIK^{+}06] Kenichi Ozaki, Masahiro Isono, Takayuki Kawahara, Shigeo Iida, Takuma Kudo, and Kenji Fukuyama. A mechanistic approach to evaluation of umbrella species as conservation surrogates. Un método mecanicista para la evaluación de especies sombrilla como substitutos de conservación. Conservation Biology, 20(5):1507–1515, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ostfeld:2000:BDR

[OK00] Richard S. Ostfeld and Felicia Keesing. Biodiversity and disease risk: the case of Lyme disease. Biodiversidad y riesgo de enfermedades: El caso de la enfermedad de Lyme. Conservation Biology, 14(3):722–728, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Odell:2001:SMS

[OK01a] Eric A. Odell and Richard L. Knight. Songbird and mediumsized mammal communities associated with exurban development in Pitkin County, Colorado. Comunidades de aves canoras y mamíferos Pequeños asociadas con desarrollo exurbano en el Condado Pitkin, Colorado. Conservation Biology, 15(4):1143– 1150, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OKeefe:2001:APV

[O’K01b] John O’Keefe. An alternative point of view. Conservation Biology, 15(5):1469–1470, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olden:2003:SSA

[Old03] Julian D. Olden. A species-specific approach to modeling biological communities and its potential for conservation. Un enfoque especie–específico al modelaje de comunidades biológicas y su potencial para la conservación. Conservation Biology, 17 (3):854–863, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olsson:2000:AEE

[Ols00] Gunilla A. Olsson. Assessing Europe’s environment. Conservation Biology, 14(2):590–591, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olsson:2007:GOS

[Ols07] Ola Olsson. Genetic origin and success of reintroduced white storks. Origen genético y Éxito de cigüeñas blancas reintroducidas. Conservation Biology, 21(5):1196–1206, October 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olupot:2000:MDA

[Olu00] William Olupot. Mass differences among male mangabey monkeys inhabiting logged and unlogged forest compartments diferencias en masa corporal entre machos de monos mangabey residentes de compartimentos de bosques talados y no talados. Conservation Biology, 14(3):833–843, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ouellet:2005:HEW

[OMP^{+}05] Martin Ouellet, Igor Mikaelian, Bruce D. Pauli, Jean Rodrigue, and David M. Green. Historical evidence of widespread chytrid infection in North American amphibian populations. Evidencia histórica de quitridiomicosis generalizada en poblaciones de anfibios de Norte América. Conservation Biology, 19(5):1431– 1440, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

ONeill:2007:TCM

[O’N07] Kate O’Neill. Territories of chance meet territories of choice. Conservation Biology, 22(1):1658–1659, December 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2000:P

[Orr00a] David Orr. 2020: a proposal. Conservation Biology, 14(2):338– 341, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2000:IPT

[Orr00b] David Orr. Ideasclerosis: Part two. Conservation Biology, 14(6): 1571–1573, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2000:IPO

[Orr00c] David W. Orr. Ideasclerosis: Part one. Conservation Biology, 14 (4):926–928, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2001:L

[Orr01a] David W. Orr. Leverage. Conservation Biology, 15(6):1480– 1482, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2001:LR

[Orr01b] David W. Orr. A literature of redemption. Conservation Biology, 15(2):305–307, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2001:RTC

[Orr01c] David W. Orr. Rewriting the ten commandments of American politics. Conservation Biology, 15(4):821–824, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2002:EVM

[Orr02a] David W. Orr. The events of 9-11: a view from the margin. Conservation Biology, 16(2):288–290, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2002:LS

[Orr02b] David W. Orr. The labors of Sisyphus. Conservation Biology, 16 (4):857–860, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2003:CN

[Orr03a] David Orr. The constitution of nature. Conservation Biology, 17 (6):1478–1484, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2003:D

[Orr03b] David W. Orr. Diversity. Conservation Biology, 17(4):948–951, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2003:WNS

[Orr03c] David W. Orr. Walking north on a southbound train. Conservation Biology, 17(2):348–351, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2004:CRS

[Orr04a] David W. Orr. The corruption (and redemption) of science. Conservation Biology, 18(4):862–865, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2004:HHT

[Orr04b] David W. Orr. Hope in hard times. Conservation Biology, 18 (2):295–298, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2004:OL

[Orr04c] David W. Orr. Orr’s Laws. Conservation Biology, 18(6):1457– 1460, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2004:ORH

[Orr04d] David W. Orr. Other responses to “Hope”. Conservation Biology, 18(2):301, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2005:AVE

[Orr05a] David W. Orr. Armageddon versus extinction. Conservation Biology, 19(2):290–292, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2005:FME

[Orr05b] David W. Orr. Free marketeers on the edge: a response to McCoy and Atwood. Conservation Biology, 19(4):1321–1322, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2005:R

[Orr05c] David W. Orr. A response. Conservation Biology, 19(6):1697– 1698, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2006:FS

[Orr06a] David W. Orr. Framing sustainability. Conservation Biology, 20(2):265–268, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2006:RL

[Orr06b] David W. Orr. The right to life. Conservation Biology, 20 (4):935–939, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2006:T

[Orr06c] David W. Orr. The trial. Conservation Biology, 20(6):1570– 1573, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2007:CC

[Orr07a] David W. Orr. The carbon connection. Conservation Biology, 21(2):289–292, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2007:OHD

[Orr07b] David W. Orr. One hundred days of climate action. Conservation Biology, 21(4):907–911, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2007:OHH

[Orr07c] David W. Orr. Optimism and hope in a hotter time. Conservation Biology, 21(6):1392–1395, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2008:LUC

[Orr08a] David W. Orr. Land use and climate change. Conservation Biology, 22(6):1372–1374, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2008:PS

[Orr08b] David W. Orr. The psychology of survival. Conservation Biology, 22(4):819–822, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2009:BCC

[Orr09a] David W. Orr. Baggage: the case for climate mitigation. Conservation Biology, 23(4):790–793, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2009:RPU

[Orr09b] David W. Orr. Retrospect and prospect: the unbearable lightness of conservation. Conservation Biology, 23(6):1349–1351, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:2009:SL

[Orr09c] David W. Orr. Shelf life. Conservation Biology, 23(2):248–251, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orians:2001:WCB

[OS01] Gordon H. Orians and Michael E. Soulé. Whither conservation biology research? Conservation Biology, 15(4):1187–1188, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oppel:2004:RSE

[OS04] Steffen Oppel and Martin Stock. Reconsidering species extinctions in national parks: Reply to Berger. Conservation Biology, 18(3):845–846, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Owen-Smith:2007:PHE

[OS07] Norman Owen-Smith. Protagonists of healthy ecosystems. Conservation Biology, 21(2):888, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Osofsky:2000:CMV

[Oso00] Steven A. Osofsky. Conservation medicine: a veterinary perspective. Conservation Biology, 14(2):336–337, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ohl-Schacherer:2007:SSH

[OSSK^{+}07] Julia Ohl-Schacherer, Glenn H. Shepard, Jr., Hillard Kaplan, Carlos A. Peres, Taal Levi, and Douglas W. Yu. The sustainability of subsistence hunting by Matsigenka native communities in Manu National Park, Peru. La sustentabilidad de la cacería de subsistencia de comunidades nativas Matsigenka en el Parque Nacional Manú, Perú. Conservation Biology, 21(5):1174–1185, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ovadia:2003:RHV

[Ova03] Ofer Ovadia. Ranking hotspots of varying sizes: a lesson from the nonlinearity of the species-area relationship. Conservation Biology, 17(5):1440–1441, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Owen:2008:WWE

[Owe08] Dave Owen. Water war: Explaining the Klamath Controversy. Conservation Biology, 22(6):1664–1666, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ogada:2003:LDA

[OWOF03] Mordecai O. Ogada, Rosie Woodroffe, Nicholas O. Oguge, and Laurence G. Frank. Limiting depredation by African carnivores: the role of livestock husbandry. Limitando la depredación de carnívoros africanos: el papel de la crianza de ganado. Conservation Biology, 17(6):1521–1530, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pauchard:2004:IEL

[PA04] Aníbal Pauchard and Paul B. Alaback. Influence of elevation, land use, and landscape context on patterns of alien plant invasions along roadsides in protected areas of South-Central Chile. Influencia de la altitud, uso de suelo y contexto del paisaje sobre invasiones de plantas exóticas. Conservation Biology, 18(1): 238–248, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parr:2006:PMB

[PA06a] Catherine L. Parr and Alan N. Andersen. Patch mosaic burning for biodiversity conservation: a critique of the pyrodiversity paradigm. Quema de mosaico de parches para la conservación de biodiversidad: una crítica del paradigma de la pirodiversidad. Conservation Biology, 20(6):1610–1619, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

PostVanDerBug:2006:HCA

[PA06b] Max Post Van Der Bug and J. Tyre Andrew. An honest and complete assessment. Conservation Biology, 21(2):1828, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pelicice:2008:FPF

[PA08] Fernando Mayer Pelicice and Angelo Antonio Agostinho. Fishpassage facilities as ecological traps in large neotropical rivers. Instalaciones para el paso de peces como trampas ecológicas en ríos neotropicales grandes. Conservation Biology, 22(1):180–188, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Philpott:2008:BLL

[PAA^{+}08] Stacy M. Philpott, Wayne J. Arendt, Inge Armbrecht, Peter Bichier, Thomas V. Diestch, Caleb Gordon, Russell Greenberg, Ivette Perfecto, Roberto Reynoso-Santos, Lorena Soto-Pinto, Cesar Tejeda-Cruz, Guadalupe Williams-Linera, Jorge Valenzuela, and José Manuel Zolotoff. Biodiversity loss in Latin American coffee landscapes: Review of the evidence on ants, birds, and trees. Pérdida de biodiversidad en paisajes cafetaleros en Latinoamérica: Revisión de la evidencia en hormigas, aves y Árboles. Conservation Biology, 22(5):1093–1105, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paterson:2008:MAT

[PAB^{+}08] James S. Paterson, Miguel B. Araújo, Pam M. Berry, Jake M. Piper, and Mark D. A. Rounsevell. Mitigation, adaptation, and the threat to biodiversity. Conservation Biology, 22(5):1352– 1355, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paehlke:2005:SBC

[Pae05] Robert Paehlke. Sustainability as a bridging concept. Conservation Biology, 19(1):36–38, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Price:2004:UGW

[PAHM04] Peter W. Price, Warren G. Abrahamson, Mark D. Hunter, and George Melika. Using gall wasps on oaks to test broad ecological concepts. Utilización de avispas en robles para probar conceptos ecológicos generales. Conservation Biology, 18(5):1405–1416, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pathikonda:2009:IDC

[PAHM09] Sharmila Pathikonda, Azmy S. Ackleh, Karl H. Hasenstein, and Susan Mopper. Invasion, disturbance, and competition: Modeling the fate of coastal plant populations. Invasión, perturbación y competencia: Modelando el destino de poblaciones de plantas costeras. Conservation Biology, 23(1):164–173, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palmer:2005:EVE

[Pal05] Clare Palmer. Environmental virtue ethics then and now. Conservation Biology, 19(5):1674–1675, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Panzer:2002:CPB

[Pan02] Ron Panzer. Compatibility of prescribed burning with the conservation of insects in small, isolated prairie reserves. Compatibilidad de incendios reglamentados con la conservación de insectos en reservas de praderas Pequeñas y aisladas. Conservation Biology, 16(5):1296–1307, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pandey:2003:CRC

[Pan03] Deep Narayan Pandey. Cultural resources for conservation science. Conservation Biology, 17(2):633–635, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pandit:2009:OFW

[Pan09] M. K. Pandit. Other factors at work in the melting Himalaya: Follow-up to Xu et al. Conservation Biology, 23(6):1346–1347, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [XGS^{+}09].

Paquette:2007:ICH

[Paq07] Sébastien Rioux Paquette. Importance of the “Crocodile hunter” phenomenon. Conservation Biology, 21(1):6, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parker:2001:IEE

[Par01] Ingrid M. Parker. Invasion ecology: Echoes of elton in the twenty-first century. Conservation Biology, 15(3):806–808, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parker:2002:TFG

[Par02] V. Thomas Parker. Transforming forests with good intentions. Conservation Biology, 16(5):1444–1446, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parejko:2003:PES

[Par03] Ken Parejko. Pliny the Elder’s Silphium: First recorded species extinction. Conservation Biology, 17(3):925–927, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parmesan:2008:VCC

[Par08] Camille Parmesan. Visualization of climate-change basics. Conservation Biology, 22(3):805–806, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Passell:2000:CTN

[Pas00] Howard Passell. A common thread to a new global paradigm. Conservation Biology, 14(4):924–925, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paton:2000:DBP

[Pat00] David C. Paton. Disruption of bird-plant pollination systems in Southern Australia. Conservation Biology, 14(5):1232–1234, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paton:2008:SC

[Pat08] Peter W. C. Paton. Superman of conservation. Conservation Biology, 22(5):1362–1363, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paul:2008:CAU

[Pau08] Ellen Paul. Clarification on amendment to U.S. migratory bird treaty act. Conservation Biology, 22(3):508, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Powell:2004:HLC

[PB04] George V. N. Powell and Robin D. Bjork. Habitat linkages and the conservation of tropical biodiversity as indicated by seasonal migrations of three-wattled bellbirds. Conexiones de hábitat y la conservación de la biodiversidad tropical indicada por migraciones estacionales de Procnias tricarunculata. Conservation Biology, 18(2):500–509, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pressey:2008:OTE

[PB08] Robert L. Pressey and Madeleine C. Bottrill. Opportunism, threats, and the evolution of systematic conservation planning. Conservation Biology, 22(5):1340–1345, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Persha:2009:MDM

[PB09] Lauren Persha and Tom Blomley. Management decentralization and montane forest conditions in Tanzania. Descentralización del manejo y condiciones del bosque montano en Tanzania. Conservation Biology, 23(6):1485–1496, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pullin:2009:CFE

[PBC^{+}09] Andrew S. Pullin, András Báldi, Ozgun Emre Can, Martin Dieterich, Vassiliki Kati, Barbara Livoreil, Gabor Lövei, Barbara Mihók, Owen Nevin, Nuria Selva, and Isabel Sousa-Pinto. Conservation focus on Europe: Major conservation policy issues that need to be informed by conservation science. Enfoque de conservación en Europa: Temas mayores de políticas de conservación que requieren información de la ciencia de la conservación. Conservation Biology, 23(4):818–824, August 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Penman:2009:POM

[PBK09] Trent D. Penman, Doug L. Binns, and Rodney P. Kavanagh. Patch-occupancy modeling as a method for monitoring changes in forest floristics: a case study in Southeastern Australia. Modelado de la ocupación de fragmentos como método para el monitoreo de cambios en la Florística de bosques: Un estudio de caso en el sureste de Australia. Conservation Biology, 23(3):740–749, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Potvin:2000:MHS

[PBL00] François Potvin, Louis Bélanger, and Kim Lowell. Marten habitat selection in a clearcut boreal landscape. Selección de hábitat por la marta en un paisaje boreal talado. Conservation Biology, 14(3):844–857, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peery:2004:ADP

[PBN^{+}04] M. Zachariah Peery, Steven R. Beissinger, Scott H. Newman, Esther B. Burkett, and Tony D. Williams. Applying the declining population paradigm: Diagnosing causes of poor reproduction in the marbled murrelet. Aplicación del paradigma de la población en Disminucioń: Diagnosis de las causas de la reproducción deficiente de Brachyramphus marmoratus. Conservation Biology, 18(4):1088–1098, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parry:2009:HST

[PBP09] Luke Parry, Jos Barlow, and Carlos A. Peres. Hunting for sustainability in tropical secondary forests. Cacería para la sustentabilidad en bosques tropicales secundarios. Conservation Biology, 23(5):1270–1280, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Philpott:2007:FTE

[PBRG07] Stacy M. Philpott, Peter Bichier, Robert Rice, and Russell Greenberg. Field-testing ecological and economic benefits of coffee certification programs. Pruebas en campo de los beneficios ecológicos y económicos de los programas de certificación de café. Conservation Biology, 21(4):975–985, August 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:2003:API

[PBS03] Ben L. Phillips, Gregory P. Brown, and Richard Shine. Assessing the potential impact of cane toads on Australian snakes. Evaluación del impacto potencial de sapos sobre serpientes Australianas. Conservation Biology, 17(6):1738–1747, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palsboll:2006:DRL

[PBSR06] Per J. Palsbøll, Martine Bérubé, Hans J. Skaug, and Caroline Raymakers. DNA registers of legally obtained wildlife and derived products as means to identify illegal takes. Registros de ADN de vida silvestre y productos derivados obtenidos legalmente como medio para identificar capturas ilegales. Conservation Biology, 33(5):1284–1293, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pressey:2001:RSA

[PC01] R. L. Pressey and R. M. Cowling. Reserve selection algorithms and the real world. Conservation Biology, 15(1):275–277, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paton:2002:UPP

[PC02] Peter W. C. Paton and William B. Crouch III. Using the phenology of pond-breeding amphibians to develop conservation strategies. Uso de la fenología de anfibios que se reproducen en estanques para desarrollar estrategias de conservación. Conservation Biology, 16(1):194–204, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pillon:2007:TEE

[PC07] Yohan Pillon and Mark W. Chase. Taxonomic exaggeration and its effects on orchid conservation. La exageración taxonómica y sus efectos sobre la conservación de orquídeas. Conservation Biology, 21(1):263–265, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Picco:2008:ACL

[PC08] Angela M. Picco and James P. Collins. Amphibian commerce as a likely source of pathogen pollution. El comercio de anfibios como una probable fuente de contaminación por patógenos. Conservation Biology, 22(6):1582–1589, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2003:SPT

[PCC03] Garry D. Peterson, Graeme S. Cumming, and Stephen R. Carpenter. Scenario planning: a tool for conservation in an uncertain world. Planificación de un escenario: una herramienta para la conservación en un mundo incierto. Conservation Biology, 17 (2):358–366, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pain:2003:CET

[PCD^{+}03] D. J. Pain, A. A. Cunningham, P. F. Donald, J. W. Duckworth, D. C. Houston, T. Katzner, J. Parry-Jones, C. Poole, V. Prakash, P. Round, and R. Timmins. Causes and effects of temporospatial declines of Gyps vultures in Asia. Conservation Biology, 17(3):661–671, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ponder:2001:EMC

[PCFC01] W. F. Ponder, G. A. Carter, P. Flemons, and R. R. Chapman. Evaluation of museum collection data for use in biodiversity assessment. Análisis del uso de datos de colecciones de museo en la evaluación de la biodiversidad. Conservation Biology, 15(3): 648–657, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pergams:2004:LCA

[PCHN04] Oliver R. W. Pergams, Brian Czech, J. Christopher Haney, and Dennis Nyberg. Linkage of conservation activity to trends in the U.S. economy. Vinculación de la conservación con tendencias de la economía de E.U.A. Conservation Biology, 18(6):1617–1623, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Park:2003:CSE

[PCL^{+}03] Young-Seuk Park, Jianbo Chang, Sovan Lek, Wenxuan Cao, and Sebastien Brosse. Conservation strategies for endemic fish species threatened by the Three Gorges Dam. Estrategias de conservación para especies de peces endémicas amenazadas por la Presa Three Gorges. Conservation Biology, 17(6):1748–1758, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2008:HLC

[PCL08] M. Nils Peterson, Xiaodong Chen, and Jianguo Liu. Household location choices: Implications for biodiversity conservation. Selecciones de ubicación de viviendas: Implicaciones para la conservación de la biodiversidad. Conservation Biology, 22(4):912– 921, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Poulsen:2009:BSC

[PCME09] J. R. Poulsen, C. J. Clark, G. Mavah, and P. W. Elkan. Bushmeat supply and consumption in a tropical logging concession in Northern Congo. Acopio y consumo de carne silvestre en una concesión maderera tropical en el norte de Congo. Conservation Biology, 23(6):1597–1608, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pierce:2006:QRE

[PCVC06] Simon Pierce, Roberta M. Ceriani, Mauro Villa, and Bruno Cerabolini. Quantifying relative extinction risks and targeting intervention for the orchid flora of a natural park in the European prealps. Cuantificación de los riesgos relativos de extinción y de la Intervención para la flora de orquídeas de un parque natural en los Prealpes europeos. Conservation Biology, 21(2):1804–1810, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pennock:2000:DPS

[PD00] David S. Pennock and Walter W. Dimmick. Distinct population segments and Congressional intent: Reply to Waples. Conservation Biology, 14(2):567–569, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [WABT08].

Peterson:2003:CIM

[PD03a] James T. Peterson and Jason Dunham. Combining inferences from models of capture efficiency, detectability, and suitable habitat to classify landscapes for conservation of threatened bull trout. Combinación de inferencias de modelos de eficiencia de captura, detectabilidad y hábitat adecuado para clasificar paisajes para la conservación de Salvelinus confluentus. Conservation Biology, 17(4):1070–1077, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Philpott:2003:CCG

[PD03b] Stacy M. Philpott and Thomas Dietsch. Coffee and conservation: a global context and the value of Farmer involvement. Conservation Biology, 17(6):1844–1846, December 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palomares:2000:ILF

[PDF^{+}00] F. Palomares, M. Delibes, P. Ferreras, J. M. Fedriani, J. Calzada, and E. Revilla. Iberian lynx in a fragmented landscape: Predispersal, dispersal, and postdispersal habitats. El lince Ibérico en un paisaje fragmentado: Hábitats seleccionados antes, durante y después de la dispersión. Conservation Biology, 14(3):809–818, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:2000:WPL

[PDH00] Francis E. Putz, Dennis P. Dykstra, and Rudolf Heinrich. Why poor logging practices persist in the tropics. Porque persisten las prácticas madereras deficientes en los trópicos. Conservation Biology, 14(4):951–956, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pascal:2008:MOT

[PDLS08] Michel Pascal, Bertrand Richer De Forges, Hervé Le Guyader, and Daniel Simberloff. Mining and other threats to the New Caledonia biodiversity hotspot. Conservation Biology, 22(2): 498–499, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pitman:2007:WAA

[PDS^{+}07] Nigel C. A. Pitman, María Del Carmen Loyola Azáldegui, Karina Salas, Gabriela T. Vigo, and David A. Lutz. Written accounts of an Amazonian landscape over the last 450 years. Documentos sobre un paisaje amazónico durante los Últimos 450 años. Conservation Biology, 21(1):253–262, February 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Patten:2001:CVR

[PE01] Michael A. Patten and Richard A. Erickson. Conservation value and rankings of exotic species. Conservation Biology, 15(4):817– 818, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pearman:2001:CVI

[Pea01] Peter B. Pearman. Conservation value of independently evolving units: Sacred cow or testable hypothesis? Conservation Biology, 15(3):780–783, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Posadas:2001:UPD

[PEC01] Paula Posadas, Daniel R. Miranda Esquivel, and Jorge V. Crisci. Using phylogenetic diversity measures to set priorities in conservation: an example from Southern South America. Utilización de medidas de diversidad filogenética para definir prioridades en conservación: un ejemplo del sur de América del sur. Conservation Biology, 15(5):1325–1334, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pedynowski:2003:PEM

[Ped03] Dena Pedynowski. Prospects for ecosystem management in the crown of the continent ecosystem, Canada–United States: Survey and recommendations. Perspectivas para el manejo de ecosistemas en el ecosistema corona del continente, Canadá– Estados Unidos: Estudio y recomendaciones. Conservation Biology, 17(5):1261–1269, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pederson:2007:CSI

[Ped07] Judith Pederson. Connecting species invasions with ecological theory. Conservation Biology, 21(4):1127–1128, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pemberton:2001:CVD

[Pem01] Robert W. Pemberton. The conflicting values and differing priorities of biological control. Conservation Biology, 15(4):1192– 1194, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peres:2000:ESH

[Per00] Carlos A. Peres. Effects of subsistence hunting on vertebrate community structure in Amazonian forests. Efectos de la cacería de subsistencia sobre la estructura de la comunidad de vertebrados en bosques amazónicos. Conservation Biology, 14(1): 240–253, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peres:2001:SES

[Per01] Carlos A. Peres. Synergistic effects of subsistence hunting and habitat fragmentation on Amazonian forest vertebrates. Efectos sinergistas de la cacería de subsistencia y la fragmentación del hábitat sobre vertebrados de bosques en la Amazonía. Conservation Biology, 15(6):1490–1505, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perkins:2002:LCD

[Per02] John P. Perkins. Like a cool drink of water. Conservation Biology, 16(3):852–853, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peres:2005:WWN

[Per05a] Carlos A. Peres. Why we need megareserves in Amazonia. Porque necesitamos megareservas amazónicas. Conservation Biology, 19(3):728–733, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perez:2005:WSC

[Pér05b] Hector E. Pérez. What students can do to improve graduate education in conservation biology. Conservation Biology, 19(6): 2033–2035, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perry:2001:IHB

[PFL01] William L. Perry, Jeffrey L. Feder, and David M. Lodge. Implications of hybridization between introduced and resident Orconectes crayfishes. Implicaciones de la hibridación de langostinos de aqua dulce Orconectes introducidos y residentes. Conservation Biology, 15(6):1656–1666, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phalan:2009:COP

[PFR^{+}09] Ben Phalan, Emily B. Fitzherbert, Sylvain Rafflegeau, Matthew J. Struebig, and Aude Verwilghen. Conservation in oil-palm landscapes. Conservation Biology, 23(2):244–245, April 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parker:2007:ITB

[PGD^{+}07] Ingrid M. Parker, Eduardo C. García, Katrina M. Dlugosch, Pete Holloran, Sarah M. Swope, Kris B. Hulvey, and Robert Clark. Invasions as a tool for basic research. Conservation Biology, 21(5):1376–1377, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pellet:2004:CAI

[PGP04] Jérôme Pellet, Antoine Guisan, and Nicolas Perrin. A concentric analysis of the impact of urbanization on the threatened European tree frog in an agricultural landscape. Un análisis concéntrico del impacto de la urbanización sobre la rana arborícola europea amenazada en un paisaje agrícola. Conservation Biology, 18(6):1599–1606, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parks:2002:RSL

[PH02] S. A. Parks and A. H. Harcourt. Reserve size, local human density, and mammalian extinctions in U.S. protected areas tamaño de reserva, densidad humana local y extinción de mamíferos en Áreas protegidas de los Estados Unidos. Conservation Biology, 16(3):800–808, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:2000:PTK

[Phi00] Stephen S. Phillips. Population trends and the koala conservation debate tendencias. Poblacionales y el debate sobre la conservación del koala. Conservation Biology, 14(3):650–659, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peacock:2005:PDF

[PHW05] Evan Peacock, Wendell R. Haag, and Melvin L. Warren, Jr. Prehistoric decline in freshwater mussels coincident with the advent of maize agriculture. Declinación prehistórica en mejillones de agua dulce coincidente con la llegada de la agricultura de maíz. Conservation Biology, 19(2):547–551, April 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Plathong:2000:ESG

[PIH00] Sakanan Plathong, Graeme J. Inglis, and Michael E. Huber. Effects of self-guided snorkeling trails on corals in a tropical marine park. Efectos de senderos para el buceo libre con esnorquel en los corales dentro de un parque marino. Conservation Biology, 14(6):1821–1830, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pimm:2006:PGI

[Pim06] Stuart Pimm. The perfect gift to inspire your students and make your mother love you even more. Conservation Biology, 21(3): 923, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pimentel:2007:AP

[Pim07a] David Pimentel. An alien perspective. Conservation Biology, 21 (4):1128–1129, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pimm:2007:ASD

[Pim07b] Stuart L. Pimm. Africa: Still the “Dark Continent”. Conservation Biology, 21(3):567–569, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parendes:2000:RLA

[PJ00] Laurie A. Parendes and Julia A. Jones. Role of light availability and dispersal in exotic plant invasion along roads and streams in the H. J. Andrews Experimental Forest, Oregon. Disponibilidad de luz y mechanismos de dispersión e invasión de plantas exóticas a lo largo de caminos y arroyos en el Bosque Experimental H. J. Andrews, Oregon. Conservation Biology, 14(1):64–75, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perrings:2006:BAL

[PJB^{+}06] Charles Perrings, Louise Jackson, Kamal Bawa, Lijbert Brussaard, Stephen Brush, Tom Gavin, Roberto Papa, Unai Pascual, and Peter De Ruiter. Biodiversity in agricultural landscapes: Saving natural capital without losing interest. Conservation Biology, 20(2):263–264, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pysek:2003:INA

[PJK03] Petr Pyšek, Vojtěch Jarošík, and Tomáš Kučera. Inclusion of native and alien species in temperate nature reserves: an historical study from Central Europe. Inclusión de especies nativas y exóticas en reservas naturales templadas: un estudio histórico de Europa Central. Conservation Biology, 17(5):1414–1424, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pedersen:2007:IDE

[PJNA07] Amy B. Pedersen, Kate E. Jones, Charles L. Nunn, and Sonia Altizer. Infectious diseases and extinction risk in wild mammals. Enfermedades infecciosas y riesgo de extinción en mamíferos silvestres. Conservation Biology, 21(5):1269–1279, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pitman:2002:ERE

[PJW^{+}02] Nigel C. A. Pitman, Peter M. Jørgensen, Robert S. R. Williams, Susana León-Yánez, and Renato Valencia. Extinction-rate estimates for a modern neotropical flora. Estimaciones de tasas de extinción para una flora moderna neotropical. Conservation Biology, 16(5):1427–1431, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perry:2004:FSF

[PJYO04] David A. Perry, Huang Jing, Andrew Youngblood, and Doug R. Oetter. Forest structure and fire susceptibility in volcanic landscapes of the Eastern High Cascades, Oregon. Estructura del bosque y susceptibilidad al fuego en paisajes volcánicos de Cascadas Altas del Este, Oregon. Conservation Biology, 18(4):913– 926, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pullin:2001:ECP

[PK01] Andrew S. Pullin and Teri M. Knight. Effectiveness in conservation practice: Pointers from medicine and public health. Efectividad de la conservación Práctica: Indicadores de medicina y salud Pública. Conservation Biology, 15(1):50–54, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pullin:2005:ACM

[PK05] Andrew S. Pullin and Teri M. Knight. Assessing conservation Management’s evidence base: a survey of management-plan compilers in the United Kingdom and Australia. Evaluación de la base de evidencias de la gestión de conservación: un sondeo de compiladores de planes de gestión en el Reino Unido y Australia. Conservation Biology, 19(6):1989–1996, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perhans:2008:CGR

[PKB^{+}08] Karin Perhans, Claes Kindstrand, Mattias Boman, Line Boberg Djupström, Lena Gustafsson, Leif Mattsson, Leif Martin Schroeder, Jan Weslien, and Sofie Wikberg. Conservation goals and the relative importance of costs and benefits in reserve selection. Metas de conservación y la importancia relativa de los costos y beneficios en la selección de reservas. Conservation Biology, 22(5):1331–1339, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primack:2000:DPR

[PKM00] Richard Primack, Hiromi Kobori, and Seiwa Mori. Dragonfly pond restoration promotes conservation awareness in Japan. Conservation Biology, 14(5):1553–1554, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peres:2003:ENR

[PL03] Carlos A. Peres and Iain R. Lake. Extent of nontimber resource extraction in tropical forests: Accessibility to game vertebrates by hunters in the Amazon Basin. Alcance de la extracción de recursos no maderables en bosques tropicales: Acceso a vertebrados de caza por cazadores en la Cuenca del Amazonas. Conservation Biology, 17(2):521–535, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2006:BCG

[PL06] David L. Peterson and Jeremy Littell. Biological change in the global greenhouse. Conservation Biology, 20(1):255, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pergams:2000:CMA

[PLA00] Oliver R. W. Pergams, Robert C. Lacy, and Mary V. Ashley. Conservation and management of Anacapa Island deer mice. Conservación y manejo del ratón Peromyscus maniculatus anacapae de la Isla Anacapa. Conservation Biology, 14(3): 819–832, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pfaller:2009:NSS

[PLB09] Joseph B. Pfaller, Colin J. Limpus, and Karen A. Bjorndal. Nest-site selection in individual loggerhead turtles and consequences for doomed-egg relocation. Selección de sitios de nidación de individuos de tortugas de Carey y sus consecuencias para la reubicación de huevos vulnerables. Conservation Biology, 23(1):72–80, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perez-Losada:2009:CAS

[PLBBJC09] M. Pérez-Losada, G. Bond-Buckup, C. G. Jara, and K. A. Crandall. Conservation assessment of Southern South American freshwater ecoregions on the basis of the distribution and genetic diversity of crabs from the genus Aegla. Evaluación de la conservación de ecoregiones dulceacuícolas del sur de Sudamérica con base en la distribución y diversidad genética de cangrejos del género Aegla. Conservation Biology, 23(3):692–702, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2005:WLT

[PLL^{+}05] M. Nils Peterson, Roel R. Lopez, Edward J. Laurent, Philip A. Frank, Nova J. Silvy, and Jianguo Liu. Wildlife loss through domestication: the case of endangered key deer. Pérdida de vida silvestre por domesticación: El caso de Odocoileus virginianus clavium. Conservation Biology, 19(3):939–944, June 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pyke:2005:CGM

[PM05] Christopher R. Pyke and Jaymee Marty. Cattle grazing mediates climate change impacts on ephemeral wetlands. El apacentamiento de ganado influye en los impactos del cambio climático sobre humedales Efímeros. Conservation Biology, 19(5):1619– 1625, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pautasso:2007:BER

[PM07] Marco Pautasso and Michael L. Mckinney. The botanist effect revisited: Plant species richness, county area, and human population size in the United States. El efecto botánico revisado: Riqueza de especies de plantas, superficie del condado y tamaño de la población humana en los Estados Unidos. Conservation Biology, 21(5):1333–1340, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pidgeon:2001:RAC

[PMBN01] A. M. Pidgeon, N. E. Mathews, R. Benoit, and E. V. Nordheim. Response of avian communities to historic habitat change in the Northern Chihuahuan desert. Respuesta de las comunidades de aves a cambios históricos de hábitat en el norte del desierto de Chihuahua. Conservation Biology, 15(6):1772–1789, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Poiani:2001:ICP

[PMC01] Karen A. Poiani, Michael D. Merrill, and Kim A. Chapman. Identifying conservation-priority areas in a fragmented Minnesota landscape based on the umbrella species concept and selection of large patches of natural vegetation. Identificación de areas de conservación prioritarias en un paisaje fragmentado en Minnesota con base en el concepto de especie sombrilla y la selección de manchones de vegetación natural. Conservation Biology, 15(2):513–522, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pejchar:2007:EPC

[PMC^{+}07] Liba Pejchar, Peter M. Morgan, Margaret R. Caldwell, Carl Palmer, and Gretchen C. Daily. Evaluating the potential for conservation development: Biophysical, economic, and institutional perspectives. Evaluación del potencial de la desarrollo para la conservación: Perspectivas biofísicas, económicas e institucionales. Conservation Biology, 21(1):69–78, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Posadas:2004:WTA

[PMEC04] Paula Posadas, Daniel Rafael Miranda-Esquivel, and Jorge V. Crisci. On words, tests, and applications: Reply to Faith et al. Conservation Biology, 18(1):262–266, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pineda:2005:FBD

[PMEH05] Eduardo Pineda, Claudia Moreno, Federico Escobar, and Gonzalo Halffter. Frog, bat, and dung beetle diversity in the cloud forest and coffee agroecosystems of Veracruz, Mexico. Diversidad de ranas, murciélagos y escarabajos del estiércol en el bosque de niebla y agrosistemas de café en Veracruz, México. Conservation Biology, 19(2):400–410, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phinney:2001:URS

[PMKDS01] Jonathan T. Phinney, Frank Muller-Karger, Phil Dustan, and Jack Sobel. Using remote sensing to reassess the mass mortality of Diadema antillarum 1983–1984. Utilización de percepción remota para reanalizar la mortandad masiva de Diadema antillarum de 1983–1984. Conservation Biology, 15(4):885–891, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peters:2006:ESL

[PMZ06] Sandra L. Peters, Jay R. Malcolm, and Barbara L. Zimmerman. Effects of selective logging on bat communities in the Southeastern Amazon. Efectos de la tala selectiva sobre comunidades de murciélagos en el sureste de la Amazonía. Conservation Biology, 20(5):1410–1421, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pergams:2005:EPL

[PN05] Oliver R. W. Pergams and Dennis W. Nyberg. Evaluating the predicted local extinction of a once-common mouse. Evaluación del pronóstico de la extinción local de un ratón anteriormente común. Conservation Biology, 19(4):1312–1317, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Podolsky:2001:GVM

[Pod01] Robert H. Podolsky. Genetic variation for morphological and allozyme variation in relation to population size in Clarkia dudleyana, an endemic annual. Variación genética en la variación morfológica y alozímica en relación con el tamaño poblacional de Clarkia dudleyana, una especie anual endémica. Conservation Biology, 15(2):412–423, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pollock:2006:DPD

[Pol06] Jacob F. Pollock. Detecting population declines over large areas with presence-absence, time-to-encounter, and count survey methods. Detección de declinaciones poblacionales en Áreas extensas mediante métodos muestreo de presencia-ausencia, tiempo al encuentro y conteo. Conservation Biology, 21(3):882– 892, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Poonswad:2000:CAB

[Poo00] Pilai Poonswad. Cataloguing avian biodiversity. Conservation Biology, 14(2):588–589, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pope:2008:ACA

[Pop08] Karen L. Pope. Assessing changes in amphibian population dynamics following experimental manipulations of introduced fish. Evaluación de cambios en la dinámica poblacional de anfibios después de manipulaciones experimentales de peces introducidos. Conservation Biology, 22(6):1572–1581, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Portman:2008:MSR

[Por08] Michelle E. Portman. Man and the sea revisited. Conservation Biology, 22(3):808–809, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Povilitis:2001:TRN

[Pov01] Tony Povilitis. Toward a robust natural imperative for conservation. Conservation Biology, 15(2):533–535, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Power:2006:PTS

[Pow06] Thomas Michael Power. Public timber supply, market adjustments, and local economies: Economic assumptions of the northwest forest plan. Existencia de madera pública, ajustes de mercado y economías locales: Suposiciones económicas del plan forestal del noroeste. Conservation Biology, 20(2):341–350, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

President:2001:AC

[PP01] Reed Noss Past President and Malcolm Hunter President. From assemblage to community. Conservation Biology, 15(5):1201– 1202, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petit:2003:EIH

[PP03] Lisa J. Petit and Daniel R. Petit. Evaluating the importance of human-modified lands for neotropical bird conservation. Evaluación de la importancia de tierras modificadas por humanos para la conservación de aves neotropicales. Conservation Biology, 17 (3):687–694, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perry:2008:IIB

[PP08a] Dan Perry and Gad Perry. Improving interactions between animal rights groups and conservation biologists. Mejorando las interacciones entre grupos Pro derechos de los animales y biólogos de la consevación. Conservation Biology, 22(1):27–35, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perry:2008:ARC

[PP08b] Gad Perry and Dan Perry. Animal rights and conservation. Conservation Biology, 22(4):816–818, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pestes:2008:BDA

[PPBW08] Lynsey R. Pestes, Randall M. Peterman, Michael J. Bradford, and Chris C. Wood. Bayesian decision analysis for evaluating management options to promote recovery of a depleted salmon population. Análisis de decisión Bayesiano para la evaluación de opciones de manejo para promover la recuperación de una población de salmón disminuida. Conservation Biology, 22(2): 351–361, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paoli:2001:EEA

[PPLS01] Gary D. Paoli, David R. Peart, Mark Leighton, and Ismayadi Samsoedin. An ecological and economic assessment of the nontimber forest product gaharu wood in Gunung Palung National Park, West Kalimantan, Indonesia. Evaluación ecológica y económica de la madera gaharu, un producto forestal no maderable, en el Parque Nacional Gunung Palung, Kalimantan Occidental, Indonesia. Conservation Biology, 15(6):1721–1732, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2005:CMC

[PPP05] M. Nils Peterson, Markus J. Peterson, and Tarla Rai Peterson. Conservation and the myth of consensus. Conservación y el mito del consenso. Conservation Biology, 19(3):762–767, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2006:WCN

[PPP06] M. Nils Peterson, Markus J. Peterson, and Tarla Rai Peterson. Why conservation needs dissent. Conservation Biology, 20(2): 576–578, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putzel:2008:CTT

[PPPV08] Louis Putzel, Christine Padoch, and Miguel Pinedo-Vasquez. The Chinese timber trade and the logging of Peruvian Amazonia. Conservation Biology, 22(6):1659–1661, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Povilitis:2006:BIM

[PPRB06] Anthony Povilitis, David R. Parsons, Michael J. Robinson, and C. Dusti Becker. The bureaucratically imperiled Mexican wolf. Conservation Biology, 20(4):942–945, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pusey:2007:CLT

[PPW^{+}07] Anne E. Pusey, Lilian Pintea, Michael L. Wilson, Shadrack Kamenya, and Jane Goodall. The contribution of long-term research at Gombe National Park to chimpanzee conservation. La Contribución de la investigación de largo plazo en el Parque Nacional Gombe a la conservación del chimpancé. Conservation Biology, 21(3):623–634, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pruett:2009:ABP

[PPW09] Christin L. Pruett, Michael A. Patten, and Donald H. Wolfe. Avoidance behavior by prairie grouse: Implications for development of wind energy. Conducta elusiva por urogallos de pradera: Implicaciones para el desarrollo de energía eólica. Conservation Biology, 23(5):1253–1259, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phifer:2001:CPG

[PR01a] Paul Phifer and Paul Roebuck. The complexity of population growth: Reply to Pletscher and Schwartz. Conservation Biology, 15(6):1810–1811, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [PS00b] and reply [PS01b].

Putz:2001:BTC

[PR01b] Francis E. Putz and Claudia Romero. Biologists and timber certification. Conservation Biology, 15(2):313–314, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2003:UEN

[PR03] A. Townsend Peterson and C. Richard Robins. Using ecologicalniche modeling to predict barred owl invasions with implications for spotted owl conservation. Uso de modelos de Nicho– Ecológico para predecir invasiones de Strix varia y sus implicaciones para la conservación de Strix occidentalis. Conservation Biology, 17(4):1161–1165, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prager:2008:CLF

[PR08] Michael H. Prager and Andrew A. Rosenberg. Closing the loop in fishery management: the importance of instituting regular independent management review. Conservation Biology, 22(4): 830–831, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prato:2005:AUM

[Pra05] Tony Prato. Accounting for uncertainty in making species protection decisions. Dando cuenta de la incertidumbre en la toma de decisiones sobre la protección de especies. Conservation Biology, 19(3):806–814, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pressey:2004:CPB

[Pre04] R. L. Pressey. Conservation planning and biodiversity: Assembling the best data for the job. Conservation Biology, 18(6): 1677–1681, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pretty:2006:WWT

[Pre06] Jules Pretty. Where the wild things were. Conservation Biology, 33(5):1325–1326, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Preston:2009:CL

[Pre09] Christopher J. Preston. Carson’s legacy. Conservation Biology, 23(2):511–512, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primack:2001:PAA

[Pri01a] Richard B. Primack. Publish again in another language. Conservation Biology, 15(1):290–291, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primack:2001:R

[Pri01b] Richard B. Primack. Response. Conservation Biology, 15(6): 1477–1478, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primack:2003:ECB

[Pri03] Richard B. Primack. Evaluating conservation biology textbooks. Conservation Biology, 17(5):1202–1203, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pahkala:2002:LSE

[PRL^{+}02] Maarit Pahkala, Katja Räsänen, Anssi Laurila, Ulf Johanson, Lars Olof Björn, and Juha Merilä. Lethal and sublethal effects of UV-B/pH synergism on common frog embryos. Efectos letales y subletales del sinergismo B-UV/pH en embriones de rana común. Conservation Biology, 16(4):1063–1073, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parker:2003:EAU

[PRL03] Ingrid M. Parker, Joseph Rodriguez, and Michael E. Loik. An evolutionary approach to understanding the biology of invasions: Local adaptation and general-purpose genotypes in the weed Verbascum thapsus. Una aproximación evolutiva para entender la biología de las invasiones: Adaptación local y genotipos de Propó-sito general en la hierba Verbascum thapsus. Conservation Biology, 17(1):59–72, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:2006:RSA

[PRMV06] Oliver L. Phillips, Sam Rose, Abel Monteagudo Mendoza, and Percy Núñez Vargas. Resilience of Southwestern Amazon forests to anthropogenic edge effects. Resiliencia de bosques del suroeste de la Amazonía a efectos de borde antropogénicos. Conservation Biology, 20(6):1698–1710, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Proffitt:2004:LAS

[Pro04] W. Trexler Proffitt, Jr. Legitimacy and adoption of a scientific biological control program: an institutional analysis of hoddle. Conservation Biology, 18(1):58–61, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pacheco:2000:GSM

[PS00a] Luis F. Pacheco and Javier A. Simonetti. Genetic structure of a mimosoid tree deprived of its seed disperser, the spider monkey. Estructura genética de un arbol Mimósido privado de su dispersor de semillas, el marimono. Conservation Biology, 14 (6):1766–1775, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pletscher:2000:TPG

[PS00b] Daniel H. Pletscher and Michael K. Schwartz. The tyranny of population growth. Conservation Biology, 14(6):1918–1919, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [PR01a] and reply [PS01b].

Perez-Salicrup:2001:CTS

[PS01a] Diego R. Pérez-Salicrup. A conservation tool: Software to identify tropical tree species. Conservation Biology, 15(2):544–546, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pletscher:2001:RDC

[PS01b] Daniel H. Pletscher and Michael K. Schwartz. Rearranging the deck chairs on the Malthusian ship: Reply to Phifer and Roebuck. Conservation Biology, 15(6):1812–1813, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pretty:2004:SCB

[PS04] Jules Pretty and David Smith. Social capital in biodiversity conservation and management. Capital social en la conservación y gestión de la biodiversidad. Conservation Biology, 18(3):631– 638, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Padoch:2005:CCP

[PS05] Christine Padoch and Robin R. Sears. Conserving concepts: in praise of sustainability. Conservation Biology, 19(1):39–41, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pullin:2006:GSR

[PS06] Andrew S. Pullin and Gavin B. Stewart. Guidelines for systematic review in conservation and environmental management. Directrices para la revisión sistemática en gestión ambiental y de conservación. Conservation Biology, 20(6):1647–1656, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parker:2005:EAE

[PSBG05] Timothy H. Parker, Brooke M. Stansberry, C. Dustin Becker, and Philip S. Gipson. Edge and area effects on the occurrence of migrant forest songbirds. Efectos de borde y de Área sobre la ocurrencia de aves canoras migratorias. Conservation Biology, 19(4):1157–1167, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peer:2005:VCC

[PSF05] Guy Pe’er, David Saltz, and Karin Frank. Virtual corridors for conservation management. Corredores virtuales para la gestión de la conservación. Conservation Biology, 19(6):1997–2003, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Penn:2000:DFK

[PSG^{+}00] Angela M. Penn, William B. Sherwin, Greg Gordon, Daniel Lunney, Alistair Melzer, and Robert C. Lacy. Demographic forecasting in koala conservation. Predicción demográfica de la conservación del koala. Conservation Biology, 14(3):629–638, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pinsky:2009:RWS

[PSGA09] Malin L. Pinsky, Dane B. Springmeyer, Matthew N. Goslin, and Xanthippe Augerot. Range-wide selection of catchments for Pacific salmon conservation. Selección de sitios de captación en toda el Área de distribución para la conservación de salmón del Pacífico. Conservation Biology, 23(3):680–691, June 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pyke:2008:CPF

[PTP^{+}08] Christopher R. Pyke, Roxanne Thomas, Read D. Porter, Jessica J. Hellmann, Jeffrey S. Dukes, David M. Lodge, and Gabriela Chavarria. Current practices and future opportunities for policy on climate change and invasive species. Prácticas actuales y oportunidades futuras para las políticas sobre cambio climático y especies invasoras. Conservation Biology, 22(3): 585–592, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Purvis:2003:IHD

[Pur03] Andy Purvis. Implications of the “Homogecene debt”. Conservation Biology, 21(6):337–338, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:2003:CPC

[Put03] Francis E. Putz. Companies partnering with communities. Conservation Biology, 17(2):645–646, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:2005:BER

[Put05] Francis E. Putz. Black Earth from red desert and green Hell. Conservation Biology, 19(3):978–979, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:2007:PPG

[Put07] Francis E. Putz. Poverty & power, geography & corruption, biodiversity & population pressure. Conservation Biology, 22 (1):1661–1662, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perfecto:2002:QAM

[PV02] Ivette Perfecto and John Vandermeer. Quality of agroecological matrix in a tropical montane landscape: Ants in coffee plantations in Southern Mexico. Calidad de la matriz agroecológica en un paisaje montañoso tropical: Hormigas en plantaciones de café en el sur de México. Conservation Biology, 16(1):174–182, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petit:2006:EPS

[PV06] Eric Petit and Nathaniel Valiere. Estimating population size with noninvasive capture-mark-recapture data. Estimación del tamaño poblacional con datos de captura-marca-recaptura no invasivos. Conservation Biology, 33(5):1062–1073, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perot:2009:PDB

[PV09a] Aurore Pérot and Marc-André Villard. Putting density back into the habitat-quality equation: Case study of an open-nesting forest bird. Regresando la densidad a la ecuación de calidad del hábitat: Caso de estudio de una ave forestal de nidos abiertos. Conservation Biology, 23(6):1550–1557, December 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perry:2009:FRB

[PV09b] Gad Perry and Dan Vice. Forecasting the risk of brown tree snake dispersal from Guam: a mixed transport-establishment model. Prodicción del riesgo de dispersión de Boiga irregularis desde Guam: un modelo mixto de transporte–establecimiento. Conservation Biology, 23(4):992–1000, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Puth:2001:BCC

[PW01] Linda M. Puth and Karen A. Wilson. Boundaries and corridors as a continuum of ecological flow control: Lessons from rivers and streams. Linderos y corredores como un control de flujo ecológico continuo: Lecciones de ríos y arroyos. Conservation Biology, 15(1):21–30, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pykala:2000:MHE

[Pyk00] Juha Pykälä. Mitigating human effects on European biodiversity through traditional animal husbandry. Mitigación de los efectos humanos sobre la biodiversidad europea mediante la producción tradicional de animales. Conservation Biology, 14(3):705–712, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pykala:2004:ENF

[Pyk04] Juha Pykälä. Effects of new forestry practices on rare epiphytic macrolichens. Efectos de prácticas silvícolas nuevas sobre macrolíquenes epifíticos. Conservation Biology, 18(3):831–838, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pyke:2005:ASC

[Pyk05] Christopher R. Pyke. Assessing suitability for conservation action: Prioritizing interpond linkages for the California tiger salamander. Evaluación de aptitud para acciones de conservación: Priorización de conexiones intercharcas para Ambystoma californiense. Conservation Biology, 19(2):492–503, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pyne:2004:PRS

[Pyn04] Stephen J. Pyne. Pyromancy: Reading stories in the flames piromancia: Leyendo historias en las llamas. Conservation Biology, 18(4):874–877, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peres:2001:PPP

[PZ01] Carlos A. Peres and Barbara Zimmerman. Perils in parks or parks in peril? Reconciling conservation in Amazonian reserves with and without use. Conservation Biology, 15(3):793–797, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Packer:2005:CGP

[PZG^{+}05] Laurence Packer, Amro Zayed, Jennifer C. Grixti, Luisa Ruz, Robin E. Owen, Felipe Vivallo, and Haroldo Toro. Conservation genetics of potentially endangered mutualisms: Reduced levels of genetic variation in specialist versus generalist bees. Genética de la conservación de mutualismos potencialmente en peligro: Bajos niveles de variación genética en abejas especialistas versus generalistas. Conservation Biology, 19(1):195–202, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quintana-Ascencio:2003:FEP

[QAMW03] Pedro Francisco Quintana-Ascencio, Eric S. Menges, and Carl W. Weekley. A fire-explicit population viability analysis of Hypericum cumulicola in Florida rosemary scrub. Análisis de viabilidad poblacional con incendios Explícitos para Hypericum cumulicola en Breñal de romero de Florida. Conservation Biology, 17(2):433–449, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Qian:2007:RBP

[Qia07] Hong Qian. Relationships between plant and animal species richness at a regional scale in China. Relaciones de riqueza de especies entre taxa de plantas y animales a una escala regional en China. Conservation Biology, 21(4):937–944, August 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quayle:2005:CSB

[QR05] James F. Quayle and Leah R. Ramsay. Conservation status as a biodiversity trend indicator: Recommendations from a decade of listing species at risk in British Columbia. El estatus de conservación como un indicador de tendencias en biodiversidad: Recomendaciones de una Década de enlistar especies en riesgo en Columbia Británica. Conservation Biology, 19(4):1306–1311, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quayle:2006:BIB

[QR06] James F. Quayle and Leah R. Ramsay. Biodiversity indicators based on trends in conservation status: Advancing the science. Conservation Biology, 20(2):582–583, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quayle:2007:TSB

[QRF07] James F. Quayle, Leah R. Ramsay, and David F. Fraser. Trend in the status of breeding bird fauna in British Columbia, Canada, based on the IUCN red list index method. Tendencia en el estatus de la fauna de aves residentes en Columbia Británica, con base en el método del Índice de la lista roja de IUCN. Conservation Biology, 21(5):1241–1247, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberge:2004:UUS

[RA04a] Jean-Michel Roberge and Per Angelstam. Usefulness of the umbrella species concept as a conservation tool. Utilidad del concepto de especie paraguas como herramienta de conservación. Conservation Biology, 18(1):76–85, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Romero:2004:ICO

[RA04b] Claudia Romero and Germán I. Andrade. International conservation organizations and the fate of local tropical forest conservation initiatives. Conservation Biology, 18(2):578–580, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Romero:2004:L

[RA04c] Claudia Romero and Germán I. Andrade. Letters: 3. Conservation Biology, 18(6):1454–1455, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roback:2005:JUM

[RA05] Paul J. Roback and Robert A. Askins. Judicious use of multiple hypothesis tests. Uso juicioso de pruebas de hipótesis múltiples. Conservation Biology, 19(1):261–267, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2009:PES

[RA09] Kent H. Redford and William M. Adams. Payment for ecosystem services and the challenge of saving nature. Conservation Biology, 23(4):785–787, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rabosky:2009:SBM

[Rab09] Daniel L. Rabosky. Speciation in birds and more. Conservation Biology, 23(2):506–508, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Raffensperger:2003:CEP

[Raf03] Carolyn Raffensperger. Constitutional experiments: Protecting the environment and future generations. Conservation Biology, 17(6):1487–1488, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Root:2003:MAE

[RAG03] Karen V. Root, H. Resit Akçakaya, and Lev Ginzburg. A multispecies approach to ecological valuation and conservation. Una aproximación multiespecífica para la valoración ecológica y la conservación. Conservation Biology, 17(1):196–206, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rains:2000:AOF

[Rai00] Mark Cable Rains. Avoiding an ozymandian fate. Conservation Biology, 14(2):587–588, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rains:2003:RGR

[Rai03] Mark Cable Rains. The role of groundwater in resource conservation efforts. Conservation Biology, 17(3):933–934, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redpath:2004:UDM

[RAL^{+}04] S. M. Redpath, B. E. Arroyo, F. M. Leckie, P. Bacon, N. Bayfield, R. J. Gutiérrez, and S. J. Thirgood. Using decision modeling with stakeholders to reduce human–wildlife conflict: a raptor–grouse case study. Utilización de modelos de decisión con grupos de interés para reducir conflictos humanos-vida silvestre: un estudio de caso rapaz–codorniz. Conservation Biology, 18 (2):350–359, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Raman:2001:ESB

[Ram01] T. R. Shankar Raman. Effect of slash-and-burn shifting cultivation on rainforest birds in Mizoram, Northeast India. Efecto de la rotación a tumba y quema para cultivo en las aves del bosque lluvioso en Mizoram, India Nororiental. Conservation Biology, 15(3):685–698, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Randall:2000:ALL

[Ran00] Jan A. Randall. An awful lot to learn. Conservation Biology, 14 (6):1928–1929, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roux:2008:ICS

[RANM08] Dirk J. Roux, Peter J. Ashton, Jeanne L. Nel, and Heather M. Mackay. Improving cross-sector policy integration and cooperation in support of freshwater conservation. Conservation Biology, 22(6):1382–1387, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Raphael:2006:CMM

[Rap06] Martin G. Raphael. Conservation of the marbled murrelet under the northwest forest plan. Conservación de Brachyramphus marmoratus mediante el plan forestal del noroeste, E.U.A. Conservation Biology, 20(2):297–305, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2004:DBD

[RAR^{+}04] W. Douglas Robinson, George R. Angehr, Tara R. Robinson, Lisa J. Petit, Daniel R. Petit, and Jeffrey D. Brawn. Distribution of bird diversity in a vulnerable neotropical landscape. Distribución de la diversidad de aves en un paisaje neotropical vulnerable. Conservation Biology, 18(2):510–518, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rau:2002:BC

[Rau02] Jaime R. Rau. Biodiversidad y su cuantificación. (Spanish) [Biodiversity and its quantification]. Conservation Biology, 16(6): 1666–1668, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2007:UPG

[RAW07] David G. Roberts, David J. Ayre, and Robert J. Whelan. Urban plants as genetic reservoirs or threats to the integrity of bushland plant populations. Plantas urbanas como reservorios genéticos o amenazas a la integridad de poblaciones de plantas de matorrales xerófitos. Conservation Biology, 21(2):842–852, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ray:2006:FCM

[Ray06a] G. Carleton Ray. Future course for marine mammal conservation research? Conservation Biology, 21(2):1823–1825, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ray:2006:LTM

[Ray06b] G. Carleton Ray. Living tools for marine conservation. Conservation Biology, 20(1):260–262, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rylands:2005:BPA

[RB05] Anthony B. Rylands and Katrina Brandon. Brazilian protected areas. Áreas protegidas Brasileñas. Conservation Biology, 19(3): 612–618, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rondinini:2006:DUE

[RB06] Carlo Rondinini and Luigi Boitani. Differences in the umbrella effects of African amphibians and mammals based on two estimators of the area of occupancy. Diferencias en el efecto sombrilla de anfibios y mamíferos africanos con base en dos estimadores del Área de ocupación. Conservation Biology, 20(1): 170–179, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rondinini:2007:SCP

[RB07] Carlo Rondinini and Luigi Boitani. Systematic conservation planning and the cost of tackling conservation conflicts with large carnivores in Italy. Planificación sistemática de la conservación y el costo de afrontar conflictos de conservación con carnívoros mayores en Italia. Conservation Biology, 21(6):1455– 1462, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rothley:2004:CSS

[RBG^{+}04] Kristina D. Rothley, Claire N. Berger, Catherine Gonzalez, Elisabeth M. Webster, and Daniel I. Rubenstein. Combining strategies to select reserves in fragmented landscapes. Combinación de estrategias para seleccionar reservas en paisajes fragmentados. Conservation Biology, 18(4):1121–1131, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rominger:2006:BSM

[RBG06] Eric M. Rominger, Vernon C. Bleich, and Elise J. Goldstein. Bighorn sheep, mountain lions, and the ethics of conservation. Conservation Biology, 20(5):1341, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rogers-Bennett:2002:USE

[RBHKK02] Laura Rogers-Bennett, Peter L. Haaker, Konstantin A. Karpov, and David J. Kushner. Using spatially explicit data to evaluate marine protected areas for abalone in Southern California. Utilización de datos espacialmente Explícitos para evaluar Áreas marinas protegidas para abalón en California meridional. Conservation Biology, 16(5):1308–1317, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Riley:2005:EUD

[RBK^{+}05] Seth P. D. Riley, Gary T. Busteed, Lee B. Kats, Thomas L. Vandergon, Lena F. S. Lee, Rosi G. Dagit, Jacob L. Kerby, Robert N. Fisher, and Raymond M. Sauvajot. Effects of urbanization on the distribution and abundance of amphibians and invasive species in Southern California streams. Efectos de la urbanización sobre la distribución y abundancia de anfibios y especies invasoras en arroyos del sur de California. Conservation Biology, 19(6):1894–1907, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regan:2005:CER

[RBM^{+}05] Tracey J. Regan, Mark A. Burgman, Michael A. Mccarthy, Lawrence L. Master, David A. Keith, Georgina M. Mace, and Sandy J. Andelman. The consistency of extinction risk classification protocols. La consistencia de los protocolos de clasificación del riesgo de extinción. Conservation Biology, 19(6):1969–1977, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rogers-Bennett:2001:IBM

[RBP01] Laura Rogers-Bennett and John S. Pearse. Indirect benefits of marine protected areas for juvenile abalone. Beneficios indirectos de Áreas marinas protegidas para juveniles de abulón. Conservation Biology, 15(3):642–647, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reeves:2006:PLR

[RBRB06] Gordon H. Reeves, Peter A. Bisson, Bruce E. Rieman, and Lee E. Benda. Postfire logging in riparian areas. Cosecha de madera post fuego en Áreas ribereñas. Conservation Biology, 20 (4):994–1004, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rahel:2008:MAS

[RBT08] Frank J. Rahel, Britta Bierwagen, and Yoshinori Taniguchi. Managing aquatic species of conservation concern in the face of climate change and invasive species. Manejo de especies acuáticas de interés para la conservación ante el cambio climático y las especies invasoras. Conservation Biology, 22(3):551–561, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regan:2000:FST

[RC00] Helen M. Regan and Mark Colyvan. Fuzzy sets and threatened species classification. Conservation Biology, 14(4):1197–1199, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Recher:2004:CBP

[RC04] Harry F. Recher and Michael Calver. Conservation biology in the Pacific. Conservation Biology, 18(6):1449–1451, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reid:2005:ICP

[RC05] John Reid and Wilson Cabral De Sousa, Jr. Infrastructure and conservation policy in Brazil. Infraestructura y políticas de conservación en Brasil. Conservation Biology, 19(3):740–746, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rouget:2006:DLS

[RCL^{+}06] Mathieu Rouget, Richard M. Cowling, Amanda T. Lombard, Andrew T. Knight, and Graham I. H. Kerley. Designing largescale conservation corridors for pattern and process. Diseño de corredores de conservación de gran escala para patrones y procesos. Conservation Biology, 20(2):549–561, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roman-Cuesta:2006:EPA

[RCMV06] Rosa María Román-Cuesta and Jordi Martínez-Vilalta. Effectiveness of protected areas in mitigating fire within their boundaries: Case study of Chiapas, Mexico. Efectividad de las Áreas protegidas en la mitigación de fuego dentro de sus límites: un estudio de caso en Chipas, México. Conservation Biology, 33(5): 1074–1086, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2000:UPM

[RCP00] Dina L. Roberts, Robert J. Cooper, and Lisa J. Petit. Use of premontane moist forest and shade coffee agroecosystems by army ants in Western Panama. Utilización de bosque Húmedo premontano y agroecosistemas cafetaleros de sombra por hormigas arrieras en Panama occidental. Conservation Biology, 14(1): 192–199, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2003:MCL

[RCS^{+}03] Kent H. Redford, Peter Coppolillo, Eric W. Sanderson, Gustavo A. B. Da Fonseca, Eric Dinerstein, Craig Groves, Georgina Mace, Stewart Maginnis, Russell A. Mittermeier, Reed Noss, David Olson, John G. Robinson, Amy Vedder, and Michael Wright. Mapping the conservation landscape. Clasificando los enfoques utilizadas en la conservación. Conservation Biology, 17 (1):116–131, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rohlf:2005:LEE

[RD05] Daniel J. Rohlf and David S. Dobkin. Legal ecology: Ecosystem function and the law. Conservation Biology, 19(5):1344–1348, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rudel:2009:CDD

[RDAL09] Thomas K. Rudel, Ruth Defries, Gregory P. Asner, and William F. Laurance. Changing drivers of deforestation and new opportunities for conservation. Cambio en los factores de deforestación y nuevas oportunidades de conservación. Conservation Biology, 23(6):1396–1405, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ricketts:2001:CBM

[RDEF01] Taylor H. Ricketts, Gretchen C. Daily, Paul R. Ehrlich, and John P. Fay. Countryside biogeography of moths in a fragmented landscape: Biodiversity in native and agricultural habitats. Biogeografía rural de polillas en un paisaje fragmentado: Biodiversidad en hábitats nativos y agrícolas. Conservation Biology, 15 (2):378–388, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Read:2006:BMM

[RDN06] Andrew J. Read, Phebe Drinker, and Simon Northridge. Bycatch of marine mammals in U.S. and global fisheries captura incidental de mamíferos en pesquerías de E.U.A. y globales. Conservation Biology, 20(1):163–169, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ries:2001:CVR

[RDW01] Leslie Ries, Diane M. Debinski, and Michelle L. Wieland. Conservation value of roadside prairie restoration to butterfly communities. Valor de la conservación de la restauración de praderas ruderales para comunidades de mariposas. Conservation Biology, 15(2):401–411, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosenbaum:2000:UNA

[REC^{+}00] Howard C. Rosenbaum, Mary G. Egan, Phillip J. Clapham, Robert L. Brownell, Jr., Sobia Malik, Moira W. Brown, Bradley N. White, Peter Walsh, and Rob Desalle. Utility of North Atlantic right whale museum specimens for assessing changes in genetic diversity. Utilidad de los especímenes de museo de la ballena franca septentrional Eubalaena glacialis del Atlántico norte para la estimación de cambios en la diversidad genética. Conservation Biology, 14(6):1837–1842, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2005:RBP

[Ree05] David H. Reed. Relationship between population size and fitness. Relación entre tamaño poblacional y adaptabilidad. Conservation Biology, 19(2):563–568, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reid:2005:CRT

[Rei05a] Donald G. Reid. Conservation on the right track. Conservation Biology, 19(3):981–982, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reid:2005:SWN

[Rei05b] Walter V. Reid. Science: Who needs it? Conservation Biology, 19(5):1341–1343, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Remes:2003:EEH

[Rem03] Vladimír Remeš. Effects of exotic habitat on nesting success, territory density, and settlement patterns in the blackcap (Sylvia atricapilla). Efectos de hábitat Exótico sobre el Éxito de anidación, densidad territorial y patrones de establecimiento de Sylvia atricapilla. Conservation Biology, 17(4):1127–1133, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regosin:2000:CBW

[RF00] Jonathan V. Regosin and Michelle Frankel. Conservation biology and western religious teachings. Conservation Biology, 14(1): 322–324, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2003:CBF

[RF03] David H. Reed and Richard Frankham. Correlation between fitness and genetic diversity. Correlación entre adaptabilidad y diversidad genética. Conservation Biology, 17(1):230–237, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regnier:2009:KRL

[RFB09] Claire Régnier, Benoît Fontaine, and Philippe Bouchet. Not knowing, not recording, not listing: Numerous unnoticed mollusk extinctions. No conocer, no registrar, no enlistar: Numerosas extinciones de moluscos no detectadas. Conservation Biology, 23(5):1214–1221, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:2001:EUD

[RFD01] Carlie A. Rodriguez, Karl W. Flessa, and David L. Dettman. Effects of upstream diversion of Colorado River water on the estuarine bivalve mollusc Mulinia coloradoensis. Efectos de la desviación de las aguas del Río Colorado en los moluscos estuarinos bivalvos Mulinia coloradoensis. Conservation Biology, 15 (1):249–258, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodrigues:2002:RCP

[RG02] Ana S. L. Rodrigues and Kevin J. Gaston. Rarity and conservation planning across geopolitical units. Rareza y planeación de la conservación a Través de unidades Geopolíticas. Conservation Biology, 16(3):674–682, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2004:MOS

[RG04a] S. C. Roberts and L. M. Gosling. Manipulation of olfactory signaling and mate choice for conservation breeding: a case study of harvest mice. Manipulación de la señalización olfativa y selección de pareja para la reproducción con fines de conservación: un estudio de caso de Micromys minutus. Conservation Biology, 18(2):548–556, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2004:PPP

[RG04b] John G. Robinson and Joshua R. Ginsberg. Parks, people, and pipelines. Conservation Biology, 18(3):607–608, June 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reeves:2006:RBC

[RG06] Randall R. Reeves and Nicholas J. Gales. Realities of baiji conservation. Conservation Biology, 20(3):626–628, June 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Renken:2004:EFM

[RGF^{+}04] Rochelle B. Renken, Wendy K. Gram, Debra K. Fantz, Stephen C. Richter, Timothy J. Miller, Kevin B. Ricke, Bradley Russell, and Xiaoyin Wang. Effects of forest management on amphibians and reptiles in Missouri Ozark forests. Efectos de la gestión de bosques sobre anfibios y reptiles en los bosques Ozark, Missouri. Conservation Biology, 18(1):174–188, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rolstad:2002:UIS

[RGGS02] Jørund Rolstad, Ivar Gjerde, Vegard S. Gundersen, and Magne Sætersdal. Use of indicator species to assess forest continuity: a critique. Conservation Biology, 16(1):253–257, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Romagosa:2009:CLV

[RGW09] Christina M. Romagosa, Craig Guyer, and Michael C. Wooten. Contribution of the live-vertebrate trade toward taxonomic homogenization. Contribución del comercio de vertebrados vivos a la homogenización Taxonómica. Conservation Biology, 23(4): 1001–1007, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robertson:2001:BBT

[RH01] David P. Robertson and R. Bruce Hull. Beyond biology: toward a more public ecology for conservation. Más allá de la biología: Hacia una ecología más Pública para la conservación. Conservation Biology, 15(4):970–979, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robertson:2008:EMW

[RH08] Morgan Robertson and Nicholas Hayden. Evaluation of a market in wetland credits: Entrepreneurial wetland banking in Chicago. Evaluación de un mercado de créditos de humedales en Chicago. Conservation Biology, 22(3):636–646, June 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Remis:2009:TSA

[RH09] Melissa J. Remis and Rebecca Hardin. Transvalued species in an African forest. Especies sobrestimadas en un bosque africano. Conservation Biology, 23(6):1588–1596, December 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rachowicz:2005:NEP

[RHA^{+}05] Lara J. Rachowicz, Jean-Marc Hero, Ross A. Alford, John W. Taylor, Jess A. T. Morgan, Vance T. Vredenburg, James P. Collins, and Cheryl J. Briggs. The novel and endemic pathogen hypotheses: Competing explanations for the origin of emerging infectious diseases of wildlife. La hipótesis del patógeno incipiente y Endémico: Explicaciones opuestas del origen de enfermedades infecciosas emergentes en la vida silvestre. Conservation Biology, 19(5):1441–1448, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Radeloff:2005:RSS

[RHS05] Volker C. Radeloff, Roger B. Hammer, and Susan I. Stewart. Rural and suburban sprawl in the U.S. Midwest from 1940 to 2000 and its relation to forest fragmentation. Expansión urbana y suburbana en el medio Oeste de E.U.A. de 1940 a 2000 y su relación con la fragmentación de bosques. Conservation Biology, 19(3):793–805, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rwego:2008:GBT

[RIBGG08] Innocent B. Rwego, Gilbert Isabirye-Basuta, Thomas R. Gillespie, and Tony L. Goldberg. Gastrointestinal bacterial transmission among humans, mountain gorillas, and livestock in Bwindi Impenetrable National Park, Uganda. Transmisión de bacterias gastrointestinales entre humanos, gorilas de montaña y ganado en el Parque Nacional Bwindi Impenetrable, Uganda. Conservation Biology, 22(6):1600–1607, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ricciardi:2004:BCB

[Ric04a] Anthony Ricciardi. Biological conveyor belts. Conservation Biology, 18(6):1699–1700, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ricketts:2004:TFF

[Ric04b] Taylor H. Ricketts. Tropical forest fragments enhance pollinator activity in nearby coffee crops. Fragmentos de bosque tropical incrementan la actividad de polinizadores en cultivos de café cercanos. Conservation Biology, 18(5):1262–1271, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ricciardi:2005:REC

[Ric05] Anthony Ricciardi. Rapid evolutionary change in homogocene. Conservation Biology, 19(5):1672–1673, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ricciardi:2007:MBI

[Ric07] Anthony Ricciardi. Are modern biological invasions an unprecedented form of global change? `?Las invasiones biológicas modernas son una forma de cambio global sin precedente? Conservation Biology, 21(2):329–336, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rigg:2001:OEM

[Rig01] Catherine M. Rigg. Orchestrating ecosystem management: Challenges and lessons from Sequoia National Forest. Instrumentación del manejo de ecosistemas: Desafíos y lecciones del Bosque Nacional Sequoia. Conservation Biology, 15(1):78–90, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Riley:2000:TDS

[Ril00] Stephen Riley. Technology and the demise of salmon. Conservation Biology, 14(3):914–916, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Risebrough:2002:CCR

[Ris02] Robert W. Risebrough. California Condor Recovery Program: Response to Beissinger. Conservation Biology, 16(4):1156–1157, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Bei01].

Rahmig:2009:GBR

[RJW09] Corina J. Rahmig, William E. Jensen, and Kimberly A. With. Grassland bird responses to land management in the largest remaining tallgrass prairie. Respuestas de aves de pastizal al manejo de tierra en el remanente más extenso de tallgrass prairie. Conservation Biology, 23(2):420–432, April 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rubbo:2005:ABD

[RK05] Michael J. Rubbo and Joseph M. Kiesecker. Amphibian breeding distribution in an urbanized landscape. Distribución de la reproducción de anfibios en un paisaje urbanizado. Conservation Biology, 19(2):504–511, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richardson:2006:SMR

[RKEJP06] Elizabeth A. Richardson, Michel J. Kaiser, Gareth Edwards-Jones, and Hugh P. Possingham. Sensitivity of marine-reserve design to the spatial resolution of socioeconomic data. Sensibilidad del diseño de reservas marinas a la resolución espacial de datos socioeconómico. Conservation Biology, 33(5):1191–1202, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rappole:2003:CC

[RKR03a] John H. Rappole, David I. King, and Jorge H. Vega Rivera. Coffee and conservation. Conservation Biology, 21(6):334–336, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rappole:2003:CCI

[RKR03b] John H. Rappole, David I. King, and Jorge H. Vega Rivera. Coffee and conservation III: Reply to Philpott and Dietsch. Conservation Biology, 17(6):1847–1849, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rand:2006:SAS

[RL06] Tatyana A. Rand and Svata M. Louda. Spillover of agriculturally subsidized predators as a potential threat to native insect herbivores in fragmented landscapes. El excedente de depredadores subsidiados por la agricultura como una amenaza potencial para los insecto herbívoros nativos en paisajes fragmentados. Conservation Biology, 20(6):1720–1729, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Randi:2000:MDV

[RLC^{+}00] Ettore Randi, Vittorio Lucchini, Mads Fjeldsø Christensen, Nadia Mucci, Stephan M. Funk, Gaudenz Dolf, and Volker Loeschcke. Mitochondrial DNA variability in Italian and East European wolves: Detecting the consequences of small population size and hybridization. Variabilidad en el ADN mitocondrial de lobos italianos y de Europa oriental: Detección de las consecuencias de un tamaño poblacional pequeño y de la Hibridización. Conservation Biology, 14(2):464–473, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rissman:2007:CEB

[RLC^{+}07] Adena R. Rissman, Lynn Lozier, Tosha Comendant, Peter Kareiva, Joseph M. Kiesecker, M. Rebecca Shaw, and Adina M. Merenlender. Conservation easements: Biodiversity protection and private use. Concesiones para conservación: Protección de biodiversidad y uso privado. Conservation Biology, 21(3):709– 718, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2001:CAA

[RLK^{+}01] Julie A. Robinson, Kamlesh P. Lulla, Minoru Kashiwagi, Maggie Suzuki, M. Duane Nellis, Charles E. Bussing, Warren J. Lee Long, and Len J. McKenzie. Conservation applications of astronaut photographs of Earth: Tidal-flat loss (Japan), elephant effects on vegetation (Botswana), and seagrass and mangrove monitoring (Australia). Usos en la conservación de fotografías tomadas por astronautas: Pérdida de zonas de mareas (Japón), efectos de elefantes en la vegetación (Botswana), y monitoreo de pastos marinos y manglares (Australia). Conservation Biology, 15(4):876–884, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rutledge:2001:SDE

[RLXL01] Daniel T. Rutledge, Christopher A. Lepczyk, Jialong Xie, and Jianguo Liu. Spatiotemporal dynamics of endangered species hotspots in the United States. Dinámica espaciotemporal de areas críticas para especies en peligro en los Estados Unidos. Conservation Biology, 15(2):475–487, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Restani:2001:ACU

[RM01] Marco Restani and John M. Marzluff. Avian conservation under the Endangered Species Act: Expenditures versus recovery priorities. Conservación de aves bajo el acta de especies en peligro: Prioridades entre gastos versus recuperación. Conservation Biology, 15(5):1292–1299, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosser:2002:OSE

[RM02] Alison M. Rosser and Sue A. Mainka. Overexploitation and species extinctions. Conservation Biology, 16(3):584–586, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reid:2003:TCB

[RM03] Walter V. Reid and Georgina M. Mace. Taking conservation biology to new levels in environmental decision-making. Conservation Biology, 17(4):943–945, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redding:2006:IEM

[RM06] David W. Redding and Arne Ø. Mooers. Incorporating evolutionary measures into conservation prioritization. Incorporación de Parámetros evolutivos a la definición de prioridades de conservación. Conservation Biology, 20(6):1670–1678, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2002:EIP

[RMD^{+}02] J. Michael Reed, L. Scott Mills, John B. Dunning, Jr., Eric S. Menges, Kevin S. McKelvey, Robert Frye, Steven R. Beissinger, Marie-Charlotte Anstett, and Philip Miller. Emerging issues in population viability analysis. Uso y temas emergentes del análisis de viabilidad poblacional. Conservation Biology, 16 (1):7–19, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rohr:2007:DMP

[RMK07] Jason R. Rohr, Carolyn G. Mahan, and Ke Chung Kim. Developing a monitoring program for invertebrates: Guidelines and a case study desarrollo de un programa de monitoreo de invertebrados: Directrices y un estudio de caso. Conservation Biology, 21(2):422–433, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reunanen:2000:MBF

[RMN00] Pasi Reunanen, Mikko Mönkkönen, and Ari Nikula. Managing boreal forest landscapes for flying squirrels manejo de paisajes en bosques boreales para ardillas voladoras. Conservation Biology, 14(1):218–226, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reist-Marti:2003:WAC

[RMSG^{+}03] S. B. Reist-Marti, H. Simianer, J. Gibson, O. Hanotte, and J. E. O. Rege. Weitzman’s approach and conservation of breed diversity: an application to African cattle breeds. El método de Weitzman y la conservación de la diversidad de razas: una aplicación a razas de ganado Africano. Conservation Biology, 17(5): 1299–1311, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rice:2004:L

[RN04] Richard E. Rice and Eduard T. Niesten. Letters: 2. Conservation Biology, 18(6):1452–1454, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosenfield:2004:RHV

[RNL^{+}04] Jonathan A. Rosenfield, Stacy Nolasco, Steven Lindauer, Claudette Sandoval, and Astrid Kodric-Brown. The role of hybrid vigor in the replacement of pecos pupfish by its hybrids with sheepshead minnow. El papel de vigor Híbrido en el reemplazo de Cyprinodon pecosensis por sus Híbridos con C. variegatus. Conservation Biology, 18(6):1589–1598, December 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ramstad:2007:SCC

[RNP^{+}07] Kristina M. Ramstad, N. J. Nelson, G. Paine, D. Beech, A. Paul, P. Paul, F. W. Allendorf, and C. H. Daugherty. Species and cultural conservation in New Zealand: Maori traditional ecological knowledge of tuatara. Conservación cultural y de especies en Nueva Zelanda: Conocimiento ecológico tradicional Maori de la tuatara. Conservation Biology, 21(2):455–464, April 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robertson:2004:PSF

[RO04] Kevin M. Robertson and Thomas E. Ostertag. Problems with Schurbon and Fauth’s test of effects of prescribed burning on amphibian diversity. Conservation Biology, 18(4):1154–1155, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SF03].

Rahel:2008:AEC

[RO08] Frank J. Rahel and Julian D. Olden. Assessing the effects of climate change on aquatic invasive species. Evaluación de los efectos del cambio climático sobre especies acuáticas invasoras. Conservation Biology, 22(3):521–533, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robbins:2000:CAM

[Rob00] Christopher S. Robbins. Comparative analysis of management regimes and medicinal plant trade monitoring mechanisms for American ginseng and goldenseal. Análisis comparativo de regímenes de manejo y monitoreos del mercado de plantas medicinales para el ginseng americano y la hidrastia. Conservation Biology, 14(5):1422–1434, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2004:LEM

[ROB^{+}04] David H. Reed, Julian J. O’Grady, Barry W. Brook, Jonathan D. Ballou, and Richard Frankham. Large estimates of minimum viable population sizes. Conservation Biology, 18(5):1179, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2006:DTT

[Rob06a] John Roberts. Diversity in three tenses. Conservation Biology, 20(5):1560–1561, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2006:CBR

[Rob06b] John G. Robinson. Conservation Biology and real-world conservation. Conservation Biology y la conservación en el mundo real. Conservation Biology, 20(3):658–669, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robison:2009:CDP

[Rob09] Bruce H. Robison. Conservation of deep pelagic biodiversity. Conservación de la biodiversidad pelágica profunda. Conservation Biology, 23(4):847–858, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:2003:PAB

[Rod03] Jon Paul Rodríguez. Un paso adelante para la biologí a de la conservación en Latinoamérica. (Spanish) [A step forward for conservation biology in Latin America]. Conservation Biology, 17 (6):1855–1857, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodrigues:2005:CBR

[Rod05] Miguel Trefaut Rodrigues. The conservation of Brazilian reptiles: Challenges for a megadiverse country. La conservación de reptiles brasileños: Retos para un país megadiverso. Conservation Biology, 19(3):659–664, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:2009:I

[Rod09] Jon Paul Rodríguez. Introduction. Conservation Biology, 23 (4):797–798, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rolston:2000:FSL

[Rol00] Holmes Rolston III. Four spikes, last chance. Conservation Biology, 14(2):584–585, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rolston:2006:DRH

[Rol06a] Holmes Rolston III. Disenchanting the rhetoric: Human uniqueness and human responsibility. Conservation Biology, 20(6): 1576–1578, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rolston:2006:FER

[Rol06b] Holmes Rolston III. Finite ethics: Revolution or threshold? Conservation Biology, 20(5):1558–1560, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Romero:2003:DTC

[Rom03] Aldemaro Romero. Death and taxes: the case of the depletion of pearl oyster beds in Sixteenth-Century Venezuela. Muerte e impuestos: El caso del agotamiento de los bancos de ostras perliferas en Venezuela en el Siglo XVI. Conservation Biology, 17 (4):1013–1023, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roman:2007:SG

[Rom07] Joe Roman. The selfless geneticist. Conservation Biology, 21(5): 1378–1379, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Romero:2009:KCC

[Rom09] Claudia Romero. Keeping up with climate-change policies. Conservation Biology, 23(5):1334–1336, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roosenburg:2000:HTD

[Roo00] Willem M. Roosenburg. Hypothesis testing, decision theory, and common sense in resource management. Conservation Biology, 14(4):1208–1210, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosenfeld:2002:LFA

[Ros02] Jordan S. Rosenfeld. Logical fallacies in the assessment of functional redundancy. Conservation Biology, 16(3):837–839, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosales:2006:EGB

[Ros06] Jon Rosales. Economic growth and biodiversity loss in an age of tradable permits. Crecimiento económico y pérdida de biodiversidad en tiempos de permisos comerciables. Conservation Biology, 33(5):1042–1050, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosales:2008:EGC

[Ros08] Jon Rosales. Economic growth, climate change, biodiversity loss: Distributive justice for the Global North and South. Crecimiento económico, cambio climático, pérdida de biodiversidad: Justicia distributiva para el Norte y el Sur. Conservation Biology, 22(6): 1409–1417, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rothstein:2001:LAC

[Rot01] Stephen I. Rothstein. A landscape approach to conserving North America’s avifauna. Conservation Biology, 15(6):1827–1828, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roubik:2000:PSS

[Rou00] David W. Roubik. Pollination system stability in tropical America. Conservation Biology, 14(5):1235–1236, October 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Royle:2004:MAI

[Roy04] J. Andrew Royle. Modeling abundance index data from anuran calling surveys. Modelaje de datos de Índices de abundancia a partir de muestreos de llamados de anuros. Conservation Biology, 18(5):1378–1385, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rondinini:2007:SSS

[RP07] Carlo Rondinini and Robert L. Pressey. Special Section: Systematic Conservation Planning in the European Landscape: Conflicts, Environmental Changes, and the Challenge of Countdown 2010. Conservation Biology, 21(6):1404–1405, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Revilla:2001:ECE

[RPD01] Eloy Revilla, Francisco Palomares, and Miguel Delibes. Edgecore effects and the effectiveness of traditional reserves in conservation: Eurasian badgers in Doñana National Park. Efectos de borde–núcleo y la efectividad de las reservas tradicionales en la conservación: el Tejón Eurasiático en el Parque Nacional Doñana. Conservation Biology, 15(1):148–158, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ramsey:2009:QES

[RPM09] David S. L. Ramsey, John Parkes, and Scott A. Morrison. Quantifying eradication success: the removal of feral pigs from Santa Cruz Island, California. Cuantificación del Éxito de erradicación: La remoción de cerdos cimarrones de la Isla Santa Cruz, California. Conservation Biology, 23(2):449–459, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rivard:2000:CSR

[RPP^{+}00] Donald H. Rivard, Jean Poitevin, Daniel Plasse, Michel Carleton, and David J. Currie. Changing species richness and composition in Canadian National Parks. Cambios en la riqueza y composición de especies en Parques Nacionales canadienses. Conservation Biology, 14(4):1099–1109, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reaser:2000:CBG

[RPT00] Jamie K. Reaser, Rafe Pomerance, and Peter O. Thomas. Coral bleaching and global climate change: Scientific findings and policy recommendations. Blanqueamiento de coral y cambio climático global: Información científica y recomendaciones políticas. Conservation Biology, 14(5):1500–1511, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richter:2000:PFR

[RR00] Brian D. Richter and Holly E. Richter. Prescribing flood regimes to sustain riparian ecosystems along meandering rivers. Prescripción de regímenes de inundación para mantener ecosistemas riparios a lo largo de ríos sinuosos. Conservation Biology, 14(5): 1467–1478, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2001:ANS

[RR01] Scott K. Robinson and W. Douglas Robinson. Avian nesting success in a selectively harvested north temperate deciduous forest. Éxito de nidación de aves en un bosque deciduo templado selectivamente cosechado. Conservation Biology, 15(6):1763–1771, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2006:PS

[RRA06] Kent H. Redford, John G. Robinson, and William M. Adams. Parks as shibboleths. Conservation Biology, 20(1):1–2, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:2006:PCB

[RRCOM^{+}06] Jon Paul Rodríguez, Kathryn M. Rodríguez-Clark, María A. Oliveira-Miranda, Tatjana Good, and Alejandro Grajal. Professional capacity building: the missing agenda in conservation priority setting. Conservation Biology, 20(5):1340, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rogers:2009:PED

[RRH^{+}09] David A. Rogers, Thomas P. Rooney, Todd J. Hawbaker, Volker C. Radeloff, and Donald M. Waller. Paying the extinction debt in Southern Wisconsin forest understories. Pago de la deuda de extinción en sotobosques de bosques del sur de Wisconsin. Conservation Biology, 23(6):1497–1506, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rao:2002:SRP

[RRK02] Madhu Rao, Alan Rabinowitz, and Saw Tun Khaing. Status review of the protected-area system in Myanmar, with recommendations for conservation planning. Revisión del estatus del sistema de Áreas protegidas en mynamar con recomendaciones para la planificación de la conservación. Conservation Biology, 16(2):360–368, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2000:EHN

[RS00] Kent H. Redford and Steven E. Sanderson. Extracting humans from nature. Conservation Biology, 14(5):1362–1364, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roldan:2001:PMI

[RS01] Alejandra I. Roldán and Javier A. Simonetti. Plant-mammal interactions in tropical Bolivian forests with different hunting pressures. Interacciones planta-animal en bosques tropicales bolivianos bajo distintas presiones de cacería. Conservation Biology, 15(3):617–623, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2002:LWE

[RS02a] Robert N. Reed and Richard Shine. Lying in wait for extinction: Ecological correlates of conservation status among Australian elapid snakes. Esperando la extinción: Correlaciones ecológicas de estatus de conservación entre serpientes Elápidas Australianas. Conservation Biology, 16(2):451–461, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodgers:2002:BZD

[RS02b] James A. Rodgers, Jr. and Stephen T. Schwikert. Buffer-zone distances to protect foraging and loafing waterbirds from disturbance by personal watercraft and outboard-powered boats. Distancias de amortiguamiento para proteger a aves acuáticas en actividades de forraje y descanso de la perturbación por embarcaciones personales y motores fuera de borda. Conservation Biology, 16(1):216–224, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rothermel:2002:EIL

[RS02c] Betsie B. Rothermel and Raymond D. Semlitsch. An experimental investigation of landscape resistance of forest versus oldfield habitats to emigrating juvenile amphibians. Una investigación experimental de la resistencia de paisaje de bosque versus hábitats de campos viejos sobre anfibios juveniles emigrantes. Conservation Biology, 16(5):1324–1332, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2003:RC

[RS03] Kent Redford and M. A. Sanjayan. Retiring Cassandra. Conservation Biology, 17(6):1473–1474, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Root:2006:CCC

[RS06] Terry L. Root and Stephen H. Schneider. Conservation and climate change: the challenges ahead. Conservation Biology, 20 (3):706–708, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rondinini:2005:HSM

[RSB05] Carlo Rondinini, Simon Stuart, and Luigi Boitani. Habitat suitability models and the shortfall in conservation planning for African vertebrates. Modelos de aptitud de hábitat y la deficiente planificación de la conservación de vertebrados africanos. Conservation Biology, 19(5):1488–1497, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robert:2003:VAI

[RSC03] Alexandre Robert, François Sarrazin, and Denis Couvet. Variation among individuals, demographic stochasticity, and extinction: Response to Kendall and Fox. Conservation Biology, 17 (4):1166–1169, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robert:2004:RAV

[RSCL04] Alexandre Robert, François Sarrazin, Denis Couvet, and Stéphane Legendre. Releasing adults versus young in reintroductions: Interactions between demography and genetics. Liberación de adultos versus juveniles en reintroducciones: Interacciones entre demografía y genética. Conservation Biology, 18 (4):1078–1087, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ribon:2003:BEA

[RSD03] Rômulo Ribon, José Eduardo Simon, and Geraldo Theodoro De Mattos. Bird extinctions in Atlantic forest fragments of the Viçosa Region, Southeastern Brazil. Extinciones de aves en fragmentos de bosque Atlántico de la región de Viçosa, Sureste de Brasil. Conservation Biology, 17(6):1827–1839, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Riley:2003:EUH

[RSF^{+}03] Seth P. D. Riley, Raymond M. Sauvajot, Todd K. Fuller, Eric C. York, Denise A. Kamradt, Cassity Bromley, and Robert K. Wayne. Effects of urbanization and habitat fragmentation on bobcats and coyotes in Southern California. Efectos de la urbanización y la fragmentación del hábitat sobre gatos silvestres (Lynx rufus) y coyotes en el sur de California. Conservation Biology, 17(2):566–576, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ringsby:2006:DCE

[RSJE06] Thor Harald Ringsby, Bernt-Erik SæTher, Henrik Jensen, and Steinar Engen. Demographic characteristics of extinction in a small, insular population of house sparrows in Northern Norway. Características demográficas de la extinción en una pequeña población insular de gorriones domésticos en el norte de Noruega. Conservation Biology, 20(6):1761–1767, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ringsby:2007:E

[RSJE07] Thor Harald Ringsby, Bernt-Erik SæTher, Henrik Jensen, and Steinar Engen. Errata. Conservation Biology, 20(6):566, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:2005:CAN

[RSPM05] Jon Paul Rodríguez, Javier A. Simonetti, Andrea Premoli, and Miguel Ângelo Marini. Conservation in Austral and neotropical America: Building scientific capacity equal to the challenges. Conservation Biology, 19(3):969–972, June 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2005:BTB

[RSW05] Catherine J. Robinson, Dermot Smyth, and Peter J. Whitehead. Bush tucker, bush pets, and bush threats: Cooperative management of feral animals in Australia’s Kakadu National Park.

Alimento, mascotas y amenazas en el matorral: Manejo cooperativo de animales cimarrones en el Parque Nacional Kakadu en Australia. Conservation Biology, 19(5):1385–1391, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ralls:2000:BPM

[RT00] Katherine Ralls and Barbara L. Taylor. Better policy and management decisions through explicit analysis of uncertainty: New approaches from marine conservation. Conservation Biology, 14 (5):1240–1242, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosenzweig:2003:EDU

[RTCR03] Michael L. Rosenzweig, Will R. Turner, Jonathan G. Cox, and Taylor H. Ricketts. Estimating diversity in unsampled habitats of a biogeographical province. Estimación de la diversidad en hábitats no muestreados de una provincia biogeográfica. Conservation Biology, 17(3):864–874, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rao:2001:IHF

[RTN01] Madhu Rao, John Terborgh, and Percy Nuñez. Increased herbivory in forest isolates: Implications for plant community structure and composition. Incremento de la herbivoría en fragmentos aislados de bosque: Implicaciones en la estructura y composición de la comunidad de plantas. Conservation Biology, 15(3):624– 633, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruane:2000:FPD

[Rua00] John Ruane. A framework for prioritizing domestic animal breeds for conservation purposes at the national level: a Norwegian case study. Marco de trabajo para priorizar razas de animales domésticos con Propósitos de conservación a nivel nacional: Un caso de estudio noruego. Conservation Biology, 14(5): 1385–1393, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rubinoff:2001:ECG

[Rub01] Daniel Rubinoff. Evaluating the California gnatcatcher as an umbrella species for conservation of Southern California coastal sage scrub. Evaluando a la perlita de California como una especie sombrilla para la conservación del matorral de salvia costero del sur de California. Conservation Biology, 15(5):1374–1383, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rubinoff:2006:DBE

[Rub06a] Daniel Rubinoff. DNA barcoding evolves into the familiar. Conservation Biology, 20(5):1548–1549, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rubinoff:2006:UMD

[Rub06b] Daniel Rubinoff. Utility of mitochondrial DNA barcodes in species conservation. Utilidad de los códigos de barras de AND mitocondrial en la conservación de especies. Conservation Biology, 33(5):1026–1033, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rudolph:2003:MFM

[Rud03] Victor J. Rudolph. Modifying forest management for biodiversity. Conservation Biology, 17(5):1463–1464, October 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rudel:2006:STF

[Rud06] Thomas K. Rudel. Shrinking tropical forests, human agents of change, and conservation policy. Contracción de los bosques tropicales, agentes de cambio humanos y políticas de conservación. Conservation Biology, 20(6):1604–1609, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruesink:2005:GAF

[Rue05] Jennifer L. Ruesink. Global analysis of factors affecting the outcome of freshwater fish introductions. Análisis global de los factores que afectan las repercusiones de las introducciones peces dulceacuícolas. Conservation Biology, 19(6):1883–1893, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruttenberg:2001:EAF

[Rut01] Benjamin I. Ruttenberg. Effects of artisanal fishing on marine communities in the Galápagos Islands. Efectos de la pesca artesanal en comunidades marinas en las Islas Galápagos. Conservation Biology, 15(6):1691–1699, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rull:2008:BOC

[RVNM08] Valentí Rull, Teresa Vegas, Sandra Nogué, and Encarni Montoya. Bureaucratic obstruction of conservation science in the Guayana Highlands. Conservation Biology, 22(3):508–509, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rull:2009:CUN

[RVVNH09] Valentí Rull, Teresa Vegas-Vilarrúbia, Sandra Nogué, and Otto Huber. Conservation of the unique neotropical vascular flora of the Guayana Highlands in the face of global warming. Conservation Biology, 23(5):1323–1327, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Read:2000:SMM

[RW00] Andrew J. Read and Paul R. Wade. Status of marine mammals in the United States. Estado de los mamíferos marinos en los Estados Unidos. Conservation Biology, 14(4):929–940, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reinert:2002:SHP

[RW02] Thomas R. Reinert and Kimberly A. Winter. Sustainability of harvested pacú (Colossoma macropomum) populations in the Northeastern Bolivian Amazon. Sustentabilidad de poblaciones cosechadas de pacú (Colossoma macropomum) en el noreste del Amazonas boliviano. Conservation Biology, 16(5):1344–1351, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roemer:2003:CCT

[RW03] Gary W. Roemer and Robert K. Wayne. Conservation in conflict: the tale of two endangered species. Conservación en conflicto: el cuento de dos especies en peligro. Conservation Biology, 17(5):1251–1260, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reeves:2006:ACS

[RWBG06] Gordon H. Reeves, Jack E. Williams, Kelly M. Burnett, and Kirsten Gallo. The aquatic conservation strategy of the northwest forest plan. La estrategia de conservación acuática del plan forestal del noroeste. Conservation Biology, 20(2):319–329, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rhodes:2006:ANA

[RWJRR06] Monika Rhodes, Grant W. Wardell-Johnson, Martin P. Rhodes, and Ben Raymond. Applying network analysis to the conservation of habitat trees in urban environments: a case study from Brisbane, Australia. Aplicación del análisis de redes en la conservación de Árboles en ambientes urbanos: un estudio de caso de Brisbane, Australia. Conservation Biology, 21(3):861–870, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rhodes:2006:MSD

[RWM^{+}06] Jonathan R. Rhodes, Thorsten Wiegand, Clive A. Mcalpine, John Callaghan, Daniel Lunney, Michiala Bowen, and Hugh P. Possingham. Modeling species’ distributions to improve conservation in semiurban landscapes: Koala case study. Modelaje de las distribuciones de especies para mejorar la conservación en paisaje semiurbanos: Caso de estudio koala. Conservation Biology, 20(2):449–459, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rooney:2004:BIH

[RWRW04] Thomas P. Rooney, Shannon M. Wiegmann, David A. Rogers, and D. M. Waller. Biotic impoverishment and homogenization in unfragmented forest understory communities. Empobrecimiento biótico y homogenización en comunidades de sotobosque en bosques no fragmentados. Conservation Biology, 18(3):787– 798, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richards:2002:ALS

[RWS02] William H. Richards, David O. Wallin, and Nathan H. Schumaker. An analysis of late-seral forest connectivity in Western Oregon, U.S.A. Análisis de la conectividad de un bosque seral tardío en el oeste de Oregon, E.U.A. Conservation Biology, 16 (5):1409–1421, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryan:2006:RCC

[Rya06] S. J. Ryan. The role of culture in conservation planning for small or endangered populations. Conservation Biology, 33(5): 1321–1324, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stockwell:2004:RAC

[SA04] Craig A. Stockwell and Mary V. Ashley. Rapid adaptation and conservation. Conservation Biology, 18(1):272–273, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saatchi:2001:EFL

[SAA^{+}01] S. Saatchi, D. Agosti, K. Alger, J. Delabie, and J. Musinsky. Examining fragmentation and loss of primary forest in the Southern Bahian Atlantic forest of Brazil with radar imagery. Exámen de la fragmentación y de la pérdida de bosque primario en el bosque del Atlántico de Bahía del sur en Brasil utilizando imágenes de radar. Conservation Biology, 15(4):867–875, August 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sutherland:2009:OHQ

[SAA^{+}09] W. J. Sutherland, W. M. Adams, R. B. Aronson, R. Aveling, T. M. Blackburn, S. Broad, G. Ceballos, I. M. Côté, R. M. Cowling, G. A. B. Da Fonseca, E. Dinerstein, P. J. Ferraro, E. Fleishman, C. Gascon, M. Hunter, Jr., J. Hutton, P. Kareiva, A. Kuria, D. W. Macdonald, K. Mackinnon, F. J. Madgwick, M. B. Mascia, J. Mcneely, E. J. Milner-Gulland, S. Moon, C. G. Morley, S. Nelson, D. Osborn, M. Pai, E. C. M. Parsons, L. S. Peck, H. Possingham, S. V. Prior, A. S. Pullin, M. R. W. Rands, J. Ranganathan, K. H. Redford, J. P. Rodriguez, F. Seymour, J. Sobel, N. S. Sodhi, A. Stott, K. Vance-Borland, and A. R. Watkinson. One hundred questions of importance to the conservation of global biological diversity. Cien preguntas de importancia para la conservación de la diversidad biológica global. Conservation Biology, 23(3):557–567, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sadler:2005:RBB

[Sad05] Rohan J. Sadler. Rule breaking by building a frame for restoration. Conservation Biology, 19(3):977–978, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sagoff:2007:CCE

[Sag07] Mark Sagoff. On the compatibility of a conservation ethic with biological science. Sobre la compatibilidad de una Ética de conservación con la ciencia biológica. Conservation Biology, 21(2): 337–345, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanchez-Azofeifa:2005:NIR

[SAKQ^{+}05] G. Arturo Sánchez-Azofeifa, Margaret Kalacska, Mauricio Quesada, Julio C. Calvo-Alvarado, Jafet M. Nassar, and Jon Paul Rodríguez. Need for integrated research for a sustainable future in tropical dry forests. Conservation Biology, 19(2):285–286, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salzman:2002:WAE

[Sal02] James Salzman. A welcome addition to the ESA literature. Conservation Biology, 16(1):278–280, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salzman:2005:MRE

[Sal05] James Salzman. A “Must read” on ecosystem services. Conservation Biology, 19(2):582–583, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sale:2007:PPT

[Sal07] Peter F. Sale. Parasites and passerines tell different tales. Conservation Biology, 21(1):281–282, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Samways:2006:IEI

[Sam06] Michael J. Samways. Insect extinctions and insect survival. Conservation Biology, 20(1):245–246, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seddon:2007:DSR

[SAM07] Philip J. Seddon, Doug P. Armstrong, and Richard F. Maloney. Developing the science of reintroduction biology. Desarrollando la ciencia de biología de la reintroducción. Conservation Biology, 21(2):303–312, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sangster:2000:TSA

[San00] George Sangster. Taxonomic stability and avian extinctions. Conservation Biology, 14(2):579–581, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanford:2006:EAV

[San06] Monte P. Sanford. Extirpating the agriculture versus conservation dichotomy. Conservation Biology, 20(1):253–254, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanchez-Azofeifa:2007:CRP

[SAPRB07] G. Arturo Sánchez-Azofeifa, Alexander Pfaff, Juan Andres Robalino, and Judson P. Boomhower. Costa Rica’s payment for environmental services program: Intention, implementation, and impact. El programa de pago de servicios ambientales de Costa Rica: Intención, implementación e impacto. Conservation Biology, 21(5):1165–1173, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarukhan:2006:CBV

[Sar06] José Sarukhán. Conservation biology: Views from the ecological sciences. Conservation Biology, 20(3):674–676, June 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarkar:2007:PPP

[Sar07] Sahotra Sarkar. Pragmatic pluralism in protecting nature. Conservation Biology, 22(1):1660, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seddon:2009:RAC

[SAS^{+}09] Philip J. Seddon, Doug P. Armstrong, Pritpal Soorae, Frederic Launay, Sally Walker, Carlos R. Ruiz-Miranda, Sanjay Molur, Heather Koldewey, and Devra G. Kleiman. The risks of assisted colonization. Conservation Biology, 23(4):788–789, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sayre:2003:EGD

[Say03] Nathan F. Sayre. The end of the grazing debate. Conservation Biology, 17(4):1186–1188, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sayre:2005:IEL

[Say05] Nathan F. Sayre. Interacting effects of landownership, land use, and endangered species on conservation of Southwestern U.S. rangelands. Interacción de efectos de la propiedad de tierras, uso del suelo y especies en peligro sobre la conservación de llanuras del suroeste. Conservation Biology, 19(3):783–792, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sayre:2009:BAR

[Say09] Nathan F. Sayre. Bad abstractions: Response to Sullivan. Conservation Biology, 23(4):1050–1052, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Semlitsch:2003:BCB

[SB03] Raymond D. Semlitsch and J. Russell Bodie. Biological criteria for buffer zones around wetlands and riparian habitats for amphibians and reptiles. Criterios biológicos para zonas de amortiguamiento alrededor de hábitats de humedales y riparios para anfibios y reptiles. Conservation Biology, 17(5):1219–1228, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silliman:2004:SDD

[SB04] Brian R. Silliman and Mark D. Bertness. Shoreline development drives invasion of Phragmites australis and the loss of plant diversity on New England salt marshes. El desarrollo costero impulsa la invasión de Phragmites australis y la pérdida de diversidad de plantas en marismas de Nueva Inglaterra. Conservation Biology, 18(5):1424–1434, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skewes:2006:BCC

[SB06] Elizabeth Skewes and P. Dee Boersma. Boot camp for coastalmarine conservation. Conservation Biology, 20(1):257–259, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stevens:2008:IHQ

[SB08] Virginie M. Stevens and Michel Baguette. Importance of habitat quality and landscape connectivity for the persistence of endangered natterjack toads. Importancia de la calidad del hábitat y de la conectividad del paisaje para la persistencia de Bufo calamita en peligro. Conservation Biology, 22(5):1194–1204, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Slik:2008:WDC

[SBB^{+}08] J. W. Ferry Slik, Caroline S. Bernard, Floris C. Breman, Marloes Van Beek, Agus Salim, and Douglas Sheil. Wood density as a conservation tool: Quantification of disturbance and identification of conservation-priority areas in tropical forests. Densidad de madera como una herramienta de conservación: Cuantificación de la perturbación e identificación de las Áreas prioritarias para la conservación en bosques tropicales. Conservation Biology, 22(5):1299–1308, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seoane:2005:EEO

[SBDD05] Javier Seoane, Javier Bustamante, and Ricardo Díaz-Delgado. Effect of expert opinion on the predictive ability of environmental models of bird distribution. Efecto de la opinión de experto en la capacidad predictiva de modelos de distribución de aves usando predictores ambientales. Conservation Biology, 19(2):512– 522, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skelly:2000:R

[SBH00] David K. Skelly, Karen H. Beard, and Nicolas J. Hengartner. Reply [to Crist]. Conservation Biology, 14(5):1224–1225, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Cri00].

Sodhi:2006:BHL

[SBK^{+}06] Navjot S. Sodhi, Thomas M. Brooks, Lian Pin Koh, Greg Acciaioli, Maribeth Erb, Alan Khee-Jin Tan, Lisa M. Curran, Peter Brosius, Tien Ming Lee, Jason M. Patlis, Melvin Gumal, and Robert J. Lee. Biodiversity and human livelihood crises in the Malay Archipelago. Conservation Biology, 21(2):1811–1813, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saunders:2006:UPS

[SBM06] Carol D. Saunders, Amara T. Brook, and Olin Eugene Myers, Jr. Using psychology to save biodiversity and human wellbeing. Conservation Biology, 20(3):702–705, June 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stewart:2007:EIR

[SBP07] Romola R. Stewart, Ian R. Ball, and Hugh P. Possingham. The effect of incremental reserve design and changing reservation goals on the long-term efficiency of reserve systems. El efecto del diseño incremental de reservas y del cambio de metas de la reserva sobre la eficiencia a largo plazo de los sistemas de reservas. Conservation Biology, 21(2):346–354, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:2009:PPC

[SBS^{+}09] Nick Salafsky, Stuart H. M. Butchart, Daniel Salzer, Alison J. Stattersfield, Rachel Neugarten, Craig Hilton-Taylor, Ben Collen, Lawrence L. Master, Sheila O’Connor, and David Wilkie. Pragmatism and practice in classifying threats: Reply to Balmford et al. Conservation Biology, 23(2):488–493, April 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stouffer:2006:LTL

[SBSL06] Philip C. Stouffer, Richard O. Bierregaard, Jr., Cheryl Strong, and Thomas E. Lovejoy. Long-term landscape change and bird abundance in Amazonian rainforest fragments cambio a largo plazo en el paisaje y la abundancia de aves en fragmentos de bosque lluvioso en la Amazonia. Conservation Biology, 33(5): 1212–1223, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Summerville:2003:SVS

[SBVC03] Keith S. Summerville, Michael J. Boulware, Joseph A. Veech, and Thomas O. Crist. Spatial variation in species diversity and composition of forest Lepidoptera in eastern deciduous forests of North America. Variación espacial de la diversidad y composición de especies de Lepidópteros en bosques deciduos orientales de Norte América. Conservation Biology, 17(4):1045–1057, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Santiago:2000:ART

[SC00] Enrique Santiago and Armando Caballero. Application of reproduction technologies to the conservation of genetic resources. Aplicación de Tecnologías de reproducción para la conservación de recursos genéticos. Conservation Biology, 14(6):1831–1836, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schultz:2005:PSC

[SC05] Cheryl B. Schultz and Elizabeth E. Crone. Patch size and connectivity thresholds for butterfly habitat restoration. Tamaño del fragmento y umbrales de conectividad para la restauración del hábitat de mariposas. Conservation Biology, 19(3):887–896, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sagarin:2009:BTC

[SC09] Raphael D. Sagarin and Larry B. Crowder. Breaking through the crisis in marine conservation and management: Insights from the philosophies of Ed Ricketts. Progresando en la crisis de la conservación y el manejo marino: Perspectivas desde las Filosofías de Ed Ricketts. Conservation Biology, 23(1):24–30, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:2001:STE

[SCB^{+}01] N. Salafsky, H. Cauley, G. Balachander, B. Cordes, J. Parks, C. Margoluis, S. Bhatt, C. Encarnacion, D. Russell, and R. Margoluis. A systematic test of an enterprise strategy for community-based biodiversity conservation. Prueba sistemática de estrategias empresariales para la conservación de la biodiversidad basada en la comunidad. Conservation Biology, 15(6): 1585–1595, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stork:2009:VRT

[SCC^{+}09] Nigel E. Stork, Jonathan A. Coddington, Robert K. Colwell, Robin L. Chazdon, Christopher W. Dick, Carlos A. Peres, Sean Sloan, and Kathy Willis. Vulnerability and resilience of tropical forest species to land-use change vulnerabilidad y resiliencia de especies de bosques tropicales al cambio de uso de suelo. Conservation Biology, 23(6):1438–1447, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

SadovyDeMitcheson:2008:GBS

[SCD^{+}08] Yvonne Sadovy De Mitcheson, Andrew Cornish, Michael Domeier, Patrick L. Colin, Martin Russell, and Kenyon C. Lindeman. A global baseline for spawning aggregations of reef fishes una referencia global para agregaciones de desove de peces de arrecifes. Conservation Biology, 22(5):1233–1244, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schultz:2000:UIA

[Sch00] Cheryl B. Schultz. Using invertebrates to assess sustainable farming. Conservation Biology, 14(6):1929–1930, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schierenbeck:2002:FSB

[Sch02] Kristina A. Schierenbeck. Four stars for a book on fragmented populations. Conservation Biology, 16(1):274–275, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schaefer:2003:LTR

[Sch03a] James A. Schaefer. Long-term range recession and the persistence of caribou in the taiga. Recesión de rangos a largo plazo y la persistencia de caribú en la taiga. Conservation Biology, 17 (5):1435–1439, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:2003:NPP

[Sch03b] Kenneth A. Schmidt. Nest predation and population declines in Illinois songbirds: a case for mesopredator effects. Depredación de nidos y declinación poblacional de aves canoras en Illinois: Validación del efecto de los meso-depredadores. Conservation Biology, 17(4):1141–1150, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidtz:2005:URA

[Sch05] David Schmidtz. Using, respecting, and appreciating nature? Conservation Biology, 19(5):1673–1674, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwartz:2006:HCS

[Sch06] Mark W. Schwartz. How conservation scientists can help develop social capital for biodiversity. Conservation Biology, 20(5):1550– 1552, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schoen:2007:DSD

[Sch07] Daniel J. Schoen. Demon seed, demon weed. Conservation Biology, 21(1):283, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:2010:EID

[Sch10] Benedikt R. Schmidt. Estimating the impact of disease in species threatened by amphibian chytrid fungus: Comment on Murray et al. Conservation Biology, 25(6):897–899, June 2010. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [MSSM09].

Saterson:2004:DER

[SCJ^{+}04] Kathryn A. Saterson, Norman L. Christensen, Robert B. Jackson, Randall A. Kramer, Stuart L. Pimm, Martin D. Smith, and Jonathan B. Wiener. Disconnects in evaluating the relative effectiveness of conservation strategies. Conservation Biology, 18 (3):597–599, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stoner:2007:AEP

[SCM^{+}07] Chantal Stoner, Tim Caro, Simon Mduma, Charles Mlingwa, George Sabuni, and Markus Borner. Assessment of effectiveness of protection strategies in Tanzania based on a decade of survey data for large herbivores. Evaluación de efectividad de estrategias de conservación en Tanzania con base en una Década de datos de muestreos de herbívoros mayores. Conservation Biology, 21(3):635–646, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shivji:2002:GIP

[SCP^{+}02] Mahmood Shivji, Shelley Clarke, Melissa Pank, Lisa Natanson, Nancy Kohler, and Michael Stanhope. Genetic identification of pelagic shark body parts for conservation and trade monitoring. Identificación genética de las partes del cuerpo de un Tiburón Pelágico para la conservación y monitoreo de su comercialización. Conservation Biology, 16(4):1036–1047, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steinmetz:2006:CCL

[SCS06] Robert Steinmetz, Wanlop Chutipong, and Naret Seuaturien. Collaborating to conserve large mammals in Southeast Asia. Colaboración para la conservación de mamíferos mayores en el Sureste de Asia. Conservation Biology, 20(5):1391–1401, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strittholt:2001:IRA

[SD01] James R. Strittholt and Dominick A. Dellasala. Importance of roadless areas in biodiversity conservation in forested ecosystems: Case study of the Klamath–Siskiyou ecoregion of the United States. Importancia de Áreas sin caminos para la conservación de la biodiversidad en ecosistemas forestales: Estudio de caso de la ecoregión Klamath–Siskiyou, E.U.A. Conservation Biology, 15(6):1742–1754, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seigel:2002:TAP

[SD02] Richard A. Seigel and C. Kenneth Dodd, Jr. Translocations of amphibians: Proven management method or experimental technique? Conservation Biology, 16(2):552–554, April 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steffan-Dewenter:2003:IHA

[SD03] Ingolf Steffan-Dewenter. Importance of habitat area and landscape context for species richness of bees and wasps in fragmented orchard meadows. La importancia del area del hábitat y del contexto del paisaje para la riqueza de especies de abejas y avispas en praderas de huertos fragmentadas. Conservation Biology, 17(4):1036–1044, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strittholt:2006:SMO

[SDJ06] James R. Strittholt, Dominick A. Dellasala, and Hong Jiang. Status of mature and old-growth forests in the Pacific Northwest. Estatus de los bosques maduros y viejos en el Pacífico Noroeste. Conservation Biology, 20(2):363–374, April 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Su:2004:BSR

[SDJK04] Jeffrey C. Su, Diane M. Debinski, Mark E. Jakubauskas, and Kelly Kindscher. Beyond species richness: Community similarity as a measure of cross-taxon congruence for coarse-filter conservation. Más allá de la riqueza de especies: Similitud de comunidades como una medida de congruencia trans-taxones para la conservación de grano grueso. Conservation Biology, 18(1): 167–173, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snelder:2007:PMO

[SDL07] Ton H. Snelder, Katie L. Dey, and John R. Leathwick. A procedure for making optimal selection of input variables for multivariate environmental classifications. Un procedimiento para la selección Óptima de variables para clasificaciones ambientales multivariadas. Conservation Biology, 21(2):365–375, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shelden:2001:DCC

[SDRO01] Kim E. W. Shelden, Douglas P. Demaster, David J. Rugh, and Annette M. Olson. Developing classification criteria under the U.S. Endangered Species Act: Bowhead whales as a case study. Desarrollo de criterios de clasificación bajo el Acta de Especies En Peligro de EE.UU.: La ballena franca Balaena mysticetus como caso de estudio. Conservation Biology, 15(5):1300–1307, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [Tay03b].

Soule:2003:EEC

[SEBM03] Michael E. Soulé, James A. Estes, Joel Berger, and Carlos Martinez Del Rio. Ecological effectiveness: Conservation goals for interactive species. Efectividad ecológica: Metas de conservación para especies interactivas. Conservation Biology, 17(5):1238– 1250, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seidler:2004:RLT

[Sei04] Reinmar Seidler. Roads and the land: Two giants in uneasy embrace. Conservation Biology, 18(3):850–852, June 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seidensticker:2007:FS

[Sei07] John Seidensticker. A fearful symmetry. Conservation Biology, 21(2):559–560, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Semlitsch:2002:CEB

[Sem02] Raymond D. Semlitsch. Critical elements for biologically based recovery plans of aquatic-breeding amphibians. Elementos Críticos para planes de recuperación con bases biológicas de anfibios acuáticos. Conservation Biology, 16(3):619–629, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seymour:2004:PSS

[Sey04] Mike Seymour. Partnerships to support sustainable development and conservation: the West-East pipeline project, China. Conservation Biology, 18(3):613–615, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:2000:SOF

[SF00] Michael E. Soulé and Richard Frankham. Sir Otto Frankel: Memories and tributes. Conservation Biology, 14(2):582–583, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schurbon:2003:EPB

[SF03] Jamie M. Schurbon and John E. Fauth. Effects of prescribed burning on amphibian diversity in a Southeastern U.S. National Forest. Efectos de quemas prescritas sobre la diversidad de anfibios en un Parque Nacional en el Sureste de E.U.A. Conservation Biology, 17(5):1338–1349, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [RO04], response [MDJP04], and reply citeSchurbon:2004:FFF.

Schurbon:2004:FFF

[SF04] Jamie M. Schurbon and John E. Fauth. Fire as friend and foe of amphibians: a reply. Conservation Biology, 18(4):1156–1159, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SF03, MDJP04].

Solymos:2005:CPB

[SF05] Péter Sólymos and Zoltán Fehér. Conservation prioritization based on distribution of land snails in Hungary. Priorización de la conservación con base en la distribución de caracoles terrestres en Hungría. Conservation Biology, 19(4):1084–1094, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sasaki:2009:REW

[SFD09] Kiyoshi Sasaki, Stanley F. Fox, and David Duvall. Rapid evolution in the wild: Changes in body size, life-history traits, and behavior in hunted populations of the Japanese mamushi snake. Evolución rápida en la naturaleza: Cambios en el tamaño corporal, las características de la historia de vida y la conducta en poblaciones cazadas de la serpiente japonesa mamushi. Conservation Biology, 23(1):93–102, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simons:2000:EGS

[SFS00] Theodore R. Simons, George L. Farnsworth, and Susan A. Shriner. Evaluating Great Smoky Mountains National Park as a population source for the wood thrush. Evaluación del Parque Nacional Great Smoky Mountains como una población fuente para el zorzalito americano Hylocichla mustelina. Conservation Biology, 14(4):1133–1144, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:2005:DPU

[SFS05] Benedikt R. Schmidt, Reiner Feldmann, and Michael Schaub. Demographic processes underlying population growth and decline in Salamandra salamandra. Procesos demográficos subyacentes en el crecimiento y declinación poblacional de Salamandra salamandra. Conservation Biology, 19(4):1149–1156, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Slooten:2000:AUR

[SFT00] Elisabeth Slooten, David Fletcher, and Barbara L. Taylor. Accounting for uncertainty in risk assessment: Case study of Hector’s dolphin mortality due to gillnet entanglement. Consideración de la incertidumbre en la evaluación de riesgos: Caso de estudio de la mortalidad del delfín de Hector debido a enmalles en redes agalleras. Conservation Biology, 14(5):1264–1270, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shen:2008:PCL

[SFX^{+}08] Guozhen Shen, Chaoyang Feng, Zongqiang Xie, Zhiyun Ouyang, Junqing Li, and Marty Pascal. Proposed conservation landscape for giant pandas in the Minshan Mountains, China. Propuesta de paisaje para la conservación de pandas gigantes en las Montañas Minshan, China. Conservation Biology, 22(5):1144–1153, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schlaepfer:2001:EEL

[SG01] Martin A. Schlaepfer and Thomas A. Gavin. Edge effects on lizards and frogs in tropical forest fragments. Efectos de borde sobre lagartijas y ranas en fragmentos de bosque tropical. Conservation Biology, 15(4):1079–1090, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steen:2004:ERS

[SG04] David A. Steen and James P. Gibbs. Effects of roads on the structure of freshwater turtle populations. Efectos de caminos sobre la estructura de poblaciones de tortugas Dulceaucícolas. Conservation Biology, 18(4):1143–1148, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmeller:2008:NRE

[SGB^{+}08] Dirk S. Schmeller, Bernd Gruber, Eduardas Budrys, Erik Framsted, Szabolcs Lengyel, and Klaus Henle. National responsibilities in European species conservation: a methodological review. Responsabilidades nacionales en la conservación de especies europeas: una revisión metodológica. Conservation Biology, 22(3): 593–601, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schiesari:2007:BBR

[SGG07] Luis Schiesari, Britta Grillitsch, and Heinz Grillitsch. Biogeographic biases in research and their consequences for linking amphibian declines to pollution. Sesgos biogeográficos en la investigación y sus consecuencias para asociar las declinaciones de anfibios a la contaminación. Conservation Biology, 21(2):465– 471, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suter:2002:CTU

[SGH02] Werner Suter, Roland F. Graf, and Ruedi Hess. Capercaillie (Tetrao urogallus) and avian biodiversity: Testing the umbrellaspecies concept. Capercaillie (Tetrao urogallus) y biodiversidad de aves: Probando el concepto de especie sombrilla. Conservation Biology, 16(3):778–788, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sjogren-Gulve:2007:CBA

[SGLB^{+}07] Per Sjögren-Gulve, Elisabeth Långström, András Báldi, Pierre Ibisch, Vassiliki Kati, Barbara Livoreil, and Nuria Selva. Conservation Biology and the 300th anniversary of the birth of Carl Linnaeus. Conservation Biology, 21(4):905–906, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schaub:2007:UIM

[SGSA07] Michael Schaub, Olivier Gimenez, Antoine Sierro, and Raphaël Arlettaz. Use of integrated modeling to enhance estimates of population dynamics obtained from limited data. Utilización de modelos integrados para elevar estimasciones de la dinámica poblacional obtenida con datos limitados. Conservation Biology, 21(4):945–955, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2002:FDR

[SH02] Loren M. Smith and David A. Haukos. Floral diversity in relation to playa wetland area and watershed disturbance. Diversidad floral con relación al Área de humedal playero y la perturbación de la cuenca hidrológica. Conservation Biology, 16 (4):964–974, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schultz:2003:UPV

[SH03] Cheryl B. Schultz and Paul C. Hammond. Using population viability analysis to develop recovery criteria for endangered insects: Case study of the Fender’s blue butterfly. Utilización del análisis de viabilidad poblacional para desarrollar criterios de recuperación de insectos en peligro: Estudio de caso de la mariposa azul de fender. Conservation Biology, 17(5):1372–1385, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sethi:2009:RHD

[SH09] Pia Sethi and Henry F. Howe. Recruitment of hornbill-dispersed trees in hunted and logged forests of the Indian Eastern Himalaya. Reclutamiento de Árboles dispersados por Cálaos en bosques con cacería y tala en Himalaya Oriental Hindú. Conservation Biology, 23(3):710–718, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shafer:2001:CBT

[Sha01] Craig L. Shafer. Conservation biology trailblazers: George Wright, Ben Thompson, and Joseph Dixon. Pioneros de la biología de la conservación. Conservation Biology, 15(2):332–344, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaffer:2004:ARW

[Sha04] H. Bradley Shaffer. The amphibians, reptiles and a whole lot more. Conservation Biology, 18(5):1440–1447, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwartz:2008:SSC

[SHB08] Mark W. Schwartz, Malcolm L. Hunter, Jr., and P. Dee Boersma. Scientific societies in the 21st Century: a membership crisis. Conservation Biology, 22(5):1087–1089, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sheil:2001:CBM

[She01] Douglas Sheil. Conservation and biodiversity monitoring in the tropics: Realities, priorities, and distractions. Conservation Biology, 15(4):1179–1182, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shebitz:2002:PTH

[She02] Daniela Shebitz. Piecing together history. Conservation Biology, 16(2):565–566, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shier:2006:EFS

[Shi06] D. M. Shier. Effect of family support on the success of translocated black-tailed prairie dogs. Efecto del soporte familiar sobre el Éxito de perros de la pradera translocados. Conservation Biology, 21(2):1780–1790, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmeller:2009:AVB

[SHJ^{+}09] Dirk S. Schmeller, Pierre-Yves Henry, Romain Julliard, Bernd Gruber, Jean Clobert, Frank Dziock, Szabolcs Lengyel, Piotr Nowicki, Eszter Déri, Eduardas Budrys, Tiiu Kull, Kadri Tali, Bianca Bauch, Josef Settele, Chris Van Swaay, Andrej Kobler, Valerija Babij, Eva Papastergiadou, and Klaus Henle. Advantages of volunteer-based biodiversity monitoring in Europe ventajas del monitoreo de biodiversidad basado en voluntarios en Europa. Conservation Biology, 23(2):307–316, April 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwartz:2000:PPF

[SHM00] Mark W. Schwartz, Sharon M. Hermann, and Phillip J. Van Mantgem. Population persistence in Florida Torreya: Comparing modeled projections of a declining coniferous tree. Persistencia poblacional de Torreya taxifolia: Comparaciones del declive de un arbol conífero basadas en proyecciones de modelos. Conservation Biology, 14(4):1023–1033, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siren:2004:ISH

[SHM04] Anders Sirén, Peter Hambäck, and José Machoa. Including spatial heterogeneity and animal dispersal when evaluating hunting: a model analysis and an empirical assessment in an Amazonian community. Inclusión de heterogeneidad espacial y dispersión de animales al evaluar cacería: un análisis de modelo y una estimación empírica en una comunidad amazónica. Conservation Biology, 18(5):1315–1329, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steinitz:2005:PRP

[SHT^{+}05] Ofer Steinitz, Joseph Heller, Asaf Tsoar, Dotan Rotem, and Ronen Kadmon. Predicting regional patterns of similarity in species composition for conservation planning. Predicción de patrones regionales de similitud en la composición de especies para planificación de conservación. Conservation Biology, 19(6):1978–1988, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spies:2006:COG

[SHYH06] Thomas A. Spies, Miles A. Hemstrom, Andrew Youngblood, and Susan Hummel. Conserving old-growth forest diversity in disturbance-prone landscapes. Conservación de la diversidad de bosques viejos en paisajes propensos a la perturbación. Conservation Biology, 20(2):351–362, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siebert:2000:PFM

[Sie00] Stephen F. Siebert. Primary forests and medicinal plants. Conservation Biology, 14(3):596, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siebert:2004:DEC

[Sie04] Stephen F. Siebert. Demographic effects of collecting rattan cane and their implications for sustainable harvesting. Efectos demográficos de la Recolección de Cañas de Ratán y sus implicaciones para la cosecha sostenible. Conservation Biology, 18 (2):424–431, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silva:2005:BPA

[Sil05] Marina Silva. The Brazilian protected areas program. Conservation Biology, 19(3):608–611, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simberloff:2003:HMI

[Sim03] Daniel Simberloff. How much information on population biology is needed to manage introduced species? `?Cuánta información sobre biología poblacional se necesita para manejar especies introducidas? Conservation Biology, 17(1):83–92, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarkar:2005:EES

[SJF^{+}05] Sahotra Sarkar, James Justus, Trevon Fuller, Chris Kelley, Justin Garson, and Michael Mayfield. Effectiveness of environmental surrogates for the selection of conservation area networks. Efectividad de sustitutos ambientales para la selección de redes de Áreas de conservación. Conservation Biology, 19(3): 815–825, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skelly:2007:ERC

[SJP^{+}07] David K. Skelly, Liana N. Joseph, Hugh P. Possingham, L. Kealoha Freidenburg, Thomas J. Farrugia, Michael T. Kinnison, and Andrew P. Hendry. Evolutionary responses to climate change. Conservation Biology, 21(5):1353–1355, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salas:2002:SFS

[SK02a] Leonardo A. Salas and John B. Kim. Spatial factors and stochasticity in the evaluation of sustainable hunting of tapirs. Factores espaciales y estocasticidad en la evaluación de la cacería sustentable de tapires. Conservation Biology, 16(1):86–96, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stanley:2002:ARL

[SK02b] Thomas R. Stanley and Fritz L. Knopf. Avian responses to late-season grazing in a shrub-willow floodplain. Respuestas de aves al pastoreo tardío en una llanura de inundación con arbustos y sauces. Conservation Biology, 16(1):225–231, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Safi:2004:CAS

[SK04] Kamran Safi and Gerald Kerth. A comparative analysis of specialization and extinction risk in temperate-zone bats. Un análisis comparativo de la Especialización y el riesgo de extinción en murciélagos de zonas templadas. Conservation Biology, 18(5): 1293–1303, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Safina:2008:CBT

[SK08] Carl Safina and Dane H. Klinger. Collapse of bluefin tuna in the Western Atlantic. Conservation Biology, 22(2):243–246, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snodgrass:2000:RAI

[SKBB00] Joel W. Snodgrass, Mark J. Komoroski, A. Lawrence Bryan, Jr., and Joanna Burger. Relationships among isolated wetland size, hydroperiod, and amphibian species richness: Implications for wetland regulations. Relaciones entre el tamaño de humedales aislados, el hidroperíodo y la riqueza de especies de anfibios: Implicaciones para la regulación de humedales. Conservation Biology, 14(2):414–419, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sayama:2003:SPF

[SKBY03] Hiroki Sayama, Les Kaufman, and Yaneer Bar-Yam. Spontaneous pattern formation and genetic diversity in habitats with irregular geographical features. Formación Espontánea de patrones y diversidad genética en hábitats con características geográficas irregulares. Conservation Biology, 17(3):893–900, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Starnes:2000:SES

[SKP00] Susan M. Starnes, Caroline A. Kennedy, and James W. Petranka. Sensitivity of embryos of Southern Appalachian amphibians to ambient solar UV-B radiation. Sensitividad de embriones de anfibios de los Apalaches del Sur a la radiación UV-B solar ambiental. Conservation Biology, 14(1):277–282, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarrazin:2000:DAR

[SL00a] François Sarrazin and Stéphane Legendre. Demographic approach to releasing adults versus young in reintroductions. Aproximación demográfica a la liberación de adultos contra juveniles en reintroducciones. Conservation Biology, 14(2):488– 500, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sodhi:2000:ICB

[SL00b] Navjot S. Sodhi and L. H. Liow. Improving conservation biology research in Southeast Asia. Conservation Biology, 14(4):1211– 1212, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Severns:2008:ICN

[SL08] Paul M. Severns and Aaron Liston. Intraspecific chromosome number variation: a neglected threat to the conservation of rare plants. Variación Intraespecífica en el número de cromosomas: una amenaza desatendida en la conservación de plantas raras. Conservation Biology, 22(6):1641–1647, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sekercioglu:2007:PFB

[SLB^{+}07] Cagan H. Sekercioglu, Scott R. Loarie, Federico Oviedo Brenes, Paul R. Ehrlich, and Gretchen C. Daily. Persistence of forest birds in the Costa Rican agricultural countryside. Persistencia de aves en el paisaje agrícola de Costa Rica. Conservation Biology, 21(2):482–494, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schierenbeck:2004:RA

[SLC^{+}04] Kristina A. Schierenbeck, Glen Lubcke, John Campbell, Michelle Cederborg, and Herman Gray. Refreshing approaches. Conservation Biology, 18(5):1441–1443, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sheth:2008:EDE

[SLCJ08] Seema Nayan Sheth, Lúcia G. Lohmann, Trisha Consiglio, and Iván Jiménez. Effects of detectability on estimates of geographic range size in bignonieae. Efectos de la detectabilidad sobre estimaciones del rango Geográfico de bignonieae. Conservation Biology, 22(1):200–211, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Southwell:2008:SPV

[SLCW08] Darren M. Southwell, Alex M. Lechner, Terry Coates, and Brendan A. Wintle. The sensitivity of population viability analysis to uncertainty about habitat requirements: Implications for the management of the endangered southern brown bandicoot. La sensibilidad del análisis de viabilidad poblacional a la incertidumbre sobre requerimientos de hábitat: Implicaciones para el manejo de Isoodon obesulus. Conservation Biology, 22(4):1045– 1054, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schilthuizen:2005:EKF

[SLELA05] Menno Schilthuizen, Thor-Seng Liew, Berjaya Bin Elahan, and Isabelle Lackman-Ancrenaz. Effects of karst forest degradation on pulmonate and prosobranch land snail communities in Sabah, Malaysian Borneo. Efectos de la degradación del bosque sobre comunidades de caracoles terrestres pulmonados y prosobranquios en Sabah, Malasia. Conservation Biology, 19(3):949–954, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sweanor:2000:CDP

[SLH00] Linda L. Sweanor, Kenneth A. Logan, and Maurice G. Hornocker. Cougar dispersal patterns, metapopulation dynamics, and conservation. Patrones de dispersión de jaguares, dinámica metapoblacional y conservación. Conservation Biology, 14(3):798–808, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sodhi:2006:LTA

[SLKP06] Navjot S. Sodhi, Tien Ming Lee, Lian Pin Koh, and Dewi M. Prawiradilaga. Long-term avifaunal impoverishment in an isolated tropical woodlot empobrecimiento de la avifauna a largo plazo en un terreno boscoso tropical. Conservation Biology, 20 (3):772–779, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sloan:2002:HAW

[Slo02] N. A. Sloan. History and application of the wilderness concept in marine conservation. Historia y aplicación del concepto de “Silvestre” en conservación marina. Conservation Biology, 16 (2):294–305, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siitonen:2005:EFE

[SLS05] Paula Siitonen, Antti Lehtinen, and Mikko Siitonen. Effects of forest edges on the distribution, abundance, and regional persistence of wood-rotting fungi. Efectos de los bordes de bosque sobre la distribución, abundancia y persistencia regional de hongos de madera en Pudrición. Conservation Biology, 19(1):250– 260, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2004:Lb

[SLTG04] Paul P. Smith, Porter P. Lowry II, Jonathan Timberlake, and Janice S. Golding. Letter. Conservation Biology, 18(1):8–9, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Song:2001:SPS

[SM01] Samantha J. Song and R. Michael M’Gonigle. Science, power, and system dynamics: the political economy of conservation biology. Ciencia, poder y dinámica de sistemas: la economía política de la biología de la conservación. Conservation Biology, 15(4):980–989, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmeller:2007:DGE

[SM07a] Dirk S. Schmeller and Juha Merilä. Demographic and genetic estimates of effective population and breeding size in the amphibian Rana temporaria. Estimaciones demográficas y genéticas del tamaño poblacional efectivo y reproductivo en el anfibio Rana temporaria. Conservation Biology, 21(1):142–151, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soberon:2007:CST

[SM07b] Jorge Soberon and Rodrigo A. Medellín. Categorization systems of threatened species. Conservation Biology, 21(5):1366–1367, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schleuning:2009:HCP

[SM09] Matthias Schleuning and Diethart Matthies. Habitat change and plant demography: Assessing the extinction risk of a formerly common grassland perennial. Cambio de hábitat y demografía de plantas: Evaluación del riesgo de extinción de una planta perenne de pastizal anteriormente común. Conservation Biology, 23(1):174–183, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shutler:2000:BCP

[SMC00] Dave Shutler, Adele Mullie, and Robert G. Clark. Bird communities of prairie uplands and wetlands in relation to farming practices in Saskatchewan. Comunidades de aves de praderas de altiplano y humedales en relación con prácticas de labranza en Saskatchewan, Canadá. Conservation Biology, 14(5):1441–1451, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Solis-Montero:2005:SCP

[SMFPCA05] Lislie Solis-Montero, Alejandro Flores-Palacios, and Andrea Cruz-Angón. Shade-coffee plantations as refuges for tropical wild orchids in Central Veracruz, Mexico. Plantaciones cafetaleras de sombra como refugio de orquídeas silvestres en Veracruz Central, México. Conservation Biology, 19(3):908–916, June 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sher:2000:CBN

[SMG00] Anna A. Sher, Diane L. Marshall, and Steven A. Gilbert. Competition between native Populus deltoides and invasive Tamarix ramosissima and the implications for reestablishing flooding disturbance competencia entre la especie nativa Populus deltoides y la invasiva Tamarix ramosissim a y las implicaciones para reintegrar perturbaciones causadas por inundaciones. Conservation Biology, 14(6):1744–1754, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sinclair:2007:LTE

[SMH^{+}07] A. R. E. Sinclair, Simon A. R. Mduma, J. Grant C. Hopcraft, John M. Fryxell, Ray Hilborn, and Simon Thirgood. Long-term ecosystem dynamics in the Serengeti: Lessons for conservation. Dinámica del ecosistema a largo plazo en el Serengeti: Lecciones para la conservación. Conservation Biology, 21(3):580–590, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2004:EYW

[Smi04a] Andrew T. Smith. Everything you wanted to know about conservation biology. Conservation Biology, 18(1):283–284, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2004:La

[Smi04b] Tyler Smith. Letter. Conservation Biology, 18(1):7–8, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2005:MMA

[Smi05] Douglas W. Smith. Mixed messages about opportunistic carnivores. Conservation Biology, 19(5):1676–1678, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shekelle:2007:WEG

[SMI^{+}07] Myron Shekelle, Rudolf Meier, Mochamad Indrawan, Ibnu Maryanto, Agus Salim, Jatna Supriatna, and Noviar Andayani. When “Not Extinct” is not good news: Conservation in the Sangihe Islands. Conservation Biology, 21(1):4–5, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2009:EFC

[Smi09] Thomas B. Smith. Equating forest conservation with hornbill conservation. Conservation Biology, 23(3):782–783, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stratford:2000:MKP

[SMLB00] Elaine Stratford, Nicole Mazur, Daniel Lunney, and David Bennett. Managing the koala problem: Interdisciplinary perspectives manejo del problema del koala: Perspectivas interdisciplinarias. Conservation Biology, 14(3):610–618, June 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanford:2009:EUD

[SMM09] Monte P. Sanford, Patricia N. Manley, and Dennis D. Murphy. Effects of urban development on ant communities: Implications for ecosystem services and management. Efectos del desarrollo urbano sobre comunidades de hormigas: Implicaciones para los servicios y manejo del ecosistema. Conservation Biology, 23(1): 131–141, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwartzman:2000:RTF

[SMN00] Stephan Schwartzman, Adriana Moreira, and Daniel Nepstad. Rethinking tropical forest conservation: Perils in parks. Reconsideración de la conservación de los bosques tropicales: Peligros en parques. Conservation Biology, 14(5):1351–1357, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:2002:IPC

[SMRR02] Nick Salafsky, Richard Margoluis, Kent H. Redford, and John G. Robinson. Improving the practice of conservation: a conceptual framework and research agenda for conservation science. Mejoramiento de la Práctica de la conservación: un marco conceptual y un programa para la ciencia de la conservación. Conservation Biology, 16(6):1469–1479, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stem:2005:MEC

[SMSB05] Caroline Stem, Richard Margoluis, Nick Salafsky, and Marcia Brown. Monitoring and evaluation in conservation: a review of trends and approaches. Monitoreo y evaluación en conservación: una revisión de tendencias y métodos. Conservation Biology, 19(2):295–309, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stocks:2007:ICG

[SMT07] Anthony Stocks, Benjamin Mcmahan, and Peter Taber. Indigenous, colonist, and government impacts on Nicaragua’s Bosawas Reserve. Impactos de indígenas, colonizadores y gobierno sobre la Reserva Bosawas de Nicaragua. Conservation Biology, 21(6): 1495–1505, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saunders:2002:FPA

[SMV02] D. L. Saunders, J. J. Meeuwig, and A. C. J. Vincent. Freshwater protected areas: Strategies for conservation. Áreas protegidas de agua dulce: Estrategias para la conservación. Conservation Biology, 16(1):30–41, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2007:RLB

[SMW^{+}07] Robert J. Smith, Robert D. J. Muir, Matthew J. Walpole, Andrew Balmford, and Nigel Leader-Williams. Revisiting the links between governance and biodiversity loss. Conservation Biology, 21(4):899–900, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwartzman:2000:ATF

[SNM00] Stephan Schwartzman, Daniel Nepstad, and Adriana Moreira. Arguing tropical forest conservation: People versus parks. Conservation Biology, 14(5):1370–1374, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suazo-Ortuno:2008:ECD

[SOADMR08] Ireri Suazo-Ortuño, Javier Alvarado-Díaz, and Miguel Martínez-Ramos. Effects of conversion of dry tropical forest to agricultural mosaic on herpetofaunal assemblages. Efectos de la Conversión de bosque tropical seco a mosaico agrícola sobre ensambles Herpetofaunísticos. Conservation Biology, 22(2):362–374, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sodhi:2006:SHB

[Sod06] Navjot S. Sodhi. Sorrow and hope for biodiversity in a degraded landscape. Conservation Biology, 20(1):262, February 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soderstrom:2000:ILS

[SP00] Bo Söderström and Tomas Pärt. Influence of landscape scale on farmland birds breeding in semi-natural pastures influencia de la escala del paisaje sobre aves que anidan en pastizales seminaturales. Conservation Biology, 14(2):522–533, April 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:2009:BCI

[SP09] Paige M. Schmidt and Markus J. Peterson. Biodiversity conservation and indigenous land management in the era of selfdetermination. Conservación de la biodiversidad y manejo de tierras indígenas en la era de la autodeterminación. Conservation Biology, 23(6):1458–1466, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spear:2000:CMC

[Spe00] John R. Spear. Conservation medicine: the changing view of biodiversity. Conservation Biology, 14(6):1913–1917, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spector:2002:BCP

[Spe02] Sacha Spector. Biogeographic crossroads as priority areas for biodiversity conservation. Intersecciones biogeográficas como Áreas prioritarias para la conservación de la biodiversidad. Conservation Biology, 16(6):1480–1487, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spear:2004:MIR

[Spe04] John R. Spear. Minimum-impact research. Conservation Biology, 18(4):861, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Speth:2007:AOI

[Spe07] James Gustave Speth. All for one: the immediacy of fighting climate change. Conservation Biology, 21(4):914, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saura:2008:PPA

[SPFF^{+}08] María Saura, AndréS Pérez-Figueroa, Jesús Fernández, Miguel A. Toro, and Armando Caballero. Preserving population allele frequencies in ex situ conservation programs. Preservación de las frecuencias alélicas en programas de conservación ex situ. Conservation Biology, 22(5):1277–1287, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schulte:2005:OHF

[SPM05] Lisa A. Schulte, Anna M. Pidgeon, and David J. Mladenoff. One hundred fifty years of change in forest bird breeding habitat: Estimates of species distributions. Ciento cincuenta años de cambios en el hábitat para reproducción de aves de bosque: Estimación de la distribución de especies. Conservation Biology, 19(6):1944–1956, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanderson:2002:PSS

[SRC^{+}02] Eric W. Sanderson, Kent H. Redford, Cheryl-Lesley B. Chetkiewicz, Rodrigo A. Medellin, Alan R. Rabinowitz, John G. Robinson, and Andrew B. Taber. Planning to save a species: the jaguar as a model. Planeación para salvar una especie: El jaguar como modelo. Conservation Biology, 16(1):58–72, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shelden:2003:EBW

[SRDG03] Kim E. W. Shelden, David J. Rugh, Douglas P. DeMaster, and Leah R. Gerber. Evaluation of bowhead whale status: Reply to Taylor. Conservation Biology, 17(3):918–920, June 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Semlitsch:2007:SAA

[SRH^{+}07] Raymond D. Semlitsch, Travis J. Ryan, Kevin Hamed, Matt Chatfield, Bethany Drehman, Nicole Pekarek, Mike Spath, and Angie Watland. Salamander abundance along road edges and within abandoned logging roads in Appalachian forests. Abundancia de salamandras a lo largo de bordes de caminos y en caminos madereros abandonados en bosques Apalaches. Conservation Biology, 21(1):159–167, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Syphard:2009:CTD

[SRHS09] Alexandra D. Syphard, Volker C. Radeloff, Todd J. Hawbaker, and Susan I. Stewart. Conservation threats due to human-caused increases in fire frequency in Mediterranean-climate ecosystems. Amenazas a la conservación debido a incrementos en la frecuencia de incendios causados por humanos en ecosistemas de clima Mediterráneo. Conservation Biology, 23(3):758–769, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:2007:PAS

[SRL^{+}07] J. Michael Scott, Janet L. Rachlow, Robert T. Lackey, Anna B. Pidgorna, Jocelyn L. Aycrigg, Gabrielle R. Feldman, Leona K. Svancara, David A. Rupp, David I. Stanish, and R. Kirk Steinhorst. Policy advocacy in science: Prevalence, perspectives, and implications for conservation biologists. Conservation Biology, 21(1):29–35, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saltz:2000:ESU

[SRR00] David Saltz, Mary Rowen, and Daniel I. Rubenstein. The effect of space-use patterns of reintroduced Asiatic wild ass on effective population size. Efecto de los patrones de uso del espacio en el tamaño poblacional efectivo en asnos asiáticos silvestres reintroducidos. Conservation Biology, 14(6):1852–1861, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salomon:2001:IHE

[SRSP01] Anne K. Salomon, Jennifer L. Ruesink, Brice X. Semmens, and Robert T. Paine. Incorporating human and ecological communities in marine conservation: an alternative to Zacharias and Roff. Conservation Biology, 15(5):1452–1455, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [ZR00].

Saltz:2006:IIE

[SRW06] David Saltz, Daniel I. Rubenstein, and Gary C. White. The impact of increased environmental stochasticity due to climate change on the dynamics of Asiatic wild ass. El impacto del incremento de la estocasticidad ambiental debida al cambio climático sobre la dinámica de Equus hemionus. Conservation Biology, 20 (5):1402–1409, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanderson:2008:EFN

[SRW^{+}08] Eric W. Sanderson, Kent H. Redford, Bill Weber, Keith Aune, Dick Baldes, Joel Berger, Dave Carter, Charles Curtin, James Derr, Steve Dobrott, Eva Fearn, Craig Fleener, Steve Forrest, Craig Gerlach, C. Cormack Gates, John E. Gross, Peter Gogan, Shaun Grassel, Jodi A. Hilty, Marv Jensen, Kyran Kunkel, Duane Lammers, Rurik List, Karen Minkowski, Tom Olson, Chris Pague, Paul B. Robertson, and Bob Stephenson. The ecological future of the North American bison: Conceiving longterm, large-scale conservation of wildlife. El futuro ecológico del bisonte de Norte América: Concepción de la conservación de vida silvestre a largo plazo y a gran escala. Conservation Biology, 22(2):252–266, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shkedy:2000:CCC

[SS00] Yehoshua Shkedy and David Saltz. Characterizing core and corridor use by Nubian ibex in the Negev Desert, Israel. Caracterización del uso de zona núcleo y corredores por Íbices de Nubia en el desierto del Negev, Israel. Conservation Biology, 14(1): 200–206, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stratford:2001:RFG

[SS01] Jeffrey A. Stratford and Philip C. Stouffer. Reduced feather growth rates of two common birds inhabiting Central Amazonian forest fragments. Reducción de la tasas de crecimiento de lumas de dos aves canoras de fragmentos de la Amazonía Central. Conservation Biology, 15(3):721–728, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shetty:2002:PHB

[SS02a] Sohan Shetty and Richard Shine. Philopatry and homing behavior of sea snakes (Laticauda colubrina) from two adjacent islands in Fiji. Filopatría y conducta de búsqueda del hogar en la serpiente marina (Laticauda colubrina) en dos islas adyacentes en Fiji. Conservation Biology, 16(5):1422–1426, October 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stier:2002:KPO

[SS02b] Sam C. Stier and Stephen F. Siebert. The Kyoto protocol: an opportunity for biodiversity restoration forestry. Conservation Biology, 16(3):575–576, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stier:2003:L

[SS03a] Sam C. Stier and Stephen F. Siebert. Letters. Conservation Biology, 17(1):5, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Swarthout:2003:EEH

[SS03b] Elliott C. H. Swarthout and Robert J. Steidl. Experimental effects of hiking on breeding Mexican spotted owls. Efectos experimentales del excursionismo sobre búhos manchados mexicanos reproductores. Conservation Biology, 17(1):307–315, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silvano:2005:CBA

[SS05] Débora L. Silvano and Magno V. Segalla. Conservation of Brazilian amphibians. Conservación de anfibios brasileños. Conservation Biology, 19(3):653–658, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stankey:2006:FSA

[SS06] George H. Stankey and Bruce Shindler. Formation of social acceptability judgments and their implications for management of rare and little-known species. Generación de juicios de aceptabilidad social y sus implicaciones para la gestión de especies raras y poco conocidas. Conservation Biology, 20(1):28–37, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Searcy:2008:CBA

[SS08a] Christopher A. Searcy and H. Bradley Shaffer. Calculating biologically accurate mitigation credits: Insights from the California tiger salamander. Cálculo de créditos de mitigación biológicamente precisos: Perspectivas para la salamandra tigre de California. Conservation Biology, 22(4):997–1005, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Serfass:2008:DD

[SS08b] Thomas L. Serfass and Sadie S. Stevens. Detailed devotion. Conservation Biology, 22(3):806–807, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shears:2008:EBC

[SSB^{+}08] Nick T. Shears, Franz Smith, Russ C. Babcock, Clinton A. J. Duffy, and Eduardo Villouta. Evaluation of biogeographic classification schemes for conservation planning: Application to New Zealand’s coastal marine environment. Evaluación de los esquemas de clasificación biogeográfica para la planificación de la conservación: Aplicación en el ambiente marino costero de Nueva Zelanda. Conservation Biology, 22(2):467–481, April 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sekercioglu:2008:CCE

[SSFL08] Cagan H. Sekercioglu, Stephen H. Schneider, John P. Fay, and Scott R. Loarie. Climate change, elevational range shifts, and bird extinctions. Cambio climático, desplazamiento de rangos altitudinales y extinciones de aves. Conservation Biology, 22(1): 140–150, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stowe:2001:TBF

[SSG01] Johnny P. Stowe, Jr., Elise Vernon Schmidt, and Deborah Green. Toxic burials: The final insult. Conservation Biology, 15 (6):1817–1819, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shi:2005:IHS

[SSK^{+}05] Hua Shi, Ashibindu Singh, Shashi Kant, Zhiliang Zhu, and Eric Waller. Integrating habitat status, human population pressure, and protection status into biodiversity conservation priority setting. Conservation Biology, 19(4):1273–1285, August 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2006:ERI

[SSL06] Katherine F. Smith, Dov F. Sax, and Kevin D. Lafferty. Evidence for the role of infectious disease in species extinction and endangerment. Evidencia del papel de enfermedades infecciosas en la extinción y puesta en peligro de especies. Conservation Biology, 20(5):1349–1357, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:2003:RGC

[SSM03] Michael L. Scott, Susan K. Skagen, and Michael F. Merigliano. Relating geomorphic change and grazing to avian communities in riparian forests. Relación entre el cambio geomórfico y el pastoreo con las comunidades de aves en bosques riparios. Conservation Biology, 17(1):284–296, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:2008:SLB

[SSS^{+}08] Nick Salafsky, Daniel Salzer, Alison J. Stattersfield, Craig Hilton-Taylor, Rachel Neugarten, Stuart H. M. Butchart, Ben Collen, Neil Cox, Lawrence L. Master, Sheila O’Connor, and David Wilkie. A standard lexicon for biodiversity conservation: Unified classifications of threats and actions. Un lexicón estándar para la conservación de biodiversidad: Clasificaciones unificadas de amenazas y acciones. Conservation Biology, 22(4):897–911, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comment [BCK^{+}09].

Schmiegelow:2006:RSL

[SSSK06] Fiona K. A. Schmiegelow, David P. Stepnisky, Curtis A. Stambaugh, and Matti Koivula. Reconciling salvage logging of boreal forests with a natural-disturbance management model. Reconciliación de la cosecha de salvamento de bosques boreales con un modelo de gestión de perturbación natural. Conservation Biology, 20(4):971–983, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sigel:2006:ACR

[SSY06] Bryan J. Sigel, Thomas W. Sherry, and Bruce E. Young. Avian community response to lowland tropical rainforest isolation: 40 years of change at La Selva Biological Station, Costa Rica. Respuesta de la comunidad aviar al aislamiento de selva tropical inundable: 40 años de cambios en la Estación Biológica La Selva, Costa Rica. Conservation Biology, 20(1):111–121, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spencer:2005:EAI

[ST05] Ricky-John Spencer and Michael B. Thompson. Experimental analysis of the impact of foxes on freshwater turtle populations. Análisis experimental del impacto de zorros sobre poblaciones de tortugas dulceacuícolas. Conservation Biology, 19(3):845– 854, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stallard:2001:PEF

[Sta01] Robert F. Stallard. Possible environmental factors underlying amphibian decline in Eastern Puerto Rico: Analysis of U.S. government data archives. Posibles factores ambientales subyacentes en la disminución de anfibios en Puerto Rico oriental: Análisis de archivos de datos del gobierno de los Estados Unidos. Conservation Biology, 15(4):943–953, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stacey:2003:WF

[Sta03] Peter Stacey. Woodpecker fundamentals. Conservation Biology, 17(3):934–937, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Starzomski:2009:PPT

[Sta09] Brian Starzomski. Patterns, processes, and the textural discontinuity hypothesis. Conservation Biology, 23(4):1062–1063, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shivik:2003:NTM

[STC03] John A. Shivik, Adrian Treves, and Peggy Callahan. Nonlethal techniques for managing predation: Primary and secondary repellents. Técnicas no letales para gestión de la depredación: Repelentes primarios y secundarios. Conservation Biology, 17(6): 1531–1537, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steininger:2001:CFT

[STE^{+}01] M. K. Steininger, C. J. Tucker, P. Ersts, T. J. Killeen, Z. Villegas, and S. B. Hecht. Clearance and fragmentation of tropical deciduous forest in the Tierras Bajas, Santa Cruz, Bolivia. Tala y fragmentación del bosque tropical deciduo de Tierras Bajas, Santa Cruz, Bolivia. Conservation Biology, 15(4):856–866, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stergiou:2005:TLS

[Ste05] Konstantinos I. Stergiou. Time-lagged scientific dialogues. Conservation Biology, 19(3):983–985, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stewart:2006:BS

[Ste06a] Scott R. Stewart. Biodiversity and soils. Conservation Biology, 20(2):594, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stewart:2006:SP

[Ste06b] Susan I. Stewart. Sprawl’s pastiche. Conservation Biology, 21 (2):1825–1828, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sterling:2007:CET

[Ste07] Eleanor J. Sterling. Conservation education treasure trove. Conservation Biology, 21(2):893–894, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steffe:2008:DEM

[Ste08a] Aldo S. Steffe. Dependence of effective marine park zoning on survey design, data quality, and community acceptance: Response to Lynch. Conservation Biology, 22(2):482–485, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stephens:2008:TMC

[Ste08b] Philip A. Stephens. Two may be company but three is seldom a crowd: Allee effects in ecology and conservation. Conservation Biology, 22(6):1662–1664, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siitonen:2002:MSO

[STL02] Paula Siitonen, Antti Tanskanen, and Antti Lehtinen. Method for selection of old-forest reserves. Método de selección de reservas de bosques maduros. Conservation Biology, 16(5):1398– 1408, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stockwell:2004:CBB

[Sto04] Craig A. Stockwell. Conservation biology for biologists. Conservation Biology, 18(1):284–286, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stergiou:2009:FMF

[STP09] Konstantinos I. Stergiou, Athanassios C. Tsikliras, and Daniel Pauly. Farming up Mediterranean food webs. Conservation Biology, 23(1):230–232, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Staples:2005:RBV

[STS05] David F. Staples, Mark L. Taper, and Bradley B. Shepard. Risk-based viable population monitoring monitoreo de poblaciones viables con base en riesgos. Conservation Biology, 19(6): 1908–1916, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Samways:2005:ERF

[STT05] M. J. Samways, S. Taylor, and W. Tarboton. Extinction reprieve following alien removal. Conservation Biology, 19(4):1329–1330, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stuart:2001:RE

[Stu01] John Stuart. Redwoods.edu. Conservation Biology, 15(3):809– 810, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sherwin:2000:ACI

[STWH00] William B. Sherwin, Peter Timms, Jonathan Wilcken, and Bronwyn Houlden. Analysis and conservation implications of koala genetics. Análisis y consecuencias de la conservación genética del koala. Conservation Biology, 14(3):639–649, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suarez:2002:CCC

[Sua02] Andrew V. Suarez. Conservation of the common chameleon. Conservation Biology, 16(6):1665, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sullivan:2009:GPL

[Sul09] Brian K. Sullivan. The greening of public lands grazing in the Southwestern U.S.A. Conservation Biology, 23(4):1047–1049, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Surman:2002:CC

[Sur02] Chris Surman. A compendium of corals. Conservation Biology, 16(5):1438–1439, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sloan:2007:FBC

[SVBR07] N. A. Sloan, Ken Vance-Borland, and G. Carleton Ray. Fallen between the cracks: Conservation linking land and sea. Conservation Biology, 21(4):897–898, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sahley:2007:BSL

[SVV07] Catherine Teresa Sahley, Jorge Torres Vargas, and Jesus Sanchez Valdivia. Biological sustainability of live shearing of vicuña in Peru. La sostenibilidad biológica del trasquilado de vicuñas vivas en Perú. Conservation Biology, 21(1):98–105, February 2007.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanudo-Wilhelmy:2007:HCO

[SW07] Sergio A. Sañudo-Wilhelmy. Health of the coastal ocean. Conservation Biology, 21(4):1130, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Swaisgood:2009:ZDB

[Swa09] Ronald R. Swaisgood. Zoos dream of becoming conservation NGOs. Conservation Biology, 23(5):1338–1340, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stephens:2008:PRH

[SWB^{+}08] Scott E. Stephens, Johann A. Walker, Darin R. Blunck, Aneetha Jayaraman, David E. Naugle, James K. Ringelman, and Aaron J. Smith. Predicting risk of habitat conversion in native temperate grasslands. Predicción del riesgo de Conversión de hábitat en pastizales templados nativos. Conservation Biology, 22(5):1320–1330, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sieving:2000:DCF

[SWD00] Kathryn E. Sieving, Mary F. Willson, and Toni L. De Santo. Defining corridor functions for endemic birds in fragmented south-temperate rainforest. Definición de las funciones de un corredor para especies de aves endémicas en un bosque lluvioso templado del sur fragmentado. Conservation Biology, 14(4): 1120–1132, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Swenson:2008:WKA

[Swe08] Jon E. Swenson. What is known about mountain goats and a model for studying large mammals. Conservation Biology, 22 (4):1082–1083, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schrott:2005:DLA

[SWK05] Gregory R. Schrott, Kimberly A. With, and Anthony W. King. Demographic limitations of the ability of habitat restoration to rescue declining populations. Limitaciones demográficas de la capacidad de la restauración de hábitat para rescatar poblaciones en declinación. Conservation Biology, 19(4):1181–1193, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2000:CEU

[SWR00] Geoffrey R. Smith, Mary Ann Waters, and Jessica E. Rettig. Consequences of embryonic UV-B exposure for embryos and tadpoles of the plains leopard frog. Consecuencias de la exposición Embriónica a la UV-B para embriones y renacuajos de la rana leopardo de las planicies. Conservation Biology, 14(6): 1903–1907, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skelly:2003:EDD

[SYWR03] David K. Skelly, Kerry L. Yurewicz, Earl E. Werner, and Rick A. Relyea. Estimating decline and distributional change in amphibians. Estimación de la declinación y cambio distribucional de anfibios. Conservation Biology, 17(3):744–751, June 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwartzman:2005:CAI

[SZ05] Stephan Schwartzman and Barbara Zimmerman. Conservation alliances with indigenous peoples of the Amazon. Alianzas de conservación con indígenas del Amazonas. Conservation Biology, 19(3):721–727, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanchez-DeLeon:2003:RNE

[SZBR03] Yaniria Sanchez-De Leon, Xiaoming Zou, Sonia Borges, and Honghua Ruan. Recovery of native earthworms in abandoned tropical pastures. Recuperación de lombrices de tierra nativas en pastizales tropicales abandonados. Conservation Biology, 17 (4):999–1006, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Singer:2001:RPS

[SZS01] Francis J. Singer, Linda C. Zeigenfuss, and Leslie Spicer. Role of patch size, disease, and movement in rapid extinction of bighorn sheep. El papel del tamaño del parche, enfermedades y movimientos en la rápida extinción del borrego cimarrón. Conservation Biology, 15(5):1347–1354, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:2006:DWT

[TA06] Peter O. Thomas and Mark R. Albert. Data on wildlife trade. Conservation Biology, 20(3):597–598, June 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Todd:2008:RRG

[TA08] Brian D. Todd and Kimberly M. Andrews. Response of a reptile guild to forest harvesting respuesta de un gremio de reptiles a la cosecha de bosques. Conservation Biology, 22(3):753–761, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2004:EMA

[TAJ04] Jonathan R. Thompson, Mark D. Anderson, and K. Norman Johnson. Ecosystem management across ownerships: the potential for collision with antitrust laws. Gestión de ecosistemas a través de propiedades: El potencial de colisión con leyes antimonopolio. Conservation Biology, 18(6):1475–1481, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tallamy:2004:DAP

[Tal04] Douglas W. Tallamy. Do alien plants reduce insect biomass? Conservation Biology, 18(6):1689–1692, December 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tutin:2001:CBF

[TAP^{+}01] Caroline E. G. Tutin, Marc Ancrenaz, Jorge Paredes, Myriam Vacher-Vallas, Carmen Vidal, Benoît Goossens, Michael W. Bruford, and Aliette Jamart. Conservation biology framework for the release of wild-born orphaned chimpanzees into the Conkouati Reserve, Congo. Marco de trabajo de conservación biológica para la liberación de chimpancés huérfanos silvestres nacidos en la Reserva Conkouati, Congo. Conservation Biology, 15(5):1247–1257, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tasch:2009:RBO

[Tas09] Woody Tasch. The reality behind our economic myth. Conservation Biology, 23(2):256, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2003:GDT

[Tay03a] Charles E. Taylor. Genetic deterioration in threatened and endangered species. Conservation Biology, 17(5):1468–1469, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2003:WBC

[Tay03b] Martin Taylor. Why the Bering–Chukchi–Beaufort Seas bowhead whale is endangered: Response to Shelden et al. Conservation Biology, 17(3):915–917, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SDRO01].

Thibault:2001:SHR

[TB01] Marc Thibault and Sonia Blaney. Sustainable human resources in a protected area in Southwestern Gabon. Recursos humanos sostenibles en una rea protegida en el suroeste de Gabón. Conservation Biology, 15(3):591–595, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thibault:2003:OIU

[TB03] Marc Thibault and Sonia Blaney. The oil industry as an underlying factor in the bushmeat crisis in Central Africa. La industria petrolera como un factor subyacente de la crisis de carne de monte en África Central. Conservation Biology, 17(6): 1807–1813, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2004:DPA

[TB04] Frank R. Thompson III and Dirk E. Burhans. Differences in predators of artificial and real songbird nests: Evidence of bias in artificial nest studies diferencias entre nidos artificiales y reales de aves canoras: Evidencia de sesgos en estudios de nidos artificiales. Conservation Biology, 18(2):373–380, April 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2000:EDE

[TCSD00] Barbara L. Taylor, Susan J. Chivers, Stephanie Sexton, and Andrew E. Dizon. Evaluating dispersal estimates using mtDNA data: Comparing analytical and simulation approaches. Evaluación de estimaciones de dispersión usando datos de ADNmt: Comparación de estrategias analíticas y de simulación. Conservation Biology, 14(5):1287–1297, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Templeton:2002:SGB

[Tem02] Alan R. Templeton. The Speke’s gazelle breeding program as an illustration of the importance of multilocus genetic diversity in conservation biology: Response to Kalinowski et al. Conservation Biology, 16(4):1151–1155, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Terborgh:2000:FTF

[Ter00] John Terborgh. The fate of tropical forests: a matter of stewardship. Conservation Biology, 14(5):1358–1361, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Terborgh:2004:RSW

[Ter04] John Terborgh. Reflections of a scientist on the world parks Congress. Conservation Biology, 18(3):619–620, June 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Terborgh:2005:L

[Ter05] John Terborgh. Letters: 2. Conservation Biology, 19(1):5–6, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trombulak:2000:REE

[TF00] Stephen C. Trombulak and Christopher A. Frissell. Review of ecological effects of roads on terrestrial and aquatic communities. Revisión de los efectos de carreteras en comunidades terrestres y acuáticas. Conservation Biology, 14(1):18–30, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tallis:2008:LTM

[TFG08] Heather Tallis, Zach Ferdaña, and Elizabeth Gray. Linking terrestrial and marine conservation planning and threats analysis. Concatenación del análisis de amenazas y la planificación de la conservación terrestre y marina. Conservation Biology, 22(1): 120–130, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:2006:NFP

[TFGJ06] Jack Ward Thomas, Jerry F. Franklin, John Gordon, and K. Norman Johnson. The northwest forest plan: Origins, components, implementation experience, and suggestions for change. El plan forestal del noroeste: Oigenes, componentes, experiencia de implementación y sugerencias de cambio. Conservation Biology, 20(2):277–287, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tempel:2004:FRF

[TG04] Douglas J. Tempel and R. J. Gutiérrez. Factors related to fecal corticosterone levels in California spotted owls: Implications for assessing chronic stress. Factores relacionados a niveles de corticosterona fecal en búhos Strix occidentalis occidentalis: Implicaciones para la evaluación de estrés Crónico. Conservation Biology, 18(2):538–547, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tabarelli:2005:LFR

[TG05] Marcello Tabarelli and Claude Gascon. Lessons from fragmentation research: Improving management and policy guidelines for biodiversity conservation. Lecciones de la investigación sobre fragmentación: Mejoramiento de las directrices políticas y de gestión para la conservación de biodiversidad. Conservation Biology, 19(3):734–739, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tissot:2003:EAC

[TH03] Brian N. Tissot and Leon E. Hallacher. Effects of aquarium collectors on coral reef fishes in Kona, Hawaii. Efectos de colectores de acuario sobre los peces de arrecifes de coral en Kona, Hawai. Conservation Biology, 17(6):1759–1768, December 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Theobald:2003:TCA

[The03] David M. Theobald. Targeting conservation action through assessment of protection and exurban threats. Orientación de acciones de conservación a Través de evaluación de la protección y amenazas exurbanas. Conservation Biology, 17(6):1624–1637, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:2000:PPP

[Tho00] Gregory A. Thomas. Private property and public benefit: Habitat conservation planning for endangered species. Conservation Biology, 14(1):327–328, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thornhill:2003:SSC

[Tho03] Alan Thornhill. Social scientists and conservation biologists join forces. Conservation Biology, 17(6):1475–1476, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomson:2005:ECM

[Tho05a] Diane Thomson. Experiments in conservation: More tools in the toolbox. Conservation Biology, 19(6):2039–2040, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomson:2005:MMT

[Tho05b] Diane M. Thomson. Matrix models as a tool for understanding invasive plant and native plant interactions modelos matriciales como una herramienta para entender las interacciones entre plantas invasoras y plantas nativas. Conservation Biology, 19(3):917–928, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2009:SWE

[Tho09] Jonathan R. Thompson. Salvaging what, exactly? Conservation Biology, 23(5):1333–1334, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treves:2003:HCC

[TK03a] Adrian Treves and K. Ullas Karanth. Human-carnivore conflict and perspectives on carnivore management worldwide. Conflicto entre humanos y carnívoros y perspectivas de la gestión mundial de carnívoros. Conservation Biology, 17(6):1491–1499, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treves:2003:I

[TK03b] Adrian Treves and K. Ullas Karanth. Introduction. Conservation Biology, 17(6):1489–1490, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trakhtenbrot:2006:EEC

[TK06] Ana Trakhtenbrot and Ronen Kadmon. Effectiveness of environmental cluster analysis in representing regional species diversity. Efectividad del análisis de cluster ambiental en la representación de la diversidad regional de especies. Conservation Biology, 33 (5):1087–1098, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Toivanen:2007:BLS

[TK07] Tero Toivanen and Janne S. Kotiaho. Burning of logged sites to protect beetles in managed boreal forests. Quema de sitios talados para la protección de escarabajos en bosques boreales. Conservation Biology, 21(6):1562–1572, December 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tushabe:2006:NAB

[TKB^{+}06] Herbert Tushabe, James Kalema, Achilles Byaruhanga, Josephine Asasira, Paul Ssegawa, Andrew Balmford, Tim Davenport, Jon Fjeldså, Ib Friis, Deborah Pain, Derek Pomeroy, Paul Williams, and Charles Williams. A nationwide assessment of the biodiversity value of Uganda’s important bird areas network. Una evaluación nacional del valor de la biodiversidad de la red de Áreas de importancia para aves. Conservation Biology, 20(1):85–99, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trenham:2002:ATP

[TM02] Peter C. Trenham and David M. Marsh. Amphibian translocation programs: Reply to Seigel and Dodd. Conservation Biology, 16(2):555–556, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tallmon:2004:EEI

[TM04] David A. Tallmon and L. Scott Mills. Edge effects and isolation: Red-backed voles revisited. Efectos de borde y aislamiento: Clethrionomys californicus segunda parte. Conservation Biology, 18(6):1658–1664, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thirgood:2007:I

[TMKL07] Simon J. Thirgood, Simon A. R. Mduma, Julius D. Keyyu, and M. Karen Laurenson. Introduction. Conservation Biology, 21 (3):576–579, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Teder:2007:MBD

[TMR^{+}07] Tiit Teder, Mari Moora, Elle Roosaluste, Kristjan Zobel, Meelis Pärtel, Urmas Kõljalg, and Martin Zobel. Monitoring of biological diversity: a common-ground approach. Monitoreo de la diversidad biológica: un método con denominador común. Conservation Biology, 21(2):313–317, April 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomson:2007:PBS

[TNFH07] James R. Thomson, Ralph Mac Nally, Erica Fleishman, and Greg Horrocks. Predicting bird species distributions in reconstructed landscapes. Predicción de la distribución de especies de aves en paisajes reconstruidos. Conservation Biology, 21(3): 752–766, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ticktin:2002:EVH

[TNRJ02] Tamara Ticktin, Patrick Nantel, Fernando Ramirez, and Timothy Johns. Effects of variation on harvest limits for nontimber forest species in Mexico. Efectos de la variación sobre los límites de cosecha de especies forestales no maderables en México. Conservation Biology, 16(3):691–705, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treves:2004:PHC

[TNTH^{+}04] Adrian Treves, Lisa Naughton-Treves, Elizabeth K. Harper, David J. Mladenoff, Robert A. Rose, Theodore A. Sickley, and Adrian P. Wydeven. Predicting human-carnivore conflict: a spatial model derived from 25 years of data on wolf predation on livestock. Predicción de conflicto humano–carnívoro: un modelo espacial basado en 25 años de datos de depredación de ganado por lobos. Conservation Biology, 18(1):114–125, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tomimatsu:2002:EFF

[TO02] Hiroshi Tomimatsu and Masashi Ohara. Effects of forest fragmentation on seed production of the understory herb Trillium camschatcense. Efectos de la fragmentación del bosque en la producción de semillas de la hierba de sotobosque Trillium camschatcense. Conservation Biology, 16(5):1277–1285, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Todd:2000:PDF

[Tod00] Charles R. Todd. Perspectives on the definition of fuzzy sets: a reply to Regan and colyvan. Conservation Biology, 14(4):1200– 1201, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tolisano:2007:ANN

[Tol07] James Tolisano. Artists as the new naturalists: a response and expansion to Jacobson et al. Conservation Biology, 21(5):1135– 1136, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trombulak:2004:PCB

[TOR^{+}04] Stephen C. Trombulak, Kristian S. Omland, Julie A. Robinson, Jeffrey J. Lusk, Thomas L. Fleischner, Glenn Brown, and Meg Domroese. Principles of conservation biology: Recommended guidelines for conservation literacy from the Education Committee of the Society for Conservation Biology. Conservation Biology, 18(5):1180–1190, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Townsend:2003:IPC

[Tow03] Colin R. Townsend. Individual, population, community, and ecosystem consequences of a fish invader in New Zealand streams consecuencias de un pez invasor sobre individuos, poblaciones, comunidades y ecosistema en arroyos de Nueva Zelanda. Conservation Biology, 17(1):38–47, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tabarelli:2005:COB

[TPS^{+}05] Marcelo Tabarelli, Luiz Paulo Pinto, José M. C. Silva, Márcia Hirota, and Lúcio Bedê. Challenges and opportunities for biodiversity conservation in the Brazilian Atlantic forest. Retos y oportunidades para la conservación de biodiversidad en el bosque Atlántico brasileño. Conservation Biology, 19(3):695– 700, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Towns:2007:RTS

[TPT^{+}07] David R. Towns, G. Richard Parrish, Claudine L. Tyrrell, Graham T. Ussher, Alison Cree, Donald G. Newman, A. (Tony) H. Whitaker, and Ian Westbrooke. Responses of tuatara (Sphenodon punctatus) to removal of introduced Pacific rats from islands. Respuesta del tuatara (Sphenodon punctatus) a la remoción de ratas del Pacífico introducidas en islas. Conservation Biology, 21(4):1021–1031, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Towns:2003:IVI

[TPW03] David R. Towns, G. Richard Parrish, and Ian Westbrooke. Inferring vulnerability to introduced predators without experimental demonstration: Case study of Suter’s skink in New Zealand. Infiriendo vulnerabilidad a depredadores introducidos sin Demostración experimental: Estudio de caso de Oligosoma suteri en Nueva Zelanda. Conservation Biology, 17(5):1361–1371, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tordoff:2001:RGB

[TR01] Harrison B. Tordoff and Patrick T. Redig. Role of genetic background in the success of reintroduced peregrine falcons. Papel de la condición genética en el exito de halcones peregrinos reintroducidos. Conservation Biology, 15(2):528–532, April 2001. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Traveset:2005:DPL

[TR05] Anna Traveset and Nuria Riera. Disruption of a plant–lizard seed dispersal system and its ecological effects on a threatened endemic plant in the Balearic Islands. Interrupción de un sistema de dispersión de semillas planta–lagartija y sus efectos ecológicos sobre una planta endémica amenazada en las Islas Baleares. Conservation Biology, 19(2):421–431, April 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trewhella:2005:EEC

[TRCC^{+}05] W. J. Trewhella, K. M. Rodriguez-Clark, N. Corp, A. Entwistle, S. R. T. Garrett, E. Granek, K. L. Lengel, M. J. Raboude, P. F. Reason, and B. J. Sewall. Environmental education as a component of multidisciplinary conservation programs: Lessons from conservation initiatives for critically endangered fruit bats in the Western Indian Ocean. Educación ambiental como un componente de programas multidisciplinarios de conservación: Lecciones de iniciativas de conservación para murciélagos Frugívoros en peligro crítico en el Océano Índico Occidental. Conservation Biology, 19(1):75–85, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treves:2004:HED

[Tre04] Adrian Treves. Human and ecological dimensions of conservation controversy. Conservation Biology, 18(4):1164–1165, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thirgood:2000:RRG

[TRNH00] Simon Thirgood, Steve Redpath, Ian Newton, and Peter Hudson. Raptors and red grouse: Conservation conflicts and management solutions. Rapaces y lagopo rojo: Conflictos de conservación y soluciones de manejo. Conservation Biology, 14(1): 95–104, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trombulak:2000:MND

[Tro00] Stephen C. Trombulak. Misunderstanding neo-Darwinism: a reaction to Daly. Conservation Biology, 14(4):1202–1203, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Dal00a].

Trombulak:2006:WW

[Tro06] Stephen C. Trombulak. Worth the wait. Conservation Biology, 21(3):927–928, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Toro:2000:GAC

[TRSR00] Miguel A. Toro, Jaime Rodrigañez, Luis Silio, and Carmen Rodriguez. Genealogical analysis of a closed herd of black hairless Iberian pigs. Análisis Genealógico de una manada aislada del cerdo Ibérico lampiño negro. Conservation Biology, 14(6): 1843–1851, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tsutsui:2003:CSP

[TS03] Neil D. Tsutsui and Andrew V. Suarez. The colony structure and population biology of invasive ants. Estructura de una colonia y biología poblacional de hormigas invasoras. Conservation Biology, 17(1):48–58, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomson:2006:UPC

[TS06] Diane M. Thomson and Mark W. Schwartz. Using population count data to assess the effects of changing river flow on an endangered riparian plant. Utlización de datos de conteo poblacional para evaluar los efectos del cambio de flujo en un Río sobre una planta Ribereña en peligro. Conservation Biology, 33 (5):1132–1142, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tallamy:2009:RLU

[TS09] Douglas W. Tallamy and Kimberley J. Shropshire. Ranking lepidopteran use of native versus introduced plants. Clasificación del uso de plantas nativas versus introducidas por Lepidópteros. Conservation Biology, 23(4):941–947, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2006:CRF

[TSG06] Barbara E. Taylor, David E. Scott, and J. Whitfield Gibbons. Catastrophic reproductive failure, terrestrial survival, and persistence of the marbled salamander. Fracaso reproductivo catastrófico, supervivencia terrestre y persistencia de la salamandra Ambystoma opacum. Conservation Biology, 20(3):792– 801, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Takacs:2006:EES

[TSH06] David Takacs, Daniel F. Shapiro, and William D. Head. From Is to Should: Helping students translate conservation biology into conservation policy. Conservation Biology, 20(5):1342–1348, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:2001:I

[TSJ01] Woody Turner, Eleanor J. Sterling, and Anthony C. Janetos. Introduction. Conservation Biology, 15(4):832–834, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trevino:2007:TAO

[TSKD07] Heather S. Trevino, Amy L. Skibiel, Tim J. Karels, and F. Stephen Dobson. Threats to avifauna on oceanic islands. Amenazas a la avifauna de islas Oceánicas. Conservation Biology, 21(1):125–132, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trevino:2008:ICA

[TSKD08] Heather S. Trevino, Amy L. Skibiel, Tim J. Karels, and F. Stephen Dobson. Importance of causal analysis of threats to oceanic avifaunas: Reply to Blackburn et al. Conservation Biology, 22(2):495–497, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tomillo:2008:EIH

[TSP^{+}08] Pilar Santidrián Tomillo, Vincent S. Saba, Rotney Piedra, Frank V. Paladino, and James R. Spotila. Effects of illegal harvest of eggs on the population decline of leatherback turtles in Las Baulas Marine National Park, Costa Rica. Efectos de la cosecha ilegal de huevos sobre la declinación poblacional de tortugas Laúd en el parque nacional marino las baulas, Costa Rica. Conservation Biology, 22(5):1216–1224, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2003:RCI

[TTB03] Colin J. Thompson, Ben J. P. Thompson, and Mark A. Burgman. Risks from competitively inferior immigrant populations: Implications of mass effects for species conservation. Riesgos de las poblaciones inmigrantes con competitividad inferior: Implicaciones de los efectos de masa para la conservación de especies. Conservation Biology, 17(3):901–905, June 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tenhumberg:2004:LWC

[TTSP04] Brigitte Tenhumberg, Andrew J. Tyre, Katriona Shea, and Hugh P. Possingham. Linking wild and captive populations to maximize species persistence: Optimal translocation strategies. Combinación de poblaciones silvestres y cautivas para maximizar la persistencia de especies: Estrategias de translocación Óptima. Conservation Biology, 18(5):1304–1314, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tucker:2004:SHH

[Tuc04] Mary Evelyn Tucker. Scanning the horizon for hope. Conservation Biology, 18(2):299–300, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Twedt:2006:SED

[TUE06] Daniel J. Twedt, William B. Uihlein III, and A. Blaine Elliott. A spatially explicit decision support model for restoration of forest bird habitat. Un modelo de soporte de decisiones espacialmente explícito para la restauración del hábitat de aves de bosque. Conservation Biology, 20(1):100–110, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:2006:CSF

[Tur06] James Morton Turner. Conservation science and forest service policy for roadless areas. Ciencia de la conservación y políticas del servicio forestal para Áreas sin caminos. Conservation Biology, 20(3):713–722, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tutin:2003:NPH

[Tut03] Caroline E. G. Tutin. Nine parts hard work, dedication, and good science, one part luck. Conservation Biology, 17(2):639– 642, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tapio:2006:PCN

[TVB^{+}06] I. Tapio, S. Värv, J. Bennewitz, J. Maleviciute, E. Fimland, Z. Grislis, T. H. E. Meuwissen, I. Miceikiene, I. Olsaker, H. Viinalass, J. Vilkki, and J. Kantanen. Prioritization for conservation of Northern European cattle breeds based on analysis of microsatellite data. Definición de prioridades para la conservación de razas de ganado del norte de Europa con base en el análisis de datos microsatelitales. Conservation Biology, 21(2): 1768–1779, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tellez-Valdes:2003:PAC

[TVD03] Oswaldo Téllez-Valdés and Patricia Di Vila-Aranda. Protected areas and climate change: a case study of the cacti in the Tehuacán–Cuicatlán Biosphere Reserve, México areas. Protegidas y cambio climático: un estudio de caso de los cactos en la Reserva de la Biosfera Tehuacán–Cuicatlán, México. Conservation Biology, 17(3):846–853, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:2006:ADR

[TW06] Will R. Turner and David S. Wilcove. Adaptive decision rules for the acquisition of nature reserves. Reglas de decisión adaptativas para la adquisición de reservas naturales. Conservation Biology, 20(2):527–537, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2000:IUM

[TWDB00] Barbara L. Taylor, Paul R. Wade, Douglas P. De Master, and Jay Barlow. Incorporating uncertainty into management models for marine mammals. Incorporación de la incertidumbre en modelos de manejo para mamíferos marinos. Conservation Biology, 14(5):1243–1252, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2000:CPA

[TWGH00] Paul M. Thompson, Ben Wilson, Kate Grellier, and Philip S. Hammond. Combining power analysis and population viability analysis to compare traditional and precautionary approaches to conservation of coastal cetaceans. Combinación de análisis de poder y análisis de viabilidad poblacional para comparar estrategias tradicionales y preventivas para la conservación de cetáceos. Conservation Biology, 14(5):1253–1263, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:2006:AER

[TWS06] Will R. Turner, David S. Wilcove, and Hilary M. Swain. Assessing the effectiveness of reserve acquisition programs in protecting rare and threatened species. Evaluación de la efectividad de los programas de adquisición de reservas en la protección de especies raras y amenazadas. Conservation Biology, 20(6): 1657–1669, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treves:2009:PPI

[TWW09] Adrian Treves, R. B. Wallace, and S. White. Participatory planning of interventions to mitigate human–wildlife conflicts. Planificación participativa de intervenciones para mitigar conflictos entre humanos y vida silvestre. Conservation Biology, 23(6): 1577–1587, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Underwood:2008:ESB

[UCAGG08] Jared G. Underwood, Claudia J. Hernandez Camacho, David Aurioles-Gamboa, and Leah R. Gerber. Estimating sustainable bycatch rates for California sea lion populations in the Gulf of California. Estimación de tasas de captura incidental sustentables para poblaciones de lobos marinos en el Golfo de California. Conservation Biology, 22(3):701–710, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Underwood:2009:EGN

[UKC^{+}09] Emma C. Underwood, Kirk R. Klausmeyer, Robin L. Cox, Sylvia M. Busby, Scott A. Morrison, and M. Rebecca Shaw. Expanding the global network of protected areas to save the imperiled Mediterranean biome. Expansión de la red global de Áreas protegidas para salvar al bioma Mediterráneo en peligro. Conservation Biology, 23(1):43–52, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Uryu:2001:MCF

[UMT^{+}01] Yumiko Uryu, Olaf Malm, Iain Thornton, Ian Payne, and David Cleary. Mercury contamination of fish and its implications for other wildlife of the Tapajós Basin, Brazilian Amazon. Contaminación de peces por mercurio y sus implicaciones para otras especies de vida silvestre de la Cuenca Tapajós, del Amazonas Brasileño. Conservation Biology, 15(2):438–446, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Urban:2006:CBC

[Urb06a] Mark C. Urban. Conservation beyond community. Conservation Biology, 33(5):1330–1332, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Urban:2006:RFT

[Urb06b] Mark C. Urban. Road facilitation of trematode infections in snails of Northern Alaska Los caminos facilitan las infecciones de trematodos en caracoles del norte de Alaska. Conservation Biology, 33(5):1143–1149, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Uthicke:2004:SGL

[UWB04] S. Uthicke, D. Welch, and J. A. H. Benzie. Slow growth and lack of recovery in overfished holothurians on the Great Barrier Reef: Evidence from DNA fingerprints and repeated large-scale surveys. Crecimiento lento y falta de recuperación de Holotúreos sobreexplotados en la Gran Barrera Arrecifal: Evidencia de impresiones de ADN y muestreos repetidos a gran escala. Conservation Biology, 18(5):1395–1404, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanWensveen:2004:GGG

[Van04] Louke M. Van Wensveen. Genes, Genesis and God: Values and their origins in natural and human history. Conservation Biology, 18(2):590–591, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanDyke:2005:BHE

[Van05] Fred Van Dyke. Between heaven and Earth — evangelical engagement in conservation. Conservation Biology, 19(6):1693– 1696, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanHoutan:2006:CVS

[Van06] Kyle S. Van Houtan. Conservation as virtue: a scientific and social process for conservation ethics. La conservación como virtud: un proceso científico y social para la Ética de la conservación. Conservation Biology, 20(5):1367–1372, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vance-Borland:2006:OBO

[VB06] Ken Vance-Borland. Oregon’s big outside. Conservation Biology, 20(2):588–590, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vernesi:2008:WCC

[VBB^{+}08] Cristiano Vernesi, Michael W. Bruford, Giorgio Bertorelle, Elena Pecchioli, Annapaola Rizzoli, and Heidi C. Hauffe. Where’s the conservation in conservation genetics? Conservation Biology, 22(3):802–804, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Volis:2005:RCL

[VBOV05] Sergei Volis, Gil Bohrer, Gerard Oostermeijer, and Peter Van Tienderen. Regional consequences of local population demography and genetics in relation to habitat management in Gentiana pneumonanthe. Consecuencias regionales de la demografía y genética de una población local en relación con la gestión de hábitat de Gentiana pneumonanthe. Conservation Biology, 19 (2):357–367, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vance-Borland:2008:MSW

[VBRNP08] Ken Vance-Borland, Dirk Roux, Jeanne Nel, and Bob Pressey. From the mountains to the sea: Where is freshwater conservation in the SCB agenda? Conservation Biology, 22(3):505–507, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vale:2008:EFI

[VCHBP08] Mariana M. Vale, Mario Cohn-Haft, Scott Bergen, and Stuart L. Pimm. Effects of future infrastructure development on threat status and occurrence of Amazonian birds. Efectos del futuro desarrollo de infraestructura sobre el estatus de amenaza y ocurrencia de aves amazónicas. Conservation Biology, 22(4): 1006–1015, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vallet:2001:EIC

[VDHD01] Carole Vallet, Jean-Claude Dauvin, Dominique Hamon, and Christine Dupuy. Effect of the introduced common slipper shell on the suprabenthic biodiversity of the subtidal communities in the Bay of Saint-Brieuc. Efecto de la introducción de Crepidula fornicata sobre la biodiversidad Suprabéntica de las comunidades submareas en la Bahía de Saint-Brieuc (Canal Inglés del Oeste). Conservation Biology, 15(6):1686–1690, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanderHaegen:2000:SBR

[VDP00] W. Matthew Vander Haegen, Frederick C. Dobler, and D. John Pierce. Shrubsteppe bird response to habitat and landscape variables in Eastern Washington, U.S.A. Respuesta de aves de la estepa arbustiva a variables del hábitat y el paisaje en la región este de Washington, E.U.A. Conservation Biology, 14(4): 1145–1160, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Veech:2000:CSA

[Vee00] Joseph A. Veech. Choice of species-area function affects identification of hotspots. La selección de la función especie-area afecta la identificación de zonas clave. Conservation Biology, 14(1): 140–147, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Veech:2006:CLO

[Vee06] Joseph A. Veech. A comparison of landscapes occupied by increasing and decreasing populations of grassland birds. Una comparación de paisajes ocupados por poblaciones crecientes y decrecientes de aves de pradera. Conservation Biology, 20(5): 1422–1432, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanRossum:2002:CLT

[VEST02] Fabienne Van Rossum, Ghizlane Echchgadda, Ildiko Szabadi, and Ludwig Triest. Commonness and long-term survival in fragmented habitats: Primula elatior as a study case plantas comunes y supervivencia a largo plazo en habitats fragmentados: Primula elatior como estudio de caso. Conservation Biology, 16 (5):1286–1295, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanWilgen:2004:RSF

[VGB^{+}04] B. W. Van Wilgen, N. Govender, H. C. Biggs, D. Ntsala, and X. N. Funda. Response of savanna fire regimes to changing firemanagement policies in a large African national park. Respuesta de los regimenes de fuegos de sabana a cambios en las políticas de gestión de fuego en un parque nacional africano. Conservation Biology, 18(6):1533–1540, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vickery:2002:ESP

[Vic02] Peter D. Vickery. Effects of the size of prescribed fire on insect predation of northern blazing star, a rare grassland perennial. Efectos del tamaño de fuego prescrito sobre la depredación de insectos de una especie perenne rara de pastizal. Conservation Biology, 16(2):413–421, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Villasenor:2005:HTS

[VIMMO05] Jose L. Villaseñor, Guillermo Ibarra-Manríquez, Jorge A. Meave, and Enrique Ortíz. Higher taxa as surrogates of plant biodiversity in a megadiverse country. Taxa superiores como indicadores indirectos de biodiversidad de plantas en un país megadiverso. Conservation Biology, 19(1):232–238, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vitek:2004:AES

[Vit04] Bill Vitek. Abandon environmentalism for the sake of the revolution. Conservation Biology, 18(6):1463–1464, December 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanKleunen:2007:ESC

[VJ07] Mark Van Kleunen and Steven D. Johnson. Effects of selfcompatibility on the distribution range of invasive European plants in North America. Efectos de la autocompatibilidad sobre el rango de distribución de plantas europeas invasoras en Norte América. Conservation Biology, 21(6):1537–1544, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Villard:2009:BPD

[VJ09] Marc-André Villard and Bengt Gunnar Jonsson. Biodiversity as patient: Diagnoses and treatment. Conservation Biology, 23(1): 3–4, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VonDerLippe:2007:LDD

[VK07] Moritz Von Der Lippe and Ingo Kowarik. Long-distance dispersal of plants by vehicles as a driver of plant invasions. La dispersión de plantas a larga distancia por vehículos como un agente de invasiones de plantas. Conservation Biology, 21(4):986–996, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanDriesche:2001:LIL

[VL01] Jason Van Driesche and Marcus Lane. Local institutions and local conservation. Conservation Biology, 15(4):1195–1197, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valone:2002:TPG

[VMBC02] Thomas J. Valone, Marc Meyer, James H. Brown, and Robert M. Chew. Timescale of perennial grass recovery in desertified arid grasslands following livestock removal. Escala temporal de recuperación de pasto perenne en pastizales Áridos desérticos posterior a la remoción de ganado. Conservation Biology, 16(4):995–1002, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vanderwel:2007:MAB

[VMM07] Mark C. Vanderwel, Jay R. Malcolm, and Stephen C. Mills. A meta-analysis of bird responses to uniform partial harvesting across North America. Un meta análisis de las respuestas de aves a la cosecha parcial uniforme en Norte América. Conservation Biology, 21(5):1230–1240, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vanreusel:2007:TSD

[VMV07] Wouter Vanreusel, Dirk Maes, and Hans Van Dyck. Transferability of species distribution models: a functional habitat approach for two regionally threatened butterflies. Transferibilidad de modelos de distribución de especies: un enfoque de hábitat funcional para dos mariposas regionalmente amenazadas. Conservation Biology, 21(1):201–212, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vucetich:2006:NDL

[VNP06] John A. Vucetich, Michael P. Nelson, and Michael K. Phillips. The normative dimension and legal meaning of Endangered and Recovery in the U.S. Endangered Species Act. La dimensión normativa y el significado legal de en peligro y recuperación en el Acta de Especies en Peligro de E.U.A. Conservation Biology, 20(5):1383–1390, October 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [NVP07].

Vaughan:2003:IQD

[VO03] I. P. Vaughan and S. J. Ormerod. Improving the quality of distribution models for conservation by addressing shortcomings in the field collection of training data. Mejorando la calidad de los modelos de distribución para la conservación enfocando en las limitaciones en la Recolección a campo de datos de capacitación. Conservation Biology, 17(6):1601–1611, December 2003. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Villard:2004:DPA

[VP04] Marc-André Villard and Tomas Pärt. Don’t put all your eggs in real nests: a sequel to faaborg. Conservation Biology, 18(2): 371–372, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valladares-Padua:2006:IKI

[VP06] Claudio Valladares-Padua. Importance of knowledge-intensive economic development to conservation of biodiversity in developing countries. Conservation Biology, 20(3):700–701, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vandermeer:2007:AMF

[VP07] John Vandermeer and Ivette Perfecto. The agricultural matrix and a future paradigm for conservation. Conservation Biology, 21(1):274–277, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanHeezik:2005:SCG

[VS05a] Yolanda Van Heezik and Philip John Seddon. Structure and content of graduate wildlife management and conservation biology programs: an international perspective. Conservation Biology, 19(1):7–14, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vickery:2005:RMN

[VS05b] Peter D. Vickery and W. Gregory Shriver. The road map to North American bird conservation. Conservation Biology, 19 (6):2044–2046, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valentine:2009:EWM

[VS09] Leonie E. Valentine and Lin Schwarzkopf. Effects of weedmanagement burning on reptile assemblages in Australian tropical savannas. Efectos de la quema de hierbas sobre ensambles de reptiles en sabanas tropicales de Australia. Conservation Biology, 23(1):103–113, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vilas:2006:RCI

[VSAG06] Carlos Vilas, Eduardo San Miguel, Rafaela Amaro, and Carlos Garcia. Relative contribution of inbreeding depression and eroded adaptive diversity to extinction risk in small populations of shore campion. Contribución relativa de la depresión endogámica y la diversidad adaptativa erosionada al riesgo de extinción de poblaciones Pequeñas de Silene littorea. Conservation Biology, 20(1):229–238, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valladares:2006:HFE

[VSC06] Graciela Valladares, Adriana Salvo, and Luciano Cagnolo. Habitat fragmentation effects on trophic processes of insect-plant food webs. Efectos de la fragmentación del hábitat sobre procesos tróficos de redes alimenticias insecto-planta. Conservation Biology, 20(1):212–217, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vessby:2002:SRC

[VSGS02] Karolina Vessby, Bo Söderström, Anders Glimskär, and Birgitta Svensson. Species-richness correlations of six different taxa in Swedish seminatural grasslands correlaciones de riqueza de especies de seis diferentes taxones en pastizales seminaturales en Suecia. Conservation Biology, 16(2):430–439, April 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanOosterhout:2007:GCM

[VSH^{+}07] Cock Van Oosterhout, Alan M. Smith, Bernd Hänfling, Indar W. Ramnarine, Ryan S. Mohammed, and Joanne Cable. The guppy as a conservation model: Implications of parasitism and inbreeding for reintroduction success. El guppy como modelo de conservación: Implicaciones del parasitismo y la endogamia para el Éxito de reintroducción. Conservation Biology, 21(6):1573–1583, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanOosterhout:2008:CC

[VSH^{+}08] Cock Van Oosterhout, Alan M. Smith, Bernd Hänfling, Indar W. Ramnarine, Ryan S. Mohammed, and Joanne Cable. Cover caption. Conservation Biology, 21(6):i, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanNieuwstadt:2001:ECL

[VSK01] Mark G. L. Van Nieuwstadt, Douglas Sheil, and Kuswata Kartawinata. The ecological consequences of logging in the burned forests of East Kalimantan, Indonesia. Conservation Biology, 15(4):1183–1186, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vergeer:2004:ISP

[VSO04] Philippine Vergeer, Esther Sonderen, and N. Joop Ouborg. Introduction strategies put to the test: Local adaptation versus heterosis. Estrategias de introducción puestas a prueba: Adaptación local versus heterosis. Conservation Biology, 18(3): 812–821, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vanderlaan:2009:EVA

[VT09] Angelia S. M. Vanderlaan and Christopher T. Taggart. Efficacy of a voluntary area to be avoided to reduce risk of lethal vessel strikes to endangered whales. Eficacia de un Área de exclusión voluntaria para reducir el riesgo de golpes letales de barcos a ballenas en peligro. Conservation Biology, 23(6):1467–1474, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Villa:2002:ZMP

[VTA02] Ferdinando Villa, Leonardo Tunesi, and Tundi Agardy. Zoning marine protected areas through spatial multiple-criteria analysis: the case of the Asinara Island National Marine Reserve of Italy. Zonificación de Áreas marinas protegidas mediante el análisis de criterios espaciales múltiples: el caso de la Reserva Nacional Marina de la Isla Asinara en Italia. Conservation Biology, 16(2):515–526, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vellak:2009:PPE

[VTR^{+}09] Ain Vellak, Eva-Liis Tuvi, Ülle Reier, Rein Kalamees, Elle Roosaluste, Martin Zobel, and Meelis Pärtel. Past and present effectiveness of protected areas for conservation of naturally and anthropogenically rare plant species. Efectividad pasada y presente de las Áreas protegidas para la conservación de especies de plantas raras natural y antropogénicamente. Conservation Biology, 23(3):750–757, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Volpe:2000:ENR

[VTRG00] John P. Volpe, Eric B. Taylor, David W. Rimmer, and Barry W. Glickman. Evidence of natural reproduction of aquacultureescaped Atlantic salmon in a coastal British Columbia river. Evidencia de la reproducción natural de salmones del Atlántico escapados de instalaciones de acuicultura en un río de la Costa de British Columbia. Conservation Biology, 14(3):899–903, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanDyck:2009:DCW

[VVMV09] Hans Van Dyck, Arco J. Van Strien, Dirk Maes, and Chris A. M. Van Swaay. Declines in common, widespread butterflies in a landscape under intense human use. Declinaciones en mariposas comunes, con distribución amplia, en un paisaje bajo uso humano intenso. Conservation Biology, 23(4):957–965, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanBuskirk:2004:EFB

[VW04] Josh Van Buskirk and Yvonne Willi. Enhancement of farmland biodiversity within set-aside land. Mejora de la biodiversidad en tierras cultivadas dentro de terrenos de reserva. Conservation Biology, 18(4):987–994, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [KB05].

VanBuskirk:2005:MAF

[VW05] Josh Van Buskirk and Yvonne Willi. Meta-analysis of farmland biodiversity within set-aside land: Reply to Kleijn and Báldi. Conservation Biology, 19(3):967–968, June 2005. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vucetich:2000:PVE

[VWQI00] John A. Vucetich, Thomas A. Waite, Linda Qvarnemark, and Siri Ibargüen. Population variability and extinction risk. Variabilidad poblacional y riesgo de extinción. Conservation Biology, 14(6):1704–1714, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walters:2000:AP

[WA00] Christina Walters and Michael F. Antolin. The agricultural paradox. Conservation Biology, 14(3):912–914, June 2000. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woodhams:2005:ECR

[WA05] Douglas C. Woodhams and Ross A. Alford. Ecology of chytridiomycosis in rainforest stream frog assemblages of tropical Queensland. Ecología de la quitridiomicosis en ensambles de ranas de selva lluviosa en Queensland tropical. Conservation Biology, 19(5):1449–1459, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waples:2007:BFE

[WABT07a] Robin S. Waples, Peter B. Adams, James Bohnsack, and Barbara L. Taylor. A biological framework for evaluating whether a species is threatened or endangered in a significant portion of its range. Un marco biológico para evaluar si una especie está amenazada o en peligro en una Porción significativa de su distribución. Conservation Biology, 21(4):964–974, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waples:2007:NR

[WABT07b] Robin S. Waples, Peter B. Adams, James Bohnsack, and Barbara L. Taylor. Normativity redux. Conservation Biology, 22(1): 1649–1650, December 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waples:2008:LVS

[WABT08] Robin S. Waples, Peter B. Adams, James Bohnsack, and Barbara L. Taylor. Legal viability, societal values, and SPOIR: Response to D’Elia et al. Conservation Biology, 22(4):1075–1077, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [DZM08] and reply [PD00].

Wade:2000:BMC

[Wad00] Paul R. Wade. Bayesian methods in conservation biology. Métodos Bayesianos en conservación biológica. Conservation Biology, 14(5):1308–1316, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wade:2002:EYE

[Wad02] Paul R. Wade. Everything you ever wanted to know about bottlenose dolphins. Conservation Biology, 16(4):1171–1173, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walsh:2000:SSD

[Wal00] Peter D. Walsh. Sample size for the diagnosis of conservation units. Tamaño de muestra para el diagnóstico de unidades de conservación. Conservation Biology, 14(5):1533–1537, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallace:2003:SIC

[Wal03] Richard L. Wallace. Social influences on conservation: Lessons from U.S. recovery programs for marine mammals influencias sociales en la conservación: Lecciones de los programas de recuperación de mamíferos marinos en los Estados Unidos. Conservation Biology, 17(1):104–115, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallisdevries:2004:QCA

[Wal04] Michiel F. Wallisdevries. A quantitative conservation approach for the endangered butterfly Maculinea alcon. Un método cuantitativo de conservación para la mariposa Maculinea alcon en peligro. Conservation Biology, 18(2):489–499, April 2004. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waller:2006:RVC

[Wal06] Donald M. Waller. Re-visioning conservation. Conservation Biology, 20(2):587–588, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:2009:PAM

[Wal09] Kendra L. Walker. Protected-area monitoring dilemmas: a new tool to assess success. Dilemas del monitoreo de Áreas protegidas: Una herramienta nueva para evaluar el Éxito. Conservation Biology, 23(5):1294–1303, October 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2004:AOM

[Wan04] Jinliang Wang. Application of the one-migrant-per-generation rule to conservation and management. Aplicación de la regla de Un–Migrante-por–Generación en conservación y gestión. Conservation Biology, 18(2):332–343, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ward:2000:ZVP

[War00] Paul I. Ward. Zoo visitor preferences: Reply to Balmford. Conservation Biology, 14(4):1196, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Bal00].

Ward:2005:RID

[War05] Michael P. Ward. The role of immigration in the decline of an isolated migratory bird population. El papel de la Inmigración en la declinación de una población aislada de aves migratorias. Conservation Biology, 19(5):1528–1536, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

West:2006:APS

[WB06] Paige West and Dan Brockington. An anthropological perspective on some unexpected consequences of protected areas. Conservation Biology, 20(3):609–616, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiles:2003:IBT

[WBBA03] Gary J. Wiles, Jonathan Bart, Robert E. Beck, Jr., and Celestino F. Aguon. Impacts of the brown tree snake: Patterns of decline and species persistence in Guam’s avifauna. Impactos de la culebra arbórea parda: Patrones de declinación y persistencia de especies en la avifauna de Guam. Conservation Biology, 17(5):1350–1360, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weatherhead:2002:SVL

[WBDP02] P. J. Weatherhead, G. Blouin-Demers, and K. A. Prior. Synchronous variation and long-term trends in two populations of black rat snakes. Variación sincrónica y tendencias a largo plazo en dos poblaciones de culebras ratoneras negras. Conservation Biology, 16(6):1602–1608, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Warkentin:2009:EFE

[WBSB09] Ian G. Warkentin, David Bickford, Navjot S. Sodhi, and Corey J. A. Bradshaw. Eating frogs to extinction. Conservation Biology, 23(4):1056–1059, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wielgus:2007:VAR

[WBSG07] Jeffrey Wielgus, Ford Ballantyne IV, Enric Sala, and Leah R. Gerber. Viability analysis of reef fish populations based on limited demographic information. Análisis de la viabilidad de poblaciones de peces arrecifales con base en información demográfica limitada. Conservation Biology, 21(2):447–454, April 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wintle:2005:UDL

[WBV^{+}05] Brendan A. Wintle, Sarah A. Bekessy, Lisa A. Venier, Jennie L. Pearce, and Ryan A. Chisholm. Utility of dynamic-landscape metapopulation models for sustainable forest management. Utilización de modelos metapoblacionales en paisajes dinámicos para el manejo sustentable de bosques. Conservation Biology, 19 (6):1930–1943, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:2005:PBD

[WBW05] Brian G. Walker, P. Dee Boersma, and John C. Wingfield. Physiological and behavioral differences in Magellanic penguin chicks in undisturbed and tourist-visited locations of a colony. Diferencias fisiológicas y conductuales en pollos de pingüinos en sitios no perturbados y visitados por turistas. Conservation Biology, 19(5):1571–1577, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:2006:HAM

[WBW06] Brian G. Walker, P. Dee Boersma, and John C. Wingfield. Habituation of adult Magellanic penguins to human visitation as expressed through behavior and corticosterone secretion. Habituación de pingüinos de Magallanes adultos a la visita de humanos expresada a través de comportamiento y Secreción de corticoesterona. Conservation Biology, 20(1):146–154, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wasser:2008:CIT

[WCD^{+}08] Samuel K. Wasser, William Joseph Clark, Ofir Drori, Emily Stephen Kisamo, Celia Mailand, Benezeth Mutayoba, and Matthew Stephens. Combating the illegal trade in African elephant ivory with DNA forensics. Combate del comercio ilegal de marfil de elefante africano con ADN forense. Conservation Biology, 22(4): 1065–1071, August 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weathers:2001:FEN

[WCP01] Kathleen C. Weathers, Mary L. Cadenasso, and Steward T. A. Pickett. Forest edges as nutrient and pollutant concentrators: Potential synergisms between fragmentation, forest canopies, and the atmosphere. Bordes de bosques como concentradores de nutrientes y contaminantes: Sinergismos potenciales entre la Fragmentación, el dosel del bosque y la atmósfera. Conservation Biology, 15(6):1506–1514, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watkins:2003:EFR

[WCPB03] Radley Z. Watkins, Jiquan Chen, Jim Pickens, and Kimberley D. Brosofske. Effects of forest roads on understory plants in a managed hardwood landscape. Efecto de caminos sobre las plantas del sotobosque en un paisaje forestal bajo manejo. Conservation Biology, 17(2):411–419, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2009:BVC

[WCX09] Paul Williams, Sheila Colla, and Zhenghua Xie. Bumblebee vulnerability: Common correlates of winners and losers across three continents vulnerabilidad de abejorros: Correlaciones comunes de ganadores y perdedores en tres continentes. Conservation Biology, 23(4):931–940, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Welsh:2001:CUP

[WD01] Hartwell H. Welsh, Jr. and Sam Droege. A case for using plethodontid salamanders for monitoring biodiversity and ecosystem integrity of North American forests. Caso sobre el uso de salamandras plethodóntidas en el monitoreo de la biodiversidad y la integridad del ecosistema de bosques de Norteamérica. Conservation Biology, 15(3):558–569, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willson:2003:EHD

[WD03] John D. Willson and Michael E. Dorcas. Effects of habitat disturbance on stream salamanders: Implications for buffer zones and watershed management. Efectos de la perturbación de hábitat sobre salamandras de arroyo: Implicaciones para zonas de amortiguamiento y manejo de cuencas. Conservation Biology, 17(3):763–771, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2004:I

[WD04] Jack E. Williams and Dominick A. Dellasala. Introduction. Conservation Biology, 18(4):872–873, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watling:2006:FIS

[WD06] James I. Watling and Maureen A. Donnelly. Fragments as islands: a synthesis of faunal responses to habitat patchiness. Los fragmentos como islas: una síntesis de las respuestas de la fauna a la heterogeneidad de hábitats. Conservation Biology, 33(5): 1016–1025, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilcox:2009:NCF

[WD09] C. Wilcox and C. Josh Donlan. Need for a clear and fair evaluation of biodiversity offsets for fisheries bycatch. Conservation Biology, 23(3):770–772, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watzold:2006:EEM

[WDA^{+}06] Frank Wätzold, Martin Drechsler, Claire W. Armstrong, Stefan Baumgärtner, Volker Grimm, Andreas Huth, Charles Perrings, Hugh P. Possingham, Jason F. Shogren, Anders Skonhoft, Jana Verboom-Vasiljev, and Christian Wissel. Ecological-economic modeling for biodiversity management: Potential, pitfalls, and prospects. Modelado ecológico–económico para la gestión de biodiversidad: Potencial, escollos y prospectos. Conservation Biology, 33(5):1034–1041, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wikramanayake:2002:EAR

[WDL^{+}02] Eric Wikramanayake, Eric Dinerstein, Colby Loucks, David Olson, John Morrison, John Lamoreux, Meghan McKnight, and Prashant Hedao. Ecoregions in ascendance: Reply to Jepson and Whittaker. Conservation Biology, 16(1):238–243, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitelaw:2007:PLW

[WE07] Graham S. Whitelaw and Paul F. J. Eagles. Planning for long, wide conservation corridors on private lands in the Oak Ridges Moraine, Ontario, Canada. Planificación para corredores largos y anchos en terrenos privados en la Morrena Oak Ridges, Ontario, Canadá. Conservation Biology, 21(3):675–683, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weber:2001:CPR

[Web01] Ewald Weber. Current and potential ranges of three exotic goldenrods (Solidago) in Europe. Rangos actuales y potenciales de tres especies exóticas de solidago en Europa. Conservation Biology, 15(1):122–128, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weber:2004:AAD

[Web04] Bill Weber. The arrogance of America’s designer ark. Conservation Biology, 18(1):1–3, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Webb:2005:EH

[Web05] Campbell O. Webb. Engineering hope. Conservation Biology, 19 (1):275–277, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wedekind:2002:SSL

[Wed02] Claus Wedekind. Sexual selection and life-history decisions: Implications for supportive breeding and the management of captive populations. Selección sexual y decisiones de historia de vida: Implicaciones sobre la reproducción de apoyo y el manejo de poblaciones cautivas. Conservation Biology, 16(5):1204–1211, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weldon:2006:HCR

[Wel06] Aimee J. Weldon. How corridors reduce indigo bunting nest success. Los corredores reducen el Éxito de anidación de Passerina cyanea. Conservation Biology, 33(5):1300–1305, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2009:TOD

[WELCH09] Jordan T. Watson, Timothy E. Essington, Cleridy E. Lennert-Cody, and Martin A. Hall. Trade-offs in the design of fishery closures: Management of silky shark bycatch in the Eastern Pacific Ocean tuna fishery. Ventajas y desventajas del diseño de cierres de pesquerías: Manejo de la captura incidental de Carcharhinus falciformis en la pesquería de atún en el este del Océano Pacífico. Conservation Biology, 23(3):626–635, June 2009. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wenger:2008:USP

[Wen08] Seth J. Wenger. Use of surrogates to predict the stressor response of imperiled species. Utilización de sustitutos para predecir la respuesta de especies en peligro a factores estresantes. Conservation Biology, 22(6):1564–1571, December 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Western:2003:CSA

[Wes03] David Western. Conservation science in Africa and the role of international collaboration. Conservation Biology, 17(1):11–19, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wesolowski:2005:VCH

[Wes05] Tomasz Wesołowski. Virtual conservation: How the European union is turning a blind eye to its vanishing primeval forests. Conservation Biology, 19(5):1349–1358, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2001:AFS

[WFP01] James Watson, David Freudenberger, and David Paull. An assessment of the focal-species approach for conserving birds in variegated landscapes in Southeastern Australia. Evaluación de la aproximación de especie focal para la conservación de aves en paisajes variegados en el sureste de Australia. Conservation Biology, 15(5):1364–1373, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkie:2001:IPE

[WG01] David S. Wilkie and Ricardo A. Godoy. Income and price elasticities of bushmeat demand in Lowland Amerindian societies Elasticidades de insumo y precios de demanda en sociedades Amerindias de tierras bajas. Conservation Biology, 15(3):761– 769, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weyrauch:2006:EIB

[WG06] Shauna L. Weyrauch and Thomas C. Grubb, Jr. Effects of the interaction between genetic diversity and UV-B radiation on wood frog fitness. Efectos de la Interacción entre diversidad genética y radiaciones UV-B sobre la adaptabilidad de Rana sylvatica. Conservation Biology, 20(3):802–810, June 2006. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wood:2008:ECU

[WG08] Chris C. Wood and Mart R. Gross. Elemental conservation units: Communicating extinction risk without dictating targets for protection. Unidades de conservación elementales: Comunicando el riesgo de extinción sin imponer objetivos para protección. Conservation Biology, 22(1):36–47, February 2008. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkin:2007:EEG

[WGGS07] Teddy A. Wilkin, Dany Garant, Andrew G. Gosler, and Ben C. Sheldon. Edge effects in the great tit: Analyses of long-term data with GIS techniques. Efectos de borde sobre Parus major: Análisis de datos de largo plazo con técnicas de SIG. Conservation Biology, 21(5):1207–1217, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winfree:2007:EHD

[WGK07] Rachael Winfree, Terry Griswold, and Claire Kremen. Effect of human disturbance on bee communities in a forested ecosystem. Efecto de la perturbación humana sobre comunidades de abejas en un ecosistema boscoso. Conservation Biology, 21(1):213–223, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walpole:2001:PPT

[WGW01] Matthew J. Walpole, Harold J. Goodwin, and Kari G. R. Ward. Pricing policy for tourism in protected areas: Lessons from Komodo National Park, Indonesia. Política tarifaria para el turismo en areas protegidas: Lecciones del Parque Nacional Komodo, Indonesia. Conservation Biology, 15(1):218–227, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolf:2001:EHS

[WH01] Amy T. Wolf and Susan P. Harrison. Effects of habitat size and patch isolation on reproductive success of the serpentine morning glory. Efectos del tamaño del hábitat y el aislamiento de parches en el Éxito reproductivo de la convolvulaceae serpentina Calystegia collina. Conservation Biology, 15(1):111–121, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2005:PCC

[WHA^{+}05] Paul Williams, Lee Hannah, Sandy Andelman, Guy Midgley, Miguel Araújo, Greg Hughes, Lisa Manne, Enrique Martinez-Meyer, and Richard Pearson. Planning for climate change: Identifying minimum-dispersal corridors for the Cape Proteaceae. Planificación para el cambio climático: Identificación de corredores de dispersión Mínima para las Proteaceae del Cabo. Conservation Biology, 19(4):1063–1074, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whelan:2002:MFR

[Whe02] Robert J. Whelan. Managing fire regimes for conservation and property protection: an Australian response. Conservation Biology, 16(6):1659–1661, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wheelwright:2003:SAC

[Whe03] Nathaniel T. Wheelwright. State-of-the-art conservation literature. Conservation Biology, 17(5):1469–1471, October 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolf:2000:IHP

[WHH00] Amy T. Wolf, Susan P. Harrison, and James L. Hamrick. Influence of habitat patchiness on genetic diversity and spatial structure of a serpentine endemic plant. Influencia de la fragmentación del hábitat sobre la diversidad genética y estructura espacial de una planta endémica. Conservation Biology, 14(2): 454–463, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whigham:2002:ESW

[Whi02] Dennis Whigham. Ecology for the strong, the weak, and the connoisseur. Conservation Biology, 16(1):276–278, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitaker:2003:UFT

[Whi03] Darroch M. Whitaker. The use of full-time volunteers and interns by natural-resource professionals. Conservation Biology, 21(6):330–333, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitten:2006:CPF

[Whi06] Tony Whitten. Cerulean paradise-flycatcher not extinct: Subject of the first cover lives. Conservation Biology, 21(3):918–920, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:2008:CR

[Whi08] Courtney White. The 21st century ranch. Conservation Biology, 22(6):1380–1381, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wickramasinghe:2004:ASR

[WHJJ04] Liat P. Wickramasinghe, Stephen Harris, Gareth Jones, and Nancy Vaughan Jennings. Abundance and species richness of nocturnal insects on organic and conventional farms: Effects of agricultural intensification on bat foraging. Abundancia y riqueza de especies de insectos nocturnos en granjas Orgánicas y convencionales: Efectos de la intensificación agrícola sobre el forrajeo de murciélagos. Conservation Biology, 18(5):1283–1292, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitten:2001:CBD

[WHM01] Tony Whitten, Derek Holmes, and Kathy MacKinnon. Conservation biology: a displacement behavior for academia? Conservation Biology, 15(1):1–3, February 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiegand:2002:ESS

[WHS02] Kerstin Wiegand, Klaus Henle, and Stephen D. Sarre. Extinction and spatial structure in simulation models. Extinción y estructura espacial en modelos de simulación. Conservation Biology, 16(1):117–128, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiegand:2006:ECH

[Wie06] Thorsten Wiegand. Ecological consequences of habitat loss from a population and landscape perspective. Conservation Biology, 20(2):590–592, April 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiens:2007:DBW

[Wie07] John Wiens. The dangers of black-and-white conservation. Conservation Biology, 21(5):1371–1372, October 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willers:2000:RCN

[Wil00a] Bill Willers. A response to “Current Normative Concepts in Conservation” by Callicott et al. Conservation Biology, 14(2): 570–572, April 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2000:FCB

[Wil00b] Edward O. Wilson. On the future of Conservation Biology. Conservation Biology, 14(1):1–3, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilhere:2002:AMH

[Wil02a] George F. Wilhere. Adaptive management in habitat conservation plans. Manejo adaptivo en planes de conservación de hábitat. Conservation Biology, 16(1):20–29, February 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willers:2002:PDB

[Wil02b] Bill Willers. Pleistocene diversity as a biological “Baseline”. Conservation Biology, 16(2):564–565, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2002:CGE

[Wil02c] Barry L. Williams. Conservation genetics, extinction, and taxonomic status: a case history of the regal fritillary. Genética de la conservación, extinción y estatus taxonómico: un caso de la historia de la fritilaria real. Conservation Biology, 16(1):148– 157, February 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willson:2002:FSM

[Wil02d] Mary F. Willson. The fallacy of the superfluous male. Conservation Biology, 16(2):557–559, April 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2003:UPP

[Wil03] Paul H. Wilson. Using population projection matrices to evaluate recovery strategies for Snake River spring and summer Chinook salmon. Análisis matricial de la recuperación de salmón en el Río Snake. Conservation Biology, 17(3):782–794, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilhere:2008:HME

[Wil08a] George F. Wilhere. The how-much-is-enough myth. Conservation Biology, 22(3):514–517, June 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2008:BEC

[Wil08b] Robyn S. Wilson. Balancing emotion and cognition: a case for decision aiding in conservation efforts. Equilibrio entre emoción y conocimiento: an caso para asistencia a decisiones en esfuerzos de conservación. Conservation Biology, 22(6):1452–1460, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilhere:2009:SNR

[Wil09a] George F. Wilhere. Society needs a range of alternatives: a reply to Villard and Jonsson. Conservation Biology, 23(1):4– 5, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkie:2009:SHT

[Wil09b] David S. Wilkie. Strange histories and true confessions. Conservation Biology, 23(1):236–237, February 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winner:2004:SPT

[Win04] Langdon Winner. Science policy for true believers. Conservation Biology, 18(4):866, August 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

With:2002:LEI

[Wit02] Kimberly A. With. The landscape ecology of invasive spread. La ecología de paisaje de extensiones invasoras. Conservation Biology, 16(5):1192–1203, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

With:2003:HEC

[Wit03] Kimberly A. With. How to Excel in conservation biology. Conservation Biology, 17(3):931–933, June 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Webb:2009:HGT

[WK09] Edward L. Webb and Md. Enamul Kabir. Home gardening for tropical biodiversity conservation. Conservation Biology, 23(6): 1641–1644, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilcove:2004:UER

[WL04] David S. Wilcove and Joon Lee. Using economic and regulatory incentives to restore endangered species: Lessons learned from three new programs. Utilización de incentivos económicos y normativos para restaurar especies en peligro: Lecciones aprendidas de tres programas nuevos. Conservation Biology, 18(3):639–645, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willott:2000:ESL

[WLCS00] S. J. Willott, D. C. Lim, S. G. Compton, and S. L. Sutton. Effects of selective logging on the butterflies of a Bornean rainforest. Efectos del talado selectivo sobre las mariposas en un bosque tropical lluvioso de Borneo. Conservation Biology, 14(4):1055– 1065, August 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williamson:2000:ATM

[WLO^{+}00] G. Bruce Williamson, William F. Laurance, Alexandre A. Oliveira, Patricia Delamônica, Claude Gascon, Thomas E. Lovejoy, and Luciano Pohl. Amazonian tree mortality during the 1997 El Niño drought. Mortandad de arboles del Amazonas durante la sequía de El Niño de 1997. Conservation Biology, 14(5): 1538–1542, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2008:TED

[WLS^{+}08] Dajun Wang, Sheng Li, Shan Sun, Hao Wang, Ai Chen, Shengzhi Li, Juan Li, and Zhi Lu. Turning earthquake disaster into long-term benefits for the panda. Conservation Biology, 22(5):1356–1360, October 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2001:TFN

[WM01] Yeqiao Wang and Debra K. Moskovits. Tracking fragmentation of natural communities and changes in land cover: Applications of Landsat data for conservation in an urban landscape (Chicago Wilderness). Rastreo de la fragmentación de comunidades naturales y de cambios en la cobertura de suelo: Aplicaciones de datos de Landsat para la conservación en un paisaje urbano (Chicago Wilderness). Conservation Biology, 15(4):835– 843, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Welch:2005:IHV

[WM05] Nathan E. Welch and James A. Macmahon. Identifying habitat variables important to the rare Columbia spotted frog in Utah (U.S.A.): an information-theoretic approach. Identificación de variables del hábitat importantes para Rana luteiventris en Utah (E.U.A.): un método Teórico–Informático. Conservation Biology, 19(2):473–481, April 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wellnitz:2002:SWG

[WMCB02] Todd Wellnitz, Nicole MacRury, Ana Child, and Delwin Benson. Spreading the wealth: Graduate students and educational outreach. Conservation Biology, 16(2):560–563, April 2002. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wikramanayake:2004:DCL

[WMD^{+}04] Eric Wikramanayake, Meghan Mcknight, Eric Dinerstein, Anup Joshi, Bhim Gurung, and David Smith. Designing a conservation landscape for tigers in human-dominated environments. Diseño de un paisaje de conservación para tigres en ambientes dominados por humanos. Conservation Biology, 18(3):839–844, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkie:2006:PPA

[WMD^{+}06] David S. Wilkie, Gilda A. Morelli, Josefien Demmer, Malcolm Starkey, Paul Telfer, and Matthew Steil. Parks and people: Assessing the human welfare effects of establishing protected areas for biodiversity conservation. Conservation Biology, 20 (1):247–249, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wright:2009:FTS

[WMLS09] S. Joseph Wright, Helene C. Muller-Landau, and Jan Schipper. The future of tropical species on a warmer planet. El futuro de especies tropicales en un planeta más caliente. Conservation Biology, 23(6):1418–1426, December 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waltert:2004:ELU

[WMM04] Matthias Waltert, Ani Mardiastuti, and Michael Mühlenberg. Effects of land use on bird species richness in Sulawesi, Indonesia. Efectos del uso de suelo sobre la riqueza de especies de aves en Sulawesi, Indonesia. Conservation Biology, 18(5):1339–1346, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilsey:2005:PPE

[WMP05] Brian J. Wilsey, Leanne M. Martin, and H. Wayne Polley. Predicting plant extinction based on species-area curves in prairie fragments with high beta richness. Predicción de la extinción de plantas con base en curvas Especie-Área en fragmentos de pradera con riqueza beta alta. Conservation Biology, 19(6): 1835–1841, December 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiley:2008:EVC

[WMPC08] David N. Wiley, Just C. Moller, Richard M. Pace III, and Carole Carlson. Effectiveness of voluntary conservation agreements: Case study of endangered whales and commercial whale watching. Efectividad de los acuerdos de conservación voluntarios: un caso de estudio de ballenas en peligro y la observación comercial de ballenas. Conservation Biology, 22(2):450–457, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willson:2001:PPR

[WMS^{+}01] Mary F. Willson, Joan L. Morrison, Kathryn E. Sieving, Toni L. De Santo, Leonard Santisteban, and Iván Díaz. Patterns of predation risk and survival of bird nests in a Chilean agricultural landscape. Patrones de riesgo de depredación y supervivencia de nidos de aves en un paisaje agrícola chileno. Conservation Biology, 15(2):447–456, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wintle:2003:UBM

[WMVK03] B. A. Wintle, M. A. Mccarthy, C. T. Volinsky, and R. P. Kavanagh. The use of Bayesian model averaging to better represent uncertainty in ecological models. El uso de promedios de modelo Bayesiano para mejorar la representación de la incertidumbre en modelos ecológicos. Conservation Biology, 17(6):1579–1590, December 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolff:2000:CEW

[Wol00] Wim J. Wolff. Causes of extirpations in the Wadden Sea, an estuarine area in the Netherlands. Causas de extirpaciones en el Mar Wadden, un area estuarina en los Países Bajos. Conservation Biology, 14(3):876–885, June 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolman:2006:MMC

[Wol06] Abel G. Wolman. Measurement and meaningfulness in conservation science. Medición y significado en la ciencia de la conservación. Conservation Biology, 20(6):1626–1634, December 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woodroffe:2000:CAM

[Woo00] Rosie Woodroffe. Conservationists ain’t misbehavin’. Conservation Biology, 14(6):1927–1928, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woodruff:2001:AFA

[Woo01] David S. Woodruff. The archetypal flagship adrift. Conservation Biology, 15(4):1189–1192, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woodwell:2003:WET

[Woo03] George M. Woodwell. World enough and time? Conservation Biology, 17(2):356–357, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Westphal:2003:ADT

[WP03] Michael I. Westphal and Hugh P. Possingham. Applying a decision-theory framework to landscape planning for biodiversity: Follow-up to Watson et al. Conservation Biology, 21(6): 327–329, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:2008:IRM

[WPS08] Susan Walker, Robbie Price, and R. T. Theo Stephens. An index of risk as a measure of biodiversity conservation achieved through land reform. Un Índice de riesgo como una medida de la conservación de la biodiversidad llevada a cabo con reforma de terrenos. Conservation Biology, 22(1):48–59, February 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2001:GEM

[WR01] Jinliang Wang and Nils Ryman. Genetic effects of multiple generations of supportive breeding. Efectos genéticos de múltiples generaciones de reproducción de apoyo. Conservation Biology, 15(6):1619–1631, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wright:2007:ASD

[Wri07] David H. Wright. The advocacy and science divide. Conservation Biology, 21(4):901–902, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wise:2002:MPC

[WRL02] Carla A. Wise, Tom A. Ranker, and Yan B. Linhart. Modeling problems in conservation genetics with Brassica rapa: Genetic variation and fitness in plants under mild, stable conditions. Problemas de modelado en génetica de conservación con Brassica rapa: Variación genética y adaptación de plantas bajo condiciones moderadas y estables. Conservation Biology, 16(6): 1542–1554, December 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiegand:2005:EHL

[WRM05] Thorsten Wiegand, Eloy Revilla, and Kirk A. Moloney. Effects of habitat loss and fragmentation on population dynamics. Efectos de la pérdida y fragmentación del hábitat sobre la dinámica poblacional. Conservation Biology, 19(1):108–121, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wisdom:2002:MES

[WRW^{+}02] Michael J. Wisdom, Mary M. Rowland, Barbara C. Wales, Miles A. Hemstrom, Wendel J. Hann, Martin G. Raphael, Richard S. Holthausen, Rebecca A. Gravenmier, and Terrell D. Rich. Modeled effects of sagebrush-steppe restoration on greater sage-grouse in the interior Columbia Basin, U.S.A. Efectos modelados de la restauración de la estepa artemisa sobre urogallos en el interior de la Cuenca del Columbia, E.U.A. Conservation Biology, 16(5):1223–1231, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

West:2003:RRC

[WS03] Jordan M. West and Rodney V. Salm. Resistance and resilience to coral bleaching: Implications for coral reef conservation and management. Resistencia y capacidad de recuperación al blanqueado de corales: Implicaciones para la conservación y el manejo de los arrecifes de coral. Conservation Biology, 17 (4):956–967, August 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ward:2004:CAC

[WS04] Michael P. Ward and Scott Schlossberg. Conspecific attraction and the conservation of territorial songbirds. Atracción conespecífica y la conservación de aves canoras territoriales. Conservation Biology, 18(2):519–525, April 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkie:2005:RPW

[WSA^{+}05] David S. Wilkie, Malcolm Starkey, Kate Abernethy, Ernestine Nstame Effa, Paul Telfer, and Ricardo Godoy. Role of prices and wealth in consumer demand for bushmeat in Gabon, Central Africa. Papel de los precios y riqueza en la demanda de consumo de carne silvestre en Gabón, África Central. Conservation Biology, 19(1):268–274, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2009:BSA

[WSBP09] Duncan S. Wilson, Margo A. Stoddard, Matthew G. Betts, and Klaus J. Puettmann. Bayesian small area models for assessing wildlife conservation risk in patchy populations. Modelos Bayesianos de Área reducida para evaluar el riesgo de conservación de vida silvestre en poblaciones fragmentadas. Conservation Biology, 23(4):982–991, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weigle:2005:ARI

[WSCP05] Shannon M. Weigle, L. David Smith, James T. Carlton, and Judith Pederson. Assessing the risk of introducing exotic species via the live marine species trade. Evaluación del riesgo de introducciones de especies exóticas a través del comercio de especies marinas vivas. Conservation Biology, 19(1):213–223, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Witte:2008:EAF

[WSKH08] Carmel L. Witte, Michael J. Sredl, Andrew S. Kane, and Laura L. Hungerford. Epidemiologic analysis of factors associated with local disappearances of native ranid frogs in Arizona. Análisis epidemiológico de factores asociados con desapariciones locales de ranas nativas en Arizona. Conservation Biology, 22 (2):375–383, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wimberly:2000:SHV

[WSLW00] Michael C. Wimberly, Thomas A. Spies, Colin J. Long, and Cathy Whitlock. Simulating historical variability in the amount of old forests in the Oregon Coast Range. Simulación de variabilidad histórica en la cantidad de bosque maduro en la Cordillera Costera de Oregon. Conservation Biology, 14(1):167–180, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:2009:JCE

[WSPL09] Paul H. Wilson, Howard A. Schaller, Charles E. Petrosky, and John Loomis. Judging cost-effectiveness of management of Snake River salmon: Response to Halsing and Moore. Conservation Biology, 23(2):475–478, April 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitman:2007:MET

[WSQP07] Karyl L. Whitman, Anthony M. Starfield, Henley Quadling, and Craig Packer. Modeling the effects of trophy selection and environmental disturbance on a simulated population of African lions. Modelación de los efectos de la selección de presas y de la perturbación ambiental en una población simulada de leones africanos. Conservation Biology, 21(3):591–601, June 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkie:2000:RDC

[WSR^{+}00] David Wilkie, Ellen Shaw, Fiona Rotberg, Gilda Morelli, and Philippe Auzel. Roads, development, and conservation in the Congo Basin. Carreteras, desarrollo y conservación en la Cuenca del Congo. Conservation Biology, 14(6):1614–1622, December 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Warman:2004:SSR

[WSS^{+}04] Leanna D. Warman, A. R. E. Sinclair, G. G. E. Scudder, Brian Klinkenberg, and Robert L. Pressey. Sensitivity of systematic reserve selection to decisions about scale, biological data, and targets: Case study from Southern British Columbia. Sensibilidad de la selección sistemática de reservas a decisiones sobre escala, datos biológicos y objetivos: Estudio de caso del Sur de Columbia Británica. Conservation Biology, 18(3):655–666, June 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wright:2001:NPN

[WTEH^{+}01] Timothy F. Wright, Catherine A. Toft, Ernesto Enkerlin-Hoeflich, Jaime Gonzalez-Elizondo, Mariana Albornoz, Adriana Rodríguez-Ferraro, Franklin Rojas-Suárez, Virginia Sanz, Ana Trujillo, Steven R. Beissinger, Vicente Berovides A., Xiomara Gálvez A., Ann T. Brice, Kim Joyner, Jessica Eberhard, James Gilardi, S. E. Koenig, Scott Stoleson, Paulo Martuscelli, J. Michael Meyers, Katherine Renton, Angélica M. Rodríguez, Ana C. Sosa-Asanza, Francisco J. Vilella, and James W. Wiley. Nest poaching in neotropical parrots saqueo de nidos de loros neotropicales. Conservation Biology, 15(3):710– 720, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallace:2008:AES

[WTEK08] George N. Wallace, David M. Theobald, Tawnya Ernst, and Katherine King. Assessing the ecological and social benefits of private land conservation in Colorado. Evaluación de los beneficios ecológicos y sociales de la conservación de terrenos privados en Colorado. Conservation Biology, 22(2):284–296, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolf:2001:PRE

[WTR01] Diana E. Wolf, Naoki Takebayashi, and Loren H. Rieseberg. Predicting the risk of extinction through hybridization. Predicción del riesgo de extinción por hibridación. Conservation Biology, 15(4):1039–1053, August 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiersma:2005:BDN

[WU05] Yolanda F. Wiersma and Dean L. Urban. Beta diversity and nature reserve system design in the Yukon, Canada. Diversidad beta y diseño del sistema de reservas naturales en el Yukon, Canadá. Conservation Biology, 19(4):1262–1272, August 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wunder:2007:EPE

[Wun07] Sven Wunder. The efficiency of payments for environmental services in tropical conservation. La eficiencia de los pagos por servicios ambientales en la conservación trópicos. Conservation Biology, 21(1):48–58, February 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wasson:2003:DNS

[WWH03] Kerstin Wasson, Andrea Woolfolk, and Kim Hayes. Developing national strategies to combat biological invasions. Conservation Biology, 21(6):338–340, February 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wisdom:2002:PGS

[WWR^{+}02] Michael J. Wisdom, Barbara C. Wales, Mary M. Rowland, Martin G. Raphael, Richard S. Holthausen, Terrell D. Rich, and Victoria A. Saab. Performance of greater sage-grouse models for conservation assessment in the interior Columbia Basin, U.S.A. funcionamiento de modelos para la evaluación de la conservación de urogallos en la cuenca Columbia interior, E.U.A. Conservation Biology, 16(5):1232–1242, October 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wright:2000:PAM

[WZD^{+}00] S. Joseph Wright, Horacio Zeballos, Iván Domínguez, Marina M. Gallardo, Marta C. Moreno, and Roberto Ibáñez. Poachers alter mammal abundance, seed dispersal, and seed predation in a neotropical forest. Los cazadores alteran la abundancia, dispersión de semillas y depredación de semillas en un bosque neotropical. Conservation Biology, 14(1):227–239, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2006:CBN

[WZW^{+}06] Ding Wang, Xianfeng Zhang, Kexiong Wang, Zhuo Wei, Bernd Würsig, Gillian T. Braulik, and Susie Ellis. Conservation of the baiji: No simple solution. Conservation Biology, 20(3):623–625, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Xu:2009:MHC

[XGS^{+}09] Jianchu Xu, R. Edward Grumbine, Arun Shrestha, Mats Eriksson, Xuefei Yang, Yun Wang, and Andreas Wilkes. The melting Himalayas: Cascading effects of climate change on water, biodiversity, and livelihoods. Las Himalaya se derriten: Efectos en cascada del cambio climático sobre el agua, la biodiversidad y los medios de vida. Conservation Biology, 23(3):520–530, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [Pan09].

Xu:2007:REP

[XM07] Jianchu Xu and David R. Melick. Rethinking the effectiveness of public protected areas in Southwestern China. Repensando la efectividad de Áreas protegidas públicas en el suroeste de China. Conservation Biology, 21(2):318–328, April 2007. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yahner:2004:WT

[Yah04] Richard H. Yahner. The work of taxonomy. Conservation Biology, 18(1):6–7, February 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yang:2006:COB

[YBWZ06] Guang Yang, Michael W. Bruford, Fuwen Wei, and Kaiya Zhou. Conservation options for the baiji: Time for realism? Conservation Biology, 20(3):620–622, June 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yakimowski:2007:TPP

[YE07] Sarah B. Yakimowski and Christopher G. Eckert. Threatened peripheral populations in context: Geographical variation in population frequency and size and sexual reproduction in a clonal woody shrub. Poblaciones periféricas amenazadas en contexto: Variación geográfica en la frecuencia y tamaño poblacional y la reproducción sexual de un arbusto leñoso clonal. Conservation Biology, 21(2):811–822, June 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:2001:CRG

[YH01] Michael K. Young and Amy L. Harig. A critique of the recovery of greenback cutthroat trout. Crítica a la recuperación de la trucha de dorso verde. Conservation Biology, 15(6):1575–1584, December 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yamaura:2008:RIA

[YKIO08] Yuichi Yamaura, Takayuki Kawahara, Shigeo Iida, and Kenichi Ozaki. Relative importance of the area and shape of patches to the diversity of multiple taxa. Importancia relativa de la superficie y la forma de fragmentos para la diversidad de múltiples taxa. Conservation Biology, 22(6):1513–1522, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yates:2005:RIR

[YL05] Colin J. Yates and Philip G. Ladd. Relative importance of reproductive biology and establishment ecology for persistence of a rare shrub in a fragmented landscape. Importancia relativa de la biología reproductiva y la ecología para la persistencia de una hierba rara en un paisaje fragmentado. Conservation Biology, 19(1):239–249, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:2001:PDP

[YLR^{+}01] Bruce E. Young, Karen R. Lips, Jamie K. Reaser, Roberto Ibáñez, Antonio W. Salas, J. Rogelio Cedeño, Luis A. Coloma, Santiago Ron, Enrique La Marca, John R. Meyer, Antonio Muñoz, Federico Bolaños, Gerardo Chaves, and David Romo. Population declines and priorities for amphibian conservation in Latin America. Disminución de Poblaciones y prioridades para la conservación de anfibios en Latinoamérica. Conservation Biology, 15(5):1213–1223, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yount:2000:AP

[You00] J. David Yount. “Accidentally on purpose”. Conservation Biology, 14(1):4, February 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:2001:DUC

[You01] Kyle A. Young. Defining units of conservation for intraspecific biodiversity: Reply to Dimmick et al. Conservation Biology, 15(3):784–787, June 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:2002:MCK

[You02] Kenneth R. Young. Minding the children: Knowledge transfer and the future of sustainable agriculture. Conservation Biology, 16(4):855–856, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:2005:FMC

[You05] Carlos Eduardo Frickmann Young. Financial mechanisms for conservation in Brazil. Mecanismos financieros para la conservación en Brasil. Conservation Biology, 19(3):756–761, June 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yule:2006:RES

[Yul06] Jeffrey V. Yule. Resurrection ecology in the service of America’s evolutionary legacy. Conservation Biology, 33(5):1326– 1328, August 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zink:2000:GTC

[ZBABR00] Robert M. Zink, George F. Barrowclough, Jonathan L. Atwood, and Rachelle C. Blackwell-Rago. Genetics, taxonomy, and conservation of the threatened California gnatcatcher. Genética, taxonomía, y conservación de la perlita de California amenazado de extinción. Conservation Biology, 14(5):1394–1405, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zacharias:2005:SVM

[ZG05] Mark A. Zacharias and Edward J. Gregr. Sensitivity and vulnerability in marine environments: an approach to identifying vulnerable marine areas. Sensibilidad y vulnerabilidad en ambientes marinos: una metodología para identificar Áreas marinas vulnerables. Conservation Biology, 19(1):86–97, February 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zurlini:2002:SAP

[ZGR02] Giovanni Zurlini, Luigi Grossi, and Orazio Rossi. Spatialaccumulation pattern and extinction rates of Mediterranean flora as related to species confinement to habitats in preserves and larger areas. Patrón de acumulación espacial y tasas de extinción de flora Mediterránea relacionada con el confinamiento de especies a hábitats en reservas y Áreas mayores. Conservation Biology, 16(4):948–963, August 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zabel:2008:CPP

[ZHM08] Astrid Zabel and Karin Holm-Müller. Conservation performance payments for carnivore conservation in Sweden. Conservation Biology, 22(2):247–251, April 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhou:2004:ITS

[ZJ04] Zhihua Zhou and Zhigang Jiang. International trade status and crisis for snake species in China estatus del comercio internacional y crisis para especies de serpientes en China. Conservation Biology, 18(5):1386–1394, October 2004. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zint:2002:ECB

[ZKNL02] Michaela Zint, Anita Kraemer, Heather Northway, and Miyoun Lim. Evaluation of the Chesapeake Bay Foundation’s conservation education programs. Evaluación de los programas de educación en conservación de la Fundación Chesapeake Bay. Conservation Biology, 16(3):641–649, June 2002. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zonneveld:2003:OSD

[ZLM03] C. Zonneveld, T. Longcore, and C. Mulder. Optimal schemes to detect the presence of insect species. Planes Óptimos para detectar la presencia de especies de insectos. Conservation Biology, 17(2):476–487, April 2003. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zavaleta:2008:EEU

[ZMS^{+}08] Erika Zavaleta, Daniel C. Miller, Nick Salafsky, Erica Fleishman, Michael Webster, Barry Gold, David Hulse, Mary Rowen, Gary Tabor, and Jack Vanderryn. Enhancing the engagement of U.S. private foundations with conservation science. Incremento del compromiso de las fundaciones privadas de E.U.A. con la ciencia de la conservación. Conservation Biology, 22(6):1477–1484, December 2008. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zacharias:2000:HEA

[ZR00] Mark A. Zacharias and John C. Roff. A hierarchical ecological approach to conserving marine biodiversity. Estrategia ecológica por Jerarquías para conservar la biodiversidad marina. Conservation Biology, 14(5):1327–1334, October 2000. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [SRSP01].

Zacharias:2001:ZRV

[ZR01] Mark A. Zacharias and John C. Roff. Zacharias and Roff vs. Salomon et al.: Who adds more value to marine conservation efforts? Conservation Biology, 15(5):1456–1458, October 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zorn:2001:EMP

[ZSG01] Paul Zorn, William Stephenson, and Paul Grigoriev. An ecosystem management program and assessment process for Ontario National Parks. Programa de manejo del ecosistema y proceso de evaluación del Parque Nacional Ontario. Conservation Biology, 15(2):353–362, April 2001. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zabel:2006:IBC

[ZSMW06] Richard W. Zabel, Mark D. Scheuerell, Michelle M. Mcclure, and John G. Williams. The interplay between climate variability and density dependence in the population viability of Chinook salmon. La interacción entre variabilidad climática y denso dependencia en la viabilidad poblacional de Oncorhynchus tshawytscha. Conservation Biology, 20(1):190–200, February 2006. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zidon:2009:BCS

[ZSSM09] Royi Zidon, David Saltz, Laurence S. Shore, and Uzi Motro. Behavioral changes, stress, and survival following reintroduction of Persian fallow deer from two breeding facilities. Cambios conductuales, estrés y supervivencia después de la reintroducción de Dama mesopotamica proveniente de dos criaderos. Conservation Biology, 23(4):1026–1035, August 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zarin:2007:BRF

[ZSVL07] Daniel J. Zarin, Mark D. Schulze, Edson Vidal, and Marco Lentini. Beyond reaping the first harvest: Management objectives for timber production in the Brazilian Amazon. Más allá de la primera cosecha: Objetivos de manejo para la producción de madera en la Amazonía brasileña. Conservation Biology, 21 (4):916–925, August 2007. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zydelis:2009:CMM

[ŽWGW09] Ramūnas Žydelis, Bryan P. Wallace, Eric L. Gilman, and Timothy B. Werner. Conservation of marine megafauna through minimization of fisheries bycatch. Conservación de la megafauna marina medianate la minimización de la captura incidental de pesquerías. Conservation Biology, 23(3):608–616, June 2009. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zedler:2005:RC

[ZZ05] Paul H. Zedler and Joy B. Zedler. Reality check. Conservation Biology, 19(5):1675–1676, October 2005. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).